A Complete Bibliography of Electronic Journal of Probability

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

20 October 2023 Version 1.19

Title word cross-reference

(2 + 1) [Ler20]. (T) [Sch06]. + [RT09]. 0 [Hei13]. 1 [BS20c, Boi11, DFS21, Fag12, Hei13, KN17, SBB^{+}21, Unt19]. 1 + 1 [CN18, SS13]. 1/6 [NRS10]. 2 [AS18, Ber06, CCFR09, DKNS13, Din12, FG19, GL18, Gro20, HRW12, LSW01, PYZ22, Sei10, Zha21]. 3 [CCD13, CHNR22, Fer16, FR01, GGY18, Gal07, Maf20, Oda06, SY14]. = [RT09]. α [BR22, EG15, GSS21, HJ09, Han14, Nan06, PS22a, XZ10b]. b [Ber15]. β [AG13, DLV21, Lup21, Web16]. BB [DDT14]. C^{1} [KKH11]. C^{1,1} [PS22a]. d [BST16, Hol06, NN21]. Δ + Δ^{α/2} [CKS12]. ℓ^{p} [Dom22]. ℓ_{∞}^{N} [Osȩ13]. F [JN10]. \frac{1}{4} < H < \frac{1}{2} [BJQS15]. G [Dol12, Lin13, Son12, XZ10a, BPT16]. G(n, p) [AJL14]. GF[2][2] [DPWZ14]. H1 [RV07]. H = 1/2 [BDR22]. H > 1/2 [DS07a]. k [Ber11, CYY22, FR13, Tra15]. K_{4} [Kol22]. κ [WW13]. κ, ρ [WW13]. L [Pai16]. L^{2} [GT07, Hut20, MR12, Web15]. L^{p} [HL14, Ose11, OZ21]. L_{p} [Kim05, MR01]. L_{q} [LLS06]. L_{q}((0, τ], L_{p}) [Kry00]. L_{q}(L_{p}) [CKLL13]. Λ [CCSWB22, Lab14, LS06, LT15c, LZ12]. m [GN03, Sch08]. R^{d} [MV00, DRM14]. Z [BGRS10, LR20, Uch20, VT08]. Z^{2} [KSZ01]. Z^{d} [SCZ13, Ger20]. Z_{+}^{d} [Lev13]. H^{1} [Mon15]. N [BM14, BC08b, PRT22, Eld14]. p [AMP04, BV12, Che13a, FV11, Pin22, Sam09]. ∂u/∂t = ±∂^{N}u/∂x^{N} [Lac07]. Φ_{3}^{4} [BG21a, EX22]. q [ASD21, BV12, OP13, Pes22, Sza10, Vet15]. R^{2} [RM09]. R_{+} [BD03]. ρx^{α}/t^{β} [MV08]. s [Fra09]. SO_{R}(N) [DMN15]. U [MW19, Zha22]. V [SHW20]. ε [HLP10]. W_{2}^{1} [KKH11]. W_{1,+} [Hil14]. Ξ [Lim10]. Z [BP06, dT21]. Z^{2} [AH08].

-actions [Lev13]. -adic [Sam09]. -ary [Ber15, NN21]. -bounds [OZ21]. -branching [Pai16]. -Brownian [Lin13]. -card-minimum [Tra15]. -CIR [Han14]. -Coalescent [Lim10, LS06]. -coalescents [CCSWB22, LT15c]. -colourings [Gal07]. -concave [Fra09]. -Cycles [Ber11]. -Delaunay [GKT22]. -dimension [CGZ06]. -Dimensional [CCFR09, Ber06, DFS21, Eld14, Gro20, Ler20]. -Dirac [dT21]. -Doubly [JN10]. -Dyson [Lup21]. -ensemble [Web16]. -ensembles [AG13, DLV21]. -Estimates [Kim05]. -Expectations [Dol12]. [?]FlemingLabbe:2014:FFV,Liu:2012:CSP. -fold [CYY22]. -Framework [XZ10a]. -geometrically [DDT14]. -GREM [FG19]. -Hahn [Vet15]. -Inequalities [Ose11]. -interpolation [Hil14]. -invariant [dT21]. -law [Hei13]. -Mappings [AMP04]. -martingales [Son12]. -nearest [FR13]. -Normal [Sza10]. -order [GN03]. -particle [BM14, PRT22]. -percolation [Kol22]. -phase [Web15]. -regular [BST16]. -regularity [CKLL13]. -Separation [HLP10]. -shifted [Pin22]. -SLE [Zha21]. -Smooth [Sei10]. -Spaces [LLS06]. -spin [BS22b, Che13a]. -Stable [XZ10b, EG15, HJ09, PS22a]. -state [GGY18]. -statistics [MW19, SHW20, Zha22]. -sub-exponential [GSS21]. -supermartingale [BPT16]. -TASEP [Pes22]. -Theory [KKH11, MR01]. -time [Nan06]. -Variation [FV11]. -voter [ASD21]. -Wasserstein [Boi11]. -weighted [OP13].

1-2 [Li14]. 1d [Shc22].

2d [GNZ21].

3d [STV22].

stable[BR22].

Abelian [HL07]. Above [vdHS10, DSWB15, Hol08, vdHS04]. Absolute [BV16, CG18, NNP13, Cri21, SSS13]. absolutely [DRM14]. Absorbing [ST17b, Tag16]. Absorbing-state [ST17b, Tag16]. Abstract [Doh99, Fd00]. account [Lej02]. accumulation [TOR15]. Accurate [KN07]. acting [Sam17]. actions [Lev13]. activated [Tag16, ST17b]. Acyclic [EL07a]. Adaptive [BBM20, Vih11]. add [CT22a]. add-one [CT22a]. Addendum [EK23]. Additive [BV04, HM01, Mie01, Uch07, BS20b, Bel20b, CM20, Che19, CK22, GHM22, Gao17, Jan22, LO20, MR15, Pèn21, Var19]. adic [Sam09]. adjoint [Ho22]. Adsorption [CMT08]. advection [DG14]. advection-dispersion [DG14]. Affine [KL18, BJQS15, KRST13, Sén21]. after [Ful05]. Age [GS07, Che17b]. age-dependent [Che17b]. Age-Ordered [GS07]. Ages [Lam01]. Aggregation [STZ11, Coh12, Haa17, HS12]. Aging [Bus16, FG19, Gay12]. agree [Red17]. agrees [Bav19]. Airy [ACvM11, DJM16]. Aizenman [BK09, Che13a]. Aldous [Lin16]. algebra [Chh15, Duf10]. Algorithm [Nor10, Vih11, Gol18b, LP08, OP13, OC19]. algorithms [BCG21, CL21b, SND14, Tur15]. Allele [BG08, GPPW16]. Allen [BDW17, BOS22, EGL22, HRW12, OWW14]. Allocation [Pen09]. Almost [BR05, GR15, KSZ98, LM96, LV11, Tak21, Ada20, AP21, AP23, IV13, Mas22]. Along [Fd06, AD14d, Beg05, Die16, LMM04, MP20]. Alpha [BBC^{+}05, CFW09, HJ07]. alpha-gamma [CFW09]. Alpha-Stable [BBC^{+}05]. alphabet [Gil16]. Alternative [DW19]. Ambjørn [BJM14]. amenable [MR20, Sam09]. American [BT05, CB09, KNR21]. Among [BP07, Čer11, Fag10, BS22a, BS20c, Bis19, CO21, DS12a, Fag08, GP12, GPV09, GJ17, Mou12]. Amplitude [BM09b]. Analogues [Uch10]. Analysis [BV04, BBM20, BdSGP09, Dav08, DG96, FG96, Fuk11, HK20, RNP10, STZ11, SP19, Win02, AGMQ22, AF20, Ber23, Coh12, DG99, Fag12, FJ12, GK14, HT10a, KK18, LSW22, MR13, SY17, SF06]. Analytic [Fd01, MT22, GIP16]. Analytical [DS04]. Anatomy [CK23]. Ancestor [Tay07]. ancestry [BČD16]. Anchored [GJ14, Hut23]. Andersen [BDT19]. Anderson [CHNT17, CHSS18, Dre08, FM14, GS11, MT04, NXZ22]. angles [Yao13]. anharmonic [Xu20]. Animals [MW10]. anisotropic [BN21, Ler20, MVX20, Per20b]. anisotropy [Ber21b]. Annealed [Van19, Sob12]. annealing [JM22]. Annihilating [TZ11, AS12, BBJJ22, BN15]. annihilation [DFPSL19]. annular [Wer12]. Annulus [Zha06]. Anomalous [BM21, MSC15, BC22]. answer [Lin16]. Anticipating [LMCV12]. antiferromagnetic [GGY18]. any [DP17, DVZ18]. Apollonian [DHS19]. applicable [Völ14]. Application [AI09, CGGR10, Eng10, KL05, Pas11, ST06, AP17, AP21, BB96, Bor08, BPP22, CD12a, CS17, DDT14, DHS19, FM17, MT22, Szn16, Zha13b, LV11]. Applications [BP04, BL05, Dal99, DS04, GMT05, GLW05, Pet08, PM08, Ada08, AW09, BGK^{+}20, BE21, BR12b, BH21, Ber23, Ber15, BM20, BJJK21, Do21, DM14, Fag12, FGL15, FOW22, Fra09, GL23, GS16, Hal22, HMP12, Hei22, Hug19, JT21, KT17, Lê20, Pau14, PS21, QZJ14, SV09, SCZ13, Vel22, Win15, Zha19, RTW12]. Applied [TW11, JK19]. Approach [BdSGP09, BFG99, CMT08, CF11, Dey11, AH21, BBP20, BBFS20, BHJ22, DP13a, DGP09, DMR09b, DLT21, FMNS15, FIP15, GI12, GM15, GIP16, HR19, HP21, HP17, JL15, JS18, KW21, Mar18, Men14a, Nut12, PKTZ15, Shc22, Ste21]. Approaches [SSG11]. approximate [CL21b, Fd98, Lac22]. Approximating [BR13, Kry99]. Approximation [BBM22, BR16, BJKM10, BN11, BC20, EL11, GN03, GW18, PR11, Pen09, Pos05, Vil11, AL17b, BC04, BX19, BCG21, BM20, BM22b, BCG20, DRM14, ET15, Fan19, FF13, FKHL19, GNV22, Gau14, Gol18b, HM22, IJ21, KM17, LP19a, Lou22, MR15, PRR20, PS21, PT13, PS18b, Roe08, Tur15, Vil15, Zha14]. Approximations [Sch05, BCKH22, BL12a, BDD14, BF17, DJQS13, DK21, Kli14, KS17b, Lej22, MT12b, MVJ14, Slo14]. Approximative [FR12, HJvWW20]. arbitrarily [Wan14]. Arbitrary [DGP19, KSZ01, DH21, FO18, MW08]. area [CIW19, JL07, Pro21]. areas [JL07]. argmin [PT18]. arising [BP08, CCSWB22, DPP15, HNX19]. arithmetic [BY09, BAKL19, DNPR19, Maf20, ZZ13]. Arm [LSW01, Hul15, WZ17]. arms [Yao13]. Arratia [Kon17]. Array [Arc97, Arc99]. ary [Ber15, NN21]. ASEP [CT20, DZ22, Mar20, Nej21]. aspects [Ric15]. assassination [Bor08]. asset [BPR15, CCC15]. Associated [BCS06, BG08, HY10, Kry04, Mie01, SS10, BDFR15a, CHK17, DZ15, Jui18, Küh20]. Assortative [EL11]. Assortativity [BJK13]. assumptions [BM19b]. Asymmetric [BS02, GK00, FV12, Gay12, OT19, RR13, SS18b, Wan22]. Asymptotic [AD19, AF03, AH05, BW21, BDD^{+}18, BW05, BdSGP09, Ber06, BG04, BCMP09, BC08b, Cha11, DPU15, DR15a, EL07a, Fér13, FG19, Fuk11, GNV22, Gil11, Gil16, Gro17, GR17, HKN20, HJ15b, HZ19, KK11, Lii11, LSW22, LM07b, MR10, Mél19, ML10, Mun11, NY19, PR05, SSW04, Tan06, Tar15, Uch07, vB09, CR09, CHZ22, CGZ06, Dav98, Gol18b, HT10a, JK19, KM12, Kon17, KS04, RS21a, Tan19]. Asymptotically [FMNS15, BDG08, GMPW19, Pin12]. Asymptotics [AMP04, BK01, BG08, CH10, Dad17, DS11c, DP13c, DSWB15, FF14, Haa17, Kol01, KM05, LM07a, MWZZ02, Wei07, AP16, AB16a, AP21, BST22, BBI09, BF08, CHNT17, CHSS18, CGG22, DR15b, EG15, FH19, FTZ20, Gao17, HS17b, Ina16, KGS13, MN18b, OO15, Pet12, Rob15]. Atoms [BM22a]. attachment [BL12b, DM09, Der16, EM14, FI22, FJ20, Jor13, LO20, LP19b, Sén21]. Attracting [GP20, KK12]. attraction [Pin18]. Attractive [BGRS10, Jan16]. Attractiveness [Bor11]. Attractor [Bes99].

Attractors [DS11a, CC98]. automata [GL22a, MST19]. Autoregressive [BdSGP09, DHKW22, Zer18]. Average [FF98, BY09, FJ12]. average-case [FJ12]. Averaged [RASY17, BEP09]. Averages [Bas08, Pan05, BOHP19, BOPT20, BPZ22, DG15, Kul08]. Averaging [FK21, GNZ21, NZ20]. avoid [DWW19]. Avoiding [AW97, BT19, CMT15, CP16, CSW22, Ham19]. Avoiding-probabilities [AW97].

backbone [BBDS23, BCDG13]. backtracking [LR18]. Backward [APS21, CZ05, HO03, KMP06, LT22, BT20, CY21, DM16, HP21, LRTY20, MB14, Tur15]. Bakry [DP13a, Oça21]. Balanced [DM11, GPT22, JP18]. ballistic [AP16, BS20c, BT19, Bou13, BC22, GR15, RS19, Sch19b]. ballisticity [HS17a, RR22, Tou09]. Balls [Pen09, AFS20, BEGG14, CP21, Lin14]. Banach [ALM21, AN14, DMN13, Kun13, Lê23, Mas07, Sei10]. band [Au18, Shc22]. bandit [LP08]. bandwidth [DE06]. bank [BCKWB20, GdHO22]. banks [Nan23]. Bargmann [GV20]. barrier [Fag12]. barriers [HH06, KR21b]. Based [Doh99, HT10b, BBM20, BPR12, Ber21b, CKS21, DT20, Jan19, JR14, LT15b]. bases [CLT21, TT20]. Baxter [Bor22]. bays [Ang22]. BBS [FG20]. Be [Lam07, CO08]. before [ABOPS16]. behaved [Ber21a]. Behavior [BC08b, BC10, CCR05, CD06, DG03, GKW99, MR10, Tan06, Zho03, BDD^{+}18, BD08, CGZ06, DFS21, Fér13, FG19, HKN20, Kon17, MSC15, OW20, Tak21, Wag06]. Behaviour [Ber06, Don04, Mou01, CR09, CCDO21, DR15a, HZ19, MS21]. Belief [Kha14]. Bellman [SS19b]. Below [BG11, FR13]. beneficial [GPPW16]. Bercovici [CD05a]. Berg [MR13]. Bernoulli [DFF^{+}21, DE09, LM07a, Tan23]. Bernstein [Maj07, PS18a]. Bernstein-gamma [PS18a]. Berry [BHT03, BG13a, BOPT20, BCG18, ET14, PS22b, SSW04, Zha22]. Besicovitch [LM23]. Besicovitch-stability [LM23]. Besov [AF20, CKLL13, dRM22]. Bessel [Ale11, Alt21, CCD13, KL22, Lar15, PY99]. Bessel-like [Ale11]. best [EL14, FR12]. best-response [EL14]. Beta [BBC^{+}05, Chh15, Döb15, FH13, LR10, DMR16b, Duf10, Has14, KSW14, SW14]. Beta-Coalescents [BBC^{+}05]. beta-ensembles [DMR16b]. Beta-Fleming [FH13]. Beta-gamma [Chh15, Duf10]. Between [IT00, Sak09, ADH13, AJKH15, BC14, BCR07, BJN13, BPP22, CMMS23, De 01, Eis05, HMK16, KS19, KPV04, LL21, RSS22, RS12b]. Beurling [LL04]. beyond [BT15a, BG13b, BKM21]. bi [BS22c]. bi-infinite [BS22c]. Biased [BC22, BS22a, BS20c, CHK18, DS12b, Gei00, HS15a, LM19b, Tag16].

Biasing [BBM20]. Bifurcating [BdSGP09]. bifurcation [CTV21]. bifurcators [Shi17]. big [BPZ22]. Biharmonic [Nis01]. Bijection [CD05a, BJM14, Leh22]. bijections [GJ14]. Billiards [BBG19]. Binary [CKMR05, HJ15a]. binomial [Hal22, Pin16a, Roe08]. bins [CP21]. Bipartite [Wei07, BS14, BdT14, Gal07, dHNT18]. birth [Bor08, CSC15, EL16, Vil15]. birth-and-assassination [Bor08]. Birthday [CP99]. Births [Bor11, DPTV12]. bisection [MNS16]. Bismut [BG21b]. Bistable [BBM10, EP22]. bit [BO19, FJ12]. Bivariate [BAKL19]. block [BW21, KL21]. blocks [FF14]. blow [LS21a]. blow-ups [LS21a]. blowup [CHKK19, GLM^{+}17a]. Bochner [DP13a]. bodies [GW18]. Bolthausen [BG08, GM05, Sch12]. Boltzmann [BLM15, Vel22]. bond [Gol18a]. Bondesson [ALM10]. Boolean [DM23, Tyk07]. Bootstrap [BGH^{+}14, GH09, GS17a, Hol06, Bav19, CHSW22, MN21]. Borel [Fra09]. Borodin [FW17]. both [GM08]. Bouchaud [Gay12]. Bound [Vih11, BC04, BG13a, BDvdHS22, Car20, DKRV17, LS14, Mas22, MM15, Nag12, Pin12, Sub13]. Boundaries [HP17, EGW12]. Boundary [AL98, BG11, BK05, Kim06, KW11, Lot01, LW01a, Nis01, WZ17, AM04, Alt21, BBM22, Bia08, BKM21, CJK18, DFPSL19, DGK21, FZ23, FTT14, GL18, GV17, GM15, Hon20, Hug19, KOS21, Lin13, RS21a, Zha21]. Bounded [BP07, CDK06, LMM04, BA12, BG21b, CKLL13]. boundedness [Hut20]. Bounding [MR08, BLM09, DR14, DGV17]. Bounds [BHT03, Bob22, Boi11, Dal08, Fol11, GMT06, Jon11, KN07, ML10, SSW04, Sob12, ABBHK22, AJKH15, BP16, BCG18, CS17, CJK12, DP20, DM23, Dir15, ET14, EHMT17, Fan19, Gol18b, GM14a, JK19, LNO15, LL21, MW20a, OZ21, PS22b, RX13, RR96, Tan19, Zha22, dS14]. branches [Cou18, DM13, Kar20]. Branching [AR06, AI09, BB11a, BG11, Ber97, BBC^{+}05, DG96, DLW01, DEF^{+}02, DG03, DGW04, Eng10, EZ16, EFP17, FZ12, FL18, GGR19, GKW99, HMN11, Jag16, KS03, KM11, Lam07, Su14, Win02, Zäh01, ZZ15, AM15, AD14d, ABK13, AS12, BN15, BMS13, BP08, BM21, BM14, BSW21, BD14b, BGT09, Bon22, BM19b, BH14, CFW09, CHL18, CHK19, DFPSL19, DT20, DKLT22, FMH22, FH13, FMM19, Fou19, GJ17, GLM17b, GSL21, HK18, HKN20, HMT23, HM22, HZ19, HNX19, HL14, IKM21, KGS13, KZ08, LZ15, LW20a, LZ21, LP15, Mal15a, MS21, Mül08, OR17, Pag17, Pai16, PRR20, PRT22, Rap23, Rob15, RS21b, Stu20]. branching-ruin [CHK19]. branching-selection [GSL21]. Branching-stable [ZZ15]. Brascamp [Men14b]. breaking [BDI20, GH15, HMR14, MR09]. Bridge [AMP04, BIK07, CCD13, Hab19, Pit99, PY99, PT15]. Bridges [BW12, CC13, CX13, CM17, FV12, Fit98, LM07a, LPT15]. Brown [Ho22]. Brownian [Gro13, HJ09, AW97, ABK06, AFL23, AMP04, AW14, ABT15, ABPR11, ABK13, AL17b, AFP15, Ata01, BY09, BK18, BR20, BB96, Bas23, BC08a, BZ12, BGK^{+}20, BW21, BDD^{+}18, BM21, BA12, BDR22, Bon22, Bor22, BD13, BH14, BW17, BD03, Bur09, CIW19, CM17, Chh15, CC22b, CMT19, CP20, CH05, DS07a, DeB08, DeB09, DFPSL19, DMN15, DPR03, DS17, DS06, DS11a, DM14, DMP05, Eld14, ERY14, EFP17, ÉM12, FV12, FSW15, FV17, FMW06, Fra10, Gal22, GJW21, GR14, GNV22, GN03, GK07, GL22b, Gro10, Gro11, GB14, Haj11, HH06, HJ07, HK18, Hol08, Hon18, Hon20, Hu00, Hug19, Igl05, Ina16, JL07, JLML10, KM18, KW11, KS17a, LZ15, Law96b, Law00, LL22, Lin13, LZ12, Lou22, LPT15, Lup21, MK12, MS99, Nan06, NRS05, NR17]. Brownian [Nil20, NP08, NRS10, NZ14, NM04, OP22, Pai16, Pas11, PPPY01, Pic06, Pit99, PT15, PT18, QW18, RGV14, Rob15, RS21b, Sar15, ST17a, SS13, SSS09, SS15, Suz19, TOR15, Tsi06, TV03, Ven19, War07, Web16, Wei09, XKW06, Zam08, Zho03, vdHS10, vSW18]. BSDE [ALR23, BCC19, PPS18]. BSDEs [AID07, CB08, DQS16, FJX11, GS16, HH06, HR19, KTPZ15, Kli12, KR21b, Luo20, PKTZ15, RM13]. Bucy [DKT18]. Budd [BJM14]. Bulk [JV11, TV11, EY12]. Burgers [Bak16, BGS16, Sep96]. burglar [Dum12]. burning [GJ14].

càdlàg [ŁOPS21]. Cahn [BDW17, BOS22, BSW10, EGL22, HRW12, OWW14]. Cahn-type [BOS22]. Calculus [EP98, KMP06, NV09, Pet08, BM06, ČR22, Hal22]. Can [LM19a, Vih11, CO08]. Cannings [BCPW21]. canonical [BR22, LT15b, YP12, Yan22]. Capacity [XZ10a, AS20]. card [Spi22, Tra15]. Carlo [ADW21, BDD09, BJJK21, DR14, JK19]. cascading [FG19]. Case [BK05, IT00, KN07, Led04, Led05, OW04, Sza10, Tel01, AD14b, AD14a, Ban18, BK18, BCPW21, BKM21, FSW15, FJ12, FK21, GI12, GLT21, KS17a, KM15, Mic17, OO21, RS21a, SS18a, Zha21]. cases [GBM22]. Catalyst [FMW06]. Catalysts [KS03, GdHM07, GdHM09]. Catalytic [DEF^{+}02, GKW99, BP08, SF06]. catastrophes [BMS13, BARS19, MS21]. Cauchy [Ber97, Hu00, KR20, MP05]. Causal [HS14, Bud19]. cavity [Che13b]. Cayley [BT15a, HKLR19]. cellular [GL22a, MST19]. Center [Eng10]. centered [Roe08]. Central [BL12a, BRW20, BLO13, BGG14, Che13b, DS12b, IKN20, Jan22, JT21, MR12, BR18a, BDD14, BST22, CT22a, CD15, DMR09a, DLR19, HI23, HMMN15, HJ13, KT17, Lam21b, Mou12, NXZ22, Owa17, RSSZ19, SYY21, AP23, BJ09, BDD09, BBL11, Far11, HJ11, Lev13, NY09, Pan05, PP11]. centrality [BH23]. Certain [Cho05, GN03, KST02, Lyn00, MWZZ02]. Chain [BDD09, Gne07, DR14, DMR09b, ES19, GM22, KOS21, MVJ14, Xu20]. chaining [Dir15]. Chains [BP04, Fau02, FM04, FS01, JW01, JN10, PR11, SV05, SW11, Ada08, BJN13, BCC^{+}22, BC18, CDM12, CSC15, CL13, DDT14, DKN18, DS07b, EM19, EHMT17, GMPW19, HMP12, He22, LLM18, MR15, Oli12, Pau15, PSSS17, RLR22, SCZ09, Shk13, Tow15, Web21, Wor99, Zer18]. changed [DS17, DM14]. Changes [Ste99, CHK17]. Chaos [CE13, Ada20, Ber23, CP21, CC22a, Die20, Dje22, DM14, ELL21, GBM22, JS17, LT22, LM07b, NNP13, Pow18, Web15]. chaoses [ALM21]. Chaotic [BD13]. Characterisation [DG09]. Characteristic [Sch19a, Cos23, Web15]. Characteristics [Zei10, Hol12]. Characterization [DLX20, Fer16, IMP07, Kuw09, Ber21a, HL07, Kal21]. characterizations [Fuk99]. characterized [Pit99]. Characterizing [CN18]. Chavel [Pas11]. chemical [DHS17, EP23, GK14]. Chinese [RLR22]. choice [FR12]. Choices [Jun15]. chordal [FZ23]. Chung [Hu00]. CIR [Han14]. circles [Cha15]. Circulant [BHS09, LŚ22a]. Circular [Ada11, BCZ18, CES21, Lam21b, Web16, XYY17]. claims [CB09, NN13]. Clark [PT13]. Class [ALM10, AdHR11, BV04, BDD09, DKM01, Doh99, JN10, Kli11, LL04, NY09, NY10, BDG08, CD06, CL14, Dal21, EKYY13, GLM^{+}17a, Hab19, Han14, HNX19, KKO13, Kun13, Ler20, LZ21, Sam09, SS19b, SYY21, SS15]. Classes [Ada11, CO08, BBFS20, Kho96, MST19, Tsi06]. Classical [AID07, DS11b, Led04, Led05, GM22]. classification [BFMM13]. CLE [WW13]. clever [Dum12]. Clique [DGLU11]. clock [Gay12]. clocks [TY23]. Close [KSZ01]. Close-Packed [KSZ01]. closed [BBFS20]. cloud [DFS15]. CLT [AM15, AL13, BWZ09, BCP18, BP07, DDT14, DS11b, GP12, IV13, QZJ14, Yao13]. Cluster [BP06, CFT15, Fag10, FG96, She10, Ber14, BBDS23, BCDG13, CCK22, Cha15, DH15, DCLM18, HS22, Ric18]. Cluster-Formation [FG96]. Clustering [BY13, FvdHMS21, Pin22, Zho03, BJK13, BP14b]. Clusters [HH05, HLP10, BBAP06, Fag08, HB09, Hut23, Li14]. Coagulation [MWZZ02, Wag06]. Coalescence [Ber04, SS20, SS21]. Coalescences [FMM19]. Coalescent [AL98, BG08, BLS18, JK11, Lim10, Mie01, TV09, BCKWB20, FJS15, FRLS20, Guf18b, KSW14, KL21, LS06, MR17a, Sch12]. Coalescents [BP99, BBC^{+}05, GY07, Sch00, BDLSJ18, CCSWB22, DM13, LT15c]. Coalescing [NRS05, TZ11, AS12, Bel20a]. Coalesent [GM05]. Coarse [Ton09]. coarsening [DKNS13]. Coding [DS06, Spi20]. codings [HS22]. coefficient [JK19, KR21b, KN21]. Coefficients [CN00, Kim05, Zha11, BEP09, BR12b, BDP22, CL14, CW23, Do21, Dur14, Eto06, FS17, HKY14, Hua21, KM17, KS17b, Nil20, Rom18, SWY12, Tre16, YP12, Yan22]. Coexistence [CMP10, DH06]. coin [EVW20]. coin-turning [EVW20]. Collapse [Vih11]. collector [DP13c]. Colliding [KOR01, ABOPS16, FV12, GM14b]. Collision [BGdH11, EP98, PR05]. Collisions [CC11, DGP18, GY07, HH22, Sch00, DM13, FSW15, Sar15]. color [FS20b, Zha14]. Colored [Stu03, NXZ22, RS13b, TTV14]. colourings [Gal07]. column [RLR22]. Comb [Ber06, CC11, CCFR09, HM14, HS12]. combination [KS20]. Combinatorial [BJ09, BN11]. Coming [Lim10]. comment [EY12]. Common [Pet10, Tay07, DH21, HI23, LS21a]. Compact [DPR03, DS11b, ADH13, Ful12, Kli14, KL18, LZ12, Mél19]. Compactness [MM20, MM18]. Comparing [OO21, Sim14]. Comparison [DMS09, GHT03, HKL^{+}98, PM08, DMR16a, Do21, Hei22, Hut12]. comparisons [CK20, FJ12]. compensated [SW19]. Compensator [IMP07]. Competing [Arg08, DMP05, EP98, FM03, Kli11, Shk09, CEY12, Sar15, ST17a]. Competition [KPV04, BvdHK15, BSW21, Fou19, GM08]. Complement [DeB08]. Complete [BW05, FM04, FM14, BKLM19, Che14, EGvdHN20, Fox18, Gay12, GS17a, Kha14, KR08, NN21]. Completely [CLT21, JK15, Kwa19, Naj20]. completeness [CL14]. Complex [RV07, Su06, DNPR19, FvdHMS21, HHN16, HK18, HIN17]. component [CP23, Gol18a, LT15c]. Components [Win08, BDvdHS20, BJN13, FR13, FG20, Hay21, Tyk07]. composition [MN18b]. compositions [GI12, HTV22]. Compound [BJKM10]. compression [BO19]. compulsive [ALS15]. Computable [RR22]. Computation [BM06, Del03]. Computing [CSC15]. Concatenation [Wer21, GGR19]. Concave [Gro11, Fra09, Gro13, OP22, Str19]. Concavity [Law00]. Concentration [BP16, BBGT20, Fra09, FLM20, GSS21, HX08, KMP06, NV09, Pau15, RC09, Var19, AKPS19, AZ13, ACdCdH19, BLM09, CHZ22, DHS14a, DKT18, GSS19, GM14a, HT21, JG13, KZ20, SZ19, TC14, Zha18]. Concrete [MS98]. condensate [Der16]. Condensation [FI22, AD14a, FJ20, GRV13]. condensing [CGG22]. Condition [BB11b, Sch06, AP17, BBM22, BD12, BV12, DMR09b, Hut20, MX08, RM13]. Conditional [Bry96, DW10, ELL21, EL07b, KS17a, DEM13, FF14, Kur98, LS21a]. Conditioned [JW01, Lam07, AD14b, AD14a, BDR21, CC13, CR08, CD05b, CD08a, CC22b, CK08, DVW14, DWW19, FJ22, FJ23, PRU20, Shi12]. conditioning [GT07]. Conditions [CC11, HS17a, Nis01, WSH09, APS21, AM04, Alt21, AL20, Bia08, BJJK21, CJK18, FR18, GL18, Lin13, NN23, SX19, Suz19, Tur15, Web21]. conductance [DF20, Fle18]. Conductances [BP07, Bou10, Čer11, Fag10, BS20c, Bis19, Fag08, GP12, GN21, Mou12]. Cone [Zha09, BO13, DW19, GR14]. cones [BPW18, DMZ08]. configuration [BvdHK15, BDvdHS22, CS15, DvdHvLS17, DMS17, EFMN22]. configurations [DLR20]. Confinement [CIW19, Sak09]. confining [DLR20]. conformal [AHS23, CD12a]. Conjecture [HL08, Pas11, TV11, CHHM23, EY12, HJ19]. conjugate [KR08]. Connection [Eis05, CT22a, GHMM22, IJ21, MM21]. Connections [FM04, Ott22]. Connectivity [BC10, AS18, BH19, FR13, IJ21, RS13a]. consecutive [Naj20, Pin22]. consensus [HH14]. Conservative [EV22, Kra13, Zam08]. conserving [ES19]. Consistency [MNS16, BJN13, DE06]. Consistent [CGR21, FR18, Rob15]. consisting [DLR20]. Constant [Kry99, Zha09, AT16, Fou08, GT17, JK19, RT18, Seg09]. Constants [RT10, BDD14, GMPT17]. Constrained [AB02, BD03, CK19, Ko20]. Constraints [KW10, AKTKK19, SP14]. constructed [Haa17, RW18]. Constructing [Lej22]. construction [CCD13, FPRW20, GH15, Guf18a, HN22, Mar18, Mic20]. Constructions [ABBG10, DW19]. Constructive [LP14]. Contact [PZ22, Val10, vdHS04, vdHS10, AEGR15, CS15, DLZ14, DY20, HUVV22, Hua20, LR20, MVY13, MV15, MV16, SS19a, SS14, Su14]. containment [Kli14]. content [FLM20, PS22a]. contestants [AS13]. Contiguity [Ban18]. Continuation [Fd01, Ang22]. Continuity [CPZ22, GMPT17, LW20b, Tur07, BV16, CG18, Cri21, DJQS13, HS20, HKY14, HHNT15, HNX19, LS22b, MM20, NNP13, RT18, SSS13]. Continuous [Ber97, Fol11, Fou19, Lam07, Led04, MP07, Zho03, AD20, BS20a, BMS13, Bus16, CB09, CGR06, CW23, CMMS23, DEM13, FH13, FMM19, GHM22, GT17, Hua20, Lac22, LW20a, LZ21, MS21, MR09, OZ21, RR96, VY16, Web21]. Continuous-Sites [Zho03]. Continuous-State [Lam07, Fou19, FH13, FMM19, LW20a, LZ21, MS21]. continuous-time [CGR06, CMMS23, Web21]. Continuum [ADV10, BEV10, CD14, HM04, Jan16, AS18, BW17, DH15, KR21c, Tyk07]. contour [SS18a, SS19b]. contraction [EM19]. contractions [LRP13]. contractive [Mon15, PR20, Zer18]. Control [BM05, DLT21, OW04, BCC19, DZ15, Dje22, Nut12]. controllability [LO21]. Controlled [Kry99, KS01]. converge [AFL23, CM17, Che19]. Convergence [AEGR15, Arc99, BHPZ11, BCP18, BJQS10, BMSV06, BR11, BL05, BM20, BGS16, Boi11, Car21, Cho05, CT07, CL13, CHK12, Gay12, GBM22, Hol08, HL14, JW01, JPP19, JM22, Kli11, KW00, Lej02, LR11, NRS05, Oça21, OC19, Pel99, SW14, SW11, Stu03, Su01, Suz19, vdHS10, Ada20, AdlRG21, AZ13, BCP20, BP13a, BBGH12, Bel20a, Ben21, BF20, BJJK21, BDP15, CX13, CW21a, Cos23, CP20, DMR09a, DDM12, Gan13, Gia08, GT07, HBS17, Hut12, Jal21, JP18, KOP13, LLM18, LM15, Mon15, NNP13, PR20, PRW14, Pow18, RLR22, RR96, YP12, Nan23]. convergent [INS21, IMN23]. Converging [Lej02]. Convex [CSŠ22, KMP06, Men07, AGL^{+}12, AKMV20, BBJJ22, Bob07, DMZ08, Eld14, FLM20, GW18, GBM22, MW18, Pau14, ST22, Tug12]. Convolution [CDMFF11, AN14, BHM13, KP22]. Convolutions [Sei10]. cookie [BS08a, KP16]. cooling [ACdCdH19]. copies [BBGH12]. copolymer [BP15]. copolymers [Ton09]. copulae [BJN13]. copy [PW23]. core [dHNT18]. Corner [BCS06, EJS21, SS20, SS21]. Corrections [CD08a, OPR22]. corrector [LNO15]. Correlated [BV04, FK02, MCS03, Aco14, AC19, BP15, Che17a, FJS15, HHN16]. Correlation [CJK12, MS13, Ton07, BPZ12, FW17, Gia08, Hei22, MN18a, YP12, Yan22]. Correlation-length [CJK12]. Correlations [BFG99, GS11, BST16, CGL18, DPTV12]. Corrigendum [BK19, FJ23]. Cost [DDS18, CT22a, JS22, Rie17]. Costs [BT05, CB09]. Coulomb [Ame21, Ber21a]. Countable [BP04, FM07, HKLR19, Tsi06]. countably [BL12a, TOR15]. counterexample [Hei13]. counterexamples [KK20a]. counterparts [BN14]. counting [Owa17]. counts [BG09]. Coupled [BBM10, Lyn00, BF22, DLX20, HJT20]. Coupling [BK18, BFG99, JM15, KP04, Pas11, Szn16, vR04, BCC^{+}22, DMR09b, HN22, Jui18, LST19, RC09]. couplings [BG21b, JM20, Kuw09, Pau15, RS12b]. coupon [DP13c]. Covariance [Su06, Unt19, Vih11, AEK^{+}14, CYY22, Gro17, Men14b, PT13, Yan19, You13]. Cover [DPR03, CGPV13, Din12, HW10, Zha18]. covering [IKN20, JJK^{+}17, PZ20]. Cramér [HSW09, AP17, BB11b, Naj01, PS19]. Creeps [GM11]. crevices [EJS21]. Criteria [Sha15, CV23, DMR09a, DDM12, KK20a, Kra13, Oça21, RR22]. criterion [FM17, Mül08]. Critical [ABBG10, BT10, BvdHvL10, DvdHvLS17, Hol08, Jon01, KSZ98, KM11, LZ15, LSW01, MW05, MVX20, PR11, Pow18, RT10, vdHS04, vdHS10, AO13, AJL14, AA23, BDF18, BFGM19, BP15, BPT17, BDvdHS20, BDvdHS22, CPZ22, CDCH16, CS20, CDC13, Cip13, CD12b, CHK19, DHS17, GS17b, JW18, JS17, KS19, KS17a, KM15, KL21, Lin14, LP14, Max06, Nol08, OW20, Rap23, Red17, Ric15, RM09, SS19a, Spi20, Tan23, Ton09, Yao13]. criticality [DMM17, RT09]. critically [LM19b]. Crossing [BC08b, CDCH16, Sav08, LW21, Ort12]. crossings [AL13, ŁG14, Wer12]. CRT [GMS19]. Crump [SS18a, SS19b]. Crump-Mode-Jagers [SS18a]. crystal [Eza13, Ler20]. CSBP [DL14]. Cube [BCS06]. cubic [AFL23, OPR22, STV22]. cumulants [LRP13]. Curie [BBI09, BMP21, CK18a, FKK22]. curse [HJvWW20]. curtain [HN22, Jui18]. Curvature [Suz19, BG21b, EHMT17, EFP17, Web21]. curvature-dimension [Web21]. Curve [AS08, Sav08, BR13, Zha21]. Curves [GK00, BP15, CW21a, TY20]. Curvilinear [Fd06]. Cusp [DJM16]. Cusp-Airy [DJM16]. cusps [CK18b]. Cut [BP20, Law96a, BW17, Hou17, Law96b]. Cut-off [BP20, Hou17]. Cutoff [CSC08, DKN18, LS11b, BHP21, CSC15, HP18]. cutoffs [HLP16]. cuts [DMS17]. Cutting [Dev11]. Cycle [BU11, Mer08, Hou17, Muk16, SZ15b]. Cycles [Ber11, Jon11, GS17a]. cyclic [MN18b, PT07]. cylinder [BBGT20, BST22, Szn09]. cylinders [BT15b, HLP19, Ven19]. cylindric [CMT19]. Cylindrical [VY16, BR22, CHZ20, FM21, KR20, Ram22].

d [KN17, AS18, BS20c, CHNR22, DKNS13, Din12, Fag12, Fer16, FR01, GL18, HRW12, LSW01, Maf20, Oda06, PYZ22, SY14, SBB^{+}21, Unt19]. d=3 [BT10]. Damped [AF03, LM96, Gro20]. Data [Cha11, HL08, LS11a, BO19, BPR12, CG20, RS12a]. Dawson [CR08]. death [CSC15, EL16, Vil15]. Deaths [Bor11, DPTV12]. Decay [BST16, BFG99, RA08, DP13a, Gia08, GR15, KP21, Ott22]. decaying [KN17]. Decomposability [Bou11]. Decomposable [HY10, AH19b, AH19a]. Decomposition [EG18, Ada20, ACJ21, BPT16, DH13, RSZ18]. Decompositions [Eis19, DW16b, FL18, KOP13]. decorated [BBFS20]. decorrelation [HMP12]. Decoupling [AS19]. Deep [Kyp16]. default [HJvWW20]. Defaults [LQ11]. defect [AY15]. defects [HMR14]. Deffuant [HH14]. deficiency [DPWZ14]. Defined [Ram06, BS21]. Deformations [BGGM11, CDMFF11, Mai07]. deformed [McK21]. Degenerate [BP04, BP08, Kry99, Kry04, OW04, Puh04, Zha13b, ADS16b, BBM22, BA12, BGS16, CCDO21, CD15, DP14, Nag12, SZ15a, Tan13]. Degenerating [Kim06, Lot01]. Degree [BM11, BH23, BYY07, DM09, BL12b, CS15, CP23, Lod22]. Degrees [vdHHZ07, BvdHK15, BC17, DvdHvLS17, Jan09, MW08]. Delaunay [GKT22, AE19, DG09, Rou15]. Delay [KW10, NNT08, DZ15]. deletion [HS13, MR17a]. denaturation [Leg21]. Dense [CHKK19, Ban18, BI20, KR21a, Tsi06, dHMRS18]. Densities [DLW01, DS11c, KK11, HKY14, KP22, Mat23, MVJ14, MVY13, NNP13, Sch09]. Density [BK01, BCMP09, BBL11, BFMM13, CMN97, FH23, NV09, Son06, BQSS19, BC14, BCKH22, BMM20, CMN98, CO18, DE06, FO18, FL13, GM08, Gup12, JL15, MN08, Rom18, SZ15a, SW14]. Dependence [Tan06, ARS08, BX19, BJN13, DW18b, Fan19, IT99, Kif17, Kli14, MT12b]. dependency [Fér18]. Dependent [Ada11, BN15, BB10, BHS09, Bry96, DLW01, DG03, EL10, HL08, Pel99, Sch05, BPR12, BSW21, BO13, BV12, Che17b, CKX09, Dal21, DH13, Del09, DÜ22, GHM22, GSS19, GHMM22, GSL21, Hua21, Kal21, Kul08, Lev17, Pau14, Pes22, Pin16b, Roe08, Sch09, SHW20, Str19, Wan14]. Depinning [CP09]. deposition [ACSS16]. Depth [EFMN22]. Derivative [HW21, BG21b, GS16, Pow18]. Derivatives [BDR22, vB09, DGV17, HJvWW20]. Derived [BP99, CP99, PY96, EI16]. descendants [KW21]. Descent [CZ05, Kam18, Tar15]. Descents [Ful05]. describing [For22]. Detecting [BP13b, MP07, Pin13]. determinant [BM19a]. Determinantal [RV07, BO13, MR13]. determinants [BW21]. determination [Lin16]. Determined [SS00, QW18]. Determinism [XKW06]. Deterministic [AR06, ADW21, BN14, BBCC21, CMMS23, CHT19, CHNR18, Die20, He22, LP13, MS15, MT22, RTW12, XYY17]. Deviation [BM09a, MCS03, Puh04, Wan05, AL09, BCV21, CK18a, DZ22, FF13, FF15, GPP13, Tan19]. Deviations [AS20, BD03, CLR05, DE09, Fau02, Fen07, GMT05, GHT03, Gau07, HM01, KM05, LR10, LLS06, LS99, LC05, Mat11, Tan06, Vos08, APW20, Aug16, BBP20, BB11a, BR20, BCR06, Ber19, Bob07, BDG18, BD18, But16, CD19, CR14, CGPV13, Cor21, DM14, DLR20, FS20a, FH19, GW09, GLM17b, Gro17, GM20, GM22, HSW09, HS17b, Hus22, Kra20, Lau12, LP15, Mai07, McK21, MT22, Mül17, OT19, Orr20, Ort12, OPR22, Par20, Per21, Pet12, PO09, RASY17, SCZ13, ST22, Spi15, Ver19, Wor99, Zha20]. diagram [HK18]. diagrams [FSX20, Hor20]. diameter [DFS15, MW18]. Dickman [BN11, BM20, CSZ19, Gol18b, Pin18]. dictionary [DS12a]. Did [PPPY01]. Difference [IT00, LM96, CO21]. different [BBGH12, GM08]. Differentiability [And11, AID07, Bur09, DS11c, IdRS19, SSS13]. differentiable [Tre16]. Differential [AW09, BP04, Beg08, BO10, HO03, KS97, KKH11, KW10, Led04, LM96, MP05, PR20, Rho10, SS10, Var19, APS21, AM04, ADMU20, BCD20, BWY15, BKN19, BP08, Bas14, BZ12, BMBP18, BS17, CL14, Che19, CY21, CJK18, Dal18, FK09, For22, GV17, HS14, Ina16, KL12, KN21, Lej09, Lej22, LMCV12, Lin13, MT12a, MB14, Nut12, OZ21, SWY12, Tur15, WYZ20, ZRLB19, Zha13b, dRM22]. diffraction [MT12a]. diffusing [GL22b]. Diffusion [AB02, BBJJ22, BK01, BDF18, BS02, Dar01, EL11, FK02, Kol01, Kry99, Kry04, LT15c, LS99, MR10, Tay07, Zha11, AJKH15, AD15, ACSS16, AH21, BR12b, DG20b, EZ16, Eto06, FJS15, Gap06, GV17, HKN20, JR14, JK19, KK20b, KS20, KS17b, Kuw09, LZ13, Mar18, MW20a, Pin16b, Sal22, Sha15, Spi15, Tre16]. Diffusion-limited [BBJJ22]. Diffusions [BB10, BR11, FG96, FPRW20, GLW05, KP04, KSX04, Ram06, Sch06, Vil11, Vos08, BCC19, BT20, BW21, BC21, BD19, BPP22, BDG18, CCDO21, CR08, CD15, CK18b, Cri21, CLSF19, Feh17, FS20a, FTT14, FMH22, GPP13, Hab19, Hob15, HS17b, Hut12, JT21, KM17, KSX17, LZ23, LSF20, ŁG14, Mül17, ORS20, PR23, PM08, Pro12, Rie17, Slo14, SF06, Tug12]. Diffusive [CCR05, HMN11, AG13, Bel20a, HS23, Pin19]. digit [MS12]. dilations [Fra09]. dilute [BB18b, BMP21]. Dimension [AS08, AMP10, BT10, Bou10, CLR05, CB08, DHP11, Hol08, Suz19, vdHS04, vdHS10, BTW22, BDW17, CCD13, CCK22, CGZ06, Cip13, DP20, DP17, DVZ18, GKKR18, Gol18a, Hut23, Law96b, LL22, Lou22, LP14, NRS17, Rap23, Rou18, SS13, Seg09, Web21, Zha13a]. Dimensional [AC96, BRS10, BB10, Čer11, CMP10, CCFR09, Del03, DGLU11, FM04, FV11, GZ10, Gra96, Nis01, Pet10, Pos05, Ros97, Sak09, TZ11, Uch10, Vos08, AP16, AKMV20, AGT22, AZ15b, AT16, BCC19, BK18, BP12, BH21, BHTW22, Ber06, Bis19, BS08b, COPV05, CDM12, CN18, CYY22, CGPV13, DP13b, DFS21, Dre08, DR15b, Eld14, Eto06, Fou08, Gan22, GdHM09, GL22a, Gro20, Hug19, HS23, JM20, KS17a, KS20, KP21, Kwa12, Lam21a, LZ13, LLM18, Ler20, LS22b, LM19b, MT12a, Max06, MR08, Mic17, MB14, Muc21, NN23, OW04, Pet12, PO09, QZJ14, RTW12, SS19a, SYY21, Val07, Yam23, Yan22]. dimensionality [HJvWW20]. Dimensions [BR07, CD08b, FZ07, Hol06, BHMK21, BGH22, BK16, BK19, BHdS^{+}19, Gan13, Hei13, JLS13, Lau12, LS19, MR09, Nol08, Pha15, Ros23, SSS13, VY16]. dimers [BdT14, Li14]. Dini [CW23]. Diophantine [LR22]. diploid [BLS18, PW23]. Dirac [dT21]. direct [KK20b]. Directed [AY15, BCDG13, CMT19, Kur08, Zha09, AO13, AZ13, ARS08, BS20a, Bat18, Bat21, BGH22, BTW22, BBDS23, CP16, CN17, FSV14, LO21, Nic21, VT08]. direction [GR17, Jos09]. directional [KP22]. Directions [PM08]. Dirichlet [Alt21, AZ15b, AH05, BU11, Bou13, BC20, CDM12, CGG22, Cho05, DZ21, Fen07, Fle18, Foo09, KM18, MWZZ02, MP05, RGV14, SE06, Sch19a, Sha11, Tou09]. Disagreement [Chr19, Jon01]. disasters [GJ17]. Disastrous [Mou01]. Disconnection [Nit18, LS14]. Discontinuity [BC08b]. Discontinuous [Kim05, AP12, BEP09, Eto06, GL18, Nil20]. Discounted [Gap06, INS21, IMN23]. Discrepancy [DR14, Su01]. Discrete [BR06, BT05, Bou11, CDK06, DKM01, Dey11, KOR01, Led05, Lou22, PY96, Sak09, Win08, AZ15b, Ben21, Cha15, CCO20, JK14, KS04, Nic21, Nit18, PY17, SZ21, Szn09, Tur15]. Discrete-Time [CDK06, SZ21]. Discretionary [ZRLB19]. discretization [DM22, MT12a]. discretized [Zha19]. disintegration [Pin09]. disjoint [BGH22]. disks [AHS23]. disorder [BB18b, CD12b, CN17, FSV14, Leg21, Nic21, RS19]. disordered [AO13, CSZ19, ES19, Leg21]. dispersal [For22]. dispersion [DG14]. displacement [Mal15a, Rob15]. Dissipative [Bes99]. dissipativity [Sal22]. Distance [Boi11, Sch05, Sch09, ADH13, BR16, BD08, BPP22, ČP12, DHS17, For19, Pau14, RX13, SZ19]. Distances [DMM17, KS19, Lii11, vdHHZ07, BC14, DMR09a]. distinct [Zhu22]. Distorted [LL22]. distributed [BCP18, IPR09, OO15, SR13, Zem08]. Distribution [BM11, BN11, BHS09, BP06, CDMFF11, FR13, Fen07, GHM22, KK11, LC10, Lii11, SS12, WY10, Win02, AZ15a, BCP18, BP21, Bon22, CYY22, CK08, DRM14, DH21, Döb15, FO18, FH19, GNV22, HC20, HJ15b, Hua21, Jag16, JPP19, KM12, Lin14, LP15, MdBL09, PRT22, PY17, Ram22, Red17, RZ14, RS21b, SS18a, Sch17a, Vil15, Wan20, YP12, Yan22, Zhu22]. Distributional [ABBG10, BS01, CF11, KPS12, Gol18b]. Distributions [ALM10, CP99, CGGR10, FM07, KW10, Lac03, Lam07, Led04, Led05, LZ10, PY96, SSW04, ST06, Sza10, Adh22, Au18, BSV20, DMZ08, DEM13, DR15a, Duf10, DLR20, EG18, GPP13, Has14, Hor20, HW21, KS17a, Kur98, LLM18, LS21b, LM07b, Mar20, Mat23, Pin18, Pin22, ST17a]. disturbance [Bel20a]. Divergence [KKH11, CW23, WYZ20]. diversity [For22]. Divide [FS20b]. Divisibility [SSG11, GGR19, Has14, RM09]. Divisible [ALM10, ST06, WY10, BOPP21, DEM13, Eis19, Mat23]. DLA [JLS13, PRZ20]. DNA [Leg21]. do [FR01]. Dobrushin [SV05]. Doeblin [BHMM19]. Domain [And11, DeB09, Kim06, Lot01, EGL22, Pin18, QW18]. Domains [KKH11, Kim05, Men07, AGT22, Bas96, Ben21, CKS12, CKLL13, DHS14b, Feh17, Lou22, Ram22]. Dominance [HT10b]. Domination [War10, HM14, Pin06, Pin16a, Szn09]. Domino [Nor10]. Doob [BPT16]. dormancy [CCSWB22]. Double [FS17, BN15]. Doubly [JN10, CDC13, DQS16]. Down [FW08, Lim10, EGW12, RLR22]. drainage [ARS08]. Drift [ABPR11, Gro11, JLML10, Vil11, Vos08, AP12, BMBP18, BFGM19, BD14a, BBP12, BPP22, DG20a, DVW14, Dre08, FIP15, GR14, GMPW19, Gro13, HKY14, KL22, KN21, LM19a, MV08, PR20, Pin19, PS19, Whi07, ZRLB19, dRM22]. Drifted [DS07a]. Drifts [Zha11, IdRS19]. Driven [BJQS10, Gau07, Hau05, Lac07, Lej02, AM04, ADMU20, BCKH22, BP20, BOHP19, BZ12, BK08, BH21, BHM13, CHNT17, Che19, CHZ20, DH19, DS17, EG15, GHM22, GNV22, GS16, HH06, HS15b, Ina16, KS17b, KR20, LMCV12, Lin13, MPX22, NNT08, NXZ22, PR20, Rie17, SCW22, SZ15a, Wan22, dRM22, dS14]. driver [APS21]. drop [Rou18]. Drunkard [Su01]. dual [Coh12, Mic20].

Duality [BP09, Mic17, AD20, AC17, ACR21, BFGM19, CGR21, Fou19, GdHO22]. duals [FM20]. Dunkl [Vel22]. duration [LST21a]. Dying [Kol01]. Dynamic [AdHR11, BW05, CDK06, BCC19, BV16, BČD16, CCD13, HH22, HUVV22, LR20, RASY17, RSS20]. Dynamical [CGL18, KS97, Nej21, AH08, Bis19, CHNR22, CFT15, EX22, GM14a, HMP12, Hor20, MT22]. Dynamics [BFG99, CCR05, CMSS16, DP05, FT11, Fra10, GRV13, Pet11, Sch12, AM19, AB20, AB13, BN21, BBM20, Bia08, BGS16, Bor14, BO13, BMP21, CP21, CD12b, DKNS13, DSB19, EGvdHN20, EL14, FG19, HMN23, KKO13, OC19, STV22, TT22, TN22, Unt19, Völ14, WW09, dBD09, dHNT12, dHNT18]. Dynkin [DG18, DQS16, HH06]. Dyson [EY12, TV11, Lup21, War07, Web16, vSW18].

East [CM22]. Edge [AC19, HW10, LZ10, MR08, Yan19, AEGR15, AA23, Car20, FV17, HHN16, Hol18, Nak22, Tak21]. Edge-reinforced [MR08, AA23, Tak21]. edge-traversal [Nak22]. edges [HUVV22, LR20, SS18a, VT08]. Edgeworth [Fuk11, AP17, HRW07]. Effect [Hor20, CSW22, For19, GL18, RS13a]. effects [EGL22]. Efficient [HM22]. eigenfunctions [Maf20]. Eigenvalue [BB96, BP21, GK00, Aug16, BA12, GM20, Hus22, Kar15, Mai07, SW14]. Eigenvalues [BGGM11, CDMFF11, Eva02, JN06, JN07, LV11, BPZ12, BBCC21, DMR16b, Dub18, LMS22, LSW12, LL21, McK21, VS15]. Eigenvector [BHY17, BL22, Dub21, Fle18]. Eigenvectors [LO21, BGG14, CDM12]. element [Ho22]. Elementary [GA08]. elements [DS12a, Ful12, PW23]. Elliptic [AAC09, CB08, CN00, LRMT11, CO22, CW23, Ho22, HS17a, OR14, RR22, RS12a]. elliptical [DFS15]. Elworthy [BG21b]. Embeddable [Dav10]. embedded [CO08, PT15]. embedding [FIP15, Hob15]. embeddings [BT21, BG16]. Emergence [Ber11]. Émery [DP13a, Oça21]. Empirical [Arc97, Arc99, Boi11, Luç11, VS15, Ada08, BPR12, BC18, But16, DRM14, DDT14, FS20a, JPP19, JT21, LM15, LP15, Red17, YP12]. Emprirical [BD03]. empty [GS17b]. encounters [PRU20]. endogeny [RSS22]. endomorphisms [Lev13]. endpoints [BGH22]. Ends [LMS08]. Energy [Kur08, BLM15, BS22b, DLR20, ES19, HK18, Ko20, Yan23]. Engelbert [Kur07]. enlargement [JL15]. enlargements [GJW21]. Enriched [Fd06, Die20]. Ensemble [JV11, dHMRS18, AZ15a, Web16, vSW18]. Ensembles [KOR01, Led04, Led05, LR10, AG13, CCO20, DFF^{+}21, DMR16b, Dub21, DLV21, FKK22, FW17, GM22, JS18, Lam18, Mai07, SW14]. Entanglement [Cou22, GH10]. Entrance [AL98]. Entries [Ada11, HX08, Che17a, CES21, KZ20, Nem17, SR13, Str19]. Entropic [CCGR22, CE13, Mon15, Nit18]. entropies [BDFR15a]. Entropy [DP13a, DW16a, DS06, FF13, Gil11, Goz12, Web21, Win10, BCC^{+}22, CGR06, Gil16, HJ19, KZ20, PL16, RASY17, RS12a, Sam17, TC14]. Entropy-information [Web21]. Envelope [CM06]. Environment [BB11a, Far11, HMN11, MR08, Mou01, Pet10, SE06, Sch06, ST06, AP16, AY15, AD15, And21, ACdCdH19, BCM19, BP15, BTW22, BV16, Bou13, BDG18, BD18, BDI^{+}19, BDI20, DP13b, DGP19, DFS21, Feh17, For19, GPV09, GLM17b, GPT22, Hei13, HS17a, HNX19, HL14, HS23, Jos09, KSV22, KS17a, KGS13, MS15, MRS16, Mül08, Mül17, Nan23, OS20, OR17, RR22, RASY17, RSS20, Sch19b, Tou09, Tra18]. Environments [AdHR11, KSX04, Mat11, Stu03, BS15, BHTW22, BČD16, DGP18, FZ12, Gay12, GR17, HH22, HZ19, KSX17, MX07, PR12, Spi15]. epidemic [BR13, Par20]. epidemics [DY22, GL22b]. Equation [AR06, AF03, BJQS10, Bes99, BM09b, BSW10, CD08b, HO03, Lac07, LC10, MCS03, SS10, Stu03, AB13, Bak16, BQSS19, Bas96, Bas14, BZ12, BR12b, Bel20b, BDW17, BGS16, CHNR22, Che19, CK20, CY21, CK22, CN17, CG20, DW18a, DMN15, DLR19, DJQS13, DM22, EG15, EGL22, Fd98, Gro20, GNZ21, GBM22, HRW12, HMN23, HIN17, Ina16, LM19a, Lin20, LW20b, MV00, MB14, MP20, MN08, NN23, OWW14, PR19, RS13b, Sal22, SSS13, Sep96, Su20, TTV14, dBD09]. Equations [AW09, AAC09, BP04, Beg08, BO10, CN00, Dey11, FR01, Gau07, Get99, KS97, KW10, Kol01, KS01, Lac03, LM96, LS11a, Lot01, MR10, Mas07, MP05, Oda06, Ond10, PR20, RA08, Rho10, Var19, APS21, AM04, ADMU20, BCD20, BWY15, BR12a, BKN19, BP08, BP13a, BLM15, BMBP18, BR12b, BOS22, BMZ08, BDFR15b, BF17, BKM21, BS17, CC98, CL14, Che19, CKLL13, CJK18, DG14, DP20, Dal18, DG20b, DGV17, FFPV17, FK09, For22, GV17, HY18, HS14, HJT20, HP21, HHNT15, Kal21, Kra20, KL12, KN21, Kur07, LZ23, Lej09, Lej22, LMCV12, LPSŽ15, LPSŽ18, Lin13, MT12a, MW20a, MB14, NNT08, Nut12, OO21, SZ19, SWY12, Tur15, WYZ20, ZRLB19, Zha13b, Zha20, dRM22]. equidistribution [Jun15]. Equilibria [DGW04, DG99]. Equilibrium [Ber04, CMT08, ES19, MW05, Pos05, XZ23, Zäh01, AEGR15, BP13a, CEHRB18, GdHO22, MV15, Nan23, PR20, Xu20]. Equivalence [PL16, GJW21, SX19, Tsi06, dHMRS18]. equivalent [HLP16, KP22]. equivalents [CHNR18]. Erased [KL05, AGMQ22, BLV19, CW21a, Law20, LS19, LS22b, Mas09]. Erdos [BHLN15, CFT15, Kif17, Kol22, RT09]. Ergodic [BC12, CSC08, GK07, Han14, ABK13, ACdCdH19, DDT14, MR15]. Ergodicity [HM11, KK12, MST19, CKNP21, CHSW22, Eza13, Kra13, LW20a, RR13]. erosion [FRLS20]. Errata [AL17a, AL19]. Erratic [DFS21]. Erratum [Gro13, HJ09, HJ13, JN07, KSX17, SS21]. error [GNV22, JK19, KM17]. Escape [FL15, Bas23, CC22b, PRU20]. Escaping [Wei09]. ESDs [CHNR18]. Esseen [BCG18, PS22b, Zha22, BOPT20, BHT03, BG13a, ET14, SSW04]. Estimate [IT00, LL04, DF20, DP14, Men14b, Nag12, PS19, Zha19]. Estimates [AH05, BBM10, BMSV06, Kim05, KM18, KL05, Sch05, Sei10, Son06, Str19, Tel01, Uch07, AGL^{+}12, ADS16b, BB18b, BCG^{+}13, BG13b, CKS12, CHZ20, CW23, CCGR22, CJK12, DR14, FF13, Foo09, GL23, GS17b, HW21, JL07, LS19, MT12a, PZ13, Sam09, Sch09, SS20, SS21, Shk13]. Estimating [You13]. Estimation [Ber21b, BCMP09, KSX04, BCG21, BCV21, KSX17]. Estimator [Cha11]. estimators [DE06]. Euclidean [DW16a, Hei14]. Euler [AJKH15, Bes99, BPP22, DG20a, Gro20, JK19, KM17, ML10, MV00]. Eulerian [Car21]. evaluated [BDR22]. Evaluation [LS99]. Eve [DL14]. events [BHLN15, Hol18]. Evolution [BEV10, CT18, CL98, EL07a, HS23, Kol01, Val07, BB18a, BCM19, BDT19, BKM21, CM22, DPP15, FZ23, HJT20, LPSŽ15, LPSŽ18, Sch17a, Zam08]. Evolutionary [EL14, EL16, Dur14]. evolutions [DM09, Rom18, Wer12]. Evolving [Shk09, CT18, DY22, EJS21, KSW14, Sch12]. Ewens [FF14, FF15, Fér13]. Exact [BF20, DZ21, KK11, PR05]. example [LR18]. Examples [JW01, Sch06, RS19, WW09]. exceedance [BI20]. Exceedingly [OT19]. Exceptional [CHHM23, AH08]. Exchangeable [Ber04, For20, GS07, BLS18, Döb15, DK21, Guf18b, Lin16, Zhu22]. Excitable [Gra96, RTW12]. Excited [MS15, Tra18, Vol03, BS09, KZ08, KZ14, KP16, Pet12, Pha15]. Exclusion [BCS06, BS02, Doh99, CP19, Fag08, For16, GS20, GdHM07, GdHM09, Hay21, HS15b, MPX22, Nag12, OT19, RSS20, Sch19b, TT22, Wan22, dS14]. Exclusions [BFN11]. Excursion [Dav08, Gar04, BMM20, JL07, Pit99, Pro21, Zam08]. Excursions [Ale11, KZ14, Hut12, LM07a, LR22, Pic06, QW18]. Existence [BQSS19, BSY22, CK18b, FZ23, HY18, HC20, Hau05, KW10, KL15, Küh20, Lac12, LSF20, Max06, MŞ19, Ott22, PPS18, RM09, RT18, XZ10b, BS22c, ÉM12, KTPZ15, Nil20, RM13, Rom18]. Exit [AAC09, DeB08, DeB09, Feh17, HMP12, KL05, Men07, Tug12, GR14, GR15, KS20, MM20, Tou09]. Expanders [LS11b, AČ20b]. Expansion [DS11a, Fuk11, Igl05, BZ12, HHN16, KS04, NY19]. Expansions [HRW07, LP09, AP17, BB96, Dav98]. expansive [AD19]. Expectation [EL07b, AKTKK19, ERY14, SR13]. Expectations [Dol12, Coh12]. Expected [LZ10, BF20, Spi20]. expert [CO21]. Explicit [CCD13, LS11b, KS17b, KN21]. exploration [EFMN22]. explosion [KST02, LZ21]. Exponent [Law00, LSW01, GS17b, Hul15, Mas09]. Exponential [AF03, BS20b, Del09, FGL15, Hua20, LS21b, LW20a, LM96, LQ11, PR11, PRR20, Sei10, SHW20, XZ10b, Zha18, BCR07, BLR21, Chh15, FMNS15, FM14, GSS21, GR15, HL14, Kra20, KPS12, LP22, PR20, PS12, PS18a, SS20, SS21, Win15]. Exponentials [MS09]. Exponents [Mou01, CMMS23, Dre08, FSV14, MM15, WZ17]. Extended [Dje22, Get99, PS12, Ram06, DKT18, EK22, EK23]. Extending [Dal99, MP20]. Extension [AH21, CN17, LW20b]. extensions [BD13]. Extensive [FJ20]. External [JK11, Ame21, CW21b, DM13]. Extinct [Lam07]. Extinction [BBP12, Bor14, KM11, Val10, BMS13, FL18, KL22, MX07, Par20]. extinctions [TV09]. extremal [BH14, DMS17]. Extreme [BGGM11, NR17, BPZ12, BSV20, DMR16b, FG19, FY20, Kul08, McK21]. Extremes [Abe18, CK22, HLP10, HPZ11, PY17, AZ15b]. Eyring [BDW17].

Faces [KT21a, MW08]. facilitated [BDT19]. factor [Mas22]. factorisation [Kyp16]. Factorization [GHLT03, BCC^{+}22]. Factorizations [HJ11, HJ13, PS12]. failure [BI20]. Families [DDSJ10, AM08, IPR09]. Family [ALM10, CFW09, DW18b, JM20]. Fast [Kol01, BR12b, BDG18, DLZ14, FK21, Gup12]. favourite [CS98]. FBM [Fd01]. FBSDE [DLX20, FIP15]. FCLT [Zem08]. features [Ang22, SHW20]. Feller [CR08, Küh18, PR23]. Ferguson [Sha11]. Ferromagnetic [De 01]. Few [Pak98, Pak99]. Feynman [AD20, Son06, Ste21]. Field [COPV05, DLT21, Sak09, AČ20a, AČ20b, Ame21, AZ15b, AS18, AGMQ22, BCD20, BMBP18, Ben18, BBI09, Che17b, CD12b, CK18a, CW21b, CK23, CTV21, DH19, DKRV17, DLR19, DM23, DG15, Dje22, DR15b, EG18, ELL21, FM14, FMH22, Hul15, JK14, Jeg20, JT21, Kua14, Lac22, LST19, MSS20, Mül17, Nit18, Orr20, PO09, Riv21, SP19, Szn16]. Fields [EL07b, LRMT11, LLS06, Aco14, ACR21, BOPP21, BMM20, CS17, CM20, Che13b, DLMX21, DEM13, FM17, GM17a, He22, HPX22, SX19, SSH22, Ven19]. fifteen [MR17b]. Filament [FG05]. filter [DMR09b]. filtered [KR21b]. Filtering [HL08, KSX04, All21, CKX09, KSX17, RvH15]. filters [DKT18]. filtration [APS21]. filtrations [GJW21, JL15]. Finance [Pet08]. financial [AW09, HJvWW20]. FIND [SND14]. Finding [LP19b, BH23]. Fine [Bal14, LRP13, Rob15]. Finetti’s [IPR09]. Finetti’s-type [IPR09]. Finitary [HS22, Spi20, CPZ22]. Finite [BGGM11, BS01, BvdHvL10, HRW07, MT97, Pak98, PR04, vdHHZ07, vdHS10, AČ20b, Ada20, Ang22, AH19b, BR12a, BDFR15b, CKS21, CHK12, DS12a, DvdHvLS17, DWW19, GJ14, Gap06, Gil16, He22, HPX22, Kul08, LLM18, NN21, Oli12, Pak99, SCZ09, SCW22, Sob12, Son12, Spi20, Wan20, Zhu22]. finite-dimensional [LLM18]. Finite-Range [vdHS10]. finite-time [Ang22]. Finitely [Lev17, SS00, Cor21]. finiteness [AL20]. Fire [BP06, Gra14, Max06]. fires [CFT15, RT09]. First [AJL14, BT15a, CN19, DS12a, GM08, GN03, Lac07, Leh22, Lyn00, Uch11, Zha09, ADS16a, AB16a, AH19b, AT16, DHS14a, DW16a, DR15a, EGvdHN20, ERY14, EFMN22, Gan22, GT17, MS12, Nak22, RP12, RT18]. First-Order [Lyn00]. First-passage [GM08, Leh22, ADS16a, DW16a, Nak22, RP12]. Fisher [Bob22, Dal21, DDSJ10, GLW05, SF06, Tay07]. fitness [Dal21, Der16, FJ20, LO20, Sch17a]. FitzHugh [BK19, BK16, BMZ08]. Fixation [CEY12, DKNS13, FMR22, GPPW16]. Fixed [AR06, BvdHK15, BK05, KM15, Muk16, BS21, CHHM23, JSS17, PKTZ15, PR12, Tar15, Thé20]. fixed-point [PKTZ15]. FK [CDCH16]. FK-Ising [CDCH16]. Flats [EJS21]. Fleming [BEK21, CDE18, CP20, EVY20, BBP12, BB09, DG99, DGP13, FM07, FH13, GK06, Guf18a, Guf18b, Gup12, Han14, KL22, Lab14, LZ12, TN22]. Flickering [BB09]. Flights [Ken20, SBB^{+}21]. Flow [Bur09, CL98, AP12, BG21b, EFP17, HW10, HK20, Kon17, Lej22, RT18]. Flows [DS11a, FK02, Haj11, Zha11, vB09, Bel20a, HR14, HM15a, HM15b, Lab14, MSS20]. fluctuating [BEK21]. Fluctuation [CM20, CES21, FSV14, Kwa19, ACR21, HLP19, JT21]. Fluctuations [AS13, AP23, BCS06, BGGM11, BGT11, BC18, Che17b, FF98, FMW06, Gan22, GZ10, Jos09, LSW12, Luç11, NN23, Pos05, Su06, dCM21, BAKL19, Ber14, BM19a, BC22, CD12b, ES19, Hab19, HS20, IKM21, KR21a, LRP13, Lam18, Unt19, Xu20]. fluids [Fer16]. flux [BB18b]. Fock [GV20]. Fokker [GBM22, BR12b, Zha13b]. fold [CYY22]. folding [BHTW22]. foliation [BHM18]. Følner [DGK21]. forbidden [Jos09]. Force [BBM20]. forced [Yam23]. forces [KOS21]. forcing [Bak16]. Forest [BP06, ARS08, CFT15, Gra14, JSS17, Max06, RT09]. Forest-Fire [BP06, Gra14, Max06]. Forests [LMS08, HBS17, SS18a, SS19b, Tan21]. Forgetting [DMR09b, LS21b]. form [CW23, HLY14, RGV14]. formalism [CR14]. Formation [FG96, HH14, LR18]. forms [CB09, CHK17, Foo09, QZJ14]. Formula [FZ07, Fd01, NV09, TV03, AD20, Bel20b, BCPW21, BG21b, Che13a, FKHL19, MP20, NZ14]. Formulae [TZ11, ŁOPS21]. formulas [KT21b, PP15]. formulation [BC21, BL19]. Fortuin [GS17b]. Forward [CZ05, KMP06, KS01, BT20, BDFR15b, DM16, MB14]. Fourier [GIP16, SHW20]. Fourth [DVZ18]. Fractal [BK01, BC08b, BC10]. Fractals [KPX00, DKN18, HY18, Zha13a]. Fractional [Beg08, BO09, BO10, BD14a, BL05, BSV20, CKM03, DS07a, DS06, Gau07, GN03, Igl05, LRMT11, LA09, LS11a, MS09, MNV11, TV03, BJQS15, Bal14, BOPP21, BC08a, BZ12, BDR22, CHSS18, CZ21, DG14, DS17, DM22, GNV22, HPX22, Ina16, Nil20, NRS10, NZ14, PT07, Ton09, TTV14, Var19, Ven19]. fractional-colored [TTV14]. Fragmentation [AS02, Bas06, Ber04, MWZZ02, BB15, Dad17, Ged19, Shi17, Ste13, Wie15]. Fragmentations [BG04, HM04, Mie01, SW19]. Frames [LP09]. Framework [XZ10a, EP23, KQ12]. Fréchet [BLO13, Sim14]. Fredholm [FTZ20]. Fredrickson [BDT19]. Free [BM11, BS22b, CDMFF11, Gil11, Has14, Ko20, Sak09, AČ20a, AČ20b, AZ15b, AH18, AS18, Ben18, CLT21, CK23, Cor21, DFPSL19, DMM17, DR15b, GM17a, IJ21, Jeg20, Kua14, LST19, Nit18, Szn16]. freeness [Gro17]. freezing [BC18, CGL18]. Frequencies [GS07]. Frequency [BG08, CB16]. Frequent [BR07]. fringe [HJ15a, Jan22]. Frog [AMP02, BDD^{+}18, CN19, HJJ19, MR20]. Frogs [Her18, MW20b]. Front [BDT19, CM22]. frontier [ABK13, FN20]. Frustrated [De 01]. Full [Bat21, DM22, NN21]. Full-path [Bat21]. fully [DLX20, KTPZ15, Tan13]. fully-nonlinear [KTPZ15]. Function [AAC09, KN07, Ben21, CW23, DRM14, DW19, FZ23, JS22, LS19, Uch20, Zha21]. Functional [BJ09, BB11b, BT20, BDD09, BP07, CGGR10, FPT19, Kif17, KP16, LR11, Owa17, Wan20, BWY15, BSY22, BS17, Dav98, DK21, DGP09, JG13, LP14, SWY12, Vel22]. Functionals [BJKM10, CMN97, HM01, KK11, LS99, MT02a, RNP10, SY05, BP16, BBGT20, BOHP19, BOPT20, BS20b, BLR21, BST22, BDR22, CT22a, CMN98, Chh15, DDS18, DLR20, ET14, ERY14, Fuk99, Gao17, Jan22, JT21, KT17, KPS12, MR15, NZ20, PS12, PS18a, Pèn21, Pic06, PS18b, PS22b, Sch19a, SY19]. Functions [Arc99, Kry99, Kry04, Uch07, Uch10, vdHS10, AH21, AKM12, BC14, BLM09, FH23, FW17, GS16, GSS19, GMS19, HJ15a, HLY14, KL15, KR08, Lis22, Mas22, MS13, Naj20, PS18a]. fundamental [BPR15]. Further [HH14]. Fuzzy [War10].

G [Duf10]. Galton [AD14b, AD14a, AD19, AA23, BS15, BGH^{+}14, BDR21, CK08, DS12b, Dev11, DHS19, Far11, FJ22, FJ23, GS20, GGS21, Gei00, GN21, Joh19, Lin14, MPR19, MW20b, PR11, Rou18, Su14, Thé20]. gambler [ALS15]. game [AS13, BK08, DLR19, Fox18, HH06]. Games [DLT21, DG18, DQS16, Dur14, EL14, EL16, Spi22]. Gamma [Wan20, CFW09, Chh15, Duf10, ET15, IT99, OO15, PS18a, Gau14]. gamma-distributed [OO15]. gamma-mixed [IT99]. Gap [KW11, MW05, SW11, CMMS23, KKO13, Nag12, ST17a]. gaps [LL21, PY17]. gas [Ame21, BN21, Unt19]. gases [Ber21a]. Gaussian [AČ20a, AČ20b, Aco14, ALM21, AL21, AZ15b, AS18, Aug16, BQSS19, BJQS10, BCR07, Bas08, Beg05, BMM20, Ben18, BH21, BP15, Ber23, Ber14, BM22b, BM19a, CC22a, CS17, CHNT17, CK23, DLMX21, Dav98, Dom22, DR15b, DW16b, Eis05, EG15, ERSY10, Fol11, FS20b, FIP15, GZ10, Gro17, GM17a, GKT22, HHNT15, HLP10, HPZ11, IKM21, JPP19, Jeg20, JS17, Kua14, LRP13, LR22, Lam21a, LLS06, LL11, Lii11, LA09, LSW22, LV11, LST19, LP09, Mai07, MS13, Nit18, NZ20, PR20, Pen07, Pin12, Pow18, RX13, Rie17, Riv21, RS21b, RM09, SX19, Sak09, Su06, SND14, Szn16, Tel01, Web15, Zha14, vdHS04]. Gaussians [AC19, BC20]. gelation [Wag06]. Gelfand [WW09]. GEM [PY17]. gene [BP14a]. Genealogies [EP22, BJJK21]. Genealogy [HM04, PRT22, Joh19, Sch17b]. General [BM05, CV23, DLW01, DS04, IT00, Ste13, APS21, Alt21, Bat18, BPT16, CP21, Coh12, CHSW22, CKS21, CG20, DMR16a, DE06, DMR09b, EM19, EKYY13, EX22, HHNT15, HL12, IKM21, JM22, KRST13, KR21b, KQ12, KOP13, Kur07, Lej22, LW20a, NN23, Pin22, PZ22, Red17, SS18a]. Generalised [BB09, ZRLB19]. generalizations [Döb15]. Generalized [ALR23, DQS16, FZ07, JL15, AEK^{+}14, DL22, GT07, GM17a, Guf18b, Han14, HMN23, LSF20, Leg21, Pin18, RS09, Zha22]. generated [AP12, Ber18, BDI20, CHNR22, Cor21, DDS18, DLR20, MV15]. Generating [KN07]. Generation [DKM01, KS97]. generations [AD14d]. Generator [Get99, Kli12]. Generators [FJX11, Pak98, Luo20, Pak99]. generic [Adh22, Dir15]. genes [BP14a]. genetic [For22, HP08, PW23]. Genus [Bet10]. geodesic [Die16, Ver19]. geodesics [BGH22, Bud19]. Geometric [CL98, DGLU11, Jor13, Pen07, RR13, SW11, BP16, BY09, BST22, BS21, CIW19, Cou18, FR13, FY20, LRP13, Mél19]. geometrically [DDT14, Die20, MR15]. Geometry [BC17, PP11, Ray14, Sén21, XKW06, BO19, PS21]. getting [ABOPS16]. GFF [LW21]. Ghirlanda [Arg08]. Giant [Ber11, Ber14, CK23, Gol18a]. Gibbs [AD20, DG09, DLR20, EG18, FM04, GS07, GM22, HKLR19, HM22, KO08, Lac22, Vel22]. Gibbsian [DFF^{+}21, HMN23, Jan16]. Gibbsianness [KO08]. Ginibre [Dub18]. Ginzburg [BBP20, HIN17]. Girsanov [BG21a]. Gittins [CT22b]. given [CP23, Jan09, LM07a, Lod22]. giving [CK15]. glass [Bäu19, BT17, CW21b, CGL18]. glasses [Adh22, AC17, Ko20]. Glauber [Bia08, BMP21, KKO13, TT22, Völ14]. Glauber-exclusion [TT22]. Global [BDvdHS22, CKS12, Sal22, Unt19]. globally [Luo20]. Globules [Fra10]. Go [PPPY01]. GOE [FO18]. Goldie [ALM10]. goodness [KO08]. gossip [BR18a]. Governed [Beg08, BO10, Lac03, LC10]. Gradient [HKLR19, KT21b, CHZ20, DG20b, DP14, GPP13]. graining [Ton09]. grains [Bur15]. Gram [AEK17]. Graph [BYY07, DE09, KSZ01, AE19, AGMQ22, BMH22, BST16, BKLM19, Che14, EGvdHN20, FR13, Fox18, Gay12, Gib08, KR21a, Kha14, Kol22, LW21, Mél19, MŞ19]. Graph-Homomorphisms [BYY07]. Graphs [ABBG10, BM11, BS01, BvdHvL10, DH06, DHP11, DGLU11, Jor11, Tel01, Win10, vdHHZ07, AČ20a, AČ20b, AFL23, BC12, BT15a, BT21, BPT17, BHLN15, Bia08, BJK13, CDC13, CK23, CP23, CHK12, DCLM18, DY22, Fér18, FY20, Gal07, GJ14, GS17a, HR14, Hil14, HBS17, HP17, IKN20, Jan09, KP13, LRP13, LO21, MVY13, MV16, Neu07, ORS20, PL16, PZ22, RT09, Sob12, Tan21, Tan22, Tei09, Tur07, dHMRS18, dHNT18]. gravity [AFS20]. great [KT21a]. Greedy [RP12]. Greeks [BM06]. Green [AAC09, CW23, Uch10, Ben21, FZ23, Uch07, Zha21]. GREM [FG19]. grid [BI20]. Gromov [ADH13, GM17b, dCM21]. Grothendieck [Dom22]. ground [Bäu19]. group [Ber18, Mél19]. Groups [BIK07, DS11b, Jea05, Pak98, PR04, RT10, BC08a, Cor21, Ful12, HL07, Pak99, SCW22, Sam09, Sam17]. growing [AAJ21, BH23, DHS14b]. grown [Tur07]. grows [FIO^{+}21]. Growth [AID07, BCS06, FJX11, HMN11, RT10, Shi17, vB09, Ass18, BS08a, BF08, CHL18, CFT15, Dad17, Do21, EJS21, Eza13, Ged19, Hua20, IdRS19, Ler20, Mas09, Pag17, Pin16b, PRW14, Seg09, SS20, SS21]. Growth-fragmentation [Shi17, Dad17, Ged19]. GUE [JN07, JN06]. Guerra [Arg08]. Guerras [BK09].

Haar [LA09]. Hadamard [GHLT03]. Hahn [Vet15]. Hairer [Yan23]. Haldane [BCPW21]. Half [LC10, Law00, MP05, BR18b, CKS12, Gra14, GM17b, KK18, MN18a, Par22, Ray14, Ric15, Ric18, Tak21]. Half-Line [LC10, KK18, Tak21]. half-planar [BR18b, Ric15, Ric18]. half-plane [Gra14, GM17b]. Half-Space [MP05, Par22]. half-space-like [CKS12]. Hamilton [Kra20, MP20]. Hamiltonian [Kol01]. Hammersley [CD06]. Hamming [MŞ19]. Hanson [KZ20]. Hanson-Wright [KZ20]. Hard [Sak09, Chr19, DLV21, FV17, HHN16, Hol18, dHNT18]. hard-core [dHNT18]. hard-edge [FV17]. hard-sphere [Chr19]. hard-to-soft [DLV21]. hardcore [BST16]. Hardy [KW21]. Harmonic [GLM17b, GMS19, Jea05, Kry04, DW19, ES19, KS04, Lin14, Rou18]. Harnack [DS17, Foo09, HB09, MK12, SWY12]. harnesses [BW12]. Harris [SS19b]. Hausdorff [ADH13, GM17b, AS08, BHMK21, BGH22, Gal22, Law96b, Rou18, Zha13a]. having [KK20b]. Hawkes [Che17b]. Heat [ADS16b, BJQS10, BG21b, Fol11, Foo09, Lac03, LC10, MCS03, Stu03, AB13, Bel20b, CKS12, Che19, CK20, CK22, CN17, DP20, DJQS13, DM22, Ful12, HY18, HHNT15, KM18, LW20b, MN08, NN23, OO21, PS22a, RGV14, RS13b, SZ19, Su20, TTV14]. Heat-Type [Lac03, SZ19]. Heavy [BW05, CGGR10, Čer11, DHS19, BB11a, BOPT20, BGG14, BTW22, BDvdHS20, BDvdHS22, BPZ22, Bob07, CO22, DW12, Fle18, GMPW19, HZ19, HMMW12]. heavy-tail [BTW22]. Heavy-Tailed [Čer11, BOPT20, BDvdHS20, BDvdHS22, BPZ22, CO22, Fle18, GMPW19, HZ19, HMMW12]. Hedging [BT05, CB09]. Height [BdT14, SS18a, SS19b, ABBHK22, DÜ22, Lis22]. heights [Haa17, Uch20]. Hermite [CD12a]. Hermitian [BBCC21, CHNR18, Jal21, LO21, Nem17, OS11]. heterogeneities [For19]. heteroskedastic [CHZ22]. hexagonal [AJNT22]. Heyde [BM19b]. hidden [All21, DMR09b, HMP12]. Hierarchical [BGT11, DGW04, KMT12]. Hierarchically [DG99]. hierarchies [For20]. High [Cip13, CD08b, DS06, DGLU11, DR15b, HLP10, LC10, Pan05, SYY21, AT16, BL12b, BC17, CE13, Ott22, Rap23, SS19a, SSS13, Völ14, Yan22]. High-Dimensional [DGLU11, DR15b, SYY21, SS19a, Yan22]. High-Order [LC10]. Higher [ACR21, Beg08, BO10, FV11, GSS19, KR21a, Lac03, Uch10, vdHS10, BHdS^{+}19, JLS13, Lin20, LSW22, Pha15]. Higher-Order [Beg08, BO10, Lac03, KR21a, LSW22]. Higher-Point [vdHS10]. highly [BEGG14]. Hilbert [BC20, Dal18, Ose11]. Hilbert-Space-Valued [Ose11]. Hilbert-Valued [BC20]. Hill [Cha11]. Hilliard [BSW10]. hiring [Jan19]. Hirsch [Hu00]. Historical [EP98, Jac10, Kli14, MT12b]. hit [LM19a]. hitchhiking [HP08]. Hitting [JJK^{+}17, JK15, Lac07, MT02b, SY05, Uch11, DP20, ERY14, FMNS15, KKPW14, Kwa12, Muc21, Oli12]. Hölder [CHZ20, HKY14, HHNT15, HNX19, LS22b, MP05]. Hölderian [BFG99, Zem08]. Homeomorphism [Zha11]. Homogeneous [Dal99, FS01, FG03, SV05, BQSS19, BLM15, GMW19, Gia08, HMMW12, dBD09]. Homogenisation [Per20b]. Homogenization [BEP09, CN00, Lej02, LS11a, Rho10, Fag08, GM17a, GPT22, LNO15]. Homomorphisms [BYY07]. Hopf [CTV21]. horizon [ALR23, Ang22, Gap06]. horizontal [BP14a]. Hörmander [Hab19]. Hörmander’s [GH19]. Horn [DeB08]. Horton [BDR21]. Hot [Ata01]. hull [Eld14, MW18]. hulls [CSŠ22]. Hurst [DS07a, NRS10]. hybrid [EFP17, EGL22]. Hydrodynamic [BGRS10, BMH22, Fag10, FMW06, KOS21, Ler20, SS18b, Fag08, FSX20]. hydrodynamics [ORS20, PYZ22, RSS20]. Hyperbolic [LRMT11, LW01b, MW10, Xu20, BT15b, CFG17, CM17, DSB19, FY20, FvdHMS21, HMMN15, Tyk07]. Hypercontractivity [BWY15, HKL^{+}98]. hypercube [GA08, Pin13]. hypergeometric [OS20]. hypergraph [MN21]. hypergraphs [CP19]. Hyperplanes [AC07]. hypersphere [KT21a]. hyperviscous [Fer16]. hypocoercivity [ADW21]. Hypoelliptic [HS17b, CCDO21, Hab19]. Hypoellipticity [HM11, Mic17]. hypothesis [GJW21]. i.i.d [BBCC21, Jos09, KS19]. I.I.D. [Shk09]. identical [BCP18, Fit98, LZ10]. identically [IPR09, SR13, Zem08]. Identities [Arg08, Doh99, Chh15]. identity [LST21b, Par22]. IDLA [She10]. II [SS19b, AČ20b, DF20, GKT22, GS17b, HM15b, IT00, IMN23, Ram22, Sch17b]. iid [Jon01]. III [Nan23]. immigration [BD14b, FH13]. implies [HLP16]. importance [BHLN15]. impossible [GM08]. Improved [Doh99, FH19]. impulsive [BMZ08]. incipient [Ric18]. Including [IT00]. Inclusion [Doh99, BDG17, GRV13]. Incomplete [KW00, LQ11]. increasing [AM08, BBJJ22, CR09, DH21, Sep96, Tra15]. increment [RS21a]. Increments [BV04, CH05, Shk09, BOGP12, BOHP19, Gan22, GMPW19, HMMW12, MR12, SX19, SZ21]. Independence [DP05, BCC^{+}22, Lac22, Tar15]. Independent [HKL^{+}98, HRW07, KPV04, OS11, AB16a, BBCC21, Dol16, FG20, HJ07, HJ09, Jal21, Nem17, Pin18, Pit99, PS18b, PS22b, Zem08]. Index [DS07a, HM04, Mun11, BJQS15]. Indexed [Arc99, Ven19]. Indices [GS07]. individual [CKS21]. individual-based [CKS21]. Induced [Bou11, AS18, AKM12, HM15a, HM15b]. induction [CGR20]. inelastic [BLM15]. Inequalities [CGGR10, Doh99, Eis13, KMP06, NV09, Ose11, AS19, ACdCdH19, BP16, BBGT20, BRW20, BT20, BE21, BN14, BCG20, BP14b, CKNP21, Del09, DS17, FGL15, FF13, FPT19, Foo09, Fra09, GSS21, GM14a, Goz12, JG13, KZ20, Kur07, Liu20, MK12, MW19, Oça21, Osȩ13, Pau15, RC09, Rie17, Sam17, SY19, SZ19, SWY12, SHW20, Ste21, Tan19, Var19, Vel22, Wan20, Web21, Win15]. Inequality [DP05, De 01, Maj07, Ada08, DZ15, FOW22, HB09, KW21, Liu20, MR13, Pau14, Riv21, Völ14]. inert [Whi07]. infected [CSW22, DY22, SS14]. Infection [HJJ19, KPV04, BS22a, CSW22]. Infinite [Arc20, BLM15, BH21, CC11, CB08, DEF^{+}02, DH15, FM03, MB14, SSG11, XZ10b, vdHHZ07, AD14b, BCC19, BG09, BvdHK15, BP12, BB15, BR18b, BLV19, Bor14, BC17, BS22c, BFMM13, CS15, Cou18, CK08, DW16b, Fag08, GJ14, Gan13, Ger20, Gib08, GGR19, GLT21, GM17b, HJ07, HJ09, HS22, Has14, KR21c, LB18, Law20, Lin14, Max06, MN14, Neu07, Ric18, RTW12, RM09, ST17a, Uch20, VY16]. infinite-dimensional [BCC19, BP12, RTW12]. infinite-range [HS22]. Infinitely [ALM10, ST06, WY10, BOPP21, BP14a, BPZ22, BO13, DGP19, DEM13, Eis19, Mat23]. Infinity [Lim10, Sav08, Cos23, Die16, Fou19]. Influence [Leg21, FI22]. Information [BJKM10, LM08, LV11, Bob22, CCC15, FLM20, Web21]. Inhomogeneous [ADS16a, BvdHvL10, BGT09, CCDO21, CT20, EJS21, ÉM12, FZ12, Pet20, RS21b, SCZ09]. inhomogeneously [RW18]. Initial [LS11a, BGT09, BD12, CG20, DMR09b, NN23].

Innovations [HL08]. Inputs [BHS09]. Insensitivity [Tan06]. Insertion [HS13, Lev17]. Insertions [Sub13]. insider [CX13]. instantaneous [BA12]. integer [FS17]. Integers [HJ11, HJ13, KZ08, KZ14, Mas22, RS09]. Integrability [BO11, Hob15, Tou09]. integrable [HMMW12]. Integral [Dal99, LS99, ŁG14, PT07, SY05, Bav19, HP21, Kal21, NRS10, WYZ20]. integral-differential [WYZ20]. Integrals [Cho05, FZ07, Fd06, GN03, Lej02, BK18, CD12a, Dav98, GNV22, GT07, NY19, Pic06]. integrated [GLT21, TT20]. Integration [FKHL19, BKN19, BOGP12, BR22, CT15, DMN13, FM21, GIP16, VY16]. Integro [MP05, Zha13b]. Integro-Differential [MP05, Zha13b]. Interacting [BFN11, BR11, BDD09, BGdH11, CP16, DG96, DKM01, FG96, GLW05, Hut12, ORS20, Vil11, BC21, BBM20, BT19, BDFR15a, BF17, Che17b, DP13a, DG99, DMP05, EHMT17, ELL21, FTZ20, FMH22, Gup12, JT21, KL21, KR21c, LZ23, Mül17, Orr20, SS18b, SS15, XZ23]. Interaction [DG09, Eng10, Die20, DLR20]. Interactions [XZ10b, AM19, BN21, EZ16, Jan16, Mül17, SP19]. interchange [BKLM19, MŞ19]. Interface [MT97, Pos05, Val10, CP09]. Interfaces [BMSV06, Mal15a, Val07]. Interior [Kry04]. Interlacement [Tei09, ČP12]. interlacements [CPZ22, EK22, EK23, LS14, RS13a, Szn09]. interlacing [Kua14, War07]. Intermediate [BGdH11, CN17, Nic21, BDI20, FSV14, Lac12]. Intermittence [FK09]. Intermittency [GdHM07, GdHM09, GLT21, HHNT15]. intermittent [CJK12]. Internal [HS12, JLS13, Mun11, ČP12]. Interpolation [AC07, Zer11, Hil14]. interpretation [HS14]. Intersected [AS08]. intersecting [NR17]. Intersection [BR05, Law00, PSSS17, AB16a, BCR06, BJK13, Lau12, MS99]. Intersections [Ros97, GW09, Shi12]. intertwining [RLR22, War07]. interval [BA12, DWW19, FPRW20]. Intervals [Bas06, Jun15]. Intricacies [BJN13]. Intrinsic [Dar01, Gol18a, vR04, SSH22]. Invariance [And21, BR05, GMW19, Bis19, CC13, DDM12, GB14, HS17a, Rou15]. Invariant [Ata01, BT21, BS17, FGS16, FTT14, Led04, Led05, MWZZ02, OWW14, Rap23, Rou18, Ald14, AG13, BSY22, BM20, CR14, FG20, Fit21, FK21, GPP13, HJP22, HS17b, Kam18, Lam18, dT21]. Invasion [MPR19]. Inverse [MNV11, SV09]. Inversions [Tra15]. Inverting [LST19, LST21b]. Inviscid [Bak16, BM09a]. involutions [Ber18, RZ14]. iPod [Lan15]. Irregular [Beg05, BMBP18, CL14, DG20a, Zha13b]. ISE [Del03]. Ising [AKPS19, AJNT22, AC96, BdT14, CDCH16, CDC13, FM17, HL12, Spi20, War10, dT21]. Island [Hut09]. islands [CJK12]. Isolated [ABPR11, Pen09, FY20, HMMN15]. isolation [For19]. Isomorphism [EK22, FR14, EK23]. Isomorphisms [Lup21]. isoperimetric [GMPT17, Hut23]. Isoperimetry [BCR07, CGGR10, Bob07, Gol18a]. isoradial [DCLM18]. Isotropic [AEK^{+}14, CL98, vB09, Feh17]. Itô [AF03, Bel20b, FZ07, Fd01, NZ14, TV03]. Iterated [AW09, Arc97, HKL^{+}98, KP04, LR11, Ros97, AL09, BCR06, GW09, IKM21, Nan06, NP08]. iterative [DGV17]. IV [BR11].

Jacobi [Kra20, MP20]. Jacod [GJW21]. Jagers [SS18a, SS19b]. Johansson [CHHM23]. joining [Ham19]. Joint [BBGH12, GHS21, LC10, Lam01, QZJ14, Lod22, Orr20]. Jong [DP17]. juggling [ABCN15]. Jump [AB02, DB12, Ger20, KW11, PM08, BCKH22, BCC19, BM06, Col09, CZ21, DH19, FTT14, Fou08, KT21b, KS20, LSF20, SCW22]. Jump-Diffusion [AB02]. Jump-diffusions [PM08, BCC19, FTT14, LSF20]. jump-like [KS20]. jumping [Fou08]. Jumps [Ber97, Bor11, HO03, LMM04, Mat11, Mie01, ADMU20, BA12, BPZ22, DQS16, GP12, He22, JK15, KTPZ15, Kwa19, PPS18, PKTZ15, RS21a, Zha13b].

K-sat [Pan14]. Kac [AD20, BP13a, CE13, COPV05, IT00, PO09, Son06, Ste21]. Kac-like [BP13a]. Kalman [DKT18]. Karlin [Ass18]. Kasteleyn [GS17b]. Kato [BG21b]. Kawasaki [BN21, CCR05, dHNT12]. Keller [BF17]. Kernel [ERSY10, AD14c, ADS16b, CKS12, DE06, Foo09, Ful12, HHN16, Ina16, RGV14, Unt19]. Kernels [FG03, Fol11, BE21, KM18]. Kesten [BDP15, MR13]. kick [Bak16]. Killed [BG11, Uch10, BLR21, FKHL19, Kwa12, Muc21]. killing [LZ15]. Kinds [SW11]. Kinetic [ABT15, LS11a, BP13a, BKM21, CCGR22, FFPV17, JM22, Per20b]. kinetic-type [BKM21]. Kingman [FRLS20, JK11, LT15c]. Kirkpatrick [Che13b, Pan05]. Knight [LST21b]. Knock [FW08]. known [CSW22]. Kolmogorov [BDFR15b, KP04]. Korteweg [dBD09]. KPZ [CG20, CHHM23, Ler20, MB14, NN23, Nej21, PR19, SS15]. Kraichnan [HK20]. Kramers [BG13b, BDW17]. Krylov [MR01, Zha19].

label [Lod22]. labels [AS18]. Ladder [Sel06, Uch20]. ladlag [CB09]. Laguerre [AT09, JV11, SW14]. Lambda [GY07, BEK21, CDE18, CP20, EVY20]. Lambda-Coalescents [GY07]. Lambda-Fleming-Viot [BEK21, EVY20]. laminations [KM16]. Lamplighter [PR04, DKN18]. Landau [BBP20, DMN15, HIN17]. landscape [BGH22]. landscapes [Dal21, Tug12]. Langevin [BHP21, EG15, HMN23, HS17b, Ram22]. languages [Gil16]. Laplace [Dav98, GHLT03, Maf20, PRR20]. Laplacian [dT21]. Large [ABK06, APW20, Aug16, AdHR11, BR20, BM09a, Bob07, BF08, BCV21, BG03, BC08b, BD03, BDG18, But16, CLR05, Cor21, CS98, DLR20, Fau02, FF15, Fen07, FS20a, GPP13, GMT05, Gau07, GM20, GM22, HHN16, HM01, Hei22, HX08, Hus22, Jac10, KM05, Lau12, LP15, LC05, Mai07, MCS03, Mat11, McK21, Orr20, Ort12, Pet12, Puh04, SCZ13, Seg09, ST22, Tan06, Ver19, Vos08, Wan05, Zha20, BBP20, BB11a, BEGG14, Bav19, BC21, Ber19, BS08b, BD18, CD19, CR14, CYY22, CGPV13, DZ22, DDS18, Feh17, FH19, GLM17b, Gro17, HS17b, Kif17, Kra20, Mas22, MT22, MW05, Mül17, OR14, OT19, OWW14, OPR22, PO09, QZJ14, RASY17, Sob12, Spi15, SZ15b, Su20, SP19, Tar15, Thé20]. Large- [BC08b]. large-dimensional [QZJ14]. Large-range [Seg09]. Largest [CP23, Aug16, GT23, GM20, Hus22, Mai07]. last [Fit21, HS20]. Late [DPR03]. Lattice [Hol08, KSZ98, AJNT22, BN21, CMT15, CMSS16, CN18, CR14, CHT19, EL14, EL16, KMT12, SS19a, STV22]. Lattices [IMP07, MW10, BFMM13, Din12, HS12]. Law [AW09, Arc97, AdHR11, BHPZ11, BU11, BHM13, Cho05, HY10, Jac10, Lam01, AEK17, BPZ12, BCZ18, BP12, BC21, BLR21, BG13b, BDW17, CES21, CKS21, DH21, EKYY13, Fer16, FKK22, Fou08, Hei13, IKM21, Kif17, KP21, MVY13, Nem17, PY99, QW18, XYY17]. Laws [Ada11, BG04, DM11, Hu00, Nan06, Ros97, AEK^{+}14, AH19b, AH19a, BBG19, BCR06, CES22, DS07b, ÉM12, Feh17, GW09, HJ15a, KP16, Mal15b, Sim14]. layer [CN17, LW20b, Nic21]. layered [BHP21, DF20]. leading [Ber21a]. Learning [BRS10]. Least [Gro11, CL19, CL21a, Gro13, Ngu12]. Leaves [Pen20]. Lebesgue [FZ07]. left [HN22, NV13, Pin16a]. left-curtain [HN22]. left-tail [Pin16a]. leftmost [RLR22]. lemma [BL19, Lê20]. lemmas [BCP20]. Length [Del03, JK11, Mun11, CS20, CGL08, CJK12, DH21, HI23, ZZ13]. Lengths [BU11, LZ10, Ton07]. lent [BD13]. letters [DS12a]. Level [AČ20a, AČ20b, NM04, AL20, Ber21b, CMT15, CK23, DR15b, DLV21, ŁG14, Nit18, Szn16]. Level-set [AČ20a, AČ20b, Szn16]. level-sets [CK23]. levels [BDI20]. Lévy [CD08a, ADV10, AH18, AF20, Bal14, BB11b, BP20, BOHP19, BK08, Ber97, BI20, BR22, BHM13, CD05b, CM20, CK22, CLT21, DH19, Dav08, Don04, DWW19, EI16, EG15, FL18, FZ10, FPRW20, GS16, GM11, JK15, Kho96, KK11, Küh20, KS17b, KR20, KPS12, Kwa12, Kwa19, KR08, LMCV12, LR11, Mal15b, MM18, MM20, MR12, MT02a, Mie01, MVJ14, MM21, PS12, PS18a, PYZ22, Pet08, PT18, SV09, Sav08, SS00, SZ15a, SBB^{+}21, TY23, TT20, WY10]. Lévy-derived [EI16]. Lévy-type [Küh20]. Li [BG21b]. Lie [BC08a, Chh15, Ful12, Mél19]. Lie-theoretic [Chh15]. Lieb [Men14b]. Lifetimes [Lam01, BG13b]. like [Ale11, BP13a, BF08, CKS12, DÜ22, KS20, SCZ13, SY17, Son06]. Limit [Ale11, AP23, AH18, BGRS10, BHLT01, BJ09, BB11b, BDD09, BPR12, BBL11, BC08b, CP99, CR08, Col09, CHSW22, CCFR09, DW10, Dol16, EL09, Fag10, Far11, FMW06, FY20, GK06, HJ15a, HJ11, INS21, IMN23, JV11, Kar15, Lev13, NY09, PPSV21, Pan05, Pèn21, PP11, RTW12, Ste99, SBB^{+}21, TT20, Vil11, Wan05, AHS20, AB16b, AA23, BMH22, BL12a, BR18a, BLO13, BDF18, BOHP19, Bav19, BGG14, BDD14, BS20c, BBDS23, BST22, BJM14, BT19, BPW18, Bor22, BF22, CT22a, Che13b, CD15, CT20, DMR09a, DLR19, DS12b, EP23, EVW20, Fag08, FMH22, Gup12, GKT22, GM17b, GHS21, HB09, Hay21, HT10a, Her20, Hor20, HI23, HMMN15, HJ13, IKN20, Jan22, JSS17, JR14, JT21, KOS21, KP16, KT17, Lam21b]. limit [Ler20, LS22b, Lin20, ŁG14, MSS20, MR12, MS99, Mou12, NXZ22, Orr20, Owa17, PST17, Ram22, RSSZ19, SS13, SYY21, SND14, VFV15, Vet15, Xu20]. limited [ACSS16, BBJJ22]. Limiting [ABBG10, BD08, BHS09, Mou01, Red17, Yan22, BM14, CHNR18, DH21, FH19, FL13]. Limits [BS01, Bet10, BvdHvL10, BDFR15a, Luç11, Tow15, Wei07, AD14b, AD14a, ACdCdH19, BS15, BCM19, Bus16, CC22a, CP16, DKLT22, DMP05, EVY20, FSX20, Gei00, GS17b, LT15c, LW21, Nic21, ORS20, Pag17, RW18, SS18a, SS18b]. Limsup [KPX00, Zha13a]. Limts [Pen07]. Line [AS08, LC10, Uch10, WY10, AY15, DFF^{+}21, Fit21, GH15, Goz12, Ken20, KK18, LST21b, Pen20, PR19, Tak21]. line-breaking [GH15]. Linear [AF03, BWZ09, BG11, CD08b, DM03, Lot01, MP05, NY09, NY10, Ond10, RA08, Web16, AW14, Bal14, BFGM19, BL19, CB09, CMMS23, DH19, DS04, DMS09, DM22, DLT21, ET14, EG15, FIP15, HMT23, IdRS19, JT21, JS18, Kar15, KS17b, KN21, LMCV12, MR17a, Mer08, Nil20, Sal22, dS14]. Linearities [BM09b]. linearization [AL17b]. linearly [AA23, Tak21]. Liouville [AFS20, CFG17, DKRV17, DFS21, PL16, RGV14]. Lipschitz [Bas96, CKLL13, DSWB15, JG13, LR11]. Lipschitzian [SWY12]. Local [AD14b, AD14a, Ale11, AB16a, AB16b, AEK17, AH05, BM11, BS22a, BR05, Ber97, BBDS23, BGdH11, BDFR15b, CD19, CCC15, CLR05, CD15, CH05, DVW14, EP98, EL07b, FZ10, FKK22, GH09, Hu00, IKM16, JSS17, Kur08, ŁOPS21, MP07, MW20a, MX07, Nem17, Ond10, Pag17, PP11, PP15, TY23, Tel01, TV03, XYY17, XKW06, Abe18, AEK^{+}14, AS18, BX19, BCR06, Ber19, Ber21b, BČD16, Che13b, CES22, DFPSL19, EKYY13, Fan19, Foo09, For19, GS17b, HB09, Hon18, Hon20, Hug19, LZ23, Lau12, ŁG14, Lup21, MR12, MT12a, MT22, MS99, Mül17, OW20, ORS20, SV09, SP19, Tsi06, Tur15, vSW18, Dol16, EGvdHN20]. localisation [FM14]. Localization [AD15, BS20a, BR12a, Bat18, COPV05, Fle18, LV11, NY10, Ame21, Bat21, CZ21, OR17]. localized [Bel20a]. locally [ADH13, BDG08, Roe08, Sam17]. Location [Dar01, SS12]. loci [FJS15]. Loewner [FZ23, Wer12]. Log [AI09, FOW22, Aco14, Str19, Unt19]. log-concave [Str19]. log-correlated [Aco14]. log-gas [Unt19]. Log-Sobolev [FOW22]. Log-Type [AI09]. Logarithm [AW09, Arc97, Ros97, BCR06, GW09, IKM21, Nan06]. Logarithmic [BP16, BN11, BYY07, DP05, Vel22, Win08, CGL18, DHKW22, Ste21]. Logconcave [HLY14]. lognormal [BDD14]. Long [CCR05, CFG09, CCDO21, CM17, EGvdHN20, FK02, Gao17, IT99, KMT12, MY03, MS21, AKMV20, CCK22, DW18b, For22, Hul15, Kul08, MSC15, SCW22, SS18b, SP14]. long-jump [SCW22]. Long-Memory [MY03]. Long-Range [CFG09, FK02, IT99, KMT12, CCK22, DW18b, Hul15, Kul08, SS18b, SP14]. Long-time [CCR05, CCDO21]. longest [DH21, GL22a, HI23, Sep96, Tra15]. Longtime [DG03, GKW99]. look [BS08b]. lookdown [Lab14]. Looking [HY10]. Loop [AGMQ22, Cha15, KL05, AS19, AHS23, BLV19, BELR21, CW21a, FR14, Law20, LS19, LS22b, LST19, Mas09, vdBCL18]. Loop-Erased [KL05, AGMQ22, BLV19, CW21a, Law20, LS19, LS22b, Mas09]. loop-soup [LST19]. loops [BdT14]. looptrees [Arc20, CK14]. Lotka [AM19, CMP10]. Low [AC96, OR14, BT17, FG19, RS19, Völ14, dHNT12]. Lower [Bou10, CM06, Nag12, Sam09, Vih11, BDvdHS22, DP20, LS14, Sub13]. Lushnikov [Jac10]. Lyapunov [AKM12, CMMS23, Dre08, MM15].

Made [War02]. magnetohydrodynamics [Yam23]. Majorant [Gro11, Gro13, OP22]. majority [AB20]. Malliavin [SSS13, BM06, DP14, ET15, GS16, Hal22, NV09, Pet08]. Mallows [Pin22]. Manageable [Arc99]. Mandelbrot [BC08b]. Manhattan [CHT19]. manifold [CY21]. Manifolds [DPR03, vR04, ABT15, BG21b, CT18, FTT14, KT21b, MM21, PZ22]. many [BL12a, BP14a, BPZ22, BO13, DGP19, Kha14]. many-to-one [Kha14]. Map [Ram06, BCV21, GS17b]. Mappings [AMP04, ALM10, EL07a, LR11]. Maps [Wei07, AM08, BJM14, BS14, BC17, Bud19, Car20, GMS19, Leh22, MN14, MW08, Ric15]. Marchenko [Ada11]. Marcus [Jac10]. marginals [AJKH15, BY09]. mark [KL15]. Marked [Far11, FPT19, FPRW20, Hal22, KL15]. Market [KW00, LQ11, BM06]. markets [AW09, KNR21]. Markov [Ada08, All21, ADW21, BR16, BL12a, BP04, BE21, BKO16, BS20b, BDD09, BJN13, BCC^{+}22, BO13, BC18, BDFR15b, CR09, CDM12, CM06, CGR06, CSC08, CFW09, CK19, CL13, Dav10, DDT14, DPU15, DR14, DS07b, DMR09b, DM16, EM19, Eis05, EK22, EK23, EHMT17, FR18, Fau02, FS01, Fit98, FG03, Fuk99, Gao17, GMPW19, Ger20, HMP12, He22, HM01, JW01, JN10, KM05, KR21c, Kra20, Kur98, LLM18, LS21b, LP13, MR15, MVJ14, Oli12, Pag17, Pau15, PR11, RC09, RTW12, RR96, SCZ09, SV05, Shk13, SW11, Sza10, Tow15, Uch07, Web21, Wor99, Zer18]. Markovian [APS21, BBG19, BDG18, CO18, FR14, Jea05, Küh20, Kuw09, Nut12, OP22, RM13, Tra18, Tur15]. Martingale [BN14, BdSGP09, Dal99, HM01, Kur98, KS01, SV05, Sha99, XZ10a, BY09, CK15, CK19, CPS23, Fuk99, IKM21, JM20, KK20a, Küh18, Küh20, Kun13]. Martingales [AI09, CKMR05, EL07b, HY10, Jui18, MS98, Ose11, Sha99, AS13, BHJ22, CL05, FGL15, GJW21, JM15, Kar20, Osȩ13, OZ21, Pin06, Son12, VY16]. Marton [Pau15]. Maruyama [BPP22, DG20a]. Mass [Eng10, BDvdHS22, BHJ22, FGS16, For20, Gib08, HS17b, TV09]. mass-bound [BDvdHS22]. mass-structure [For20]. massive [CW21a, dT21]. master [DLR19]. matching [AGT22, BM22a]. Matchings [DH02, HL12, Kha14]. mated [GMS19]. mated-CRT [GMS19]. mathematical [LT15a]. Mating [EL11]. Matrices [Ada11, BGGM11, BHS09, CDMFF11, DM11, Dav10, DE09, ERSY10, GK00, HX08, OS11, Su06, TV11, Zei10, AEK^{+}14, AEK17, Au18, Aug16, BWZ09, BPZ12, BCZ18, BBGH12, BBCC21, BGG14, BP21, BS08b, BHY17, BM19a, BP09, CO22, Che17a, CL19, CL21a, CES21, CES22, CYY22, CHNR18, Cos23, DPWZ14, EY12, EKYY13, FTZ20, FL13, GT23, Gro17, GM20, HHN16, Hei22, Hus22, Jal21, Kar15, LŚ22a, LMS22, LSW12, LL21, LO21, McK21, Nem17, Ngu12, OR14, SR13, Shc22, Str19, TC14, Yan19, You13]. Matrix [BG03, CD05a, Lar15, MM18, AG13, BD19, BL22, CHZ22, DP14, FW17, Hol18, HT21, JPP19, MW19, Shc22, TC14, Web15, XYY17]. Matrix-valued [Lar15, BD19, JPP19, MW19]. max [DEM13, EI16, Mer08]. max-infinitely [DEM13]. max-plus [Mer08]. max-stable [EI16]. Maxima [BHLT01, BHT03, HKL^{+}98, Gap06]. Maximal [Mal15a, Nak22, BDvdHS20, CO21, Kuw09, OZ21, RS21a, Rob15, Stu20, ZZ13]. Maximality [Jon01]. Maximization [LQ11]. maximizer [CHHM23]. maximizing [PR12]. Maximum [BM05, DMS09, DMZ14, DZ15, Gro10, HSW09, JLML10, Lac03, Aco14, PY99]. Maxwellian [DMN15]. McGregor [Ass18]. McKean [BSY22, BK08, HW21, KN21, LS21a, TN22, Tug12]. Mean [DM23, DLT21, vdHHZ07, AGMQ22, BCD20, BMBP18, CS15, Che17b, CD12b, CTV21, DLR19, DG15, Dje22, EG18, ELL21, EFP17, FMH22, Hul15, JK14, Jos09, JT21, Lac22, MSS20, MV20, Mül17, Orr20, Pin09, RM09, SP19]. Mean-field [DM23, AGMQ22, BMBP18, Che17b, CTV21, EG18, FMH22, Hul15, JK14, JT21, MSS20, Mül17, Orr20, SP19]. meander [Pit99]. meanders [LM07a]. means [BLO13]. measurable [BR12b, NN13].

Measure [CC98, Dal99, DEF^{+}02, Hau05, HX08, Jea05, JG13, Luç11, MM17, Sha11, WY10, ADH13, AL20, BG21a, BM22a, BC18, But16, CLSF19, Fit21, Gal22, Ho22, Hon20, KL15, LSF20, Lin14, Mic20, OWW14, RS12a, SV09, SY17, Suz19]. Measure-Valued [Sha11, MM17, CLSF19, Mic20]. measured [LB18]. Measures [Ber04, BB09, Boi11, BD03, CDK06, Dav08, DG09, FM04, FG03, MT02a, MWZZ02, Pen07, RV10, AGL^{+}12, AN14, AD20, BSY22, BHS22, Bob07, BS17, CT18, CLT21, CO08, DW18b, ET14, EG18, FGS16, FG20, FS20a, Fra09, FLM20, FK21, HMK16, HKLR19, HT10a, HL07, Hil14, HTV22, HM22, HS17b, JT21, Kra13, KO08, Lac22, MR13, Rap23, Rou18, SCW22, Sch19a, Vel22, Wan20]. mechanical [KOS21]. Mechanics [EL10, EJU14]. mechanism [EX22]. Media [Gra96, Luç11, BR12a, HKN20, KK20b, Zha20]. Medium [GKW99, CP09, CD15, DP13b, SBB^{+}21]. meet [DGP19]. Mehta [EY12, TV11]. membrane [Cip13, KS20, RTW12]. Memory [MY03, CGL08, GPP13]. mercy [LS21a]. Merging [SCZ09]. Mertens [BPT16]. Mesoscopic [Lam18, Lam21b]. message [CL21b, LM08]. Metastability [AJNT22, AC96, BN21, BDG17, BMP21, CS15, Hol06, dHNT12, dHNT18, BBI09, FMNS15]. Metastable [BBM10, LLM18, MVY13, BG13b, LT15a]. Method [CZ05, DS11b, EL10, Ful05, RNP10, AH19b, AH19a, BMZ13, BRW20, BBM20, BG16, BS08b, BD13, CGR20, CPS23, Döb15, DK21, ET15, Ful12, Gau14, Ham19, JK19, KZ20, Rom18, Tan13, Web16]. Methods [BDD09, TW11, ADW21, CK15, DL22, HT21, Pau15]. Metric [BDvdHS20, ADH13, AFS20, GHS21, HR14, JJK^{+}17, Jor13, KL15, KO08, LW21, OC19, Suz19]. metrics [AS18, BP13a]. Metropolis [Vih11]. Meyer [BPT16, ŁOPS21]. microscopic [BGS16, CW21b, Hor20, Sep96]. Migration [Zho03, FGS16]. mild [BC21]. Milstein [KN21]. Milstein-type [KN21]. Minima [HKL^{+}98, Tsi06]. Minimal [Vil15, DMR09a, DM16, HBS17]. Minimax [LM15]. minimization [CCC15, LM15]. Minimum [Lac03, LZ10, Kha14, Tra15]. Minkowski [BHMK21]. minorant [AKMV20]. Minors [JN06, JN07]. Minus [Gro10]. Mirror [Pas11]. mixed [BS22b, BP14b, Che13a, FMR22, IT99, LM08]. mixed-up [LM08]. Mixing [CS20, CP19, FS01, Gan07, GMT06, HPX22, Hou17, Jon11, KP13, MR17b, Oli12, PR04, Sch19b, WSH09, XZ10b, BV16, Bia08, BCC^{+}22, Car20, CHK12, DDT14, DPP15, GL18, GM14a, GR17, Hei13, HP18, PSSS17, Sub13, TT22]. mixture [Ger20]. Mode [SS18a, SS19b]. Model [AMP02, AC96, BCS06, BRS10, BEV10, BMSV06, BT10, Bou10, COPV05, CMP10, DH06, EL11, FG05, GS11, Hut09, Lan15, LV11, MT04, Pan05, Pen09, Zho03, AKPS19, AE19, ASD21, AG13, AJNT22, BBP20, BvdHK15, BG13a, BDF18, BP14a, BDD^{+}18, BST16, BD14b, BBI09, BDT19, BO13, BK09, BMP21, CS15, CN19, CE13, CDCH16, CFW09, Che13b, CHNT17, CHSS18, CS20, Chr19, CDC13, Cip13, CK18a, CW21b, CTV21, CD14, CM22, Dal21, DH15, DvdHvLS17, DMS17, DG15, DMR16a, DG18, Dre08, DCLM18, EJS21, EFMN22, Eza13, FF14, FF15, FM14, Fle18, For19, FvdHMS21, FMR22, FJ20, FM17, GGY18, Gay12, GT17, Gra14, GS17b, Hab19, HH14, HS22, HK18, HKLR19, HJJ19, HL12, IJ21, KL22, KL21, LNO15, Leg21, Ler20, Li14]. model [Lin20, LM19b, LM07b, MdBL09, MR20, MNS16, NRS17, NXZ22, OO15, Pan14, PW23, PO09, Sch17a, Sch17b, Seg09, Sep96, SS20, SS21, Spi20, Tyk07, Val07, Wag06, Wie15, dCM21, dT21]. Models [BL05, BM09a, BC10, DM11, DG96, Gra96, GLW05, KW00, LV11, PZ22, Ton07, War10, All21, Ban18, BS22b, BM06, BP21, BP15, BELR21, BDvdHS20, BDvdHS22, BLS18, BCPW21, BF08, BF22, CT22a, CGZ06, CX13, CN18, Che13a, CD12b, CKS21, Fag12, FR18, Gap06, GL18, Gia08, GHMM22, Han14, HS22, HS12, KR21a, Kli14, LP19a, LS21b, LO20, Max06, MT12b, Ott22, QZJ14, RTW12, RR13, SP19, Wor99]. Moderate [BCR06, DE09, GW09, HSW09, LS99, Par20, Per21, Wor99, CK18a, Die20]. moderately [BCPW21, BDI^{+}19]. Modified [Hol06, Kon17]. Modular [Bav19]. moduli [Bav19]. Modulus [HS20]. molecules [DMN15]. Moment [AGL^{+}12, AI09, Bry96, CS17, CF11, DTV18, GM14a, JP18, KN07, KGS13, LNO15, LC05, MW19, AH19b, AL20, BCKH22, BRW20, BS08b, BCV21, DvdHvLS17, DVZ18]. Moment-Transfer [CF11]. Moments [ALM21, BvdHvL10, CB16, CCO20, Del03, HKL^{+}98, Lii11, Zei10, AB16b, CE13, CZ21, DP13c, GLM17b, LT15b, SS19a, Ton09]. Monetary [CDK06]. mono [Nan23]. mono-type [Nan23]. Monoids [Jea05]. Monopoles [Lac07]. monotone [HS22, JK15, Kli12, KR21b, Kwa19]. Monotonic [FJX11, PZ20]. Monotonicity [For16, Pha15, BMM20, HJ19, MX08]. Monotonous [Kam18]. Monte [ADW21, BDD09, BJJK21, DR14, JK19]. Moran [GLW05]. most [Bas23, HS15a]. Motion [ABPR11, Ata01, BD03, DS07a, DLW01, DeB08, DeB09, DPR03, DS06, FMW06, GN03, GK07, Gro10, Gro11, Haj11, Hol08, Igl05, JLML10, KW11, Law00, Pas11, ST06, vdHS10, AW97, AW14, ABT15, ABK13, AL17b, AFP15, BY09, Bal14, BB96, Bas23, BC08a, BW21, BM21, BA12, BDR22, Bon22, BD13, BH14, CDE18, Chh15, CC22b, CP20, DMN15, Eld14, EFP17, EGL22, ÉM12, FV17, GR14, GNV22, Gro13, GB14, HH06, HK18, Hon18, Hon20, Hug19, Ina16, KK12, LZ15, Law96b, Lin13, LZ12, Lou22, LPT15, Lup21, MS99, Nan06, Nil20, NP08, NRS10, NZ14, OP22, Pai16, Pit99, PT15, PT18, RGV14, Rob15, RS21b, TOR15, Web16, vSW18]. Motions [Bur09, LR11, NM04, ABK06, BR20, BZ12, BGK^{+}20, BO09, DFPSL19, DS17, DM14, DMP05, FSW15, HJ07, HJ09, KM18, LL22, MK12, Per20b, PT07, SS15, Suz19, War07]. Moving [Bas08, BOHP19, BOPT20, KS20, Kul08]. Multi [BK01, CMT15, GZ10, Bal14, BS09, BS22b, CP09, CES22, CN17, DS12b, FZ23, LW20b, Mar20, Max06, Nic21]. Multi-Dimensional [GZ10, Max06]. multi-excited [BS09]. Multi-Fractal [BK01]. multi-interface [CP09]. multi-layer [CN17, LW20b, Nic21]. Multi-level [CMT15]. multi-point [FZ23]. multi-resolvent [CES22]. multi-species [BS22b]. multi-type [DS12b, Mar20]. Multicolor [LMM04]. Multidimensional [DG14, GK07, RV10, Sza10, AJKH15, BDP22, BP14b, BK09, CM22, Cri21, HR19, Mül08, Tre16, dRM22]. Multifractal [BV04, Ber23, RV10, SY17]. multilevel [JK19]. Multilocus [EL11]. multinomial [BG13a]. Multiparameter [BOGP12, NM04, Zem08]. Multiple [Bor11, DLMX21, DG96, Sch00, Win02, BSW21, DG99, DDT14, DM13, EP23, FK21, GK14, Kar20, Ko20]. multiplication [CR14]. multiplication-invariant [CR14]. Multiplicative [AL98, MR10, BJQS15, Ber23, HHNT15, KL21, KL12, Lam21a, LPSŽ15, MR17a, Mas22, Naj20, PYZ22, Sal22, Web15, dBD09]. Multiplicatively [KM05]. Multipoles [Lac07]. multiscale [HK12, HS17b, Spi15]. multislice [FOW22]. Multitype [DG03, KM11, Val10, BM21, CR08]. Multivariate [ABCN15, BX19, DS11b, HT10b, KT17, Maj07, SY19, AH19b, AH19a, Ber19, BJN13, DK21, KM15, MT22]. Müntz [AW14]. Mutant [AMP10]. mutation [Ber15, CKS21, DG99, HP08]. mutations [BD14b, CB16, Sch08]. Muttalib [FW17]. Mutually [DEF^{+}02].

Nagumo [BK19, BK16, BMZ08]. naming [Fox18]. Narrow [Spi09]. Natural [Ben18]. Navier [CC98, CHNR22, FR01, Oda06]. Near [DMR16b, Lii11, Nol08, AO13, DG15, HMT23, RZ14]. Near-critical [Nol08, AO13]. Near-extreme [DMR16b]. near-involutions [RZ14]. near-supercritical [DG15]. Nearest [MW05, Fag12, FR13]. nearest-neighbor [Fag12]. Nearly [PR11]. Necessary [AL20, FR18]. need [PR12]. Negative [Ber97, HM04, Tan06, BK08, DVW14, KR08, Lin16]. negatively [KZ08]. Neighbor [Lii11, Fag12]. neighbors [CSW22]. neighbour [FR13]. neighbourhood [FI22]. Neighbours [BRS10]. neither [HLP16]. Nerman [IKM21]. nested [BDLSJ18, BDI20]. net [SSS09]. network [BDvdHS20, Bor08]. Networks [AL07, AL19, Ald14, AGH12, ARS08, BH23, BR06, BP08, BMZ08, CHL18, CEHRB18, DM09, DMM17, DHS19, EM14, EP23, FvdHMS21, GK14, HC20, AL17a]. Neumann [BMZ13, BBM22]. neurons [CTV21, ELL21]. Neutral [LM96, BD14b, CDM12, CB16, DPP15, For22]. Never [Lam07].

Newtonian [KST02]. nilpotent [BT15a, IKN20, SCW22]. nine [PPS18]. No [BT17, Mie01, AB16b, Ber97, Shi12]. nodal [AP21]. nodes [HJ15b]. Noise [AB18a, BJQS10, BL22, GMT05, Gau07, HM15a, HM15b, LV11, MR10, MCS03, MN18a, RV07, Stu03, War02, AM04, AKM12, AF20, BJQS15, BQSS19, BHP21, BH21, BMZ08, BHM13, BDG18, CHNR22, CHNT17, CK22, CKX09, DG20b, DG20a, DW18a, EG15, ES19, GHM22, GV20, HH06, Hei14, HT10a, HIN17, JR14, KL12, Lam21a, LS21a, LPSŽ15, MST19, Orr20, OWW14, RS13a, RS13b, Sal22, Var19, Yam23, dBD09]. Noise-induced [HM15a, HM15b, AKM12]. noises [DH19, HHNT15, NXZ22, PYZ22, SZ15a]. noisy [LM23]. Non [BV04, BTW22, BM09b, BFG99, BS22c, CD08b, CHNR18, CEHRB18, DM03, DFPSL19, DS04, DS11c, FV12, Gol18b, HMMW12, JW01, JK19, Kim06, Kol01, KST02, KOR01, LZ10, Ond10, SV05, Spi09, Tan19, XKW06, dS14, APS21, BCP18, Ban18, BC04, Bäu19, BBCC21, Ber14, BD14b, BL19, CHZ22, CS17, CW23, ET14, EG15, Foo09, Fou08, FIP15, GMW19, GM14b, GBM22, HMT23, HS17a, IPR09, IKN20, Jak97, Jal21, Jor13, JG13, JT21, KL22, KM17, LR18, LO21, ML10, MR20, Muc21, Nem17, NR17, Nut12, OS11, Ort12, RM09, SR13, SWY12, Xu20, Yan23, Zem08]. non-amenable [MR20]. Non-asymptotic [Gol18b, JK19, CHZ22]. non-backtracking [LR18]. Non-Colliding [KOR01, FV12, GM14b]. non-constant [Fou08]. Non-convergence [JW01]. non-crossing [Ort12]. Non-Determinism [XKW06]. Non-directed [BTW22]. non-divergence [CW23]. non-elliptic [HS17a]. Non-equilibrium [CEHRB18, Xu20]. Non-existence [BS22c]. Non-explosion [KST02]. non-extinction [KL22]. non-Gaussian [Ber14, CS17]. Non-Hamiltonian [Kol01]. Non-Hermitian [CHNR18, BBCC21, Jal21, LO21, Nem17, OS11]. Non-Hölderian [BFG99]. Non-Homogeneous [SV05, HMMW12, GMW19]. Non-identical [LZ10]. non-identically [SR13, Zem08]. non-integrable [HMMW12]. non-intersecting [NR17]. Non-Linear [CD08b, DM03, Ond10, DS04, BL19, ET14, EG15, FIP15, HMT23, JT21]. Non-Linearities [BM09b]. non-Lipschitz [JG13]. non-Lipschitzian [SWY12]. non-local [Foo09]. non-Markovian [APS21, Nut12]. non-neutral [BD14b]. non-reconstruction [Ban18]. non-Skorohod [Jak97]. Non-Smooth [Kim06, Spi09, KM17]. non-stationarity [Yan23]. Non-Stationary [BV04]. non-symmetric [IKN20, Muc21]. Non-trivial [dS14]. non-uniform [BC04, Jor13]. non-uniqueness [Bäu19]. nonamenable [Bia08, Hut20]. noncompact [Bak16]. Nonconservative [BG04]. nonconventional [Kif17]. nondecreasing [BDP22]. nonendogeny [RSS22]. Nonexistence [Ven19]. nonhomogeneous [DS07b]. Nonintersecting [BD12]. Nonlinear [CKX09, DG20b, Gau07, HL08, HT21, KSX17, LP19a, MT12b, BMZ13, BBM22, BHP21, BP21, BDFR15b, BF17, DMR16a, FK09, KTPZ15, KNR21, Kli14, LW20a, LZ21, MS21, OO21, RvH15, Tan13, Yoo98]. nonlinearities [LPSŽ18]. Nonliner [KSX04]. Nonmonotonic [DH06]. nonnegative [DVW14, Eis19, Pin16a]. Nonparametric [KSX04, KSX17]. Nonsimple [NRS05]. nonstationary [KOP13]. nonunimodular [Tan21, Tan22]. Norm [BG03, IT00, MR12, PZ13, SR13]. Normal [Pen09, Sza10, Bas96, BCG20, DPU15, Fan19, Pin06, Pin16a]. normalisation [Pow18]. normalised [MM18]. Normality [BCMP09, Cha11]. normalization [Kho96]. Normalized [Wan05, ERY14, Fau02, HSW09, KM12, Per21]. normalizing [BDD14]. norming [BM19b]. Norms [LŚ22a, BCP18, Str19]. Note [BHPZ11, Dev11, FV11, MR01, Mou01, ADH13, AKPS19, CZ21, LT15b, RM13, Tan23]. Novel [Hol12]. nuclear [FM20]. null [Pro12]. Number [BHLT01, BHT03, BGT11, DP05, DGLU11, GY07, HJ07, HJ09, AD15, BAKL19, BDR21, CHK19, Cou18, FF14, FJ12, LM07a, LMS22, MS12, PW23, Tyk07, Zhu22]. Numbers [AdHR11, Jac10, BC21, Kif17, Pin22]. Numerical [Dol12, DG18].

obeying [AW09]. observables [JG13]. observation [CKX09]. observe [BI20]. Obstacle [HO03, DMZ14, MX08, RS12a]. obstacles [GPV09]. occupancy [BG09, BG13a, BDI20]. Occupation [Boi11, BGT09, DS07b, PR11, SY05, BB96, HTV22, PRR20, SY17]. occurrence [DS12a]. Occurring [EL10]. Ocone [PT13]. ODEs [BFGM19]. off [BP20, Hou17]. offspring [BSW21, CK08, Lin14, Stu20]. often [DGP19]. Oldest [DDSJ10]. Olkin [HJ19]. once [CHK18]. once-reinforced [CHK18]. One [AMP10, BRS10, BB10, BHTW22, COPV05, CF11, DHP11, Del03, FM04, GL22a, KS97, KS20, LSW01, LS19, MT12a, Nis01, OW04, OR17, Pet10, PO09, Pos05, TZ11, Vos08, AP16, AKMV20, BO19, Bis19, BDT19, CT22a, DMR16a, Dre08, Eto06, FSW15, Fou08, HLP16, Hug19, Hul15, HS23, JM20, Kha14, KS17a, Kwa12, Lam21a, LZ13, LM19b, Mai07, Mic17, Muc21, NRS17, Pet12, Seg09]. One-Arm [LSW01, Hul15]. one-bit [BO19]. one-dimension [NRS17]. One-Dimensional [BB10, Del03, FM04, Nis01, Pet10, Pos05, BHTW22, COPV05, GL22a, KS20, MT12a, OW04, PO09, AP16, AKMV20, Bis19, Dre08, Eto06, Fou08, Hug19, HS23, Kwa12, Lam21a, LZ13, LM19b, Mic17, Muc21, Pet12]. One-point [LS19, OR17]. One-Sided [CF11, KS97, FSW15]. ones [Has14]. Online [Spi22]. Only [Zha09]. open [PS22a]. opening [EGL22]. Operator [DLV21, IT00]. Operators [Dal08, Led04, AH21, CW23, KN17, Küh20, Lam21a, Vel22, dT21]. Opinion [AM19]. opinions [CO21]. Optimal [AJKH15, BB10, BKO16, CES22, DP20, Igl05, KQ12, KL12, Pin09, Pin16a, Ang22, BM19b, CT07, DZ15, DLT21, Gap06, HLY14, JS22, LZ13, RS12b, SND14, Tan19]. Optimality [BW05]. Optimisation [DT20]. Optimisation-based [DT20]. optimization [DLX20]. Optimizing [Pin19, Hei14]. optional [KR21b]. Options [BT05, KNR21]. oracle [Win15]. Order [Beg08, BO10, Bor11, DM03, GN03, Lac03, LC10, Lyn00, MT02a, TW11, ACR21, BBJJ22, Ber21a, CW23, DM14, GLM^{+}17a, GSS19, JR14, KR21a, KTPZ15, LRTY20, LSW22, MW19, OPR22, SY19, SZ15b]. order-based [JR14]. Ordered [Bas06, DW10, DW12, EK08, GS07, HJ11, Mie01, HJ13, LB18, RLR22]. organized [Max06, RT09]. orientations [AJL14, DP13b]. Oriented [KSV22, BMH22, BBDS23, BCDG13, CHT19, Lac12, SS19a, SS13]. origin [Kwa12, Muc21]. Ornstein [AM15, BP20, BB15, CKM03, FK02, IT99, MY03, MS09, RSSZ19]. Orthant [PM08]. Orthogonal [KOR01, Led04, Led05, Ose11, ACR21, CCO20, LMS22, LP22]. Oscillations [AP16]. oscillators [CEHRB18]. Other [Ste99, Ang22, HS22, HLP16, JL07, Pau14, Pro21]. Otto [Men14a]. outdegrees [Thé20]. Outlier [BBCC21]. Overcoming [HJvWW20]. overdamped [Ram22]. overlap [Bon22, Jag16]. Overlapping [Jon11, Kar15]. Overlaps [CW21b, Ko20]. overshoots [MV20]. own [FIO^{+}21].

p [BS22b]. p.c.f [HY18]. Packed [KSZ01]. Pacman [Ben21]. pairing [OW20]. pairs [Döb15, DK21]. pairwise [HMP12, LS21b]. Palm [Win02, HL07]. paper [FMR22]. Parabola [Gro10]. Parabolic [Dey11, GS11, Gro11, HB09, Hau05, JLML10, Kim06, LRMT11, Lot01, Mas07, MT04, BP12, BG13b, CHNT17, CHSS18, CHKK19, CKLL13, CJK12, Dre08, FM14, FK09, Gro13, MT12a, NXZ22, Tan13]. parabolicity [BV12]. paradox [EL09, EL12]. Parallel [BW05, WSH09]. Parameter [BB10, Dar01, CHK19, DZ21, NRS10]. Parameter-Dependent [BB10]. Parameters [Fen07, CPZ22, RR13]. parametrization [Ben18]. Pareto [FN20, OR17]. Parisi [Che13a, MP20]. Parrondo [EL09, EL12]. Parseval [LP09]. part [Foo09, Ram22, IKN20, SCZ09]. Partial [KKH11, BWY15, BOPT20, CJK18, FK09, For22, KS20, KL12, WYZ20]. Partially [BS02, CP16]. Particle [Arg08, BGRS10, Bor11, CJK18, DKM01, GK06, MW05, PPPY01, Shk09, Vil11, AM15, AD20, BDD14, BM14, BSW21, BBP12, BD13, BDFR15a, BF17, CGR21, CD06, Che14, CGG22, DMP05, EZ16, FO18, FTZ20, GS20, GM14b, GSL21, HNX19, HS23, KL22, KR21c, Kra13, Kua14, Orr20, PRT22, RS21b, SS18b, SYY21, Wan22]. Particles [DP05, AS12, BO13, Die20, Sar15, ST17a, Wie15, dHNT12]. Particular [Doh99]. partition [Ban18]. partitioning [AGMQ22]. Partitions [Bas06, GS07, Pet11, CL13, EJU14, FPRW20, Hei14, Lab14, Ort12]. partner [BDF18]. parts [FKHL19]. Passage [Zha09, ADS16a, AT16, BT15a, CN19, DHS14a, DW16a, DR15a, EGvdHN20, Fit21, Gan22, GM08, GT17, HS20, Leh22, Nak22, RP12, RT18]. passing [CL21b]. passive [HS23]. past [MT12b]. pasting [Wer21]. Pastur [Ada11]. Pata [CD05a]. Path [BDP22, BDG08, CK18a, DGP13, FZ10, Fd00, GMT05, Hua21, LMM04, Mül17, Mun11, ABK06, AD20, Bat21, Bav19, FL18, Kal21, LR18, MR18, Wer12]. Path-by-path [BDP22]. path-dependent [Kal21]. Path-distribution [Hua21]. Path-properties [DGP13]. path-repellent [ABK06]. Path-space [CK18a]. Paths [Fd06, BKN19, BF08, BD12, BK23, CMT15, CO18, CL05, EGvdHN20, HTV22, ŁOPS21, NNT08, NR17, PP15, PZ20, RX13, SY14]. Pathwise [And11, Guf18a, BH21, BG21b, FMNS15, GIP16, PP15, RS13b]. Patlak [BF17]. pattern [AGH12, HP08]. patterned [BBGH12]. patterns [For19, WW09]. pausing [Hor20]. PDE [BMZ13, BBM22, DMS09, Lyn00, MX08]. PDE’s [Nil20, BEP09, BFGM19, CB08, DMZ14, HM11, HK12, Tan13, Yoo98]. PDMP [Mon15]. peacock [BHJ22]. peacocks [Jui18]. peaks [LM07a, Su20]. peanosphere [GHS21]. Pearcey [ACvM11, HHN16]. Penalization [BMZ13]. penalizations [SV09, TY23]. penalized [LP08]. Penalizing [Pro12]. Percolation [AB20, BC08b, BC10, DG13, DG15, Fag10, Gar04, Gra96, GH09, HH05, HBS17, Hol06, Jon01, KSZ98, KSZ01, LSW01, MW10, MN14, She10, Zer11, Zha09, AČ20a, AČ20b, ADS16a, AB18a, AdlRG21, AS19, AT16, BEGG14, BT15a, BBAP06, BT17, BPT17, Ber14, Ber15, BBDS23, BCDG13, BY13, BGH^{+}14, BC22, CPZ22, CCK22, Chr19, CD14, Cou22, DHS14a, DW16a, DHS17, DM23, DSWB15, DR15b, EGvdHN20, Fit21, Gan22, GV20, GMPT17, Gol18a, GT17, GS17a, GHS21, HB09, HMP12, HS20, HM14, Hul15, Hut20, Hut23, Jan16, Jan09, JW18, KSV22, Kol22, KMT12, Lac12, LNO15, Leh22, Lis22, LM19b, MPR19, MN21, MVX20, Nak22, Nol08, Ott22, Ray14, Ric15, Riv21, RT18, SS19a, SS13, SY14, Sob12, Szn16, Tan23, Tei09, Tur07, Tyk07]. percolation [Van19, Yao13]. Perfect [FH10]. performance [LM15]. perimeter [MW18]. period [GL22a]. Periodic [Tak97, CDC13, DFS21, GL22a, HKN20, KP16, PT07]. permanental [Eis13, FR14, MR18]. Permutation [Zei10, BCZ18]. Permutations [BU11, BT19, BBFS20, Bor22, Fér13, Kam18, Muk16, Per20a, Pin22, Sam17, SZ15b, Tar15, Tra15]. permuton [Bor22]. Perpetual [SY05]. Perpetuities [AI09, FH10, INS21, IMN23]. Persistence [BPZ22, DHKW22, CSW22]. perspective [HJ07, HJ09, LT15a]. Perturbation [ERSY10, MP07, MR13]. Perturbations [AF03, LM96, NRS17, CGL08, DSB19, OR14, XZ23]. perturbative [Hei13]. perturbed [ES19, OPR22]. Pfaffian [Pet11, TZ11]. Pfaffians [FTZ20, MS13]. Phase [AMP02, DNPR19, De 01, DMS17, HMN11, MV16, RvH15, Ber21a, BELR21, FL13, GL18, HK18, Hua20, Lac12, LM15, LO20, MŞ19, Nej21, RSS22, Tag16, Web15, vSW18]. Phases [BGdH11, COPV05]. Phenomena [LS11b, FI22]. Phenomenon [CSC08, BP20, BHP21, MS12]. phenotypic [BB18a]. phi [TC14]. phi-entropy [TC14]. phylogenetic [BLO13]. PIDEs [KTPZ15]. Piece [Kry99]. Piece-Wise [Kry99]. piecewise [ADW21, Die20, LP13, RTW12]. piecewise-deterministic [ADW21]. pillowcase [RZ14]. Pinning [AB18b, BT10, BP15, CMT15, DMR16a, Gia08]. Pitman [FF14, FF15]. Place [DeB08, DeB09, Lac07]. Planar [BHLT01, BHT03, BS01, Law00, Wei07, AM08, AFL23, Ame21, BCR06, BR18b, BO09, BMM20, Ben21, BT21, BS14, BC17, Car20, CDCH16, CDC13, GS17b, HM15a, HM15b, HMMN15, HP17, Jos09, Leh22, Mas09, MW18, MN14, MW08, Ray14, Ric15, Ric18, Riv21, Yao13]. Planck [BR12b, GBM22, Zha13b]. Plane [DEF^{+}02, Law00, Luç11, BJM14, BDP22, CT15, Gra14, GM17b, Maf20, Pen20, PRU20]. planes [DSWB15]. planted [Ban18, MNS16]. plaquette [CS20]. Plaquettes [GH10]. plasticity [BB18a]. Player [FW08]. Plus [LV11, Mer08]. Poincaré [BCG20, CKNP21, Liu20, Oça21, SY19]. Poincaré-type [CKNP21]. Point [AR06, Ata01, BHM18, BT10, CFG09, CX13, HL07, IMP07, LW01a, SS15, Uch11, vdHS10, Bas14, Ben18, BY13, CHHM23, Die16, DMR16a, FZ23, Fit21, FPT19, FKHL19, Gei00, HS13, Jan16, JSS17, LS19, OR17, PKTZ15, QW18, RM09, Sch09, TOR15, ZZ15]. Point-interacting [SS15]. Point-shift [BHM18]. point-to-line [Fit21]. Points [BR07, BK05, CMN97, CMN98, DPRZ99, DPR03, AZ15a, AKTKK19, BS21, Cip13, DLMX21, FY20, Jeg20, JK15, KS19, KM15, Law96b, Muk16, OW20, Red17, Ros23, SSS09]. pointwise [DL22, MT12a]. Poisson [AS02, AM04, AD14c, AZ15b, BP16, BBGT20, BC04, BJKM10, BO09, BO10, BD14a, BST22, BU11, BM20, BGT09, BP14b, BT15b, CGG22, DZ21, Dav08, DHP11, DM23, DMN13, DVZ18, ET14, Fen07, GMT05, Gei00, GZ10, HH06, Hau05, Her20, HLP19, IJ21, Ken20, KN17, LRP13, MR10, MWZZ02, MNV11, PS21, Pin16a, Sch05, SY19, Tyk07, War02]. Poisson-Type [BO10]. Poissonian [CB16, HS20, KS17a, VFV15]. Poland [Leg21]. Policies [Kry99]. Policy [BW05]. Pólya [JP18, MM17]. polygon [Ham19]. polygons [AHS20, Ham19]. Polyhedra [vR04]. polymatroid [HW10]. Polymer [CMT08, Ton07, CP09, CN18, HS20, OO15]. Polymers [Kur08, AO13, AZ13, AY15, BS20a, Bat18, Bat21, BTW22, BHTW22, BS22c, CIW19, CN17, FSV14, IV13, Nic21]. polymorphic [BB18a]. Polynimials [Led05]. Polynomial [Bou10, FJX11, KOR01, RT10, BK18, BM14, Car20, CCO20, Cos23, CLSF19, Do21, EX22, LPSŽ18, PRT22, Web15]. Polynomial-Growth [FJX11]. Polynomially [SS00]. Polynomials [Led04, Zei10, AKPS19, AP21, Ass18, ACR21, AL13, BBCC21, But16, Do21, FS17, GSS21, IKM16, KS19, LP22, OW20, Red17]. polyominoes [RP12, ST22]. polytopes [GW18]. Pooling [BW05]. Population [Stu03, BCM19, BL12a, BD14b, BČD16, BLS18, CCSWB22, GPPW16, Gup12, Kli14, MT12b, OC19, PW23, QZJ14, Sch17a, Sch17b]. Populations [DGW04, BB18a, GdHO22, GL22b, Nan23, TV09]. porous [BR12a, Zha20]. portfolio [DLX20]. Portmanteau [BP14b]. Posedness [Gau07, MR10, GdHO22, HP21, PYZ22, Tre16]. position [Pes22]. position-dependent [Pes22]. Positive [Bet10, CMN97, CM06, LC10, Mie01, Par22, ABK06, BKO16, CC13, CMN98, CD05b, CD08a, CM20, GM08, KK18, Pro21, Sim14]. Positively [KZ08]. Positivity [GT17, KW10, LW20b]. Post [Wag06]. Post-gelation [Wag06]. posterior [KO08]. Potential [DS11c, Ber21a, DW18a, EX22, FM14, GA08, KS04, MM15, Uch20]. potential-theoretic [Ber21a]. potentials [CR14, DLR20, JM22, KK20b, KKO13, KN17]. Potts [GGY18, GL18, Val07, War10]. Power [BOPP21, KP21, CKS21, DÜ22, MS13, MVY13, NP08, PS19]. Power-law [KP21]. power-like [DÜ22]. Powers [DS11b, Dub18, FJ22, FJ23, Ful12]. Precise [BD18, Tan06]. predator [Bor14]. predator-prey [Bor14]. Predictable [CD12a]. prediction [BKO16]. Preferential [Der16, BL12b, DM09, EM14, FI22, FJ20, Jor13, LO20, Sén21]. presence [HMR14]. prey [Bor14]. price [PP15]. prices [CCC15]. pricing [BPR15, HJvWW20]. prime [Mas22]. Principal [BA12, Ber97, CH05, Hu00, Fle18]. Principle [BM05, BR05, BM09a, DMS09, HM01, MCS03, AC17, Aug16, BCC19, BPZ22, Bis19, BCV21, CC13, DZ22, DDM12, DMZ14, DZ15, Do21, GPP13, GMW19, GT07, GB14, HS17a, PO09, Rou15]. Principles [GMT05, And21, CK18a, FF15]. Priors [Ber21a]. Probabilistic [BH21, CMMS23, HJ07, HJ09, KTPZ15]. Probabilities [BC08b, CP99, HH05, KL05, Uch07, AW97, AS13, ABCN15, BLV19, BPT17, CDCH16, DP20, DVW14, JJK^{+}17, LW21, Tan22]. Probability [BHPZ11, BG11, CLSF19, Dev11, PT13, Sha11, SW19, ST06, AJL14, BC14, BSY22, Bob07, Bor14, CLT21, CEY12, FL15, GR15, Hil14, JW18, KW21, LSF20, MR13, Sob12]. Problem [Ata01, CP99, LS21a, LW01a, MP05, AJL14, AGT22, AL09, BHJ22, CS17, CCGR22, CM22, CPS23, DZ15, Dje22, Dom22, DP13c, FIP15, Hob15, Jan19, KQ12, KR20, Lin16, ŁG14, Tug12]. Problems [BM05, BB10, DTV18, KS01, Puh04, BCC19, CMT15, CK15, CK19, DFPSL19, FR12, HW10, KK20a, Küh18, Küh20, Kun13, Kur98, Pau14, RS12a]. Process [ACvM11, BCS06, BS02, BC08b, BP06, CFG09, Dar01, DDSJ10, FF98, FG03, GM11, Jac10, LC10, Lam07, MNV11, Nis01, PR11, Sav08, Sch05, Sha11, Spi09, Ste99, Tay07, Val10, vdHS04, vdHS10, ALS15, AEGR15, AD14d, ADW21, BHM18, BMH22, BR18a, BCG^{+}13, BB15, BK08, BD14a, BM14, BDG17, BI20, BEK21, BKLM19, Bor08, BH14, BHM13, CS15, CD06, CX13, CSW22, CDE18, CZ21, CP20, CFT15, CS98, DL22, DPTV12, DGP13, DPP15, DLZ14, DY20, DJM16, EI16, EVY20, EL16, FV12, FV17, GS20, Gay12, Gei00, GLM17b, GRV13, HUVV22, HL14, Hua20, HS15b, IT99, IKM21, Kam18, KS20, KR21c, Kra20, KR20, KKPW14, Kyp16, LMCV12, LR20, LZ12, MM18, MM20, MPX22, MŞ19, MVY13, MV16, Nag12, OT19]. process [PT18, QW18, Ram22, RSS20, RLR22, SS19a, Sch19b, Stu20, SND14, TN22, Vil15, Wie15, Zem08, Zha14, dS14, Ger20]. Processes [AT09, AL07, AL19, Arc97, Arc99, AB02, BB11a, BB11b, BK01, BV04, Beg08, BO10, Beg05, BdSGP09, Ber04, Ber97, BB09, BD03, CM06, CLR05, CSC08, CDK06, DG03, DPRZ99, Don04, Eis05, Eng10, Fag10, FJX11, Fen07, FZ10, FM07, FT11, GMT05, HM01, HLP10, HPZ11, IMP07, Kli11, KM05, Kry99, Kry04, Lac03, Lac07, LC10, Lej02, Lim10, LA09, LS99, LP09, MY03, Mat11, MS09, MWZZ02, Men07, Mie01, RV07, SS00, Son06, Stu03, Uch07, Whi07, Win02, Ada08, AL21, AAJ21, AH18, AL09, BBGT20, Bal14, BMS13, BS15, BCM19, BL12a, BP20, BEV12, BOGP12, BGK^{+}20, BB15, BKO16, BO09, BS20b, BH21, BPR12, Ber15, BST22, BM20, BD14b, BY13, BR22, BPW18, BO13, BDG08, BDFR15b, CR09, CCDO21, CR08, CB09]. processes [CD05b, CD08a, CGR06, CKM03, Che17b, CGL08, Col09, CHK19, CP19, CHK17, DG14, Dad17, DG99, DDT14, DPU15, DB12, DWW19, DT20, DW16b, Eis13, Eis19, EK22, EK23, EI16, Eto06, EGW12, Fag08, FGS16, FJS15, FL18, Fit98, FR14, FPT19, FPRW20, FH13, FMM19, Fou19, FKHL19, FIP15, Fuk99, Gao17, Ger20, GLT21, Guf18a, Guf18b, Hal22, HJ07, HJ09, Han14, Hay21, HKN20, HMT23, HT10a, HS13, HZ19, Jan16, JP18, JPP19, JR14, JK15, KT21b, KRST13, KK20b, KS20, KZ08, KOP13, Kra20, KL18, Küh18, Kur98, Kuw09, KPS12, KK18, Kwa12, Kwa19, KR08, Lab14, Lar15, LT15b, Lev17, LW20a, LZ21, LSW22, LP13, MSS20, MM17, Mal15b, MR12, MR18, MS21, Mic20, MVJ14, MM21, MV15, Muc21, Owa17]. processes [PPSV21, PS22a, PS12, PS18a, Pin16b, PZ22, PT18, PS19, Pro21, RC09, RSSZ19, Rie17, RTW12, RR96, RS12b, SS12, SS18a, SS19b, Sch09, SY17, Sha15, SZ21, Shi17, Spi15, SS14, Su14, TY23, TT20, TV09, Tre16, Wan22, Wer21, ZZ15, Zer18, AL17a]. Product [BG03, EG18, KP13, Kra13, LMS22, Nem17, XYY17]. Products [Eva02, FS01, Gil11, OS11, AN14, CYY22, Cor21, GS17a, Jal21, KK20b, SCZ13]. Profile [CKMR05, Ged19, GMT06, CHNR18, MW08]. profiles [HS20, Hus22]. progeny [Bor14]. programming [BCC19, DLT21]. progressions [BAKL19, ZZ13]. progressive [JL15]. Prohorov [dCM21]. Projections [GJW21, CD12a]. projective [DDM12]. Prokhorov [ADH13, GM17b]. Prokhorov-uniform [GM17b]. Proof [SV05, BM19b, HJ19, MT22, RX13]. Propagating [AKM12]. Propagation [CP21, Die20, All21, BR12a, Dje22, ELL21, GBM22, Kha14, LT22]. Properties [ABBG10, AB02, BC10, GK07, Mas07, MT02b, Zha06, AD19, AP12, BH21, BDG08, BS21, CT18, DGP13, Eld14, For16, GLM^{+}17a, Han14, HBS17, KPS12, Lod22, MM20, MR18, Ott22, RS13a]. Property [Lej02, Sha99, SW11, XZ10a, BOS22, DGK21, DM16, DL14, HKY14, LZ12, PL16]. Proportional [BT05, CB09]. protected [HJ15b]. prove [AKM12]. prudent [PST17]. Pruning [ADV10]. Pseudo [Beg08, Lac03, Lac07, LC10, Nis01]. Pseudo-Processes [Beg08, Lac03, Lac07, LC10]. pure [DH19]. PushASEP [BF08, Fit21]. PushTASEP [Pet20]. puzzle [MR17b]. PWIT [EGvdHN20].

quadrangulation [GM17b]. Quadrangulations [Bet10, BR18b]. Quadratic [AID07, Beg05, BBL11, BM09b, CB08, PKTZ15, TTV14, AGT22, BW12, HR19, Liu20, Luo20, LP14]. Quantile [KN07, AFP15]. quantisation [AN14]. Quantitative [BB18b, DP17, EM19, FMH22, GPT22, Lac22, NXZ22, RR96, Mou12]. Quantization [DS06, LP14]. quantum [AFS20, AHS23, DKRV17]. Quasi [BN11, Coh12, DMS09, FM07, Lam07, Ram22, STZ11, AEGR15, CV23, DG99, DR14, DFS21, HC20, Nut12, Oça21, Vil15]. quasi-equilibria [DG99]. quasi-equilibrium [AEGR15]. Quasi-linear [DMS09]. Quasi-Logarithmic [BN11]. quasi-Monte [DR14]. quasi-periodic [DFS21]. quasi-stationarity [CV23, Oça21]. Quasi-Stationary [Lam07, Ram22, HC20, Vil15]. Quasi-sure [Coh12, STZ11, Nut12]. Quasiclassics [Kol01]. Quasiderivatives [Kry04]. quasilinear [DMZ14]. Quastel [Yan23]. quaternionic [BP09]. Quenched [DF20, Luç11, Rou15, Spi15, AdlRG21, AP16, AB18b, GP12, Jos09, RASY17, RS13a]. Quermass [CD14]. Queues [KOR01]. QuickQuant [FH23]. Quickselect [Gol18b]. QuickSort [FH19, FJ12]. QuickVal [FH23].

Rademacher [KT17, Pin12, RNP10]. Radial [BGK^{+}20, APW20, CMT19, VFV15]. Radius [MW08, Su20]. Ramification [BK23]. Random [ADV10, Ada11, ABBG10, ABOPS16, AFL23, AMP04, AL07, AL19, Ale11, AI09, AC07, Ass18, ARS08, AdHR11, ACdCdH19, BM11, BHLT01, BB11a, BR07, BGGM11, BYY07, BIK07, BG11, Ber11, BT10, Ber06, BOS22, Bet10, BU11, BvdHvL10, BB09, BGdH11, BČD16, BP07, BHdS^{+}19, BGT11, BG03, Bou10, Bry96, BDI^{+}19, CP99, CT22a, CFG09, CHNR22, COPV05, Čer11, CLR05, CF11, CC11, CGL08, CB08, CN00, CO08, CK08, CCFR09, CK14, DH19, DW10, DM09, DTV18, DGLU11, DP13b, DH02, DE09, DSB19, EK08, EL10, EJU14, Eva02, EL07a, Fag08, Fag10, Far11, FF98, FG03, Fol11, GP12, GHT03, Gil11, GM05, GKW99, GM15, HM04, HO03, Hau05, HH05, HMN11, HdHS^{+}15, HKL^{+}98, HX08, HRW07, HPZ11, HS15b, HJ11, Jag16]. Random [Jea05, KS97, Ken20, KS03, KPX00, KOR01, KSX04, KL05, KPV04, Law96a, LL04, LRMT11, LLS06, LL11, LS11a, LMM04, LR11, Luç11, MS12, Mat11, MRS16, MR08, MV15, MT97, MT02b, Mun11, Naj01, Neu07, NRS05, NV13, OS11, OS20, Pak98, Pak99, Pel99, Pen07, PR04, Pet10, PY96, RV10, Ros97, SE06, SCW22, Sch06, Sel06, Sha99, ST06, ST17b, SP14, Ste99, Stu03, Tak97, Ton07, Uch10, Uch11, Vol03, Win08, Win10, Zäh01, vdHHZ07, Abe18, Ada20, ABBHK22, AP16, ACJ21, ALR23, Ald14, AL17a, AY15, AJL14, AKMV20, AEK17, AS19, AGT22, AH08, AD14d, AD15, ADS16b, And21, AA23, AGH12, AP21, AS20, Au18, AL13, Bak16, BN15, BS22a, BCP18, BH23, BCM19, BAKL19, BCZ18, BS08a, BS09, BCR06]. random [BOPP21, BBGH12, BO09, BD14a, BARS19, BLV19, BP21, BC12, BT21, BM14, BD08, BP13b, Ber19, BS20c, BTW22, BHTW22, BPT17, Ber18, Ber14, BV16, BBDS23, BJM14, BT19, BELR21, BHLN15, BST16, BBI09, BCDG13, Bis19, BJK13, BL22, BBFS20, BS08b, BHP22, Bou13, BHY17, BM19b, BC22, BW17, BP09, Bud19, BD18, BDI20, Bus16, But16, CPZ22, CO22, CCK22, CMT15, CC13, Che17a, CL19, CL21a, CW21a, Che14, CHL18, CD15, CES21, CK18a, CLT21, CHK18, CHK19, CHSW22, CHNR18, Cor21, CP23, Cou18, CHK12, CS98, CSŠ22, DLMX21, DNPR19, DPWZ14, DZ21, DS12a, DGK21, DDS18, DS12b, DW12, DVW14, DW19, DH15, DH21, DF20, DGP18, DGP19, Do21, Dol16, DFS21, DGP09, DEM13, DCLM18, DKLT22, DW18b, ET14]. random [EK22, EK23, EV22, EKYY13, Eto06, Fag12, FR13, FZ12, Feh17, FS17, Fér13, FIO^{+}21, FTZ20, Fle18, For19, FW17, FY20, FL13, FM17, GPV09, GJ17, GP20, Gay12, GS16, GKKR18, GMW19, Gib08, Gil16, GN21, GSS19, GSS21, GT23, GHMM22, GLM17b, GL22a, GW18, Gro17, GB14, GR17, GPP12, GM20, GPT22, GS17b, GHS21, Haa17, HMK16, HH22, HS22, HPX22, Hei13, HJP22, HM22, Hol18, HS17a, HJ15a, HZ19, Hor20, HI23, HMMW12, HS15a, HNX19, HL14, HJ13, IPR09, IKM16, IKN20, IJ21, Jal21, Jan09, JJK^{+}17, Jeg20, Jos09, JS18, KS19, KT21a, Kam18, KR21a, KSV22, Kha14, KK20b, Kol22, KP13, KGS13, KZ08, KZ14, KP16, KN17, KSX17, LM07a, LRP13, LR22, LT15b, Law20, LR18, Leh22, LM23, LS14, LS19]. random [LS22b, LRTY20, LMS22, LSW12, Lod22, LL21, LP15, LM19b, LP22, LO21, Mal15a, Mas09, Mas22, MS15, McK21, MW18, Mél19, MPX22, MV08, MW08, MV20, MN21, MVY13, MM15, Mou12, MV16, Muk16, Mül08, Mül17, MSC15, MX07, Naj20, Nan23, Nem17, NN21, Ngu12, NNP13, OO15, OR14, OW20, OR17, OPR22, Par22, Pau14, PRR20, Pèn21, PRT22, Per21, Per20a, Per20b, Pet12, Pha15, PL16, PO09, Pin09, Pin16a, Pin13, Pin18, Pin19, PY17, PT18, PS18b, PS22b, PR12, PZ20, QZJ14, RS21a, RS09, RR22, Rap23, RASY17, RT09, RS13a, Red17, Ric15, Roe08, RW18, RP12, Rou15, SX19, SCZ13, Sam09, Sch19b, SR13, Shc22, SY14, SHW20, Shi12, Spi15, SBB^{+}21, SZ15b, Str19, Su14, Sub13, Szn09, Tag16]. random [Tak21, Tar15, Ton09, Tou09, Tra15, Tra18, TC14, Tsi06, Uch20, Ver19, VT08, Wan14, Web15, XYY17, Yan22, You13, Zem08, Zha13a, dS14]. random-cluster [DCLM18, HS22]. random-to-random [Sub13]. Randomised [Jor11]. randomized [BCC19, LP19a, LŚ22a]. randomly [BMH22, CHT19, KM12, LM08, Tur07, Zha14]. Range [BIK07, CFG09, DP05, Fag10, FK02, Win10, vdHS10, AAJ21, AS20, BMH22, CCK22, DGK21, DW18b, For22, GKKR18, HS22, Hul15, IT99, KMT12, Kul08, MSC15, Seg09, SS18b, SP14]. Ranges [BP99]. Rank [BGGM11, DPWZ14, GSL21, Jan19, Mai07, OR14]. rank-based [Jan19]. Rapid [BH19, TT22]. rapidly [AAJ21]. rare [Ber15, BHLN15, Hol18]. Rate [BHPZ11, BS08a, BR11, CC22b, Igl05, Jal21, AdlRG21, Bas23, BF20, BDP15, CKS21, Fou08, HL14, LM15, Ott22, RS21b]. Rate-Optimal [Igl05]. Rates [Ben21, DMR09a, DDM12, RA08, AZ13, AAJ21, CL13, EM19, Gia08, GBM22, Nag12, Pin16b, RR96]. Ratio [WY10]. ray [Die16, LST21b]. Rayleigh [Ken20]. reach [MW05]. Reaction [MR10, EP23, GK14, HC20, Hay21, MW20a, Sal22]. reaction-diffusion [Sal22]. Real [AS08, Do21, BS21, BP09, CES21, For20, IKM16, LMS22, LP22, PR19, Shc22]. realisations [FW17]. Rebirth [GK07]. recentered [Aco14]. recombination [DPP15]. Reconstructing [LMM04]. reconstruction [Ban18, FR18]. Record [GS07, SSW04, FN20]. Records [Gne07, HT10b]. recovery [LM15, LM08]. rectangles [CMSS16, CDCH16]. rectangular [GT23, Gro17]. Recurrence [BS09, BS01, BS14, BHP22, CFG09, CF11, GHMM22, KK20b, RT10, Sel06, Zer18, CZ21, DHS14b, KP21, Pin16b, Sha15]. Recurrent [HM01, Tel01, Win10, DGP18, HP08, KP16, Pro12, TV09, Uch20]. recurring [CSW22]. Recursive [BO10, GM05, MSS20, BB15, Ber14, GM15, HJ15a, Lod22, Sén21]. redistribution [EL12]. reduced [GS17b]. reducible [BM21]. reduction [DW16a, HL12]. reference [Bat18]. Refined [MN18b]. Reflected [AL17b, Bur09, HO03, Kli12, KR21b, PM08, Ram06, Rho10, Sav08, ALR23, BMZ13, DQS16, FL18, Slo14]. Reflecting [Ata01, BD03, KSX04, Pas11, CK18b, HH06, KSX17, Lin13]. Reflection [And11, PM08, Spi09, BBG19, Bas96, Fou19, GV17, KS20]. Regenerative [GI12, Lam01, PRW14]. regime [BDvdHS22, BC22, CT22a, DG15, FSV14, JM15, KL21, NN23, Nej21, OWW14, RS19, Sha15, TT22]. regime-switching [JM15, Sha15]. regimes [BP14b, Völ14]. Region [Pan05, BM22b]. region-stabilizing [BM22b]. Regions [BHLT01, BHT03, DH06, Pin16b]. regression [Wor99]. Regular [DEM13, HMN11, KM05, AČ20a, AČ20b, BS09, BST16, BHS22, Col09, CK23, DRM14, Gal07, Gil16, Kal21, LZ23, LT15b, MV16]. regularisation [DG20a]. Regularity [BK16, BK19, FFPV17, Hau05, KRST13, MN08, SZ15a, Wer12, AF20, Bal14, BFGM19, BG21b, CHZ20, CKLL13, DL22, HY18, HTV22, KL12, Pha15]. Regularization [BCP20, Nil20]. Regularizing [Mas07]. regularly [RS21a]. regulating [BEV12]. regulation [BČD16, Gup12]. Reinforced [Ger20, Sel06, AA23, Col09, CHK18, CZ21, LR18, MR08, RS09, Tak21, Zha14]. Related [AR06, ALM10, BGT11, BD03, FS01, Haj11, Hu00, Sza10, BW21, BO09, BGT09, DG14, ÉM12, FTZ20, HH06, Has14, HBS17, LSW12, MT12a, MV08, Pro21, Zer18]. relations [Van19]. Relative [CGR06, BDFR15a]. Relativistic [IT00]. Relaxation [BFN11, KP13]. removal [GL22b]. Renewal [FT11, Gia08, GPP12, Ton07, AB16b, AB18b, Ber19, BDP15, CSZ19, CD19, GI12, Hol12]. renewals [AB16a, Ber19, HUVV22]. renormalisation [BK16, BK19]. Renormalizability [DKRV17]. Renormalization [Hon18, SF06]. Renormalizations [Zäh01]. Renormalized [BR05, BCR06]. Rényi [BHLN15, CFT15, Kif17, Kol22, RT09]. repellent [ABK06]. repelling [Che14, Dum12, VT08]. Replacement [DM11]. replica [Pan14]. Representation [CD05a, CB09, FJX11, GLW05, MS98, BCC19, BOGP12, BdT14, CD12a, Guf18b, KTPZ15, Mél19, Pic06, PT13, RS12a, Son12]. Representations [Bry96, Ger20, LA09, ČR22, Coh12, CJK18, DT20, FS20b, MR17a, PT07]. Reproducing [Jor11]. Repulsion [CMT08, Nit18]. Rescaled [Che19, Kli11]. Rescaling [CP20, EVY20]. Residual [Lam01]. Resistance [CHL18, CHK17]. Resolution [DS06]. resolvent [CES22, Kra20]. Resource [BW05]. respect [BR22, FM21, NRS10]. response [EL14]. restaurant [RLR22]. restricted [CCC15]. Restriction [Maf20]. Result [AI09, Dav08, DS04, DJQS13, Dje22, Kli14]. Results [BMSV06, HUVV22, KM11, PM08, AP21, Ban18, BLS18, EP23, Eza13, FMR22, HH14, IPR09, PPS18, PRW14, RS13b, Sch17a, Sch17b]. Return [HH05, Tan22, Sob12]. Reversal [DS07a, Bel20a, HMT23, LP13]. reversals [BKLM19]. reversibility [FMNS15]. Reversible [CFG09, Sha11, BDG17, DPU15, PSSS17, Shk13]. Reversing [BW17]. revisit [But16]. revisited [BP16, BM19b, CPS23, Men14a]. revolving [BHP22]. reward [HLY14]. Reznikoff [Men14a]. Riccati [BD19]. Ricci [BG21b, EHMT17]. Richardson [DH06]. Riemann [GNV22]. Riemannian [ABT15, BGK^{+}20, BG21b, CY21, Fd00, Suz19, vR04]. Riesz [Ber21a]. Riffle [Ful05]. right [FH19, Wer21]. right-tail [FH19]. Rigid [MR17a]. Rigorous [Sch17a, Sch17b, FMR22]. rising [DP13c]. Risk [CDK06, JS22, CCC15, LM15]. risk-minimization [CCC15]. Risk-sensitive [JS22]. risks [HJvWW20]. Robinson [OP13]. Robust [All21, Gan07, EM14]. rock [FMR22]. rock-paper-scissors [FMR22]. role [CD12b]. Root [BCS06, BH23, FR18, Nas22]. Rooted [Wei07]. roots [Do21, FS17, IKM16, OW20]. Rosenblatt [DL22]. rotational [HMR14, MR09]. rotationally [HJP22]. rotations [Car20]. Rotators [Luç11, JK14]. Rough [Dey11, FZ10, Nil20, BCD20, BK23, CHNR22, CHNT17, CO18, CL05, DM22, Ina16, Lej09, Lej22, NNT08, RX13, Rom18, Wer12]. Row [Arc99]. Rowlinson [AE19]. ruin [CHK19]. rules [GL22a, HLY14]. rumor [Bor08]. running [Gol18b]. RWRE [GR15, RS19].

S.D.E.s [Fou08]. S.P.D.E [Dal99, DM03]. Salem [AP21, AP23]. salesman [Pau14]. same [BY09]. Sample [GMT05, HRW07, MR18, Su06, AEK^{+}14, BPZ12, BPZ22, CYY22, Gro17, Hei22, Yan22, dCM21]. sampler [AD20]. Samples [BHLT01]. Sampling [Gal07, BHLN15, FF14, FF15, Pit99, PY17]. sand [Bur15]. sandpile [MdBL09]. Sandpiles [ST17b]. Sasamoto [BF22]. sat [Pan14]. Satisfying [BB11b, CF11]. sausages [GW09]. Scale [Ald14, BM11, DG96, Win02, BM20, CS20, Dal18, DG99, DMM17, IJ21, KR08, MS99]. scale-free [IJ21]. Scale-invariant [Ald14, BM20]. scales [EP23, FG19, GK14, SW14]. Scaling [AHS20, AA23, BCM19, BS20c, Bet10, BT19, BvdHvL10, BF22, DKLT22, GS17b, GM17b, JV11, LW21, PST17, RW18, VFV15, Wei07, vdHS04, AO13, BS15, BJM14, CP16, EVW20, GHS21, LS22b, MVX20, SS13, SS18a, Van19, Xu20]. Scattering [BS02, LSW22]. sceneries [DGP09]. Scenery [LMM04, MP07, BP13b, DF20, GB14, GPP12, Pèn21, Per21]. Scheduling [BW05]. Scheme [ML10, AJKH15, BG09, BPP22, BDI20, Che13a, DG20a, DG18, KM17, KN21]. Schemes [BFN11, Lam01, BK09, Dol12, DW16b, FF13]. Schensted [OP13]. Scheraga [Leg21]. Schramm [FZ23, Wer12]. Schrödinger [CCGR22, DW18a, Gau07, Get99, IT00, KN17, Lam21a]. scissors [FMR22]. scores [BM22b]. scotching [Bor08]. SDE [AP12, DP14, LRTY20]. SDE’s [BDP22, And11, BMZ13, BCKH22, BSY22, BP12, CCDO21, CHZ20, CZ05, DS17, DB12, GHM22, GL23, GLM^{+}17a, HKY14, HW21, Hua21, IdRS19, JK19, KS17b, SZ15a, Zha11, Zha13b]. Search [CKMR05, HJ15b, HJ15a, Pin19]. Second [AL19, AN14, DM03, GN03, KTPZ15, LRTY20, MT02a, OPR22, AL20, CD06, CW23, DM14, GLM^{+}17a, SY19]. second-order [DM14]. Sectorial [XKW06]. seed [BCKWB20, GdHO22, LP19b, Nan23]. seed-bank [GdHO22]. seed-banks [Nan23]. Seen [KSZ98, AEGR15, Cos23, Die16, EI16, FL18, LMM04, SS14]. Segel [BF17]. segregation [Hay21]. Seiberg [DKRV17]. selection [AL17b, BEK21, BCPW21, CK19, DG99, Dur14, EVY20, GSL21, Sch17a, Sch17b]. selective [CEY12]. Self [BEV12, BC08a, BR11, BG04, Bou11, BS21, BKM21, CM06, Eng10, Fau02, FT11, HM04, HY10, HSW09, PY96, VT08, Wan05, AH19b, AH19a, BCR06, BKO16, BBM20, BT19, BLM09, BDG08, CR09, CMT15, CP16, Dad17, DB12, Dum12, DW16b, Ged19, Ham19, HT10a, Ho22, KM12, Kho96, KK12, Lau12, Max06, Neu07, Per21, RT09, SZ21, SSH22]. self-adjoint [Ho22]. self-attracting [KK12]. self-avoiding [BT19, CMT15, Ham19]. self-bounding [BLM09]. Self-Decomposable [HY10, AH19b, AH19a]. Self-Interacting [BR11, BBM20]. Self-Interaction [Eng10]. self-intersection [BCR06, Lau12]. self-normalization [Kho96]. Self-Normalized [Wan05, Fau02, HSW09, KM12, Per21]. self-organized [Max06, RT09]. Self-regulating [BEV12]. Self-repelling [VT08, Dum12]. Self-Similar [CM06, FT11, PY96, BG04, BS21, BKM21, HM04, BKO16, CR09, Dad17, DB12, DW16b, Ged19, Neu07, SZ21]. Self-similarity [BC08a]. self-stabilizing [HT10a]. Semi [Bou11, CL05, Bor22, Cou18, Fuk99, Nic21, PR20]. semi-Baxter [Bor22]. semi-contractive [PR20]. semi-discrete [Nic21]. semi-infinite [Cou18]. semi-martingale [Fuk99]. Semi-martingales [CL05]. Semi-Self-Decomposability [Bou11]. semicircle [EKYY13, FKK22]. Semicircular [BHPZ11, CDMFF11]. Semiclassical [Dav08]. Semigroup [IT00, CT18, HT21]. Semigroups [Bou11, Lam21a, Mas07, Son06, BR16]. Semilinear [AAC09, HM11, Mas07, BEP09, BBM22, GH19, KL12, MX08, RS12a]. semimartingale [ČR22]. Semimartingales [Bas08, FM20, GN03, ADMU20, CG18, FM21, PZ13]. Seminorms [BO11]. Seneta [BM19b]. sense [GHS21]. sensitive [GV20, JS22]. Sensitivity [GK14, AB18a, BL22, For16, HP18]. separable [Yan19]. Separation [HLP10, BOS22, HLP16]. sequence [BB18a, CP23, Zhu22]. Sequences [Ton07, CHK12, DS12a, DRM14, DHKW22, DE06, Kul08, Lin16, MS12, PS18b, Tow15]. sequential [BJJK21, DDT14]. Series [GHT03, Igl05, LA09, CD12a, MS13]. Server [BW05]. sesquilinear [QZJ14]. Set [HRW07, Win08, AČ20a, AČ20b, AS19, AH08, BMM20, ČP12, CB09, PT15, Szn16, Tar15]. Sets [Gar04, Lam01, NM04, PY96, Sha99, AS18, AL20, BAKL19, Ber21b, CK23, DR15b, Fra09, JJK^{+}17, Nit18, PS22a, Tsi06]. setting [Bak16, FMR22]. several [BBGH12, DGP18, QZJ14]. sewing [BL19, Lê20, Lê23]. Shadow [BHJ22]. Shaken [STV22]. Shape [HLP19, BEGG14, EJS21, Eza13, FM14, MW18]. shapes [Hor20]. shared [BGH22]. Sharp [BMSV06, BG13b, BELR21, BBI09, CGZ06, FTZ20, Ose11, Osȩ13, OZ21, Kol22, MN21]. shattering [Wie15]. Sheet [TV03, XKW06, BY09]. Shell [BM09a, HJP22]. Shepp [HJ19]. Sherrington [Che13b, Pan05]. shift [BHM18, HL07, PT15]. shifted [Pin22]. Shocks [BS02]. Short [Ina16, For22, SS18a, SW14]. Shot [GMT05]. show [FR01]. shuffle [Sub13]. Shuffles [Ful05, Jon11]. Shuffling [Nor10]. Sided [CF11, KS97, BLV19, FSW15, Law20, MM20, Shi12, Son06]. sigma [DMR16a]. signal [CKX09]. signed [LST19]. Similar [CM06, FT11, PY96, BKO16, BG04, BDG08, BS21, BKM21, CR09, Dad17, DB12, DW16b, Ged19, HM04, Neu07, SZ21]. similarity [BC08a, SSH22]. Simple [BS02, BIK07, Ber06, Boi11, BJJK21, CC11, CO08, CCFR09, Fol11, HH05, Law96a, LW01b, Abe18, AFP15, BDLSJ18, Bud19, CCK22, CS98, GdHM09, HS15b, MPX22, Nag12, PZ20, RSS20, RX13, Rom18, Sch19b, dS14]. simplices [HJP22]. Simplified [ČR22]. simply [DDS18]. Sims [BK09, Che13a]. Simulated [WSH09]. Simulation [FH10, HL12, DZ21, Eto06]. Simultaneous [Sch00, BCKWB20, Sar15]. Sine [ERSY10, AD14c]. Sine-Kernel [ERSY10]. Single [AMP10, Uch11]. Singular [BM05, KS01, MT04, SCZ09, Zha11, APS21, CL19, CL21a, DLR20, GL23, GT23, HKY14, Hua21, Ngu12, Vel22]. Singularities [FR01, DNPR19, HMMN15]. singularity [Cri21]. SIRSN [Ken20]. Site [KSZ98, BB18b, Bas23, GHS21, SS14]. Sites [Zho03, AD14d, CS98, Gib08, HS15a]. six [Lin20]. Size [BP06, Gei00, OWW14]. size-biased [Gei00]. sizes [FJ22, FJ23]. SK [BK09]. skew [AN14, Bor22, ÉM12]. skewed [TOR15]. skewness [BR18b]. Skorohod [CT15, Jak97, ŁG14, NY19]. Skorokhod [Bas96, FIP15, Hob15, Ram06]. slabs [DKNS13, GR15]. SLE [FZ23, WW13, AS08, APW20, AHS23, Ben18, CW21a, FL15, Kar20, TY20, WZ17, Yea22, Zha06, Zha21]. Slepian [PT15]. slope [Ton09]. slow [GI12]. Slowdown [Ber11, AP16, Pet12]. slowing [Pes22]. slowly [GM14a]. Small [AH05, BG09, DM14, Don04, ERSY10, GHT03, Hab19, Kol01, LR10, LLS06, MS99, AL09, BR06, BHP21, BPP22, BDG18, Dur14, HT10a, HS17b, JR14, LT15c, MM15, Orr20, OWW14, RS13a]. small-noise [Orr20]. Small-time [AH05, Don04, Hab19]. smallest [GT23]. Smooth [And11, CMN97, Kim06, Sei10, Spi09, AL21, CMN98, GW18, KM17, MM21]. Smoothing [BK05, EX22, KM15, LS21b]. Smoothness [BMM20, Fou08, Kry04, Oda06, CKLL13]. snails [GL22b]. Snake [AS02, War02, Ham19]. snakes [AW97]. Sobolev [BP16, DP05, FOW22, HTV22, MX08, Ond10, Ste21, Vel22, Zha11]. social [BH19]. Sock [Ste99]. Sock-Sorting [Ste99]. soft [DLV21, GPV09]. Sojourn [Lac03, LC10]. Solid [Pos05]. Soliton [dBD09, FG20]. Solution [SS10, AB13, BZ12, GL22a, MX08, Pan14, Yan23]. Solutions [AW09, FR01, KS01, LW01b, Oda06, RA08, BMZ13, BR12a, BLM15, BMBP18, BKM21, DH19, DGV17, FR12, Fd98, GM14b, Gro20, HJT20, Hob15, HNX19, KK20a, Küh20, LZ23, Nil20, RM13, RS12a, Sal22]. solvability [Yoo98]. solvable [Luo20]. Solving [BCD20, HR19]. Some [AM08, BK01, CC11, DM03, Eza13, Gar04, ML10, PZ13, Puh04, SSG11, Shk13, Ton07, WW09, Zha06, Ada11, AL09, BP21, DS07b, ÉM12, Fér13, Foo09, Fou08, IPR09, MST19, RS09, RW18, Rou18, Tan23, Yoo98, Zem08]. Sorting [Ste99, AGH12]. SOS [HKLR19]. soup [AS19, LST19]. soups [FR14, vdBCL18]. souvlaki [CFG17]. Space [DG96, FJX11, Fd00, FG96, GZ10, HJT20, MP05, Ose11, BS20a, BJQS15, BQSS19, BLO13, BK16, BK19, BF08, BP14b, BT15b, CKS12, CK18a, CKS21, DH19, DG99, DG13, DVZ18, DMR09b, ET15, FPRW20, HY18, Her20, HIN17, Jak97, KR21b, KL18, LRP13, LL22, Par22, Pet20, Tyk07, Wan20, Yam23]. space-like [BF08]. Space-Time [DG96, HJT20, BQSS19, DH19, HY18, HIN17, Yam23]. Spaces [AH05, FM07, FS01, Kry00, LLS06, LC05, Mas07, MP05, Ond10, Sei10, ADH13, ALM21, AN14, AD20, CM17, Coh12, CL13, Dal18, DMN13, EM19, FM20, HT19, JJK^{+}17, Jor13, KRST13, KL15, Kun13, LSF20, Lê23, MN18a, Pet08, PZ22, Suz19]. spacetime [PT15]. spacings [VS15]. Spanning [LZ10, LMS08, HBS17, Tan21].

Sparse [Cos23, BJK13, BL22, BHY17, BDI^{+}19, CL19, DPWZ14, GT23, LL21, MRS16, ORS20]. Spatial [BEV10, BU11, CHNT17, CKNP21, DLW01, DG03, Dur14, Eng10, GdHO22, Hay21, Nan23, Oda06, SS19a, Wei07, Win02, vB09, Ald14, AB13, BT19, BEK21, CDE18, CP20, DP20, DW18a, EVY20, Gup12, LS06, PRT22, Wag06]. Spatial-segregation [Hay21]. Spatially [Dal99, MCS03, DH13]. SPDE [CT20, DS04, Kim05]. SPDEs [BK19, HNX19, Alt21, BJQS15, BP12, BG13b, BK16, BM09b, CS17, CHKK19, CKNP21, CJK12, DH19, GH19, Hau05, Kim06, Kli11, Kry00, MR01, Par22]. Special [KR08, SSS09, AH08, KKO13]. Species [EP98, FM03, Kli11, BS22b, GM08]. Spectral [BWZ09, BOGP12, BBL11, BHS09, Bou10, CCK22, CH10, Fag12, GMT06, HX08, KW11, KKO13, KM05, KK18, Kwa12, Led04, Muc21, PS22a, SX19, SW11, BCC^{+}22, CYY22, FL13, JPP19, Nag12, Pau15]. spectrally [BK08, CM20, KR08, Pro21]. Spectrum [BG08, CO22, DE09, CB16, Mél19, NN21]. Speed [BP13a, Lim10, BR12a, BvdHK15, BH14, CHK18, GS20, GN21, LM19b, MPX22, Pha15, PRU20, PR12]. speeds [BvdHK15, GM08]. Sphere [Su01, AFL23, AZ15a, Chr19, Gal22, Su20]. Spheres [GH10, Cou22]. spherical [Adh22, AZ15a, BS22b, Che13a, CW21b, Dub21, HT19, Ko20]. spherically [Per20a]. Spiked [CDMFF11, LV11, QZJ14]. spikes [Ang22, EJS21]. spiking [CTV21]. Spin [Adh22, vdBCL18, AC17, BS22b, Bäu19, BT17, BDT19, CR14, Che13a, CW21b, CGL18, Ko20, Lin20]. spin-glass [Bäu19]. spinal [HMT23]. spine [AD14b]. Spins [Lis22, Pan05, BK09, Dom22]. Split [Mun11, Hol12]. Splitting [HT19, CB16, LB18, Tan13]. Spohn [BF22]. Spontaneous [HMR14, MR09]. Spots [Ata01]. Spread [AD14d, vdHS04, BS22a, HJJ19]. Spread-out [vdHS04]. Spreading [KPV04, BM21]. square [CS20]. squares [RM09]. Stability [AF03, AB02, DM16, GL23, LM96, MV20, PM08, BD19, BDFR15b, DKT18, HR19, Hei14, LM23, MN18a, SCZ09, vSW18]. stabilizability [MdBL09]. Stabilization [PRZ20, AKM12, HM15a, HM15b]. stabilizing [BM22b, CT22a, HT10a]. Stable [BP99, BBC^{+}05, BB09, CLR05, CH10, DHP11, DPRZ99, FMW06, FH13, GT07, HMK16, Her20, HX08, LR11, MY03, MNV11, Men07, RSSZ19, Son06, XZ10b, Arc20, Bal14, BHP21, BPW18, CK14, CSŠ22, DMZ08, EI16, EG15, FS20b, GH15, HJ07, HJ09, KM16, KKPW14, KK18, Kyp16, MT22, Muc21, Nas22, PS22a, PT07, PS19, Pro21, SCZ13, SY17, Sim14, Wor99, ZZ15, dRM22]. stable-driven [dRM22]. Stable-like [Son06, SY17]. stacks [KP16]. staircase [BD12]. stalkers [Wei09]. Standard [Bou10]. Starr [BK09, Che13a]. starting [MW05]. State [DG03, Lam07, BMS13, BDFR15b, DMR09b, EM19, FH13, FMM19, Fou19, GGY18, KRST13, KL18, LW20a, LZ21, MS21, ST17b, Tag16]. state-space [DMR09b]. States [DEF^{+}02, Bäu19, BGT09, CEHRB18]. stationarity [CV23, HMT23, Oça21, SSH22, Yan23]. Stationary [BV04, FM07, FR01, Gro20, HT10a, HLP10, KW10, Kra13, KS01, LZ23, Lam07, Mar20, MT97, OW04, Oda06, ST17a, AL21, BOGP12, BOHP19, BC12, BF22, CN18, DPU15, FSW15, HMK16, HC20, Hei13, MdBL09, PL16, Pin19, Ram22, RS12b, SX19, SS12, SZ21, Vil15]. Statistical [EL10, EJU14]. Statistically [FR01]. Statistics [AZ15b, BWZ09, FG05, Kur08, AD14c, BHY17, Dub21, Fér13, JS18, Kar15, KN17, MW19, NR17, PRR20, SHW20, Web16, Zha22]. stay [CC13, CD05b, CD08a, DVW14]. steady [CEHRB18]. Steepest [CZ05]. Stein [AH19b, AH19a, BRW20, BG16, CGR20, Dal08, DS11b, Döb15, DGV17, DK21, EL10, ET15, Ful05, Ful12, Gau14, PS18b, RNP10, Web16]. Stein-Type [Dal08]. Stepping [Zho03, For19]. Stepping-Stone [Zho03]. Steutel [ALM10]. stick [BDI20]. stick-breaking [BDI20]. Sticky [HMMN15, PR23, BR20, Bas14, GV17]. Stieltjes [FZ07]. Stirring [Bur15, DPTV12]. Stochastic [AM04, AR06, AF03, AW09, BM05, BJQS10, BKN19, BP04, BFGM19, BCG21, Bes99, BGdH11, BSW10, BR22, BMZ08, Bor11, Bur09, CK20, Cho05, CD08b, CKS21, CL98, Dal99, Dal18, DMS09, EP98, FZ07, FM21, For22, Fuk11, Gau07, GV17, HM11, HK12, Haj11, HR14, HO03, HM14, HIN17, HHNT15, JN10, KS97, KKH11, KW10, KMP06, KST02, KL21, Lê23, Lej02, LPSŽ15, LPSŽ18, LM96, Lin13, Lot01, LW01a, MR10, MCS03, MV00, Ond10, PR20, PP11, Pet11, RNP10, RA08, Rho10, RS12a, SS10, Sei10, Sha11, STZ11, Ste99, Stu03, TW11, Var19, War10, WYZ20, Zha11, APS21, ADMU20, AB13, AH21, BO19, BQSS19, BCC19, BWY15, BR12a, BP08, Bas14, BZ12, BMBP18, BK08, Bel20b, BDW17, BOS22, BGS16, BHJ22, BV12]. stochastic [BHM13, BS17, CC98, CD12a, ČR22, CL14, Che19, CY21, CK22, CKLL13, CGL08, CN17, CJK18, CHK17, DG14, DP20, DMZ14, DG15, DZ15, DMN13, EGL22, FF13, FFPV17, Fer16, FK09, GT07, Gro20, GM17a, GIP16, GNZ21, GK14, HRW12, HY18, HS14, HC20, HJT20, HP21, Kal21, KOP13, KL12, KR20, KN21, Kur07, LNO15, LP19a, Lê20, LMCV12, LM19a, Lin20, LW20b, MSS20, MM18, MT12a, Mer08, MW20a, MB14, MN08, NN23, Nut12, OO21, OWW14, Pau14, PYZ22, PS21, Pic06, RTW12, RR13, Rom18, Sal22, SSS13, SZ19, SWY12, Su20, TTV14, Tur15, VY16, Wie15, XZ23, Yoo98, ZRLB19, Zha20, dRM22, ST17b]. Stokes [CC98, CHNR22, FR01, Oda06]. Stone [Zho03, For19]. Stopped [Pel99]. Stopping [Ang22, AKTKK19, BB10, CT07, DLT21, Gap06, HLY14, JS22, KQ12, LZ13, ZRLB19]. story [BS20c]. Strahler [BDR21]. strange [Pin18]. strategies [Jan19]. stratified [DP13b, HK20]. Stratonovich [BK18, CT15, DJQS13, NRS10]. Stratonovitch [Fd98]. stretched [IV13]. stretching [BHTW22]. Strict [De 01, Law00, Ose11, Pet11, LW20b]. Strictly [DMZ08]. String [MT02b]. strips [HMMW12]. Strong [BGRS10, BMBP18, Ber19, CL14, CCFR09, DM11, DRM14, GM14b, Mal15b, MR15, Sha99, Vos08, BN21, BG16, BBP12, Bor22, CD19, DDM12, LM19a, Sal22]. strong-Baxter [Bor22]. Strongly [FG03, Tel01, GPPW16]. structural [PRW14]. structure [BDvdHS20, For20, HK12, Muk16, OP22]. structured [GPPW16]. Structures [BN11, BK01, BK16, BK19, BC20]. studies [Mic17]. study [BCKH22, CV23, CY21, DL22, OO21]. Sub [Bou13, PR20, Tel01, BGK^{+}20, BS20c, GSS21, Gro17, Spi20]. Sub-ballistic [Bou13, BS20c]. sub-critical [Spi20]. Sub-exponential [PR20]. Sub-Gaussian [Tel01, Gro17]. sub-Riemannian [BGK^{+}20]. Subadditivity [TC14]. subballistic [DGP19]. Subcritical [AEGR15, SS14, CC22a, CP23, Lau12, NN23, Stu20]. Subdiffusive [DHS14a, Fag12]. subdivided [TV09]. Subdivisions [Beg05]. subexponentiality [SV09]. Subexponentialiy [Mat23]. Subgaussian [AZ13]. Subgeometric [ADW21]. subgraph [Owa17]. subject [MST19, TV09]. sublinear [CN19, Coh12, DM09]. Sublinearity [Cou18]. Submitted [FF98]. subordinate [KM18, MK12]. subordinated [Kul08]. subordinates [OZ21]. Subordinator [MNV11, CSZ19]. Subordinators [BP99, DS11c, DR15a]. subsequence [Tra15]. subsequences [DH21, HI23, Kam18, Sep96]. substitution [BBFS20]. substitution-closed [BBFS20]. subtree [FJ22, FJ23]. subtrees [DHS19]. suffice [AKTKK19]. Sufficient [CC11, WSH09, AL20, FR18, KK20a]. Sum [HRW07, Wan14, BCZ18, CL21a, FJ22, FJ23, GM20, Ho22]. summation [Zem08]. Sums [Arc99, HSW09, Kul08, LL11, Pel99, SSW04, BOPT20, BM22b, BCG18, BCG20, Del09, Dol16, KM12, Kif17, Pin16a, Pin18, Roe08]. sup [BDR22, DW18b]. sup-functionals [BDR22]. sup-measures [DW18b]. Super [BP04, FMW06, Hol08, vdHS10, AW97, CP20, DMP05, Hon18, Hon20, Hug19, IdRS19, KS17b, KN21, Pin06, Pin22, RSSZ19, Sal22]. Super-Brownian [FMW06, Hol08, vdHS10, AW97, CP20, DMP05, Hon18, Hon20, Hug19]. super-clustering [Pin22]. super-linear [IdRS19, KS17b, KN21, Sal22]. Super-Markov [BP04]. SuperBrownian [CDE18]. Supercritical [AI09, Bud19, Fag10, She10, AS19, BEGG14, BBDS23, BC22, CK23, DG15, GMPT17, Gol18a, GLM17b, Hut23, IKM21, LST21a, SS13]. supermarket [LM07b]. supermartingale [BPT16]. Supermedian [FG03]. superprocess [Kli14, MT12b]. Superprocesses [DLW01, FM03, Kol01, BGT09, DH13, MX07, RSZ18]. superquadratic [RM13]. superquantiles [BCG21]. Superreplication [NN13]. supersolutions [DM16]. supOU [GLT21]. Support [BCMP09, BSW10, Del03, Gau07, Kal21, CO18, LZ12, TY20, Unt19]. suprema [Ada08, LT15b]. supremum [EI16, SS12]. Sure [BR05, LM96, LV11, Ada20, AP21, AP23, Coh12, IV13, Mas22, Nut12, STZ11, Tak21]. Surface [FF98, Ass18, HLP19]. Survival [AMP10, BG11, GPV09, BS22a, KS17a]. survive [CK08]. Susceptible [DY22]. SUSY [Shc22]. sweeps [CEY12]. switching [BCKWB20, JM15, Sha15]. symbiotic [DY20]. Symbol [SS10]. Symmetric [CO08, DPRZ99, KN07, Men07, RSS20, Roe08, Abe18, Ber18, DH19, Fuk99, GdHM07, GdHM09, GRV13, HJ07, HJ09, IKN20, JK15, Kwa12, MPX22, Muc21, Ngu12, Pan14, Shc22, TT20, dS14]. symmetries [ADMU20]. symmetry [HMR14, MR09]. Synchronization [JK14]. System [BBM10, BW05, GK06, Lyn00, LW01b, MW05, AM15, CHNR22, CGL18, Die20, FK21, KR21c, MVX20, OWW14, Yam23]. Systems [Arg08, AS12, BGRS10, BP04, BGdH11, Bor11, De 01, DKM01, KS97, KM11, KKH11, KST02, LA09, MR01, NY09, NY10, Pet10, Shk09, TZ11, Vil11, XZ10b, ABK06, BBJJ22, BHP21, BSW21, BBP12, BGT09, BPT16, BDFR15a, CSZ19, CGR21, CR14, CT18, CMMS23, CGG22, DP13a, DMP05, ELL21, FTZ20, GM14a, GM14b, GSL21, HJ07, HJ09, HNX19, Ko20, Kra13, Kua14, MT22, Mer08, Orr20, PYZ22, ST17a, SS18b, SYY21, XZ23, vdBCL18]. Sznitman [BG08, GM05, Sch12].

tacnode [FV12, FV17]. Tagged [GK06, GS20, Wan22]. Tail [BHS22, Bou10, Dir15, JL07, LL21, WY10, Ada08, BTW22, DZ22, DF20, FH19, Pin16a]. Tailed [CGGR10, Čer11, BOPT20, BGG14, BDvdHS20, BDvdHS22, BPZ22, CO22, Fle18, GMPW19, HZ19, HMMW12]. tails [Aug16, BB11a, BM14, Bob07, CG20, DW12, DHKW22, Gro17, HMP12, PRT22, Pin12]. Talagrand [Riv21, SZ19]. tampering [Pin13]. Tanaka [ŁOPS21, PP15]. Tandem [KOR01]. Tangent [SSH22]. target [Pin19]. TASEP [BB18b, FO18, GGS21, Pes22, Vet15]. Taylor [BZ12]. TCP [BCG^{+}13]. technique [AH21, BHLN15, DGV17]. telegraph [Lin20]. Temperature [Pan05, ABK06, BT17, CDC13, FG19, Ott22, Völ14, dHNT12]. Temperatures [AC96, Bon22]. Tempering [WSH09]. Temporal [CHSS18, GL22a]. tensor [CYY22]. Terminal [CB08, APS21, JS22, LRTY20, RM13]. Terms [BBL11, DLX20, WYZ20]. ternary [HJ15b]. tessellation [GKT22]. tessellations [HT19, KT21a]. text [LM08]. Their [MT02a, BE21, Ber04, Ber15, BS21, DG03, DGLU11, EGW12, FL18, HJT20, OZ21, Pag17, Uch10, Zei10]. Theorem [BSW10, DMS09, FJX11, Naj01, SV05, Ame21, AGH12, ABK13, BG21a, BR18a, BOPT20, BGG14, BDD14, BPR15, BDP15, But16, CD19, CO18, CD15, CT22b, DMR09a, DLR19, DS12b, GH19, HB09, Hei22, HLP19, HI23, HJ13, JT21, Kur07, Lam21b, Mou12, Owa17, TY20, AP23, BJ09, BDD09, Dol16, Far11, HJ11, Lev13, NY09, Pan05]. Theorems [Ale11, BHLT01, BB11b, CCFR09, DW10, GHT03, GLW05, Ste99, Wan05, AB16b, AH18, BLO13, Ber19, BPR12, BBDS23, CT22a, CSZ19, CR08, Che13b, Col09, CHSW22, DP17, DVZ18, EK22, EK23, EL09, FR14, GPP12, Her20, HMMN15, INS21, IMN23, IKN20, Jan22, Kar15, KT17, ŁG14, MR12, MS99, NXZ22, PPSV21, Pèn21, RSSZ19, RTW12, SYY21, SBB^{+}21, TT20, BBL11, PP11]. theoretic [Ber21a, Chh15]. Theory [HM11, KKH11, KM05, MR01, AB16b, BOHP19, BC12, BST22, CM20, CL05, DKRV17, DLR19, FW17, FY20, FKHL19, GA08, GI12, GM15, HJP22, Hol12, Kwa12, Kwa19, Mél19, Muc21, Wer12, Yan23]. thermal [KOS21]. Thermodynamic [CR14, CT22a]. Thick [DPRZ99, Jeg20, Ros23]. Thin [ACJ21, HJP22, CMSS16]. Thin-shell [HJP22]. Thinnings [Sch05, Sch09]. Third [BvdHvL10, DvdHvLS17]. Thoma [BO13]. Three [AC96, Ros97, SW11, Yam23, BK16, BK19, GdHM09, LS19, LS22b]. Three-dimensional [Yam23, GdHM09, LS22b]. Threshold [BW05, BC08b, Hol06, BI20, FR13, FS20b, HLY14, Kol22, MN21, MSC15, Seg09, Tur07]. thresholds [MNS16, Tan23]. Tightness [Aco14, DFF^{+}21, Ros23, FM17]. Tilings [Nor10, LM23, RP12]. tilted [DSWB15]. tilting [SW19]. tilts [CIW19]. Time [AAC09, Bel20a, BGdH11, BT05, CDK06, CB08, CHK17, CH05, DM03, DS07a, DG96, DPR03, FZ10, FG96, Fol11, GS11, GMT06, GM11, HMT23, Jon11, Lac03, Lac07, LC10, MW05, PR05, Ste99, TV03, Uch11, Zha09, ACJ21, ALR23, AJKH15, Ang22, AH05, AD20, AT16, BQSS19, Bia08, BF08, BO13, BPP22, BF17, Bus16, CN19, CCR05, Car20, CCDO21, CGR06, CMMS23, CGPV13, DH19, DG99, DS17, DS07b, DM14, Don04, DR15a, EP23, ÉM12, FZ12, FL18, FG19, FMR22, GPV09, Gao17, GR14, GMPT17, Gol18b, GT17, GPPW16, GBM22, GK14, Hab19, HY18, HMP12, HPX22, HJT20, Hon20, Hor20, HIN17, Hug19, Ina16, KQ12, KP13, KKPW14, LRTY20, LP13, Lup21, MM20, MS21, Mer08, MR17b, Nak22, Nan06, NZ14, Pin16b, RR96, SV09, SCZ09]. time [SZ21, Sub13, TY23, Tur15, Web21, Yam23]. time-changed [DS17, DM14]. Time-changes [CHK17]. time-dependent [BO13, Pin16b]. time-inhomogeneous [CCDO21]. time-marginals [AJKH15]. time-nonhomogeneous [DS07b]. Time-reversal [Bel20a]. time-scales [EP23, GK14]. Time-Space [FG96]. Times [BBM10, BR05, Ber97, CLR05, EP98, FS01, Gra96, Hu00, Law96a, Men07, PR11, PR04, SY05, Abe18, ACJ21, AH08, BB96, BCR06, BGT09, CGR06, CSC15, CP19, CT07, CHHM23, CHK12, Din12, ERY14, FMNS15, HP18, Hon18, JK15, Lau12, LT15c, ŁOPS21, MR12, MT12a, MS99, Oli12, PSSS17, PP15, Pit99, RP12, Sch19b, Tou09, Zha18]. Toda [GM22]. Toeplitz [LSW12]. tolerance [HS13]. top [BL22]. Topological [IMP07, CDCH16]. topologies [FS20a]. topology [GM17b, Jak97, Yea22]. toral [Lev13]. tori [JJK^{+}17]. Torpid [WSH09]. Torus [AC96, Win08, ASD21, CGPV13]. Total [BCG^{+}13, BPP22, HLP16, JK11, BR16, BCP20, BX19, BHP21, Bor14]. Totally [LB18, OT19]. tournament [JS18]. trace [LLM18, QW18]. Traces [AT09, DS11b, Ful12]. tracking [JM15]. Tractable [FJS15]. Tracy [BPZ12, FO18, OO15, Vet15]. trading [CX13]. Traffic [Au18, BW05]. trail [BP13b]. Trait [BSW21, CKS21, Kli14]. trait-dependence [Kli14]. Trait-dependent [BSW21]. trajectories [KS17a]. Transaction [BT05]. Transfer [CF11, BP14a, Shc22]. Transform [BK05, AFP15, KM15, LSW22, LPT15]. Transforms [GHLT03, AW14, KO08]. Transience [BBAP06, CFG09, Hut23, KST02, Pin16b, PRU20, BS09, DHS14b, FL15, GHMM22, Mül08, MW20b, Sha15, Zer18]. Transience/recurrence [Pin16b]. Transient [BGdH11, DPRZ99, Win10, AD15, BS08a, GKKR18, Tei09]. Transition [AMP02, BK01, BBM10, BLV19, Ber11, De 01, Gar04, Mas07, Uch07, BELR21, DLV21, FL13, Hol18, LO20, MVJ14, MŞ19, MV16, Nej21, RSS22, ST17b, Tag16, vSW18]. transitions [Ber21a, DMS17, GL18, Hua20, RvH15]. transitive [Tan21, Tan22]. translated [BC04]. Transport [FF13, Goz12, Sam17, AJKH15, BHJ22, CHNR22, Tan19, Win15]. Transport-Entropy [FF13, Goz12, Sam17]. Transportation [BE21, Rie17, Liu20]. Transportation-cost [Rie17]. transportation-variance [Liu20]. transports [HMK16]. Transpositions [Ber11]. transverse [Gan22]. trap [ABK06, Fag12, Gay12]. trapped [ABOPS16]. Travelling [Lyn00, TW11, Pau14]. traversal [Nak22]. trawl [PPSV21, TT20]. Tree [ADV10, Dev11, Far11, HM04, BS09, BB15, BBFS20, BL12b, BW17, CB16, DGP13, DPP15, DKLT22, FIO^{+}21, GGY18, Guf18a, Guf18b, HKLR19, Lod22, MSS20, Nas22, PRW14]. tree-valued [DGP13, DPP15, DKLT22, Guf18a, Guf18b]. Trees [BDLSJ18, CP99, CKMR05, CH10, DH02, DÜ22, GM05, Hol08, LZ10, LL11, MW10, Mun11, Tak97, Vol03, Wei07, AČ20a, Abe18, AD14b, AD14a, AD19, ABBHK22, AO13, And21, AA23, ARS08, BHMM19, BLO13, Bäu19, Ber14, Ber15, BELR21, BGH^{+}14, Bor14, BDR21, BS21, BFMM13, CFW09, CM17, Col09, CHK18, CHK19, Cou18, CK08, DDS18, DS12b, DHS19, FR18, FJ22, FJ23, For20, FI22, GS20, GGS21, Gei00, Ger20, GN21, GGR19, GH15, GM15, Guf18b, Haa17, Her18, HJJ19, Hol12, HJ15b, HJ15a, HS15a, Hua20, Hut12, Jan22, Joh19, LZ23, LB18, Lin14, LP19b, MPR19, MR20, MW20b, NN21, Pag17, RW18, Rou18, SS19a, Sén21, Ste13, Su14, Thé20]. Triangle [DG09]. Triangular [Arc97, Arc99, KSZ98]. triangularly [Luo20]. Triangulating [KM16]. triangulation [Ric18]. triangulations [AHS20, CMSS16, Car21, GHS21, Ray14, Rou15]. Trickle [EGW12]. Trickle-down [EGW12]. Tricolor [SY14]. Tridiagonal [GK00]. tries [Jan22]. Trigonometric [Igl05, LA09, AP21, AL13, IKM16]. trimmed [Mal15b]. Triple [Ros97, Sar15]. trivial [dS14]. Triviality [HRW12]. Truncated [Cha11, DPTV12, KN07, Dub21, LMS22]. Tsetlin [WW09]. Tubes [Doh99, Spi09]. turning [EVW20]. turnpike [CCGR22]. Tutte [Leh22]. Two [AS18, Ata01, BR07, BPR15, BP99, Čer11, Che14, CMP10, DH06, DDSJ10, FZ07, Fen07, FW08, Gra96, Lam01, LW01a, MR08, Sak09, Shi12, Son06, Tur15, Uch10, Wei07, Zha21, AB16a, AGT22, AZ15b, BK18, BLV19, BDW17, Bon22, BHP22, BH14, BPP22, Bur15, CDM12, CGPV13, DZ21, DP13b, Gan22, Gol18a, GM20, HH06, Hay21, Hei13, HJ15b, KR21b, KP21, Law20, MM20, MR09, NN23, Nol08, Pin09, Ros23, Val07, Zha14, dHNT12]. two-color [Zha14]. Two-curve [Zha21]. Two-Dimensional [Čer11, CMP10, Gra96, MR08, AGT22, AZ15b, BK18, CDM12, CGPV13, DP13b, Gan22, KP21, NN23]. Two-Dimensions [FZ07]. two-parameter [DZ21]. Two-Player [FW08]. Two-Point [Ata01, LW01a]. two-protected [HJ15b]. Two-sided [Shi12, Son06, BLV19, Law20, MM20]. two-speed [BH14]. Two-Type [DH06, Wei07]. two-value [Pin09]. Two-valued [AS18]. Type [AI09, BO10, BHS09, Dal08, DH06, GK06, HSW09, KKH11, Lac03, LS99, Naj01, Wei07, AM19, BOPP21, BB15, BM06, Bel20b, BOS22, BBP12, BK23, BF17, BKM21, CHZ22, CCDO21, CKNP21, DS12b, DB12, EG15, FKK22, Gan13, Hab19, HP21, IPR09, KZ20, KL22, KKO13, Küh20, KN21, Luo20, Mar20, Men14b, Nan23, NZ14, PYZ22, SZ19, Slo14, Zha13b]. type-I [HP21]. types [BL12a, dHNT12]. typical [BKN19, BCG18, PP15, SS14].

UFG [CCDO21]. UFG-type [CCDO21]. Uhlenbeck [AM15, BP20, BB15, CKM03, FK02, IT99, MY03, MS09, RSSZ19]. UIPQ [Die16]. ultrametric [DG13, vSW18]. unbalanced [Hay21]. Unbounded [DeB09, Men07, Vil11, XZ10b, Ada08, BPP22, CD15, CGL08, GP12, Gao17, JS22, KZ20, Lou22, RM13, Tyk07]. uncertainty [All21, CT22b, DG18, NN13]. unconstrained [BO19]. uncoupled [Bus16]. underlying [LZ12]. underprojective [DMR09a]. understanding [BHLN15]. Unequal [CP99, BvdHK15]. unfolded [VS15]. unfolding [Der16]. unified [EP23, HP21, HP17]. Uniform [BBM10, BR18b, BF17, GP12, KZ20, LMS08, AZ15a, BC04, BG13a, BV16, BJM14, DDS18, DE06, GBM22, GM17b, Jor13, LP19b, MN14, Per20a, PST17, Pit99, Ric18, SZ19, Tan21]. Uniformly [KN07, Win10]. Unimodular [AL07, AL19, BHMK21, AL17a, BT21, BPT17]. Unique [HM11, HC20, Yoo98]. Uniqueness [Bas96, Bäu19, BR12b, BHM13, GGY18, Hau05, JS17, KW10, MdBL09, Son12, BP12, BFGM19, BDP22, CPZ22, CHHM23, CK15, CK18b, GLM^{+}17a, KK20a, PPS18, Pow18, RS13b]. unitary [BW21, Dub21, Lam18, Web15]. Universal [ABBHK22, BC10, Ric15, Shk13]. Universality [BDvdHS20, Che17a, CL19, CL21a, CL21b, DTV18, DCLM18, ERSY10, FO18, TV11, AC19, EY12, EX22, IKM16, SS19b, SS15, Yan19, Yan23]. up-down [RLR22]. updating [EL16]. upon [Kwa12, Muc21]. Upper [BB11a, Bob22, Dal08, DZ22, Fol11, vdHS04, vdHS10, Car20, Gra14, Mas22, MM15]. Upper-tail [DZ22]. ups [LS21a]. Urn [DM11, MV08, Pen09, DW16b, LP19a, MM17, SP19]. Urn-related [MV08]. urns [MN18b, Zha14]. use [Hol12]. used [FJ12]. Using [BCKH22, Kry99, BX19, BRW20, BM06, BD14b, Jui18]. UST [Kar20]. Utility [LQ11].

Vacant [Win08, AS19]. valleys [AD15]. Value [Ber97, CH05, Jon01, Kry99, LW01a, BDF18, CL19, CL21a, GT23, Ngu12, Pin09]. Valued [BC20, Ose11, Sha11, AS18, BD19, CLSF19, DGP13, DPP15, DKLT22, Guf18a, Guf18b, JPP19, Lar15, LSF20, MM17, Mic20, MW19]. Values [BC08b, Hu00, SSW04, CT07, HKLR19, Kul08, Naj20, Osȩ13, SCZ09, Zhu22]. Vanishes [Zha09]. vanishing [GN21]. variability [HTV22]. Variable [ABPR11, MPX22, PM08, CGL08]. Variables [Bry96, CF11, EL10, HKL^{+}98, HRW07, Naj01, Pel99, Ada20, BCP18, DZ21, Del09, GSS19, GSS21, IPR09, LT15b, Pau14, Pin09, Pin16a, Pin18, PS22b, Roe08, Wan14, Zem08]. Variance [BST22, BGT11, FM03, Gau14, HJ09, HPZ11, LP22, OW04, vdHHZ07, BvdHK15, CN19, CHNR18, CK08, DPU15, DR14, DWW19, ET15, GLT21, HJ07, Hus22, Kul08, Lin14, Liu20, Tan19, Uch20, Völ14]. Variance-Gamma [Gau14, ET15]. Variation [FV11, BR16, BCP20, BX19, BCG^{+}13, BHP21, BHS22, BPP22, GI12, HLP16, LB18, PW23, Son12]. Variational [AID07, DG09, DZ15]. Variations [AP21, Beg05, BBL11, BOPP21, NP08, TTV14]. various [TY23]. varying [BS15, LL22, Lou22, LP22, RS21a]. Vector [BG03, Dom22]. vectors [Dol16, FS20b, NNP13, RM09, Yan22]. veering [NV13]. Velocity [Pai16]. Version [Maj07, OP13, Zha19]. versions [BPR15]. versus [Ber21a, DHS14b, GHMM22]. Vertex [CZ21, LR18, Col09, Jan09, Lin20, Ger20]. vertex-reinforced [Col09]. Vertices [BL12b, Gro11, Gro13, Thé20, Lod22]. Vervaat [FH10, LPT15]. Very [AC96]. Via [KN07, AHS23, BMZ13, BG21a, BJKM10, BC21, BdSGP09, Ber97, BJM14, BdT14, ČR22, CGR20, CKNP21, DP13a, Dir15, EL12, FL18, GHLT03, Gau14, GMT06, Ham19, HT21, KZ20, KSX04, KSX17, Kra20, Lab14, Nil20, Ram06, RC09, SHW20, Yan23]. view [Ben18, FKHL19]. viewpoint [LP19a]. violates [CHHM23]. Viot [BBP12, BB09, BEK21, CDE18, CP20, DG99, DGP13, EVY20, FM07, FH13, GK06, Guf18a, Guf18b, Gup12, Han14, KL22, Lab14, LZ12, TN22]. Viot-type [BBP12, KL22]. Virgin [Hut09]. Viscosity [CK15]. visited [AD14d, AD15, Bas23, Gib08, HS15a]. Vlasov [BSY22, GBM22, HW21, KN21, LS21a, TN22, Tug12]. Volatility [Fuk11, NN13, RR13]. Volterra [AM19, AF03, BH21, BK23, CMP10, HP21, Kal21, RA08]. Volterra-type [BK23]. Volume [AFS20, DH15, Gal22, GW09, Max06, Pag17, Spi20]. Volumetric [Eld14]. Voronoi [AB18a, AdlRG21, RP12, Van19]. Vortex [FG05]. Voter [BMSV06, ASD21, NRS17]. voting [DLZ14]. Vries [dBD09]. VRJP [KP21]. vs [BHTW22, RASY17]. Vulnerability [EM14].

Waiting [Sch08, CGR06]. Walk [BIK07, BG11, BT10, Ber06, BP07, Čer11, CO08, CCFR09, Far11, GKW99, HH05, KS03, KL05, Law96a, LMM04, MR08, Sel06, Su01, Tak97, Vol03, Win08, Win10, AP16, AKMV20, AS19, AH08, AD14d, AS20, ACdCdH19, BN15, BS08a, BS09, Bat18, BARS19, BM14, BD08, BS20c, Ber18, BV16, BBDS23, BT19, BCDG13, BHdS^{+}19, Bou13, BC22, Bud19, BD18, Bus16, CCK22, CC13, CP16, CHK18, CHT19, CS98, DGK21, DS12b, DW19, DF20, DP13b, EVW20, EV22, Eto06, FIO^{+}21, GPV09, GJ17, GKKR18, Gib08, GN21, GPP12, HPX22, Hei13, HdHS^{+}15, HM22, HS17a, HS15b, Jeg20, Jos09, JS18, KP13, Law20, LS19, LS22b, LP15, LM19b, Mal15a, Mas09, MS15, MW18, MPX22, MV08, Mou12, Mül08, OR17, PRR20, PRT22, Per21]. walk [PST17, PR12, PZ20, RS21a, RS09, RASY17, SP14, Szn09, Tag16, Tra18, Uch20, VT08, dS14]. walker [BP13b]. walkers [DGP18, DGP19]. Walking [DHS14b]. Walks [Ale11, AI09, AdHR11, BR07, BGdH11, BGT11, CFG09, CLR05, CC11, DW10, DH02, EK08, Fol11, Gil11, HMN11, Jag16, Jea05, KOR01, KPV04, LL04, NRS05, Pak98, PR04, Pet10, Ros97, SE06, Uch10, Uch11, Zäh01, Abe18, ABOPS16, ADS16b, And21, AA23, AFP15, BS22a, BCR06, BLV19, Ber19, BČD16, Bis19, BHdS^{+}19, BHP22, BM19b, BDI^{+}19, CMT15, CW21a, Che14, CHSW22, Cor21, CK08, CHK12, CSŠ22, DW12, DVW14, DFS21, DGP09, DKLT22, Fag08, Fag12, FZ12, GP12, GP20, GMW19, Gil16, GR17, HH22, Ham19, HdHS^{+}15, Hou17, HMMW12, HS15a, IKN20, KGS13, KZ08, KZ14, KP16, LM07a, LR22, LR18, MRS16, MV20, MM15, MV15, MSC15, NN21, Neu07, NV13, OS20, OPR22, Pak99, Par22, Pèn21, Pet12, Pha15, PT18]. walks [PRU20, RR22, Rap23, Rou15, SCZ13, SCW22, Sam09, Shi12, Su14, Tak21, Tou09, Ver19, ST17b]. wall [Fit21]. Walls [Sak09, BS20c]. Walsh [Haj11]. Wasserstein [BP13a, Boi11, Fan19, LSF20, Mar18, OC19]. Wasserstein-2 [Fan19]. Watanabe [Kur07, CR08]. Watson [AD14b, AD14a, AD19, AA23, BS15, BGH^{+}14, BDR21, CK08, DS12b, Dev11, DHS19, Far11, FJ22, FJ23, GS20, GGS21, Gei00, GN21, Joh19, Lin14, MPR19, MW20b, PR11, Rou18, Su14, Thé20]. Wave [CD08b, Ond10, BQSS19, GNZ21, LM19a, OO21, SSS13]. Wavelet [AF20, DL22, KOP13]. Waves [Lyn00, TW11, DNPR19, EP22]. Weak [ADMU20, Arc99, BJQS10, Bes99, EX22, KM17, LW01b, Win15, AP17, BMZ13, Bel20a, BV16, BJN13, BCPW21, BL19, CX13, DJQS13, Hab19, Kif17, NRS10, ORS20, PO09]. Weakly [Pel99, BDFR15a, CT20, Del09, DLR20, EHMT17, FJS15, GSS19, Tre16, Xu20]. wealth [EL12]. web [SS13, VFV15, NRS05]. webs [CMT19]. Wedge [CC11, PRZ20]. Wedges [Ata01, BBAP06]. Weibull [FM14]. weight [DÜ22, GHMM22, HS20, Kha14]. weight-dependent [GHMM22]. Weighted [AR06, DE06, Fér18, MP05, Naj01, NP08, Pan05, BCG18, BCG20, FS20a, For20, KM12, MT12a, OP13, RS21a, Sén21, Tei09, Wan20]. Weights [AR06, Tan21, ADS16b, KP21, LP22, OO15, SS20, SZ15b, SS21]. Weiss [BBI09, BMP21, CK18a, FKK22]. welding [AHS23]. Well [MR10, PYZ22, Tre16, Ber21a, FMR22, GdHO22, HP21, Gau07]. well-behaved [Ber21a]. well-mixed [FMR22]. Well-posedness [PYZ22, Tre16, GdHO22, HP21, Gau07]. wells [BS20c]. wetting [CGZ06, LT15a]. Where [PPPY01]. which [CO08, GM11]. White [BJQS10, XZ10b, AF20, DH19, DW18a, HIN17, Yam23]. Whittaker [Che19]. whole [PPS18]. Wick [LPSŽ18]. Wick-polynomial [LPSŽ18]. Widder [CD12a]. wide [EGL22]. Widom [AE19, BPZ12, FO18, OO15, Vet15]. Width [MT97, ABBHK22]. Wiener [CS98, ET15, GW09, MT02a, Mun11, NNP13, Pic06, Spi09]. Wigner [AEK^{+}14, Aug16, BWZ09, BM19a, CDMFF11, CES22, ERSY10, EY12, McK21, TV11]. Williams [DH13, RSZ18]. winding [Yao13]. windings [BW21, BHP22]. window [AO13, DvdHvLS17]. winning [AS13]. Wise [Kry99]. Wishart [CHZ22, HHN16]. Wishart-type [CHZ22]. within [BDLSJ18, DHS14b]. Without [Vih11, Aug16, CE13, FMNS15]. Wong [CS17, Gan13, Slo14]. Wong-Zakai [CS17]. word [DS12a, GS17b]. Words [KSZ98, DH21, HI23, KSZ01]. world [BR06]. Wreath [Eva02, FS01, KP13, SCZ13]. Wright [Dal21, DDSJ10, KZ20, SF06, Tay07]. Wrighter [GLW05].

XOR [BdT14].

Yaglom [BPW18, Vil11]. Yamada [Kur07]. yards [PPS18]. Young [FSX20, Hor20]. Yule [Ber15].

Z [Val10]. Zakai [CS17, Gan13, Slo14]. Zeitouni [But16]. Zelditch [But16]. Zero [AAJ21, DP05, Fag10, Sav08, BMH22, GMPW19, KK20b, KKPW14, LM19a, Mal15b, MM18, MV20, PL16, Pin09, PT15, RM09]. zero-mean [Pin09]. Zero-range [AAJ21]. zeroes [KS19]. Zeros [AL21, ABPR11, LP22, MS13, Red17]. zeroth [Ber21a]. zerothorder [Ber21a]. Zhang [MdBL09]. zones [EFP17, EGL22]. Zooming [BI20, Nas22]. Zooming-in [BI20]. Zygmund [AP21, AP23].

References

Andriopoulos:2023:SLL

[AA23] George Andriopoulos and Eleanor Archer. Scaling limit of linearly edge-reinforced random walks on critical Galton–Watson trees. Electronic Journal of Probability, 28(??):1–64, 2023. CO-

            DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/
            journals/electronic-journal-of-probability/volume-28/
            issue-none/Scaling-limit-of-linearly-edge-reinforced-
            random-walks-on-critical/10.1214/23-EJP901.full.

Atar:2009:ETG

[AAC09] Rami Atar, Siva Athreya, and Zhen-Qing Chen. Exit time, green function and semilinear elliptic equations. Electronic Journal of Probability, 14:3:50–3:71, 2009. CODEN ???? ISSN 1083-6489.

            URL http://ejp.ejpecp.org/article/view/597.

Andjel:2021:ZRP

[AAJ21] Enrique Andjel, Inés Armendáriz, and Milton Jara. Zerorange processes with rapidly growing rates. Electronic Journal of Probability, 26(18):1–29, February 2021. CODEN

???? ISSN 1083-6489. URL https://projecteuclid.org/
journals/electronic-journal-of-probability/volume-26/
issue-none/Zero-range-processes-with-rapidly-growing-
rates/10.1214/21-EJP670.full.

Atar:2002:SPC

[AB02] Rami Atar and Amarjit Budhiraja. Stability properties of constrained jump-diffusion processes. Electronic Journal of Probability, 7:22:1–22:31, 2002. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/121.

Assing:2013:SDS

[AB13] Sigurd Assing and James Bichard. On the spatial dynamics of the solution to the stochastic heat equation. Electronic Journal of Probability, 18:70:1–70:32, 2013. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/2797.

Alexander:2016:LAF

[AB16a] Kenneth S. Alexander and Quentin Berger. Local asymptotics for the first intersection of two independent renewals. Electronic Journal of Probability, 21(??):68:1–68:20, ???? 2016. CODEN

???? ISSN 1083-6489. URL https://projecteuclid.org/
euclid.ejp/1480561217.

Alexander:2016:LLT

[AB16b] Kenneth S. Alexander and Quentin Berger. Local limit theorems and renewal theory with no moments. Electronic Journal of Probability, 21(??):66:1–66:18, ???? 2016. CODEN ????

ISSN 1083-6489. URL https://projecteuclid.org/euclid.
ejp/1480129233.

Ahlberg:2018:NSV

[AB18a] Daniel Ahlberg and Rangel Baldasso. Noise sensitivity and Voronoi percolation. Electronic Journal of Probability, 23(??): 108:1–108:21, ???? 2018. CODEN ???? ISSN 1083-6489. URL

https://projecteuclid.org/euclid.ejp/1540865371.

Alexander:2018:PRQ

[AB18b] Kenneth S. Alexander and Quentin Berger. Pinning of a renewal on a quenched renewal. Electronic Journal of Probability, 23(??): 6:1–6:48, ???? 2018. CODEN ???? ISSN 1083-6489. URL https:

//projecteuclid.org/euclid.ejp/1518426053.

Amir:2020:PMD

[AB20] Gideon Amir and Rangel Baldasso. Percolation in majority dynamics. Electronic Journal of Probability, 25(??):13:1–13:18, ???? 2020. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1580374825.

Addario-Berry:2010:CRG

[ABBG10] Louigi Addario-Berry, Nicolas Broutin, and Christina Goldschmidt. Critical random graphs: Limiting constructions and distributional properties. Electronic Journal of Probability, 15: 25:741–25:775, 2010. CODEN ???? ISSN 1083-6489. URL http:

//ejp.ejpecp.org/article/view/772.

Addario-Berry:2022:UHW

[ABBHK22] Louigi Addario-Berry, Anna Brandenberger, Jad Hamdan, and Céline Kerriou. Universal height and width bounds for random trees. Electronic Journal of Probability, 27(??):1–24, 2022. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/
journals/electronic-journal-of-probability/volume-27/
issue-none/Universal-height-and-width-bounds-for-random-
trees/10.1214/22-EJP842.full.

Ayyer:2015:MJP

[ABCN15] Arvind Ayyer, Jérémie Bouttier, Sylvie Corteel, and François Nunzi. Multivariate juggling probabilities. Electronic Journal of Probability, 20(??):5:1–5:29, ???? 2015. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/3495.

Abe:2018:ELT

[Abe18] Yoshihiro Abe. Extremes of local times for simple random walks on symmetric trees. Electronic Journal of Probability, 23(??): 40:1–40:41, ???? 2018. CODEN ???? ISSN 1083-6489. URL

https://projecteuclid.org/euclid.ejp/1525852817.

Adams:2006:LSP

[ABK06] Stefan Adams, Jean-Bernard Bru, and Wolfgang Koenig. Large systems of path-repellent Brownian motions in a trap at positive temperature. Electronic Journal of Probability, 11:19:460–19:485, 2006. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.

org/article/view/330.

Arguin:2013:ETF

[ABK13] Louis-Pierre Arguin, Anton Bovier, and Nicola Kistler. An ergodic theorem for the frontier of branching Brownian motion. Electronic Journal of Probability, 18:53:1–53:25, 2013. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/
view/2082.

Addario-Berry:2016:RWC

[ABOPS16] Louigi Addario-Berry, Roberto I. Oliveira, Yuval Peres, and Perla Sousi. Random walks colliding before getting trapped. Electronic Journal of Probability, 21(??):42:1–42:19, ???? 2016. CO-DEN ???? ISSN 1083-6489. URL https://projecteuclid.

org/euclid.ejp/1469199632.

Antunovic:2011:IZB

[ABPR11] Tonci Antunovic, Krzysztof Burdzy, Yuval Peres, and Julia Ruscher. Isolated zeros for Brownian motion with variable drift. Electronic Journal of Probability, 16:65:1793–65:1814, 2011. CO-DEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/927.

Angst:2015:KBM

[ABT15] Jürgen Angst, Ismaël Bailleul, and Camille Tardif. Kinetic Brownian motion on Riemannian manifolds. Electronic Journal of Probability, 20(??):110:1–110:40, ???? 2015. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/4054.

Arous:1996:MTD

[AC96] Gérard Ben Arous and Raphaël Cerf. Metastability of the three dimensional Ising model on a torus at very low temperatures. Electronic Journal of Probability, 1:10:1–10:55, 1996. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/
view/10;   http://www.math.washington.edu/~ejpecp/EjpVol1/
paper10.abs.html.

Arias-Castro:2007:IRH

[AC07] Ery Arias-Castro. Interpolation of random hyperplanes. Electronic Journal of Probability, 12:38:1052–38:1071, 2007. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/
view/435.

Auffinger:2017:DPS

[AC17] Antonio Auffinger and Wei-Kuo Chen. A duality principle in spin glasses. Electronic Journal of Probability, 22(??):61:1–61:17, ???? 2017. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1500689052.

Adhikari:2019:EUC

[AC19] Arka Adhikari and Ziliang Che. Edge universality of correlated Gaussians. Electronic Journal of Probability, 24(??):44:1–44:25, ???? 2019. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1556179228.

Abacherli:2020:LSPa

[AČ20a] Angelo Abächerli and Jiří Černý. Level-set percolation of the Gaussian free field on regular graphs I: regular trees. Electronic Journal of Probability, 25(??):65:1–65:24, ???? 2020. CODEN

???? ISSN 1083-6489. URL https://projecteuclid.org/
euclid.ejp/1592445679.

Abacherli:2020:LSPb

[AČ20b] Angelo Abächerli and Jiří Černý. Level-set percolation of the Gaussian free field on regular graphs II: finite expanders. Electronic Journal of Probability, 25(??):1–39, 2020. CODEN

???? ISSN 1083-6489. URL https://projecteuclid.org/
journals/electronic-journal-of-probability/volume-25/
issue-none/Level-set-percolation-of-the-Gaussian-free-
field-on-regular/10.1214/20-EJP532.full.

Avena:2019:RWC

[ACdCdH19] Luca Avena, Yuki Chino, Conrado da Costa, and Frank den Hollander. Random walk in cooling random environment: ergodic limits and concentration inequalities. Electronic Journal of Probability, 24(??):38:1–38:35, ???? 2019. CODEN ????

ISSN 1083-6489. URL https://projecteuclid.org/euclid.
ejp/1554775418.

Aksamit:2021:TTR

[ACJ21] Anna Aksamit, Tahir Choulli, and Monique Jeanblanc. Thin times and random times’ decomposition. Electronic Journal of Probability, 26(??):1–22, 2021. CODEN ???? ISSN

1083-6489.  URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-26/issue-none/

Thin-times-and-random-times-decomposition/10.1214/20-
EJP569.full.

Acosta:2014:TRM

[Aco14] Javier Acosta. Tightness of the recentered maximum of logcorrelated Gaussian fields. Electronic Journal of Probability, 19 (??):90:1–90:25, ???? 2014. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/3170.

Ayala:2021:HOF

[ACR21] Mario Ayala, Gioia Carinci, and Frank Redig. Higher order fluctuation fields and orthogonal duality polynomials. Electronic Journal of Probability, 26(??):1–35, 2021. CODEN

???? ISSN 1083-6489. URL https://projecteuclid.org/
journals/electronic-journal-of-probability/volume-26/
issue-none/Higher-order-fluctuation-fields-and-orthogonal-
duality-polynomials/10.1214/21-EJP586.full.

Asselah:2016:DLD

[ACSS16] Amine Asselah, Emilio N. M. Cirillo, Benedetto Scoppola, and Elisabetta Scoppola. On diffusion limited deposition. Electronic Journal of Probability, 21(??):19:1–19:29, ???? 2016. CODEN

???? ISSN 1083-6489. URL https://projecteuclid.org/
euclid.ejp/1456499641.

Adler:2011:PAP

[ACvM11] Mark Adler, Mattia Cafasso, and Pierre van Moerbeke. From the Pearcey to the Airy process. Electronic Journal of Probability, 16:36:1048–36:1064, 2011. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/898.

Abraham:2014:LLCb

[AD14a] Romain Abraham and Jean-François Delmas. Local limits of conditioned Galton–Watson trees: the condensation case. Electronic Journal of Probability, 19:56:1–56:29, 2014. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/3164.

Abraham:2014:LLCa

[AD14b] Romain Abraham and Jean-François Delmas. Local limits of conditioned Galton–Watson trees: the infinite spine case. Electronic Journal of Probability, 19:2:1–2:19, 2014. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/2747.

Allez:2014:SKP

[AD14c] Romain Allez and Laure Dumaz. From sine kernel to Poisson statistics. Electronic Journal of Probability, 19(??):114:1–114:25, ???? 2014. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/3742.

Andreoletti:2014:SVS

[AD14d] Pierre Andreoletti and Pierre Debs. Spread of visited sites of a random walk along the generations of a branching process. Electronic Journal of Probability, 19:42:1–42:22, 2014. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
2790.

Andreoletti:2015:LNV

[AD15] Pierre Andreoletti and Alexis Devulder. Localization and number of visited valleys for a transient diffusion in random environment. Electronic Journal of Probability, 20(??):56:1–56:58, ???? 2015. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/3173.

Abraham:2019:APE

[AD19] Romain Abraham and Jean-François Delmas. Asymptotic properties of expansive Galton–Watson trees. Electronic Journal of Probability, 24(??):15:1–15:51, ???? 2019. CODEN ????

ISSN 1083-6489. URL https://projecteuclid.org/euclid.
ejp/1550826098.

Arnaudon:2020:DFP

[AD20] Marc Arnaudon and Pierre Del Moral. A duality formula and a particle Gibbs sampler for continuous time Feynman–Kac measures on path spaces. Electronic Journal of Probability, 25(??):1–54, 2020. CODEN ???? ISSN

1083-6489.  URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-25/issue-none/
A-duality-formula-and-a-particle-Gibbs-sampler-for-continuous/
10.1214/20-EJP546.full.

Adamczak:2008:TIS

[Ada08] Radoslaw Adamczak. A tail inequality for suprema of unbounded empirical processes with applications to Markov chains. Electronic Journal of Probability, 13:34:1000–34:1034, 2008. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/
view/521.

Adamczak:2011:MPC

[Ada11] Radoslaw Adamczak. On the Marchenko–Pastur and circular laws for some classes of random matrices with dependent entries. Electronic Journal of Probability, 16:37:1065–37:1095, 2011. CO-DEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/899.

Adamczak:2020:ASC

[Ada20] Radosław Adamczak. On almost sure convergence of random variables with finite chaos decomposition. Electronic Journal of Probability, 25(??):1–28, 2020. CODEN ????

ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-25/issue-none/
On-almost-sure-convergence-of-random-variables-with-finite-
chaos/10.1214/20-EJP538.full.

Abraham:2013:NGH

[ADH13] Romain Abraham, Jean-François Delmas, and Patrick Hoscheit. A note on the Gromov–Hausdorff–Prokhorov distance between (locally) compact metric measure spaces. Electronic Journal of Probability, 18:14:1–14:21, 2013. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/2116.

Adhikari:2022:SDG

[Adh22] Arka Adhikari. Spin distributions for generic spherical spin glasses. Electronic Journal of Probability, 27(??):1–43, 2022. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/
journals/electronic-journal-of-probability/volume-27/
issue-none/Spin-distributions-for-generic-spherical-spin-
glasses/10.1214/22-EJP755.full.

Avena:2011:LLN

[AdHR11] Luca Avena, Frank den Hollander, and Frank Redig. Law of large numbers for a class of random walks in dynamic random environments. Electronic Journal of Probability, 16:21:587–21:617, 2011. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/866.

Ahlberg:2021:RCQ

[AdlRG21] Daniel Ahlberg, Daniel de la Riva, and Simon Griffiths. On the rate of convergence in quenched Voronoi percolation. Electronic Journal of Probability, 26(18):1–26, February 2021. CODEN

???? ISSN 1083-6489. URL https://projecteuclid.org/
journals/electronic-journal-of-probability/volume-26/
issue-none/On-the-rate-of-convergence-in-quenched-Voronoi-
percolation/10.1214/21-EJP712.full.

Albeverio:2020:WSS

[ADMU20] Sergio Albeverio, Francesco C. De Vecchi, Paola Morando, and Stefania Ugolini. Weak symmetries of stochastic differential equations driven by semimartingales with jumps. Electronic Journal of Probability, 25(??):44:1–44:34, ???? 2020. CODEN ????

ISSN 1083-6489. URL https://projecteuclid.org/euclid.
ejp/1585965704.

Ahlberg:2016:IFP

[ADS16a] Daniel Ahlberg, Michael Damron, and Vladas Sidoravicius. Inhomogeneous first-passage percolation. Electronic Journal of Probability, 21(??):4:1–4:19, ???? 2016. CODEN ???? ISSN 1083-6489.

URL https://projecteuclid.org/euclid.ejp/1454514664.

Andres:2016:HKE

[ADS16b] Sebastian Andres, Jean-Dominique Deuschel, and Martin Slowik. Heat kernel estimates for random walks with degenerate weights. Electronic Journal of Probability, 21(??):33:1–33:21, ???? 2016. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.

org/euclid.ejp/1461007174.

Abraham:2010:PLC

[ADV10] Romain Abraham, Jean-François Delmas, and Guillaume Voisin. Pruning a Lévy continuum random tree. Electronic Journal of Probability, 15:46:1429–46:1473, 2010. CODEN ???? ISSN 1083-

6489. URL http://ejp.ejpecp.org/article/view/802.

Andrieu:2021:SHP

[ADW21] Christophe Andrieu, Paul Dobson, and Andi Q. Wang. Subgeometric hypocoercivity for piecewise-deterministic Markov process Monte Carlo methods. Electronic Journal of Probability, 26(18):1–26, February 2021. CODEN ???? ISSN

1083-6489.  URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-26/issue-none/
Subgeometric-hypocoercivity-for-piecewise-deterministic-
Markov-process-Monte-Carlo-methods/10.1214/21-EJP643.
full.

Adams:2019:WRM

[AE19] Stefan Adams and Michael Eyers. The Widom–Rowlinson model on the Delaunay graph. Electronic Journal of Probability, 24(??): 114:1–114:41, ???? 2019. CODEN ???? ISSN 1083-6489. URL

https://projecteuclid.org/euclid.ejp/1570759239.

Andjel:2015:SCP

[AEGR15] Enrique Andjel, François Ezanno, Pablo Groisman, and Leonardo T. Rolla. Subcritical contact process seen from the edge: Convergence to quasi-equilibrium. Electronic Journal of Probability, 20(??):32:1–32:16, ???? 2015. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/3881.

Alex:2014:ILL

[AEK^{+}14] Bloemendal Alex, László Erdős, Antti Knowles, Horng-Tzer Yau, and Jun Yin. Isotropic local laws for sample covariance and generalized Wigner matrices. Electronic Journal of Probability, 19: 33:1–33:53, 2014. CODEN ???? ISSN 1083-6489. URL http://

ejp.ejpecp.org/article/view/3054.

Alt:2017:LLR

[AEK17] Johannes Alt, László Erdős, and Torben Krüger. Local law for random Gram matrices. Electronic Journal of Probability, 22 (??):25:1–25:41, ???? 2017. CODEN ???? ISSN 1083-6489. URL

https://projecteuclid.org/euclid.ejp/1488942016.

Appleby:2003:EAS

[AF03] John Appleby and Alan Freeman. Exponential asymptotic stability of linear Itô–Volterra equation with damped stochastic perturbations. Electronic Journal of Probability, 8:22:1–22:22, 2003. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/179.

Aziznejad:2020:WAB

[AF20] Shayan Aziznejad and Julien Fageot. Wavelet analysis of the Besov regularity of Lévy white noise. Electronic Journal of Probability, 25(??):1–38, 2020. CODEN ???? ISSN

1083-6489.  URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-25/issue-none/
Wavelet-analysis-of-the-Besov-regularity-of-L%c3%a9vy-
white-noise/10.1214/20-EJP554.full.

Albenque:2023:RCP

[AFL23] Marie Albenque, Éric Fusy, and Thomas Lehéricy. Random cubic planar graphs converge to the Brownian sphere. Electronic Journal of Probability, 28(??):1–54, 2023. CODEN

???? ISSN 1083-6489. URL https://projecteuclid.org/
journals/electronic-journal-of-probability/volume-28/
issue-none/Random-cubic-planar-graphs-converge-to-the-
Brownian-sphere/10.1214/23-EJP912.full.

Assaf:2015:QTS

[AFP15] Sami Assaf, Noah Mills Forman, and Jim Pitman. The quantile transform of simple walks and Brownian motion. Electronic Journal of Probability, 20(??):90:1–90:39, ???? 2015. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/
view/3479.

Ang:2020:VMB

[AFS20] Morris Ang, Hugo Falconet, and Xin Sun. Volume of metric balls in Liouville quantum gravity. Electronic Journal of Probability, 25(??):1–50, 2020. CODEN ???? ISSN

1083-6489.  URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-25/issue-none/
Volume-of-metric-balls-in-Liouville-quantum-gravity/10.
1214/20-EJP564.full.

Allez:2013:DMM

[AG13] Romain Allez and Alice Guionnet. A diffusive matrix model for invariant β-ensembles. Electronic Journal of Probability, 18:62:1– 62:30, 2013. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/2073.

Angel:2012:PTR

[AGH12] Omer Angel, Vadim Gorin, and Alexander Holroyd. A pattern theorem for random sorting networks. Electronic Journal of Probability, 17:99:1–99:16, 2012. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/2448.

Adamczak:2012:MEC

[AGL^{+}12] Radosław Adamczak, Olivier Guédon, Rafał Latała, Alexander Litvak, Krzysztof Oleszkiewicz, Alain Pajor, and Nicole Tomczak-Jaegermann. Moment estimates for convex measures. Electronic Journal of Probability, 17:101:1–101:19, 2012. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/
view/2150.

Avena:2022:LEP

[AGMQ22] Luca Avena, Alexandre Gaudillière, Paolo Milanesi, and Matteo Quattropani. Loop-erased partitioning of a graph: mean-field analysis. Electronic Journal of Probability, 27(??):1–35, 2022. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.

org/journals/electronic-journal-of-probability/volume-
27/issue-none/Loop-erased-partitioning-of-a-graph-mean-
field-analysis/10.1214/22-EJP792.full.

Ambrosio:2022:QRM

[AGT22] Luigi Ambrosio, Michael Goldman, and Dario Trevisan. On the quadratic random matching problem in two-dimensional domains. Electronic Journal of Probability, 27(??):1–35, 2022. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/
journals/electronic-journal-of-probability/volume-27/
issue-none/On-the-quadratic-random-matching-problem-in-
two-dimensional-domains/10.1214/22-EJP784.full.

Ariyoshi:2005:STA

[AH05] Teppei Ariyoshi and Masanori Hino. Small-time asymptotic estimates in local Dirichlet spaces. Electronic Journal of Probability, 10:37:1236–37:1259, 2005. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/286.

Amir:2008:SSE

[AH08] Gideon Amir and Christopher Hoffman. A special set of exceptional times for dynamical random walk on Z^{2}. Electronic Journal of Probability, 13:63:1927–63:1951, 2008. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/571.

Arizmendi:2018:LTF

[AH18] Octavio Arizmendi and Takahiro Hasebe. Limit theorems for free Lévy processes. Electronic Journal of Probability, 23(??): 101:1–101:36, ???? 2018. CODEN ???? ISSN 1083-6489. URL

https://projecteuclid.org/euclid.ejp/1538618571.

Arras:2019:SMMb

[AH19a] Benjamin Arras and Christian Houdré. On Stein’s method for multivariate self-decomposable laws. Electronic Journal of Probability, 24(??):128:1–128:63, ???? 2019. CODEN ???? ISSN

1083-6489. URL https://projecteuclid.org/euclid.ejp/
1573268592.

Arras:2019:SMMa

[AH19b] Benjamin Arras and Christian Houdré. On Stein’s method for multivariate self-decomposable laws with finite first moment. Electronic Journal of Probability, 24(??):29:1–29:33, ???? 2019. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.

org/euclid.ejp/1553565780.

Assing:2021:ETF

[AH21] Sigurd Assing and John Herman. Extension technique for functions of diffusion operators: a stochastic approach. Electronic Journal of Probability, 26(18):1–32, February 2021. CODEN

???? ISSN 1083-6489. URL https://projecteuclid.org/
journals/electronic-journal-of-probability/volume-26/
issue-none/Extension-technique-for-functions-of-diffusion-
operators--a-stochastic/10.1214/21-EJP624.full.

Albenque:2020:SLT

[AHS20]  Marie Albenque, Nina Holden, and Xin Sun.  Scaling

limit of triangulations of polygons. Electronic Journal of Probability, 25(??):1–43, 2020. CODEN ???? ISSN

1083-6489.  URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-25/issue-none/
Scaling-limit-of-triangulations-of-polygons/10.1214/20-
EJP537.full.

Ang:2023:SLC

[AHS23] Morris Ang, Nina Holden, and Xin Sun. The SLE loop via conformal welding of quantum disks. Electronic Journal of Probability, 28(??):1–20, 2023. CODEN ???? ISSN

1083-6489.  URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-28/issue-none/
The-SLE-loop-via-conformal-welding-of-quantum-disks/10.
1214/23-EJP914.full.

Alsmeyer:2009:LTM

[AI09] Gerold Alsmeyer and Alex Iksanov. A log-type moment result for perpetuities and its application to martingales in supercritical branching random walks. Electronic Journal of Probability, 14: 10:289–10:313, 2009. CODEN ???? ISSN 1083-6489. URL http:

//ejp.ejpecp.org/article/view/596.

Ankirchner:2007:CVD

[AID07] Stefan Ankirchner, Peter Imkeller, and Goncalo Dos Reis. Classical and variational differentiability of BSDEs with quadratic growth. Electronic Journal of Probability, 12:53:1418–53:1453, 2007. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.

org/article/view/462.

Alfonsi:2015:OTB

[AJKH15] Aurélien Alfonsi, Benjamin Jourdain, and Arturo Kohatsu-Higa. Optimal transport bounds between the time-marginals of a multidimensional diffusion and its Euler scheme. Electronic Journal of Probability, 20(??):70:1–70:31, ???? 2015. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/4195.

Alm:2014:FCP

[AJL14] Sven Erick Alm, Svante Janson, and Svante Linusson. First critical probability for a problem on random orientations in G(n, p). Electronic Journal of Probability, 19:69:1–69:14, 2014. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/
view/2725.

Apollonio:2022:MIM

[AJNT22] Valentina Apollonio, Vanessa Jacquier, Francesca Romana Nardi, and Alessio Troiani. Metastability for the Ising model on the hexagonal lattice. Electronic Journal of Probability, 27(??):1–48, 2022. CODEN ???? ISSN

1083-6489.  URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-27/issue-none/
Metastability-for-the-Ising-model-on-the-hexagonal-lattice/
10.1214/22-EJP763.full.

Athreya:2012:PLF

[AKM12] Avanti Athreya, Tiffany Kolba, and Jonathan Mattingly. Propagating Lyapunov functions to prove noise-induced stabilization. Electronic Journal of Probability, 17:96:1–96:38, 2012. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/
view/2410.

Alsmeyer:2020:HLC

[AKMV20] Gerold Alsmeyer, Zakhar Kabluchko, Alexander Marynych, and Vladislav Vysotsky. How long is the convex minorant of a one-dimensional random walk? Electronic Journal of Probability, 25(??):1–22, 2020. CODEN ???? ISSN

1083-6489.  URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-25/issue-none/
How-long-is-the-convex-minorant-of-a-one-dimensional/
10.1214/20-EJP497.full.

Adamczak:2019:NCP

[AKPS19] Radosław Adamczak, Michał Kotowski, Bartłomiej Polaczyk, and Michał Strzelecki. A note on concentration for polynomials in the Ising model. Electronic Journal of Probability, 24(??):42:1–42:22, ???? 2019. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1555466612.

Ankirchner:2019:SEC

[AKTKK19] Stefan Ankirchner, Nabil Kazi-Tani, Maike Klein, and Thomas Kruse. Stopping with expectation constraints: 3 points suffice. Electronic Journal of Probability, 24(??):66:1–66:16, ???? 2019. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.

org/euclid.ejp/1561687599.

Aldous:1998:EBM

[AL98] David Aldous and Vlada Limic. The entrance boundary of the multiplicative coalescent. Electronic Journal of Probability, 3: 3:1–3:59, 1998. CODEN ???? ISSN 1083-6489. URL http://

ejp.ejpecp.org/article/view/25.

Aldous:2007:PUR

[AL07] David Aldous and Russell Lyons. Processes on unimodular random networks. Electronic Journal of Probability, 12:54:1454– 54:1508, 2007. CODEN ???? ISSN 1083-6489. URL http://

ejp.ejpecp.org/article/view/463. See errata [AL17a, AL19].

Aurzada:2009:SDP

[AL09] Frank Aurzada and Mikhail Lifshits. On the small deviation problem for some iterated processes. Electronic Journal of Probability, 14:68:1992–68:2010, 2009. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/689.

Azais:2013:CCR

[AL13] Jean-Marc Azaïs and José León. CLT for crossings of random trigonometric polynomials. Electronic Journal of Probability, 18: 68:1–68:17, 2013. CODEN ???? ISSN 1083-6489. URL http://

ejp.ejpecp.org/article/view/2403.

Aldous:2017:EPU

[AL17a] David Aldous and Russell Lyons. Errata to “Processes on unimodular random networks”. Electronic Journal of Probability, 22(??):51:1–51:4, ???? 2017. CODEN ???? ISSN

1083-6489. URL https://projecteuclid.org/euclid.ejp/

1498010464. See [AL07, AL19].

Arnaudon:2017:RBM

[AL17b] Marc Arnaudon and Xue-Mei Li. Reflected Brownian motion: selection, approximation and linearization. Electronic Journal of Probability, 22(??):31:1–31:55, ???? 2017. CODEN ????

ISSN 1083-6489. URL https://projecteuclid.org/euclid.
ejp/1490407496.

Aldous:2019:SEP

[AL19] David Aldous and Russell Lyons. Second errata to “Processes on Unimodular Random Networks”. Electronic Journal of Probability, 24(??):25:1–25:2, ???? 2019. CODEN ???? ISSN

1083-6489. URL https://projecteuclid.org/euclid.ejp/

1553565776. See [AL07, AL17a].

Azais:2020:NSC

[AL20] Jean-Marc Azaïs and José R. León. Necessary and sufficient conditions for the finiteness of the second moment of the measure of level sets. Electronic Journal of Probability, 25(??):1–15, 2020. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.

org/journals/electronic-journal-of-probability/volume-
25/issue-none/Necessary-and-sufficient-conditions-for-
the-finiteness-of-the-second/10.1214/20-EJP508.full.

Ancona:2021:ZSS

[AL21] Michele Ancona and Thomas Letendre. Zeros of smooth stationary Gaussian processes. Electronic Journal of Probability, 26(18):1–81, February 2021. CODEN ???? ISSN

1083-6489.  URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-26/issue-none/
Zeros-of-smooth-stationary-Gaussian-processes/10.1214/
21-EJP637.full.

Aldous:2014:SIR

[Ald14] David Aldous. Scale-invariant random spatial networks. Electronic Journal of Probability, 19:15:1–15:41, 2014. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
2920.

Alexander:2011:ELL

[Ale11] Kenneth Alexander. Excursions and local limit theorems for Bessel-like random walks. Electronic Journal of Probability, 16: 1:1–1:44, 2011. CODEN ???? ISSN 1083-6489. URL http://

ejp.ejpecp.org/article/view/848.

Allan:2021:RFP

[All21] Andrew L. Allan. Robust filtering and propagation of uncertainty in hidden Markov models. Electronic Journal of Probability, 26(18):1–37, February 2021. CODEN

???? ISSN 1083-6489. URL https://projecteuclid.org/
journals/electronic-journal-of-probability/volume-26/
issue-none/Robust-filtering-and-propagation-of-uncertainty-
in-hidden-Markov-models/10.1214/21-EJP633.full.

Aoyama:2010:NFM

[ALM10] Takahiro Aoyama, Alexander Lindner, and Makoto Maejima. A new family of mappings of infinitely divisible distributions related to the Goldie–Steutel–Bondesson class. Electronic Journal of Probability, 15:35:1119–35:1142, 2010. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/791.

Adamczak:2021:MGC

[ALM21] Radosław Adamczak, Rafał Latała, and Rafał Meller. Moments of Gaussian chaoses in Banach spaces. Electronic Journal of Probability, 26(??):1–36, 2021. CODEN ???? ISSN

1083-6489.  URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-26/issue-none/
Moments-of-Gaussian-chaoses-in-Banach-spaces/10.1214/
20-EJP567.full.

Aksamit:2023:GBR

[ALR23] Anna Aksamit, Libo Li, and Marek Rutkowski. Generalized BSDE and reflected BSDE with random time horizon. Electronic Journal of Probability, 28(??):1–41, 2023. CODEN

???? ISSN 1083-6489. URL https://projecteuclid.org/
journals/electronic-journal-of-probability/volume-28/
issue-none/Generalized-BSDE-and-reflected-BSDE-with-random-
time-horizon/10.1214/23-EJP927.full.

Aldous:2015:CGP

[ALS15] David Aldous, Daniel Lanoue, and Justin Salez. The compulsive gambler process. Electronic Journal of Probability, 20(??):35:1– 35:18, ???? 2015. CODEN ???? ISSN 1083-6489. URL http://

ejp.ejpecp.org/article/view/3582.

Altman:2021:BSG

[Alt21] Henri Elad Altman. Bessel SPDEs with general Dirichlet boundary conditions. Electronic Journal of Probability, 26(18):1–36, February 2021. CODEN ???? ISSN

1083-6489.  URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-26/issue-none/
Bessel-SPDEs-with-general-Dirichlet-boundary-conditions/
10.1214/21-EJP632.full.

Alabert:2004:SDE

[AM04] Aureli Alabert and Miguel Angel Marmolejo. Stochastic differential equations with boundary conditions driven by a Poisson noise. Electronic Journal of Probability, 9:9:230–254, 2004. CO-DEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/157.

Albenque:2008:SFI

[AM08] Marie Albenque and Jean-François Marckert. Some families of increasing planar maps. Electronic Journal of Probability, 13: 56:1624–56:1671, 2008. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/563.

Adamczak:2015:COU

[AM15] Radosław Adamczak and Piotr Miłoś. CLT for Ornstein– Uhlenbeck branching particle system. Electronic Journal of Probability, 20(??):42:1–42:35, ???? 2015. CODEN ???? ISSN 1083-

6489. URL http://ejp.ejpecp.org/article/view/4233.

Aleandri:2019:ODL

[AM19] Michele Aleandri and Ida G. Minelli. Opinion dynamics with Lotka–Volterra type interactions. Electronic Journal of Probability, 24(??):122:1–122:31, ???? 2019. CODEN ???? ISSN

1083-6489. URL https://projecteuclid.org/euclid.ejp/
1573030843.

Ameur:2021:LTP

[Ame21] Yacin Ameur. A localization theorem for the planar Coulomb gas in an external field. Electronic Journal of Probability, 26(18):1–21, February 2021. CODEN ???? ISSN

1083-6489.  URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-26/issue-none/
A-localization-theorem-for-the-planar-Coulomb-gas-in-
an/10.1214/21-EJP613.full.

Alves:2002:PTF

[AMP02] Oswaldo Alves, Fabio Machado, and Serguei Popov. Phase transition for the frog model. Electronic Journal of Probability, 7: 16:1–16:21, 2002. CODEN ???? ISSN 1083-6489. URL http://

ejp.ejpecp.org/article/view/115.

Aldous:2004:BBA

[AMP04] David Aldous, Gregory Miermont, and Jim Pitman. Brownian bridge asymptotics for random p-mappings. Electronic Journal of Probability, 9:3:37–3:56, 2004. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/186.

Andjel:2010:SSM

[AMP10] Enrique Andjel, Judith Miller, and Etienne Pardoux. Survival of a single mutant in one dimension. Electronic Journal of Probability, 15:14:386–14:408, 2010. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/769.

Applebaum:2014:SQS

[AN14] David Applebaum and Jan Neerven. Second quantisation for skew convolution products of measures in Banach spaces. Electronic Journal of Probability, 19:11:1–11:17, 2014. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
3031.

Andres:2011:PDS

[And11] Sebastian Andres. Pathwise differentiability for SDEs in a smooth domain with reflection. Electronic Journal of Probability, 16: 28:845–28:879, 2011. CODEN ???? ISSN 1083-6489. URL http:

//ejp.ejpecp.org/article/view/872.

Andriopoulos:2021:IPR

[And21] George Andriopoulos. Invariance principles for random walks in random environment on trees. Electronic Journal of Probability, 26(18):1–38, February 2021. CODEN ???? ISSN

1083-6489.  URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-26/issue-none/
Invariance-principles-for-random-walks-in-random-environment-
on-trees/10.1214/21-EJP687.full.

Angelis:2022:SSC

[Ang22] Tiziano De Angelis. Stopping spikes, continuation bays and other features of optimal stopping with finite-time horizon. Electronic Journal of Probability, 27(??):1–41, 2022. CODEN

???? ISSN 1083-6489. URL https://projecteuclid.org/
journals/electronic-journal-of-probability/volume-27/
issue-none/Stopping-spikes-continuation-bays-and-other-
features-of-optimal-stopping/10.1214/21-EJP733.full.

Alberts:2013:NCS

[AO13] Tom Alberts and Marcel Ortgiese. The near-critical scaling window for directed polymers on disordered trees. Electronic Journal of Probability, 18:19:1–19:24, 2013. CODEN ???? ISSN 1083-

6489. URL http://ejp.ejpecp.org/article/view/2036.

Aryasova:2012:PFG

[AP12] Olga Aryasova and Andrey Pilipenko. On properties of a flow generated by an SDE with discontinuous drift. Electronic Journal of Probability, 17:106:1–106:20, 2012. CODEN ???? ISSN 1083-

6489. URL http://ejp.ejpecp.org/article/view/2138.

Ahn:2016:OQS

[AP16] Sung Won Ahn and Jonathon Peterson. Oscillations of quenched slowdown asymptotics for ballistic one-dimensional random walk in a random environment. Electronic Journal of Probability, 21 (??):16:1–16:27, ???? 2016. CODEN ???? ISSN 1083-6489. URL

https://projecteuclid.org/euclid.ejp/1456412956.

Angst:2017:WCC

[AP17] Jürgen Angst and Guillaume Poly. A weak Cramér condition and application to Edgeworth expansions. Electronic Journal of Probability, 22(??):59:1–59:24, ???? 2017. CODEN ????

ISSN 1083-6489. URL https://projecteuclid.org/euclid.
ejp/1500516021.

Angst:2021:VSZ

[AP21] Jürgen Angst and Guillaume Poly. Variations on Salem– Zygmund results for random trigonometric polynomials: application to almost sure nodal asymptotics. Electronic Journal of Probability, 26(18):1–36, February 2021. CODEN

???? ISSN 1083-6489. URL https://projecteuclid.org/
journals/electronic-journal-of-probability/volume-26/
issue-none/Variations-on-SalemZygmund-results-for-random-
trigonometric-polynomials--application/10.1214/21-EJP716.
full.

Angst:2023:FSZ

[AP23] Jürgen Angst and Guillaume Poly. Fluctuations in Salem– Zygmund almost sure Central Limit Theorem. Electronic Journal of Probability, 28(??):1–40, 2023. CODEN ????

ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-28/issue-none/
Fluctuations-in-SalemZygmund-almost-sure-Central-Limit-
Theorem/10.1214/23-EJP931.full.

Ahmadi:2021:BSD

[APS21] Mahdi Ahmadi, Alexandre Popier, and Ali Devin Sezer. Backward stochastic differential equations with non-Markovian singular terminal conditions for general driver and filtration. Electronic Journal of Probability, 26(18):1–27, February 2021. CODEN

???? ISSN 1083-6489. URL https://projecteuclid.org/
journals/electronic-journal-of-probability/volume-26/
issue-none/Backward-stochastic-differential-equations-
with-non-Markovian-singular-terminal-conditions/10.1214/
21-EJP619.full.

Ang:2020:LDR

[APW20] Morris Ang, Minjae Park, and Yilin Wang. Large deviations of radial SLE. Electronic Journal of Probability, 25(??):1–13, 2020. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.

org/journals/electronic-journal-of-probability/volume-
25/issue-none/Large-deviations-of-radial-SLE_infty-/10.
1214/20-EJP502.full.

Alsmeyer:2006:SFP

[AR06] Gerold Alsmeyer and Uwe Rösler. A stochastic fixed point equation related to weighted branching with deterministic weights. Electronic Journal of Probability, 11:2:27–2:56, 2006. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/
view/296.

Arcones:1997:LIL

[Arc97] Miguel A. Arcones. The law of the iterated logarithm for a triangular array of empirical processes. Electronic Journal of Probability, 2:5:1–5:39, 1997. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/19.

Arcones:1999:WCR

[Arc99] Miguel A. Arcones. Weak convergence for the row sums of a triangular array of empirical processes indexed by a manageable triangular array of functions. Electronic Journal of Probability, 4:7:1–7:17, 1999. CODEN ???? ISSN 1083-6489. URL http://

ejp.ejpecp.org/article/view/44.

Archer:2020:ISL

[Arc20] Eleanor Archer. Infinite stable looptrees. Electronic Journal of Probability, 25(??):11:1–11:48, ???? 2020. CODEN ????

ISSN 1083-6489. URL https://projecteuclid.org/euclid.
ejp/1580267011.

Arguin:2008:CPS

[Arg08] Louis-Pierre Arguin. Competing particle systems and the Ghirlanda–Guerra identities. Electronic Journal of Probability, 13:69:2101–69:2117, 2008. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/579.

Athreya:2008:RDT

[ARS08] Siva Athreya, Rahul Roy, and Anish Sarkar. Random directed trees and forest — drainage networks with dependence. Electronic Journal of Probability, 13:71:2160–71:2189, 2008. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
580.

Abraham:2002:PSF

[AS02] Romain Abraham and Laurent Serlet. Poisson snake and fragmentation. Electronic Journal of Probability, 7:17:1–17:15, 2002. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/116.

Alberts:2008:HDS

[AS08] Tom Alberts and Scott Sheffield. Hausdorff dimension of the SLE curve intersected with the real line. Electronic Journal of Probability, 13:40:1166–40:1188, 2008. CODEN ???? ISSN 1083-

6489. URL http://ejp.ejpecp.org/article/view/515.

Athreya:2012:SBA

[AS12] Siva Athreya and Jan Swart. Systems of branching, annihilating, and coalescing particles. Electronic Journal of Probability, 17: 80:1–80:32, 2012. CODEN ???? ISSN 1083-6489. URL http://

ejp.ejpecp.org/article/view/2003.

Aldous:2013:FMW

[AS13] David Aldous and Mykhaylo Shkolnikov. Fluctuations of martingales and winning probabilities of game contestants. Electronic Journal of Probability, 18:47:1–47:17, 2013. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/2422.

Aru:2018:TVL

[AS18] Juhan Aru and Avelio Sepúlveda. Two-valued local sets of the 2D continuum Gaussian free field: connectivity, labels, and induced metrics. Electronic Journal of Probability, 23(??):61:1– 61:35, ???? 2018. CODEN ???? ISSN 1083-6489. URL https:/

/projecteuclid.org/euclid.ejp/1529460159.

Alves:2019:DIS

[AS19] Caio Alves and Artem Sapozhnikov. Decoupling inequalities and supercritical percolation for the vacant set of random walk loop soup. Electronic Journal of Probability, 24(??):110:1–110:34, ???? 2019. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1570068174.

Asselah:2020:DCR

[AS20] Amine Asselah and Bruno Schapira. Deviations for the capacity of the range of a random walk. Electronic Journal of Probability, 25(??):1–28, 2020. CODEN ???? ISSN

1083-6489.  URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-25/issue-none/
Deviations-for-the-capacity-of-the-range-of-a-random/
10.1214/20-EJP560.full.

Agarwal:2021:VMT

[ASD21] Pooja Agarwal, Mackenzie Simper, and Rick Durrett. The q-voter model on the torus. Electronic Journal of Probability, 26(18):1–33, February 2021. CODEN ???? ISSN

1083-6489.  URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-26/issue-none/
The-q-voter-model-on-the-torus/10.1214/21-EJP682.full.

Assiotis:2018:RSG

[Ass18] Theodoros Assiotis. Random surface growth and Karlin– McGregor polynomials. Electronic Journal of Probability, 23(??): 106:1–106:81, ???? 2018. CODEN ???? ISSN 1083-6489. URL

https://projecteuclid.org/euclid.ejp/1540260052.

Abreu:2009:TLP

[AT09] Victor Perez Abreu and Constantin Tudor. On the traces of Laguerre processes. Electronic Journal of Probability, 14:76:2241– 76:2263, 2009. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/702.

Auffinger:2016:TCH

[AT16] Antonio Auffinger and Si Tang. On the time constant of high dimensional first passage percolation. Electronic Journal of Probability, 21(??):24:1–24:23, ???? 2016. CODEN ???? ISSN

1083-6489. URL https://projecteuclid.org/euclid.ejp/
1459880112.

Atar:2001:IWT

[Ata01] Rami Atar. Invariant wedges for a two-point reflecting Brownian motion and the “hot spots” problem. Electronic Journal of Probability, 6:18:1–18:19, 2001. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/91.

Au:2018:TDR

[Au18] Benson Au. Traffic distributions of random band matrices. Electronic Journal of Probability, 23(??):77:1–77:48, ???? 2018. CO-DEN ???? ISSN 1083-6489. URL https://projecteuclid.

org/euclid.ejp/1536717736.

Augeri:2016:LDP

[Aug16] Fanny Augeri. Large deviations principle for the largest eigenvalue of Wigner matrices without Gaussian tails. Electronic Journal of Probability, 21(??):32:1–32:49, ???? 2016. CODEN ????

ISSN 1083-6489. URL https://projecteuclid.org/euclid.
ejp/1461007173.

Abraham:1997:APB

[AW97] Romain Abraham and Wendelin Werner. Avoiding-probabilities for Brownian snakes and super-Brownian motion. Electronic Journal of Probability, 2:3:1–3:27, 1997. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/17.

Appleby:2009:SSD

[AW09] John Appleby and Huizhong Wu. Solutions of stochastic differential equations obeying the law of the iterated logarithm, with applications to financial markets. Electronic Journal of Probability, 14:33:912–33:959, 2009. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/642.

Alili:2014:MLT

[AW14] Larbi Alili and Ching-Tang Wu. Müntz linear transforms of Brownian motion. Electronic Journal of Probability, 19:36:1– 36:15, 2014. CODEN ???? ISSN 1083-6489. URL http://

ejp.ejpecp.org/article/view/2424.

Alexander:2015:DPR

[AY15] Kenneth S. Alexander and Gökhan Yıldırım. Directed polymers in a random environment with a defect line. Electronic Journal of Probability, 20(??):6:1–6:20, ???? 2015. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/3379.

Alexander:2013:SCR

[AZ13] Kenneth Alexander and Nikolaos Zygouras. Subgaussian concentration and rates of convergence in directed polymers. Electronic Journal of Probability, 18:5:1–5:28, 2013. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/2005.

Alishahi:2015:SEU

[AZ15a] Kasra Alishahi and Mohammadsadegh Zamani. The spherical ensemble and uniform distribution of points on the sphere. Electronic Journal of Probability, 20(??):23:1–23:27, ???? 2015. CO-DEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/3733.

Arguin:2015:PDS

[AZ15b] Louis-Pierre Arguin and Olivier Zindy. Poisson–Dirichlet statistics for the extremes of the two-dimensional discrete Gaussian free field. Electronic Journal of Probability, 20(??):59:1–59:19, ???? 2015. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/3077.

Ben-Ari:2012:PEB

[BA12] Iddo Ben-Ari. Principal eigenvalue for Brownian motion on a bounded interval with degenerate instantaneous jumps. Electronic Journal of Probability, 17:87:1–87:13, 2012. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
1791.

Bakhtin:2016:IBE

[Bak16] Yuri Bakhtin. Inviscid Burgers equation with random kick forcing in noncompact setting. Electronic Journal of Probability, 21(??): 37:1–37:50, ???? 2016. CODEN ???? ISSN 1083-6489. URL

https://projecteuclid.org/euclid.ejp/1463683782.

Barhoumi-Andreani:2019:BFN

[BAKL19] Yacine Barhoumi-Andréani, Christoph Koch, and Hong Liu. Bivariate fluctuations for the number of arithmetic progressions in random sets. Electronic Journal of Probability, 24(??):145:1– 145:32, ???? 2019. CODEN ???? ISSN 1083-6489. URL https:

//projecteuclid.org/euclid.ejp/1577502322.

Balanca:2014:FRL

[Bal14] Paul Balança. Fine regularity of Lévy processes and linear (multi)fractional stable motion. Electronic Journal of Probability, 19(??):101:1–101:37, ???? 2014. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/3393.

Banerjee:2018:CNR

[Ban18] Debapratim Banerjee. Contiguity and non-reconstruction results for planted partition models: the dense case. Electronic Journal of Probability, 23(??):18:1–18:28, ???? 2018. CODEN ????

ISSN 1083-6489. URL https://projecteuclid.org/euclid.
ejp/1519354947.

Ben-Ari:2019:RWC

[BARS19] Iddo Ben-Ari, Alexander Roitershtein, and Rinaldo B. Schinazi. A random walk with catastrophes. Electronic Journal of Probability, 24(??):28:1–28:21, ???? 2019. CODEN ???? ISSN 1083-6489.

URL https://projecteuclid.org/euclid.ejp/1553565779.

Bass:1996:USE

[Bas96] Richard F. Bass. Uniqueness for the Skorokhod equation with normal reflection in Lipschitz domains. Electronic Journal of Probability, 1:11:1–11:29, 1996. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/11;   http:/
/www.math.washington.edu/~ejpecp/EjpVol1/paper11.abs.
html.

Basdevant:2006:FOP

[Bas06] Anne-Laure Basdevant. Fragmentation of ordered partitions and intervals. Electronic Journal of Probability, 11:16:394–16:417, 2006. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.

org/article/view/323.

Basse:2008:GMA

[Bas08] Andreas Basse. Gaussian moving averages and semimartingales. Electronic Journal of Probability, 13:39:1140–39:1165, 2008. CO-DEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/526.

Bass:2014:SDE

[Bas14] Richard Bass. A stochastic differential equation with a sticky point. Electronic Journal of Probability, 19:32:1–32:22, 2014. CO-DEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/2350.

Bass:2023:REM

[Bas23] Richard F. Bass. The rate of escape of the most visited site of Brownian motion. Electronic Journal of Probability, 28(??): 1–12, 2023. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/journals/electronic-journal-of-probability/
volume-28/issue-none/The-rate-of-escape-of-the-most-visited-
site-of/10.1214/23-EJP916.full.

Bates:2018:LDP

[Bat18] Erik Bates. Localization of directed polymers with general reference walk. Electronic Journal of Probability, 23(??):30:1–30:45, ???? 2018. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1522375270.

Bates:2021:FPL

[Bat21] Erik Bates. Full-path localization of directed polymers. Electronic Journal of Probability, 26(18):1–24, February 2021. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/
journals/electronic-journal-of-probability/volume-26/
issue-none/Full-path-localization-of-directed-polymers/
10.1214/21-EJP641.full.

Baumler:2019:UNU

[Bäu19] Johannes Bäumler. Uniqueness and non-uniqueness for spin-glass ground states on trees. Electronic Journal of Probability, 24(??): 77:1–77:17, ???? 2019. CODEN ???? ISSN 1083-6489. URL

https://projecteuclid.org/euclid.ejp/1563264040.

Baverez:2019:MBA

[Bav19] Guillaume Baverez. Modular bootstrap agrees with the path integral in the large moduli limit. Electronic Journal of Probability, 24(??):144:1–144:22, ???? 2019. CODEN ???? ISSN 1083-6489.

URL https://projecteuclid.org/euclid.ejp/1576810976.

Bass:1996:EEB

[BB96] Richard F. Bass and Krzysztof Burdzy. Eigenvalue expansions for Brownian motion with an application to occupation times. Electronic Journal of Probability, 1:3:1–3:19, 1996. CO-DEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.

org/article/view/3;   http://www.math.washington.edu/
~ejpecp/EjpVol1/paper3.abs.html.

Birkner:2009:GSF

[BB09] Matthias Birkner and Jochen Blath. Generalised stable Fleming– Viot processes as flickering random measures. Electronic Journal of Probability, 14:84:2418–84:2437, 2009. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/712.

Bank:2010:PDO

[BB10] Peter Bank and Christoph Baumgarten. Parameter-dependent optimal stopping problems for one-dimensional diffusions. Electronic Journal of Probability, 15:64:1971–64:1993, 2010. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/
view/835.

Bansaye:2011:ULD

[BB11a] Vincent Bansaye and Christian Böinghoff. Upper large deviations for branching processes in random environment with heavy tails. Electronic Journal of Probability, 16:69:1900–69:1933, 2011.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/933.

Barczy:2011:FLT

[BB11b] Matyas Barczy and Jean Bertoin. Functional limit theorems for Lévy processes satisfying Cramér’s condition. Electronic Journal of Probability, 16:73:2020–73:2038, 2011. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/930.

Baur:2015:FPI

[BB15] Erich Baur and Jean Bertoin. The fragmentation process of an infinite recursive tree and Ornstein–Uhlenbeck type processes. Electronic Journal of Probability, 20(??):98:1–98:20, ???? 2015. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/3866.

Baar:2018:PES

[BB18a] Martina Baar and Anton Bovier. The polymorphic evolution sequence for populations with phenotypic plasticity. Electronic Journal of Probability, 23(??):72:1–72:27, ???? 2018. CODEN

???? ISSN 1083-6489. URL https://projecteuclid.org/
euclid.ejp/1532678635.

Bahadoran:2018:QEF

[BB18b] C. Bahadoran and T. Bodineau. Quantitative estimates for the flux of TASEP with dilute site disorder. Electronic Journal of Probability, 23(??):44:1–44:44, ???? 2018. CODEN ????

ISSN 1083-6489. URL https://projecteuclid.org/euclid.
ejp/1525852821.

Berger:2006:TPC

[BBAP06] Noam Berger, Itai Benjamini, Omer Angel, and Yuval Peres. Transience of percolation clusters on wedges. Electronic Journal of Probability, 11:25:655–25:669, 2006. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/345.

Birkner:2005:ASB

[BBC^{+}05] Matthias Birkner, Jochen Blath, Marcella Capaldo, Alison Etheridge, Martin Möhle, Jason Schweinsberg, and Anton Wakolbinger. Alpha-stable branching and beta-coalescents. Electronic Journal of Probability, 10:9:303–9:325, 2005. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/241.

Belinschi:2021:OEN

[BBCC21] Serban Belinschi, Charles Bordenave, Mireille Capitaine, and Guillaume Cébron. Outlier eigenvalues for non-Hermitian polynomials in independent i.i.d. matrices and deterministic matrices. Electronic Journal of Probability, 26(18):1–37, February 2021. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.

org/journals/electronic-journal-of-probability/volume-
26/issue-none/Outlier-eigenvalues-for-non-Hermitian-polynomials-
in-independent-iid-matrices/10.1214/21-EJP666.full.

Bethuelsen:2023:LLT

[BBDS23] Stein Andreas Bethuelsen, Matthias Birkner, Andrej Depperschmidt, and Timo Schlüter. Local limit theorems for a directed random walk on the backbone of a supercritical oriented percolation cluster. Electronic Journal of Probability, 28(??):1–54, 2023. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.

org/journals/electronic-journal-of-probability/volume-
28/issue-none/Local-limit-theorems-for-a-directed-random-
walk-on-the/10.1214/23-EJP924.full.

Borga:2020:DTA

[BBFS20] Jacopo Borga, Mathilde Bouvel, Valentin Féray, and Benedikt Stufler. A decorated tree approach to random permutations in substitution-closed classes. Electronic Journal of Probability, 25 (??):67:1–67:52, ???? 2020. CODEN ???? ISSN 1083-6489. URL

https://projecteuclid.org/euclid.ejp/1592618469.

Barnes:2019:BMR

[BBG19] Clayton Barnes, Krzysztof Burdzy, and Carl-Erik Gauthier. Billiards with Markovian reflection laws. Electronic Journal of Probability, 24(??):147:1–147:32, ???? 2019. CODEN ????

ISSN 1083-6489. URL https://projecteuclid.org/euclid.
ejp/1577761457.

Basu:2012:JCS

[BBGH12] Riddhipratim Basu, Arup Bose, Shirshendu Ganguly, and Rajat Hazra. Joint convergence of several copies of different patterned random matrices. Electronic Journal of Probability, 17: 82:1–82:33, 2012. CODEN ???? ISSN 1083-6489. URL http://

ejp.ejpecp.org/article/view/1970.

Baci:2020:CIF

[BBGT20] Anastas Baci, Carina Betken, Anna Gusakova, and Christoph Thäle. Concentration inequalities for functionals of Poisson cylinder processes. Electronic Journal of Probability, 25(??): 1–27, 2020. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/journals/electronic-journal-of-probability/
volume-25/issue-none/Concentration-inequalities-for-functionals-
of-Poisson-cylinder-processes/10.1214/20-EJP529.full.

Bianchi:2009:SAM

[BBI09] Alessandra Bianchi, Anton Bovier, and Dmitry Ioffe. Sharp asymptotics for metastability in the random field Curie–Weiss model. Electronic Journal of Probability, 14:53:1541–53:1603, 2009. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.

org/article/view/673.

Bahl:2022:DLA

[BBJJ22] Riti Bahl, Philip Barnet, Tobias Johnson, and Matthew Junge. Diffusion-limited annihilating systems and the increasing convex order. Electronic Journal of Probability, 27(??):1–19, 2022. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/
journals/electronic-journal-of-probability/volume-27/
issue-none/Diffusion-limited-annihilating-systems-and-
the-increasing-convex-order/10.1214/22-EJP808.full.

Bierme:2011:CLT

[BBL11] Hermine Biermé, Aline Bonami, and José R. Leon. Central Limit Theorems and quadratic variations in terms of spectral density. Electronic Journal of Probability, 16:13:362–13:395, 2011. CO-DEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/862.

Barret:2010:UEM

[BBM10] Florent Barret, Anton Bovier, and Sylvie Méléard. Uniform estimates for metastable transition times in a coupled bistable system. Electronic Journal of Probability, 15:12:323–12:345, 2010. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/751.

Benaim:2020:AAB

[BBM20] Michel Benaïm, Charles-Edouard Bréhier, and Pierre Monmarché. Analysis of an adaptive biasing force method based on self-interacting dynamics. Electronic Journal of Probability, 25(??):1–28, 2020. CODEN ???? ISSN

1083-6489.  URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-25/issue-none/
Analysis-of-an-Adaptive-Biasing-Force-method-based-on-
self/10.1214/20-EJP490.full.

Bahlali:2022:ADS

[BBM22] Khaled Bahlali, Brahim Boufoussi, and Soufiane Mouchtabih. Approximation of a degenerate semilinear PDE with a nonlinear Neumann boundary condition. Electronic Journal of Probability, 27(??):1–21, 2022. CODEN ???? ISSN

1083-6489.  URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-27/issue-none/
Approximation-of-a-degenerate-semilinear-PDE-with-a-nonlinear-
Neumann/10.1214/22-EJP823.full.

Bieniek:2012:EFV

[BBP12] Mariusz Bieniek, Krzysztof Burdzy, and Soumik Pal. Extinction of Fleming–Viot-type particle systems with strong drift. Electronic Journal of Probability, 17:11:1–11:15, 2012. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
1770.

Banerjee:2020:NAL

[BBP20] Sayan Banerjee, Amarjit Budhiraja, and Michael Perlmutter. A new approach to large deviations for the Ginzburg–Landau model. Electronic Journal of Probability, 25(??):26:1–26:51, ???? 2020. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1582254382.

Barbour:2004:NUB

[BC04] Andrew Barbour and Kwok Choi. A non-uniform bound for translated Poisson approximation. Electronic Journal of Probability, 9:2:18–2:36, 2004. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/182.

Baudoin:2008:SSF

[BC08a] Fabrice Baudoin and Laure Coutin. Self-similarity and fractional Brownian motion on Lie groups. Electronic Journal of Probability, 13:38:1120–38:1139, 2008. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/530.

Broman:2008:LLC

[BC08b] Erik Broman and Federico Camia. Large-N limit of crossing probabilities, discontinuity, and asymptotic behavior of threshold values in Mandelbrot’s fractal percolation process. Electronic Journal of Probability, 13:33:980–33:999, 2008. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/511.

Broman:2010:UBC

[BC10] Erik Broman and Federico Camia. Universal behavior of connectivity properties in fractal percolation models. Electronic Journal of Probability, 15:44:1394–44:1414, 2010. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/805.

Benjamini:2012:ETS

[BC12] Itai Benjamini and Nicolas Curien. Ergodic theory on stationary random graphs. Electronic Journal of Probability, 17:93:1–93:20, 2012. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.

org/article/view/2401.

Bally:2014:DBP

[BC14] Vlad Bally and Lucia Caramellino. On the distances between probability density functions. Electronic Journal of Probability, 19(??):110:1–110:33, ???? 2014. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/3175.

Budd:2017:GIP

[BC17] Timothy Budd and Nicolas Curien. Geometry of infinite planar maps with high degrees. Electronic Journal of Probability, 22 (??):35:1–35:37, ???? 2017. CODEN ???? ISSN 1083-6489. URL

https://projecteuclid.org/euclid.ejp/1492588824.

Bouguet:2018:FEM

[BC18] Florian Bouguet and Bertrand Cloez. Fluctuations of the empirical measure of freezing Markov chains. Electronic Journal of Probability, 23(??):2:1–2:31, ???? 2018. CODEN ????

ISSN 1083-6489. URL https://projecteuclid.org/euclid.
ejp/1516093310.

Bourguin:2020:AHV

[BC20] Solesne Bourguin and Simon Campese. Approximation of Hilbert-valued Gaussians on Dirichlet structures. Electronic Journal of Probability, 25(??):1–30, 2020. CODEN ????

ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-25/issue-none/
Approximation-of-Hilbert-Valued-Gaussians-on-Dirichlet-
structures/10.1214/20-EJP551.full.

Bechtold:2021:LLN

[BC21] Florian Bechtold and Fabio Coppini. A law of large numbers for interacting diffusions via a mild formulation. Electronic Journal of Probability, 26(18):1–27, February 2021. CODEN

???? ISSN 1083-6489. URL https://projecteuclid.org/
journals/electronic-journal-of-probability/volume-26/
issue-none/A-law-of-large-numbers-for-interacting-diffusions-
via-a/10.1214/21-EJP671.full.

Bowditch:2022:BRW

[BC22] Adam M. Bowditch and David A. Croydon. Biased random walk on supercritical percolation: anomalous fluctuations in the ballistic regime. Electronic Journal of Probability, 27(??):1–22, 2022. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.

org/journals/electronic-journal-of-probability/volume-
27/issue-none/Biased-random-walk-on-supercritical-percolation-
-anomalous-fluctuations-in/10.1214/22-EJP794.full.

Bandini:2019:BRR

[BCC19] Elena Bandini, Fulvia Confortola, and Andrea Cosso. BSDE representation and randomized dynamic programming principle for stochastic control problems of infinite-dimensional jumpdiffusions. Electronic Journal of Probability, 24(??):81:1–81:37, ???? 2019. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1568080857.

Blanca:2022:MMC

[BCC^{+}22] Antonio Blanca, Pietro Caputo, Zongchen Chen, Daniel Parisi, Daniel Štefankovič, and Eric Vigoda. On mixing of Markov chains: coupling, spectral independence, and entropy factorization. Electronic Journal of Probability, 27(??):1–42, 2022. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/
journals/electronic-journal-of-probability/volume-27/
issue-none/On-mixing-of-Markov-chains--coupling-spectral-
independence-and/10.1214/22-EJP867.full.

Birkner:2016:RWD

[BČD16] Matthias Birkner, Jiří Černý, and Andrej Depperschmidt. Random walks in dynamic random environments and ancestry under local population regulation. Electronic Journal of Probability, 21 (??):38:1–38:43, ???? 2016. CODEN ???? ISSN 1083-6489. URL

https://projecteuclid.org/euclid.ejp/1464269713.

Bailleul:2020:SMF

[BCD20] Ismaël Bailleul, Rémi Catellier, and François Delarue. Solving mean field rough differential equations. Electronic Journal of Probability, 25(??):21:1–21:51, ???? 2020. CODEN ????

ISSN 1083-6489. URL https://projecteuclid.org/euclid.
ejp/1581044444.

Birkner:2013:DRW

[BCDG13] Matthias Birkner, Jiri Cerny, Andrej Depperschmidt, and Nina Gantert. Directed random walk on the backbone of an oriented percolation cluster. Electronic Journal of Probability, 18:80:1– 80:35, 2013. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/2302.

Bardet:2013:TVE

[BCG^{+}13] Jean-Baptiste Bardet, Alejandra Christen, Arnaud Guillin, Florent Malrieu, and Pierre-André Zitt. Total variation estimates for the TCP process. Electronic Journal of Probability, 18:10:1– 10:21, 2013. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/1720.

Bobkov:2018:BEB

[BCG18] S. G. Bobkov, G. P. Chistyakov, and F. Götze. Berry–Esseen bounds for typical weighted sums. Electronic Journal of Probability, 23(??):92:1–92:22, ???? 2018. CODEN ???? ISSN 1083-6489.

URL https://projecteuclid.org/euclid.ejp/1536976980.

Bobkov:2020:PIN

[BCG20] S. G. Bobkov, G. P. Chistyakov, and F. Götze. Poincaré inequalities and normal approximation for weighted sums. Electronic Journal of Probability, 25(??):1–31, 2020. CODEN

???? ISSN 1083-6489. URL https://projecteuclid.org/
journals/electronic-journal-of-probability/volume-25/
issue-none/Poincar%c3%a9-inequalities-and-normal-approximation-
for-weighted-sums/10.1214/20-EJP549.full.

Bercu:2021:SAA

[BCG21] Bernard Bercu, Manon Costa, and Sébastien Gadat. Stochastic approximation algorithms for superquantiles estimation. Electronic Journal of Probability, 26(18):1–29, February 2021. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/
journals/electronic-journal-of-probability/volume-26/
issue-none/Stochastic-approximation-algorithms-for-superquantiles-
estimation/10.1214/21-EJP648.full.

Bally:2022:UMA

[BCKH22] Vlad Bally, Lucia Caramellino, and Arturo Kohatsu-Higa. Using moment approximations to study the density of jump driven SDEs. Electronic Journal of Probability, 27(??):1–21, 2022. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/
journals/electronic-journal-of-probability/volume-27/
issue-none/Using-moment-approximations-to-study-the-density-
of-jump-driven/10.1214/22-EJP785.full.

Blath:2020:SBC

[BCKWB20] Jochen Blath, Adrián González Casanova, Noemi Kurt, and Maite Wilke-Berenguer. The seed bank coalescent with simultaneous switching. Electronic Journal of Probability, 25(??):27:1– 27:21, ???? 2020. CODEN ???? ISSN 1083-6489. URL https:/

/projecteuclid.org/euclid.ejp/1582254383.

Bansaye:2019:SLP

[BCM19] Vincent Bansaye, Maria-Emilia Caballero, and Sylvie Méléard. Scaling limits of population and evolution processes in random environment. Electronic Journal of Probability, 24(??):19:1– 19:38, ???? 2019. CODEN ???? ISSN 1083-6489. URL https:/

/projecteuclid.org/euclid.ejp/1552013626.

Biau:2009:AND

[BCMP09] Gérard Biau, Benoit Cadre, David Mason, and Bruno Pelletier. Asymptotic normality in density support estimation. Electronic Journal of Probability, 14:91:2617–91:2635, 2009. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
722.

Bally:2018:CDN

[BCP18] Vlad Bally, Lucia Caramellino, and Guillaume Poly. Convergence in distribution norms in the CLT for non identical distributed random variables. Electronic Journal of Probability, 23(??):45:1– 45:51, ???? 2018. CODEN ???? ISSN 1083-6489. URL https:/

/projecteuclid.org/euclid.ejp/1527213726.

Bally:2020:RLC

[BCP20] Vlad Bally, Lucia Caramellino, and Guillaume Poly. Regularization lemmas and convergence in total variation. Electronic Journal of Probability, 25(??):74:1–74:20, ???? 2020. CODEN ????

ISSN 1083-6489. URL https://projecteuclid.org/euclid.
ejp/1593828035.

Boenkost:2021:HFC

[BCPW21] Florin Boenkost, Adrián González Casanova, Cornelia Pokalyuk, and Anton Wakolbinger. Haldane’s formula in Cannings models: the case of moderately weak selection. Electronic Journal of Probability, 26(??):1–36, 2021. CODEN

???? ISSN 1083-6489. URL https://projecteuclid.org/
journals/electronic-journal-of-probability/volume-26/
issue-none/Haldanes-formula-in-Cannings-models-thecaseofmoderately-
weak-selection/10.1214/20-EJP572.full.

Bass:2006:MDL

[BCR06] Richard Bass, Xia Chen, and Jay Rosen. Moderate deviations and laws of the iterated logarithm for the renormalized selfintersection local times of planar random walks. Electronic Journal of Probability, 11:37:993–37:1030, 2006. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/362.

Barthe:2007:IBE

[BCR07] Franck Barthe, Patrick Cattiaux, and Cyril Roberto. Isoperimetry between exponential and Gaussian. Electronic Journal of Probability, 12:44:1212–44:1237, 2007. CODEN ???? ISSN 1083-

6489. URL http://ejp.ejpecp.org/article/view/441.

Balazs:2006:CRF

[BCS06] Marton Balazs, Eric Cator, and Timo Seppalainen. Cube root fluctuations for the corner growth model associated to the exclusion process. Electronic Journal of Probability, 11:42:1094– 42:1132, 2006. CODEN ???? ISSN 1083-6489. URL http://

ejp.ejpecp.org/article/view/366.

Botero:2021:LDP

[BCV21] Alonso Botero, Matthias Christandl, and Péter Vrana. Large deviation principle for moment map estimation. Electronic Journal of Probability, 26(18):1–23, February 2021. CODEN

???? ISSN 1083-6489. URL https://projecteuclid.org/
journals/electronic-journal-of-probability/volume-26/
issue-none/Large-deviation-principle-for-moment-map-estimation/
10.1214/21-EJP636.full.

Basak:2018:CLS

[BCZ18] Anirban Basak, Nicholas Cook, and Ofer Zeitouni. Circular law for the sum of random permutation matrices. Electronic Journal of Probability, 23(??):33:1–33:51, ???? 2018. CODEN ????

ISSN 1083-6489. URL https://projecteuclid.org/euclid.
ejp/1524880977.

Budhiraja:2003:LDE

[BD03] Amarjit Budhiraja and Paul Dupuis. Large deviations for the emprirical measures of reflecting Brownian motion and related constrained processes in R_{+}. Electronic Journal of Probability, 8: 16:1–16:46, 2003. CODEN ???? ISSN 1083-6489. URL http://

ejp.ejpecp.org/article/view/154.

Berestycki:2008:LBD

[BD08] Nathanael Berestycki and Rick Durrett. Limiting behavior for the distance of a random walk. Electronic Journal of Probability, 13:14:374–14:395, 2008. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/490.

Breuer:2012:NPS

[BD12] Jonathan Breuer and Maurice Duits. Nonintersecting paths with a staircase initial condition. Electronic Journal of Probability, 17: 60:1–60:24, 2012. CODEN ???? ISSN 1083-6489. URL http://

ejp.ejpecp.org/article/view/1902.

Bouleau:2013:CEL

[BD13] Nicolas Bouleau and Laurent Denis. Chaotic extensions and the lent particle method for Brownian motion. Electronic Journal of Probability, 18:56:1–56:16, 2013. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/1838.

Beghin:2014:FPP

[BD14a] Luisa Beghin and Mirko D’Ovidio. Fractional Poisson process with random drift. Electronic Journal of Probability, 19(??): 122:1–122:26, ???? 2014. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/3258.

Bi:2014:PMN

[BD14b] Hongwei Bi and Jean-François Delmas. A population model with non-neutral mutations using branching processes with immigration. Electronic Journal of Probability, 19:62:1–62:23, 2014. CO-DEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/2939.

Buraczewski:2018:PLD

[BD18] Dariusz Buraczewski and Piotr Dyszewski. Precise large deviations for random walk in random environment. Electronic Journal of Probability, 23(??):114:1–114:26, ???? 2018. CODEN ????

ISSN 1083-6489. URL https://projecteuclid.org/euclid.
ejp/1542942365.

Bishop:2019:SMV

[BD19] Adrian N. Bishop and Pierre Del Moral. On the stability of matrix-valued Riccati diffusions. Electronic Journal of Probability, 24(??):84:1–84:40, ???? 2019. CODEN ???? ISSN 1083-6489.

URL https://projecteuclid.org/euclid.ejp/1568080863.

Bercu:2009:FCL

[BDD09] Bernard Bercu, Pierre Del Moral, and Arnaud Doucet. A functional Central Limit Theorem for a class of interacting Markov chain Monte Carlo methods. Electronic Journal of Probability, 14:73:2130–73:2155, 2009. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/701.

Berard:2014:LCL

[BDD14] Jean Bérard, Pierre Del Moral, and Arnaud Doucet. A lognormal central limit theorem for particle approximations of normalizing constants. Electronic Journal of Probability, 19(??):94:1–94:28, ???? 2014. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/3428.

Beckman:2018:ABB

[BDD^{+}18] Erin Beckman, Emily Dinan, Rick Durrett, Ran Huo, and Matthew Junge. Asymptotic behavior of the Brownian frog model. Electronic Journal of Probability, 23(??):104:1–104:19, ???? 2018. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1540000928.

Basak:2018:DLP

[BDF18] Anirban Basak, Rick Durrett, and Eric Foxall. Diffusion limit for the partner model at the critical value. Electronic Journal of Probability, 23(??):102:1–102:42, ???? 2018. CODEN ????

ISSN 1083-6489. URL https://projecteuclid.org/euclid.
ejp/1539309901.

Budhiraja:2015:LRE

[BDFR15a] Amarjit Budhiraja, Paul Dupuis, Markus Fischer, and Kavita Ramanan. Limits of relative entropies associated with weakly interacting particle systems. Electronic Journal of Probability, 20(??):80:1–80:22, ???? 2015. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/4003.

Budhiraja:2015:LSK

[BDFR15b] Amarjit Budhiraja, Paul Dupuis, Markus Fischer, and Kavita Ramanan. Local stability of Kolmogorov forward equations for finite state nonlinear Markov processes. Electronic Journal of Probability, 20(??):81:1–81:30, ???? 2015. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/4004.

Boufoussi:2008:PPC

[BDG08] Brahim Boufoussi, Marco Dozzi, and Raby Guerbaz. Path properties of a class of locally asymptotically self similar processes. Electronic Journal of Probability, 13:29:898–29:921, 2008. CO-DEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/505.

Bianchi:2017:MRI

[BDG17] Alessandra Bianchi, Sander Dommers, and Cristian Giardinà. Metastability in the reversible inclusion process. Electronic Journal of Probability, 22(??):70:1–70:34, ???? 2017. CODEN ????

ISSN 1083-6489. URL https://projecteuclid.org/euclid.
ejp/1505268101.

Budhiraja:2018:LDS

[BDG18] Amarjit Budhiraja, Paul Dupuis, and Arnab Ganguly. Large deviations for small noise diffusions in a fast Markovian environment. Electronic Journal of Probability, 23(??):112:1–112:33, ???? 2018. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1540951492.

Buraczewski:2019:RWM

[BDI^{+}19] Dariusz Buraczewski, Piotr Dyszewski, Alexander Iksanov, Alexander Marynych, and Alexander Roitershtein. Random walks in a moderately sparse random environment. Electronic Journal of Probability, 24(??):69:1–69:44, ???? 2019. CODEN

???? ISSN 1083-6489. URL https://projecteuclid.org/
euclid.ejp/1561687602.

Buraczewski:2020:ILN

[BDI20] Dariusz Buraczewski, Bohdan Dovgay, and Alexander Iksanov. On intermediate levels of nested occupancy scheme in random environment generated by stick-breaking I. Electronic Journal of Probability, 25(??):1–24, 2020. CODEN

???? ISSN 1083-6489. URL https://projecteuclid.org/
journals/electronic-journal-of-probability/volume-25/
issue-none/On-intermediate-levels-of-nested-occupancy-
scheme-in-random-environment/10.1214/20-EJP534.full.

Blancas:2018:TWT

[BDLSJ18] Airam Blancas, Jean-Jil Duchamps, Amaury Lambert, and Arno Siri-Jégousse. Trees within trees: simple nested coalescents. Electronic Journal of Probability, 23(??):94:1–94:27, ???? 2018. CO-DEN ???? ISSN 1083-6489. URL https://projecteuclid.

org/euclid.ejp/1537257886.

Buraczewski:2015:RCK

[BDP15] Dariusz Buraczewski, Ewa Damek, and Tomasz Przebinda. On the rate of convergence in the Kesten renewal theorem. Electronic Journal of Probability, 20(??):22:1–22:35, ???? 2015. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/
view/3708.

Bogso:2022:PPU

[BDP22] Antoine-Marie Bogso, Moustapha Dieye, and Olivier Menoukeu Pamen. Path-by-path uniqueness of multidimensional SDE’s on the plane with nondecreasing coefficients. Electronic Journal of Probability, 27(??):1–26, 2022. CODEN ???? ISSN

1083-6489.  URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-27/issue-none/

Path-by-path-uniqueness-of-multidimensional-SDEs-on-the-
plane/10.1214/22-EJP844.full.

Brandenberger:2021:HSN

[BDR21] Anna Brandenberger, Luc Devroye, and Tommy Reddad. The Horton–strahler number of conditioned Galton–Watson trees. Electronic Journal of Probability, 26(18):1–29, February 2021. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.

org/journals/electronic-journal-of-probability/volume-
26/issue-none/The-HortonStrahler-number-of-conditioned-
GaltonWatson-trees/10.1214/21-EJP678.full.

Bisewski:2022:DSF

[BDR22] Krzysztof Bisewski, Krzysztof Dȩbicki, and Tomasz Rolski. Derivatives of sup-functionals of fractional Brownian motion evaluated at H = 1/2. Electronic Journal of Probability, 27(??):1–35, 2022. CODEN ???? ISSN

1083-6489.  URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-27/issue-none/
Derivatives-of-sup-functionals-of-fractional-Brownian-
motion-evaluated-at/10.1214/22-EJP848.full.

Bercu:2009:AAB

[BdSGP09] Bernard Bercu, Benoîte de Saporta, and Anne Gégout-Petit. Asymptotic analysis for bifurcating autoregressive processes via a martingale approach. Electronic Journal of Probability, 14: 87:2492–87:2526, 2009. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/717.

Boutillier:2014:HRX

[BdT14] Cédric Boutillier and Béatrice de Tilière. Height representation of XOR–Ising loops via bipartite dimers. Electronic Journal of Probability, 19(??):80:1–80:33, ???? 2014. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/2449.

Blondel:2019:FEF

[BDT19] Oriane Blondel, Aurelia Deshayes, and Cristina Toninelli. Front evolution of the Fredrickson–Andersen one spin facilitated model. Electronic Journal of Probability, 24(??):1:1–1:32, ???? 2019. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.

org/euclid.ejp/1546571126.

Bhamidi:2020:UCH

[BDvdHS20] Shankar Bhamidi, Souvik Dhara, Remco van der Hofstad, and Sanchayan Sen. Universality for critical heavy-tailed network models: Metric structure of maximal components. Electronic Journal of Probability, 25(??):47:1–47:57, ???? 2020. CODEN

???? ISSN 1083-6489. URL https://projecteuclid.org/
euclid.ejp/1587693778.

Bhamidi:2022:GLM

[BDvdHS22] Shankar Bhamidi, Souvik Dhara, Remco van der Hofstad, and Sanchayan Sen. Global lower mass-bound for critical configuration models in the heavy-tailed regime. Electronic Journal of Probability, 27(??):1–29, 2022. CODEN

???? ISSN 1083-6489. URL https://projecteuclid.org/
journals/electronic-journal-of-probability/volume-27/
issue-none/Global-lower-mass-bound-for-critical-configuration-
models-in-the/10.1214/22-EJP821.full.

Berglund:2017:EKL

[BDW17] Nils Berglund, Giacomo Di Gesù, and Hendrik Weber. An Eyring–Kramers law for the stochastic Allen–Cahn equation in dimension two. Electronic Journal of Probability, 22(??):41:1– 41:27, ???? 2017. CODEN ???? ISSN 1083-6489. URL https:/

/projecteuclid.org/euclid.ejp/1493345028.

Baudoin:2021:TIM

[BE21] Fabrice Baudoin and Nathaniel Eldredge. Transportation inequalities for Markov kernels and their applications. Electronic Journal of Probability, 26(18):1–30, February 2021. CODEN

???? ISSN 1083-6489. URL https://projecteuclid.org/
journals/electronic-journal-of-probability/volume-26/
issue-none/Transportation-inequalities-for-Markov-kernels-
and-their-applications/10.1214/21-EJP605.full.

Begyn:2005:QVA

[Beg05] Arnaud Begyn. Quadratic variations along irregular subdivisions for Gaussian processes. Electronic Journal of Probability, 10: 20:691–20:717, 2005. CODEN ???? ISSN 1083-6489. URL http:

//ejp.ejpecp.org/article/view/245.

Beghin:2008:PPG

[Beg08] Luisa Beghin. Pseudo-processes governed by higher-order fractional differential equations. Electronic Journal of Probability, 13:16:467–16:485, 2008. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/496.

Basdevant:2014:SLB

[BEGG14] Anne-Laure Basdevant, Nathanaël Enriquez, Lucas Gerin, and Jean-Baptiste Gouéré. The shape of large balls in highly supercritical percolation. Electronic Journal of Probability, 19:26:1– 26:14, 2014. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/3062.

Biswas:2021:SLF

[BEK21] Niloy Biswas, Alison Etheridge, and Aleksander Klimek. The spatial Lambda-Fleming-Viot process with fluctuating selection. Electronic Journal of Probability, 26(??):1–51, 2021. CODEN

???? ISSN 1083-6489. URL https://projecteuclid.org/
journals/electronic-journal-of-probability/volume-26/
issue-none/The-spatial-Lambda-Fleming-Viot-process-with-
fluctuating-selection/10.1214/21-EJP593.full.

Bell:2020:TRC

[Bel20a] James Bell. Time-reversal of coalescing diffusive flows and weak convergence of localized disturbance flows. Electronic Journal of Probability, 25(??):1–38, 2020. CODEN ????

ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-25/issue-none/
Time-reversal-of-coalescing-diffusive-flows-and-weak-
convergence-of/10.1214/20-EJP500.full.

Bellingeri:2020:ITF

[Bel20b] Carlo Bellingeri. An Itô type formula for the additive stochastic heat equation. Electronic Journal of Probability, 25(??):2:1–2:52, ???? 2020. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1578366206.

Betz:2021:SPT

[BELR21] Volker Betz, Johannes Ehlert, Benjamin Lees, and Lukas Roth. Sharp phase transition for random loop models on trees. Electronic Journal of Probability, 26(18):1–26, February 2021. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/
journals/electronic-journal-of-probability/volume-26/
issue-none/Sharp-phase-transition-for-random-loop-models-
on-trees/10.1214/21-EJP677.full.

Benoist:2018:NPS

[Ben18] Stéphane Benoist. Natural parametrization of SLE: the Gaussian free field point of view. Electronic Journal of Probability, 23(??): 103:1–103:16, ???? 2018. CODEN ???? ISSN 1083-6489. URL

https://projecteuclid.org/euclid.ejp/1539828067.

Benes:2021:RCP

[Ben21] Christian Beneš. Rates of convergence for the planar discrete Green’s function in Pacman domains. Electronic Journal of Probability, 26(18):1–14, February 2021. CODEN

???? ISSN 1083-6489. URL https://projecteuclid.org/
journals/electronic-journal-of-probability/volume-26/
issue-none/Rates-of-convergence-for-the-planar-discrete-
Greens-function-in/10.1214/21-EJP599.full.

Bahlali:2009:HSP

[BEP09] Khaled Bahlali, A. Elouaflin, and Etienne Pardoux. Homogenization of semilinear PDEs with discontinuous averaged coefficients. Electronic Journal of Probability, 14:18:477–18:499, 2009. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/627.

Bertoin:1997:CPV

[Ber97] Jean Bertoin. Cauchy’s principal value of local times of Lévy processes with no negative jumps via continuous branching processes. Electronic Journal of Probability, 2:6:1–6:12, 1997. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
20.

Berestycki:2004:EFC

[Ber04] Julien Berestycki. Exchangeable fragmentation–coalescence processes and their equilibrium measures. Electronic Journal of Probability, 9:25:770–25:824, 2004. CODEN ???? ISSN 1083-

6489. URL http://ejp.ejpecp.org/article/view/227.

Bertacchi:2006:ABS

[Ber06] Daniela Bertacchi. Asymptotic behaviour of the simple random walk on the 2-dimensional comb. Electronic Journal of Probability, 11:45:1184–45:1203, 2006. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/377.

Berestycki:2011:EGC

[Ber11] Nathanael Berestycki. Emergence of giant cycles and slowdown transition in random transpositions and k-cycles. Electronic Journal of Probability, 16:5:152–5:173, 2011. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/850.

Bertoin:2014:NGF

[Ber14] Jean Bertoin. On the non-Gaussian fluctuations of the giant cluster for percolation on random recursive trees. Electronic Journal of Probability, 19:24:1–24:15, 2014. CODEN ???? ISSN 1083-

6489. URL http://ejp.ejpecp.org/article/view/2822.

Berzunza:2015:YPR

[Ber15] Gabriel Berzunza. Yule processes with rare mutation and their applications to percolation on b-ary trees. Electronic Journal of Probability, 20(??):43:1–43:23, ???? 2015. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/3789.

Bernstein:2018:RWS

[Ber18] Megan Bernstein. A random walk on the symmetric group generated by random involutions. Electronic Journal of Probability, 23(??):26:1–26:28, ???? 2018. CODEN ???? ISSN 1083-6489.

URL https://projecteuclid.org/euclid.ejp/1521079339.

Berger:2019:SRT

[Ber19] Quentin Berger. Strong renewal theorems and local large deviations for multivariate random walks and renewals. Electronic Journal of Probability, 24(??):46:1–46:47, ???? 2019. CODEN

???? ISSN 1083-6489. URL https://projecteuclid.org/
euclid.ejp/1557453645.

Berman:2021:PLW

[Ber21a] Robert J. Berman. Priors leading to well-behaved Coulomb and Riesz gases versus zeroth-order phase transitions — a potential-theoretic characterization. Electronic Journal of Probability, 26(18):1–49, February 2021. CODEN ???? ISSN

1083-6489.  URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-26/issue-none/
Priors-leading-to-well-behaved-Coulomb-and-Riesz-gases-
versus/10.1214/21-EJP700.full.

Berzin:2021:ELA

[Ber21b] Corinne Berzin. Estimation of local anisotropy based on level sets. Electronic Journal of Probability, 26(18):1–72, February 2021. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.

org/journals/electronic-journal-of-probability/volume-
26/issue-none/Estimation-of-local-anisotropy-based-on-
level-sets/10.1214/21-EJP721.full.

Bertacco:2023:MAG

[Ber23] Federico Bertacco. Multifractal analysis of Gaussian multiplicative chaos and applications. Electronic Journal of Probability, 28(??):1–36, 2023. CODEN ???? ISSN

1083-6489.  URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-28/issue-none/
Multifractal-analysis-of-Gaussian-multiplicative-chaos-
and-applications/10.1214/22-EJP893.full.

Bessaih:1999:SWA

[Bes99] Hakima Bessaih. Stochastic weak attractor for a dissipative Euler equation. Electronic Journal of Probability, 5:3:1–3:16, 1999. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/59.

Bettinelli:2010:SLR

[Bet10] Jérémie Bettinelli. Scaling limits for random quadrangulations of positive genus. Electronic Journal of Probability, 15:52:1594– 52:1644, 2010. CODEN ???? ISSN 1083-6489. URL http://

ejp.ejpecp.org/article/view/810.

Barton:2010:NME

[BEV10] Nick Barton, Alison Etheridge, and Amandine Véber. A new model for evolution in a spatial continuum. Electronic Journal of Probability, 15:7:162–7:216, 2010. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/741.

Barriere:2012:SRP

[BEV12] Olivier Barrière, Antoine Echelard, and Jacques Lévy Véhel. Selfregulating processes. Electronic Journal of Probability, 17:103:1– 103:30, 2012. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/2010.

Borodin:2008:LTA

[BF08] Alexei Borodin and Patrik Ferrari. Large time asymptotics of growth models on space-like paths I: PushASEP. Electronic Journal of Probability, 13:50:1380–50:1418, 2008. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/541.

Budhiraja:2017:UTI

[BF17] Amarjit Budhiraja and Wai-Tong Louis Fan. Uniform in time interacting particle approximations for nonlinear equations of Patlak–Keller–Segel type. Electronic Journal of Probability, 22 (??):8:1–8:37, ???? 2017. CODEN ???? ISSN 1083-6489. URL

https://projecteuclid.org/euclid.ejp/1485831704.

Berthet:2020:ERC

[BF20] Philippe Berthet and Jean Claude Fort. Exact rate of convergence of the expected. Electronic Journal of Probability, 25(??): 1–16, 2020. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/journals/electronic-journal-of-probability/
volume-25/issue-none/Exact-rate-of-convergence-of-the-
expected-W_2-distance-between/10.1214/19-EJP410.full.

Butelmann:2022:SLS

[BF22] Ian Butelmann and Gregorio R. Moreno Flores. Scaling limit of stationary coupled Sasamoto–Spohn models. Electronic Journal of Probability, 27(??):1–25, 2022. CODEN

???? ISSN 1083-6489. URL https://projecteuclid.org/
journals/electronic-journal-of-probability/volume-27/
issue-none/Scaling-limit-of-stationary-coupled-Sasamoto-
Spohn-models/10.1214/22-EJP819.full.

Bressaud:1999:DCN

[BFG99] Xavier Bressaud, Roberto Fernández, and Antonio Galves. Decay of correlations for non-Hölderian dynamics. A coupling approach. Electronic Journal of Probability, 4:3:1–3:19, 1999. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
40.

Beck:2019:SOS

[BFGM19] Lisa Beck, Franco Flandoli, Massimiliano Gubinelli, and Mario Maurelli. Stochastic ODEs and stochastic linear PDEs with critical drift: regularity, duality and uniqueness. Electronic Journal of Probability, 24(??):136:1–136:72, ???? 2019. CODEN ????

ISSN 1083-6489. URL https://projecteuclid.org/euclid.
ejp/1574996477.

Busic:2013:DCI

[BFMM13] Ana Bušić, Nazim Fatès, Jean Mairesse, and Irène Marcovici. Density classification on infinite lattices and trees. Electronic Journal of Probability, 18:51:1–51:22, 2013. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/2325.

Bahadoran:2011:RSI

[BFN11] Christophe Bahadoran, Jozsef Fritz, and Katalin Nagy. Relaxation schemes for interacting exclusions. Electronic Journal of Probability, 16:8:230–8:262, 2011. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/857.

Bottcher:2003:NPL

[BG03] Albrecht Böttcher and Sergei Grudsky. The norm of the product of a large matrix and a random vector. Electronic Journal of Probability, 8:7:1–7:29, 2003. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/132.

Bertoin:2004:ALN

[BG04] Jean Bertoin and Alexander Gnedin. Asymptotic laws for nonconservative self-similar fragmentations. Electronic Journal of Probability, 9:19:575–19:593, 2004. CODEN ???? ISSN 1083-

6489. URL http://ejp.ejpecp.org/article/view/215.

Basdevant:2008:AAF

[BG08] Anne-Laure Basdevant and Christina Goldschmidt. Asymptotics of the allele frequency spectrum associated with the Bolthausen– Sznitman coalescent. Electronic Journal of Probability, 13: 17:486–17:512, 2008. CODEN ???? ISSN 1083-6489. URL http:

//ejp.ejpecp.org/article/view/494.

Barbour:2009:SCI

[BG09] A. Barbour and A. Gnedin. Small counts in the infinite occupancy scheme. Electronic Journal of Probability, 14:13:365– 13:384, 2009. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/608.

Berard:2011:SPB

[BG11] Jean Bérard and Jean-Baptiste Gouéré. Survival probability of the branching random walk killed below a linear boundary. Electronic Journal of Probability, 16:14:396–14:418, 2011. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/
view/861.

Bartroff:2013:BEB

[BG13a] Jay Bartroff and Larry Goldstein. A Berry–Esséen bound for the uniform multinomial occupancy model. Electronic Journal of Probability, 18:27:1–27:29, 2013. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/1983.

Berglund:2013:SEM

[BG13b] Nils Berglund and Barbara Gentz. Sharp estimates for metastable lifetimes in parabolic SPDEs: Kramers’ law and beyond. Electronic Journal of Probability, 18:24:1–24:58, 2013. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
1802.

Bhattacharjee:2016:SES

[BG16] Chinmoy Bhattacharjee and Larry Goldstein. On strong embeddings by Stein’s method. Electronic Journal of Probability, 21 (??):15:1–15:30, ???? 2016. CODEN ???? ISSN 1083-6489. URL

https://projecteuclid.org/euclid.ejp/1456412955.

Barashkov:2021:MGT

[BG21a] Nikolay Barashkov and Massimiliano Gubinelli. The Φ_{3}^{4} measure via Girsanov’s theorem. Electronic Journal of Probability, 26(18):1–29, February 2021. CODEN ???? ISSN

1083-6489.  URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-26/issue-none/
The-%ce%a634-measure-via-Girsanovs-theorem/10.1214/21-
EJP635.full.

Braun:2021:HFR

[BG21b] Mathias Braun and Batu Güneysu. Heat flow regularity, Bismut– Elworthy–Li’s derivative formula, and pathwise couplings on Riemannian manifolds with Kato bounded Ricci curvature. Electronic Journal of Probability, 26(18):1–25, February 2021. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/
journals/electronic-journal-of-probability/volume-26/
issue-none/Heat-flow-regularity-BismutElworthyLis-derivative-
formula-and-pathwise-couplings-on/10.1214/21-EJP703.full.

Birkner:2011:CLT

[BGdH11] Matthias Birkner, Andreas Greven, and Frank den Hollander. Collision local time of transient random walks and intermediate phases in interacting stochastic systems. Electronic Journal of Probability, 16:20:552–20:586, 2011. CODEN ???? ISSN 1083-

6489. URL http://ejp.ejpecp.org/article/view/878.

Benaych-Georges:2014:CLT

[BGG14] Florent Benaych-Georges and Alice Guionnet. Central limit theorem for eigenvectors of heavy tailed matrices. Electronic Journal of Probability, 19:54:1–54:27, 2014. CODEN ???? ISSN 1083-

6489. URL http://ejp.ejpecp.org/article/view/3093.

Benaych-Georges:2011:FEE

[BGGM11] Florent Benaych-Georges, Alice Guionnet, and Mylène Maida. Fluctuations of the extreme eigenvalues of finite rank deformations of random matrices. Electronic Journal of Probability, 16: 60:1621–60:1662, 2011. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/929.

Bollobas:2014:BPG

[BGH^{+}14] Béla Bollobás, Karen Gunderson, Cecilia Holmgren, Svante Janson, and Michał Przykucki. Bootstrap percolation on Galton– Watson trees. Electronic Journal of Probability, 19:13:1–13:27, 2014. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.

org/article/view/2758.

Bates:2022:HDS

[BGH22] Erik Bates, Shirshendu Ganguly, and Alan Hammond. Hausdorff dimensions for shared endpoints of disjoint geodesics in the directed landscape. Electronic Journal of Probability, 27(??): 1–44, 2022. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/journals/electronic-journal-of-probability/
volume-27/issue-none/Hausdorff-dimensions-for-shared-
endpoints-of-disjoint-geodesics-in-the/10.1214/21-EJP706.
full.

Baudoin:2020:RPS

[BGK^{+}20] Fabrice Baudoin, Erlend Grong, Kazumasa Kuwada, Robert Neel, and Anton Thalmaier. Radial processes for sub-Riemannian Brownian motions and applications. Electronic Journal of Probability, 25(??):1–17, 2020. CODEN ????

ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-25/issue-none/
Radial-processes-for-sub-Riemannian-Brownian-motions-
and-applications/10.1214/20-EJP501.full.

Bahadoran:2010:SHL

[BGRS10] Christophe Bahadoran, Hervé Guiol, Krishnamurthi Ravishankar, and Ellen Saada. Strong hydrodynamic limit for attractive particle systems on Z. Electronic Journal of Probability, 15: 1:1–1:43, 2010. CODEN ???? ISSN 1083-6489. URL http://

ejp.ejpecp.org/article/view/728.

Blondel:2016:CSB

[BGS16] Oriane Blondel, Patrícia Gonçalves, and Marielle Simon. Convergence to the stochastic Burgers equation from a degenerate microscopic dynamics. Electronic Journal of Probability, 21(??): 69:1–69:25, ???? 2016. CODEN ???? ISSN 1083-6489. URL

https://projecteuclid.org/euclid.ejp/1480561218.

Bojdecki:2009:OTB

[BGT09] Tomasz Bojdecki, Luis Gorostiza, and Anna Talarczyk. Occupation times of branching systems with initial inhomogeneous Poisson states and related superprocesses. Electronic Journal of Probability, 14:46:1328–46:1371, 2009. CODEN ???? ISSN 1083-

6489. URL http://ejp.ejpecp.org/article/view/665.

Bojdecki:2011:NVH

[BGT11] Tomasz Bojdecki, Luis Gorostiza, and Anna Talarczyk. Number variance for hierarchical random walks and related fluctuations. Electronic Journal of Probability, 16:75:2059–75:2079, 2011. CO-DEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/937.

Bovier:2014:EPT

[BH14] Anton Bovier and Lisa Hartung. The extremal process of twospeed branching Brownian motion. Electronic Journal of Probability, 19:18:1–18:28, 2014. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/2982.

Benjamini:2019:RSC

[BH19] Itai Benjamini and Jonathan Hermon. Rapid social connectivity. Electronic Journal of Probability, 24(??):32:1–32:33, ???? 2019.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.

org/euclid.ejp/1554775411.

Benth:2021:IDP

[BH21] Fred Espen Benth and Fabian A. Harang. Infinite dimensional pathwise Volterra processes driven by Gaussian noise

— probabilistic properties and applications. Electronic Journal of Probability, 26(18):1–42, February 2021. CODEN

???? ISSN 1083-6489. URL https://projecteuclid.org/
journals/electronic-journal-of-probability/volume-26/
issue-none/Infinite-dimensional-pathwise-Volterra-processes-
driven-by-Gaussian-noise-/10.1214/21-EJP683.full.

Banerjee:2023:DCR

[BH23] Sayan Banerjee and Xiangying Huang. Degree centrality and root finding in growing random networks. Electronic Journal of Probability, 28(??):1–39, 2023. CODEN

???? ISSN 1083-6489. URL https://projecteuclid.org/
journals/electronic-journal-of-probability/volume-28/
issue-none/Degree-centrality-and-root-finding-in-growing-
random-networks/10.1214/23-EJP930.full.

Blondel:2019:RWR

[BHdS^{+}19] Oriane Blondel, Marcelo R. Hilário, Renato S. dos Santos, Vladas Sidoravicius, and Augusto Teixeira. Random walk on random walks: higher dimensions. Electronic Journal of Probability, 24 (??):80:1–80:33, ???? 2019. CODEN ???? ISSN 1083-6489. URL

https://projecteuclid.org/euclid.ejp/1567670466.

Bruckerhoff:2022:SMS

[BHJ22] Martin Brückerhoff, Martin Huesmann, and Nicolas Juillet. Shadow martingales — a stochastic mass transport approach to the peacock problem. Electronic Journal of Probability, 27(??):1–62, 2022. CODEN ???? ISSN

1083-6489.  URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-27/issue-none/
Shadow-martingales--a-stochastic-mass-transport-approach-
to-the/10.1214/22-EJP846.full.

Bhamidi:2015:IST

[BHLN15] Shankar Bhamidi, Jan Hannig, Chia Ying Lee, and James Nolen. The importance sampling technique for understanding rare events in Erdős–Rényi random graphs. Electronic Journal of Probability, 20(??):107:1–107:30, ???? 2015. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/2696.

Bai:2001:LTN

[BHLT01] Zhi-Dong Bai, Hsien-Kuei Hwang, Wen-Qi Liang, and Tsung-Hsi Tsai. Limit theorems for the number of maxima in random samples from planar regions. Electronic Journal of Probability, 6:3:1–3:41, 2001. CODEN ???? ISSN 1083-6489. URL http://

ejp.ejpecp.org/article/view/76.

Brzezniak:2013:ULS

[BHM13] Zdzisław Brzeźniak, Erika Hausenblas, and Elżbieta Motyl. Uniqueness in law of the stochastic convolution process driven by Lévy noise. Electronic Journal of Probability, 18:57:1–57:15, 2013. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.

org/article/view/2807.

Baccelli:2018:PSF

[BHM18] François Baccelli and Mir-Omid Haji-Mirsadeghi. Point-shift foliation of a point process. Electronic Journal of Probability, 23 (??):19:1–19:25, ???? 2018. CODEN ???? ISSN 1083-6489. URL

https://projecteuclid.org/euclid.ejp/1519354948.

Baccelli:2021:UHM

[BHMK21] François Baccelli, Mir-Omid Haji-Mirsadeghi, and Ali Khezeli. Unimodular Hausdorff and Minkowski dimensions. Electronic Journal of Probability, 26(18):1–64, February 2021. CODEN

???? ISSN 1083-6489. URL https://projecteuclid.org/
journals/electronic-journal-of-probability/volume-26/
issue-none/Unimodular-Hausdorff-and-Minkowski-dimensions/
10.1214/21-EJP692.full.

Baccelli:2019:DT

[BHMM19] François Baccelli, Mir-Omid Haji-Mirsadeghi, and James T. Murphy III. Doeblin trees. Electronic Journal of Probability, 24(??): 120:1–120:36, ???? 2019. CODEN ???? ISSN 1083-6489. URL

https://projecteuclid.org/euclid.ejp/1573009611.

Barrera:2021:CPT

[BHP21] Gerardo Barrera, Michael A. Högele, and Juan Carlos Pardo. The cutoff phenomenon in total variation for nonlinear Langevin systems with small layered stable noise. Electronic Journal of Probability, 26(18):1–76, February 2021. CODEN

???? ISSN 1083-6489. URL https://projecteuclid.org/
journals/electronic-journal-of-probability/volume-26/
issue-none/The-cutoff-phenomenon-in-total-variation-for-
nonlinear-Langevin-systems/10.1214/21-EJP685.full.

Bosi:2022:RWT

[BHP22] Gianluca Bosi, Yiping Hu, and Yuval Peres. Recurrence and windings of two revolving random walks. Electronic Journal of Probability, 27(??):1–22, 2022. CODEN ???? ISSN

1083-6489.  URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-27/issue-none/
Recurrence-and-windings-of-two-revolving-random-walks/
10.1214/22-EJP781.full.

Bai:2011:NRC

[BHPZ11] Zhidong Bai, Jiang Hu, Guangming Pan, and Wang Zhou. A note on rate of convergence in probability to semicircular law. Electronic Journal of Probability, 16:88:2439–88:2451, 2011. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/
view/963.

Bose:2009:LSD

[BHS09] Arup Bose, Rajat Hazra, and Koushik Saha. Limiting spectral distribution of circulant type matrices with dependent inputs. Electronic Journal of Probability, 14:86:2463–86:2491, 2009. CO-DEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/714.

Bladt:2022:TMR

[BHS22] Martin Bladt, Enkelejd Hashorva, and Georgiy Shevchenko. Tail measures and regular variation. Electronic Journal of Probability, 27(??):1–43, 2022. CODEN ???? ISSN

1083-6489.  URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-27/issue-none/
Tail-measures-and-regular-variation/10.1214/22-EJP788.
full.

Bai:2003:BEB

[BHT03] Zhi-Dong Bai, Hsien-Kuei Hwang, and Tsung-Hsi Tsai. Berry– Esséen bounds for the number of maxima in planar regions. Electronic Journal of Probability, 8:9:1–9:26, 2003. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/

Berger:2022:ODP

[BHTW22] Quentin Berger, Chien-Hao Huang, Niccolò Torri, and Ran Wei. One-dimensional polymers in random environments: stretching vs. folding. Electronic Journal of Probability, 27(??):1–45, 2022. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.

org/journals/electronic-journal-of-probability/volume-
27/issue-none/One-dimensional-polymers-in-random-environments-
stretching-vs-folding/10.1214/22-EJP862.full.

Bourgade:2017:ESS

[BHY17] Paul Bourgade, Jiaoyang Huang, and Horng-Tzer Yau. Eigenvector statistics of sparse random matrices. Electronic Journal of Probability, 22(??):64:1–64:38, ???? 2017. CODEN ????

ISSN 1083-6489. URL https://projecteuclid.org/euclid.
ejp/1502417019.

Bisewski:2020:ZLP

[BI20] Krzysztof Bisewski and Jevgenijs Ivanovs. Zooming-in on a Lévy process: failure to observe threshold exceedance over a dense grid. Electronic Journal of Probability, 25(??):1–33, 2020. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/
journals/electronic-journal-of-probability/volume-25/
issue-none/Zooming-in-on-a-L%c3%a9vy-process--failure-
to-observe/10.1214/20-EJP513.full.

Bianchi:2008:GDN

[Bia08] Alessandra Bianchi. Glauber dynamics on nonamenable graphs: boundary conditions and mixing time. Electronic Journal of Probability, 13:65:1980–65:2012, 2008. CODEN ???? ISSN 1083-

6489. URL http://ejp.ejpecp.org/article/view/574.

Benjamini:2007:RSR

[BIK07] Itai Benjamini, Roey Izkovsky, and Harry Kesten. On the range of the simple random walk bridge on groups. Electronic Journal of Probability, 12:20:591–20:612, 2007. CODEN ???? ISSN 1083-

6489. URL http://ejp.ejpecp.org/article/view/408.

Biskup:2019:IPO

[Bis19] Marek Biskup. An invariance principle for one-dimensional random walks among dynamical random conductances. Electronic Journal of Probability, 24(??):87:1–87:29, ???? 2019. CODEN

???? ISSN 1083-6489. URL https://projecteuclid.org/
euclid.ejp/1568080867.

Barbour:2009:FCC

[BJ09] Andrew Barbour and Svante Janson. A functional combinatorial Central Limit Theorem. Electronic Journal of Probability, 14: 81:2352–81:2370, 2009. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/709.

Brown:2021:SCC

[BJJK21] Suzie Brown, Paul A. Jenkins, Adam M. Johansen, and Jere Koskela. Simple conditions for convergence of sequential Monte Carlo genealogies with applications. Electronic Journal of Probability, 26(??):1–22, 2021. CODEN

???? ISSN 1083-6489. URL https://projecteuclid.org/
journals/electronic-journal-of-probability/volume-26/
issue-none/Simple-conditions-for-convergence-of-sequential-
Monte-Carlo-genealogies-with/10.1214/20-EJP561.full.

Bloznelis:2013:ACS

[BJK13] Mindaugas Bloznelis, Jerzy Jaworski, and Valentas Kurauskas. Assortativity and clustering of sparse random intersection graphs. Electronic Journal of Probability, 18:38:1–38:24, 2013. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/
view/2277.

Barbour:2010:CPA

[BJKM10] A. D. Barbour, Oliver Johnson, Ioannis Kontoyiannis, and Mokshay Madiman. Compound Poisson approximation via information functionals. Electronic Journal of Probability, 15:42:1344– 42:1369, 2010. CODEN ???? ISSN 1083-6489. URL http://

ejp.ejpecp.org/article/view/799.

Bettinelli:2014:SLU

[BJM14] Jérémie Bettinelli, Emmanuel Jacob, and Grégory Miermont. The scaling limit of uniform random plane maps, via the Ambjørn–Budd bijection. Electronic Journal of Probability, 19: 74:1–74:16, 2014. CODEN ???? ISSN 1083-6489. URL http://

ejp.ejpecp.org/article/view/3213.

Bielecki:2013:IDB

[BJN13] Tomasz Bielecki, Jacek Jakubowski, and Mariusz Niewȩgłowski. Intricacies of dependence between components of multivariate Markov chains: weak Markov consistency and weak Markov copulae. Electronic Journal of Probability, 18:45:1–45:21, 2013. CO-DEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/2238.

Bardina:2010:WCS

[BJQS10] Xavier Bardina, Maria Jolis, and Lluís Quer-Sardanyons. Weak convergence for the stochastic heat equation driven by Gaussian white noise. Electronic Journal of Probability, 15:39:1267– 39:1295, 2010. CODEN ???? ISSN 1083-6489. URL http://

ejp.ejpecp.org/article/view/792.

Balan:2015:SAM

[BJQS15] Raluca M. Balan, Maria Jolis, and Lluis Quer-Sardanyons. SPDEs with affine multiplicative fractional noise in space with index \frac{1}{4} < H < \frac{1}{2}. Electronic Journal of Probability, 20(??): 54:1–54:36, ???? 2015. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/3719.

Barlow:2001:TDA

[BK01] Martin Barlow and Takashi Kumagai. Transition density asymptotics for some diffusion processes with multi-fractal structures. Electronic Journal of Probability, 6:9:1–9:23, 2001. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
82.

Biggins:2005:FPS

[BK05] John Biggins and Andreas Kyprianou. Fixed points of the smoothing transform: the boundary case. Electronic Journal of Probability, 10:17:609–17:631, 2005. CODEN ???? ISSN 1083-

6489. URL http://ejp.ejpecp.org/article/view/255.

Baurdoux:2008:MSG

[BK08] Erik Baurdoux and Andreas Kyprianou. The McKean stochastic game driven by a spectrally negative Lévy process. Electronic Journal of Probability, 13:8:173–8:197, 2008. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/484.

Bovier:2009:ASS

[BK09] Anton Bovier and Anton Klimovsky. The Aizenman–Sims–Starr and Guerras schemes for the SK model with multidimensional spins. Electronic Journal of Probability, 14:8:161–8:241, 2009. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/611.

Berglund:2016:RSR

[BK16] Nils Berglund and Christian Kuehn. Regularity structures and renormalisation of FitzHugh–Nagumo SPDEs in three space dimensions. Electronic Journal of Probability, 21(??):18:1–18:48, ???? 2016. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1456412958. See corrigendum

[BK19].

Banerjee:2018:CPS

[BK18] Sayan Banerjee and Wilfrid Kendall. Coupling polynomial Stratonovich integrals: the two-dimensional Brownian case. Electronic Journal of Probability, 23(??):24:1–24:43, ???? 2018. CO-DEN ???? ISSN 1083-6489. URL https://projecteuclid.

org/euclid.ejp/1519722153.

Berglund:2019:CRS

[BK19] Nils Berglund and Christian Kuehn. Corrigendum to “Regularity structures and renormalisation of FitzHugh–Nagumo SPDEs in three space dimensions”. Electronic Journal of Probability, 24(??):113:1–113:22, ???? 2019. CODEN ???? ISSN

1083-6489. URL https://projecteuclid.org/euclid.ejp/

1570672858. See [BK16].

Bruned:2023:RVT

[BK23] Yvain Bruned and Foivos Katsetsiadis. Ramification of Volterratype rough paths. Electronic Journal of Probability, 28(??): 1–25, 2023. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/journals/electronic-journal-of-probability/
volume-28/issue-none/Ramification-of-Volterra-type-rough-
paths/10.1214/22-EJP890.full.

Bjornberg:2019:IPR

[BKLM19] Jakob E. Björnberg, Michał Kotowski, Benjamin Lees, and Piotr Miłoś. The interchange process with reversals on the complete graph. Electronic Journal of Probability, 24(??):108:1–108:43, ???? 2019. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1569981823.

Buraczewski:2021:SSS

[BKM21] Dariusz Buraczewski, Konrad Kolesko, and Matthias Meiners. Self-similar solutions to kinetic-type evolution equations: beyond the boundary case. Electronic Journal of Probability, 26(??):1–18, 2021. CODEN ???? ISSN

1083-6489.  URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-26/issue-none/
Self-similar-solutions-to-kinetic-type-evolution-equations-
-beyond/10.1214/20-EJP568.full.

Bartl:2019:SID

[BKN19] Daniel Bartl, Michael Kupper, and Ariel Neufeld. Stochastic integration and differential equations for typical paths. Electronic Journal of Probability, 24(??):97:1–97:21, ???? 2019. CODEN

???? ISSN 1083-6489. URL https://projecteuclid.org/
euclid.ejp/1568793790.

Baurdoux:2016:OPP

[BKO16] Erik J. Baurdoux, Andreas E. Kyprianou, and Curdin Ott. Optimal prediction for positive self-similar Markov processes. Electronic Journal of Probability, 21(??):48:1–48:24, ???? 2016. CO-DEN ???? ISSN 1083-6489. URL https://projecteuclid.

org/euclid.ejp/1469720443.

Berzin:2005:CFM

[BL05] Corinne Berzin and José León. Convergence in fractional models and applications. Electronic Journal of Probability, 10:10:326– 10:370, 2005. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/172.

Barbour:2012:CLA

[BL12a] Andrew Barbour and Malwina Luczak. Central limit approximations for Markov population processes with countably many types. Electronic Journal of Probability, 17:90:1–90:16, 2012. CO-DEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/1760.

Brightwell:2012:VHD

[BL12b] Graham Brightwell and Malwina Luczak. Vertices of high degree in the preferential attachment tree. Electronic Journal of Probability, 17:14:1–14:43, 2012. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/1803.

Brault:2019:NLS

[BL19] Antoine Brault and Antoine Lejay. The non-linear sewing lemma I: weak formulation. Electronic Journal of Probability, 24(??):

59:1–59:24, ???? 2019. CODEN ???? ISSN 1083-6489. URL

https://projecteuclid.org/euclid.ejp/1561082668.

Bordenave:2022:NST

[BL22] Charles Bordenave and Jaehun Lee. Noise sensitivity for the top eigenvector of a sparse random matrix. Electronic Journal of Probability, 27(??):1–50, 2022. CODEN

???? ISSN 1083-6489. URL https://projecteuclid.org/
journals/electronic-journal-of-probability/volume-27/
issue-none/Noise-sensitivity-for-the-top-eigenvector-
of-a-sparse-random/10.1214/22-EJP770.full.

Boucheron:2009:CSB

[BLM09] Stephane Boucheron, Gabor Lugosi, and Pascal Massart. On concentration of self-bounding functions. Electronic Journal of Probability, 14:64:1884–64:1899, 2009. CODEN ???? ISSN 1083-

6489. URL http://ejp.ejpecp.org/article/view/690.

Bassetti:2015:IES

[BLM15] Federico Bassetti, Lucia Ladelli, and Daniel Matthes. Infinite energy solutions to inelastic homogeneous Boltzmann equations. Electronic Journal of Probability, 20(??):89:1–89:34, ???? 2015. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/3531.

Barden:2013:CLT

[BLO13] Dennis Barden, Huiling Le, and Megan Owen. Central limit theorems for Fréchet means in the space of phylogenetic trees. Electronic Journal of Probability, 18:25:1–25:25, 2013. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
2201.

Behme:2021:LKE

[BLR21] Anita Behme, Alexander Lindner, and Jana Reker. On the law of killed exponential functionals. Electronic Journal of Probability, 26(18):1–35, February 2021. CODEN

???? ISSN 1083-6489. URL https://projecteuclid.org/
journals/electronic-journal-of-probability/volume-26/
issue-none/On-the-law-of-killed-exponential-functionals/
10.1214/21-EJP616.full.

Birkner:2018:CRD

[BLS18] Matthias Birkner, Huili Liu, and Anja Sturm. Coalescent results for diploid exchangeable population models. Electronic Journal of Probability, 23(??):49:1–49:44, ???? 2018. CODEN ????

ISSN 1083-6489. URL https://projecteuclid.org/euclid.
ejp/1527818427.

Benes:2019:TPI

[BLV19] Christian Beneš, Gregory F. Lawler, and Fredrik Viklund. Transition probabilities for infinite two-sided loop-erased random walks. Electronic Journal of Probability, 24(??):139:1–139:22, ???? 2019. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.

org/euclid.ejp/1575428689.

Bahlali:2005:GSM

[BM05] Seid Bahlali and Brahim Mezerdi. A general stochastic maximum principle for singular control problems. Electronic Journal of Probability, 10:30:988–30:1004, 2005. CODEN ???? ISSN 1083-

6489. URL http://ejp.ejpecp.org/article/view/271.

Bavouzet:2006:CGU

[BM06] Marie Pierre Bavouzet and Marouen Messaoud. Computation of Greeks using Malliavin’s calculus in jump type market models. Electronic Journal of Probability, 11:10:276–10:300, 2006. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/314.

Bessaih:2009:LDP

[BM09a] Hakima Bessaih and Annie Millet. Large deviation principle and inviscid shell models. Electronic Journal of Probability, 14: 89:2551–89:2579, 2009. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/719.

Blomker:2009:AES

[BM09b] Dirk Blömker and Wael Mohammed. Amplitude equation for SPDEs with quadratic non-linearities. Electronic Journal of Probability, 14:88:2527–88:2550, 2009. CODEN ???? ISSN 1083-

6489. URL http://ejp.ejpecp.org/article/view/716.

Backhausz:2011:LDD

[BM11] Agnes Backhausz and Tamas Mori. Local degree distribution in scale free random graphs. Electronic Journal of Probability, 16: 54:1465–54:1488, 2011. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/916.

Berard:2014:LPP

[BM14] Jean Bérard and Pascal Maillard. The limiting process of Nparticle branching random walk with polynomial tails. Electronic Journal of Probability, 19:22:1–22:17, 2014. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/3111.

Bourgade:2019:GFD

[BM19a] Paul Bourgade and Krishnan Mody. Gaussian fluctuations of the determinant of Wigner matrices. Electronic Journal of Probability, 24(??):96:1–96:28, ???? 2019. CODEN ???? ISSN 1083-6489.

URL https://projecteuclid.org/euclid.ejp/1568793789.

Boutaud:2019:RPS

[BM19b] Pierre Boutaud and Pascal Maillard. A revisited proof of the Seneta–Heyde norming for branching random walks under optimal assumptions. Electronic Journal of Probability, 24(??):99:1– 99:22, ???? 2019. CODEN ???? ISSN 1083-6489. URL https:/

/projecteuclid.org/euclid.ejp/1568793793.

Bhattacharjee:2020:CSI

[BM20] Chinmoy Bhattacharjee and Ilya Molchanov. Convergence to scale-invariant Poisson processes and applications in Dickman approximation. Electronic Journal of Probability, 25(??):79:1–79:20, ???? 2020. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1594432887.

Belloum:2021:ASR

[BM21] Mohamed Ali Belloum and Bastien Mallein. Anomalous spreading in reducible multitype branching Brownian motion. Electronic Journal of Probability, 26(18):1–39, February 2021. CODEN

???? ISSN 1083-6489. URL https://projecteuclid.org/
journals/electronic-journal-of-probability/volume-26/
issue-none/Anomalous-spreading-in-reducible-multitype-
branching-Brownian-motion/10.1214/21-EJP629.full.

Bencs:2022:AMM

[BM22a] Ferenc Bencs and András Mészáros. Atoms of the matching measure. Electronic Journal of Probability, 27(??):1–38, 2022. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/
journals/electronic-journal-of-probability/volume-27/
issue-none/Atoms-of-the-matching-measure/10.1214/22-EJP809.
full.

Bhattacharjee:2022:GAS

[BM22b] Chinmoy Bhattacharjee and Ilya Molchanov. Gaussian approximation for sums of region-stabilizing scores. Electronic Journal of Probability, 27(??):1–27, 2022. CODEN

???? ISSN 1083-6489. URL https://projecteuclid.org/
journals/electronic-journal-of-probability/volume-27/
issue-none/Gaussian-approximation-for-sums-of-region-
stabilizing-scores/10.1214/22-EJP832.full.

Bauer:2018:SSM

[BMBP18] Martin Bauer, Thilo Meyer-Brandis, and Frank Proske. Strong solutions of mean-field stochastic differential equations with irregular drift. Electronic Journal of Probability, 23(??):132:1–132:35, ???? 2018. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1545447917.

Balazs:2022:HLZ

[BMH22] Márton Balázs and Felix Maxey-Hawkins. Hydrodynamic limit of the zero range process on a randomly oriented graph. Electronic Journal of Probability, 27(??):1–29, 2022. CODEN

???? ISSN 1083-6489. URL https://projecteuclid.org/
journals/electronic-journal-of-probability/volume-27/
issue-none/Hydrodynamic-limit-of-the-zero-range-process-
on-a-randomly/10.1214/22-EJP753.full.

Beliaev:2020:SME

[BMM20] Dmitry Beliaev, Michael McAuley, and Stephen Muirhead. Smoothness and monotonicity of the excursion set density of planar Gaussian fields. Electronic Journal of Probability, 25(??): 1–37, 2020. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/journals/electronic-journal-of-probability/
volume-25/issue-none/Smoothness-and-monotonicity-of-the-
excursion-set-density-of-planar/10.1214/20-EJP470.full.

Bovier:2021:MDC

[BMP21] Anton Bovier, Saeda Marello, and Elena Pulvirenti. Metastability for the dilute Curie–Weiss model with Glauber dynamics. Electronic Journal of Probability, 26(18):1–38, February 2021. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/
journals/electronic-journal-of-probability/volume-26/
issue-none/Metastability-for-the-dilute-CurieWeiss-model-
with-Glauber-dynamics/10.1214/21-EJP610.full.

Bansaye:2013:ECS

[BMS13] Vincent Bansaye, Juan Carlos Pardo Millan, and Charline Smadi. On the extinction of continuous state branching processes with catastrophes. Electronic Journal of Probability, 18:106:1–106:31, 2013. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.

org/article/view/2774.

Belhaouari:2006:CRS

[BMSV06] Samir Belhaouari, Thomas Mountford, Rongfeng Sun, and Glauco Valle. Convergence results and sharp estimates for the voter model interfaces. Electronic Journal of Probability, 11: 30:768–30:801, 2006. CODEN ???? ISSN 1083-6489. URL http:

//ejp.ejpecp.org/article/view/349.

Bonaccorsi:2008:SFN

[BMZ08] Stefano Bonaccorsi, Carlo Marinelli, and Giacomo Ziglio. Stochastic FitzHugh–Nagumo equations on networks with impulsive noise. Electronic Journal of Probability, 13:49:1362–49:1379, 2008. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.

org/article/view/532.

Bahlali:2013:PMN

[BMZ13] Khaled Bahlali, Lucian Maticiuc, and Adrian Zalinescu. Penalization method for a nonlinear Neumann PDE via weak solutions of reflected SDEs. Electronic Journal of Probability, 18: 102:1–102:19, 2013. CODEN ???? ISSN 1083-6489. URL http:

//ejp.ejpecp.org/article/view/2467.

Barbour:2011:ADD

[BN11] Andrew Barbour and Bruno Nietlispach. Approximation by the Dickman distribution and quasi-logarithmic combinatorial structures. Electronic Journal of Probability, 16:29:880–29:902, 2011. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/881.

Beiglbock:2014:MID

[BN14] Mathias Beiglböck and Marcel Nutz. Martingale inequalities and deterministic counterparts. Electronic Journal of Probability, 19 (??):95:1–95:15, ???? 2014. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/3270.

Balazs:2015:DDB

[BN15] Marton Balazs and Attila Laszlo Nagy. Dependent double branching annihilating random walk. Electronic Journal of Probability, 20(??):84:1–84:32, ???? 2015. CODEN ???? ISSN 1083-

6489. URL http://ejp.ejpecp.org/article/view/4045.

Baldassarri:2021:MLG

[BN21] Simone Baldassarri and Francesca Romana Nardi. Metastability in a lattice gas with strong anisotropic interactions under Kawasaki dynamics. Electronic Journal of Probability, 26(18):1–66, February 2021. CODEN ???? ISSN

1083-6489.  URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-26/issue-none/
Metastability-in-a-lattice-gas-with-strong-anisotropic-
interactions-under/10.1214/21-EJP701.full.

Beghin:2009:FPP

[BO09] Luisa Beghin and Enzo Orsingher. Fractional Poisson processes and related planar random motions. Electronic Journal of Probability, 14:61:1790–61:1826, 2009. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/675.

Beghin:2010:PTP

[BO10] Luisa Beghin and Enzo Orsingher. Poisson-type processes governed by fractional and higher-order recursive differential equations. Electronic Journal of Probability, 15:22:684–22:709, 2010. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/762.

Basse-OConnor:2011:IS

[BO11] Andreas Basse-O’Connor. Integrability of seminorms. Electronic Journal of Probability, 16:7:216–7:229, 2011. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/853.

Borodin:2013:MDT

[BO13] Alexei Borodin and Grigori Olshanski. Markov dynamics on the Thoma cone: a model of time-dependent determinantal processes with infinitely many particles. Electronic Journal of Probability, 18:75:1–75:43, 2013. CODEN ???? ISSN 1083-6489. URL http:

//ejp.ejpecp.org/article/view/2729.

Baccelli:2019:SGU

[BO19] François Baccelli and Eliza O’Reilly. The stochastic geometry of unconstrained one-bit data compression. Electronic Journal of Probability, 24(??):138:1–138:27, ???? 2019. CODEN ????

ISSN 1083-6489. URL https://projecteuclid.org/euclid.
ejp/1575342533.

Bobkov:2007:LDI

[Bob07] Sergey Bobkov. Large deviations and isoperimetry over convex probability measures with heavy tails. Electronic Journal of Probability, 12:39:1072–39:1100, 2007. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/440.

Bobkov:2022:UBF

[Bob22] Sergey G. Bobkov. Upper bounds for Fisher information. Electronic Journal of Probability, 27(??):1–44, 2022. CODEN

???? ISSN 1083-6489. URL https://projecteuclid.org/
journals/electronic-journal-of-probability/volume-27/
issue-none/Upper-Bounds-for-Fisher-information/10.1214/
22-EJP834.full.

Basse-OConnor:2012:MPS

[BOGP12] Andreas Basse-O’Connor, Svend-Erik Graversen, and Jan Pedersen. Multiparameter processes with stationary increments: Spectral representation and integration. Electronic Journal of Probability, 17:74:1–74:21, 2012. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/2287.

Basse-OConnor:2019:LTF

[BOHP19] Andreas Basse-O’Connor, Claudio Heinrich, and Mark Podolskij. On limit theory for functionals of stationary increments Lévy driven moving averages. Electronic Journal of Probability, 24 (??):79:1–79:42, ???? 2019. CODEN ???? ISSN 1083-6489. URL

https://projecteuclid.org/euclid.ejp/1567648850.

Boissard:2011:SBC

[Boi11] Emmanuel Boissard. Simple bounds for the convergence of empirical and occupation measures in 1-Wasserstein distance. Electronic Journal of Probability, 16:83:2296–83:2333, 2011. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/
view/958.

Bonnefont:2022:ODT

[Bon22] Benjamin Bonnefont. The overlap distribution at two temperatures for the branching Brownian motion. Electronic Journal of Probability, 27(??):1–21, 2022. CODEN ???? ISSN

1083-6489.  URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-27/issue-none/
The-overlap-distribution-at-two-temperatures-for-the-
branching-Brownian/10.1214/22-EJP841.full.

Basse-OConnor:2021:PVF

[BOPP21] Andreas Basse-O’Connor, Vytautė Pilipauskaitė, and Mark Podolskij. Power variations for fractional type infinitely divisible random fields. Electronic Journal of Probability, 26(18):1–35, February 2021. CODEN ???? ISSN

1083-6489.  URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-26/issue-none/
Power-variations-for-fractional-type-infinitely-divisible-
random-fields/10.1214/21-EJP617.full.

Basse-OConnor:2020:BET

[BOPT20] Andreas Basse-O’Connor, Mark Podolskij, and Christoph Thäle. A Berry–Esseén theorem for partial sums of functionals of heavytailed moving averages. Electronic Journal of Probability, 25(??): 31:1–31:31, ???? 2020. CODEN ???? ISSN 1083-6489. URL

https://projecteuclid.org/euclid.ejp/1582858935.

Bordenave:2008:BAP

[Bor08] Charles Bordenave. On the birth-and-assassination process, with an application to scotching a rumor in a network. Electronic Journal of Probability, 13:66:2014–66:2030, 2008. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
573.

Borrello:2011:SOA

[Bor11] Davide Borrello. Stochastic order and attractiveness for particle systems with multiple births, deaths and jumps. Electronic Journal of Probability, 16:4:106–4:151, 2011. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/852.

Bordenave:2014:EPT

[Bor14] Charles Bordenave. Extinction probability and total progeny of predator-prey dynamics on infinite trees. Electronic Journal of Probability, 19:20:1–20:33, 2014. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/2361.

Borga:2022:PLS

[Bor22] Jacopo Borga. The permuton limit of strong-Baxter and semi-Baxter permutations is the skew Brownian permuton. Electronic Journal of Probability, 27(??):1–53, 2022. CODEN

???? ISSN 1083-6489. URL https://projecteuclid.org/
journals/electronic-journal-of-probability/volume-27/
issue-none/The-permuton-limit-of-strong-Baxter-and-semi-
Baxter-permutations/10.1214/22-EJP886.full.

Bertacco:2022:RSP

[BOS22] Federico Bertacco, Carlo Orrieri, and Luca Scarpa. Random separation property for stochastic Allen–Cahn-type equations. Electronic Journal of Probability, 27(??):1–32, 2022. CODEN

???? ISSN 1083-6489. URL https://projecteuclid.org/
journals/electronic-journal-of-probability/volume-27/
issue-none/Random-separation-property-for-stochastic-
Allen--Cahn-type-equations/10.1214/22-EJP830.full.

Boukhadra:2010:SSD

[Bou10] Omar Boukhadra. Standard spectral dimension for the polynomial lower tail random conductances model. Electronic Journal of Probability, 15:68:2069–68:2086, 2010. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/839.

Bouzar:2011:DSS

[Bou11] Nadjib Bouzar. Discrete semi-self-decomposability induced by semigroups. Electronic Journal of Probability, 16:39:1117– 39:1133, 2011. CODEN ???? ISSN 1083-6489. URL http://

ejp.ejpecp.org/article/view/890.

Bouchet:2013:SBR

[Bou13] Élodie Bouchet. Sub-ballistic random walk in Dirichlet environment. Electronic Journal of Probability, 18:58:1–58:25, 2013. CO-DEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/2109.

Bertoin:1999:TCD

[BP99] Jean Bertoin and Jim Pitman. Two coalescents derived from the ranges of stable subordinators. Electronic Journal of Probability, 5:7:1–7:17, 1999. CODEN ???? ISSN 1083-6489. URL http://

ejp.ejpecp.org/article/view/63.

Bass:2004:CSD

[BP04] Richard Bass and Edwin Perkins. Countable systems of degenerate stochastic differential equations with applications to super-Markov chains. Electronic Journal of Probability, 9:22:634– 22:673, 2004. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/222.

Brouwer:2006:CSD

[BP06] Rachel Brouwer and Juho Pennanen. The cluster size distribution for a forest-fire process on Z. Electronic Journal of Probability, 11:43:1133–43:1143, 2006. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/369.

Biskup:2007:FCR

[BP07] Marek Biskup and Timothy Prescott. Functional CLT for random walk among bounded random conductances. Electronic Journal of Probability, 12:49:1323–49:1348, 2007. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/456.

Bass:2008:DSD

[BP08] Richard Bass and Edwin Perkins. Degenerate stochastic differential equations arising from catalytic branching networks. Electronic Journal of Probability, 13:60:1808–60:1885, 2008. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/
view/568.

Bryc:2009:DRQ

[BP09] Włodek Bryc and Virgil Pierce. Duality of real and quaternionic random matrices. Electronic Journal of Probability, 14:17:452– 17:476, 2009. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/606.

Bass:2012:ULP

[BP12] Richard Bass and Edwin Perkins. On uniqueness in law for parabolic SPDEs and infinite-dimensional SDEs. Electronic Journal of Probability, 17:36:1–36:54, 2012. CODEN ???? ISSN 1083-

6489. URL http://ejp.ejpecp.org/article/view/2049.

Bassetti:2013:SCE

[BP13a] Federico Bassetti and Eleonora Perversi. Speed of convergence to equilibrium in Wasserstein metrics for Kac-like kinetic equations. Electronic Journal of Probability, 18:6:1–6:35, 2013. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
2054.

Berger:2013:DTR

[BP13b] Noam Berger and Yuval Peres. Detecting the trail of a random walker in a random scenery. Electronic Journal of Probability, 18: 87:1–87:18, 2013. CODEN ???? ISSN 1083-6489. URL http://

ejp.ejpecp.org/article/view/2367.

Baumdicker:2014:IMG

[BP14a] Franz Baumdicker and Peter Pfaffelhuber. The infinitely many genes model with horizontal gene transfer. Electronic Journal of Probability, 19(??):115:1–115:27, ???? 2014. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/2642.

Bourguin:2014:PIP

[BP14b] Solesne Bourguin and Giovanni Peccati. Portmanteau inequalities on the Poisson space: mixed regimes and multidimensional clustering. Electronic Journal of Probability, 19:66:1–66:42, 2014. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/2879.

Berger:2015:CCP

[BP15] Quentin Berger and Julien Poisat. On the critical curves of the pinning and copolymer models in correlated Gaussian environment. Electronic Journal of Probability, 20(??):71:1–71:35, ???? 2015. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.

org/article/view/3514.

Bachmann:2016:CBG

[BP16] Sascha Bachmann and Giovanni Peccati. Concentration bounds for geometric Poisson functionals: Logarithmic Sobolev inequalities revisited. Electronic Journal of Probability, 21(??):6:1–6:44, ???? 2016. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1454682887.

Barrera:2020:CPO

[BP20] Gerardo Barrera and Juan Carlos Pardo. Cut-off phenomenon for Ornstein–Uhlenbeck processes driven by Lévy processes. Electronic Journal of Probability, 25(??):15:1–15:33, ???? 2020. CO-DEN ???? ISSN 1083-6489. URL https://projecteuclid.

org/euclid.ejp/1580871681.

Benigni:2021:EDS

[BP21] Lucas Benigni and Sandrine Péché. Eigenvalue distribution of some nonlinear models of random matrices. Electronic Journal of Probability, 26(18):1–37, February 2021. CODEN

???? ISSN 1083-6489. URL https://projecteuclid.org/
journals/electronic-journal-of-probability/volume-26/
issue-none/Eigenvalue-distribution-of-some-nonlinear-
models-of-random-matrices/10.1214/21-EJP699.full.

Bras:2022:TVD

[BPP22] Pierre Bras, Gilles Pagès, and Fabien Panloup. Total variation distance between two diffusions in small time with unbounded drift: application to the Euler–Maruyama scheme. Electronic Journal of Probability, 27(??):1–19, 2022. CODEN

???? ISSN 1083-6489. URL https://projecteuclid.org/
journals/electronic-journal-of-probability/volume-27/
issue-none/Total-variation-distance-between-two-diffusions-
in-small-time-with/10.1214/22-EJP881.full.

Berti:2012:LTE

[BPR12] Patrizia Berti, Luca Pratelli, and Pietro Rigo. Limit theorems for empirical processes based on dependent data. Electronic Journal of Probability, 17:9:1–9:18, 2012. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/1765.

Berti:2015:TVF

[BPR15] Patrizia Berti, Luca Pratelli, and Pietro Rigo. Two versions of the fundamental theorem of asset pricing. Electronic Journal of Probability, 20(??):34:1–34:21, ???? 2015. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/3321.

Bouchard:2016:GDM

[BPT16] Bruno Bouchard, Dylan Possamaï, and Xiaolu Tan. A general Doob–Meyer–Mertens decomposition for g-supermartingale systems. Electronic Journal of Probability, 21(??):36:1–36:21, ???? 2016. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1462192627.

Beringer:2017:PCP

[BPT17] Dorottya Beringer, Gábor Pete, and Ádám Timár. On percolation critical probabilities and unimodular random graphs. Electronic Journal of Probability, 22(??):106:1–106:26, ???? 2017. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.

org/euclid.ejp/1514430042.

Bogdan:2018:YLS

[BPW18] Krzysztof Bogdan, Zbigniew Palmowski, and Longmin Wang. Yaglom limit for stable processes in cones. Electronic Journal of Probability, 23(??):11:1–11:19, ???? 2018. CODEN ????

ISSN 1083-6489. URL https://projecteuclid.org/euclid.
ejp/1518426059.

Bao:2012:TWL

[BPZ12] Zhigang Bao, Guangming Pan, and Wang Zhou. Tracy–Widom law for the extreme eigenvalues of sample correlation matrices. Electronic Journal of Probability, 17:88:1–88:32, 2012. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/
view/1962.

Bhattacharya:2022:PHT

[BPZ22] Ayan Bhattacharya, Zbigniew Palmowski, and Bert Zwart. Persistence of heavy-tailed sample averages: principle of infinitely many big jumps. Electronic Journal of Probability, 27(??):1– 25, 2022. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/journals/electronic-journal-of-probability/
volume-27/issue-none/Persistence-of-heavy-tailed-sample-
averages--principle-of-infinitely/10.1214/22-EJP774.full.

Balan:2019:EDS

[BQSS19] Raluca M. Balan, Lluís Quer-Sardanyons, and Jian Song. Existence of density for the stochastic wave equation with space-time homogeneous Gaussian noise. Electronic Journal of Probability, 24(??):106:1–106:43, ???? 2019. CODEN ???? ISSN 1083-6489.

URL https://projecteuclid.org/euclid.ejp/1569895475.

Bass:2005:ASI

[BR05] Richard Bass and Jay Rosen. An almost sure invariance principle for renormalized intersection local times. Electronic Journal of Probability, 10:4:124–4:164, 2005. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/236.

Barbour:2006:DSW

[BR06] Andrew Barbour and Gesine Reinert. Discrete small world networks. Electronic Journal of Probability, 11:47:1234–47:1283, 2006. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.

org/article/view/381.

Bass:2007:FPR

[BR07] Richard Bass and Jay Rosen. Frequent points for random walks in two dimensions. Electronic Journal of Probability, 12:1:1–1:46, 2007. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.

org/article/view/388.

Benaim:2011:SID

[BR11] Michel Benaïm and Olivier Raimond. Self-interacting diffusions IV: Rate of convergence. Electronic Journal of Probability, 16: 66:1815–66:1843, 2011. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/948.

Barbu:2012:LSS

[BR12a] Viorel Barbu and Michael Roeckner. Localization of solutions to stochastic porous media equations: finite speed of propagation. Electronic Journal of Probability, 17:10:1–10:11, 2012. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/
view/1768.

Belaribi:2012:UFP

[BR12b] Nadia Belaribi and Francesco Russo. Uniqueness for Fokker– Planck equations with measurable coefficients and applications to the fast diffusion equation. Electronic Journal of Probability, 17:84:1–84:28, 2012. CODEN ???? ISSN 1083-6489. URL http:

//ejp.ejpecp.org/article/view/2349.

Barbour:2013:AEC

[BR13] Andrew Barbour and Gesine Reinert. Approximating the epidemic curve. Electronic Journal of Probability, 18:54:1–54:30, 2013. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.

org/article/view/2557.

Bally:2016:AMS

[BR16] Vlad Bally and Clément Rey. Approximation of Markov semigroups in total variation distance. Electronic Journal of Probability, 21(??):12:1–12:44, ???? 2016. CODEN ???? ISSN 1083-6489.

URL https://projecteuclid.org/euclid.ejp/1455717196.

Barbour:2018:CLT

[BR18a] A. D. Barbour and Adrian Röllin. A central limit theorem for the gossip process. Electronic Journal of Probability, 23(??):123:1– 123:37, ???? 2018. CODEN ???? ISSN 1083-6489. URL https:

            //projecteuclid.org/euclid.ejp/1545102141.

Baur:2018:UIH

[BR18b] Erich Baur and Loïc Richier. Uniform infinite half-planar quadrangulations with skewness. Electronic Journal of Probability, 23 (??):54:1–54:43, ???? 2018. CODEN ???? ISSN 1083-6489. URL

            https://projecteuclid.org/euclid.ejp/1528358488.

Barraquand:2020:LDS

[BR20] Guillaume Barraquand and Mark Rychnovsky. Large deviations for sticky Brownian motions. Electronic Journal of Probability, 25(??):1–52, 2020. CODEN ???? ISSN

            1083-6489.  URL https://projecteuclid.org/journals/
            electronic-journal-of-probability/volume-25/issue-none/
            Large-deviations-for-sticky-Brownian-motions/10.1214/
            20-EJP515.full.

Bodo:2022:SIR

[BR22] Gergely Bodó and Markus Riedle. Stochastic integration with respect to canonical α_{s}tablecylindricalLévyprocesses.ElectronicJournalofProbability, 27

            (??) : 1−−23, 2022.CODEN????ISSN1083−6489.URLhttpsprojecteuclidorgjournalselectronicjournalofprobabilityvolume27issuenoneStochasticintegrationwithrespecttocanonicalceb1stablecylindricalLc3a9vy10121422EJP884full.

Bandyopadhyay:2010:ODL

[BRS10] Antar Bandyopadhyay, Rahul Roy, and Anish Sarkar. On the one dimensional “Learning from Neighbours” model. Electronic Journal of Probability, 15:51:1574–51:1593, 2010. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/809.

Barbour:2020:CMI

[BRW20] A. D. Barbour, Nathan Ross, and Yuting Wen. Central moment inequalities using Stein’s method. Electronic Journal of Probability, 25(??):1–21, 2020.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-25/issue-none/Central-moment-
inequalities-using-Steins-method/10.1214/20-EJP493.full.

Bryc:1996:CMR

[Bry96] Włodzimierz Bryc. Conditional moment representations for dependent random variables. Electronic Journal of Probability, 1:7:1–7:14, 1996. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/7.

Benjamini:2001:RDL

[BS01] Itai Benjamini and Oded Schramm. Recurrence of distributional limits of finite planar graphs. Electronic Journal of Probability, 6:23:1–23:13, 2001. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/96.

Belitsky:2002:DSS

[BS02] Vladimir Belitsky and Gunter Schütz. Diffusion and scattering of shocks in the partially asymmetric simple exclusion process. Electronic Journal of Probability, 7:11:1–11:21, 2002. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.

org/article/view/110.

Basdevant:2008:RGT

[BS08a] Anne-Laure Basdevant and Arvind Singh. Rate of growth of a transient cookie random walk. Electronic Journal of Probability, 13:26:811–26:851, 2008. CO-

DEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
498.

Bose:2008:ALM

[BS08b] Arup Bose and Arnab Sen. Another look at the moment method for large dimensional random matrices. Electronic Journal of Probability, 13:21:588–21:628, 2008. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/

view/501.

Basdevant:2009:RTM

[BS09] Anne-Laure Basdevant and Arvind Singh. Recurrence and transience of a multiexcited random walk on a regular tree. Electronic Journal of Probability, 14: 55:1628–55:1669, 2009. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/672.

Bjornberg:2014:RBP

[BS14] Jakob Björnberg and Sigurdur Stefánsson. Recurrence of bipartite planar maps. Electronic Journal of Probability, 19:31:1–31:40, 2014. CODEN ???? ISSN 1083-

6489. URL http://ejp.ejpecp.org/article/view/3102.

Bansaye:2015:SLG

[BS15] Vincent Bansaye and Florian Simatos. On the scaling limits of Galton–Watson processes in varying environments. Electronic Journal of Probability, 20(??): 75:1–75:36, ???? 2015. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/3812.

Butkovsky:2017:IMS

[BS17] Oleg Butkovsky and Michael Scheutzow. Invariant measures for stochastic functional differential equations. Electronic Journal of Probability, 22(??):

98:1–98:23, ???? 2017. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1510802252.

Bakhtin:2020:LDP

[BS20a] Yuri Bakhtin and Donghyun Seo. Localization of directed polymers in continuous space. Electronic Journal of Probability, 25(??):1–56, 2020. CODEN ????

ISSN 1083-6489. URL https://projecteuclid.org/journals/electronic-
journal-of-probability/volume-25/issue-none/Localization-of-directed-
polymers-in-continuous-space/10.1214/20-EJP530.full.

Behme:2020:EFM

[BS20b] Anita Behme and Apostolos Sideris. Exponential functionals of Markov additive processes. Electronic Journal of Probability, 25(??):37:1–37:25, ???? 2020.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.
ejp/1585274716.

Berger:2020:SLS

[BS20c] Quentin Berger and Michele Salvi. Scaling limit of sub-ballistic 1D random walk among biased conductances: a story of wells and walls. Electronic Journal of Probability, 25(??):30:1–30:43, ???? 2020. CODEN ???? ISSN 1083-6489. URL

https://projecteuclid.org/euclid.ejp/1582534895.

Broutin:2021:SSR

[BS21] Nicolas Broutin and Henning Sulzbach. Self-similar real trees defined as fixed points and their geometric properties. Electronic Journal of Probability, 26(18):1–50, February 2021. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-26/issue-none/Self-similar-real-trees-defined-
as-fixed-points-and-their/10.1214/21-EJP647.full.

Baldasso:2022:LSS

[BS22a] Rangel Baldasso and Alexandre Stauffer. Local survival of spread of infection among biased random walks. Electronic Journal of Probability, 27(??):1–28, 2022. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

journals/electronic-journal-of-probability/volume-27/issue-none/Local-
survival-of-spread-of-infection-among-biased-random-walks/10.1214/
22-EJP861.full.

Bates:2022:FEM

[BS22b] Erik Bates and Youngtak Sohn. Free energy in multi-species mixed p -spin spherical models. Electronic Journal of Probability, 27(??):1–75, 2022. CODEN ????

ISSN 1083-6489. URL https://projecteuclid.org/journals/electronic-
journal-of-probability/volume-27/issue-none/Free-energy-in-multi-
species-mixed-p-spin-spherical-models/10.1214/22-EJP780.full.

Busani:2022:NEB

[BS22c] Ofer Busani and Timo Seppäläinen. Non-existence of bi-infinite polymers. Electronic Journal of Probability, 27(??):1–40, 2022. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-27/issue-none/Non-existence-of-bi-infinite-
polymers/10.1214/21-EJP731.full.

Bhatnagar:2016:DCH

[BST16] Nayantara Bhatnagar, Allan Sly, and Prasad Tetali. Decay of correlations for the hardcore model on the d-regular random graph. Electronic Journal of Probability, 21(??):9:1–9:42, ???? 2016. CODEN ???? ISSN 1083-6489. URL

https://projecteuclid.org/euclid.ejp/1454682890.

Betken:2022:VAC

[BST22] Carina Betken, Matthias Schulte, and Christoph Thäle. Variance asymptotics and central limit theory for geometric functionals of Poisson cylinder processes. Electronic Journal of Probability, 27(??):1–47, 2022. CODEN ????

ISSN 1083-6489. URL https://projecteuclid.org/journals/electronic-
journal-of-probability/volume-27/issue-none/Variance-asymptotics-
and-central-limit-theory-for-geometric-functionals-of/10.1214/22-
EJP805.full.

Boudabsa:2020:FED

[BSV20] Lotfi Boudabsa, Thomas Simon, and Pierre Vallois. Fractional extreme distributions. Electronic Journal of Probability, 25(??):1–20, 2020. CODEN ????

ISSN 1083-6489. URL https://projecteuclid.org/journals/electronic-
journal-of-probability/volume-25/issue-none/Fractional-extreme-distributions/
10.1214/20-EJP520.full.

Bo:2010:STS

[BSW10] Lijun Bo, Kehua Shi, and Yongjin Wang. Support theorem for a stochastic Cahn–Hilliard equation. Electronic Journal of Probability, 15:17:484–17:525, 2010. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/

view/760.

Berzunza:2021:TDB

[BSW21] Gabriel Berzunza, Anja Sturm, and Anita Winter. Trait-dependent branching particle systems with competition and multiple offspring. Electronic Journal of Probability, 26(18):1–41, February 2021. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-26/issue-none/Trait-dependent-branching-particle-
systems-with-competition-and-multiple-offspring/10.1214/21-EJP707.
full.

Bao:2022:EIP

[BSY22] Jianhai Bao, Michael Scheutzow, and Chenggui Yuan. Existence of invariant probability measures for functional McKean–Vlasov SDEs. Electronic Journal of Probability, 27(??):1–14, 2022. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-27/issue-none/Existence-of-invariant-probability-
measures-for-functional-McKean--Vlasov-SDEs/10.1214/22-EJP773.full.

Bouchard:2005:HAO

[BT05] Bruno Bouchard and Emmanuel Teman. On the hedging of American options in discrete time with proportional transaction costs. Electronic Journal of Probability, 10:22:746–22:760, 2005. CODEN ???? ISSN 1083-6489. URL http://

ejp.ejpecp.org/article/view/266.

Berger:2010:CPR

[BT10] Quentin Berger and Fabio Toninelli. On the critical point of the random walk pinning model in dimension d=3. Electronic Journal of Probability, 15:21:654– 21:683, 2010. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/761.

Benjamini:2015:FPP

[BT15a] Itai Benjamini and Romain Tessera. First passage percolation on nilpotent Cayley graphs and beyond. Electronic Journal of Probability, 20(??):99:1–99:20, ???? 2015. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/3940.

Broman:2015:PCH

[BT15b] Erik Ivar Broman and Johan Tykesson. Poisson cylinders in hyperbolic space. Electronic Journal of Probability, 20(??):41:1–41:25, ???? 2015. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/3645.

Berger:2017:NPL

[BT17] Noam Berger and Ran J. Tessler. No percolation in low temperature spin glass. Electronic Journal of Probability, 22(??):88:1–88:19, ???? 2017. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.ejp/
1508292259.

Betz:2019:SLB

[BT19] Volker Betz and Lorenzo Taggi. Scaling limit of ballistic self-avoiding walk interacting with spatial random permutations. Electronic Journal of Probability, 24(??):74:1–74:37, ???? 2019. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1562119474.

Bartl:2020:FIF

[BT20] Daniel Bartl and Ludovic Tangpi. Functional inequalities for forward and backward diffusions. Electronic Journal of Probability, 25(??):1–22, 2020. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-25/issue-none/Functional-
inequalities-for-forward-and-backward-diffusions/10.1214/20-EJP495.
full.

Benjamini:2021:IEU

[BT21] Itai Benjamini and Ádám Timár. Invariant embeddings of unimodular random planar graphs. Electronic Journal of Probability, 26(18):1–18, February 2021.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-26/issue-none/Invariant-
embeddings-of-unimodular-random-planar-graphs/10.1214/21-EJP665.full.

Berger:2022:NDP

[BTW22] Quentin Berger, Niccolò Torri, and Ran Wei. Non-directed polymers in heavytail random environment in dimension. Electronic Journal of Probability, 27(??): 1–67, 2022. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.

org/journals/electronic-journal-of-probability/volume-27/issue-none/
Non-directed-polymers-in-heavy-tail-random-environment-in-dimension/
10.1214/22-EJP873.full.

Betz:2011:SRP

[BU11] Volker Betz and Daniel Ueltschi. Spatial random permutations and Poisson– Dirichlet law of cycle lengths. Electronic Journal of Probability, 16:41:1173– 41:1192, 2011. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/901.

Budzinski:2019:SCM

[Bud19] Thomas Budzinski. Supercritical causal maps: geodesics and simple random walk. Electronic Journal of Probability, 24(??):86:1–86:43, ???? 2019. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.ejp/
1568080866.

Burdzy:2009:DSF

[Bur09] Krzysztof Burdzy. Differentiability of stochastic flow of reflected Brownian motions. Electronic Journal of Probability, 14:75:2182–75:2240, 2009. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/700.

Burdzy:2015:STG

[Bur15] Krzysztof Burdzy. Stirring two grains of sand. Electronic Journal of Probability, 20(??):13:1–13:29, ???? 2015. CODEN ???? ISSN 1083-6489. URL http://

ejp.ejpecp.org/article/view/3845.

Busani:2016:AUC

[Bus16] Ofer Busani. Aging uncoupled continuous time random walk limits. Electronic Journal of Probability, 21(??):7:1–7:17, ???? 2016. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/euclid.ejp/1454682888.

Butez:2016:LDE

[But16] Raphaël Butez. Large deviations for the empirical measure of random polynomials: revisit of the Zeitouni–Zelditch theorem. Electronic Journal of Probability, 21(??):73:1–73:37, ???? 2016. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1481079628.

Barral:2004:MAC

[BV04] Julien Barral and Jacques Véhel. Multifractal analysis of a class of additive processes with correlated non-stationary increments. Electronic Journal of Probability, 9:16:508–16:543, 2004. CODEN ???? ISSN 1083-6489. URL http://

ejp.ejpecp.org/article/view/208.

Brzezniak:2012:SPC

[BV12] Zdzislaw Brzezniak and Mark Veraar. Is the stochastic parabolicity condition dependent on p and q ? Electronic Journal of Probability, 17:56:1–56:24, 2012.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
2186.

Bethuelsen:2016:ACW

[BV16] Stein Andreas Bethuelsen and Florian Völlering. Absolute continuity and weak uniform mixing of random walk in dynamic random environment. Electronic Journal of Probability, 21(??):71:1–71:32, ???? 2016. CODEN ???? ISSN

1083-6489. URL https://projecteuclid.org/euclid.ejp/1480688088.

Baroni:2015:FSC

[BvdHK15] Enrico Baroni, Remco van der Hofstad, and Julia Komjathy. Fixed speed competition on the configuration model with infinite variance degrees: unequal speeds.

Electronic Journal of Probability, 20(??):116:1–116:48, ???? 2015. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/3749.

Bhamidi:2010:SLC

[BvdHvL10] Shankar Bhamidi, Remco van der Hofstad, and Johan van Leeuwaarden. Scaling limits for critical inhomogeneous random graphs with finite third moments. Electronic Journal of Probability, 15:54:1682–54:1702, 2010. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/817.

Bell:2005:DSP

[BW05] Steven Bell and Ruth Williams. Dynamic scheduling of a parallel server system in heavy traffic with complete resource pooling: Asymptotic optimality of a threshold policy. Electronic Journal of Probability, 10:33:1044–33:1115, 2005.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
281.

Bryc:2012:BQH

[BW12] Włodek Bryc and Jacek Wesołowski. Bridges of quadratic harnesses. Electronic Journal of Probability, 17:105:1–105:25, 2012. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/1866.

Broutin:2017:RCT

[BW17] Nicolas Broutin and Minmin Wang. Reversing the cut tree of the Brownian continuum random tree. Electronic Journal of Probability, 22(??):80:1–80:23, ???? 2017. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.

org/euclid.ejp/1507255394.

Baudoin:2021:AWB

[BW21] Fabrice Baudoin and Jing Wang. Asymptotic windings of the block determinants of a unitary Brownian motion and related diffusions. Electronic Journal of Probability, 26(??):1–21, 2021. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-26/issue-none/Asymptotic-windings-of-the-block-
determinants-of-a-unitary-Brownian/10.1214/21-EJP600.full.

Bao:2015:HFS

[BWY15] Jianhai Bao, Feng-Yu Wang, and Chenggui Yuan. Hypercontractivity for functional stochastic partial differential equations. Electronic Journal of Probability, 20(??):93:1–93:15, ???? 2015. CODEN ???? ISSN 1083-6489. URL http://

ejp.ejpecp.org/article/view/4108.

Bai:2009:CLS

[BWZ09] Zhidong Bai, Xiaoying Wang, and Wang Zhou. CLT for linear spectral statistics of Wigner matrices. Electronic Journal of Probability, 14:83:2391–83:2417, 2009.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
705.

Barbour:2019:MAT

[BX19] A. D. Barbour and A. Xia. Multivariate approximation in total variation using local dependence. Electronic Journal of Probability, 24(??):27:1–27:35, ???? 2019. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

euclid.ejp/1553565778.

Baker:2009:BSM

[BY09] David Baker and Marc Yor. A Brownian sheet martingale with the same marginals as the arithmetic average of geometric Brownian motion. Electronic Journal of Probability, 14:52:1532–52:1540, 2009. CODEN ???? ISSN 1083-

6489. URL http://ejp.ejpecp.org/article/view/674.

Blaszczyszyn:2013:CPP

[BY13] Bartlomiej Blaszczyszyn and Dhandapani Yogeshwaran. Clustering and percolation of point processes. Electronic Journal of Probability, 18:72:1–72:20, 2013.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
2468.

Benjamini:2007:RGH

[BYY07] Itai Benjamini, Ariel Yadin, and Amir Yehudayoff. Random graphhomomorphisms and logarithmic degree. Electronic Journal of Probability, 12:32:926–32:950, 2007. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/427.

Baudoin:2012:TES

[BZ12] Fabrice Baudoin and Xuejing Zhang. Taylor expansion for the solution of a stochastic differential equation driven by fractional Brownian motions. Electronic Journal of Probability, 17:51:1–51:21, 2012. CODEN ???? ISSN 1083-

6489. URL http://ejp.ejpecp.org/article/view/2136.

Caraceni:2020:PUB

[Car20] Alessandra Caraceni. A polynomial upper bound for the mixing time of edge rotations on planar maps. Electronic Journal of Probability, 25(??):1–30, 2020.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-25/issue-none/A-polynomial-
upper-bound-for-the-mixing-time-of-edge/10.1214/20-EJP519.full.

Carrance:2021:CET

[Car21] Ariane Carrance. Convergence of Eulerian triangulations. Electronic Journal of Probability, 26(??):1–48, 2021. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-26/issue-none/Convergence-of-Eulerian-triangulations/
10.1214/21-EJP579.full.

Confortola:2008:QBR

[CB08] Fulvia Confortola and Philippe Briand. Quadratic BSDEs with random terminal time and elliptic PDEs in infinite dimension. Electronic Journal of Probability, 13:54:1529–54:1561, 2008. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/514.

Chassagneux:2009:RCL

[CB09] Jean-François Chassagneux and Bruno Bouchard. Representation of continuous linear forms on the set of ladlag processes and the hedging of American claims under proportional costs. Electronic Journal of Probability, 14:24:612–24:632, 2009. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/

view/625.

Champagnat:2016:MFS

[CB16] Nicolas Champagnat and Henry Benoit. Moments of the frequency spectrum of a splitting tree with neutral Poissonian mutations. Electronic Journal of Probability, 21(??):53:1–53:34, ???? 2016. CODEN ???? ISSN 1083-6489. URL

https://projecteuclid.org/euclid.ejp/1472830615.

Capinski:1998:MAS

[CC98] Marek Capiński and Nigel J. Cutland. Measure attractors for stochastic Navier– Stokes equations. Electronic Journal of Probability, 3:8:1–8:15, 1998. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/30.

Chen:2011:SSC

[CC11] Xinxing Chen and Dayue Chen. Some sufficient conditions for infinite collisions of simple random walks on a wedge comb. Electronic Journal of Probability, 16:49:1341–49:1355, 2011. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/907.

Caravenna:2013:IPR

[CC13] Francesco Caravenna and Loïc Chaumont. An invariance principle for random walk bridges conditioned to stay positive. Electronic Journal of Probability, 18: 60:1–60:32, 2013. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.

org/article/view/2362.

Caravenna:2022:GLS

[CC22a] Francesco Caravenna and Francesca Cottini. Gaussian limits for subcritical chaos. Electronic Journal of Probability, 27(??):1–35, 2022. CODEN ????

ISSN 1083-6489. URL https://projecteuclid.org/journals/electronic-
journal-of-probability/volume-27/issue-none/Gaussian-limits-for-subcritical-
chaos/10.1214/22-EJP798.full.

Collin:2022:REC

[CC22b] Orphée Collin and Francis Comets. Rate of escape of conditioned Brownian motion. Electronic Journal of Probability, 27(??):1–26, 2022. CODEN ????

ISSN 1083-6489. URL https://projecteuclid.org/journals/electronic-
journal-of-probability/volume-27/issue-none/Rate-of-escape-of-conditioned-
Brownian-motion/10.1214/21-EJP737.full.

Ceci:2015:LRM

[CCC15] Claudia Ceci, Alessandra Cretarola, and Katia Colaneri. Local riskminimization under restricted information on asset prices. Electronic Journal of Probability, 20(??):96:1–96:30, ???? 2015. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/3204.

Campi:2013:ECD

[CCD13] Luciano Campi, Umut Cetin, and Albina Danilova. Explicit construction of a dynamic Bessel bridge of dimension 3. Electronic Journal of Probability, 18: 30:1–30:25, 2013. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.

org/article/view/1907.

Cass:2021:LTB

[CCDO21] Thomas Cass, Dan Crisan, Paul Dobson, and Michela Ottobre. Long-time behaviour of degenerate diffusions: UFG-type SDEs and time-inhomogeneous hypoelliptic processes. Electronic Journal of Probability, 26(??):1–72, 2021. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-26/issue-none/Long-time-
behaviour-of-degenerate-diffusions--UFG-type-SDEs/10.1214/20-EJP577.
full.

Csaki:2009:SLT

[CCFR09] Endre Csáki, Miklós Csörgö, Antonia Feldes, and Pál Révész. Strong limit theorems for a simple random walk on the 2-dimensional comb. Electronic Journal of Probability, 14:82:2371–82:2390, 2009. CODEN ???? ISSN 1083-

6489. URL http://ejp.ejpecp.org/article/view/710.

Chiarini:2022:ETE

[CCGR22] Alberto Chiarini, Giovanni Conforti, Giacomo Greco, and Zhenjie Ren. Entropic turnpike estimates for the kinetic Schrödinger problem. Electronic Journal of Probability, 27(??):1–32, 2022. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-27/issue-none/Entropic-turnpike-estimates-for-
the-kinetic-Schr%c3%b6dinger-problem/10.1214/22-EJP850.full.

Can:2022:SDS

[CCK22] V. H. Can, D. A. Croydon, and T. Kumagai. Spectral dimension of simple random walk on a long-range percolation cluster. Electronic Journal of Probability, 27(??):1–37, 2022. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-27/issue-none/Spectral-dimension-of-simple-
random-walk-on-a-long-range/10.1214/22-EJP783.full.

Cohen:2020:MDO

[CCO20] Philip Cohen, Fabio Deelan Cunden, and Neil O’Connell. Moments of discrete orthogonal polynomial ensembles. Electronic Journal of Probability, 25 (??):72:1–72:19, ???? 2020. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1593568836.

Cancrini:2005:DLT

[CCR05] Nicoletta Cancrini, Filippo Cesi, and Cyril Roberto. Diffusive long-time behavior of Kawasaki dynamics. Electronic Journal of Probability, 10:7:216–7:249, 2005. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/

view/239.

Cordero:2022:TCA

[CCSWB22] Fernando Cordero, Adrián González Casanova, Jason Schweinsberg, and Maite Wilke-Berenguer. Λ-coalescents arising in a population with dormancy. Electronic Journal of Probability, 27(??):1–34, 2022. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-27/issue-none/%ce%9b-coalescents-arising-in-
a-population-with-dormancy/10.1214/22-EJP739.full.

Cabanal-Duvillard:2005:MRB

[CD05a] Thierry Cabanal-Duvillard. A matrix representation of the Bercovici–Pata bijection. Electronic Journal of Probability, 10:18:632–18:661, 2005. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/246.

Chaumont:2005:LPC

[CD05b] Loïc Chaumont and Ronald Doney. On Lévy processes conditioned to stay positive. Electronic Journal of Probability, 10:28:948–28:961, 2005. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/261. See

corrections [CD08a].

Cator:2006:BSC

[CD06] Eric Cator and Sergei Dobrynin. Behavior of a second class particle in Hammersley’s process. Electronic Journal of Probability, 11:26:670–26:685, 2006. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/340.

Chaumont:2008:CLP

[CD08a] Loïc Chaumont and Ronald Doney. Corrections to “On Lévy processes conditioned to stay positive”. Electronic Journal of Probability, 13:1:1–1:4, 2008.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/

466. See [CD05b].

Conus:2008:NLS

[CD08b] Daniel Conus and Robert Dalang. The non-linear stochastic wave equation in high dimensions. Electronic Journal of Probability, 13:22:629–22:670, 2008.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
500.

Casserini:2012:PPC

[CD12a] Matteo Casserini and Freddy Delbaen. Predictable projections of conformal stochastic integrals: an application to Hermite series and to Widder’s representation. Electronic Journal of Probability, 17:22:1–22:14, 2012. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/1883.

Collet:2012:RDD

[CD12b] Francesca Collet and Paolo Dai Pra. The role of disorder in the dynamics of critical fluctuations of mean field models. Electronic Journal of Probability, 17: 26:1–26:40, 2012. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.

org/article/view/1896.

Coupier:2014:CPQ

[CD14] David Coupier and David Dereudre. Continuum percolation for Quermass model. Electronic Journal of Probability, 19:35:1–35:19, 2014. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/2298.

Chiarini:2015:LCL

[CD15] Alberto Chiarini and Jean-Dominique Deuschel. Local central limit theorem for diffusions in a degenerate and unbounded random medium. Electronic Journal of Probability, 20(??):112:1–112:30, ???? 2015. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/4190.

Caravenna:2019:LLD

[CD19] Francesco Caravenna and Ron Doney. Local large deviations and the strong renewal theorem. Electronic Journal of Probability, 24(??):72:1–72:48, ???? 2019.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.
ejp/1561687605.

Cimasoni:2013:CTI

[CDC13] David Cimasoni and Hugo Duminil-Copin. The critical temperature for the Ising model on planar doubly periodic graphs. Electronic Journal of Probability, 18: 44:1–44:18, 2013. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.

org/article/view/2352.

Chelkak:2016:CPT

[CDCH16] Dmitry Chelkak, Hugo Duminil-Copin, and Clément Hongler. Crossing probabilities in topological rectangles for the critical planar FK-Ising model. Electronic Journal of Probability, 21(??):5:1–5:28, ???? 2016. CODEN ???? ISSN

1083-6489. URL https://projecteuclid.org/euclid.ejp/1454682886.

Chetwynd-Diggle:2018:SMS

[CDE18] Jonathan A. Chetwynd-Diggle and Alison M. Etheridge. SuperBrownian motion and the spatial Lambda–Fleming–Viot process. Electronic Journal of Probability, 23(??):71:1–71:36, ???? 2018. CODEN ???? ISSN 1083-6489. URL

https://projecteuclid.org/euclid.ejp/1532570599.

Cheridito:2006:DMR

[CDK06] Patrick Cheridito, Freddy Delbaen, and Michael Kupper. Dynamic monetary risk measures for bounded discrete-time processes. Electronic Journal of Probability, 11:3:57–3:106, 2006. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/302.

Champagnat:2012:DEN

[CDM12] Nicolas Champagnat, Persi Diaconis, and Laurent Miclo. On Dirichlet eigenvectors for neutral two-dimensional Markov chains. Electronic Journal of Probability, 17:63:1–63:41, 2012. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/1830.

Capitaine:2011:FCS

[CDMFF11] Mireille Capitaine, Catherine Donati-Martin, Delphine Féral, and Maxime Février. Free convolution with a semicircular distribution and eigenvalues of spiked deformations of Wigner matrices. Electronic Journal of Probability, 16: 64:1750–64:1792, 2011. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/934.

Carrapatoso:2013:CEC

[CE13] Kleber Carrapatoso and Amit Einav. Chaos and entropic chaos in Kac’s model without high moments. Electronic Journal of Probability, 18:78:1–78:38, 2013.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
2683.

Cuneo:2018:NES

[CEHRB18] Noé Cuneo, Jean-Pierre Eckmann, Martin Hairer, and Luc Rey-Bellet. Nonequilibrium steady states for networks of oscillators. Electronic Journal of Probability, 23(??):55:1–55:28, ???? 2018. CODEN ???? ISSN 1083-6489. URL

https://projecteuclid.org/euclid.ejp/1528358489.

Cerny:2011:TDR

[Čer11] Jiří Černý. On two-dimensional random walk among heavy-tailed conductances. Electronic Journal of Probability, 16:10:293–10:313, 2011. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/849.

Cipolloni:2021:FAC

[CES21] Giorgio Cipolloni, László Erdős, and Dominik Schröder. Fluctuation around the circular law for random matrices with real entries. Electronic Journal of Probability, 26(??):1–61, 2021. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-26/issue-none/Fluctuation-around-the-circular-
law-for-random-matrices-with-real/10.1214/21-EJP591.full.

Cipolloni:2022:OMR

[CES22] Giorgio Cipolloni, László Erdős, and Dominik Schröder. Optimal multi-resolvent local laws for Wigner matrices. Electronic Journal of Probability, 27(??):1–38, 2022. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

journals/electronic-journal-of-probability/volume-27/issue-none/Optimal-
multi-resolvent-local-laws-for-Wigner-matrices/10.1214/22-EJP838.
full.

Cuthbertson:2012:FPC

[CEY12] Charles Cuthbertson, Alison Etheridge, and Feng Yu. Fixation probability for competing selective sweeps. Electronic Journal of Probability, 17:31:1–31:36, 2012. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/

view/1954.

Chen:2011:MTA

[CF11] Che-Hao Chen and Michael Fuchs. On the moment-transfer approach for random variables satisfying a one-sided distributional recurrence. Electronic Journal of Probability, 16:30:903–30:928, 2011. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/885.

Caputo:2009:RTL

[CFG09] Pietro Caputo, Alessandra Faggionato, and Alexandre Gaudilliere. Recurrence and transience for long-range reversible random walks on a random point process. Electronic Journal of Probability, 14:90:2580–90:2616, 2009. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/721.

Carmesin:2017:LHS

[CFG17] Johannes Carmesin, Bruno Federici, and Agelos Georgakopoulos. A Liouville hyperbolic souvlaki. Electronic Journal of Probability, 22(??):36:1–36:19, ???? 2017. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

euclid.ejp/1493085635.

Crane:2015:CGD

[CFT15] Edward Crane, Nic Freeman, and Bálint Tóth. Cluster growth in the dynamical Erdős–Rényi process with forest fires. Electronic Journal of Probability, 20(??): 101:1–101:33, ???? 2015. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/4035.

Chen:2009:NFM

[CFW09] Bo Chen, Daniel Ford, and Matthias Winkel. A new family of Markov branching trees: the alpha-gamma model. Electronic Journal of Probability, 14:15:400– 15:430, 2009. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/616.

Criens:2018:ACS

[CG18] David Criens and Kathrin Glau. Absolute continuity of semimartingales. Electronic Journal of Probability, 23(??):125:1–125:28, ???? 2018. CODEN

???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.ejp/
1545188692.

Corwin:2020:KET

[CG20] Ivan Corwin and Promit Ghosal. KPZ equation tails for general initial data. Electronic Journal of Probability, 25(??):66:1–66:38, ???? 2020. CODEN

???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.ejp/
1592618468.

Chleboun:2022:PDA

[CGG22] Paul Chleboun, Simon Gabriel, and Stefan Grosskinsky. Poisson–Dirichlet asymptotics in condensing particle systems. Electronic Journal of Probability, 27(??):1–35, 2022. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/journals/electronic-journal-of-probability/volume-
27/issue-none/Poisson--Dirichlet-asymptotics-in-condensing-particle-
systems/10.1214/22-EJP882.full.

Cattiaux:2010:FIH

[CGGR10] Patrick Cattiaux, Nathael Gozlan, Arnaud Guillin, and Cyril Roberto. Functional inequalities for heavy tailed distributions and application to isoperimetry. Electronic Journal of Probability, 15:13:346–13:385, 2010. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/754.

Collet:2008:RPS

[CGL08] Pierre Collet, Antonio Galves, and Florencia Leonardi. Random perturbations of stochastic processes with unbounded variable length memory. Electronic Journal of Probability, 13:48:1345–48:1361, 2008. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/538.

Cortines:2018:DFS

[CGL18] Aser Cortines, Julian Gold, and Oren Louidor. Dynamical freezing in a spin glass system with logarithmic correlations. Electronic Journal of Probability, 23 (??):59:1–59:31, ???? 2018. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1528704077.

Comets:2013:LDC

[CGPV13] Francis Comets, Christophe Gallesco, Serguei Popov, and Marina Vachkovskaia. On large deviations for the cover time of two-dimensional torus. Electronic Journal of Probability, 18:96:1–96:18, 2013. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/2856.

Chazottes:2006:REW

[CGR06] Jean-René Chazottes, Cristian Giardina, and Frank Redig. Relative entropy and waiting times for continuous-time Markov processes. Electronic Journal of Probability, 11:40:1049–40:1068, 2006. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/374.

Chen:2020:SMI

[CGR20] Louis H. Y. Chen, Larry Goldstein, and Adrian Röllin. Stein’s method via induction. Electronic Journal of Probability, 25(??):1–49, 2020. CODEN ????

ISSN 1083-6489. URL https://projecteuclid.org/journals/electronic-
journal-of-probability/volume-25/issue-none/Steins-method-via-induction/
10.1214/20-EJP535.full.

Carinci:2021:CPS

[CGR21] Gioia Carinci, Cristian Giardinà, and Frank Redig. Consistent particle systems and duality. Electronic Journal of Probability, 26(18):1–31, February 2021.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-26/issue-none/Consistent-
particle-systems-and-duality/10.1214/21-EJP684.full.

Caravenna:2006:SAB

[CGZ06] Francesco Caravenna, Giambattista Giacomin, and Lorenzo Zambotti. Sharp asymptotic behavior for wetting models in (1+1)-dimension. Electronic Journal of Probability, 11:14:345–14:362, 2006. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/320.

Csaki:2005:IPV

[CH05] Endre Csaki and Yueyun Hu. On the increments of the principal value of Brownian local time. Electronic Journal of Probability, 10:27:925–27:947, 2005.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
269.

Croydon:2010:SAS

[CH10] David Croydon and Benjamin Hambly. Spectral asymptotics for stable trees. Electronic Journal of Probability, 15:57:1772–57:1801, 2010. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/819.

Chakrabarty:2011:ANH

[Cha11] Arijit Chakrabarty. Asymptotic normality of hill estimator for truncated data. Electronic Journal of Probability, 16:74:2039–74:2058, 2011. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/935.

Chang:2015:LCD

[Cha15] Yinshan Chang. Loop cluster on discrete circles. Electronic Journal of Probability, 20(??):2:1–2:32, ???? 2015. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/3176.

Chen:2013:ASS

[Che13a] Wei-Kuo Chen. The Aizenman–Sims–Starr scheme and Parisi formula for mixed p-spin spherical models. Electronic Journal of Probability, 18:94:1–94:14, 2013.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/

Chen:2013:CLT

[Che13b] Wei-Kuo Chen. Central limit theorems for cavity and local fields of the Sherrington–Kirkpatrick model. Electronic Journal of Probability, 18:2:1–2:25, 2013. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/

view/1763.

Chen:2014:TPR

[Che14] Jun Chen. Two particles’ repelling random walks on the complete graph. Electronic Journal of Probability, 19(??):113:1–113:17, ???? 2014. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/2669.

Che:2017:URM

[Che17a] Ziliang Che. Universality of random matrices with correlated entries. Electronic Journal of Probability, 22(??):30:1–30:38, ???? 2017. CODEN ???? ISSN

1083-6489. URL https://projecteuclid.org/euclid.ejp/1490320846.

Chevallier:2017:FMF

[Che17b] Julien Chevallier. Fluctuations for mean-field interacting age-dependent Hawkes processes. Electronic Journal of Probability, 22(??):42:1–42:49, ???? 2017. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.ejp/
1493777018.

Chen:2019:RWD

[Che19] Yu-Ting Chen. Rescaled Whittaker driven stochastic differential equations converge to the additive stochastic heat equation. Electronic Journal of Probability, 24(??):36:1–36:33, ???? 2019. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1554775415.

Chhaibi:2015:BGA

[Chh15] Reda Chhaibi. Beta-gamma algebra identities and Lie-theoretic exponential functionals of Brownian motion. Electronic Journal of Probability, 20(??):108:1– 108:20, ???? 2015. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.

org/article/view/3666.

Corwin:2023:ETW

[CHHM23] Ivan Corwin, Alan Hammond, Milind Hegde, and Konstantin Matetski. Exceptional times when the KPZ fixed point violates Johansson’s conjecture on maximizer uniqueness. Electronic Journal of Probability, 28(??):1–81, 2023.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-28/issue-none/Exceptional-
times-when-the-KPZ-fixed-point-violates-Johanssons-conjecture/10.
1214/22-EJP898.full.

Croydon:2012:CMT

[CHK12] David Croydon, Ben Hambly, and Takashi Kumagai. Convergence of mixing times for sequences of random walks on finite graphs. Electronic Journal of Probability, 17:3:1–3:32, 2012. CODEN ???? ISSN 1083-6489. URL http://

ejp.ejpecp.org/article/view/1705.

Croydon:2017:TCS

[CHK17] David Croydon, Ben Hambly, and Takashi Kumagai. Time-changes of stochastic processes associated with resistance forms. Electronic Journal of Probability, 22 (??):82:1–82:41, ???? 2017. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1507795233.

Collevecchio:2018:SOR

[CHK18] Andrea Collevecchio, Mark Holmes, and Daniel Kious. On the speed of oncereinforced biased random walk on trees. Electronic Journal of Probability, 23 (??):86:1–86:32, ???? 2018. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1536717745.

Collevecchio:2019:BRN

[CHK19] Andrea Collevecchio, Cong Bang Huynh, and Daniel Kious. The branching-ruin number as critical parameter of random processes on trees. Electronic Journal of Probability, 24(??):121:1–121:29, ???? 2019. CODEN ???? ISSN 1083-6489.

URL https://projecteuclid.org/euclid.ejp/1573030842.

Chen:2019:DBP

[CHKK19] Le Chen, Jingyu Huang, Davar Khoshnevisan, and Kunwoo Kim. Dense blowup for parabolic SPDEs. Electronic Journal of Probability, 24(??):118:1–118:33, ???? 2019. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.

org/euclid.ejp/1572314778.

Chen:2018:RGB

[CHL18] Dayue Chen, Yueyun Hu, and Shen Lin. Resistance growth of branching random networks. Electronic Journal of Probability, 23(??):52:1–52:17, ???? 2018.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.
ejp/1527818430.

Cook:2018:NHR

[CHNR18] Nicholas Cook, Walid Hachem, Jamal Najim, and David Renfrew. Non-Hermitian random matrices with a variance profile (I): deterministic equivalents and limiting ESDs. Electronic Journal of Probability, 23(??):110:1–110:61, ???? 2018. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

euclid.ejp/1540865373.

Cardona:2022:RDS

[CHNR22] Jorge Cardona, Martina Hofmanová, Torstein Nilssen, and Nimit Rana. Random dynamical system generated by the 3D Navier–Stokes equation with rough transport noise. Electronic Journal of Probability, 27(??):1–27, 2022. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-27/issue-none/Random-dynamical-
system-generated-by-the-3D-Navier--Stokes-equation/10.1214/22-EJP813.
full.

Chen:2017:SAP

[CHNT17] Xia Chen, Yaozhong Hu, David Nualart, and Samy Tindel. Spatial asymptotics for the parabolic Anderson model driven by a Gaussian rough noise. Electronic Journal of Probability, 22(??):65:1–65:38, ???? 2017. CODEN ???? ISSN

1083-6489. URL https://projecteuclid.org/euclid.ejp/1503367245.

Chorro:2005:CDL

[Cho05] Christophe Chorro. Convergence in Dirichlet law of certain stochastic integrals. Electronic Journal of Probability, 10:31:1005–31:1025, 2005. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/272.

Christoph:2019:DPH

[Chr19] Hofer-Temmel Christoph. Disagreement percolation for the hard-sphere model. Electronic Journal of Probability, 24(??):91:1–91:22, ???? 2019. CODEN

???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.ejp/
1568080871.

Chen:2018:TAF

[CHSS18] Xia Chen, Yaozhong Hu, Jian Song, and Xiaoming Song. Temporal asymptotics for fractional parabolic Anderson model. Electronic Journal of Probability, 23 (??):14:1–14:39, ???? 2018. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1519182022.

Collevecchio:2022:LTE

[CHSW22] Andrea Collevecchio, Kais Hamza, Meng Shi, and Ruth J. Williams. Limit theorems and ergodicity for general bootstrap random walks. Electronic Journal of Probability, 27(??):1–22, 2022. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-27/issue-none/Limit-theorems-and-ergodicity-
for-general-bootstrap-random-walks/10.1214/22-EJP818.full.

Collevecchio:2019:DWR

[CHT19] Andrea Collevecchio, Kais Hamza, and Laurent Tournier. A deterministic walk on the randomly oriented Manhattan lattice. Electronic Journal of Probability, 24(??):137:1–137:20, ???? 2019. CODEN ???? ISSN 1083-6489. URL https:/

/projecteuclid.org/euclid.ejp/1575342532.

Chen:2020:HRG

[CHZ20] Zhen-Qing Chen, Zimo Hao, and Xicheng Zhang. Hölder regularity and gradient estimates for SDEs driven by cylindrical. Electronic Journal of Probability, 25(??):1–23, 2020. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-25/issue-none/H%c3%b6lder-regularity-and-gradient-
estimates-for-SDEs-driven-by-cylindrical/10.1214/20-EJP542.full.

Cai:2022:NAC

[CHZ22] T. Tony Cai, Rungang Han, and Anru R. Zhang. On the non-asymptotic concentration of heteroskedastic Wishart-type matrix. Electronic Journal of Probability, 27(??):1–40, 2022. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-27/issue-none/On-the-non-asymptotic-concentration-
of-heteroskedastic-Wishart-type-matrix/10.1214/22-EJP758.full.

Cipriani:2013:HPM

[Cip13] Alessandra Cipriani. High points for the membrane model in the critical dimension. Electronic Journal of Probability, 18:86:1–86:17, 2013. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/2750.

Caputo:2019:CBP

[CIW19] Pietro Caputo, Dmitry Ioffe, and Vitali Wachtel. Confinement of Brownian polymers under geometric area tilts. Electronic Journal of Probability, 24(??): 37:1–37:21, ???? 2019. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1554775416.

Conus:2012:CLB

[CJK12] Daniel Conus, Mathew Joseph, and Davar Khoshnevisan. Correlation-length bounds, and estimates for intermittent islands in parabolic SPDEs. Electronic Journal of Probability, 17:102:1–102:15, 2012. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/2429.

Crisan:2018:PRS

[CJK18] Dan Crisan, Christopher Janjigian, and Thomas G. Kurtz. Particle representations for stochastic partial differential equations with boundary conditions. Electronic Journal of Probability, 23(??):65:1–65:29, ???? 2018. CODEN ???? ISSN

1083-6489. URL https://projecteuclid.org/euclid.ejp/1532570593.

Croydon:2008:RWG

[CK08] David Croydon and Takashi Kumagai. Random walks on Galton–Watson trees with infinite variance offspring distribution conditioned to survive. Electronic Journal of Probability, 13:51:1419–51:1441, 2008. CODEN ???? ISSN 1083-

6489. URL http://ejp.ejpecp.org/article/view/536.

Curien:2014:RSL

[CK14] Nicolas Curien and Igor Kortchemski. Random stable looptrees. Electronic Journal of Probability, 19(??):108:1–108:35, ???? 2014. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/2732.

Costantini:2015:VMG

[CK15] Cristina Costantini and Thomas Gordon Kurtz. Viscosity methods giving uniqueness for martingale problems. Electronic Journal of Probability, 20(??): 67:1–67:27, ???? 2015. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/3624.

Collet:2018:PSM

[CK18a] Francesca Collet and Richard C. Kraaij. Path-space moderate deviation principles for the random field Curie–Weiss model. Electronic Journal of Probability, 23(??):21:1–21:45, ???? 2018. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1519722150.

Costantini:2018:EUR

[CK18b] Cristina Costantini and Thomas G. Kurtz. Existence and uniqueness of reflecting diffusions in cusps. Electronic Journal of Probability, 23(??):84:1–84:21, ???? 2018. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.

org/euclid.ejp/1536717743.

Costantini:2019:MSC

[CK19] Cristina Costantini and Thomas G. Kurtz. Markov selection for constrained martingale problems. Electronic Journal of Probability, 24(??):135:1–135:31, ???? 2019. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.

org/euclid.ejp/1573700462.

Chen:2020:SCS

[CK20] Le Chen and Kunwoo Kim. Stochastic comparisons for stochastic heat equation. Electronic Journal of Probability, 25(??):1–38, 2020. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-25/issue-none/Stochastic-comparisons-for-stochastic-
heat-equation/10.1214/20-EJP541.full.

Chong:2022:ESH

[CK22] Carsten Chong and Péter Kevei. Extremes of the stochastic heat equation with additive Lévy noise. Electronic Journal of Probability, 27(??):1–21, 2022.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-27/issue-none/Extremes-of-
the-stochastic-heat-equation-with-additive-L%c3%a9vy-noise/10.1214/
22-EJP855.full.

Conchon-Kerjan:2023:AGG

[CK23] Guillaume Conchon-Kerjan. Anatomy of a Gaussian giant: supercritical level-sets of the free field on regular graphs. Electronic Journal of Probability, 28(??):1–60, 2023. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-28/issue-none/Anatomy-of-a-Gaussian-giant--
supercritical-level-sets-of/10.1214/23-EJP920.full.

Cioica:2013:RBS

[CKLL13] Petru Cioica, Kyeong-Hun Kim, Kijung Lee, and Felix Lindner. On the L_{q}(L_{p})- regularity and Besov smoothness of stochastic parabolic equations on bounded Lipschitz domains. Electronic Journal of Probability, 18:82:1–82:41, 2013. CO-

DEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
2478.

Cheridito:2003:FOU

[CKM03] Patrick Cheridito, Hideyuki Kawaguchi, and Makoto Maejima. Fractional Ornstein–Uhlenbeck processes. Electronic Journal of Probability, 8:3:1–3:14, 2003. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/

view/125.

Chauvin:2005:MPB

[CKMR05] B. Chauvin, T. Klein, J.-F. Marckert, and A. Rouault. Martingales and profile of binary search trees. Electronic Journal of Probability, 10:12:420–12:435, 2005.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
257.

Chen:2021:SES

[CKNP21] Le Chen, Davar Khoshnevisan, David Nualart, and Fei Pu. Spatial ergodicity for SPDEs via Poincaré-type inequalities. Electronic Journal of Probability, 26(18):1–37, February 2021. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-26/issue-none/Spatial-ergodicity-for-SPDEs-
via-Poincar%c3%a9-type-inequalities/10.1214/21-EJP690.full.

Chen:2012:GHK

[CKS12] Zhen-Qing Chen, Panki Kim, and Renming Song. Global heat kernel estimates for Δ + Δ^{α/2} in half-space-like domains. Electronic Journal of Probability, 17: 32:1–32:32, 2012. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.

org/article/view/1751.

Coquille:2021:SIB

[CKS21] Loren Coquille, Anna Kraut, and Charline Smadi. Stochastic individual-based models with power law mutation rate on a general finite trait space. Electronic Journal of Probability, 26(18):1–37, February 2021. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-26/issue-none/Stochastic-individual-based-models-
with-power-law-mutation-rate-on/10.1214/21-EJP693.full.

Crisan:2009:NFS

[CKX09] Dan Crisan, Michael Kouritzin, and Jie Xiong. Nonlinear filtering with signal dependent observation noise. Electronic Journal of Probability, 14:63:1863– 63:1883, 2009. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/687.

Cranston:1998:GEU

[CL98] Michael Cranston and Yves Le Jan. Geometric evolution under isotropic stochastic flow. Electronic Journal of Probability, 3:4:1–4:36, 1998. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/26.

Coutin:2005:SMR

[CL05] Laure Coutin and Antoine Lejay. Semi-martingales and rough paths theory. Electronic Journal of Probability, 10:23:761–23:785, 2005. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/162.

Crane:2013:CRM

[CL13] Harry Crane and Steven Lalley. Convergence rates of Markov chains on spaces of partitions. Electronic Journal of Probability, 18:61:1–61:23, 2013. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/2389.

Chen:2014:SCC

[CL14] Xin Chen and Xue-Mei Li. Strong completeness for a class of stochastic differential equations with irregular coefficients. Electronic Journal of Probability, 19 (??):91:1–91:34, ???? 2014. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/3293.

Che:2019:ULS

[CL19] Ziliang Che and Patrick Lopatto. Universality of the least singular value for sparse random matrices. Electronic Journal of Probability, 24(??):9:1–9:53, ???? 2019. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

euclid.ejp/1550221265.

Che:2021:ULS

[CL21a] Ziliang Che and Patrick Lopatto. Universality of the least singular value for the sum of random matrices. Electronic Journal of Probability, 26(18):1–38, February 2021. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

journals/electronic-journal-of-probability/volume-26/issue-none/Universality-
of-the-least-singular-value-for-the-sum-of/10.1214/21-EJP603.full.

Chen:2021:UAM

[CL21b] Wei-Kuo Chen and Wai-Kit Lam. Universality of approximate message passing algorithms. Electronic Journal of Probability, 26(??):1–44, 2021. CODEN ????

ISSN 1083-6489. URL https://projecteuclid.org/journals/electronic-
journal-of-probability/volume-26/issue-none/Universality-of-approximate-
message-passing-algorithms/10.1214/21-EJP604.full.

Chen:2005:LDL

[CLR05] Xia Chen, Wenbo Li, and Jay Rosen. Large deviations for local times of stable processes and stable random walks in 1 dimension. Electronic Journal of Probability, 10:16:577–16:608, 2005. CODEN ???? ISSN 1083-6489. URL http:

//ejp.ejpecp.org/article/view/260.

Cuchiero:2019:PMV

[CLSF19] Christa Cuchiero, Martin Larsson, and Sara Svaluto-Ferro. Probability measurevalued polynomial diffusions. Electronic Journal of Probability, 24(??):30:1– 30:32, ???? 2019. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1553565781.

Collet:2021:CRM

[CLT21] Francesca Collet, Fabrizio Leisen, and Steen Thorbjørnsen. Completely random measures and Lévy bases in free probability. Electronic Journal of Probability, 26(18):1–41, February 2021. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-26/issue-none/Completely-random-measures-and-
L%c3%a9vy-bases-in-free-probability/10.1214/21-EJP620.full.

Chaumont:2006:LEP

[CM06] Loic Chaumont and Juan Carlos Pardo Millan. The lower envelope of positive self-similar Markov processes. Electronic Journal of Probability, 11:49:1321– 49:1341, 2006. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/382.

Chen:2017:LBB

[CM17] Xinxin Chen and Grégory Miermont. Long Brownian bridges in hyperbolic spaces converge to Brownian trees. Electronic Journal of Probability, 22(??): 58:1–58:15, ???? 2017. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1500516020.

Chaumont:2020:FTS

[CM20] Loïc Chaumont and Marine Marolleau. Fluctuation theory for spectrally positive additive Lévy fields. Electronic Journal of Probability, 25(??):1–26, 2020.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-25/issue-none/Fluctuation-
theory-for-spectrally-positive-additive-L%c3%a9vy-fields/10.1214/
20-EJP547.full.

Couzinie:2022:FEP

[CM22] Yannick Couzinié and Fabio Martinelli. On a front evolution problem for the multidimensional East model. Electronic Journal of Probability, 27(??):1–30, 2022. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

journals/electronic-journal-of-probability/volume-27/issue-none/On-
a-front-evolution-problem-for-the-multidimensional-East-model/10.
1214/22-EJP870.full.

Chitour:2023:GBD

[CMMS23] Yacine Chitour, Guilherme Mazanti, Pierre Monmarché, and Mario Sigalotti. On the gap between deterministic and probabilistic Lyapunov exponents for continuous-time linear systems. Electronic Journal of Probability, 28(??):1–39, 2023. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

journals/electronic-journal-of-probability/volume-28/issue-none/On-
the-gap-between-deterministic-and-probabilistic-Lyapunov-exponents-
for/10.1214/23-EJP932.full.

Chaleyat-Maurel:1997:PPD

[CMN97] Mireille Chaleyat-Maurel and David Nualart. Points of positive density for smooth functionals. Electronic Journal of Probability, 3:1:1–1:8, 1997. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/23.

Chaleyat-Maurel:1998:PPD

[CMN98] Mireille Chaleyat-Maurel and David Nualart. Points of positive density for smooth functionals. Electronic Journal of Probability, 3:1:1–1:8, 1998. CODEN

???? ISSN 1083-6489. URL http://www.math.washington.edu/~ejpecp/
EjpVol3/paper1.abs.html.

Cox:2010:CTD

[CMP10] J. Theodore Cox, Mathieu Merle, and Edwin Perkins. Coexistence in a two-dimensional Lotka–Volterra model. Electronic Journal of Probability, 15: 38:1190–38:1266, 2010. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/795.

Caputo:2016:DLT

[CMSS16] Pietro Caputo, Fabio Martinelli, Alistair Sinclair, and Alexandre Stauffer. Dynamics of lattice triangulations on thin rectangles. Electronic Journal of Probability, 21(??):29:1–29:22, ???? 2016. CODEN ???? ISSN 1083-6489. URL

https://projecteuclid.org/euclid.ejp/1460652930.

Caputo:2008:AEP

[CMT08] Pietro Caputo, Fabio Martinelli, and Fabio Toninelli. On the approach to equilibrium for a polymer with adsorption and repulsion. Electronic Journal of Probability, 13:10:213–10:258, 2008. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/486.

Caputo:2015:MLP

[CMT15] Pietro Caputo, Fabio Martinelli, and Fabio Lucio Toninelli. Multi-level pinning problems for random walks and self-avoiding lattice paths. Electronic Journal of Probability, 20(??):8:1–8:29, ???? 2015. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/3849.

Coupier:2019:DCR

[CMT19] David Coupier, Jean-François Marckert, and Viet Chi Tran. Directed, cylindric and radial Brownian webs. Electronic Journal of Probability, 24(??):20:1–20:48, ???? 2019. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.

org/euclid.ejp/1553133829.

Conlon:2000:HEE

[CN00] Joseph Conlon and Ali Naddaf. On homogenization of elliptic equations with random coefficients. Electronic Journal of Probability, 5:9:1–9:58, 2000. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/65.

Corwin:2017:IDD

[CN17] Ivan Corwin and Mihai Nica. Intermediate disorder directed polymers and the multi-layer extension of the stochastic heat equation. Electronic Journal of Probability, 22(??):13:1–13:49, ???? 2017. CODEN ???? ISSN 1083-6489. URL

https://projecteuclid.org/euclid.ejp/1486090893.

Chaumont:2018:CSD

[CN18] Hans Chaumont and Christian Noack. Characterizing stationary 1 + 1 dimensional lattice polymer models. Electronic Journal of Probability, 23(??): 38:1–38:19, ???? 2018. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1525852815.

Can:2019:FPT

[CN19] Van Hao Can and Shuta Nakajima. First passage time of the frog model has a sublinear variance. Electronic Journal of Probability, 24(??):76:1–76:27, ???? 2019. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

euclid.ejp/1562292237.

Cox:2008:CMW

[CO08] Alexander Cox and Jan Obloj. Classes of measures which can be embedded in the simple symmetric random walk. Electronic Journal of Probability, 13: 42:1203–42:1228, 2008. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/516.

Chevyrev:2018:SDT

[CO18] Ilya Chevyrev and Marcel Ogrodnik. A support and density theorem for Markovian rough paths. Electronic Journal of Probability, 23(??):56:1–56:16, ???? 2018. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

euclid.ejp/1528704074.

Cichomski:2021:MDA

[CO21] Stanisław Cichomski and Adam Osȩkowski. The maximal difference among expert’s opinions. Electronic Journal of Probability, 26(18):1–17, February 2021.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-26/issue-none/The-maximal-
difference-among-experts-opinions/10.1214/21-EJP675.full.

Campbell:2022:SHT

[CO22] Andrew Campbell and Sean O’Rourke. Spectrum of heavy-tailed elliptic random matrices. Electronic Journal of Probability, 27(??):1–56, 2022. CODEN ????

ISSN 1083-6489. URL https://projecteuclid.org/journals/electronic-
journal-of-probability/volume-27/issue-none/Spectrum-of-heavy-tailed-
elliptic-random-matrices/10.1214/22-EJP849.full.

Cohen:2012:QSA

[Coh12] Samuel Cohen. Quasi-sure analysis, aggregation and dual representations of sublinear expectations in general spaces. Electronic Journal of Probability, 17:

62:1–62:15, 2012. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.

org/article/view/2224.

Collevecchio:2009:LTV

[Col09] Andrea Collevecchio. Limit theorems for vertex-reinforced jump processes on regular trees. Electronic Journal of Probability, 14:66:1936–66:1962, 2009. CO-

DEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
693.

Cassandro:2005:ODR

[COPV05] Marzio Cassandro, Enza Orlandi, Pierre Picco, and Maria Eulalia Vares. Onedimensional random field Kac’s model: Localization of the phases. Electronic Journal of Probability, 10:24:786–24:864, 2005. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/263.

Corso:2021:LDR

[Cor21] Emilio Corso. Large deviations for random walks on free products of finitely generated groups. Electronic Journal of Probability, 26(18):1–22, February 2021.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-26/issue-none/Large-deviations-
for-random-walks-on-free-products-of-finitely/10.1214/21-EJP695.full.

Coste:2023:SMC

[Cos23] Simon Coste. Sparse matrices: convergence of the characteristic polynomial seen from infinity. Electronic Journal of Probability, 28(??):1–40, 2023. CODEN ????

ISSN 1083-6489. URL https://projecteuclid.org/journals/electronic-
journal-of-probability/volume-28/issue-none/Sparse-matrices--convergence-
of-the-characteristic-polynomial-seen-from/10.1214/22-EJP875.full.

Coupier:2018:SNS

[Cou18] David Coupier. Sublinearity of the number of semi-infinite branches for geometric random trees. Electronic Journal of Probability, 23(??):37:1–37:33, ???? 2018. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

euclid.ejp/1525852814.

Couronne:2022:EPS

[Cou22] Olivier Couronné. Entanglement percolation and spheres in. Electronic Journal of Probability, 27(??):1–17, 2022. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-27/issue-none/Entanglement-percolation-and-
spheres-in-Zd/10.1214/22-EJP816.full.

Camarri:1999:LDR

[CP99] Michael Camarri and Jim Pitman. Limit distributions and random trees derived from the birthday problem with unequal probabilities. Electronic Journal of Probability, 5:2:1–2:18, 1999. CODEN ???? ISSN 1083-6489. URL http://

ejp.ejpecp.org/article/view/58.

Caravenna:2009:DPM

[CP09] Francesco Caravenna and Nicolas Pétrélis. Depinning of a polymer in a multiinterface medium. Electronic Journal of Probability, 14:70:2038–70:2067, 2009.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
698.

Cerny:2012:IDI

[ČP12] Jiří Černý and Serguei Popov. On the internal distance in the interlacement set. Electronic Journal of Probability, 17:29:1–29:25, 2012. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/1936.

Carmona:2016:IPD

[CP16] Philippe Carmona and Nicolas Pétrélis. Interacting partially directed self avoiding walk: scaling limits. Electronic Journal of Probability, 21(??):49:1–49:52, ???? 2016. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.

org/euclid.ejp/1469720444.

Connor:2019:MTE

[CP19] Stephen B. Connor and Richard J. Pymar. Mixing times for exclusion processes on hypergraphs. Electronic Journal of Probability, 24(??):73:1–73:48, ???? 2019.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.
ejp/1561687606.

Cox:2020:RSL

[CP20] J. Theodore Cox and Edwin A. Perkins. Rescaling the spatial Lambda–Fleming– Viot process and convergence to super-Brownian motion. Electronic Journal of Probability, 25(??):57:1–57:56, ???? 2020. CODEN ???? ISSN 1083-6489. URL

https://projecteuclid.org/euclid.ejp/1588644040.

Cancrini:2021:PCG

[CP21] Nicoletta Cancrini and Gustavo Posta. Propagation of chaos for a general balls into bins dynamics. Electronic Journal of Probability, 26(??):1–20, 2021.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-26/issue-none/Propagation-
of-chaos-for-a-general-balls-into-bins-dynamics/10.1214/21-EJP590.
full.

Coulson:2023:LCS

[CP23] Matthew Coulson and Guillem Perarnau. Largest component of subcritical random graphs with given degree sequence. Electronic Journal of Probability, 28(??):1–28, 2023. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-28/issue-none/Largest-component-of-subcritical-
random-graphs-with-given-degree-sequence/10.1214/23-EJP921.full.

Criens:2023:MPM

[CPS23] David Criens, Peter Pfaffelhuber, and Thorsten Schmidt. The martingale problem method revisited. Electronic Journal of Probability, 28(??):1–46, 2023.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-28/issue-none/The-martingale-
problem-method-revisited/10.1214/23-EJP902.full.

Cai:2022:CUP

[CPZ22] Zhenhao Cai, Eviatar B. Procaccia, and Yuan Zhang. Continuity and uniqueness of percolation critical parameters in finitary random interlacements. Electronic Journal of Probability, 27(??):1–46, 2022. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-27/issue-none/Continuity-and-uniqueness-of-
percolation-critical-parameters-in-finitary-random/10.1214/22-EJP824.
full.

Champagnat:2008:LTC

[CR08] Nicolas Champagnat and Sylvie Roelly. Limit theorems for conditioned multitype Dawson–Watanabe processes and Feller diffusions. Electronic Journal of Probability, 13:25:777–25:810, 2008. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/504.

Caballero:2009:ABI

[CR09] María Caballero and Víctor Rivero. On the asymptotic behaviour of increasing self-similar Markov processes. Electronic Journal of Probability, 14:31:865– 31:894, 2009. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/637.

Chazottes:2014:TFL

[CR14] Jean-René Chazottes and Frank Redig. Thermodynamic formalism and large deviations for multiplication-invariant potentials on lattice spin systems. Electronic Journal of Probability, 19:39:1–39:19, 2014. CODEN ???? ISSN 1083-

6489. URL http://ejp.ejpecp.org/article/view/3189.

Cerny:2022:SSC

[ČR22] Aleš Černý and Johannes Ruf. Simplified stochastic calculus via semimartingale representations. Electronic Journal of Probability, 27(??):1–32, 2022. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-27/issue-none/Simplified-
stochastic-calculus-via-semimartingale-representations/10.1214/21-
EJP729.full.

Criens:2021:ACS

[Cri21] David Criens. On absolute continuity and singularity of multidimensional diffusions. Electronic Journal of Probability, 26(??):1–26, 2021. CODEN ????

ISSN 1083-6489. URL https://projecteuclid.org/journals/electronic-
journal-of-probability/volume-26/issue-none/On-absolute-continuity-
and-singularity-of-multidimensional-diffusions/10.1214/20-EJP555.
full.

Csaki:1998:LFS

[CS98] Endre Csáki and Zhan Shi. Large favourite sites of simple random walk and the Wiener process. Electronic Journal of Probability, 3:14:1–14:31, 1998. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/36.

Can:2015:MCP

[CS15] Van Hao Can and Bruno Schapira. Metastability for the contact process on the configuration model with infinite mean degree. Electronic Journal of Probability, 20(??):26:1–26:22, ???? 2015. CODEN ???? ISSN 1083-6489. URL http://

ejp.ejpecp.org/article/view/3859.

Chandra:2017:MBS

[CS17] Ajay Chandra and Hao Shen. Moment bounds for SPDEs with non-Gaussian fields and application to the Wong-Zakai problem. Electronic Journal of Probability, 22(??):68:1–68:32, ???? 2017. CODEN ???? ISSN 1083-6489. URL

https://projecteuclid.org/euclid.ejp/1504922531.

Chleboun:2020:MSP

[CS20] Paul Chleboun and Aaron Smith. Mixing of the square plaquette model on a critical length scale. Electronic Journal of Probability, 25(??):1–53, 2020. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-25/issue-none/Mixing-of-
the-square-plaquette-model-on-a-critical-length/10.1214/20-EJP487.
full.

Chen:2008:CPE

[CSC08] Guan-Yu Chen and Laurent Saloff-Coste. The cutoff phenomenon for ergodic Markov processes. Electronic Journal of Probability, 13:3:26–3:78, 2008. CO-

DEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
474.

Chen:2015:CCT

[CSC15] Guan-Yu Chen and Laurent Saloff-Coste. Computing cutoff times of birth and death chains. Electronic Journal of Probability, 20(??):76:1–76:47, ???? 2015.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
4077.

Cygan:2022:CHS

[CSŠ22] Wojciech Cygan, Nikola Sandrić, and Stjepan Šebek. Convex hulls of stable random walks. Electronic Journal of Probability, 27(??):1–30, 2022. CODEN ????

ISSN 1083-6489. URL https://projecteuclid.org/journals/electronic-
journal-of-probability/volume-27/issue-none/Convex-hulls-of-stable-
random-walks/10.1214/22-EJP826.full.

Chatterjee:2022:EAK

[CSW22] Shirshendu Chatterjee, David Sivakoff, and Matthew Wascher. The effect of avoiding known infected neighbors on the persistence of a recurring infection process. Electronic Journal of Probability, 27(??):1–40, 2022. CODEN ????

ISSN 1083-6489. URL https://projecteuclid.org/journals/electronic-
journal-of-probability/volume-27/issue-none/The-effect-of-avoiding-
known-infected-neighbors-on-the-persistence/10.1214/22-EJP836.full.

Caravenna:2019:DSR

[CSZ19] Francesco Caravenna, Rongfeng Sun, and Nikos Zygouras. The Dickman subordinator, renewal theorems, and disordered systems. Electronic Journal of Probability, 24(??):101:1–101:40, ???? 2019. CODEN ???? ISSN 1083-6489.

URL https://projecteuclid.org/euclid.ejp/1568793795.

Coquet:2007:CVO

[CT07] François Coquet and Sandrine Toldo. Convergence of values in optimal stopping and convergence of optimal stopping times. Electronic Journal of Probability, 12: 8:207–8:228, 2007. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.

org/article/view/288.

Chouk:2015:SSI

[CT15] Khalil Chouk and Samy Tindel. Skorohod and Stratonovich integration in the plane. Electronic Journal of Probability, 20(??):39:1–39:39, ???? 2015. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/3041.

Cheng:2018:ESM

[CT18] Li-Juan Cheng and Anton Thalmaier. Evolution systems of measures and semigroup properties on evolving manifolds. Electronic Journal of Probability, 23 (??):20:1–20:27, ???? 2018. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1519722149.

Corwin:2020:SLW

[CT20] Ivan Corwin and Li-Cheng Tsai. SPDE limit of weakly inhomogeneous ASEP. Electronic Journal of Probability, 25(??):1–55, 2020. CODEN ????

ISSN 1083-6489. URL https://projecteuclid.org/journals/electronic-
journal-of-probability/volume-25/issue-none/SPDE-limit-of-weakly-
inhomogeneous-ASEP/10.1214/20-EJP565.full.

Can:2022:RCM

[CT22a] Van Hao Can and Khanh Duy Trinh. Random connection models in the thermodynamic regime: central limit theorems for add-one cost stabilizing functionals. Electronic Journal of Probability, 27(??):1–40, 2022. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-27/issue-none/Random-connection-models-in-the-
thermodynamic-regime--central-limit/10.1214/22-EJP759.full.

Cohen:2022:GTU

[CT22b] Samuel N. Cohen and Tanut Treetanthiploet. Gittins’ theorem under uncertainty. Electronic Journal of Probability, 27(??):1–48, 2022. CODEN ????

ISSN 1083-6489. URL https://projecteuclid.org/journals/electronic-
journal-of-probability/volume-27/issue-none/Gittins-theorem-under-
uncertainty/10.1214/22-EJP742.full.

Cormier:2021:HBM

[CTV21] Quentin Cormier, Etienne Tanré, and Romain Veltz. Hopf bifurcation in a mean-field model of spiking neurons. Electronic Journal of Probability, 26(18):1–40, February 2021. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-26/issue-none/Hopf-bifurcation-in-a-Mean-Field-
model-of-spiking-neurons/10.1214/21-EJP688.full.

Champagnat:2023:GCS

[CV23] Nicolas Champagnat and Denis Villemonais. General criteria for the study of quasi-stationarity. Electronic Journal of Probability, 28(??):1–84, 2023. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-28/issue-none/General-criteria-
for-the-study-of-quasi-stationarity/10.1214/22-EJP880.full.

Chelkak:2021:CML

[CW21a] Dmitry Chelkak and Yijun Wan. On the convergence of massive looperased random walks to massive SLE(2) curves. Electronic Journal of Probability, 26(18):1–35, February 2021. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-26/issue-none/On-the-convergence-of-massive-
loop-erased-random-walks-to/10.1214/21-EJP615.full.

Collins-Woodfin:2021:OSS

[CW21b] Elizabeth Collins-Woodfin. Overlaps of a spherical spin glass model with microscopic external field. Electronic Journal of Probability, 26(18):1–22, February 2021. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

journals/electronic-journal-of-probability/volume-26/issue-none/Overlaps-
of-a-spherical-spin-glass-model-with-microscopic-external/10.1214/
21-EJP722.full.

Chen:2023:GFE

[CW23] Zhen-Qing Chen and Jie-Ming Wang. Green function estimates for second order elliptic operators in non-divergence form with Dini continuous coefficients. Electronic Journal of Probability, 28(??):1–54, 2023. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-28/issue-none/Green-function-estimates-for-
second-order-elliptic-operators-in-non/10.1214/23-EJP925.full.

Cetin:2013:PPB

[CX13] Umut Cetin and Hao Xing. Point process bridges and weak convergence of insider trading models. Electronic Journal of Probability, 18:26:1–26:24, 2013.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
2039.

Chen:2021:SBS

[CY21] Xin Chen and Wenjie Ye. A study of backward stochastic differential equation on a Riemannian manifold. Electronic Journal of Probability, 26(18):1–31, February 2021. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

journals/electronic-journal-of-probability/volume-26/issue-none/A-
study-of-backward-stochastic-differential-equation-on-a-Riemannian/
10.1214/21-EJP649.full.

Collins:2022:SDS

[CYY22] Benoît Collins, Jianfeng Yao, and Wangjun Yuan. On spectral distribution of sample covariance matrices from large dimensional and large k-fold tensor products. Electronic Journal of Probability, 27(??):1–18, 2022. CODEN ????

ISSN 1083-6489. URL https://projecteuclid.org/journals/electronic-
journal-of-probability/volume-27/issue-none/On-spectral-distribution-
of-sample-covariance-matrices-from-large-dimensional/10.1214/22-EJP825.
full.

Cvitanic:2005:SDM

[CZ05] Jaksa Cvitanic and Jianfeng Zhang. The steepest descent method for forward– backward SDEs. Electronic Journal of Probability, 10:45:1468–45:1495, 2005.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
295.

Collevecchio:2021:NRV

[CZ21] Andrea Collevecchio and Xiaolin Zeng. A note on recurrence of the vertex reinforced jump process and fractional moments localization. Electronic Journal of Probability, 26(18):1–16, February 2021. CODEN ????

ISSN 1083-6489. URL https://projecteuclid.org/journals/electronic-
journal-of-probability/volume-26/issue-none/A-note-on-recurrence-
of-the-Vertex-reinforced-jump-process/10.1214/21-EJP609.full.

Dadoun:2017:ASS

[Dad17] Benjamin Dadoun. Asymptotics of self-similar growth-fragmentation processes. Electronic Journal of Probability, 22(??):27:1–27:30, ???? 2017. CODEN

???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.ejp/
1490061797.

Dalang:1999:EMM

[Dal99] Robert C. Dalang. Extending the martingale measure stochastic integral with applications to spatially homogeneous s.p.d.e.’s. Electronic Journal of Probability, 4:6:1–6:29, 1999. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/43.

Daly:2008:UBS

[Dal08] Fraser Daly. Upper bounds for Stein-type operators. Electronic Journal of Probability, 13:20:566–20:587, 2008. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/479.

Daletskii:2018:SDE

[Dal18] Alexei Daletskii. Stochastic differential equations in a scale of Hilbert spaces. Electronic Journal of Probability, 23(??):119:1–119:15, ???? 2018. CODEN

???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.ejp/
1544843299.

Dalmau:2021:WFM

[Dal21] Joseba Dalmau. The Wright–Fisher model for class–dependent fitness landscapes. Electronic Journal of Probability, 26(18):1–44, February 2021. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-26/issue-none/The-WrightFisher-
model-for-classdependent-fitness-landscapes/10.1214/21-EJP704.full.

Darling:2001:ILP

[Dar01] R. Darling. Intrinsic location parameter of a diffusion process. Electronic Journal of Probability, 7:3:1–3:23, 2001. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/102.

Davies:1998:LAE

[Dav98] Ian M. Davies. Laplace asymptotic expansions for Gaussian functional integrals. Electronic Journal of Probability, 3:13:1–13:19, 1998. CODEN ???? ISSN 1083-

6489. URL http://ejp.ejpecp.org/article/view/35.

Davies:2008:SAN

[Dav08] Ian Davies. Semiclassical analysis and a new result for Poisson–Lévy excursion measures. Electronic Journal of Probability, 13:45:1283–45:1306, 2008. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/513.

Davies:2010:EMM

[Dav10] E. Davies. Embeddable Markov matrices. Electronic Journal of Probability, 15:47:1474–47:1486, 2010. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/733.

Doring:2012:JTS

[DB12] Leif Döring and Matyas Barczy. Jump type SDEs for self-similar processes. Electronic Journal of Probability, 17:94:1–94:39, 2012. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/2402.

deBouard:2009:SDK

[dBD09] Anne de Bouard and Arnaud Debussche. Soliton dynamics for the Korteweg–de Vries equation with multiplicative homogeneous noise. Electronic Journal of Probability, 14:58:1727–58:1744, 2009. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/683.

Duminil-Copin:2018:URC

[DCLM18] Hugo Duminil-Copin, Jhih-Huang Li, and Ioan Manolescu. Universality for the random-cluster model on isoradial graphs. Electronic Journal of Probability, 23 (??):96:1–96:70, ???? 2018. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1537322681.

deCatelan:2021:FGP

[dCM21] Jacques de Catelan and Pierre-Loïc Méliot. Fluctuations of the Gromov– Prohorov sample model. Electronic Journal of Probability, 26(18):1–37, February 2021. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

journals/electronic-journal-of-probability/volume-26/issue-none/Fluctuations-
of-the-GromovProhorov-sample-model/10.1214/21-EJP634.full.

Dedecker:2012:RCS

[DDM12] Jérôme Dedecker, Paul Doukhan, and Florence Merlevède. Rates of convergence in the strong invariance principle under projective criteria. Electronic Journal of Probability, 17:16:1–16:31, 2012. CODEN ???? ISSN 1083-6489. URL http:

//ejp.ejpecp.org/article/view/1849.

Delmas:2018:CFL

[DDS18] Jean-François Delmas, Jean-Stéphane Dhersin, and Marion Sciauveau. Cost functionals for large (uniform and simply generated) random trees. Electronic Journal of Probability, 23(??):87:1–87:36, ???? 2018. CODEN ???? ISSN

1083-6489. URL https://projecteuclid.org/euclid.ejp/1536717746.

Delmas:2010:TOF

[DDSJ10] Jean-François Delmas, Jean-Stéphane Dhersin, and Arno Siri-Jegousse. On the two oldest families for the Wright–Fisher process. Electronic Journal of Probability, 15:26:776–26:800, 2010. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/771.

Dehling:2014:SEC

[DDT14] Herold Dehling, Olivier Durieu, and Marco Tusche. A sequential empirical CLT for multiple mixing processes with application to BB-geometrically ergodic Markov chains. Electronic Journal of Probability, 19(??):87:1–87:26, ???? 2014.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
3216.

DeSantis:2001:SIP

[De 01] Emilio De Santis. Strict inequality for phase transition between ferromagnetic and frustrated systems. Electronic Journal of Probability, 6:6:1–6:27, 2001. CO-

DEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/79.

Dony:2006:WUC

[DE06] Julia Dony and Uwe Einmahl. Weighted uniform consistency of kernel density estimators with general bandwidth sequences. Electronic Journal of Probability, 11:33:844–33:859, 2006. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/354.

Doring:2009:MDR

[DE09] Hanna Döring and Peter Eichelsbacher. Moderate deviations in a random graph and for the spectrum of Bernoulli random matrices. Electronic Journal of Probability, 14:92:2636–92:2656, 2009. CODEN ???? ISSN 1083-6489. URL http:/

/ejp.ejpecp.org/article/view/723.

DeBlassie:2008:EPB

[DeB08] Dante DeBlassie. The exit place of Brownian motion in the complement of a horn. Electronic Journal of Probability, 13:36:1068–36:1095, 2008. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/524.

DeBlassie:2009:EPB

[DeB09] Dante DeBlassie. The exit place of Brownian motion in an unbounded domain. Electronic Journal of Probability, 14:93:2657–93:2690, 2009. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/726.

Dawson:2002:MCB

[DEF^{+}02] Donald Dawson, Alison Etheridge, Klaus Fleischmann, Leonid Mytnik, Edwin Perkins, and Jie Xiong. Mutually catalytic branching in the plane: Infinite measure states. Electronic Journal of Probability, 7:15:1–15:61, 2002. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/114.

Delmas:2003:CML

[Del03] Jean-François Delmas. Computation of moments for the length of the onedimensional ISE support. Electronic Journal of Probability, 8:17:1–17:15, 2003.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
161.

Delyon:2009:EIS

[Del09] Bernard Delyon. Exponential inequalities for sums of weakly dependent variables. Electronic Journal of Probability, 14:28:752–28:779, 2009. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/636.

Dombry:2013:RCD

[DEM13] Clément Dombry and Frédéric Eyi-Minko. Regular conditional distributions of continuous max-infinitely divisible random fields. Electronic Journal of Probability, 18:7:1–7:21, 2013. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/1991.

Dereich:2016:PAF

[Der16] Steffen Dereich. Preferential attachment with fitness: unfolding the condensate. Electronic Journal of Probability, 21(??):3:1–3:38, ???? 2016. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.ejp/
1454514663.

Devroye:2011:NPC

[Dev11] Luc Devroye. A note on the probability of cutting a Galton–Watson tree. Electronic Journal of Probability, 16:72:2001–72:2019, 2011. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/952.

Deya:2011:DAR

[Dey11] Aurélien Deya. A discrete approach to rough parabolic equations. Electronic Journal of Probability, 16:55:1489–55:1518, 2011. CODEN ???? ISSN 1083-

6489. URL http://ejp.ejpecp.org/article/view/919.

Deuschel:2020:QTE

[DF20] Jean-Dominique Deuschel and Ryoki Fukushima. Quenched tail estimate for the random walk in random scenery and in random layered conductance II. Electronic Journal of Probability, 25(??):75:1–75:28, ???? 2020. CODEN ???? ISSN

1083-6489. URL https://projecteuclid.org/euclid.ejp/1593828036.

Dimitrov:2021:TBG

[DFF^{+}21] Evgeni Dimitrov, Xiang Fang, Lukas Fesser, Christian Serio, Carson Teitler, Angela Wang, and Weitao Zhu. Tightness of Bernoulli Gibbsian line ensembles. Electronic Journal of Probability, 26(18):1–93, February 2021. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-26/issue-none/Tightness-
of-Bernoulli-Gibbsian-line-ensembles/10.1214/21-EJP698.full.

DeMasi:2019:NLB

[DFPSL19] Anna De Masi, Pablo A. Ferrari, Errico Presutti, and Nahuel Soprano-Loto. Non local branching Brownian motions with annihilation and free boundary problems. Electronic Journal of Probability, 24(??):63:1–63:30, ???? 2019. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.ejp/
1561082672.

Demichel:2015:DEC

[DFS15] Yann Demichel, Ana-Karina Fermin, and Philippe Soulier. The diameter of an elliptical cloud. Electronic Journal of Probability, 20(??):27:1–27:32, ???? 2015.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/

Dolgopyat:2021:EBD

[DFS21] Dmitry Dolgopyat, Bassam Fayad, and Maria Saprykina. Erratic behavior for 1-dimensional random walks in a Liouville quasi-periodic environment. Electronic Journal of Probability, 26(18):1–36, February 2021. CODEN ????

ISSN 1083-6489. URL https://projecteuclid.org/journals/electronic-
journal-of-probability/volume-26/issue-none/Erratic-behavior-for-
1-dimensional-random-walks-in-a-Liouville/10.1214/21-EJP622.full.

Dawson:1996:MST

[DG96] Donald A. Dawson and Andreas Greven. Multiple space-time scale analysis for interacting branching models. Electronic Journal of Probability, 1:14:1–14:84, 1996. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/

view/14.

Dawson:1999:HIF

[DG99] Donald A. Dawson and Andreas Greven. Hierarchically interacting Fleming– Viot processes with selection and mutation: multiple space time scale analysis and quasi-equilibria. Electronic Journal of Probability, 4:4:1–4:81, 1999. CO-

DEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/41.

Dawson:2003:SDM

[DG03] Donald Dawson and Andreas Greven. State dependent multitype spatial branching processes and their longtime behavior. Electronic Journal of Probability, 8: 4:1–4:93, 2003. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/126.

Dereudre:2009:VCG

[DG09] David Dereudre and Hans-Otto Georgii. Variational characterisation of Gibbs measures with Delaunay triangle interaction. Electronic Journal of Probability, 14:85:2438–85:2462, 2009. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/713.

Dawson:2013:PUS

[DG13] Donald Dawson and Luis Gorostiza. Percolation in an ultrametric space. Electronic Journal of Probability, 18:12:1–12:26, 2013. CODEN ???? ISSN 1083-

6489. URL http://ejp.ejpecp.org/article/view/1789.

DOvidio:2014:MFA

[DG14] Mirko D’Ovidio and Roberto Garra. Multidimensional fractional advectiondispersion equations and related stochastic processes. Electronic Journal of Probability, 19:61:1–61:31, 2014. CODEN ???? ISSN 1083-6489. URL http://

ejp.ejpecp.org/article/view/2854.

Ding:2015:PAS

[DG15] Jian Ding and Subhajit Goswami. Percolation of averages in the stochastic mean field model: the near-supercritical regime. Electronic Journal of Probability, 20 (??):124:1–124:21, ???? 2015. CODEN ???? ISSN 1083-6489. URL http://

ejp.ejpecp.org/article/view/4111.

Dolinsky:2018:NSD

[DG18] Yan Dolinsky and Benjamin Gottesman. Numerical scheme for Dynkin games under model uncertainty. Electronic Journal of Probability, 23(??):74:1–74:20, ???? 2018. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.

org/euclid.ejp/1532678637.

Dareiotis:2020:RNE

[DG20a] Konstantinos Dareiotis and Máté Gerencsér. On the regularisation of the noise for the Euler–Maruyama scheme with irregular drift. Electronic Journal of Probability, 25(??):1–18, 2020. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-25/issue-none/On-the-regularisation-of-the-
noise-for-the-Euler--Maruyama/10.1214/20-EJP479.full.

Dareiotis:2020:NDE

[DG20b] Konstantinos Dareiotis and Benjamin Gess. Nonlinear diffusion equations with nonlinear gradient noise. Electronic Journal of Probability, 25(??):35:1–35:43, ???? 2020. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.

org/euclid.ejp/1585101794.

Deligiannidis:2021:BRR

[DGK21] George Deligiannidis, Sébastien Gouëzel, and Zemer Kosloff. The boundary of the range of a random walk and the Følner property. Electronic Journal of Probability, 26(18):1–39, February 2021. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-26/issue-none/The-boundary-of-the-range-of-
a-random-walk-and/10.1214/21-EJP667.full.

Devroye:2011:HDR

[DGLU11] Luc Devroye, András György, Gábor Lugosi, and Frederic Udina. Highdimensional random geometric graphs and their clique number. Electronic Journal of Probability, 16:90:2481–90:2508, 2011. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/967.

Dombry:2009:FAR

[DGP09] Clement Dombry and Nadine Guillotin-Plantard. A functional approach for random walks in random sceneries. Electronic Journal of Probability, 14:50:1495– 50:1512, 2009. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/659.

Depperschmidt:2013:PPT

[DGP13] Andrej Depperschmidt, Andreas Greven, and Peter Pfaffelhuber. Pathproperties of the tree-valued Fleming–Viot process. Electronic Journal of Probability, 18:84:1–84:47, 2013. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/2514.

Devulder:2018:CSW

[DGP18] Alexis Devulder, Nina Gantert, and Françoise Pène. Collisions of several walkers in recurrent random environments. Electronic Journal of Probability, 23 (??):90:1–90:34, ???? 2018. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1536717749.

Devulder:2019:AMW

[DGP19] Alexis Devulder, Nina Gantert, and Françoise Pène. Arbitrary many walkers meet infinitely often in a subballistic random environment. Electronic Journal of Probability, 24(??):100:1–100:25, ???? 2019. CODEN ???? ISSN 1083-6489.

URL https://projecteuclid.org/euclid.ejp/1568793794.

Dobler:2017:ITB

[DGV17] Christian Döbler, Robert E. Gaunt, and Sebastian J. Vollmer. An iterative technique for bounding derivatives of solutions of Stein equations. Electronic Journal of Probability, 22(??):96:1–96:39, ???? 2017. CODEN ???? ISSN

1083-6489. URL https://projecteuclid.org/euclid.ejp/1510802250.

Dawson:2004:HEB

[DGW04] Donald Dawson, Luis Gorostiza, and Anton Wakolbinger. Hierarchical equilibria of branching populations. Electronic Journal of Probability, 9:12:316–12:381, 2004. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/

view/200.

Diaconis:2002:RWT

[DH02] Persi Diaconis and Susan Holmes. Random walks on trees and matchings. Electronic Journal of Probability, 7:6:1–6:17, 2002. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/105.

Deijfen:2006:NCR

[DH06] Maria Deijfen and Olle Häggström. Nonmonotonic coexistence regions for the two-type Richardson model on graphs. Electronic Journal of Probability, 11: 13:331–13:344, 2006. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.

org/article/view/321.

Delmas:2013:WDS

[DH13] Jean-François Delmas and Olivier Hénard. A Williams decomposition for spatially dependent superprocesses. Electronic Journal of Probability, 18:37:1– 37:43, 2013. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/1801.

Dereudre:2015:IVC

[DH15] David Dereudre and Pierre Houdebert. Infinite volume continuum random cluster model. Electronic Journal of Probability, 20(??):125:1–125:24, ???? 2015.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
4718.

Dalang:2019:RFS

[DH19] Robert C. Dalang and Thomas Humeau. Random field solutions to linear SPDEs driven by symmetric pure jump Lévy space-time white noises. Electronic Journal of Probability, 24(??):60:1–60:28, ???? 2019. CODEN ???? ISSN 1083-6489.

URL https://projecteuclid.org/euclid.ejp/1561082669.

Deslandes:2021:LLL

[DH21] Clément Deslandes and Christian Houdré. On the limiting law of the length of the longest common and increasing subsequences in random words with arbitrary distribution. Electronic Journal of Probability, 26(18):1–27, February 2021.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-26/issue-none/On-the-limiting-
law-of-the-length-of-the-longest/10.1214/21-EJP612.full.

Denisov:2022:PAS

[DHKW22] Denis Denisov, Günter Hinrichs, Martin Kolb, and Vitali Wachtel. Persistence of autoregressive sequences with logarithmic tails. Electronic Journal of Probability, 27(??):1–43, 2022. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-27/issue-none/Persistence-of-autoregressive-
sequences-with-logarithmic-tails/10.1214/22-EJP879.full.

denHollander:2018:EED

[dHMRS18] F. den Hollander, M. Mandjes, A. Roccaverde, and N. J. Starreveld. Ensemble equivalence for dense graphs. Electronic Journal of Probability, 23(??): 12:1–12:26, ???? 2018. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1518426060.

denHollander:2012:MKD

[dHNT12] Frank den Hollander, Francesca Nardi, and Alessio Troiani. Metastability for Kawasaki dynamics at low temperature with two types of particles. Electronic Journal of Probability, 17:2:1–2:26, 2012. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/1693.

denHollander:2018:MHC

[dHNT18] Frank den Hollander, Francesca R. Nardi, and Siamak Taati. Metastability of hard-core dynamics on bipartite graphs. Electronic Journal of Probability, 23 (??):97:1–97:65, ???? 2018. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1537495434.

Deijfen:2011:SPG

[DHP11] Maria Deijfen, Alexander Holroyd, and Yuval Peres. Stable Poisson graphs in one dimension. Electronic Journal of Probability, 16:44:1238–44:1253, 2011.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
897.

Damron:2014:SCF

[DHS14a] Michael Damron, Jack Hanson, and Philippe Sosoe. Subdiffusive concentration in first passage percolation. Electronic Journal of Probability, 19(??):109:1– 109:27, ???? 2014. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.

org/article/view/3680.

Dembo:2014:WWG

[DHS14b] Amir Dembo, Ruojun Huang, and Vladas Sidoravicius. Walking within growing domains: recurrence versus transience. Electronic Journal of Probability, 19(??): 106:1–106:20, ???? 2014. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/3272.

Damron:2017:CDC

[DHS17] Michael Damron, Jack Hanson, and Philippe Sosoe. On the chemical distance in critical percolation. Electronic Journal of Probability, 22(??):75:1–75:43, ???? 2017. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

euclid.ejp/1505354464.

Devroye:2019:HSG

[DHS19] Luc Devroye, Cecilia Holmgren, and Henning Sulzbach. Heavy subtrees of Galton–Watson trees with an application to Apollonian networks. Electronic Journal of Probability, 24(??):2:1–2:44, ???? 2019. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/euclid.ejp/1549357219.

Dieuleveut:2016:USP

[Die16] Daphné Dieuleveut. The UIPQ seen from a point at infinity along its geodesic ray. Electronic Journal of Probability, 21(??):54:1–54:44, ???? 2016. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.ejp/
1473188081.

Diez:2020:PCM

[Die20] Antoine Diez. Propagation of chaos and moderate interaction for a piecewise deterministic system of geometrically enriched particles. Electronic Journal of Probability, 25(??):1–38, 2020. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-25/issue-none/Propagation-of-chaos-and-moderate-
interaction-for-a-piecewise-deterministic/10.1214/20-EJP496.full.

Ding:2012:CTL

[Din12] Jian Ding. On cover times for 2D lattices. Electronic Journal of Probability, 17: 45:1–45:18, 2012. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.

org/article/view/2089.

Dirksen:2015:TBG

[Dir15] Sjoerd Dirksen. Tail bounds via generic chaining. Electronic Journal of Probability, 20(??):53:1–53:29, ???? 2015. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/3760.

Djete:2022:EMF

[Dje22] Mao Fabrice Djete. Extended mean field control problem: a propagation of chaos result. Electronic Journal of Probability, 27(??):1–53, 2022. CODEN ????

ISSN 1083-6489. URL https://projecteuclid.org/journals/electronic-
journal-of-probability/volume-27/issue-none/Extended-mean-field-control-
problem--a-propagation-of-chaos/10.1214/21-EJP726.full.

Duse:2016:CAP

[DJM16] Erik Duse, Kurt Johansson, and Anthony Metcalfe. The cusp-Airy process. Electronic Journal of Probability, 21(??):57:1–57:50, ???? 2016. CODEN ???? ISSN

1083-6489. URL https://projecteuclid.org/euclid.ejp/1473424498.

Deya:2013:SHE

[DJQS13] Aurélien Deya, Maria Jolis, and Lluís Quer-Sardanyons. The Stratonovich heat equation: a continuity result and weak approximations. Electronic Journal of Probability, 18:3:1–3:34, 2013. CODEN ???? ISSN 1083-6489. URL http://

ejp.ejpecp.org/article/view/2004.

Dobler:2021:SME

[DK21] Christian Döbler and Mikołaj J. Kasprzak. Stein’s method of exchangeable pairs in multivariate functional approximations. Electronic Journal of Probability, 26(??):1–50, 2021. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-26/issue-none/Steins-method-of-exchangeable-
pairs-in-multivariate-functional-approximations/10.1214/21-EJP587.
full.

Duquesne:2022:SLT

[DKLT22] Thomas Duquesne, Robin Khanfir, Shen Lin, and Niccolò Torri. Scaling limits of tree-valued branching random walks. Electronic Journal of Probability, 27(??):1–54, 2022. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-27/issue-none/Scaling-limits-of-tree-valued-
branching-random-walks/10.1214/22-EJP741.full.

DelMoral:2001:CDG

[DKM01] P. Del Moral, M. Kouritzin, and L. Miclo. On a class of discrete generation interacting particle systems. Electronic Journal of Probability, 6:16:1–16:26, 2001. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/

view/89.

Dembo:2018:CLC

[DKN18] Amir Dembo, Takashi Kumagai, and Chikara Nakamura. Cutoff for lamplighter chains on fractals. Electronic Journal of Probability, 23(??):73:1–73:21, ???? 2018. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

euclid.ejp/1532678636.

Damron:2013:FCD

[DKNS13] Michael Damron, Hana Kogan, Charles Newman, and Vladas Sidoravicius. Fixation for coarsening dynamics in 2D slabs. Electronic Journal of Probability, 18: 105:1–105:20, 2013. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.

org/article/view/3059.

David:2017:RLQ

[DKRV17] François David, Antti Kupiainen, Rémi Rhodes, and Vincent Vargas. Renormalizability of Liouville quantum field theory at the Seiberg bound. Electronic Journal of Probability, 22(??):93:1–93:26, ???? 2017. CODEN ???? ISSN

1083-6489. URL https://projecteuclid.org/euclid.ejp/1509501716.

DelMoral:2018:SCE

[DKT18] Pierre Del Moral, Aline Kurtzmann, and Julian Tugaut. On the stability and the concentration of extended Kalman–Bucy filters. Electronic Journal of Probability, 23(??):91:1–91:30, ???? 2018. CODEN ???? ISSN 1083-6489. URL

https://projecteuclid.org/euclid.ejp/1536717750.

Duquesne:2014:EPC

[DL14] Thomas Duquesne and Cyril Labbé. On the Eve property for CSBP. Electronic Journal of Probability, 19:6:1–6:31, 2014. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/2831.

Daw:2022:WMS

[DL22] Lara Daw and Laurent Loosveldt. Wavelet methods to study the pointwise regularity of the generalized Rosenblatt process. Electronic Journal of Probability, 27(??):1–45, 2022. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-27/issue-none/Wavelet-methods-to-study-the-
pointwise-regularity-of-the-generalized/10.1214/22-EJP878.full.

Dalang:2021:MPG

[DLMX21] Robert C. Dalang, Cheuk Yin Lee, Carl Mueller, and Yimin Xiao. Multiple points of Gaussian random fields. Electronic Journal of Probability, 26(??):1– 25, 2021. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

journals/electronic-journal-of-probability/volume-26/issue-none/Multiple-
points-of-Gaussian-random-fields/10.1214/21-EJP589.full.

Delarue:2019:MEM

[DLR19] François Delarue, Daniel Lacker, and Kavita Ramanan. From the master equation to mean field game limit theory: a central limit theorem. Electronic Journal of Probability, 24(??):51:1–51:54, ???? 2019. CODEN ???? ISSN 1083-6489.

URL https://projecteuclid.org/euclid.ejp/1558576902.

Dupuis:2020:LDC

[DLR20] Paul Dupuis, Vaios Laschos, and Kavita Ramanan. Large deviations for configurations generated by Gibbs distributions with energy functionals consisting of singular interaction and weakly confining potentials. Electronic Journal of Probability, 25(??):46:1–46:41, ???? 2020. CODEN ???? ISSN 1083-6489. URL

https://projecteuclid.org/euclid.ejp/1587693777.

Dumitrescu:2021:COS

[DLT21] Roxana Dumitrescu, Marcos Leutscher, and Peter Tankov. Control and optimal stopping mean field games: a linear programming approach. Electronic Journal of Probability, 26(18):1–49, February 2021. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-26/issue-none/Control-and-optimal-stopping-
Mean-Field-Games--a-linear/10.1214/21-EJP713.full.

Dumaz:2021:OLH

[DLV21] Laure Dumaz, Yun Li, and Benedek Valkó. Operator level hard-to-soft transition for β-ensembles. Electronic Journal of Probability, 26(18):1–28, February 2021.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-26/issue-none/Operator-level-
hard-to-soft-transition-for-%ce%b2-ensembles/10.1214/21-EJP602.full.

Dawson:2001:SDS

[DLW01] Donald Dawson, Zenghu Li, and Hao Wang. Superprocesses with dependent spatial motion and general branching densities. Electronic Journal of Probability, 6:25:1–25:33, 2001. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.

org/article/view/98.

Drapeau:2020:CFC

[DLX20] Samuel Drapeau, Peng Luo, and Dewen Xiong. Characterization of fully coupled FBSDE in terms of portfolio optimization. Electronic Journal of Probability, 25 (??):24:1–24:26, ???? 2020. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1581735876.

Durrett:2014:CPF

[DLZ14] Rick Durrett, Thomas Liggett, and Yuan Zhang. The contact process with fast voting. Electronic Journal of Probability, 19:28:1–28:19, 2014. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/3021.

Dalang:2003:SNL

[DM03] Robert Dalang and Carl Mueller. Some non-linear S.P.D.E’s that are second order in time. Electronic Journal of Probability, 8:1:1–1:21, 2003. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/123.

Dereich:2009:RNS

[DM09] Steffen Dereich and Peter Mörters. Random networks with sublinear preferential attachment: Degree evolutions. Electronic Journal of Probability, 14:43:1222– 43:1267, 2009. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/647.

Dasgupta:2011:SLU

[DM11] Amites Dasgupta and Krishanu Maulik. Strong laws for urn models with balanced replacement matrices. Electronic Journal of Probability, 16:63:1723– 63:1749, 2011. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/928.

Dhersin:2013:EBC

[DM13] Jean-Stéphane Dhersin and Martin Möhle. On the external branches of coalescents with multiple collisions. Electronic Journal of Probability, 18:40:1–40:11, 2013. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/

view/2286.

Dobbs:2014:SDT

[DM14] Daniel Dobbs and Tai Melcher. Small deviations for time-changed Brownian motions and applications to second-order chaos. Electronic Journal of Probability, 19(??):85:1–85:23, ???? 2014. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/2993.

Drapeau:2016:SMP

[DM16] Samuel Drapeau and Christoph Mainberger. Stability and Markov property of forward backward minimal supersolutions. Electronic Journal of Probability, 21 (??):41:1–41:15, ???? 2016. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1466166072.

Deya:2022:FDR

[DM22] Aurélien Deya and Renaud Marty. A full discretization of the rough fractional linear heat equation. Electronic Journal of Probability, 27(??):1–41, 2022.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-27/issue-none/A-full-discretization-
of-the-rough-fractional-linear-heat-equation/10.1214/22-EJP839.full.

Dewan:2023:MFB

[DM23] Vivek Dewan and Stephen Muirhead. Mean-field bounds for Poisson–Boolean percolation. Electronic Journal of Probability, 28(??):1–24, 2023. CODEN ????

ISSN 1083-6489. URL https://projecteuclid.org/journals/electronic-
journal-of-probability/volume-28/issue-none/Mean-field-bounds-for-
Poisson-Boolean-percolation/10.1214/23-EJP923.full.

Dereich:2017:DSF

[DMM17] Steffen Dereich, Christian Mönch, and Peter Mörters. Distances in scale free networks at criticality. Electronic Journal of Probability, 22(??):77:1–77:38, ???? 2017. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

euclid.ejp/1506931227.

Dirksen:2013:PSI

[DMN13] Sjoerd Dirksen, Jan Maas, and Jan Neerven. Poisson stochastic integration in Banach spaces. Electronic Journal of Probability, 18:100:1–100:28, 2013. CO-

DEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/

Delarue:2015:LEM

[DMN15] François Delarue, Stéphane Menozzi, and Eulalia Nualart. The Landau equation for Maxwellian molecules and the Brownian motion on SO_{R}(N). Electronic Journal of Probability, 20(??):92:1–92:39, ???? 2015. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/4012.

Durrett:2005:CSB

[DMP05] Richard Durrett, Leonid Mytnik, and Edwin Perkins. Competing super-Brownian motions as limits of interacting particle systems. Electronic Journal of Probability, 10:35:1147–35:1220, 2005. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/229.

Dedecker:2009:RCM

[DMR09a] Jérôme Dedecker, Florence Merlevède, and Emmanuel Rio. Rates of convergence for minimal distances in the central limit theorem underprojective criteria. Electronic Journal of Probability, 14:35:978–35:1011, 2009. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/648.

Douc:2009:FIC

[DMR09b] Randal Douc, Eric Moulines, and Yaacov Ritov. Forgetting of the initial condition for the filter in general state-space hidden Markov chain: a coupling approach. Electronic Journal of Probability, 14:2:27–2:49, 2009. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/593.

Disertori:2016:CNS

[DMR16a] Margherita Disertori, Franz Merkl, and Silke W. W. Rolles. A comparison of a nonlinear sigma model with general pinning and pinning at one point. Electronic Journal of Probability, 21(??):27:1–27:16, ???? 2016. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/euclid.ejp/1460141798.

Donati-Martin:2016:NEE

[DMR16b] Catherine Donati-Martin and Alain Rouault. Near-extreme eigenvalues in the beta-ensembles. Electronic Journal of Probability, 21(??):52:1–52:17, ???? 2016.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.
ejp/1472142775.

Denis:2009:MPC

[DMS09] Laurent Denis, Anis Matoussi, and Lucretiu Stoica. Maximum principle and comparison theorem for quasi-linear stochastic PDE’s. Electronic Journal of Probability, 14:19:500–19:530, 2009. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/629.

Dhara:2017:PTE

[DMS17] Souvik Dhara, Debankur Mukherjee, and Subhabrata Sen. Phase transitions of extremal cuts for the configuration model. Electronic Journal of Probability, 22 (??):86:1–86:29, ???? 2017. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1507946761.

Davydov:2008:SSD

[DMZ08] Youri Davydov, Ilya Molchanov, and Sergei Zuyev. Strictly stable distributions on convex cones. Electronic Journal of Probability, 13:11:259–11:321, 2008. CO-

DEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
487.

Denis:2014:MPQ

[DMZ14] Laurent Denis, Anis Matoussi, and Jing Zhang. Maximum principle for quasilinear stochastic PDEs with obstacle. Electronic Journal of Probability, 19: 44:1–44:32, 2014. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.

org/article/view/2716.

Dalmao:2019:PSC

[DNPR19] Federico Dalmao, Ivan Nourdin, Giovanni Peccati, and Maurizia Rossi. Phase singularities in complex arithmetic random waves. Electronic Journal of Probability, 24(??):71:1–71:45, ???? 2019. CODEN ???? ISSN 1083-6489. URL

https://projecteuclid.org/euclid.ejp/1561687604.

Do:2021:RRR

[Do21] Yen Q. Do. Real roots of random polynomials with coefficients of polynomial growth: a comparison principle and applications. Electronic Journal of Probability, 26(18):1–45, February 2021. CODEN ????

ISSN 1083-6489. URL https://projecteuclid.org/journals/electronic-
journal-of-probability/volume-26/issue-none/Real-roots-of-random-
polynomials-with-coefficients-of-polynomial-growth/10.1214/21-EJP719.
full.

Dobler:2015:SME

[Döb15] Christian Döbler. Stein’s method of exchangeable pairs for the beta distribution and generalizations. Electronic Journal of Probability, 20(??):109:1–109:34, ???? 2015. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/

view/3933.

Dohmen:1999:IIE

[Doh99] Klaus Dohmen. Improved inclusion–exclusion identities and inequalities based on a particular class of abstract tubes. Electronic Journal of Probability, 4:

5:1–5:12, 1999. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.

org/article/view/42.

Dolinsky:2012:NSE

[Dol12] Yan Dolinsky. Numerical schemes for G-expectations. Electronic Journal of Probability, 17:98:1–98:15, 2012. CODEN ???? ISSN 1083-6489. URL http://

ejp.ejpecp.org/article/view/2284.

Dolgopyat:2016:LLT

[Dol16] Dmitry Dolgopyat. A Local Limit Theorem for sums of independent random vectors. Electronic Journal of Probability, 21(??):39:1–39:15, ???? 2016. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.ejp/
1465991837.

Dominguez:2022:GGP

[Dom22] Tomas Dominguez. The ℓ^{p} Gaussian–Grothendieck problem with vector spins. Electronic Journal of Probability, 27(??):1–46, 2022. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-27/issue-none/The-%e2%84%93p-Gaussian-Grothendieck-
problem-with-vector-spins/10.1214/22-EJP801.full.

Doney:2004:STB

[Don04] Ronald Doney. Small-time behaviour of Lévy processes. Electronic Journal of Probability, 9:8:209–8:229, 2004. CODEN ???? ISSN 1083-6489. URL http://

ejp.ejpecp.org/article/view/193.

DaiPra:2005:LSI

[DP05] Paolo Dai Pra and Gustavo Posta. Logarithmic Sobolev inequality for zero– range dynamics: Independence of the number of particles. Electronic Journal of Probability, 10:15:525–15:576, 2005. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/259.

DaiPra:2013:EDI

[DP13a] Paolo Dai Pra and Gustavo Posta. Entropy decay for interacting systems via the Bochner–Bakry–Émery approach. Electronic Journal of Probability, 18:52:1– 52:21, 2013. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/2041.

Devulder:2013:RWR

[DP13b] Alexis Devulder and Françoise Pène. Random walk in random environment in a two-dimensional stratified medium with orientations. Electronic Journal of Probability, 18:18:1–18:23, 2013. CODEN ???? ISSN 1083-6489. URL http://

ejp.ejpecp.org/article/view/2459.

Doumas:2013:ARM

[DP13c] Aristides Doumas and Vassilis Papanicolaou. Asymptotics of the rising moments for the coupon collector’s problem. Electronic Journal of Probability, 18:41:1– 41:15, 2013. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/1746.

Dong:2014:MMD

[DP14] Zhao Dong and Xuhui Peng. Malliavin matrix of degenerate SDE and gradient estimate. Electronic Journal of Probability, 19:73:1–73:26, 2014. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/3120.

Dobler:2017:QJT

[DP17] Christian Döbler and Giovanni Peccati. Quantitative de Jong theorems in any dimension. Electronic Journal of Probability, 22(??):2:1–2:35, ???? 2017. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.ejp/
1483585524.

Dalang:2020:OLB

[DP20] Robert C. Dalang and Fei Pu. Optimal lower bounds on hitting probabilities for stochastic heat equations in spatial dimension. Electronic Journal of Probability, 25(??):1–31, 2020. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-25/issue-none/Optimal-lower-bounds-on-hitting-
probabilities-for-stochastic-heat-equations/10.1214/20-EJP438.full.

Depperschmidt:2015:MTV

[DPP15] Andrej Depperschmidt, Étienne Pardoux, and Peter Pfaffelhuber. A mixing tree-valued process arising under neutral evolution with recombination. Electronic Journal of Probability, 20(??):94:1–94:22, ???? 2015. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/4286.

Dembo:2003:BMC

[DPR03] Amir Dembo, Yuval Peres, and Jay Rosen. Brownian motion on compact manifolds: Cover time and late points. Electronic Journal of Probability, 8:15:1– 15:14, 2003. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/139.

Dembo:1999:TPT

[DPRZ99] Amir Dembo, Yuval Peres, Jay Rosen, and Ofer Zeitouni. Thick points for transient symmetric stable processes. Electronic Journal of Probability, 4:10:1– 10:13, 1999. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/47.

DeMasi:2012:TCS

[DPTV12] Anna De Masi, Errico Presutti, Dimitrios Tsagkarogiannis, and Maria Vares. Truncated correlations in the stirring process with births and deaths. Electronic Journal of Probability, 17:6:1–6:35, 2012. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/1734.

Deligiannidis:2015:AVS

[DPU15] George Deligiannidis, Magda Peligrad, and Sergey Utev. Asymptotic variance of stationary reversible and normal Markov processes. Electronic Journal of Probability, 20(??):20:1–20:26, ???? 2015. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/3183.

Darling:2014:RDS

[DPWZ14] Richard W. R. Darling, Mathew D. Penrose, Andrew R. Wade, and Sandy L. Zabell. Rank deficiency in sparse random GF[2][2] matrices. Electronic Journal of Probability, 19(??):83:1–83:36, ???? 2014. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/2458.

Dumitrescu:2016:GDG

[DQS16] Roxana Dumitrescu, Marie-Claire Quenez, and Agnès Sulem. Generalized Dynkin games and doubly reflected BSDEs with jumps. Electronic Journal of Probability, 21(??):64:1–64:32, ???? 2016. CODEN ???? ISSN 1083-6489.

URL https://projecteuclid.org/euclid.ejp/1477395747.

Dick:2014:DEV

[DR14] Josef Dick and Daniel Rudolf. Discrepancy estimates for variance bounding Markov chain quasi-Monte Carlo. Electronic Journal of Probability, 19(??): 105:1–105:24, ???? 2014. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/3132.

Doney:2015:ABF

[DR15a] Ronald Arthur Doney and Victor Rivero. Asymptotic behaviour of first passage time distributions for subordinators. Electronic Journal of Probability, 20(??): 91:1–91:28, ???? 2015. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/3879.

Drewitz:2015:HDA

[DR15b] Alexander Drewitz and Pierre-François Rodriguez. High-dimensional asymptotics for percolation of Gaussian free field level sets. Electronic Journal of Probability, 20(??):47:1–47:39, ???? 2015. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/3416.

Drewitz:2008:LEO

[Dre08] Alexander Drewitz. Lyapunov exponents for the one-dimensional parabolic Anderson model with drift. Electronic Journal of Probability, 13:76:2283–76:2336, 2008. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/

view/586.

Dedecker:2014:SAE

[DRM14] Jérôme Dedecker, Emmanuel Rio, and Florence Merlevède. Strong approximation of the empirical distribution function for absolutely regular sequences in R^{d}. Electronic Journal of Probability, 19:9:1–9:56, 2014. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/2658.

deRaynal:2022:MSD

[dRM22] Paul-Éric Chaudru de Raynal and Stéphane Menozzi. On multidimensional stable-driven stochastic differential equations with Besov drift. Electronic Journal of Probability, 27(??):1–52, 2022. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-27/issue-none/On-multidimensional-stable-driven-
stochastic-differential-equations-with-Besov-drift/10.1214/22-EJP864.
full.

Denis:2004:GAR

[DS04] Laurent Denis and L. Stoica. A general analytical result for non-linear SPDE’s and applications. Electronic Journal of Probability, 9:23:674–23:709, 2004. CO-

DEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
223.

Dereich:2006:HRQ

[DS06] Steffen Dereich and Michael Scheutzow. High resolution quantization and entropy coding for fractional Brownian motion. Electronic Journal of Probability, 11:28:700–28:722, 2006. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/344.

Darses:2007:TRD

[DS07a] Sebastien Darses and Bruno Saussereau. Time reversal for drifted fractional Brownian motion with Hurst index H > 1/2. Electronic Journal of Probability, 12:43:1181–43:1211, 2007. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/439.

Dietz:2007:OLS

[DS07b] Zach Dietz and Sunder Sethuraman. Occupation laws for some timenonhomogeneous Markov chains. Electronic Journal of Probability, 12:23:661– 23:683, 2007. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/413.

Dimitroff:2011:AEB

[DS11a] Georgi Dimitroff and Michael Scheutzow. Attractors and expansion for Brownian flows. Electronic Journal of Probability, 16:42:1193–42:1213, 2011. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/894.

Dobler:2011:SMM

[DS11b] Christian Döbler and Michael Stolz. Stein’s method and the multivariate CLT for traces of powers on the compact classical groups. Electronic Journal of Probability, 16:86:2375–86:2405, 2011. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/960.

Doring:2011:NDA

[DS11c] Leif Döring and Mladen Savov. (non)differentiability and asymptotics for potential densities of subordinators. Electronic Journal of Probability, 16:17:470– 17:503, 2011. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/860.

DeSantis:2012:FOW

[DS12a] Emilio De Santis and Fabio Spizzichino. First occurrence of a word among the elements of a finite dictionary in random sequences of letters. Electronic Journal of Probability, 17:25:1–25:9, 2012. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/1878.

Dembo:2012:CLT

[DS12b] Amir Dembo and Nike Sun. Central limit theorem for biased random walk on multi-type Galton–Watson trees. Electronic Journal of Probability, 17:75:1– 75:40, 2012. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/2294.

dosSantos:2014:NTL

[dS14] Renato Soares dos Santos. Non-trivial linear bounds for a random walk driven by a simple symmetric exclusion process. Electronic Journal of Probability, 19: 49:1–49:18, 2014. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.

org/article/view/3159.

Deng:2017:HIS

[DS17] Chang-Song Deng and René L. Schilling. Harnack inequalities for SDEs driven by time-changed fractional Brownian motions. Electronic Journal of Probability, 22(??):71:1–71:23, ???? 2017. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1505268102.

Dorsch:2019:RPH

[DSB19] Florian Dorsch and Hermann Schulz-Baldes. Random perturbations of hyperbolic dynamics. Electronic Journal of Probability, 24(??):89:1–89:23, ???? 2019.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.
ejp/1568080869.

Drewitz:2015:ALP

[DSWB15] Alexander Drewitz, Michael Scheutzow, and Maite Wilke-Berenguer. Asymptotics for Lipschitz percolation above tilted planes. Electronic Journal of Probability, 20(??):117:1–117:23, ???? 2015. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/4251.

Driver:2020:OBR

[DT20] David P. Driver and Michael R. Tehranchi. Optimisation-based representations for branching processes. Electronic Journal of Probability, 25(??):1–15, 2020.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-25/issue-none/Optimisation-
based-representations-for-branching-processes/10.1214/20-EJP548.full.

deTiliere:2021:ZDM

[dT21] Béatrice de Tilière. The Z-dirac and massive Laplacian operators in the Zinvariant Ising model. Electronic Journal of Probability, 26(18):1–86, February 2021. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

journals/electronic-journal-of-probability/volume-26/issue-none/The-
Z-Dirac-and-massive-Laplacian-operators-in-the-Z/10.1214/21-EJP601.
full.

Dette:2018:URM

[DTV18] Holger Dette, Dominik Tomecki, and Martin Venker. Universality in random moment problems. Electronic Journal of Probability, 23(??):15:1–15:23, ???? 2018. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

euclid.ejp/1519354944.

Durhuus:2022:TPL

[DÜ22] Bergfinnur Durhuus and Meltem Ünel. Trees with power-like height dependent weight. Electronic Journal of Probability, 27(??):1–24, 2022. CODEN ????

ISSN 1083-6489. URL https://projecteuclid.org/journals/electronic-
journal-of-probability/volume-27/issue-none/Trees-with-power-like-
height-dependent-weight/10.1214/22-EJP857.full.

Dubach:2018:PGE

[Dub18] Guillaume Dubach. Powers of Ginibre eigenvalues. Electronic Journal of Probability, 23(??):111:1–111:31, ???? 2018. CODEN ???? ISSN 1083-6489. URL

https://projecteuclid.org/euclid.ejp/1540865374.

Dubach:2021:ESS

[Dub21] Guillaume Dubach. On eigenvector statistics in the spherical and truncated unitary ensembles. Electronic Journal of Probability, 26(18):1–29, February 2021.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-26/issue-none/On-eigenvector-
statistics-in-the-spherical-and-truncated-unitary-ensembles/10.1214/
21-EJP686.full.

Dufresne:2010:GDB

[Duf10] Daniel Dufresne. G distributions and the beta-gamma algebra. Electronic Journal of Probability, 15:71:2163–71:2199, 2010. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/845.

Dumaz:2012:CSR

[Dum12] Laure Dumaz. A clever (self-repelling) burglar. Electronic Journal of Probability, 17:61:1–61:17, 2012. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.

org/article/view/1758.

Durrett:2014:SEG

[Dur14] Rick Durrett. Spatial evolutionary games with small selection coefficients. Electronic Journal of Probability, 19(??):121:1–121:64, ???? 2014. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/3621.

Dhara:2017:CWC

[DvdHvLS17] Souvik Dhara, Remco van der Hofstad, Johan S. H. van Leeuwaarden, and Sanchayan Sen. Critical window for the configuration model: finite third moment degrees. Electronic Journal of Probability, 22(??):16:1–16:33, ???? 2017.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.
ejp/1487127644.

Denisov:2014:LPR

[DVW14] Denis Denisov, Vladimir Vatutin, and Vitali Wachtel. Local probabilities for random walks with negative drift conditioned to stay nonnegative. Electronic Journal of Probability, 19(??):88:1–88:17, ???? 2014. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/3426.

Dobler:2018:FMT

[DVZ18] Christian Döbler, Anna Vidotto, and Guangqu Zheng. Fourth moment theorems on the Poisson space in any dimension. Electronic Journal of Probability, 23 (??):36:1–36:27, ???? 2018. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1525312960.

Denisov:2010:CLT

[DW10] Denis Denisov and Vitali Wachtel. Conditional limit theorems for ordered random walks. Electronic Journal of Probability, 15:11:292–11:322, 2010. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/752.

Denisov:2012:ORW

[DW12] Denis Denisov and Vitali Wachtel. Ordered random walks with heavy tails. Electronic Journal of Probability, 17:4:1–4:21, 2012. CODEN ???? ISSN 1083-

6489. URL http://ejp.ejpecp.org/article/view/1719.

Damron:2016:ERE

[DW16a] Michael Damron and Xuan Wang. Entropy reduction in Euclidean first-passage percolation. Electronic Journal of Probability, 21(??):65:1–65:23, ???? 2016.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.
ejp/1479524422.

Durieu:2016:IUS

[DW16b] Olivier Durieu and Yizao Wang. From infinite urn schemes to decompositions of self-similar Gaussian processes. Electronic Journal of Probability, 21 (??):43:1–43:23, ???? 2016. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1469557136.

Debussche:2018:SES

[DW18a] Arnaud Debussche and Hendrik Weber. The Schrödinger equation with spatial white noise potential. Electronic Journal of Probability, 23(??):28:1–28:16, ???? 2018. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

euclid.ejp/1522375268.

Durieu:2018:FRS

[DW18b] Olivier Durieu and Yizao Wang. A family of random sup-measures with longrange dependence. Electronic Journal of Probability, 23(??):107:1–107:24, ???? 2018. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

euclid.ejp/1540260053.

Denisov:2019:ACH

[DW19] Denis Denisov and Vitali Wachtel. Alternative constructions of a harmonic function for a random walk in a cone. Electronic Journal of Probability, 24 (??):92:1–92:26, ???? 2019. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1568253841.

Doring:2019:LPF

[DWW19] Leif Döring, Alexander R. Watson, and Philip Weissmann. Lévy processes with finite variance conditioned to avoid an interval. Electronic Journal of Probability, 24(??):55:1–55:32, ???? 2019. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1559700305.

Durrett:2020:SCP

[DY20] Rick Durrett and Dong Yao. The symbiotic contact process. Electronic Journal of Probability, 25(??):4:1–4:21, ???? 2020. CODEN ???? ISSN 1083-6489. URL

https://projecteuclid.org/euclid.ejp/1579143695.

Durrett:2022:SIE

[DY22] Rick Durrett and Dong Yao. Susceptible–infected epidemics on evolving graphs. Electronic Journal of Probability, 27(??):1–66, 2022. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-27/issue-none/Susceptibleinfected-epidemics-
on-evolving-graphs/10.1214/22-EJP828.full.

Diomande:2015:MPO

[DZ15] Bakarime Diomande and Adrian Zalinescu. Maximum principle for an optimal control problem associated to a stochastic variational inequality with delay. Electronic Journal of Probability, 20(??):12:1–12:35, ???? 2015. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/2741.

Dassios:2021:EST

[DZ21] Angelos Dassios and Junyi Zhang. Exact simulation of two-parameter Poisson– Dirichlet random variables. Electronic Journal of Probability, 26(??):1–20, 2021.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-26/issue-none/Exact-simulation-
of-two-parameter-Poisson--Dirichlet-random-variables/10.1214/20-EJP573.
full.

Das:2022:UTL

[DZ22] Sayan Das and Weitao Zhu. Upper-tail large deviation principle for the ASEP. Electronic Journal of Probability, 27(??):1–34, 2022. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-27/issue-none/Upper-tail-large-deviation-principle-
for-the-ASEP/10.1214/21-EJP730.full.

Enriquez:2022:DFE

[EFMN22] Nathanaël Enriquez, Gabriel Faraud, Laurent Ménard, and Nathan Noiry. Depth first exploration of a configuration model. Electronic Journal of Probability, 27(??):1–27, 2022. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/journals/electronic-journal-of-probability/volume-
27/issue-none/Depth-first-exploration-of-a-configuration-model/10.
1214/22-EJP762.full.

Etheridge:2017:BBM

[EFP17] Alison Etheridge, Nic Freeman, and Sarah Penington. Branching Brownian motion, mean curvature flow and the motion of hybrid zones. Electronic Journal of Probability, 22(??):103:1–103:40, ???? 2017. CODEN ???? ISSN 1083-6489.

URL https://projecteuclid.org/euclid.ejp/1512615692.

Eon:2015:GAN

[EG15] Richard Eon and Mihai Gradinaru. Gaussian asymptotics for a non-linear Langevin type equation driven by an α-stable Lévy noise. Electronic Journal of Probability, 20(??):100:1–100:19, ???? 2015. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/4068.

Eldan:2018:DMF

[EG18] Ronen Eldan and Renan Gross. Decomposition of mean-field Gibbs distributions into product measures. Electronic Journal of Probability, 23(??):35:1–35:24, ???? 2018. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.

org/euclid.ejp/1524880979.

Etheridge:2022:EWO

[EGL22] Alison M. Etheridge, Mitchel D. Gooding, and Ian Letter. On the effects of a wide opening in the domain of the (stochastic) Allen–Cahn equation and the motion of hybrid zones. Electronic Journal of Probability, 27(??):1–53, 2022.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-27/issue-none/On-the-effects-
of-a-wide-opening-in-the-domain/10.1214/22-EJP888.full.

Eckhoff:2020:LPF

[EGvdHN20] Maren Eckhoff, Jesse Goodman, Remco van der Hofstad, and Francesca R. Nardi. Long paths in first passage percolation on the complete graph I. Local PWIT dynamics. Electronic Journal of Probability, 25(??):1–45, 2020. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-25/issue-none/Long-paths-
in-first-passage-percolation-on-the-complete-graph/10.1214/20-EJP484.
full.

Evans:2012:TPT

[EGW12] Steven Evans, Rudolf Grübel, and Anton Wakolbinger. Trickle-down processes and their boundaries. Electronic Journal of Probability, 17:1:1–1:58, 2012. CO-

DEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
1698.

Erbar:2017:RCB

[EHMT17] Matthias Erbar, Christopher Henderson, Georg Menz, and Prasad Tetali. Ricci curvature bounds for weakly interacting Markov chains. Electronic Journal of Probability, 22(??):40:1–40:23, ???? 2017. CODEN ???? ISSN 1083-6489. URL

https://projecteuclid.org/euclid.ejp/1493345027.

Engelke:2016:LDP

[EI16] Sebastian Engelke and Jevgenijs Ivanovs. A Lévy-derived process seen from its supremum and max-stable processes. Electronic Journal of Probability, 21 (??):14:1–14:19, ???? 2016. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1456246245.

Eisenbaum:2005:CBG

[Eis05] Nathalie Eisenbaum. A connection between Gaussian processes and Markov processes. Electronic Journal of Probability, 10:6:202–6:215, 2005. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/238.

Eisenbaum:2013:IPP

[Eis13] Nathalie Eisenbaum. Inequalities for permanental processes. Electronic Journal of Probability, 18:99:1–99:15, 2013. CODEN ???? ISSN 1083-6489. URL http:

//ejp.ejpecp.org/article/view/2919.

Eisenbaum:2019:DID

[Eis19] Nathalie Eisenbaum. Decompositions of infinitely divisible nonnegative processes. Electronic Journal of Probability, 24(??):109:1–109:25, ???? 2019. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.ejp/
1569981824.

Emrah:2021:FSC

[EJS21] Elnur Emrah, Christopher Janjigian, and Timo Seppäläinen. Flats, spikes and crevices: the evolving shape of the inhomogeneous corner growth model. Electronic Journal of Probability, 26(??):1–45, 2021. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-26/issue-none/Flats-spikes-and-crevices--the-
evolving-shape-of-the/10.1214/21-EJP595.full.

Ercolani:2014:RPS

[EJU14] Nicholas M. Ercolani, Sabine Jansen, and Daniel Ueltschi. Random partitions in statistical mechanics. Electronic Journal of Probability, 19(??):82:1–82:37, ???? 2014. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/3244.

Eichelsbacher:2008:ORW

[EK08] Peter Eichelsbacher and Wolfgang König. Ordered random walks. Electronic Journal of Probability, 13:46:1307–46:1336, 2008. CODEN ???? ISSN 1083-

6489. URL http://ejp.ejpecp.org/article/view/539.

Eisenbaum:2022:ITE

[EK22] Nathalie Eisenbaum and Haya Kaspi. Isomorphism theorems, extended Markov processes and random interlacements. Electronic Journal of Probability, 27(??): 1–27, 2022. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.

org/journals/electronic-journal-of-probability/volume-27/issue-none/
Isomorphism-theorems-extended-Markov-processes-and-random-interlacements/
10.1214/22-EJP887.full. See addendum [EK23].

Eisenbaum:2023:AIT

[EK23] Nathalie Eisenbaum and Haya Kaspi. Addendum to “Isomorphism theorems, extended Markov processes and random interlacements”. Electronic Journal of Probability, 28(??):1–3, 2023. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-28/issue-none/Addendum-to-Isomorphism-theorems-
extended-Markov-processes-and-random-interlacements/10.1214/23-EJP935.

full. See [EK22].

Erdos:2013:LSL

[EKYY13] László Erdős, Antti Knowles, Horng-Tzer Yau, and Jun Yin. The local semicircle law for a general class of random matrices. Electronic Journal of Probability, 18: 59:1–59:58, 2013. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.

org/article/view/2473.

Evans:2007:AEA

[EL07a] Steven Evans and Tye Lidman. Asymptotic evolution of acyclic random mappings. Electronic Journal of Probability, 12:42:1051–42:1180, 2007. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/437.

Evans:2007:ECE

[EL07b] Steven Evans and Tye Lidman. Expectation, conditional expectation and martingales in local fields. Electronic Journal of Probability, 12:17:498–17:515, 2007.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
405.

Ethier:2009:LTP

[EL09] S. Ethier and Jiyeon Lee. Limit theorems for Parrondo’s paradox. Electronic Journal of Probability, 14:62:1827–62:1862, 2009. CODEN ???? ISSN 1083-

6489. URL http://ejp.ejpecp.org/article/view/684.

Eichelsbacher:2010:SMD

[EL10] Peter Eichelsbacher and Matthias Loewe. Stein’s method for dependent random variables occurring in statistical mechanics. Electronic Journal of Probability, 15:30:962–30:988, 2010. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/777.

Etheridge:2011:DAM

[EL11] Alison Etheridge and Sophie Lemaire. Diffusion approximation of a multilocus model with assortative mating. Electronic Journal of Probability, 16:78:2122– 78:2181, 2011. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/932.

Ethier:2012:PPR

[EL12] Stewart Ethier and Jiyeon Lee. Parrondo’s paradox via redistribution of wealth. Electronic Journal of Probability, 17:20:1–20:21, 2012. CODEN ???? ISSN 1083-

6489. URL http://ejp.ejpecp.org/article/view/1867.

Evilsizor:2014:EGL

[EL14] Stephen Evilsizor and Nicolas Lanchier. Evolutionary games on the lattice: best-response dynamics. Electronic Journal of Probability, 19:75:1–75:12, 2014.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
3126.

Evilsizor:2016:EGL

[EL16] Stephen Evilsizor and Nicolas Lanchier. Evolutionary games on the lattice: death–birth updating process. Electronic Journal of Probability, 21(??): 17:1–17:29, ???? 2016. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1456412957.

Eldan:2014:VPC

[Eld14] Ronen Eldan. Volumetric properties of the convex hull of an n-dimensional Brownian motion. Electronic Journal of Probability, 19:45:1–45:34, 2014. CO-

DEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/

Erny:2021:CPC

[ELL21] Xavier Erny, Eva Löcherbach, and Dasha Loukianova. Conditional propagation of chaos for mean field systems of interacting neurons. Electronic Journal of Probability, 26(??):1–25, 2021. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-26/issue-none/Conditional-propagation-of-chaos-
for-mean-field-systems-of-interacting/10.1214/21-EJP580.full.

Etore:2012:ETI

[ÉM12] Pierre Étoré and Miguel Martinez. On the existence of a time inhomogeneous skew Brownian motion and some related laws. Electronic Journal of Probability, 17:19:1–19:27, 2012. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.

org/article/view/1858.

Eckhoff:2014:VRP

[EM14] Maren Eckhoff and Peter Mörters. Vulnerability of robust preferential attachment networks. Electronic Journal of Probability, 19:57:1–57:47, 2014. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/2974.

Eberle:2019:QCR

[EM19] Andreas Eberle and Mateusz B. Majka. Quantitative contraction rates for Markov chains on general state spaces. Electronic Journal of Probability, 24 (??):26:1–26:36, ???? 2019. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1553565777.

Englander:2010:CMS

[Eng10] Janos Englander. The center of mass for spatial branching processes and an application for self-interaction. Electronic Journal of Probability, 15:63:1938– 63:1970, 2010. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/822.

Evans:1998:CLT

[EP98] Steven N. Evans and Edwin A. Perkins. Collision local times, historical stochastic calculus, and competing species. Electronic Journal of Probability, 3:5:1– 5:120, 1998. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/27.

Etheridge:2022:GBW

[EP22] Alison Etheridge and Sarah Penington. Genealogies in bistable waves. Electronic Journal of Probability, 27(??):1–99, 2022. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-27/issue-none/Genealogies-in-bistable-waves/
10.1214/22-EJP845.full.

Enger:2023:UFL

[EP23] Timo Enger and Peter Pfaffelhuber. A unified framework for limit results in chemical reaction networks on multiple time-scales. Electronic Journal of Probability, 28(??):1–33, 2023. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-28/issue-none/A-unified-framework-for-limit-
results-in-chemical-reaction-networks/10.1214/22-EJP897.full.

Erdos:2010:USK

[ERSY10] Laszlo Erdos, Jose Ramirez, Benjamin Schlein, and Horng-Tzer Yau. Universality of sine-kernel for Wigner matrices with a small Gaussian perturbation. Electronic Journal of Probability, 15:18:526–18:604, 2010. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/768.

Elie:2014:ENB

[ERY14] Romuald Elie, Mathieu Rosenbaum, and Marc Yor. On the expectation of normalized Brownian functionals up to first hitting times. Electronic Journal of Probability, 19:37:1–37:23, 2014. CODEN ???? ISSN 1083-6489. URL http:

//ejp.ejpecp.org/article/view/3049.

Erignoux:2019:EFD

[ES19] Clément Erignoux and Marielle Simon. Equilibrium fluctuations for the disordered harmonic chain perturbed by an energy conserving noise. Electronic Journal of Probability, 24(??):143:1–143:52, ???? 2019. CODEN ???? ISSN

1083-6489. URL https://projecteuclid.org/euclid.ejp/1576810975.

Eichelsbacher:2014:NBE

[ET14] Peter Eichelsbacher and Christoph Thäle. New Berry–Esséen bounds for nonlinear functionals of Poisson random measures. Electronic Journal of Probability, 19(??):102:1–102:25, ???? 2014. CODEN ???? ISSN 1083-6489. URL http://

ejp.ejpecp.org/article/view/3061.

Eichelsbacher:2015:MSM

[ET15] Peter Eichelsbacher and Christoph Thäle. Malliavin–Stein method for variancegamma approximation on Wiener space. Electronic Journal of Probability, 20 (??):123:1–123:28, ???? 2015. CODEN ???? ISSN 1083-6489. URL http://

ejp.ejpecp.org/article/view/4136.

Etore:2006:RWS

[Eto06] Pierre Etoré. On random walk simulation of one-dimensional diffusion processes with discontinuous coefficients. Electronic Journal of Probability, 11: 9:249–9:275, 2006. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.

org/article/view/311.

Englander:2022:CRW

[EV22] János Engländer and Stanislav Volkov. Conservative random walk. Electronic Journal of Probability, 27(??):1–29, 2022. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-27/issue-none/Conservative-random-walk/10.1214/
22-EJP863.full.

Evans:2002:ERW

[Eva02] Steven Evans. Eigenvalues of random wreath products. Electronic Journal of Probability, 7:9:1–9:15, 2002. CODEN ???? ISSN 1083-6489. URL http://

ejp.ejpecp.org/article/view/108.

Englander:2020:CTW

[EVW20] János Engländer, Stanislav Volkov, and Zhenhua Wang. The coin-turning walk and its scaling limit. Electronic Journal of Probability, 25(??):3:1–3:38, ???? 2020. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

euclid.ejp/1578452592.

Etheridge:2020:RLS

[EVY20] Alison M. Etheridge, Amandine Véber, and Feng Yu. Rescaling limits of the spatial Lambda-Fleming-Viot process with selection. Electronic Journal of Probability, 25(??):1–89, 2020. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-25/issue-none/Rescaling-limits-of-the-spatial-
Lambda-Fleming-Viot-process-with/10.1214/20-EJP523.full.

Erhard:2022:WUD

[EX22] Dirk Erhard and Weijun Xu. Weak universality of dynamical Φ_{3}^{4}: polynomial potential and general smoothing mechanism. Electronic Journal of Probability, 27(??):1–43, 2022. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-27/issue-none/Weak-universality-of-dynamical-
%ce%a634--polynomial-potential-and-general/10.1214/22-EJP833.full.

Erdos:2012:CWD

[EY12] László Erdős and Horng-Tzer Yau. A comment on the Wigner–Dyson–Mehta bulk universality conjecture for Wigner matrices. Electronic Journal of Probability, 17:28:1–28:5, 2012. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/1779.

Englander:2016:BDP

[EZ16] János Engländer and Liang Zhang. Branching diffusion with particle interactions. Electronic Journal of Probability, 21(??):67:1–67:25, ???? 2016. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.ejp/
1480388424.

Ezanno:2013:SRA

[Eza13] François Ezanno. Some results about ergodicity in shape for a crystal growth model. Electronic Journal of Probability, 18:33:1–33:20, 2013. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/2177.

Faggionato:2008:RWE

[Fag08] Alessandra Faggionato. Random walks and exclusion processes among random conductances on random infinite clusters: homogenization and hydrodynamic limit. Electronic Journal of Probability, 13:73:2217–73:2247, 2008. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/591.

Faggionato:2010:HLZ

[Fag10] Alessandra Faggionato. Hydrodynamic limit of zero range processes among random conductances on the supercritical percolation cluster. Electronic Journal of Probability, 15:10:259–10:291, 2010. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/748.

Faggionato:2012:SAN

[Fag12] Alessandra Faggionato. Spectral analysis of 1D nearest-neighbor random walks and applications to subdiffusive trap and barrier models. Electronic Journal of Probability, 17:15:1–15:36, 2012. CODEN ???? ISSN 1083-6489. URL http://

ejp.ejpecp.org/article/view/1831.

Fang:2019:WBN

[Fan19] Xiao Fang. Wasserstein-2 bounds in normal approximation under local dependence. Electronic Journal of Probability, 24(??):35:1–35:14, ???? 2019. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.ejp/
1554775414.

Faraud:2011:CLT

[Far11] Gabriel Faraud. A Central Limit Theorem for random walk in a random environment on a marked Galton–Watson tree. Electronic Journal of Probability, 16: 6:174–6:215, 2011. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.

org/article/view/851.

Faure:2002:SNL

[Fau02] Mathieu Faure. Self-normalized large deviations for Markov chains. Electronic Journal of Probability, 7:23:1–23:31, 2002. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/122.

Feyel:1998:ASS

[Fd98] Denis Feyel and Arnaud de La Pradelle. On the approximate solutions of the Stratonovitch equation. Electronic Journal of Probability, 3:7:1–7:14, 1998. CO-

DEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/29.

Feyel:2000:ARP

[Fd00] D. Feyel and A. de La Pradelle. The abstract Riemannian path space. Electronic Journal of Probability, 5:11:1–11:17, 2000. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/67.

Feyel:2001:FIF

[Fd01] D. Feyel and A. de La Pradelle. The FBM Itô’s formula through analytic continuation. Electronic Journal of Probability, 6:26:1–26:22, 2001. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/99.

Feyel:2006:CIA

[Fd06] Denis Feyel and Arnaud de La Pradelle. Curvilinear integrals along enriched paths. Electronic Journal of Probability, 11:34:860–34:892, 2006. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/356.

Fehrman:2017:ELI

[Feh17] Benjamin Fehrman. Exit laws of isotropic diffusions in random environment from large domains. Electronic Journal of Probability, 22(??):63:1–63:37, ???? 2017. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

euclid.ejp/1502330523.

Feng:2007:LDD

[Fen07] Shui Feng. Large deviations for Dirichlet processes and Poisson–Dirichlet distribution with two parameters. Electronic Journal of Probability, 12:27:787–27:807, 2007. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/

view/417.

Feray:2013:ABS

[Fér13] Valentin Féray. Asymptotic behavior of some statistics in Ewens random permutations. Electronic Journal of Probability, 18:76:1–76:32, 2013. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/2496.

Ferrario:2016:CLS

[Fer16] Benedetta Ferrario. Characterization of the law for 3D stochastic hyperviscous fluids. Electronic Journal of Probability, 21(??):26:1–26:22, ???? 2016.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.
ejp/1459960920.

Feray:2018:WDG

[Fér18] Valentin Féray. Weighted dependency graphs. Electronic Journal of Probability, 23(??):93:1–93:65, ???? 2018. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1537257885.

Ferrari:1998:FSS

[FF98] P. A. Ferrari and L. R. G. Fontes. Fluctuations of a surface submitted to a random average process. Electronic Journal of Probability, 3:6:1–6:34, 1998.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
28.

Fathi:2013:TEI

[FF13] Max Fathi and Noufel Frikha. Transport-entropy inequalities and deviation estimates for stochastic approximation schemes. Electronic Journal of Probability, 18:67:1–67:36, 2013. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.

org/article/view/2586.

Favaro:2014:ANB

[FF14] Stefano Favaro and Shui Feng. Asymptotics for the number of blocks in a conditional Ewens–Pitman sampling model. Electronic Journal of Probability, 19:21:1–21:15, 2014. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.

org/article/view/2881.

Favaro:2015:LDP

[FF15] Stefano Favaro and Shui Feng. Large deviation principles for the Ewens–Pitman sampling model. Electronic Journal of Probability, 20(??):40:1–40:26, ???? 2015.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
3668.

Fedrizzi:2017:RSK

[FFPV17] Ennio Fedrizzi, Franco Flandoli, Enrico Priola, and Julien Vovelle. Regularity of stochastic kinetic equations. Electronic Journal of Probability, 22(??): 48:1–48:42, ???? 2017. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1496196076.

Fleischmann:1996:TSA

[FG96] Klaus Fleischmann and Andreas Greven. Time-space analysis of the clusterformation in interacting diffusions. Electronic Journal of Probability, 1:6:1–6:46, 1996. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/

view/6.

Fitzsimmons:2003:HRM

[FG03] Patrick Fitzsimmons and Ronald Getoor. Homogeneous random measures and strongly supermedian kernels of a Markov process. Electronic Journal of Probability, 8:10:1–10:54, 2003. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/142.

Flandoli:2005:SVF

[FG05] Franco Flandoli and Massimiliano Gubinelli. Statistics of a vortex filament model. Electronic Journal of Probability, 10:25:865–25:900, 2005. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/267.

Fontes:2019:ABA

[FG19] Luiz Renato Fontes and Véronique Gayrard. Asymptotic behavior and aging of a low temperature cascading 2-GREM dynamics at extreme time scales. Electronic Journal of Probability, 24(??):142:1–142:50, ???? 2019. CODEN ???? ISSN

1083-6489. URL https://projecteuclid.org/euclid.ejp/1576638111.

Ferrari:2020:BIM

[FG20] Pablo A. Ferrari and Davide Gabrielli. BBS invariant measures with independent soliton components. Electronic Journal of Probability, 25(??):78:1–78:26, ???? 2020. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.

org/euclid.ejp/1594432886.

Fan:2015:EIM

[FGL15] Xiequan Fan, Ion Grama, and Quansheng Liu. Exponential inequalities for martingales with applications. Electronic Journal of Probability, 20(??):1:1– 1:22, ???? 2015. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.

org/article/view/3496.

Fajfrova:2016:IMM

[FGS16] Lucie Fajfrová, Thierry Gobron, and Ellen Saada. Invariant measures of mass migration processes. Electronic Journal of Probability, 21(??):60:1–60:52, ???? 2016. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

euclid.ejp/1475266508.

Fill:2010:PSV

[FH10] James Fill and Mark Huber. Perfect simulation of Vervaat perpetuities. Electronic Journal of Probability, 15:4:96–4:109, 2010. CODEN ???? ISSN 1083-

6489. URL http://ejp.ejpecp.org/article/view/734.

Foucart:2013:SCS

[FH13] Clément Foucart and Olivier Hénard. Stable continuous-state branching processes with immigration and beta-Fleming–Viot processes with immigration.

Electronic Journal of Probability, 18:23:1–23:21, 2013. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/2024.

Fill:2019:QIR

[FH19] James Allen Fill and Wei-Chun Hung. QuickSort: improved right-tail asymptotics for the limiting distribution, and large deviations. Electronic Journal of Probability, 24(??):67:1–67:13, ???? 2019. CODEN ???? ISSN 1083-6489. URL

https://projecteuclid.org/euclid.ejp/1561687600.

Fill:2023:DFQ

[FH23] James Allen Fill and Wei-Chun Hung. Density functions for QuickQuant and QuickVal. Electronic Journal of Probability, 28(??):1–50, 2023. CODEN ????

ISSN 1083-6489. URL https://projecteuclid.org/journals/electronic-
journal-of-probability/volume-28/issue-none/Density-functions-for-
QuickQuant-and-QuickVal/10.1214/22-EJP899.full.

Fountoulakis:2022:CPP

[FI22] Nikolaos Fountoulakis and Tejas Iyer. Condensation phenomena in preferential attachment trees with neighbourhood influence. Electronic Journal of Probability, 27(??):1–49, 2022. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-27/issue-none/Condensation-phenomena-in-preferential-
attachment-trees-with-neighbourhood-influence/10.1214/22-EJP787.full.

Figueiredo:2021:RWG

[FIO^{+}21] Daniel Figueiredo, Giulio Iacobelli, Roberto Oliveira, Bruce Reed, and Rodrigo Ribeiro. On a random walk that grows its own tree. Electronic Journal of Probability, 26(??):1–40, 2021. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-26/issue-none/On-a-random-walk-that-grows-its-
own-tree/10.1214/20-EJP574.full.

Fromm:2015:FAS

[FIP15] Alexander Fromm, Peter Imkeller, and David J. Prömel. An FBSDE approach to the Skorokhod embedding problem for Gaussian processes with non-linear drift. Electronic Journal of Probability, 20(??):127:1–127:38, ???? 2015. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/3758.

Fitzsimmons:1998:MPI

[Fit98] P. J. Fitzsimmons. Markov processes with identical bridges. Electronic Journal of Probability, 3:12:1–12:12, 1998. CODEN ???? ISSN 1083-6489. URL http:

//ejp.ejpecp.org/article/view/34.

FitzGerald:2021:IMP

[Fit21] Will FitzGerald. The invariant measure of PushASEP with a wall and pointto-line last passage percolation. Electronic Journal of Probability, 26(18):1–26, February 2021. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.

org/journals/electronic-journal-of-probability/volume-26/issue-none/
The-invariant-measure-of-PushASEP-with-a-wall-and-point/10.1214/21-
EJP661.full.

Fill:2012:NBC

[FJ12] James Fill and Svante Janson. The number of bit comparisons used by Quicksort: an average-case analysis. Electronic Journal of Probability, 17:43:1–43:22, 2012. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/

view/1812.

Freeman:2020:ECM

[FJ20] Nic Freeman and Jonathan Jordan. Extensive condensation in a model of preferential attachment with fitness. Electronic Journal of Probability, 25(??): 68:1–68:42, ???? 2020. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1592964036.

Fill:2022:SPS

[FJ22] James Allen Fill and Svante Janson. The sum of powers of subtree sizes for conditioned Galton–Watson trees. Electronic Journal of Probability, 27(??):1–77, 2022. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

journals/electronic-journal-of-probability/volume-27/issue-none/The-
sum-of-powers-of-subtree-sizes-for-conditioned-GaltonWatson/10.1214/
22-EJP831.full.

Fill:2023:CSP

[FJ23] James Allen Fill and Svante Janson. Corrigendum to: The sum of powers of subtree sizes for conditioned Galton–Watson trees. Electronic Journal of Probability, 28(??):1–2, 2023. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-28/issue-none/Corrigendum-to--The-sum-of-powers-
of-subtree-sizes/10.1214/23-EJP915.full.

Fearnhead:2015:TDC

[FJS15] Paul Fearnhead, Paul Jenkins, and Yun Song. Tractable diffusion and coalescent processes for weakly correlated loci. Electronic Journal of Probability, 20(??): 58:1–58:25, ???? 2015. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/3564.

Fan:2011:RTG

[FJX11] ShengJun Fan, Long Jiang, and YingYing Xu. Representation theorem for generators of BSDEs with monotonic and polynomial-growth generators in the space of processes. Electronic Journal of Probability, 16:27:830–27:844, 2011.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
873.

Fannjiang:2002:DLR

[FK02] Albert Fannjiang and Tomasz Komorowski. Diffusion in long-range correlated Ornstein–Uhlenbeck flows. Electronic Journal of Probability, 7:20:1–20:22, 2002.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
119.

Foondun:2009:INP

[FK09] Mohammud Foondun and Davar Khoshnevisan. Intermittence and nonlinear parabolic stochastic partial differential equations. Electronic Journal of Probability, 14:21:548–21:568, 2009. CODEN ???? ISSN 1083-6489. URL http://

ejp.ejpecp.org/article/view/614.

Freidlin:2021:ACM

[FK21] M. Freidlin and L. Koralov. Averaging in the case of multiple invariant measures for the fast system. Electronic Journal of Probability, 26(18):1–17, February 2021. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

journals/electronic-journal-of-probability/volume-26/issue-none/Averaging-
in-the-case-of-multiple-invariant-measures-for-the/10.1214/21-EJP681.
full.

Frikha:2019:IPF

[FKHL19] Noufel Frikha, Arturo Kohatsu-Higa, and Libo Li. Integration by parts formula for killed processes: a point of view from approximation theory. Electronic Journal of Probability, 24(??):95:1–95:44, ???? 2019. CODEN ???? ISSN

1083-6489. URL https://projecteuclid.org/euclid.ejp/1568793788.

Fleermann:2022:LSL

[FKK22] Michael Fleermann, Werner Kirsch, and Thomas Kriecherbauer. Local semicircle law for Curie–Weiss type ensembles. Electronic Journal of Probability, 27(??):1–27, 2022. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-27/issue-none/Local-semicircle-law-for-Curie-
Weiss-type-ensembles/10.1214/22-EJP767.full.

Friesen:2013:PTL

[FL13] Olga Friesen and Matthias Löwe. A phase transition for the limiting spectral density of random matrices. Electronic Journal of Probability, 18:17:1–17:17, 2013. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/

view/2118.

Field:2015:EPT

[FL15] Laurence S. Field and Gregory F. Lawler. Escape probability and transience for SLE. Electronic Journal of Probability, 20(??):10:1–10:14, ???? 2015. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/3714.

Felipe:2018:BPS

[FL18] Miraine Dávila Felipe and Amaury Lambert. Branching processes seen from their extinction time via path decompositions of reflected Lévy processes. Electronic Journal of Probability, 23(??):98:1–98:30, ???? 2018. CODEN ???? ISSN

1083-6489. URL https://projecteuclid.org/euclid.ejp/1537841130.

Flegel:2018:LPD

[Fle18] Franziska Flegel. Localization of the principal Dirichlet eigenvector in the heavy-tailed random conductance model. Electronic Journal of Probability, 23 (??):68:1–68:43, ???? 2018. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1532570596.

Fradelizi:2020:CIC

[FLM20] Matthieu Fradelizi, Jiange Li, and Mokshay Madiman. Concentration of information content for convex measures. Electronic Journal of Probability, 25 (??):20:1–20:22, ???? 2020. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1580979618.

Fleischmann:2003:CSS

[FM03] Klaus Fleischmann and Leonid Mytnik. Competing species superprocesses with infinite variance. Electronic Journal of Probability, 8:8:1–8:59, 2003. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/136.

Fernandez:2004:CCC

[FM04] Roberto Fernandez and Gregory Maillard. Chains with complete connections and one-dimensional Gibbs measures. Electronic Journal of Probability, 9:6:145– 6:176, 2004. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/149.

Ferrari:2007:QSD

[FM07] Pablo Ferrari and Nevena Maric. Quasi stationary distributions and Fleming– Viot processes in countable spaces. Electronic Journal of Probability, 12:24:684– 24:702, 2007. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/415.

Fiodorov:2014:CLE

[FM14] Artiom Fiodorov and Stephen Muirhead. Complete localisation and exponential shape of the parabolic Anderson model with Weibull potential field. Electronic Journal of Probability, 19:58:1–58:27, 2014. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/3203.

Furlan:2017:TCR

[FM17] Marco Furlan and Jean-Christophe Mourrat. A tightness criterion for random fields, with application to the Ising model. Electronic Journal of Probability, 22(??):97:1–97:29, ???? 2017. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1510802251.

Fonseca-Mora:2020:SDN

[FM20] Christian A. Fonseca-Mora. Semimartingales on duals of nuclear spaces. Electronic Journal of Probability, 25(??):36:1–36:24, ???? 2020. CODEN ???? ISSN

1083-6489. URL https://projecteuclid.org/euclid.ejp/1585188065.

Fonseca-Mora:2021:SIR

[FM21] Christian A. Fonseca-Mora. Stochastic integration with respect to cylindrical semimartingales. Electronic Journal of Probability, 26(18):1–48, February 2021.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-26/issue-none/Stochastic-
integration-with-respect-to-cylindrical-semimartingales/10.1214/21-
EJP718.full.

Fontbona:2022:QMF

[FMH22] Joaquín Fontbona and Felipe Muñoz-Hernández. Quantitative mean-field limit for interacting branching diffusions. Electronic Journal of Probability, 27(??):1– 32, 2022. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

journals/electronic-journal-of-probability/volume-27/issue-none/Quantitative-
mean-field-limit-for-interacting-branching-diffusions/10.1214/22-
EJP874.full.

Foucart:2019:CCS

[FMM19] Clément Foucart, Chunhua Ma, and Bastien Mallein. Coalescences in continuous-state branching processes. Electronic Journal of Probability, 24(??): 103:1–103:52, ???? 2019. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1569895472.

Fernandez:2015:AEH

[FMNS15] Roberto Fernandez, Francesco Manzo, Francesca Romana Nardi, and Elisabetta Scoppola. Asymptotically exponential hitting times and metastability: a pathwise approach without reversibility. Electronic Journal of Probability, 20(??): 122:1–122:37, ???? 2015. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/3656.

Foxall:2022:FTR

[FMR22] Eric Foxall, Bilal Madani, and Adam Roemer. Fixation time of the rockpaper-scissors model: rigorous results in the well-mixed setting. Electronic Journal of Probability, 27(??):1–23, 2022. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-27/issue-none/Fixation-time-of-the-rock-paper-
scissors-model--rigorous/10.1214/22-EJP807.full.

Fleischmann:2006:HLF

[FMW06] Klaus Fleischmann, Peter Mörters, and Vitali Wachtel. Hydrodynamic limit fluctuations of super-Brownian motion with a stable catalyst. Electronic Journal of Probability, 11:29:723–29:767, 2006. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/348.

Fill:2020:PRF

[FN20] James Allen Fill and Daniel Q. Naiman. The Pareto record frontier. Electronic Journal of Probability, 25(??):1–24, 2020. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-25/issue-none/The-Pareto-record-frontier/10.
1214/20-EJP492.full.

Ferrari:2018:UGT

[FO18] Patrik L. Ferrari and Alessandra Occelli. Universality of the GOE Tracy–Widom distribution for TASEP with arbitrary particle density. Electronic Journal of Probability, 23(??):51:1–51:24, ???? 2018. CODEN ???? ISSN 1083-6489. URL

https://projecteuclid.org/euclid.ejp/1527818429.

Folz:2011:GUB

[Fol11] Matthew Folz. Gaussian upper bounds for heat kernels of continuous time simple random walks. Electronic Journal of Probability, 16:62:1693–62:1722, 2011. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/

view/926.

Foondun:2009:HKE

[Foo09] Mohammud Foondun. Heat kernel estimates and Harnack inequalities for some Dirichlet forms with non-local part. Electronic Journal of Probability, 14:11:314– 11:340, 2009. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/604.

Forsstrom:2016:MPE

[For16] Malin Palö Forsström. Monotonicity properties of exclusion sensitivity. Electronic Journal of Probability, 21(??):45:1–45:22, ???? 2016. CODEN ???? ISSN

1083-6489. URL https://projecteuclid.org/euclid.ejp/1469557138.

Forien:2019:SSM

[For19] Raphaël Forien. The stepping stone model in a random environment and the effect of local heterogeneities on isolation by distance patterns. Electronic Journal of Probability, 24(??):57:1–57:35, ???? 2019. CODEN ???? ISSN 1083-6489.

URL https://projecteuclid.org/euclid.ejp/1560391565.

Forman:2020:EHM

[For20] Noah Forman. Exchangeable hierarchies and mass-structure of weighted real trees. Electronic Journal of Probability, 25(??):1–28, 2020. CODEN ????

ISSN 1083-6489. URL https://projecteuclid.org/journals/electronic-
journal-of-probability/volume-25/issue-none/Exchangeable-hierarchies-
and-mass-structure-of-weighted-real-trees/10.1214/20-EJP522.full.

Forien:2022:SPD

[For22] Raphaël Forien. Stochastic partial differential equations describing neutral genetic diversity under short range and long range dispersal. Electronic Journal of Probability, 27(??):1–41, 2022. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-27/issue-none/Stochastic-partial-differential-
equations-describing-neutral-genetic-diversity-under-short/10.1214/
22-EJP827.full.

Fournier:2008:SLS

[Fou08] Nicolas Fournier. Smoothness of the law of some one-dimensional jumping s.d.e.s with non-constant rate of jump. Electronic Journal of Probability, 13:6:135– 6:156, 2008. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/480.

Foucart:2019:CSB

[Fou19] Clément Foucart. Continuous-state branching processes with competition: duality and reflection at infinity. Electronic Journal of Probability, 24(??):33:1–33:38, ???? 2019. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.

org/euclid.ejp/1554775412.

Filmus:2022:LSI

[FOW22] Yuval Filmus, Ryan O’Donnell, and Xinyu Wu. Log-Sobolev inequality for the multislice, with applications. Electronic Journal of Probability, 27(??):1–30, 2022. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

journals/electronic-journal-of-probability/volume-27/issue-none/Log-
Sobolev-inequality-for-the-multislice-with-applications/10.1214/22-
EJP749.full.

Foxall:2018:NGC

[Fox18] Eric Foxall. The naming game on the complete graph. Electronic Journal of Probability, 23(??):126:1–126:42, ???? 2018. CODEN ???? ISSN 1083-6489.

URL https://projecteuclid.org/euclid.ejp/1545188693.

Forman:2020:DSI

[FPRW20] Noah Forman, Soumik Pal, Douglas Rizzolo, and Matthias Winkel. Diffusions on a space of interval partitions: construction from marked Lévy processes. Electronic Journal of Probability, 25(??):1–46, 2020. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-25/issue-none/Diffusions-on-a-space-of-interval-
partitions--construction-from/10.1214/20-EJP521.full.

Flint:2019:FIM

[FPT19] Ian Flint, Nicolas Privault, and Giovanni Luca Torrisi. Functional inequalities for marked point processes. Electronic Journal of Probability, 24(??): 116:1–116:40, ???? 2019. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1570759241.

Flandoli:2001:SSS

[FR01] Franco Flandoli and Marco Romito. Statistically stationary solutions to the 3D Navier–Stokes equations do not show singularities. Electronic Journal of Probability, 6:5:1–5:15, 2001. CODEN ???? ISSN 1083-6489. URL http://

ejp.ejpecp.org/article/view/78.

Faller:2012:ASB

[FR12] Andreas Faller and Ludger Rüschendorf. Approximative solutions of best choice problems. Electronic Journal of Probability, 17:54:1–54:22, 2012. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/2172.

Falgas-Ravry:2013:DCN

[FR13] Victor Falgas-Ravry. Distribution of components in the k-nearest neighbour random geometric graph for k below the connectivity threshold. Electronic Journal of Probability, 18:83:1–83:22, 2013. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/2465.

Fitzsimmons:2014:MLS

[FR14] Patrick Fitzsimmons and Jay Rosen. Markovian loop soups: permanental processes and isomorphism theorems. Electronic Journal of Probability, 19:60:1– 60:30, 2014. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/3255.

Fan:2018:NSC

[FR18] Wai-Tong (Louis) Fan and Sebastien Roch. Necessary and sufficient conditions for consistent root reconstruction in Markov models on trees. Electronic Journal of Probability, 23(??):47:1–47:24, ???? 2018. CODEN ???? ISSN 1083-6489.

URL https://projecteuclid.org/euclid.ejp/1527213728.

Fradelizi:2009:CIC

[Fra09] Matthieu Fradelizi. Concentration inequalities for s-concave measures of dilations of Borel sets and applications. Electronic Journal of Probability, 14: 71:2068–71:2090, 2009. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/695.

Fradon:2010:BDG

[Fra10] Myriam Fradon. Brownian dynamics of globules. Electronic Journal of Probability, 15:6:142–6:161, 2010. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/739.

Foutel-Rodier:2020:KCE

[FRLS20] Félix Foutel-Rodier, Amaury Lambert, and Emmanuel Schertzer. Kingman’s coalescent with erosion. Electronic Journal of Probability, 25(??):56:1–56:33, ???? 2020. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.

org/euclid.ejp/1588644039.

Fill:2001:MTM

[FS01] James Fill and Clyde Schoolfield, Jr. Mixing times for Markov chains on wreath products and related homogeneous spaces. Electronic Journal of Probability, 6: 11:1–11:22, 2001. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.

org/article/view/84.

Feldheim:2017:DRR

[FS17] Ohad N. Feldheim and Arnab Sen. Double roots of random polynomials with integer coefficients. Electronic Journal of Probability, 22(??):10:1–10:23, ???? 2017. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

euclid.ejp/1486090890.

Ferre:2020:LDE

[FS20a] Grégoire Ferré and Gabriel Stoltz. Large deviations of empirical measures of diffusions in weighted topologies. Electronic Journal of Probability, 25(??):1–52, 2020. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

journals/electronic-journal-of-probability/volume-25/issue-none/Large-
deviations-of-empirical-measures-of-diffusions-in-weighted-topologies/
10.1214/20-EJP514.full.

Forsstrom:2020:DCR

[FS20b] Malin P. Forsström and Jeffrey E. Steif. Divide and color representations for threshold Gaussian and stable vectors. Electronic Journal of Probability, 25 (??):54:1–54:45, ???? 2020. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1588644037.

Flores:2014:FED

[FSV14] Gregorio R. Moreno Flores, Timo Seppäläinen, and Benedek Valkó. Fluctuation exponents for directed polymers in the intermediate disorder regime. Electronic Journal of Probability, 19(??):89:1–89:28, ???? 2014. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/3307.

Ferrari:2015:BMO

[FSW15] Patrik Lino Ferrari, Herbert Spohn, and Thomas Weiss. Brownian motions with one-sided collisions: the stationary case. Electronic Journal of Probability, 20 (??):69:1–69:41, ???? 2015. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/4177.

Fatkullin:2020:HLY

[FSX20] Ibrahim Fatkullin, Sunder Sethuraman, and Jianfei Xue. On hydrodynamic limits of Young diagrams. Electronic Journal of Probability, 25(??):58:1–58:44, ???? 2020. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.

org/euclid.ejp/1588644041.

Fisher:2011:SSD

[FT11] Albert Fisher and Marina Talet. The self-similar dynamics of renewal processes. Electronic Journal of Probability, 16:31:929–31:961, 2011. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/888.

Filipovic:2014:IMB

[FTT14] Damir Filipović, Stefan Tappe, and Josef Teichmann. Invariant manifolds with boundary for jump-diffusions. Electronic Journal of Probability, 19:51:1–51:28, 2014. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/

view/2882.

FitzGerald:2020:SAF

[FTZ20] Will FitzGerald, Roger Tribe, and Oleg Zaboronski. Sharp asymptotics for Fredholm Pfaffians related to interacting particle systems and random matrices. Electronic Journal of Probability, 25(??):1–15, 2020. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-25/issue-none/Sharp-asymptotics-for-Fredholm-
Pfaffians-related-tointeracting-particle-systems-and/10.1214/20-EJP512.
full.

Fukushima:1999:SMC

[Fuk99] Masatoshi Fukushima. On semi-martingale characterizations of functionals of symmetric Markov processes. Electronic Journal of Probability, 4:18:1–18:32, 1999. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/

view/55.

Fukasawa:2011:AAS

[Fuk11] Masaaki Fukasawa. Asymptotic analysis for stochastic volatility: Edgeworth expansion. Electronic Journal of Probability, 16:25:764–25:791, 2011. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/879.

Fulman:2005:SMD

[Ful05] Jason Fulman. Stein’s method and descents after riffle shuffles. Electronic Journal of Probability, 10:26:901–26:924, 2005. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/268.

Fulman:2012:SMH

[Ful12] Jason Fulman. Stein’s method, heat kernel, and traces of powers of elements of compact Lie groups. Electronic Journal of Probability, 17:66:1–66:16, 2012.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
2251.

Friz:2011:NHD

[FV11] Peter Friz and Nicolas Victoir. A note on higher dimensional p-variation. Electronic Journal of Probability, 16:68:1880–68:1899, 2011. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/951.

Ferrari:2012:NCB

[FV12] Patrik Ferrari and Bálint Vető. Non-colliding Brownian bridges and the asymmetric tacnode process. Electronic Journal of Probability, 17:44:1–44:17, 2012.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/

Ferrari:2017:HET

[FV17] Patrik L. Ferrari and Bálint Vető. The hard-edge tacnode process for Brownian motion. Electronic Journal of Probability, 22(??):79:1–79:32, ???? 2017.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.
ejp/1506931229.

Fountoulakis:2021:CHM

[FvdHMS21] Nikolaos Fountoulakis, Pim van der Hoorn, Tobias Müller, and Markus Schepers. Clustering in a hyperbolic model of complex networks. Electronic Journal of Probability, 26(??):1–132, 2021. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-26/issue-none/Clustering-in-a-hyperbolic-model-
of-complex-networks/10.1214/21-EJP583.full.

Fill:2008:TPK

[FW08] James Fill and David Wilson. Two-player knock ’em down. Electronic Journal of Probability, 13:9:198–9:212, 2008. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/485.

Forrester:2017:MBE

[FW17] Peter J. Forrester and Dong Wang. Muttalib–Borodin ensembles in random matrix theory — realisations and correlation functions. Electronic Journal of Probability, 22(??):54:1–54:43, ???? 2017. CODEN ???? ISSN 1083-6489. URL

https://projecteuclid.org/euclid.ejp/1498183245.

Fountoulakis:2020:LTI

[FY20] Nikolaos Fountoulakis and Joseph Yukich. Limit theory for isolated and extreme points in hyperbolic random geometric graphs. Electronic Journal of Probability, 25(??):1–51, 2020. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-25/issue-none/Limit-theory-for-isolated-and-
extreme-points-in-hyperbolic-random/10.1214/20-EJP531.full.

Feng:2007:GIF

[FZ07] Chunrong Feng and Huaizhong Zhao. A generalized Itô’s formula in twodimensions and stochastic Lebesgue–Stieltjes integrals. Electronic Journal of Probability, 12:57:1568–57:1599, 2007. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/468.

Feng:2010:LTR

[FZ10] Chunrong Feng and Huaizhong Zhao. Local time rough path for Lévy processes. Electronic Journal of Probability, 15:16:452–16:483, 2010. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/770.

Fang:2012:BRW

[FZ12] Ming Fang and Ofer Zeitouni. Branching random walks in time inhomogeneous environments. Electronic Journal of Probability, 17:67:1–67:18, 2012. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/2253.

Fakhry:2023:EMP

[FZ23] Rami Fakhry and Dapeng Zhan. Existence of multi-point boundary Green’s function for chordal Schramm–Loewner evolution (SLE). Electronic Journal of Probability, 28(??):1–29, 2023. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-28/issue-none/Existence-of-multi-point-boundary-
Greens-function-for-chordal-Schramm/10.1214/23-EJP936.full.

Gayrard:2008:EPT

[GA08] Véronique Gayrard and Gérard Ben Arous. Elementary potential theory on the hypercube. Electronic Journal of Probability, 13:59:1726–59:1807, 2008.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
527.

Galvin:2007:SCR

[Gal07] David Galvin. Sampling 3-colourings of regular bipartite graphs. Electronic Journal of Probability, 12:16:481–16:497, 2007. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/403.

Gall:2022:VMB

[Gal22] Jean-François Le Gall. The volume measure of the Brownian sphere is a Hausdorff measure. Electronic Journal of Probability, 27(??):1–28, 2022. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-27/issue-none/The-volume-
measure-of-the-Brownian-sphere-is-a-Hausdorff/10.1214/22-EJP837.full.

Ganapathy:2007:RM

[Gan07] Murali Ganapathy. Robust mixing. Electronic Journal of Probability, 12:10:262– 10:299, 2007. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/398.

Ganguly:2013:WZT

[Gan13] Arnab Ganguly. Wong–Zakai type convergence in infinite dimensions. Electronic Journal of Probability, 18:31:1–31:34, 2013. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/2650.

Gangopadhyay:2022:FTI

[Gan22] Ujan Gangopadhyay. Fluctuations of transverse increments in two-dimensional first passage percolation. Electronic Journal of Probability, 27(??):1–61, 2022.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-27/issue-none/Fluctuations-
of-transverse-increments-in-two-dimensional-first-passage-percolation/
10.1214/22-EJP772.full.

Gao:2017:LTA

[Gao17] Fuqing Gao. Long time asymptotics of unbounded additive functionals of Markov processes. Electronic Journal of Probability, 22(??):94:1–94:21, ???? 2017. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

euclid.ejp/1509501717.

Gapeev:2006:DOS

[Gap06] Pavel Gapeev. Discounted optimal stopping for maxima in diffusion models with finite horizon. Electronic Journal of Probability, 11:38:1031–38:1048, 2006.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
367.

Garet:2004:PTS

[Gar04] Olivier Garet. Percolation transition for some excursion sets. Electronic Journal of Probability, 9:10:255–10:292, 2004. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/196.

Gautier:2007:SNS

[Gau07] Eric Gautier. Stochastic nonlinear Schrödinger equations driven by a fractional noise. Well-posedness, large deviations and support. Electronic Journal of Probability, 12:29:848–29:861, 2007. CODEN ???? ISSN 1083-6489. URL http://

ejp.ejpecp.org/article/view/416.

Gaunt:2014:VGA

[Gau14] Robert Gaunt. Variance-Gamma approximation via Stein’s method. Electronic Journal of Probability, 19:38:1–38:33, 2014. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/3020.

Gayrard:2012:CCP

[Gay12] Véronique Gayrard. Convergence of clock process in random environments and aging in Bouchaud’s asymmetric trap model on the complete graph. Electronic Journal of Probability, 17:58:1–58:33, 2012. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/2211.

Gu:2014:IPB

[GB14] Yu Gu and Guillaume Bal. An invariance principle for Brownian motion in random scenery. Electronic Journal of Probability, 19:1:1–1:19, 2014. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/2894.

Guillin:2022:CRV

[GBM22] Arnaud Guillin, Pierre Le Bris, and Pierre Monmarché. Convergence rates for the Vlasov–Fokker–Planck equation and uniform in time propagation of chaos in non convex cases. Electronic Journal of Probability, 27(??):1–44, 2022.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-27/issue-none/Convergence-
rates-for-the-Vlasov-Fokker--Planck-equation-and-uniform/10.1214/
22-EJP853.full.

Gartner:2007:ICS

[GdHM07] Jürgen Gärtner, Frank den Hollander, and Gregory Maillard. Intermittency on catalysts: symmetric exclusion. Electronic Journal of Probability, 12:18:516– 18:573, 2007. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/407.

Gartner:2009:ICT

[GdHM09] Jürgen Gärtner, Frank den Hollander, and Grégory Maillard. Intermittency on catalysts: three-dimensional simple symmetric exclusion. Electronic Journal of Probability, 14:72:2091–72:2129, 2009. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/694.

Greven:2022:SPS

[GdHO22] Andreas Greven, Frank den Hollander, and Margriet Oomen. Spatial populations with seed-bank: well-posedness, duality and equilibrium. Electronic Journal of Probability, 27(??):1–88, 2022. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-27/issue-none/Spatial-populations-with-seed-
bank--well-posedness-duality-and/10.1214/21-EJP728.full.

Ged:2019:PSS

[Ged19] François Gaston Ged. Profile of a self-similar growth-fragmentation. Electronic Journal of Probability, 24(??):7:1–7:21, ???? 2019. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/euclid.ejp/1550199785.

Geiger:2000:PPP

[Gei00] Jochen Geiger. Poisson point process limits in size-biased Galton–Watson trees. Electronic Journal of Probability, 5:17:1–17:12, 2000. CODEN ???? ISSN 1083-

6489. URL http://ejp.ejpecp.org/article/view/73.

Gerard:2020:RVR

[Ger20] Thomas Gerard. Representations of the Vertex Reinforced Jump Process as a mixture of Markov processes on Z^{d} and infinite trees. Electronic Journal of Probability, 25(??):1–45, 2020. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-25/issue-none/Representations-of-the-Vertex-
Reinforced-Jump-Process-as-a-mixture/10.1214/20-EJP510.full.

Getoor:1999:EGS

[Get99] Ronald K. Getoor. An extended generator and Schrödinger equations. Electronic Journal of Probability, 4:19:1–19:23, 1999. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/56.

Glode:2019:BTC

[GGR19] Patric Glöde, Andreas Greven, and Thomas Rippl. Branching trees I: concatenation and infinite divisibility. Electronic Journal of Probability, 24(??):52:1–52:55, ???? 2019. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.

org/euclid.ejp/1559354444.

Gantert:2021:TGW

[GGS21] Nina Gantert, Nicos Georgiou, and Dominik Schmid. The TASEP on Galton– Watson trees. Electronic Journal of Probability, 26(18):1–38, February 2021.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-26/issue-none/The-TASEP-
on-GaltonWatson-trees/10.1214/21-EJP725.full.

Galanis:2018:USA

[GGY18] Andreas Galanis, Leslie Ann Goldberg, and Kuan Yang. Uniqueness for the 3-state antiferromagnetic Potts model on the tree. Electronic Journal of Probability, 23(??):82:1–82:43, ???? 2018. CODEN ???? ISSN 1083-6489. URL

https://projecteuclid.org/euclid.ejp/1536717741.

Gravner:2009:LBP

[GH09] Janko Gravner and Alexander Holroyd. Local bootstrap percolation. Electronic Journal of Probability, 14:14:385–14:399, 2009. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/607.

Grimmett:2010:PSE

[GH10] Geoffrey Grimmett and Alexander Holroyd. Plaquettes, spheres, and entanglement. Electronic Journal of Probability, 15:45:1415–45:1428, 2010. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/804.

Goldschmidt:2015:LBC

[GH15] Christina Goldschmidt and Bénédicte Haas. A line-breaking construction of the stable trees. Electronic Journal of Probability, 20(??):16:1–16:24, ???? 2015.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
3690.

Gerasimovics:2019:HTS

[GH19] Andris Gerasimovičs and Martin Hairer. Hörmander’s theorem for semilinear SPDEs. Electronic Journal of Probability, 24(??):132:1–132:56, ???? 2019. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.ejp/
1573614084.

Gao:2003:LTH

[GHLT03] Fuchang Gao, Jan Hannig, Tzong-Yow Lee, and Fred Torcaso. Laplace transforms via Hadamard factorization. Electronic Journal of Probability, 8:13:1– 13:20, 2003. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/151.

Galeati:2022:DDS

[GHM22] Lucio Galeati, Fabian A. Harang, and Avi Mayorcas. Distribution dependent SDEs driven by additive continuous noise. Electronic Journal of Probability, 27(??):1–38, 2022. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-27/issue-none/Distribution-dependent-SDEs-driven-
by-additive-continuous-noise/10.1214/22-EJP756.full.

Gracar:2022:RVT

[GHMM22] Peter Gracar, Markus Heydenreich, Christian Mönch, and Peter Mörters. Recurrence versus transience for weight-dependent random connection models. Electronic Journal of Probability, 27(??):1–31, 2022. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-27/issue-none/Recurrence-versus-transience-
for-weight-dependent-random-connection-models/10.1214/22-EJP748.full.

Gwynne:2021:JSL

[GHS21] Ewain Gwynne, Nina Holden, and Xin Sun. Joint scaling limit of site percolation on random triangulations in the metric and peanosphere sense. Electronic Journal of Probability, 26(18):1–58, February 2021. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-26/issue-none/Joint-scaling-limit-of-site-percolation-
on-random-triangulations-in/10.1214/21-EJP659.full.

Gao:2003:CTS

[GHT03] Fuchang Gao, Jan Hannig, and Fred Torcaso. Comparison theorems for small deviations of random series. Electronic Journal of Probability, 8:21:1–21:17, 2003. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/

view/147.

Gnedin:2012:RCC

[GI12] Alexander Gnedin and Alexander Iksanov. Regenerative compositions in the case of slow variation: A renewal theory approach. Electronic Journal of Probability, 17:77:1–77:19, 2012. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/2002.

Giacomin:2008:RCR

[Gia08] Giambattista Giacomin. Renewal convergence rates and correlation decay for homogeneous pinning models. Electronic Journal of Probability, 13:18:513– 18:529, 2008. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.

org/article/view/497.

Gibson:2008:MSV

[Gib08] Lee Gibson. The mass of sites visited by a random walk on an infinite graph. Electronic Journal of Probability, 13:44:1257–44:1282, 2008. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/531.

Gilch:2011:AER

[Gil11] Lorenz Gilch. Asymptotic entropy of random walks on free products. Electronic Journal of Probability, 16:3:76–3:105, 2011. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/841.

Gilch:2016:AER

[Gil16] Lorenz A. Gilch. Asymptotic entropy of random walks on regular languages over a finite alphabet. Electronic Journal of Probability, 21(??):8:1–8:42, ???? 2016.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.
ejp/1454682889.

Gubinelli:2016:FAA

[GIP16] Massimiliano Gubinelli, Peter Imkeller, and Nicolas Perkowski. A Fourier analytic approach to pathwise stochastic integration. Electronic Journal of Probability, 21(??):2:1–2:37, ???? 2016. CODEN ???? ISSN 1083-6489. URL

https://projecteuclid.org/euclid.ejp/1454514662.

Gamlin:2014:ABB

[GJ14] Samuel L. Gamlin and Antal A. Járai. Anchored burning bijections on finite and infinite graphs. Electronic Journal of Probability, 19(??):117:1–117:23, ????

2014. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/

view/3542.

Gantert:2017:BRW

[GJ17] Nina Gantert and Stefan Junk. A branching random walk among disasters. Electronic Journal of Probability, 22(??):67:1–67:34, ???? 2017. CODEN ???? ISSN

1083-6489. URL https://projecteuclid.org/euclid.ejp/1504922530.

Gapeev:2021:PME

[GJW21] Pavel V. Gapeev, Monique Jeanblanc, and Dongli Wu. Projections of martingales in enlargements of Brownian filtrations under Jacod’s equivalence hypothesis. Electronic Journal of Probability, 26(18):1–24, February 2021. CODEN ????

ISSN 1083-6489. URL https://projecteuclid.org/journals/electronic-
journal-of-probability/volume-26/issue-none/Projections-of-martingales-
in-enlargements-of-Brownian-filtrations-under-Jacods/10.1214/21-EJP694.
full.

Goldsheid:2000:ECA

[GK00] Ilya Goldsheid and Boris Khoruzhenko. Eigenvalue curves of asymmetric tridiagonal matrices. Electronic Journal of Probability, 5:16:1–16:28, 2000. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/72.

Grigorescu:2006:TPL

[GK06] Ilie Grigorescu and Min Kang. Tagged particle limit for a Fleming–Viot type system. Electronic Journal of Probability, 11:12:311–12:331, 2006. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/316.

Grigorescu:2007:EPM

[GK07] Ilie Grigorescu and Min Kang. Ergodic properties of multidimensional Brownian motion with rebirth. Electronic Journal of Probability, 12:48:1299–48:1322, 2007. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/

view/450.

Gupta:2014:SAS

[GK14] Ankit Gupta and Mustafa Khammash. Sensitivity analysis for stochastic chemical reaction networks with multiple time-scales. Electronic Journal of Probability, 19:59:1–59:53, 2014. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/3246.

Georgiou:2018:DRT

[GKKR18] Nicos Georgiou, Davar Khoshnevisan, Kunwoo Kim, and Alex D. Ramos. The dimension of the range of a transient random walk. Electronic Journal of Probability, 23(??):83:1–83:31, ???? 2018. CODEN ???? ISSN 1083-6489. URL

https://projecteuclid.org/euclid.ejp/1536717742.

Gusakova:2022:TDT

[GKT22] Anna Gusakova, Zakhar Kabluchko, and Christoph Thäle. The - delaunay tessellation II: the Gaussian limit tessellation. Electronic Journal of Probability, 27(??):1–33, 2022. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-27/issue-none/The-%ce%b2-Delaunay-tessellation-
II-the-Gaussian-limit-tessellation/10.1214/22-EJP782.full.

Greven:1999:LBB

[GKW99] Andreas Greven, Achim Klenke, and Anton Wakolbinger. The longtime behavior of branching random walk in a catalytic medium. Electronic Journal of Probability, 4:12:1–12:80, 1999. CODEN ???? ISSN 1083-6489. URL http://

ejp.ejpecp.org/article/view/49.

Gheissari:2018:EBC

[GL18] Reza Gheissari and Eyal Lubetzky. The effect of boundary conditions on mixing of 2D Potts models at discontinuous phase transitions. Electronic Journal of Probability, 23(??):57:1–57:30, ???? 2018. CODEN ???? ISSN 1083-6489. URL

https://projecteuclid.org/euclid.ejp/1528704075.

Gravner:2022:ODC

[GL22a] Janko Gravner and Xiaochen Liu. One-dimensional cellular automata with random rules: longest temporal period of a periodic solution. Electronic Journal of Probability, 27(??):1–23, 2022. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-27/issue-none/One-dimensional-cellular-automata-
with-random-rules--longest-temporal/10.1214/22-EJP744.full.

Grimmett:2022:BSR

[GL22b] Geoffrey R. Grimmett and Zhongyang Li. Brownian snails with removal: epidemics in diffusing populations. Electronic Journal of Probability, 27(??):1–31, 2022. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

journals/electronic-journal-of-probability/volume-27/issue-none/Brownian-
snails-with-removal-epidemics-in-diffusing-populations/10.1214/22-
EJP804.full.

Galeati:2023:SES

[GL23] Lucio Galeati and Chengcheng Ling. Stability estimates for singular SDEs and applications. Electronic Journal of Probability, 28(??):1–31, 2023. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-28/issue-none/Stability-
estimates-for-singular-SDEs-and-applications/10.1214/23-EJP913.full.

Gomez:2017:UBP

[GLM^{+}17a] Alejandro Gomez, Jong Jun Lee, Carl Mueller, Eyal Neuman, and Michael Salins. On uniqueness and blowup properties for a class of second order SDEs. Electronic Journal of Probability, 22(??):72:1–72:17, ???? 2017. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.ejp/
1505268103.

Grama:2017:HML

[GLM17b] Ion Grama, Quansheng Liu, and Eric Miqueu. Harmonic moments and large deviations for a supercritical branching process in a random environment. Electronic Journal of Probability, 22(??):99:1–99:23, ???? 2017. CODEN ???? ISSN

1083-6489. URL https://projecteuclid.org/euclid.ejp/1510802253.

Grahovac:2021:IIV

[GLT21] Danijel Grahovac, Nikolai N. Leonenko, and Murad S. Taqqu. Intermittency and infinite variance: the case of integrated supOU processes. Electronic Journal of Probability, 26(18):1–31, February 2021. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-26/issue-none/Intermittency-and-infinite-variance-
-the-case-of-integrated-supOU/10.1214/21-EJP623.full.

Greven:2005:RTI

[GLW05] Andreas Greven, Vlada Limic, and Anita Winter. Representation theorems for interacting Moran models, interacting Fisher–Wrighter diffusions and applications. Electronic Journal of Probability, 10:39:1286–39:1358, 2005. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/290.

Goldschmidt:2005:RRT

[GM05] Christina Goldschmidt and James Martin. Random recursive trees and the Bolthausen–Sznitman coalesent. Electronic Journal of Probability, 10:21:718– 21:745, 2005. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/265.

Garet:2008:FPC

[GM08] Olivier Garet and Régine Marchand. First-passage competition with different speeds: positive density for both species is impossible. Electronic Journal of Probability, 13:70:2118–70:2159, 2008. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/581.

Griffin:2011:TWL

[GM11] Philip Griffin and Ross Maller. The time at which a Lévy process creeps. Electronic Journal of Probability, 16:79:2182–79:2202, 2011. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/945.

Gouezel:2014:MBC

[GM14a] Sébastien Gouëzel and Ian Melbourne. Moment bounds and concentration inequalities for slowly mixing dynamical systems. Electronic Journal of Probability, 19(??):93:1–93:30, ???? 2014. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/3427.

Graczyk:2014:SSN

[GM14b] Piotr Graczyk and Jacek Małecki. Strong solutions of non-colliding particle systems. Electronic Journal of Probability, 19(??):119:1–119:21, ???? 2014.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
3842.

Grubel:2015:RRT

[GM15] Rudolf Grübel and Igor Michailow. Random recursive trees: a boundary theory approach. Electronic Journal of Probability, 20(??):37:1–37:22, ???? 2015.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
3832.

Gu:2017:GGF

[GM17a] Yu Gu and Jean-Christophe Mourrat. On generalized Gaussian free fields and stochastic homogenization. Electronic Journal of Probability, 22(??):28:1–28:21, ???? 2017. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.

org/euclid.ejp/1490320844.

Gwynne:2017:SLU

[GM17b] Ewain Gwynne and Jason Miller. Scaling limit of the uniform infinite half-plane quadrangulation in the Gromov–Hausdorff–Prokhorov-uniform topology. Electronic Journal of Probability, 22(??):84:1–84:47, ???? 2017. CODEN ???? ISSN

1083-6489. URL https://projecteuclid.org/euclid.ejp/1507946759.

Guionnet:2020:LDL

[GM20] Alice Guionnet and Mylène Maïda. Large deviations for the largest eigenvalue of the sum of two random matrices. Electronic Journal of Probability, 25(??):14:1–14:24, ???? 2020. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1580871680.

Guionnet:2022:LDG

[GM22] Alice Guionnet and Ronan Memin. Large deviations for Gibbs ensembles of the classical Toda chain. Electronic Journal of Probability, 27(??):1–29, 2022.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-27/issue-none/Large-deviations-
for-Gibbs-ensembles-of-the-classical-Toda-chain/10.1214/22-EJP771.
full.

Garet:2017:CTI

[GMPT17] Olivier Garet, Régine Marchand, Eviatar B. Procaccia, and Marie Théret. Continuity of the time and isoperimetric constants in supercritical percolation. Electronic Journal of Probability, 22(??):78:1–78:35, ???? 2017. CODEN ???? ISSN

1083-6489. URL https://projecteuclid.org/euclid.ejp/1506931228.

Georgiou:2019:MCH

[GMPW19] Nicholas Georgiou, Mikhail V. Menshikov, Dimitri Petritis, and Andrew R. Wade. Markov chains with heavy-tailed increments and asymptotically zero drift. Electronic Journal of Probability, 24(??):62:1–62:28, ???? 2019. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.ejp/
1561082671.

Gwynne:2019:HFM

[GMS19] Ewain Gwynne, Jason Miller, and Scott Sheffield. Harmonic functions on mated-CRT maps. Electronic Journal of Probability, 24(??):58:1–58:55, ???? 2019.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.
ejp/1561082667.

Ganesh:2005:SPL

[GMT05] Ayalvadi Ganesh, Claudio Macci, and Giovanni Torrisi. Sample path large deviations principles for Poisson shot noise processes and applications. Electronic Journal of Probability, 10:32:1026–32:1043, 2005. CODEN ???? ISSN 1083-

6489. URL http://ejp.ejpecp.org/article/view/273.

Goel:2006:MTB

[GMT06] Sharad Goel, Ravi Montenegro, and Prasad Tetali. Mixing time bounds via the spectral profile. Electronic Journal of Probability, 11:1:1–1:26, 2006. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/300.

Georgiou:2019:IPN

[GMW19] Nicholas Georgiou, Aleksandar Mijatović, and Andrew R. Wade. Invariance principle for non-homogeneous random walks. Electronic Journal of Probability, 24(??):48:1–48:38, ???? 2019. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1558145016.

Gradinaru:2003:AFS

[GN03] Mihai Gradinaru and Ivan Nourdin. Approximation at first and second order of m-order integrals of the fractional Brownian motion and of certain semimartingales. Electronic Journal of Probability, 8:18:1–18:26, 2003. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/166.

Glatzel:2021:SRW

[GN21] Tabea Glatzel and Jan Nagel. The speed of random walk on Galton–Watson trees with vanishing conductances. Electronic Journal of Probability, 26(18): 1–19, February 2021. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/journals/electronic-journal-of-probability/volume-
26/issue-none/The-speed-of-random-walk-on-Galton--Watson-trees-with/
10.1214/21-EJP645.full.

Gnedin:2007:CR

[Gne07] Alexander Gnedin. The chain records. Electronic Journal of Probability, 12: 26:767–26:786, 2007. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.

org/article/view/410.

Garino:2022:AED

[GNV22] Valentin Garino, Ivan Nourdin, and Pierre Vallois. Asymptotic error distribution for the Riemann approximation of integrals driven by fractional Brownian motion. Electronic Journal of Probability, 27(??):1–43, 2022. CODEN ????

ISSN 1083-6489. URL https://projecteuclid.org/journals/electronic-
journal-of-probability/volume-27/issue-none/Asymptotic-error-distribution-
for-the-Riemann-approximation-of-integrals-driven/10.1214/22-EJP852.
full.

Guerrero:2021:ASW

[GNZ21] Raul Bolaños Guerrero, David Nualart, and Guangqu Zheng. Averaging 2d stochastic wave equation. Electronic Journal of Probability, 26(18):1–32, February 2021. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

journals/electronic-journal-of-probability/volume-26/issue-none/Averaging-
2d-stochastic-wave-equation/10.1214/21-EJP672.full.

Gold:2018:IIG

[Gol18a] Julian Gold. Intrinsic isoperimetry of the giant component of supercritical bond percolation in dimension two. Electronic Journal of Probability, 23(??): 53:1–53:41, ???? 2018. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1527818431.

Goldstein:2018:NAD

[Gol18b] Larry Goldstein. Non-asymptotic distributional bounds for the Dickman approximation of the running time of the Quickselect algorithm. Electronic Journal of Probability, 23(??):100:1–100:13, ???? 2018. CODEN ???? ISSN 1083-6489.

URL https://projecteuclid.org/euclid.ejp/1538445816.

Gozlan:2012:TEI

[Goz12] Nathael Gozlan. Transport-entropy inequalities on the line. Electronic Journal of Probability, 17:49:1–49:18, 2012. CODEN ???? ISSN 1083-6489. URL http:

//ejp.ejpecp.org/article/view/1864.

Gallesco:2012:RWU

[GP12] Christophe Gallesco and Serguei Popov. Random walks with unbounded jumps among random conductances I: Uniform quenched CLT. Electronic Journal of Probability, 17:85:1–85:22, 2012. CODEN ???? ISSN 1083-6489. URL http://

ejp.ejpecp.org/article/view/1826.

Gaudio:2020:ARW

[GP20] Julia Gaudio and Yury Polyanskiy. Attracting random walks. Electronic Journal of Probability, 25(??):73:1–73:31, ???? 2020. CODEN ???? ISSN 1083-6489.

URL https://projecteuclid.org/euclid.ejp/1593568837.

Guillotin-Plantard:2012:RTR

[GPP12] Nadine Guillotin-Plantard and Françoise Pène. Renewal theorems for random walk in random scenery. Electronic Journal of Probability, 17:78:1–78:22, 2012.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
1843.

Gadat:2013:LDP

[GPP13] Sébastien Gadat, Fabien Panloup, and Clément Pellegrini. Large deviation principle for invariant distributions of memory gradient diffusions. Electronic Journal of Probability, 18:81:1–81:34, 2013. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/2031.

Greven:2016:FTS

[GPPW16] Andreas Greven, Peter Pfaffelhuber, Cornelia Pokalyuk, and Anton Wakolbinger. The fixation time of a strongly beneficial allele in a structured population. Electronic Journal of Probability, 21(??):61:1–61:42, ???? 2016. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.ejp/
1475586182.

Guo:2022:QHB

[GPT22] Xiaoqin Guo, Jonathon Peterson, and Hung V. Tran. Quantitative homogenization in a balanced random environment. Electronic Journal of Probability, 27(??):1–31, 2022. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-27/issue-none/Quantitative-homogenization-in-
a-balanced-random-environment/10.1214/22-EJP851.full.

Gantert:2009:STR

[GPV09] Nina Gantert, Serguei Popov, and Marina Vachkovskaia. Survival time of random walk in random environment among soft obstacles. Electronic Journal of Probability, 14:22:569–22:593, 2009. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/631.

Garbit:2014:ETC

[GR14] Rodolphe Garbit and Kilian Raschel. On the exit time from a cone for Brownian motion with drift. Electronic Journal of Probability, 19:63:1–63:27, 2014.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
3169.

Guerra:2015:AED

[GR15] Enrique Guerra and Alejandro F. Ramirez. Almost exponential decay for the exit probability from slabs of ballistic RWRE. Electronic Journal of Probability, 20(??):24:1–24:17, ???? 2015. CODEN ???? ISSN 1083-6489. URL http://

ejp.ejpecp.org/article/view/3655.

Guerra:2017:ADR

[GR17] Enrique Guerra and Alejandro F. Ramírez. Asymptotic direction for random walks in mixing random environments. Electronic Journal of Probability, 22 (??):92:1–92:41, ???? 2017. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1508810545.

Gravner:1996:PTT

[Gra96] Janko Gravner. Percolation times in two-dimensional models for excitable media. Electronic Journal of Probability, 1:12:1–12:19, 1996. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/12.

Graf:2014:FFM

[Gra14] Robert Graf. A forest-fire model on the upper half-plane. Electronic Journal of Probability, 19:8:1–8:27, 2014. CODEN ???? ISSN 1083-6489. URL http://

ejp.ejpecp.org/article/view/2625.

Groeneboom:2010:MBM

[Gro10] Piet Groeneboom. The maximum of Brownian motion minus a parabola. Electronic Journal of Probability, 15:62:1930–62:1937, 2010. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/826.

Groeneboom:2011:VLC

[Gro11] Piet Groeneboom. Vertices of the least concave majorant of Brownian motion with parabolic drift. Electronic Journal of Probability, 16:84:2334–84:2358, 2011.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/

959. See erratum [Gro13].

Groeneboom:2013:EVL

[Gro13] Piet Groeneboom. Erratum: “Vertices of the least concave majorant of Brownian motion with parabolic drift”. Electronic Journal of Probability, 18:46:1, 2013.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/

2697. See [Gro11].

Groux:2017:AFR

[Gro17] Benjamin Groux. Asymptotic freeness for rectangular random matrices and large deviations for sample covariance matrices with sub-Gaussian tails. Electronic Journal of Probability, 22(??):53:1–53:40, ???? 2017. CODEN ???? ISSN

1083-6489. URL https://projecteuclid.org/euclid.ejp/1498010466.

Grotto:2020:SSD

[Gro20] Francesco Grotto. Stationary solutions of damped stochastic 2-dimensional Euler’s equation. Electronic Journal of Probability, 25(??):69:1–69:24, ???? 2020.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.
ejp/1593137129.

Grosskinsky:2013:DCS

[GRV13] Stefan Grosskinsky, Frank Redig, and Kiamars Vafayi. Dynamics of condensation in the symmetric inclusion process. Electronic Journal of Probability, 18: 66:1–66:23, 2013. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.

org/article/view/2720.

Griffiths:2007:RIA

[GS07] Robert Griffiths and Dario Spano. Record indices and age-ordered frequencies in exchangeable Gibbs partitions. Electronic Journal of Probability, 12:40:1101– 40:1130, 2007. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/434.

Gartner:2011:TCP

[GS11] Jürgen Gärtner and Adrian Schnitzler. Time correlations for the parabolic Anderson model. Electronic Journal of Probability, 16:56:1519–56:1548, 2011.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
917.

Geiss:2016:MDR

[GS16] Christel Geiss and Alexander Steinicke. Malliavin derivative of random functions and applications to Lévy driven BSDEs. Electronic Journal of Probability, 21 (??):10:1–10:28, ???? 2016. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1455026806.

Gravner:2017:BPP

[GS17a] Janko Gravner and David Sivakoff. Bootstrap percolation on products of cycles and complete graphs. Electronic Journal of Probability, 22(??):29:1–29:20, ???? 2017. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

euclid.ejp/1490320845.

Gwynne:2017:SLC

[GS17b] Ewain Gwynne and Xin Sun. Scaling limits for the critical Fortuin–Kasteleyn model on a random planar map II: local estimates and empty reduced word exponent. Electronic Journal of Probability, 22(??):45:1–45:56, ???? 2017. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.ejp/
1494036159.

Gantert:2020:STP

[GS20] Nina Gantert and Dominik Schmid. The speed of the tagged particle in the exclusion process on Galton–Watson trees. Electronic Journal of Probability, 25 (??):71:1–71:27, ???? 2020. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1593568835.

Groisman:2021:RDB

[GSL21] Pablo Groisman and Nahuel Soprano-Loto. Rank dependent branching-selection particle systems. Electronic Journal of Probability, 26(18):1–27, February 2021.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-26/issue-none/Rank-dependent-
branching-selection-particle-systems/10.1214/21-EJP724.full.

Gotze:2019:HOC

[GSS19] Friedrich Götze, Holger Sambale, and Arthur Sinulis. Higher order concentration for functions of weakly dependent random variables. Electronic Journal of Probability, 24(??):85:1–85:19, ???? 2019. CODEN ???? ISSN 1083-6489. URL

https://projecteuclid.org/euclid.ejp/1568080865.

Gotze:2021:CIP

[GSS21] Friedrich Götze, Holger Sambale, and Arthur Sinulis. Concentration inequalities for polynomials in α-sub-exponential random variables. Electronic Journal of Probability, 26(18):1–22, February 2021. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-26/issue-none/Concentration-inequalities-for-
polynomials-in-%ce%b1-sub-exponential-random-variables/10.1214/21-
EJP606.full.

Giovanni:2007:SCG

[GT07] Peccati Giovanni and Murad Taqqu. Stable convergence of generalized L^{2} stochastic integrals and the principle of conditioning. Electronic Journal of Probability, 12:15:447–15:480, 2007. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/404.

Gouere:2017:PTC

[GT17] Jean-Baptiste Gouéré and Marie Théret. Positivity of the time constant in a continuous model of first passage percolation. Electronic Journal of Probability, 22(??):49:1–49:21, ???? 2017. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1496196077.

Gotze:2023:LSS

[GT23] F. Götze and A. Tikhomirov. On the largest and the smallest singular value of sparse rectangular random matrices. Electronic Journal of Probability, 28(??):1–18, 2023. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-28/issue-none/On-the-largest-and-the-smallest-
singular-value-of-sparse/10.1214/23-EJP919.full.

Gufler:2018:PCT

[Guf18a] Stephan Gufler. Pathwise construction of tree-valued Fleming–Viot processes. Electronic Journal of Probability, 23(??):42:1–42:58, ???? 2018. CODEN

???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.ejp/
1525852819.

Gufler:2018:REC

[Guf18b] Stephan Gufler. A representation for exchangeable coalescent trees and generalized tree-valued Fleming–Viot processes. Electronic Journal of Probability, 23 (??):41:1–41:42, ???? 2018. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1525852818.

Gupta:2012:FVL

[Gup12] Ankit Gupta. The Fleming–Viot limit of an interacting spatial population with fast density regulation. Electronic Journal of Probability, 17:104:1–104:55, 2012.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
1964.

Grothaus:2017:SDE

[GV17] Martin Grothaus and Robert Voßhall. Stochastic differential equations with sticky reflection and boundary diffusion. Electronic Journal of Probability, 22 (??):7:1–7:37, ???? 2017. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1485486107.

Garban:2020:BFP

[GV20] Christophe Garban and Hugo Vanneuville. Bargmann–Fock percolation is noise sensitive. Electronic Journal of Probability, 25(??):1–20, 2020. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-25/issue-none/Bargmann--
Fock-percolation-is-noise-sensitive/10.1214/20-EJP491.full.

Gao:2009:MDL

[GW09] Fuqing Gao and Yanqing Wang. Moderate deviations and laws of the iterated logarithm for the volume of the intersections of Wiener sausages. Electronic Journal of Probability, 14:65:1900–65:1935, 2009. CODEN ???? ISSN 1083-

6489. URL http://ejp.ejpecp.org/article/view/692.

Grote:2018:ASC

[GW18] Julian Grote and Elisabeth Werner. Approximation of smooth convex bodies by random polytopes. Electronic Journal of Probability, 23(??):9:1–9:21, ???? 2018. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

euclid.ejp/1518426057.

Gnedin:2007:NCL

[GY07] Alexander Gnedin and Yuri Yakubovich. On the number of collisions in lambdacoalescents. Electronic Journal of Probability, 12:56:1547–56:1567, 2007. CO-

DEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
464.

Giovanni:2010:MDG

[GZ10] Peccati Giovanni and Cengbo Zheng. Multi-dimensional Gaussian fluctuations on the Poisson space. Electronic Journal of Probability, 15:48:1487–48:1527, 2010. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/

view/813.

Haas:2017:AHR

[Haa17] Bénédicte Haas. Asymptotics of heights in random trees constructed by aggregation. Electronic Journal of Probability, 22(??):21:1–21:25, ???? 2017. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.ejp/
1487646307.

Habermann:2019:STF

[Hab19] Karen Habermann. Small-time fluctuations for the bridge in a model class of hypoelliptic diffusions of weak Hörmander type. Electronic Journal of Probability, 24(??):11:1–11:19, ???? 2019. CODEN ???? ISSN 1083-6489. URL

https://projecteuclid.org/euclid.ejp/1550480425.

Hajri:2011:SFR

[Haj11] Hatem Hajri. Stochastic flows related to Walsh Brownian motion. Electronic Journal of Probability, 16:58:1563–58:1599, 2011. CODEN ???? ISSN 1083-

6489. URL http://ejp.ejpecp.org/article/view/924.

Halconruy:2022:MCM

[Hal22] Hélène Halconruy. Malliavin calculus for marked binomial processes and applications. Electronic Journal of Probability, 27(??):1–39, 2022. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-27/issue-none/Malliavin-
calculus-for-marked-binomial-processes-and-applications/10.1214/22-
EJP892.full.

Hammond:2019:SAP

[Ham19] Alan Hammond. On self-avoiding polygons and walks: the snake method via polygon joining. Electronic Journal of Probability, 24(??):49:1–49:43, ???? 2019.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.
ejp/1558404407.

Handa:2014:EPC

[Han14] Kenji Handa. Ergodic properties for α-cir models and a class of generalized Fleming–Viot processes. Electronic Journal of Probability, 19:65:1–65:25, 2014.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
2928.

Hasebe:2014:FID

[Has14] Takahiro Hasebe. Free infinite divisibility for beta distributions and related ones. Electronic Journal of Probability, 19(??):81:1–81:33, ???? 2014. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/3448.

Hausenblas:2005:EUR

[Hau05] Erika Hausenblas. Existence, uniqueness and regularity of parabolic SPDEs driven by Poisson random measure. Electronic Journal of Probability, 10: 46:1496–46:1546, 2005. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/297.

Hayashi:2021:SSL

[Hay21] Kohei Hayashi. Spatial-segregation limit for exclusion processes with two components under unbalanced reaction. Electronic Journal of Probability, 26 (18):1–36, February 2021. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/journals/electronic-journal-of-probability/volume-
26/issue-none/Spatial-segregation-limit-for-exclusion-processes-with-
two-components-under/10.1214/21-EJP621.full.

Hambly:2009:PHI

[HB09] Ben Hambly and Martin Barlow. Parabolic Harnack inequality and local limit theorem for percolation clusters. Electronic Journal of Probability, 14:1:1–1:26, 2009. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/

view/587.

Hirsch:2017:PCP

[HBS17] Christian Hirsch, Tim Brereton, and Volker Schmidt. Percolation and convergence properties of graphs related to minimal spanning forests. Electronic Journal of Probability, 22(??):105:1–105:21, ???? 2017. CODEN ???? ISSN

1083-6489. URL https://projecteuclid.org/euclid.ejp/1514430041.

Hansen:2020:EUQ

[HC20] Mads Christian Hansen and Wiuf Carsten. Existence of a unique quasistationary distribution in stochastic reaction networks. Electronic Journal of Probability, 25(??):45:1–45:30, ???? 2020. CODEN ???? ISSN 1083-6489. URL

https://projecteuclid.org/euclid.ejp/1587024023.

Hilario:2015:RWR

[HdHS^{+}15] Marcelo Hilário, Frank den Hollander, Vladas Sidoravicius, Renato Soares dos Santos, and Augusto Teixeira. Random walk on random walks. Electronic Journal of Probability, 20(??):95:1–95:35, ???? 2015. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/4437.

He:2022:MCF

[He22] Jimmy He. Markov chains on finite fields with deterministic jumps. Electronic Journal of Probability, 27(??):1–17, 2022. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-27/issue-none/Markov-chains-on-finite-fields-
with-deterministic-jumps/10.1214/22-EJP757.full.

Heil:2013:SMP

[Hei13] Hadrian Heil. A stationary, mixing and perturbative counterexample to the 0–1-law for random walk in random environment in two dimensions. Electronic Journal of Probability, 18:1:1–1:33, 2013. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/1880.

Heilman:2014:EPO

[Hei14] Steven Heilman. Euclidean partitions optimizing noise stability. Electronic Journal of Probability, 19:71:1–71:37, 2014. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/3083.

Heiny:2022:LSC

[Hei22] Johannes Heiny. Large sample correlation matrices: a comparison theorem and its applications. Electronic Journal of Probability, 27(??):1–20, 2022. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-27/issue-none/Large-sample-
correlation-matrices--a-comparison-theorem-and-its/10.1214/22-EJP817.
full.

Hermon:2018:FT

[Her18] Jonathan Hermon. Frogs on trees? Electronic Journal of Probability, 23 (??):17:1–17:40, ???? 2018. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1519354946.

Herry:2020:SLT

[Her20] Ronan Herry. Stable limit theorems on the Poisson space. Electronic Journal of Probability, 25(??):1–30, 2020. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-25/issue-none/Stable-limit-theorems-on-the-
Poisson-space/10.1214/20-EJP557.full.

Heicklen:2005:RPS

[HH05] Deborah Heicklen and Christopher Hoffman. Return probabilities of a simple random walk on percolation clusters. Electronic Journal of Probability, 10: 8:250–8:302, 2005. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.

org/article/view/240.

Hamadene:2006:BTR

[HH06] Said Hamadene and Mohammed Hassani. BSDEs with two reflecting barriers driven by a Brownian motion and Poisson noise and related Dynkin game. Electronic Journal of Probability, 11:5:121–5:145, 2006. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/303.

Haggstrom:2014:FRC

[HH14] Olle Häggström and Timo Hirscher. Further results on consensus formation in the Deffuant model. Electronic Journal of Probability, 19:19:1–19:26, 2014.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
3116.

Halberstam:2022:CRW

[HH22] Noah Halberstam and Tom Hutchcroft. Collisions of random walks in dynamic random environments. Electronic Journal of Probability, 27(??):1–18, 2022.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-27/issue-none/Collisions-
of-random-walks-in-dynamic-random-environments/10.1214/21-EJP738.
full.

Hachem:2016:LCC

[HHN16] Walid Hachem, Adrien Hardy, and Jamal Najim. Large complex correlated Wishart matrices: the Pearcey kernel and expansion at the hard edge. Electronic Journal of Probability, 21(??):1:1–1:36, ???? 2016. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/euclid.ejp/1454514661.

Hu:2015:SHE

[HHNT15] Yaozhong Hu, Jingyu Huang, David Nualart, and Samy Tindel. Stochastic heat equations with general multiplicative Gaussian noises: Hölder continuity and intermittency. Electronic Journal of Probability, 20(??):55:1–55:50, ???? 2015.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
3316.

Houdre:2023:CLT

[HI23] Christian Houdré and Ümit Işlak. A central limit theorem for the length of the longest common subsequences in random words. Electronic Journal of Probability, 28(??):1–24, 2023. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-28/issue-none/A-central-limit-theorem-for-the-
length-of-the-longest/10.1214/22-EJP894.full.

Hillion:2014:IPM

[Hil14] Erwan Hillion. W_{1,+}-interpolation of probability measures on graphs. Electronic Journal of Probability, 19(??):92:1–92:29, ???? 2014. CODEN ???? ISSN 1083-

6489. URL http://ejp.ejpecp.org/article/view/3336.

Hoshino:2017:SCG

[HIN17] Masato Hoshino, Yuzuru Inahama, and Nobuaki Naganuma. Stochastic complex Ginzburg–Landau equation with space-time white noise. Electronic Journal of Probability, 22(??):104:1–104:68, ???? 2017. CODEN ???? ISSN 1083-6489.

URL https://projecteuclid.org/euclid.ejp/1513349792.

Hambly:2007:NVP

[HJ07] Ben Hambly and Liza Jones. Number variance from a probabilistic perspective: infinite systems of independent Brownian motions and symmetric alpha stable processes. Electronic Journal of Probability, 12:30:862–30:887, 2007. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/419. See

erratum [HJ09].

Hambly:2009:ENV

[HJ09] Ben Hambly and Lisa Jones. Erratum to “Number Variance from a probabilistic perspective, infinite systems of independent Brownian motions and symmetric α-stable processes”. Electronic Journal of Probability, 14:37:1074–37:1079, 2009.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/

658. See [HJ07].

Hwang:2011:CLT

[HJ11] Hsien-Kuei Hwang and Svante Janson. A Central Limit Theorem for random ordered factorizations of integers. Electronic Journal of Probability, 16:12:347– 12:361, 2011. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/ article/view/858. See erratum [HJ13].

Hwang:2013:ECL

[HJ13] Hsien-Kuei Hwang and Svante Janson. Erratum: “A central limit theorem for random ordered factorizations of integers”. Electronic Journal of Probability, 18:16:1–16:3, 2013. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.

org/article/view/2297. See [HJ11].

Holmgren:2015:LLF

[HJ15a] Cecilia Holmgren and Svante Janson. Limit laws for functions of fringe trees for binary search trees and random recursive trees. Electronic Journal of Probability, 20(??):4:1–4:51, ???? 2015. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/3627.

Holmgren:2015:ADT

[HJ15b] Cecilia Ingrid Holmgren and Svante Janson. Asymptotic distribution of twoprotected nodes in ternary search trees. Electronic Journal of Probability, 20 (??):9:1–9:20, ???? 2015. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/3577.

Hillion:2019:PSO

[HJ19] Erwan Hillion and Oliver Johnson. A proof of the Shepp–Olkin entropy monotonicity conjecture. Electronic Journal of Probability, 24(??):126:1–126:14, ???? 2019. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

euclid.ejp/1573268590.

Hoffman:2019:ISF

[HJJ19] Christopher Hoffman, Tobias Johnson, and Matthew Junge. Infection spread for the frog model on trees. Electronic Journal of Probability, 24(??):112:1–112:29, ???? 2019. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.

org/euclid.ejp/1570586692.

Heiny:2022:TST

[HJP22] Johannes Heiny, Samuel Johnston, and Joscha Prochno. Thin-shell theory for rotationally invariant random simplices. Electronic Journal of Probability, 27(??): 1–41, 2022. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.

org/journals/electronic-journal-of-probability/volume-27/issue-none/
Thin-shell-theory-for-rotationally-invariant-random-simplices/10.
1214/21-EJP734.full.

Herman:2020:STC

[HJT20] John Herman, Ifan Johnston, and Lorenzo Toniazzi. Space-time coupled evolution equations and their stochastic solutions. Electronic Journal of Probability, 25(??):1–21, 2020. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-25/issue-none/Space-time-coupled-evolution-
equations-and-their-stochastic-solutions/10.1214/20-EJP544.full.

Hutzenthaler:2020:OCD

[HJvWW20] Martin Hutzenthaler, Arnulf Jentzen, and von Wurstemberger Wurstemberger. Overcoming the curse of dimensionality in the approximative pricing of financial derivatives with default risks. Electronic Journal of Probability, 25(??):1– 73, 2020. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

journals/electronic-journal-of-probability/volume-25/issue-none/Overcoming-
the-curse-of-dimensionality-in-the-approximative-pricing-of/10.1214/
20-EJP423.full.

Hairer:2012:SPM

[HK12] Martin Hairer and David Kelly. Stochastic PDEs with multiscale structure. Electronic Journal of Probability, 17:52:1–52:38, 2012. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/1807.

Hartung:2018:PDC

[HK18] Lisa Hartung and Anton Klimovsky. The phase diagram of the complex branching Brownian motion energy model. Electronic Journal of Probability, 23(??): 127:1–127:27, ???? 2018. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1545188694.

Huang:2020:ASK

[HK20] Jingyu Huang and Davar Khoshnevisan. Analysis of a stratified Kraichnan flow. Electronic Journal of Probability, 25(??):1–67, 2020. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-25/issue-none/Analysis-of-a-stratified-Kraichnan-
flow/10.1214/20-EJP524.full.

Hitczenko:1998:HCM

[HKL^{+}98] Paweł Hitczenko, Stanisław Kwapień, Wenbo V. Li, Gideon Schechtman, Thomas Schlumprecht, and Joel Zinn. Hypercontractivity and comparison of moments of iterated maxima and minima of independent random variables. Electronic Journal of Probability, 3:2:1–2:26, 1998. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/24.

Henning:2019:GGM

[HKLR19] Florian Henning, Christof Külske, Arnaud Le Ny, and Utkir A. Rozikov. Gradient Gibbs measures for the SOS model with countable values on a Cayley tree. Electronic Journal of Probability, 24(??):104:1–104:23, ???? 2019. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.ejp/
1569895473.

Hebbar:2020:ABB

[HKN20] Pratima Hebbar, Leonid Koralov, and James Nolen. Asymptotic behavior of branching diffusion processes in periodic media. Electronic Journal of Probability, 25(??):1–40, 2020. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-25/issue-none/Asymptotic-behavior-of-branching-
diffusion-processes-in-periodic-media/10.1214/20-EJP527.full.

Hayashi:2014:HCP

[HKY14] Masafumi Hayashi, Arturo Kohatsu, and Go Yuki. Hölder continuity property of the densities of SDEs with singular drift coefficients. Electronic Journal of Probability, 19:77:1–77:22, 2014. CODEN ???? ISSN 1083-6489. URL http://

ejp.ejpecp.org/article/view/2609.

Heveling:2007:PSC

[HL07] Matthias Heveling and Gunter Last. Point shift characterization of Palm measures on Abelian groups. Electronic Journal of Probability, 12:5:122–5:137, 2007.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
394.

Heunis:2008:ICN

[HL08] Andrew Heunis and Vladimir Lucic. On the innovations conjecture of nonlinear filtering with dependent data. Electronic Journal of Probability, 13:72:2190– 72:2216, 2008. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/585.

Huber:2012:SRI

[HL12] Mark Huber and Jenny Law. Simulation reduction of the Ising model to general matchings. Electronic Journal of Probability, 17:33:1–33:15, 2012. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/1998.

Huang:2014:CER

[HL14] Chunmao Huang and Quansheng Liu. Convergence in L^{p} and its exponential rate for a branching process in a random environment. Electronic Journal of Probability, 19(??):104:1–104:22, ???? 2014. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/3388.

Huesler:2010:CHE

[HLP10] Juerg Huesler, Anna Ladneva, and Vladimir Piterbarg. On clusters of high extremes of Gaussian stationary processes with ε-separation. Electronic Journal of Probability, 15:59:1825–59:1862, 2010. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/828.

Hermon:2016:TVS

[HLP16] Jonathan Hermon, Hubert Lacoin, and Yuval Peres. Total variation and separation cutoffs are not equivalent and neither one implies the other. Electronic Journal of Probability, 21(??):44:1–44:36, ???? 2016. CODEN ???? ISSN

1083-6489. URL https://projecteuclid.org/euclid.ejp/1469557137.

Hilario:2019:STS

[HLP19] Marcelo Hilario, Xinyi Li, and Petr Panov. Shape theorem and surface fluctuation for Poisson cylinders. Electronic Journal of Probability, 24(??):68:1–68:16, ???? 2019. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.

org/euclid.ejp/1561687601.

Hsiau:2014:LRF

[HLY14] Shoou-Ren Hsiau, Yi-Shen Lin, and Yi-Ching Yao. Logconcave reward functions and optimal stopping rules of threshold form. Electronic Journal of Probability, 19(??):120:1–120:18, ???? 2014. CODEN ???? ISSN 1083-6489. URL http://

ejp.ejpecp.org/article/view/3745.

Heck:2001:PLD

[HM01] Matthias Heck and Faïza Maaouia. The principle of large deviations for martingale additive functionals of recurrent Markov processes. Electronic Journal of Probability, 6:8:1–8:26, 2001. CODEN ???? ISSN 1083-6489. URL http://

ejp.ejpecp.org/article/view/81.

Haas:2004:GSS

[HM04] Bénédicte Haas and Grégory Miermont. The genealogy of self-similar fragmentations with negative index as a continuum random tree. Electronic Journal of Probability, 9:4:57–4:97, 2004. CODEN ???? ISSN 1083-6489. URL http://

ejp.ejpecp.org/article/view/187.

Hairer:2011:THU

[HM11] Martin Hairer and Jonathan Mattingly. A theory of hypoellipticity and unique ergodicity for semilinear stochastic PDEs. Electronic Journal of Probability, 16:23:658–23:738, 2011. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/875.

Holroyd:2014:SDC

[HM14] Alexander Holroyd and James Martin. Stochastic domination and comb percolation. Electronic Journal of Probability, 19:5:1–5:16, 2014. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/2806.

Herzog:2015:NISa

[HM15a] David P. Herzog and Jonathan C. Mattingly. Noise-induced stabilization of planar flows I. Electronic Journal of Probability, 20(??):111:1–111:43, ???? 2015.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
4047.

Herzog:2015:NISb

[HM15b] David P. Herzog and Jonathan C. Mattingly. Noise-induced stabilization of planar flows II. Electronic Journal of Probability, 20(??):113:1–113:37, ???? 2015. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/

view/4048.

Ho:2022:EAB

[HM22] Fu-Hsuan Ho and Pascal Maillard. Efficient approximation of branching random walk Gibbs measures. Electronic Journal of Probability, 27(??):1–18, 2022.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-27/issue-none/Efficient-
approximation-of-branching-random-walk-Gibbs-measures/10.1214/22-
EJP800.full.

Haji-Mirsadeghi:2016:STB

[HMK16] Mir-Omid Haji-Mirsadeghi and Ali Khezeli. Stable transports between stationary random measures. Electronic Journal of Probability, 21(??):51:1–51:25, ???? 2016. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.

org/euclid.ejp/1470414022.

Huckemann:2015:SCL

[HMMN15] Stephan Huckemann, Jonathan Mattingly, Ezra Miller, and James Nolen. Sticky central limit theorems at isolated hyperbolic planar singularities. Electronic Journal of Probability, 20(??):78:1–78:34, ???? 2015. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/3887.

Hryniv:2012:NHR

[HMMW12] Ostap Hryniv, Iain MacPhee, Mikhail Menshikov, and Andrew Wade. Nonhomogeneous random walks with non-integrable increments and heavy-tailed random walks on strips. Electronic Journal of Probability, 17:59:1–59:28, 2012.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
2216.

Heil:2011:BRW

[HMN11] Hadrian Heil, Nakashima Makoto, and Yoshida Nobuo. Branching random walks in random environment are diffusive in the regular growth phase. Electronic Journal of Probability, 16:48:1318–48:1340, 2011. CODEN ???? ISSN 1083-

6489. URL http://ejp.ejpecp.org/article/view/922.

Herzog:2023:GDG

[HMN23] David P. Herzog, Jonathan C. Mattingly, and Hung D. Nguyen. Gibbsian dynamics and the generalized Langevin equation. Electronic Journal of Probability, 28(??):1–29, 2023. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-28/issue-none/Gibbsian-dynamics-and-the-generalized-
Langevin-equation/10.1214/23-EJP904.full.

Hammond:2012:ETT

[HMP12] Alan Hammond, Elchanan Mossel, and Gábor Pete. Exit time tails from pairwise decorrelation in hidden Markov chains, with applications to dynamical percolation. Electronic Journal of Probability, 17:68:1–68:16, 2012. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/2229.

Heydenreich:2014:SBR

[HMR14] Markus Heydenreich, Franz Merkl, and Silke W. W. Rolles. Spontaneous breaking of rotational symmetry in the presence of defects. Electronic Journal of Probability, 19(??):111:1–111:17, ???? 2014. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/2971.

Henry:2023:TRS

[HMT23] Benoît Henry, Sylvie Méléard, and Viet Chi Tran. Time reversal of spinal processes for linear and non-linear branching processes near stationarity. Electronic Journal of Probability, 28(??):1–27, 2023. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-28/issue-none/Time-reversal-of-spinal-processes-
for-linear-and-non-linear/10.1214/23-EJP911.full.

Hobson:2022:CLC

[HN22] David Hobson and Dominykas Norgilas. A construction of the left-curtain coupling. Electronic Journal of Probability, 27(??):1–46, 2022. CODEN ????

ISSN 1083-6489. URL https://projecteuclid.org/journals/electronic-
journal-of-probability/volume-27/issue-none/A-construction-of-the-
left-curtain-coupling/10.1214/22-EJP868.full.

Hu:2019:HCS

[HNX19] Yaozhong Hu, David Nualart, and Panqiu Xia. Hölder continuity of the solutions to a class of SPDE’s arising from branching particle systems in a random environment. Electronic Journal of Probability, 24(??):105:1–105:52, ???? 2019.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.
ejp/1569895474.

Hamadene:2003:RBS

[HO03] Said Hamadène and Youssef Ouknine. Reflected backward stochastic differential equation with jumps and random obstacle. Electronic Journal of Probability, 8:2:1–2:20, 2003. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.

org/article/view/124.

Ho:2022:BMS

[Ho22] Ching-Wei Ho. The Brown measure of the sum of a self-adjoint element and an elliptic element. Electronic Journal of Probability, 27(??):1–32, 2022. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-27/issue-none/The-Brown-
measure-of-the-sum-of-a-self-adjoint/10.1214/22-EJP840.full.

Hobson:2015:ISS

[Hob15] David Hobson. Integrability of solutions of the Skorokhod embedding problem for diffusions. Electronic Journal of Probability, 20(??):83:1–83:26, ???? 2015.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
4121.

Holroyd:2006:MTM

[Hol06] Alexander Holroyd. The metastability threshold for modified bootstrap percolation in d dimensions. Electronic Journal of Probability, 11:17:418–17:433, 2006. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/

view/326.

Holmes:2008:CLT

[Hol08] Mark Holmes. Convergence of lattice trees to super-Brownian motion above the critical dimension. Electronic Journal of Probability, 13:23:671–23:755, 2008.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
499.

Holmgren:2012:NCS

[Hol12] Cecilia Holmgren. Novel characteristics of split trees by use of renewal theory. Electronic Journal of Probability, 17:5:1–5:27, 2012. CODEN ???? ISSN 1083-

6489. URL http://ejp.ejpecp.org/article/view/1723.

Holcomb:2018:RMH

[Hol18] Diane Holcomb. The random matrix hard edge: rare events and a transition. Electronic Journal of Probability, 23(??):85:1–85:20, ???? 2018. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.ejp/
1536717744.

Hong:2018:RLT

[Hon18] Jieliang Hong. Renormalization of local times of super-Brownian motion. Electronic Journal of Probability, 23(??):109:1–109:45, ???? 2018. CODEN

???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.ejp/
1540865372.

Hong:2020:BLT

[Hon20] Jieliang Hong. On the boundary local time measure of super-Brownian motion. Electronic Journal of Probability, 25(??):1–66, 2020. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-25/issue-none/On-the-boundary-local-time-measure-
of-super-Brownian-motion/10.1214/20-EJP507.full.

Hora:2020:EMP

[Hor20] Akihito Hora. Effect of microscopic pausing time distributions on the dynamical limit shapes for random Young diagrams. Electronic Journal of Probability, 25 (??):70:1–70:21, ???? 2020. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1593137130.

Hough:2017:MCC

[Hou17] Robert Hough. Mixing and cut-off in cycle walks. Electronic Journal of Probability, 22(??):90:1–90:49, ???? 2017. CODEN ???? ISSN 1083-6489. URL

https://projecteuclid.org/euclid.ejp/1508292261.

Hermisson:2008:PGH

[HP08] Joachim Hermisson and Peter Pfaffelhuber. The pattern of genetic hitchhiking under recurrent mutation. Electronic Journal of Probability, 13:68:2069–68:2106, 2008. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/

view/577.

Hutchcroft:2017:BPG

[HP17] Tom Hutchcroft and Yuval Peres. Boundaries of planar graphs: a unified approach. Electronic Journal of Probability, 22(??):100:1–100:20, ???? 2017. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.ejp/
1511578855.

Hermon:2018:SMT

[HP18] Jonathan Hermon and Yuval Peres. On sensitivity of mixing times and cutoff. Electronic Journal of Probability, 23(??):25:1–25:34, ???? 2018. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.ejp/
1521079338.

Hernandez:2021:UAW

[HP21] Camilo Hernández and Dylan Possamaï. A unified approach to well-posedness of type-I backward stochastic Volterra integral equations. Electronic Journal of Probability, 26(18):1–35, February 2021. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-26/issue-none/A-unified-approach-to-well-posedness-
of-type-I-backward/10.1214/21-EJP653.full.

He:2022:MTF

[HPX22] Jimmy He, Huy Tuan Pham, and Max Wenqiang Xu. Mixing time of fractional random walk on finite fields. Electronic Journal of Probability, 27(??):1–15, 2022.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-27/issue-none/Mixing-time-
of-fractional-random-walk-on-finite-fields/10.1214/22-EJP858.full.

Huesler:2011:EGP

[HPZ11] Juerg Huesler, Vladimir Piterbarg, and Yueming Zhang. Extremes of Gaussian processes with random variance. Electronic Journal of Probability, 16:45:1254– 45:1280, 2011. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/904.

Hajri:2014:SFM

[HR14] Hatem Hajri and Olivier Raimond. Stochastic flows on metric graphs. Electronic Journal of Probability, 19:12:1–12:20, 2014. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/2773.

Harter:2019:SAS

[HR19] Jonathan Harter and Adrien Richou. A stability approach for solving multidimensional quadratic BSDEs. Electronic Journal of Probability, 24(??):4:1–4:51, ???? 2019. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.

org/euclid.ejp/1549616425.

Hu:2007:EES

[HRW07] Zhishui Hu, John Robinson, and Qiying Wang. Edgeworth expansions for a sample sum from a finite set of independent random variables. Electronic Journal of Probability, 12:52:1402–52:1417, 2007. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/447.

Hairer:2012:TSA

[HRW12] Martin Hairer, Marc Ryser, and Hendrik Weber. Triviality of the 2D stochastic Allen–Cahn equation. Electronic Journal of Probability, 17:39:1–39:14, 2012.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
1731.

Huss:2012:IAM

[HS12] Wilfried Huss and Ecaterina Sava. Internal aggregation models on comb lattices. Electronic Journal of Probability, 17:30:1–30:21, 2012. CODEN ???? ISSN 1083-

6489. URL http://ejp.ejpecp.org/article/view/1940.

Holroyd:2013:IDT

[HS13] Alexander Holroyd and Terry Soo. Insertion and deletion tolerance of point processes. Electronic Journal of Probability, 18:74:1–74:24, 2013. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/2621.

Hansen:2014:CIS

[HS14] Niels Richard Hansen and Alexander Sokol. Causal interpretation of stochastic differential equations. Electronic Journal of Probability, 19(??):100:1–100:24, ???? 2014. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/2891.

Hu:2015:MVS

[HS15a] Yueyun Hu and Zhan Shi. The most visited sites of biased random walks on trees. Electronic Journal of Probability, 20(??):62:1–62:14, ???? 2015. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/4051.

Huveneers:2015:RWD

[HS15b] François Huveneers and François Simenhaus. Random walk driven by simple exclusion process. Electronic Journal of Probability, 20(??):105:1–105:42, ????

2015. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/

view/3906.

Holmes:2017:CBI

[HS17a] Mark Holmes and Thomas S. Salisbury. Conditions for ballisticity and invariance principle for random walk in non-elliptic random environment. Electronic Journal of Probability, 22(??):81:1–81:18, ???? 2017. CODEN ???? ISSN

1083-6489. URL https://projecteuclid.org/euclid.ejp/1507536148.

Hu:2017:HML

[HS17b] Wenqing Hu and Konstantinos Spiliopoulosï. Hypoelliptic multiscale Langevin diffusions: large deviations, invariant measures and small mass asymptotics. Electronic Journal of Probability, 22(??):55:1–55:38, ???? 2017. CODEN

???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.ejp/
1498809677.

Hammond:2020:MCP

[HS20] Alan Hammond and Sourav Sarkar. Modulus of continuity for polymer fluctuations and weight profiles in Poissonian last passage percolation. Electronic Journal of Probability, 25(??):29:1–29:38, ???? 2020. CODEN ???? ISSN

1083-6489. URL https://projecteuclid.org/euclid.ejp/1582534894.

Harel:2022:FCR

[HS22] Matan Harel and Yinon Spinka. Finitary codings for the randomcluster model and other infinite-range monotone models. Electronic Journal of Probability, 27(??):1–32, 2022. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-27/issue-none/Finitary-codings-for-the-random-
cluster-model-and-other-infinite/10.1214/22-EJP778.full.

Huveneers:2023:EPP

[HS23] François Huveneers and François Simenhaus. Evolution of a passive particle in a one-dimensional diffusive environment. Electronic Journal of Probability, 28(??):1–31, 2023. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-28/issue-none/Evolution-of-a-passive-particle-
in-a-one-dimensional-diffusive/10.1214/22-EJP896.full.

Hu:2009:CTM

[HSW09] Zhishui Hu, Qi-Man Shao, and Qiying Wang. Cramér type moderate deviations for the maximum of self-normalized sums. Electronic Journal of Probability, 14:41:1181–41:1197, 2009. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/663.

Herrmann:2010:SMS

[HT10a] Samuel Herrmann and Julian Tugaut. Stationary measures for self-stabilizing processes: asymptotic analysis in the small noise limit. Electronic Journal of Probability, 15:69:2087–69:2116, 2010. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/842.

Hwang:2010:MRB

[HT10b] Hsien-Kuei Hwang and Tsung-Hsi Tsai. Multivariate records based on dominance. Electronic Journal of Probability, 15:60:1863–60:1892, 2010. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/825.

Hug:2019:STS

[HT19] Daniel Hug and Christoph Thäle. Splitting tessellations in spherical spaces. Electronic Journal of Probability, 24(??):24:1–24:60, ???? 2019. CODEN

???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.ejp/
1553565775.

Huang:2021:NMC

[HT21] De Huang and Joel A. Tropp. Nonlinear matrix concentration via semigroup methods. Electronic Journal of Probability, 26(??):1–31, 2021. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-26/issue-none/Nonlinear-
matrix-concentration-via-semigroup-methods/10.1214/20-EJP578.full.

Hinz:2022:SRO

[HTV22] Michael Hinz, Jonas M. Tölle, and Lauri Viitasaari. Sobolev regularity of occupation measures and paths, variability and compositions. Electronic Journal of Probability, 27(??):1–29, 2022. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-27/issue-none/Sobolev-regularity-of-occupation-
measures-and-paths-variability-and-compositions/10.1214/22-EJP797.
full.

Hu:2000:LCH

[Hu00] Yueyun Hu. The laws of Chung and Hirsch for Cauchy’s principal values related to Brownian local times. Electronic Journal of Probability, 5:10:1–10:16, 2000.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
66.

Huang:2020:EGC

[Hua20] Xiangying Huang. Exponential growth and continuous phase transitions for the contact process on trees. Electronic Journal of Probability, 25(??):77:1–77:21, ???? 2020. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.

org/euclid.ejp/1594432885.

Huang:2021:PDD

[Hua21] Xing Huang. Path-distribution dependent SDEs with singular coefficients. Electronic Journal of Probability, 26(18):1–21, February 2021. CODEN ????

ISSN 1083-6489. URL https://projecteuclid.org/journals/electronic-
journal-of-probability/volume-26/issue-none/Path-distribution-dependent-
SDEs-with-singular-coefficients/10.1214/21-EJP630.full.

Hughes:2019:BLT

[Hug19] Thomas Hughes. A boundary local time for one-dimensional super-Brownian motion and applications. Electronic Journal of Probability, 24(??):54:1–54:58, ???? 2019. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.

org/euclid.ejp/1559700304.

Hulshof:2015:OAE

[Hul15] Tim Hulshof. The one-arm exponent for mean-field long-range percolation. Electronic Journal of Probability, 20(??):115:1–115:26, ???? 2015. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/3935.

Husson:2022:LDL

[Hus22] Jonathan Husson. Large deviations for the largest eigenvalue of matrices with variance profiles. Electronic Journal of Probability, 27(??):1–44, 2022. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-27/issue-none/Large-deviations-
for-the-largest-eigenvalue-of-matrices-with-variance/10.1214/22-EJP793.
full.

Hutzenthaler:2009:VIM

[Hut09] Martin Hutzenthaler. The Virgin Island model. Electronic Journal of Probability, 14:39:1117–39:1161, 2009. CODEN ???? ISSN 1083-6489. URL http://

ejp.ejpecp.org/article/view/646.

Hutzenthaler:2012:IDT

[Hut12] Martin Hutzenthaler. Interacting diffusions and trees of excursions: convergence and comparison. Electronic Journal of Probability, 17:71:1–71:49, 2012. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/2278.

Hutchcroft:2020:BCN

[Hut20] Tom Hutchcroft. The L^{2} boundedness condition in nonamenable percolation. Electronic Journal of Probability, 25(??):1–27, 2020. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-

of-probability/volume-25/issue-none/The-L2-boundedness-condition-
in-nonamenable-percolation/10.1214/20-EJP525.full.

Hutchcroft:2023:TAI

[Hut23] Tom Hutchcroft. Transience and anchored isoperimetric dimension of supercritical percolation clusters. Electronic Journal of Probability, 28(??):1–15, 2023.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-28/issue-none/Transience-
and-anchored-isoperimetric-dimension-of-supercritical-percolation-
clusters/10.1214/23-EJP905.full.

Hilario:2022:RCP

[HUVV22] Marcelo Hilário, Daniel Ungaretti, Daniel Valesin, and Maria Eulália Vares. Results on the contact process with dynamic edges or under renewals. Electronic Journal of Probability, 27(??):1–31, 2022. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-27/issue-none/Results-on-the-contact-process-
with-dynamic-edges-or-under/10.1214/22-EJP811.full.

Hessler:2010:ECP

[HW10] Martin Hessler and Johan Wästlund. Edge cover and polymatroid flow problems. Electronic Journal of Probability, 15:72:2200–72:2219, 2010. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/846.

Huang:2021:DED

[HW21] Xing Huang and Feng-Yu Wang. Derivative estimates on distributions of McKean–Vlasov SDEs. Electronic Journal of Probability, 26(??):1–12, 2021.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-26/issue-none/Derivative-
estimates-on-distributions-of-McKean--Vlasov-SDEs/10.1214/21-EJP582.
full.

Houdre:2008:CSM

[HX08] Christian Houdré and Hua Xu. Concentration of the spectral measure for large random matrices with stable entries. Electronic Journal of Probability, 13:5:107– 5:134, 2008. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/482.

Hirsch:2010:LMA

[HY10] Francis Hirsch and Marc Yor. Looking for martingales associated to a selfdecomposable law. Electronic Journal of Probability, 15:29:932–29:961, 2010.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/

Hambly:2018:EST

[HY18] Ben Hambly and Weiye Yang. Existence and space-time regularity for stochastic heat equations on p.c.f. fractals. Electronic Journal of Probability, 23(??): 22:1–22:30, ???? 2018. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1519722151.

Hong:2019:ABH

[HZ19] Wenming Hong and Xiaoyue Zhang. Asymptotic behaviour of heavy-tailed branching processes in random environments. Electronic Journal of Probability, 24(??):56:1–56:17, ???? 2019. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1559700306.

Imkeller:2019:DSD

[IdRS19] Peter Imkeller, Gonçalo dos Reis, and William Salkeld. Differentiability of SDEs with drifts of super-linear growth. Electronic Journal of Probability, 24 (??):3:1–3:43, ???? 2019. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1549616424.

Igloi:2005:ROT

[Igl05] Endre Iglói. A rate-optimal trigonometric series expansion of the fractional Brownian motion. Electronic Journal of Probability, 10:41:1381–41:1397, 2005.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
287.

Iyer:2021:PAC

[IJ21] Srikanth K. Iyer and Sanjoy Kr Jhawar. Poisson approximation and connectivity in a scale-free random connection model. Electronic Journal of Probability, 26(18):1–23, February 2021. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-26/issue-none/Poisson-approximation-and-connectivity-
in-a-scale-free-random-connection/10.1214/21-EJP651.full.

Iksanov:2016:LUR

[IKM16] Alexander Iksanov, Zakhar Kabluchko, and Alexander Marynych. Local universality for real roots of random trigonometric polynomials. Electronic Journal of Probability, 21(??):63:1–63:19, ???? 2016. CODEN ???? ISSN 1083-6489. URL

https://projecteuclid.org/euclid.ejp/1476706888.

Iksanov:2021:GFL

[IKM21] Alexander Iksanov, Konrad Kolesko, and Matthias Meiners. Gaussian fluctuations and a law of the iterated logarithm for Nerman’s martingale in the supercritical general branching process. Electronic Journal of Probability, 26(18):1–22, February 2021. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-26/issue-none/Gaussian-fluctuations-and-a-law-
of-the-iterated-logarithm-for/10.1214/21-EJP727.full.

Ishiwata:2020:CLT

[IKN20] Satoshi Ishiwata, Hiroshi Kawabi, and Ryuya Namba. Central limit theorems for non-symmetric random walks on nilpotent covering graphs: Part I. Electronic Journal of Probability, 25(??):1–46, 2020. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-25/issue-none/Central-limit-theorems-for-non-
symmetric-random-walks-on-nilpotent/10.1214/20-EJP486.full.

Iksanov:2023:LTD

[IMN23] Alexander Iksanov, Alexander Marynych, and Anatolii Nikitin. Limit theorems for discounted convergent perpetuities II. Electronic Journal of Probability, 28(??):1–22, 2023. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-28/issue-none/Limit-theorems-for-discounted-
convergent-perpetuities-II/10.1214/23-EJP907.full.

Ivanoff:2007:CCP

[IMP07] B. Gail Ivanoff, Ely Merzbach, and Mathieu Plante. A compensator characterization of point processes on topological lattices. Electronic Journal of Probability, 12:2:47–2:74, 2007. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/390.

Inahama:2016:STK

[Ina16] Yuzuru Inahama. Short time kernel asymptotics for rough differential equation driven by fractional Brownian motion. Electronic Journal of Probability, 21 (??):34:1–34:29, ???? 2016. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1461332875.

Iksanov:2021:LTD

[INS21] Alexander Iksanov, Anatolii Nikitin, and Igor Samoilenko. Limit theorems for discounted convergent perpetuities. Electronic Journal of Probability, 26 (18):1–25, February 2021. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/journals/electronic-journal-of-probability/volume-
26/issue-none/Limit-theorems-for-discounted-convergent-perpetuities/
10.1214/21-EJP705.full.

Ibarrola:2009:FTR

[IPR09] Ricardo Vélez Ibarrola and Tomas Prieto-Rumeau. De Finetti’s-type results for some families of non identically distributed random variables. Electronic Journal of Probability, 14:4:72–4:86, 2009. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/602.

Igloi:1999:LRD

[IT99] E. Iglói and G. Terdik. Long-range dependence through gamma-mixed Ornstein–Uhlenbeck process. Electronic Journal of Probability, 4:16:1–16:33, 1999. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/

view/53.

Ichinose:2000:NED

[IT00] Takashi Ichinose and Satoshi Takanobu. The norm estimate of the difference between the Kac operator and Schrödinger semigroup II: The general case including the relativistic case. Electronic Journal of Probability, 5:5:1–5:47, 2000.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
61.

Ioffe:2013:ASC

[IV13] Dmitry Ioffe and Yvan Velenik. An almost sure CLT for stretched polymers. Electronic Journal of Probability, 18:97:1–97:20, 2013. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/2231.

Jacquot:2010:HLL

[Jac10] Stephanie Jacquot. A historical law of large numbers for the Marcus–Lushnikov process. Electronic Journal of Probability, 15:19:605–19:635, 2010. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/767.

Jagannath:2016:ODB

[Jag16] Aukosh Jagannath. On the overlap distribution of branching random walks. Electronic Journal of Probability, 21(??):50:1–50:16, ???? 2016. CODEN

???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.ejp/
1470316405.

Jakubowski:1997:NST

[Jak97] Adam Jakubowski. A non-Skorohod topology on the Skorohod space. Electronic Journal of Probability, 2:4:1–4:21, 1997. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/18.

Jalowy:2021:RCP

[Jal21] Jonas Jalowy. Rate of convergence for products of independent non-Hermitian random matrices. Electronic Journal of Probability, 26(18):1–24, February 2021.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-26/issue-none/Rate-of-convergence-
for-products-of-independent-non-Hermitian-random/10.1214/21-EJP625.
full.

Janson:2009:PRG

[Jan09] Svante Janson. On percolation in random graphs with given vertex degrees. Electronic Journal of Probability, 14:5:86–5:118, 2009. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/603.

Jansen:2016:CPG

[Jan16] Sabine Jansen. Continuum percolation for Gibbsian point processes with attractive interactions. Electronic Journal of Probability, 21(??):47:1–47:22, ???? 2016. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

euclid.ejp/1469720442.

Janson:2019:HPR

[Jan19] Svante Janson. The hiring problem with rank-based strategies. Electronic Journal of Probability, 24(??):125:1–125:35, ???? 2019. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/euclid.ejp/1573268589.

Janson:2022:CLT

[Jan22] Svante Janson. Central limit theorems for additive functionals and fringe trees in tries. Electronic Journal of Probability, 27(??):1–63, 2022. CODEN ????

ISSN 1083-6489. URL https://projecteuclid.org/journals/electronic-
journal-of-probability/volume-27/issue-none/Central-limit-theorems-
for-additive-functionals-and-fringe-trees-in/10.1214/22-EJP776.full.

Jean:2005:RWG

[Jea05] Mairesse Jean. Random walks on groups and monoids with a Markovian harmonic measure. Electronic Journal of Probability, 10:43:1417–43:1441, 2005.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
293.

Jego:2020:TPR

[Jeg20] Antoine Jego. Thick points of random walk and the Gaussian free field. Electronic Journal of Probability, 25(??):32:1–32:39, ???? 2020. CODEN ???? ISSN

1083-6489. URL https://projecteuclid.org/euclid.ejp/1582858936.

Joulin:2013:MCT

[JG13] Aldéric Joulin and Arnaud Guillin. Measure concentration through non-Lipschitz observables and functional inequalities. Electronic Journal of Probability, 18:65:1–65:26, 2013. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/2425.

Jarvenpaa:2017:HPR

[JJK^{+}17] Esa Järvenpää, Maarit Järvenpää, Henna Koivusalo, Bing Li, Ville Suomala, and Yimin Xiao. Hitting probabilities of random covering sets in tori and metric spaces. Electronic Journal of Probability, 22(??):1:1–1:18, ???? 2017. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.ejp/
1483585523.

Janson:2011:TEL

[JK11] Svante Janson and Götz Kersting. On the total external length of the Kingman coalescent. Electronic Journal of Probability, 16:80:2203–80:2218, 2011. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/955.

Jahnel:2014:SDM

[JK14] Benedikt Jahnel and Christof Külske. Synchronization for discrete mean-field rotators. Electronic Journal of Probability, 19:14:1–14:26, 2014. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/2948.

Juszczyszyn:2015:HTP

[JK15] Tomasz Juszczyszyn and Mateusz Kwaśnicki. Hitting times of points for symmetric Lévy processes with completely monotone jumps. Electronic Journal of Probability, 20(??):48:1–48:24, ???? 2015. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/3440.

Jourdain:2019:NAE

[JK19] Benjamin Jourdain and Ahmed Kebaier. Non-asymptotic error bounds for the multilevel Monte Carlo Euler method applied to SDEs with constant diffusion coefficient. Electronic Journal of Probability, 24(??):12:1–12:34, ???? 2019.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.
ejp/1550653271.

Janson:2007:TEB

[JL07] Svante Janson and Guy Louchard. Tail estimates for the Brownian excursion area and other Brownian areas. Electronic Journal of Probability, 12:58:1600– 58:1632, 2007. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/471.

Jiao:2015:GDA

[JL15] Ying Jiao and Shanqiu Li. Generalized density approach in progressive enlargement of filtrations. Electronic Journal of Probability, 20(??):85:1–85:21, ???? 2015. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/3296.

Janson:2010:MBM

[JLML10] Svante Janson, Guy Louchard, and Anders Martin-Löf. The maximum of Brownian motion with parabolic drift. Electronic Journal of Probability, 15:61:1893– 61:1929, 2010. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/830.

Jerison:2013:IDH

[JLS13] David Jerison, Lionel Levine, and Scott Sheffield. Internal DLA in higher dimensions. Electronic Journal of Probability, 18:98:1–98:14, 2013. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/3137.

Jacka:2015:CTR

[JM15] Saul Jacka and Aleksandar Mijatovic. Coupling and tracking of regime-switching martingales. Electronic Journal of Probability, 20(??):38:1–38:39, ???? 2015.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
2307.

Jourdain:2020:NFO

[JM20] B. Jourdain and W. Margheriti. A new family of one dimensional martingale couplings. Electronic Journal of Probability, 25(??):1–50, 2020. CODEN ????

ISSN 1083-6489. URL https://projecteuclid.org/journals/electronic-
journal-of-probability/volume-25/issue-none/A-new-family-of-one-dimensional-
martingale-couplings/10.1214/20-EJP543.full.

Journel:2022:CKA

[JM22] Lucas Journel and Pierre Monmarché. Convergence of the kinetic annealing for general potentials. Electronic Journal of Probability, 27(??):1–37, 2022.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-27/issue-none/Convergence-
of-the-kinetic-annealing-for-general-potentials/10.1214/22-EJP891.
full.

Johansson:2006:EGM

[JN06] Kurt Johansson and Eric Nordenstam. Eigenvalues of GUE minors. Electronic Journal of Probability, 11:50:1342–50:1371, 2006. CODEN ???? ISSN 1083-

6489. URL http://ejp.ejpecp.org/article/view/370. See erratum [JN07].

Johansson:2007:EEG

[JN07] Kurt Johansson and Eric Nordenstam. Erratum to “Eigenvalues of GUE Minors”. Electronic Journal of Probability, 12:37:1048–37:1051, 2007. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/816. See

[JN06].

Jakubowski:2010:CDS

[JN10] Jecek Jakubowski and Mariusz Nieweglowski. A class of F -doubly stochastic Markov chains. Electronic Journal of Probability, 15:56:1743–56:1771, 2010.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
815.

Johnston:2019:GGW

[Joh19] Samuel G. G. Johnston. The genealogy of Galton–Watson trees. Electronic Journal of Probability, 24(??):94:1–94:35, ???? 2019. CODEN ???? ISSN

1083-6489. URL https://projecteuclid.org/euclid.ejp/1568361635.

Jonasson:2001:DPM

[Jon01] Johan Jonasson. On disagreement percolation and maximality of the critical value for iid percolation. Electronic Journal of Probability, 6:15:1–15:13, 2001.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
88.

Jonasson:2011:MTB

[Jon11] Johan Jonasson. Mixing time bounds for overlapping cycles shuffles. Electronic Journal of Probability, 16:46:1281–46:1295, 2011. CODEN ???? ISSN 1083-

6489. URL http://ejp.ejpecp.org/article/view/912.

Jordan:2011:RRG

[Jor11] Jonathan Jordan. Randomised reproducing graphs. Electronic Journal of Probability, 16:57:1549–57:1562, 2011. CODEN ???? ISSN 1083-6489. URL http:/

/ejp.ejpecp.org/article/view/921.

Jordan:2013:GPA

[Jor13] Jonathan Jordan. Geometric preferential attachment in non-uniform metric spaces. Electronic Journal of Probability, 18:8:1–8:15, 2013. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/2271.

Joseph:2009:FQM

[Jos09] Mathew Joseph. Fluctuations of the quenched mean of a planar random walk in an i.i.d. random environment with forbidden direction. Electronic Journal of Probability, 14:44:1268–44:1289, 2009. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/655.

Janson:2018:MCB

[JP18] Svante Janson and Nicolas Pouyanne. Moment convergence of balanced Pólya processes. Electronic Journal of Probability, 23(??):34:1–34:13, ???? 2018. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.ejp/
1524880978.

Jaramillo:2019:CES

[JPP19] Arturo Jaramillo, Juan Carlos Pardo, and José Luis Pérez. Convergence of the empirical spectral distribution of Gaussian matrix-valued processes. Electronic Journal of Probability, 24(??):10:1–10:22, ???? 2019. CODEN ???? ISSN

1083-6489. URL https://projecteuclid.org/euclid.ejp/1550286034.

Jourdain:2014:SNL

[JR14] Benjamin Jourdain and Julien Reygner. The small noise limit of order-based diffusion processes. Electronic Journal of Probability, 19:29:1–29:36, 2014. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/2906.

Junnila:2017:UCG

[JS17] Janne Junnila and Eero Saksman. Uniqueness of critical Gaussian chaos. Electronic Journal of Probability, 22(??):11:1–11:31, ???? 2017. CODEN ???? ISSN

1083-6489. URL https://projecteuclid.org/euclid.ejp/1486090891.

Joyner:2018:RWA

[JS18] Christopher H. Joyner and Uzy Smilansky. A random walk approach to linear statistics in random tournament ensembles. Electronic Journal of Probability, 23(??):80:1–80:37, ???? 2018. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1536717739.

Jelito:2022:RSO

[JS22] Damian Jelito and Łukasz Stettner. Risk-sensitive optimal stopping with unbounded terminal cost function. Electronic Journal of Probability, 27(??):1–30, 2022. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

journals/electronic-journal-of-probability/volume-27/issue-none/Risk-
sensitive-optimal-stopping-with-unbounded-terminal-cost-function/
10.1214/21-EJP736.full.

Johnson:2017:LLF

[JSS17] Tobias Johnson, Anne Schilling, and Erik Slivken. Local limit of the fixed point forest. Electronic Journal of Probability, 22(??):18:1–18:26, ???? 2017.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.
ejp/1487127646.

Jourdain:2021:CLT

[JT21] Benjamin Jourdain and Alvin Tse. Central limit theorem over non-linear functionals of empirical measures with applications to the mean-field fluctuation of interacting diffusions. Electronic Journal of Probability, 26(18):1–34, February 2021. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

journals/electronic-journal-of-probability/volume-26/issue-none/Central-
limit-theorem-over-non-linear-functionals-of-empirical-measures/10.
1214/21-EJP720.full.

Juillet:2018:MAP

[Jui18] Nicolas Juillet. Martingales associated to peacocks using the curtain coupling. Electronic Journal of Probability, 23(??):8:1–8:29, ???? 2018. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.ejp/
1518426055.

Junge:2015:CIE

[Jun15] Matthew Junge. Choices, intervals and equidistribution. Electronic Journal of Probability, 20(??):97:1–97:18, ???? 2015. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/4191.

Jacquot:2011:BSL

[JV11] Stephanie Jacquot and Benedek Valko. Bulk scaling limit of the Laguerre ensemble. Electronic Journal of Probability, 16:11:314–11:346, 2011. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/854.

Jacka:2001:ECN

[JW01] Saul Jacka and Jon Warren. Examples of convergence and non-convergence of Markov chains conditioned not to die. Electronic Journal of Probability, 7: 1:1–1:22, 2001. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/100.

Janson:2018:CPP

[JW18] Svante Janson and Lutz Warnke. On the critical probability in percolation. Electronic Journal of Probability, 23(??):1:1–1:25, ???? 2018. CODEN ???? ISSN

1083-6489. URL https://projecteuclid.org/euclid.ejp/1515726029.

Kalinin:2021:SCR

[Kal21] Alexander Kalinin. Support characterization for regular path-dependent stochastic Volterra integral equations. Electronic Journal of Probability, 26(??): 1–29, 2021. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.

org/journals/electronic-journal-of-probability/volume-26/issue-none/
Support-characterization-for-regular-path-dependent-stochastic-Volterra-
integral-equations/10.1214/20-EJP576.full.

Kammoun:2018:MSD

[Kam18] Mohamed Slim Kammoun. Monotonous subsequences and the descent process of invariant random permutations. Electronic Journal of Probability, 23(??):

118:1–118:31, ???? 2018. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1543287754.

Kargin:2015:LTL

[Kar15] Vladislav Kargin. Limit theorems for linear eigenvalue statistics of overlapping matrices. Electronic Journal of Probability, 20(??):121:1–121:30, ???? 2015.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
3937.

Karrila:2020:UBM

[Kar20] Alex Karrila. UST branches, martingales, and multiple SLE(2). Electronic Journal of Probability, 25(??):1–37, 2020. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-25/issue-none/UST-branches-martingales-and-
multiple-SLE2/10.1214/20-EJP485.full.

Kendall:2020:RRF

[Ken20] Wilfrid S. Kendall. Rayleigh random flights on the Poisson line SIRSN. Electronic Journal of Probability, 25(??):1–36, 2020. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-25/issue-none/Rayleigh-Random-Flights-on-the-
Poisson-line-SIRSN/10.1214/20-EJP526.full.

Konig:2013:MAB

[KGS13] Wolfgang König, Onur Gün, and Ozren Sekulović. Moment asymptotics for branching random walks in random environment. Electronic Journal of Probability, 18:63:1–63:18, 2013. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/2212.

Khandwawala:2014:BPM

[Kha14] Mustafa Khandwawala. Belief propagation for minimum weight many-to-one matchings in the random complete graph. Electronic Journal of Probability, 19 (??):112:1–112:40, ???? 2014. CODEN ???? ISSN 1083-6489. URL http://

ejp.ejpecp.org/article/view/3491.

Khoshnevisan:1996:LCS

[Kho96] Davar Khoshnevisan. Lévy classes and self-normalization. Electronic Journal of Probability, 1:1:1–1:18, 1996. CODEN ???? ISSN 1083-

6489. URL http://ejp.ejpecp.org/article/view/1;   http://www.math.
washington.edu/~ejpecp/EjpVol1/paper1.abs.html.

Kifer:2017:FER

[Kif17] Yuri Kifer. Functional Erdős–Rényi law of large numbers for nonconventional sums under weak dependence. Electronic Journal of Probability, 22 (??):23:1–23:17, ???? 2017. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1488337348.

Kim:2005:ESD

[Kim05] Kyeong-Hun Kim. L_{p}-estimates for SPDE with discontinuous coefficients in domains. Electronic Journal of Probability, 10:1:1–1:20, 2005. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/234.

Kim:2006:PSD

[Kim06] Kyeong-Hun Kim. Parabolic SPDEs degenerating on the boundary of nonsmooth domain. Electronic Journal of Probability, 11:23:563–23:584, 2006. CO-

DEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
339.

Knopova:2011:EAD

[KK11] Victoria Knopova and Alexei Kulik. Exact asymptotic for distribution densities of Lévy functionals. Electronic Journal of Probability, 16:52:1394–52:1433, 2011.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
909.

Kleptsyn:2012:ESA

[KK12] Victor Kleptsyn and Aline Kurtzmann. Ergodicity of self-attracting motion. Electronic Journal of Probability, 17:50:1–50:37, 2012. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/2121.

Kuznetsov:2018:SAS

[KK18] Alexey Kuznetsov and Mateusz Kwaśnicki. Spectral analysis of stable processes on the positive half-line. Electronic Journal of Probability, 23(??):10:1–10:29, ???? 2018. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.

org/euclid.ejp/1518426058.

Kallsen:2020:USM

[KK20a] Jan Kallsen and Paul Krühner. On uniqueness of solutions to martingale problems — counterexamples and sufficient criteria. Electronic Journal of Probability, 25(??):1–33, 2020. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-25/issue-none/On-uniqueness-of-solutions-to-
martingale-problems--counterexamples-and/10.1214/20-EJP494.full.

Kim:2020:RDP

[KK20b] Daehong Kim and Seiichiro Kusuoka. Recurrence of direct products of diffusion processes in random media having zero potentials. Electronic Journal of Probability, 25(??):1–18, 2020. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-25/issue-none/Recurrence-of-direct-products-
of-diffusion-processes-in-random-media/10.1214/20-EJP540.full.

Kijung:2011:TSP

[KKH11] Lee Kijung and Kim Kyeong-Hun. A W_{2}^{1}-theory of stochastic partial differential systems of divergence type on C^{1} domains. Electronic Journal of Probability, 16:47:1296–47:1317, 2011. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/913.

Kondratiev:2013:SGG

[KKO13] Yuri Kondratiev, Tobias Kuna, and Natascha Ohlerich. Spectral gap for Glauber type dynamics for a special class of potentials. Electronic Journal of Probability, 18:42:1–42:18, 2013. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.

org/article/view/2260.

Kuznetsov:2014:HTZ

[KKPW14] Alexey Kuznetsov, Andreas Kyprianou, Juan Carlos Pardo, and Alexander Watson. The hitting time of zero for a stable process. Electronic Journal of Probability, 19:30:1–30:26, 2014. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/2647.

Kozdron:2005:ERW

[KL05] Michael Kozdron and Gregory Lawler. Estimates of random walk exit probabilities and application to loop-erased random walk. Electronic Journal of Probability, 10:44:1442–44:1467, 2005. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/294.

Kruse:2012:ORS

[KL12] Raphael Kruse and Stig Larsson. Optimal regularity for semilinear stochastic partial differential equations with multiplicative noise. Electronic Journal of Probability, 17:65:1–65:19, 2012. CODEN ???? ISSN 1083-6489. URL http://

ejp.ejpecp.org/article/view/2240.

Kliem:2015:EMF

[KL15] Sandra Kliem and Wolfgang Loehr. Existence of mark functions in marked metric measure spaces. Electronic Journal of Probability, 20(??):73:1–73:24, ???? 2015. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/3969.

Kruhner:2018:APC

[KL18] Paul Krühner and Martin Larsson. Affine processes with compact state space. Electronic Journal of Probability, 23(??):29:1–29:23, ???? 2018. CODEN

???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.ejp/
1522375269.

Konarovskyi:2021:SBM

[KL21] Vitalii Konarovskyi and Vlada Limic. Stochastic block model in a new critical regime and the interacting multiplicative coalescent. Electronic Journal of Probability, 26(??):1–23, 2021. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-26/issue-none/Stochastic-block-model-in-a-new-
critical-regime-and-the/10.1214/21-EJP584.full.

Kolb:2022:NEF

[KL22] Martin Kolb and Matthias Liesenfeld. On non-extinction in a Fleming–Viot-type particle model with Bessel drift. Electronic Journal of Probability, 27(??):1–28, 2022. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

journals/electronic-journal-of-probability/volume-27/issue-none/On-
non-extinction-in-a-Fleming-Viot-type-particle-model/10.1214/22-EJP866.
full.

Kliem:2011:CRC

[Kli11] Sandra Kliem. Convergence of rescaled competing species processes to a class of SPDEs. Electronic Journal of Probability, 16:22:618–22:657, 2011. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/870.

Klimsiak:2012:RBM

[Kli12] Tomasz Klimsiak. Reflected BSDEs with monotone generator. Electronic Journal of Probability, 17:107:1–107:25, 2012. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/1759.

Kliem:2014:CCR

[Kli14] Sandra Kliem. A compact containment result for nonlinear historical superprocess approximations for population models with trait-dependence. Electronic Journal of Probability, 19(??):97:1–97:13, ???? 2014. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/3506.

Kontoyiannis:2005:LDA

[KM05] Ioannis Kontoyiannis and Sean Meyn. Large deviations asymptotics and the spectral theory of multiplicatively regular Markov processes. Electronic Journal of Probability, 10:3:61–3:123, 2005. CODEN ???? ISSN 1083-6489. URL http:

//ejp.ejpecp.org/article/view/231.

Kevei:2011:CMB

[KM11] Peter Kevei and Jose Lopez Mimbela. Critical multitype branching systems: Extinction results. Electronic Journal of Probability, 16:50:1356–50:1380, 2011.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
908.

Kevei:2012:ADR

[KM12] Peter Kevei and David Mason. The asymptotic distribution of randomly weighted sums and self-normalized sums. Electronic Journal of Probability, 17: 46:1–46:21, 2012. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.

org/article/view/2092.

Kolesko:2015:FPM

[KM15] Konrad Kolesko and Sebastian Mentemeier. Fixed points of the multivariate smoothing transform: the critical case. Electronic Journal of Probability, 20 (??):52:1–52:24, ???? 2015. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/4022.

Kortchemski:2016:TSL

[KM16] Igor Kortchemski and Cyril Marzouk. Triangulating stable laminations. Electronic Journal of Probability, 21(??):11:1–11:31, ???? 2016. CODEN ???? ISSN

1083-6489. URL https://projecteuclid.org/euclid.ejp/1455559938.

Konakov:2017:WEE

[KM17] Valentin Konakov and Stéphane Menozzi. Weak error for the Euler scheme approximation of diffusions with non-smooth coefficients. Electronic Journal of Probability, 22(??):46:1–46:47, ???? 2017. CODEN ???? ISSN 1083-6489. URL

https://projecteuclid.org/euclid.ejp/1494489631.

Kim:2018:EDH

[KM18] Panki Kim and Ante Mimica. Estimates of Dirichlet heat kernels for subordinate Brownian motions. Electronic Journal of Probability, 23(??):64:1–64:45, ???? 2018. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

euclid.ejp/1532570592.

Klein:2006:CCI

[KMP06] Thierry Klein, Yutao Ma, and Nicolas Privault. Convex concentration inequalities and forward–backward stochastic calculus. Electronic Journal of Probability, 11:20:486–20:512, 2006. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/332.

Koval:2012:LRP

[KMT12] Vyacheslav Koval, Ronald Meester, and Pieter Trapman. Long-range percolation on the hierarchical lattice. Electronic Journal of Probability, 17:57:1–57:21, 2012. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/

view/1977.

Klass:2007:UAQ

[KN07] Michael Klass and Krzysztof Nowicki. Uniformly accurate quantile bounds via the truncated moment generating function: The symmetric case. Electronic Journal of Probability, 12:47:1276–47:1298, 2007. CODEN ???? ISSN 1083-

6489. URL http://ejp.ejpecp.org/article/view/452.

Kotani:2017:PSD

[KN17] Shinichi Kotani and Fumihiko Nakano. Poisson statistics for 1 d Schrödinger operators with random decaying potentials. Electronic Journal of Probability, 22(??):69:1–69:31, ???? 2017. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1504922532.

Kumar:2021:EMT

[KN21] Chaman Kumar and Neelima. On explicit Milstein-type scheme for McKean– Vlasov stochastic differential equations with super-linear drift coefficient. Electronic Journal of Probability, 26(18):1–32, February 2021. CODEN ????

ISSN 1083-6489. URL https://projecteuclid.org/journals/electronic-
journal-of-probability/volume-26/issue-none/On-explicit-Milstein-
type-scheme-for-McKeanVlasov-stochastic-differential-equations/10.
1214/21-EJP676.full.

Kim:2021:AON

[KNR21] Edward Kim, Tianyang Nie, and Marek Rutkowski. American options in nonlinear markets. Electronic Journal of Probability, 26(18):1–41, February 2021.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-26/issue-none/American-options-
in-nonlinear-markets/10.1214/21-EJP658.full.

Kulske:2008:PMG

[KO08] Christof Külske and Alex Opoku. The posterior metric and the goodness of Gibbsianness for transforms of Gibbs measures. Electronic Journal of Probability, 13:47:1307–47:1344, 2008. CODEN ???? ISSN 1083-6489. URL http://

ejp.ejpecp.org/article/view/560.

Ko:2020:FEM

[Ko20] Justin Ko. Free energy of multiple systems of spherical spin glasses with constrained overlaps. Electronic Journal of Probability, 25(??):28:1–28:34, ????

2020. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

euclid.ejp/1582254384.

Kolokoltsov:2001:SDF

[Kol01] Vassili Kolokoltsov. Small diffusion and fast dying out asymptotics for superprocesses as non-Hamiltonian quasiclassics for evolution equations. Electronic Journal of Probability, 6:21:1–21:16, 2001. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/94.

Kolesnik:2022:S

[Kol22] Brett Kolesnik. The sharp K_{4}-percolation threshold on the Erdős–Rényi random graph. Electronic Journal of Probability, 27(??):1–23, 2022. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-27/issue-none/The-sharp-
K4-percolation-threshold-on-the-Erd%c5%91sR%c3%a9nyi-random-graph/
10.1214/21-EJP710.full.

Konarovskyi:2017:ABM

[Kon17] Vitalii Konarovskyi. On asymptotic behavior of the modified Arratia flow. Electronic Journal of Probability, 22(??):19:1–19:31, ???? 2017. CODEN ???? ISSN

1083-6489. URL https://projecteuclid.org/euclid.ejp/1487386997.

Kozachenko:2013:CGW

[KOP13] Yuriy Kozachenko, Andriy Olenko, and Olga Polosmak. On convergence of general wavelet decompositions of nonstationary stochastic processes. Electronic Journal of Probability, 18:69:1–69:21, 2013. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/2234.

Konig:2001:NCR

[KOR01] Wolfgang König, Neil O’Connell, and Sébastien Roch. Non-colliding random walks, tandem queues, and discrete orthogonal polynomial ensembles. Electronic Journal of Probability, 7:5:1–5:24, 2001. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/104.

Komorowski:2021:HLC

[KOS21] Tomasz Komorowski, Stefano Olla, and Marielle Simon. Hydrodynamic limit for a chain with thermal and mechanical boundary forces. Electronic Journal of Probability, 26(??):1–49, 2021. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-26/issue-none/Hydrodynamic-limit-for-a-chain-
with-thermal-and-mechanical-boundary/10.1214/21-EJP581.full.

Kendall:2004:CIK

[KP04] Wilfrid Kendall and Catherine Price. Coupling iterated Kolmogorov diffusions. Electronic Journal of Probability, 9:13:382–13:410, 2004. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/201.

Komjathy:2013:MRT

[KP13] Júlia Komjáthy and Yuval Peres. Mixing and relaxation time for random walk on wreath product graphs. Electronic Journal of Probability, 18:71:1–71:23, 2013.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
2321.

Kosygina:2016:FLL

[KP16] Elena Kosygina and Jonathon Peterson. Functional limit laws for recurrent excited random walks with periodic cookie stacks. Electronic Journal of Probability, 21(??):70:1–70:24, ???? 2016. CODEN ???? ISSN 1083-6489. URL

https://projecteuclid.org/euclid.ejp/1480688087.

Kozma:2021:PLD

[KP21] Gady Kozma and Ron Peled. Power-law decay of weights and recurrence of the two-dimensional VRJP. Electronic Journal of Probability, 26(18):1–19, February 2021. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

journals/electronic-journal-of-probability/volume-26/issue-none/Power-
law-decay-of-weights-and-recurrence-of-the-two/10.1214/21-EJP639.
full.

Kaleta:2022:DCE

[KP22] Kamil Kaleta and Daniel Ponikowski. On directional convolution equivalent densities. Electronic Journal of Probability, 27(??):1–19, 2022. CODEN ????

ISSN 1083-6489. URL https://projecteuclid.org/journals/electronic-
journal-of-probability/volume-27/issue-none/On-directional-convolution-
equivalent-densities/10.1214/22-EJP790.full.

Kuznetsov:2012:DPE

[KPS12] Alexey Kuznetsov, Juan Carlos Pardo, and Mladen Savov. Distributional properties of exponential functionals of Lévy processes. Electronic Journal of Probability, 17:8:1–8:35, 2012. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/1755.

Kurkova:2004:ISC

[KPV04] Irina Kurkova, Serguei Popov, and M. Vachkovskaia. On infection spreading and competition between independent random walks. Electronic Journal of Probability, 9:11:293–11:315, 2004. CODEN ???? ISSN 1083-6489. URL http:

//ejp.ejpecp.org/article/view/197.

Khoshnevisan:2000:LRF

[KPX00] Davar Khoshnevisan, Yuval Peres, and Yimin Xiao. Limsup random fractals. Electronic Journal of Probability, 5:4:1–4:24, 2000. CODEN ???? ISSN 1083-

6489. URL http://ejp.ejpecp.org/article/view/60.

Kobylanski:2012:OST

[KQ12] Magdalena Kobylanski and Marie-Claire Quenez. Optimal stopping time problem in a general framework. Electronic Journal of Probability, 17:72:1–72:28, 2012. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/

view/2262.

Kyprianou:2008:SCC

[KR08] Andreas Kyprianou and Victor Rivero. Special, conjugate and complete scale functions for spectrally negative Lévy processes. Electronic Journal of Probability, 13:57:1672–57:1701, 2008. CODEN ???? ISSN 1083-6489. URL http://

ejp.ejpecp.org/article/view/567.

Kumar:2020:SCP

[KR20] Umesh Kumar and Markus Riedle. The stochastic Cauchy problem driven by a cylindrical Lévy process. Electronic Journal of Probability, 25(??):10:1–10:26, ???? 2020. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.

org/euclid.ejp/1580267010.

Kaur:2021:HOF

[KR21a] Gursharn Kaur and Adrian Röllin. Higher-order fluctuations in dense random graph models. Electronic Journal of Probability, 26(18):1–36, February 2021.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-26/issue-none/Higher-order-
fluctuations-in-dense-random-graph-models/10.1214/21-EJP708.full.

Klimsiak:2021:RBT

[KR21b] Tomasz Klimsiak and Maurycy Rzymowski. Reflected BSDEs with two optional barriers and monotone coefficient on general filtered space. Electronic Journal of Probability, 26(18):1–24, February 2021. CODEN ????

ISSN 1083-6489. URL https://projecteuclid.org/journals/electronic-
journal-of-probability/volume-26/issue-none/Reflected-BSDEs-with-
two-optional-barriers-and-monotone-coefficient-on/10.1214/21-EJP655.
full.

Kozitsky:2021:MPI

[KR21c] Yuri Kozitsky and Michael Röckner. A Markov process for an infinite interacting particle system in the continuum. Electronic Journal of Probability, 26(18):1–53, February 2021. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-26/issue-none/A-Markov-process-for-an-infinite-
interacting-particle-system-in/10.1214/21-EJP631.full.

Kraaij:2013:SPM

[Kra13] Richard Kraaij. Stationary product measures for conservative particle systems and ergodicity criteria. Electronic Journal of Probability, 18:88:1–88:33, 2013.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
2513.

Kraaij:2020:ERM

[Kra20] Richard C. Kraaij. The exponential resolvent of a Markov process and large deviations for Markov processes via Hamilton–Jacobi equations. Electronic Journal of Probability, 25(??):1–39, 2020. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-25/issue-none/The-exponential-resolvent-of-
a-Markov-process-and-large-deviations/10.1214/20-EJP539.full.

Keller-Ressel:2013:RAP

[KRST13] Martin Keller-Ressel, Walter Schachermayer, and Josef Teichmann. Regularity of affine processes on general state spaces. Electronic Journal of Probability, 18: 43:1–43:17, 2013. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.

org/article/view/2043.

Krylov:1999:AVF

[Kry99] N. V. Krylov. Approximating value functions for controlled degenerate diffusion processes by using piece-wise constant policies. Electronic Journal of Probability, 4:2:1–2:19, 1999. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.

org/article/view/39.

Krylov:2000:SS

[Kry00] N. Krylov. SPDEs in L_{q}((0, τ], L_{p}) spaces. Electronic Journal of Probability, 5: 13:1–13:29, 2000. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.

org/article/view/69.

Krylov:2004:QIS

[Kry04] N. V. Krylov. Quasiderivatives and interior smoothness of harmonic functions associated with degenerate diffusion processes. Electronic Journal of Probability, 9:21:615–21:633, 2004. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/219.

Kager:1997:GOS

[KS97] Gerald Kager and Michael Scheutzow. Generation of one-sided random dynamical systems by stochastic differential equations. Electronic Journal of Probability, 2:8:1–8:17, 1997. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/22.

Kurtz:2001:SSF

[KS01] Thomas Kurtz and Richard Stockbridge. Stationary solutions and forward equations for controlled and singular martingale problems. Electronic Journal of Probability, 6:17:1–17:52, 2001. CODEN ???? ISSN 1083-6489. URL http://

ejp.ejpecp.org/article/view/90.

Kesten:2003:BRW

[KS03] Harry Kesten and Vladas Sidoravicius. Branching random walk with catalysts. Electronic Journal of Probability, 8:5:1–5:51, 2003. CODEN ???? ISSN 1083-

6489. URL http://ejp.ejpecp.org/article/view/127.

Kozma:2004:AED

[KS04] Gady Kozma and Ehud Schreiber. An asymptotic expansion for the discrete harmonic potential. Electronic Journal of Probability, 9:1:1–1:17, 2004. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/170.

Kolb:2017:CSD

[KS17a] Martin Kolb and Mladen Savov. Conditional survival distributions of Brownian trajectories in a one dimensional Poissonian environment in the critical case. Electronic Journal of Probability, 22(??):14:1–14:29, ???? 2017. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.ejp/
1487127642.

Kumar:2017:EAL

[KS17b] Chaman Kumar and Sotirios Sabanis. On explicit approximations for Lévy driven SDEs with super-linear diffusion coefficients. Electronic Journal of Probability, 22(??):73:1–73:19, ???? 2017. CODEN ???? ISSN 1083-6489. URL

https://projecteuclid.org/euclid.ejp/1505268104.

Kabluchko:2019:DBZ

[KS19] Zakhar Kabluchko and Hauke Seidel. Distances between zeroes and critical points for random polynomials with i.i.d. zeroes. Electronic Journal of Probability, 24(??):34:1–34:25, ???? 2019. CODEN ???? ISSN 1083-6489. URL

https://projecteuclid.org/euclid.ejp/1554775413.

Kopytko:2020:ODD

[KS20] Bohdan Kopytko and Roman Shevchuk. One-dimensional diffusion processes with moving membrane: partial reflection in combination with jump-like exit of process from membrane. Electronic Journal of Probability, 25(??):41:1–41:21, ???? 2020. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.

org/euclid.ejp/1585620096.

Kolokoltsov:2002:TNE

[KST02] Vassili Kolokoltsov, R. L. Schilling, and A. Tyukov. Transience and nonexplosion of certain stochastic Newtonian systems. Electronic Journal of Probability, 7:19:1–19:19, 2002. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/118.

Kesten:2022:OPR

[KSV22] Harry Kesten, Vladas Sidoravicius, and Maria Eulália Vares. Oriented percolation in a random environment. Electronic Journal of Probability, 27(??):1–49, 2022. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

journals/electronic-journal-of-probability/volume-27/issue-none/Oriented-
percolation-in-a-random-environment/10.1214/22-EJP791.full.

Kersting:2014:EBC

[KSW14] Götz Kersting, Jason Schweinsberg, and Anton Wakolbinger. The evolving beta coalescent. Electronic Journal of Probability, 19:64:1–64:27, 2014. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/3332.

Kouritzin:2004:NFR

[KSX04] Michael Kouritzin, Wei Sun, and Jie Xiong. Nonliner filtering for reflecting diffusions in random environments via nonparametric estimation. Electronic Journal of Probability, 9:18:560–18:574, 2004. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/214. See erratum [KSX17].

Kouritzin:2017:ENF

[KSX17] Michael A. Kouritzin, Wei Sun, and Jie Xiong. Erratum: Nonlinear filtering for reflecting diffusions in random environments via nonparametric estimation. Electronic Journal of Probability, 22(??):3:1–3:2, ???? 2017. CODEN ???? ISSN

1083-6489. URL https://projecteuclid.org/euclid.ejp/1483585525. See

[KSX04].

Kesten:1998:AAW

[KSZ98] Harry Kesten, Vladas Sidoravicius, and Yu Zhang. Almost all words are seen in critical site percolation on the triangular lattice. Electronic Journal of Probability, 3:10:1–10:75, 1998. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/32.

Kesten:2001:PAW

[KSZ01] Harry Kesten, Vladas Sidoravicius, and Yu Zhang. Percolation of arbitrary words on the close-packed graph of Z^{2}. Electronic Journal of Probability, 6: 4:1–4:27, 2001. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.

org/article/view/77.

Krokowski:2017:MCL

[KT17] Kai Krokowski and Christoph Thäle. Multivariate central limit theorems for Rademacher functionals with applications. Electronic Journal of Probability, 22 (??):87:1–87:30, ???? 2017. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1508292258.

Kabluchko:2021:FRG

[KT21a] Zakhar Kabluchko and Christoph Thäle. Faces in random great hypersphere tessellations. Electronic Journal of Probability, 26(??):1–35, 2021. CODEN ????

ISSN 1083-6489. URL https://projecteuclid.org/journals/electronic-
journal-of-probability/volume-26/issue-none/Faces-in-random-great-
hypersphere-tessellations/10.1214/20-EJP570.full.

Kai:2021:GFJ

[KT21b] Hirotaka Kai and Atsushi Takeuchi. Gradient formulas for jump processes on manifolds. Electronic Journal of Probability, 26(18):1–15, February 2021.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-26/issue-none/Gradient-formulas-
for-jump-processes-on-manifolds/10.1214/21-EJP660.full.

Kazi-Tani:2015:SOB

[KTPZ15] Nabil Kazi-Tani, Dylan Possamaï, and Chao Zhou. Second order BSDEs with jumps: existence and probabilistic representation for fully-nonlinear PIDEs. Electronic Journal of Probability, 20(??):65:1–65:31, ???? 2015. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/3569.

Kuan:2014:GFF

[Kua14] Jeffrey Kuan. The Gaussian free field in interlacing particle systems. Electronic Journal of Probability, 19:72:1–72:31, 2014. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/3732.

Kuhn:2018:MPF

[Küh18] Franziska Kühn. On martingale problems and Feller processes. Electronic Journal of Probability, 23(??):13:1–13:18, ???? 2018. CODEN ???? ISSN 1083-6489.

URL https://projecteuclid.org/euclid.ejp/1518426061.

Kuhn:2020:EMS

[Küh20] Franziska Kühn. Existence of (Markovian) solutions to martingale problems associated with Lévy-type operators. Electronic Journal of Probability, 25(??): 16:1–16:26, ???? 2020. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1580871682.

Kulik:2008:SEV

[Kul08] Rafal Kulik. Sums of extreme values of subordinated long-range dependent sequences: moving averages with finite variance. Electronic Journal of Probability, 13:32:961–32:979, 2008. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/510.

Kunze:2013:CMP

[Kun13] Markus Kunze. On a class of martingale problems on Banach spaces. Electronic Journal of Probability, 18:104:1–104:30, 2013. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/2924.

Kurtz:1998:MPC

[Kur98] Thomas G. Kurtz. Martingale problems for conditional distributions of Markov processes. Electronic Journal of Probability, 3:9:1–9:29, 1998. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/31.

Kurtz:2007:YWE

[Kur07] Thomas Kurtz. The Yamada–Watanabe–Engelbert theorem for general stochastic equations and inequalities. Electronic Journal of Probability, 12:33:951– 33:965, 2007. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/431.

Kurkova:2008:LES

[Kur08] Irina Kurkova. Local energy statistics in directed polymers. Electronic Journal of Probability, 13:2:5–2:25, 2008. CODEN ???? ISSN 1083-6489. URL http:/

/ejp.ejpecp.org/article/view/475.

Kuwada:2009:CMM

[Kuw09] Kazumasa Kuwada. Characterization of maximal Markovian couplings for diffusion processes. Electronic Journal of Probability, 14:25:633–25:662, 2009. CO-

DEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
634.

Kopp:2000:CIM

[KW00] P. Kopp and Volker Wellmann. Convergence in incomplete market models. Electronic Journal of Probability, 5:15:1–15:26, 2000. CODEN ???? ISSN 1083-

6489. URL http://ejp.ejpecp.org/article/view/71.

Kinnally:2010:EUS

[KW10] Michael Kinnally and Ruth Williams. On existence and uniqueness of stationary distributions for stochastic delay differential equations with positivity constraints. Electronic Journal of Probability, 15:15:409–15:451, 2010. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/756.

Kolb:2011:SGB

[KW11] Martin Kolb and Achim Wübker. On the spectral gap of Brownian motion with jump boundary. Electronic Journal of Probability, 16:43:1214–43:1237, 2011.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
903.

Klaassen:2021:HID

[KW21] Chris A. J. Klaassen and Jon A. Wellner. Hardy’s inequality and its descendants: a probability approach. Electronic Journal of Probability, 26(18):1–34, February 2021. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

journals/electronic-journal-of-probability/volume-26/issue-none/Hardys-
inequality-and-its-descendants-a-probability-approach/10.1214/21-
EJP711.full.

Kwasnicki:2012:STS

[Kwa12] Mateusz Kwaśnicki. Spectral theory for symmetric one-dimensional Lévy processes killed upon hitting the origin. Electronic Journal of Probability, 17:83:1– 83:29, 2012. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/2013.

Kwasnicki:2019:FTL

[Kwa19] Mateusz Kwaśnicki. Fluctuation theory for Lévy processes with completely monotone jumps. Electronic Journal of Probability, 24(??):40:1–40:40, ???? 2019. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

euclid.ejp/1555034439.

Kyprianou:2016:DFS

[Kyp16] Andreas E. Kyprianou. Deep factorisation of the stable process. Electronic Journal of Probability, 21(??):23:1–23:28, ???? 2016. CODEN ???? ISSN

1083-6489. URL https://projecteuclid.org/euclid.ejp/1459880111.

Kosygina:2008:PNE

[KZ08] Elena Kosygina and Martin Zerner. Positively and negatively excited random walks on integers, with branching processes. Electronic Journal of Probability, 13:64:1952–64:1979, 2008. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/572.

Kosygina:2014:EER

[KZ14] Elena Kosygina and Martin Zerner. Excursions of excited random walks on integers. Electronic Journal of Probability, 19:25:1–25:25, 2014. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/2940.

Klochkov:2020:UHW

[KZ20] Yegor Klochkov and Nikita Zhivotovskiy. Uniform Hanson-Wright type concentration inequalities for unbounded entries via the entropy method. Electronic Journal of Probability, 25(??):22:1–22:30, ???? 2020. CODEN ???? ISSN

1083-6489. URL https://projecteuclid.org/euclid.ejp/1581130826.

Linde:2009:SRF

[LA09] Werner Linde and Antoine Ayache. Series representations of fractional Gaussian processes by trigonometric and Haar systems. Electronic Journal of Probability, 14:94:2691–94:2719, 2009. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/727.

Labbe:2014:FFV

[Lab14] Cyril Labbé. From flows of Λ-Fleming–Viot processes to lookdown processes via flows of partitions. Electronic Journal of Probability, 19:55:1–55:49, 2014.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
3192.

Lachal:2003:DST

[Lac03] Aime Lachal. Distributions of sojourn time, maximum and minimum for pseudoprocesses governed by higher-order heat-type equations. Electronic Journal of Probability, 8:20:1–20:53, 2003. CODEN ???? ISSN 1083-6489. URL http://

ejp.ejpecp.org/article/view/178.

Lachal:2007:FHT

[Lac07] Aimé Lachal. First hitting time and place, monopoles and multipoles for pseudoprocesses driven by the equation ∂u/∂t = ±∂^{N}u/∂x^{N}. Electronic Journal of Probability, 12:11:300–11:353, 2007. CODEN ???? ISSN 1083-6489. URL http:

//ejp.ejpecp.org/article/view/399.

Lacoin:2012:EIP

[Lac12] Hubert Lacoin. Existence of an intermediate phase for oriented percolation. Electronic Journal of Probability, 17:41:1–41:17, 2012. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/1761.

Lacker:2022:QAI

[Lac22] Daniel Lacker. Quantitative approximate independence for continuous mean field Gibbs measures. Electronic Journal of Probability, 27(??):1–21, 2022. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-27/issue-none/Quantitative-
approximate-independence-for-continuous-mean-field-Gibbs-measures/
10.1214/22-EJP743.full.

Lambert:2001:JLA

[Lam01] Amaury Lambert. The joint law of ages and residual lifetimes for two schemes of regenerative sets. Electronic Journal of Probability, 6:19:1–19:23, 2001. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/92.

Lambert:2007:QSD

[Lam07] Amaury Lambert. Quasi-stationary distributions and the continuous-state branching process conditioned to be never extinct. Electronic Journal of Probability, 12:14:420–14:446, 2007. CODEN ???? ISSN 1083-6489. URL http://

ejp.ejpecp.org/article/view/402.

Lambert:2018:MFU

[Lam18] Gaultier Lambert. Mesoscopic fluctuations for unitary invariant ensembles. Electronic Journal of Probability, 23(??):7:1–7:33, ???? 2018. CODEN ???? ISSN

1083-6489. URL https://projecteuclid.org/euclid.ejp/1518426054.

Lamarre:2021:SOD

[Lam21a] Pierre Yves Gaudreau Lamarre. Semigroups for one-dimensional Schrödinger operators with multiplicative Gaussian noise. Electronic Journal of Probability, 26(18):1–47, February 2021. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/journals/electronic-journal-of-probability/volume-
26/issue-none/Semigroups-for-one-dimensional-Schr%c3%b6dinger-operators-
with-multiplicative-Gaussian-noise/10.1214/21-EJP654.full.

Lambert:2021:MCL

[Lam21b] Gaultier Lambert. Mesoscopic central limit theorem for the circular. Electronic Journal of Probability, 26(??):1–33, 2021. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-26/issue-none/Mesoscopic-central-limit-theorem-
for-the-circular-beta--ensembles/10.1214/20-EJP559.full.

Lanoue:2015:IM

[Lan15] Daniel Parmet Lanoue. The iPod model. Electronic Journal of Probability, 20 (??):64:1–64:20, ???? 2015. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/3559.

Larsson:2015:MVB

[Lar15] Martin Larsson. Matrix-valued Bessel processes. Electronic Journal of Probability, 20(??):60:1–60:29, ???? 2015. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/3785.

Laurent:2012:LDS

[Lau12] Clément Laurent. Large deviations for self-intersection local times in subcritical dimensions. Electronic Journal of Probability, 17:21:1–21:20, 2012. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/1874.

Lawler:1996:CTS

[Law96a] Gregory F. Lawler. Cut times for simple random walk. Electronic Journal of Probability, 1:13:1–13:24, 1996. CODEN ???? ISSN 1083-6489. URL http://

ejp.ejpecp.org/article/view/13.

Lawler:1996:HDC

[Law96b] Gregory F. Lawler. Hausdorff dimension of cut points for Brownian motion. Electronic Journal of Probability, 1:2:1–2:20, 1996. CODEN ???? ISSN 1083-

6489. URL http://ejp.ejpecp.org/article/view/2.

Lawler:2000:SCH

[Law00] Gregory Lawler. Strict concavity of the half plane intersection exponent for planar Brownian motion. Electronic Journal of Probability, 5:8:1–8:33, 2000.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
64.

Lawler:2020:ITS

[Law20] Gregory F. Lawler. The infinite two-sided loop-erased random walk. Electronic Journal of Probability, 25(??):1–42, 2020. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-25/issue-none/The-infinite-two-sided-loop-erased-
random-walk/10.1214/20-EJP476.full.

Lambert:2018:TOM

[LB18] Amaury Lambert and Gerónimo Uribe Bravo. Totally ordered measured trees and splitting trees with infinite variation. Electronic Journal of Probability, 23 (??):120:1–120:41, ???? 2018. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1544843300.

Lozada-Chang:2005:LDM

[LC05] Li-Vang Lozada-Chang. Large deviations on moment spaces. Electronic Journal of Probability, 10:19:662–19:690, 2005. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/202.

Lachal:2010:JDP

[LC10] Aimé Lachal and Valentina Cammarota. Joint distribution of the process and its sojourn time on the positive half-line for pseudo-processes governed by highorder heat equation. Electronic Journal of Probability, 15:28:895–28:931, 2010.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
782.

Le:2020:SSL

[Lê20] Khoa Lê. A stochastic sewing lemma and applications. Electronic Journal of Probability, 25(??):38:1–38:55, ???? 2020. CODEN ???? ISSN 1083-6489. URL

https://projecteuclid.org/euclid.ejp/1585620093.

Le:2023:SSB

[Lê23] Khoa Lê. Stochastic sewing in Banach spaces. Electronic Journal of Probability, 28(??):1–22, 2023. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/journals/electronic-journal-of-probability/volume-
28/issue-none/Stochastic-sewing-in-Banach-spaces/10.1214/23-EJP918.
full.

Ledoux:2004:DOS

[Led04] Michel Ledoux. Differential operators and spectral distributions of invariant ensembles from the classical orthogonal polynomials. The continuous case. Electronic Journal of Probability, 9:7:177–7:208, 2004. CODEN ???? ISSN 1083-

6489. URL http://ejp.ejpecp.org/article/view/191.

Ledoux:2005:DIE

[Led05] Michel Ledoux. Distributions of invariant ensembles from the classical orthogonal polynimials: the discrete case. Electronic Journal of Probability, 10:34:1116– 34:1146, 2005. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/282.

Legrand:2021:IDD

[Leg21] Alexandre Legrand. Influence of disorder on DNA denaturation: the disordered generalized Poland–Scheraga model. Electronic Journal of Probability, 26(??):1– 43, 2021. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

journals/electronic-journal-of-probability/volume-26/issue-none/Influence-
of-disorder-on-DNA-denaturation--thedisordered-generalized-Poland/
10.1214/20-EJP563.full.

Lehericy:2022:FPP

[Leh22] Thomas Lehéricy. First-passage percolation in random planar maps and Tutte’s bijection. Electronic Journal of Probability, 27(??):1–50, 2022. CODEN ????

ISSN 1083-6489. URL https://projecteuclid.org/journals/electronic-
journal-of-probability/volume-27/issue-none/First-passage-percolation-
in-random-planar-maps-and-Tuttes-bijection/10.1214/21-EJP662.full.

Lejay:2002:CSI

[Lej02] Antoine Lejay. On the convergence of stochastic integrals driven by processes converging on account of a homogenization property. Electronic Journal of Probability, 7:18:1–18:18, 2002. CODEN ???? ISSN 1083-6489. URL http://

ejp.ejpecp.org/article/view/117.

Lejay:2009:RDE

[Lej09] Antoine Lejay. On rough differential equations. Electronic Journal of Probability, 14:12:341–12:364, 2009. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/613.

Lejay:2022:CGR

[Lej22] Antoine Lejay. Constructing general rough differential equations through flow approximations. Electronic Journal of Probability, 27(??):1–24, 2022. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-27/issue-none/Constructing-
general-rough-differential-equations-through-flow-approximations/
10.1214/21-EJP717.full.

Lerouvillois:2020:HLD

[Ler20] Vincent Lerouvillois. Hydrodynamic limit of a (2+1)-dimensional crystal growth model in the anisotropic KPZ class. Electronic Journal of Probability, 25(??):1– 35, 2020. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

journals/electronic-journal-of-probability/volume-25/issue-none/Hydrodynamic-
limit-of-a-21-dimensional-crystal-growth-model-in/10.1214/20-EJP473.
full.

Levin:2013:CLT

[Lev13] Mordechay Levin. Central Limit Theorem for Z_{+}^{d}-actions by toral endomorphisms. Electronic Journal of Probability, 18:35:1–35:42, 2013. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/1904.

Levy:2017:FDI

[Lev17] Avi Levy. Finitely dependent insertion processes. Electronic Journal of Probability, 22(??):32:1–32:19, ???? 2017. CODEN ???? ISSN 1083-6489. URL

https://projecteuclid.org/euclid.ejp/1491962643.

Lochowski:2014:ILL

[ŁG14] Rafał Marcin Łochowski and Raouf Ghomrasni. Integral and local limit theorems for level crossings of diffusions and the Skorohod problem. Electronic Journal of Probability, 19:10:1–10:33, 2014. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/2644.

Li:2014:MDC

[Li14] Zhongyang Li. 1-2 model, dimers and clusters. Electronic Journal of Probability, 19:48:1–48:28, 2014. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.

org/article/view/2563.

Liitiainen:2011:AMN

[Lii11] Elia Liitiäinen. Asymptotic moments of near neighbor distances for the Gaussian distribution. Electronic Journal of Probability, 16:92:2545–92:2573, 2011.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
969.

Limic:2010:SCI

[Lim10] Vlada Limic. On the speed of coming down from infinity for Ξ-coalescent processes. Electronic Journal of Probability, 15:8:217–8:240, 2010. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/742.

Lin:2013:SDE

[Lin13] Yiqing Lin. Stochastic differential equations driven by G-Brownian motion with reflecting boundary conditions. Electronic Journal of Probability, 18:9:1–9:23, 2013. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/

view/2566.

Lin:2014:HMB

[Lin14] Shen Lin. The harmonic measure of balls in critical Galton–Watson trees with infinite variance offspring distribution. Electronic Journal of Probability, 19 (??):98:1–98:35, ???? 2014. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/3498.

Lin:2016:NAP

[Lin16] Jeff Lin. A negative answer to a problem of Aldous on determination of exchangeable sequences. Electronic Journal of Probability, 21(??):30:1–30:26, ???? 2016. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

euclid.ejp/1460652931.

Lin:2020:STE

[Lin20] Yier Lin. The stochastic telegraph equation limit of the stochastic higher spin six vertex model. Electronic Journal of Probability, 25(??):1–30, 2020. CODEN ????

ISSN 1083-6489. URL https://projecteuclid.org/journals/electronic-
journal-of-probability/volume-25/issue-none/The-stochastic-telegraph-
equation-limit-of-the-stochastic-higher-spin/10.1214/20-EJP552.full.

Lis:2022:SPH

[Lis22] Marcin Lis. Spins, percolation and height functions. Electronic Journal of Probability, 27(??):1–21, 2022. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-27/issue-none/Spins-percolation-and-height-
functions/10.1214/22-EJP761.full.

Liu:2020:PIQ

[Liu20] Yuan Liu. The Poincaré inequality and quadratic transportation-variance inequalities. Electronic Journal of Probability, 25(??):1:1–1:16, ???? 2020. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.ejp/
1578020644.

Lawler:2004:BEC

[LL04] Gregory Lawler and Vlada Limic. The Beurling estimate for a class of random walks. Electronic Journal of Probability, 9:27:846–27:861, 2004. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/228.

Lifshits:2011:RGS

[LL11] Mikhail Lifshits and Werner Linde. Random Gaussian sums on trees. Electronic Journal of Probability, 16:24:739–24:763, 2011. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/871.

Lopatto:2021:TBG

[LL21] Patrick Lopatto and Kyle Luh. Tail bounds for gaps between eigenvalues of sparse random matrices. Electronic Journal of Probability, 26(18):1–26, February 2021. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

journals/electronic-journal-of-probability/volume-26/issue-none/Tail-
bounds-for-gaps-between-eigenvalues-of-sparse-random-matrices/10.
1214/21-EJP669.full.

Li:2022:DBM

[LL22] Liping Li and Shuwen Lou. Distorted Brownian motions on space with varying dimension. Electronic Journal of Probability, 27(??):1–32, 2022. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-27/issue-none/Distorted-
Brownian-motions-on-space-with-varying-dimension/10.1214/22-EJP796.
full.

Landim:2018:MMC

[LLM18] Claudio Landim, Michail Loulakis, and Mustapha Mourragui. Metastable Markov chains: from the convergence of the trace to the convergence of the finite-dimensional distributions. Electronic Journal of Probability, 23(??): 95:1–95:34, ???? 2018. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1537322680.

Lifshits:2006:SDG

[LLS06] Mikhail Lifshits, Werner Linde, and Zhan Shi. Small deviations of Gaussian random fields in L_{q}-spaces. Electronic Journal of Probability, 11:46:1204–46:1233, 2006. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/

view/379.

Liao:1996:ASE

[LM96] Xiao Xin Liao and Xuerong Mao. Almost sure exponential stability of neutral differential difference equations with damped stochastic perturbations. Electronic Journal of Probability, 1:8:1–8:16, 1996. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/8.

Labarbe:2007:ABR

[LM07a] Jean-Maxime Labarbe and Jean-François Marckert. Asymptotics of Bernoulli random walks, bridges, excursions and meanders with a given number of peaks. Electronic Journal of Probability, 12:9:229–9:261, 2007. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/397.

Luczak:2007:ADC

[LM07b] Malwina Luczak and Colin McDiarmid. Asymptotic distributions and chaos for the supermarket model. Electronic Journal of Probability, 12:3:75–3:99, 2007.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
391.

Lember:2008:IRR

[LM08] Jyri Lember and Heinrich Matzinger. Information recovery from randomly mixed-up message text. Electronic Journal of Probability, 13:15:396–15:466, 2008. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/

view/491.

Lecue:2015:MRC

[LM15] Guillaume Lecué and Shahar Mendelson. Minimax rate of convergence and the performance of empirical risk minimization in phase recovery. Electronic Journal of Probability, 20(??):57:1–57:29, ???? 2015. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/3525.

Lin:2019:CSW

[LM19a] Kevin Lin and Carl Mueller. Can the stochastic wave equation with strong drift hit zero? Electronic Journal of Probability, 24(??):14:1–14:26, ???? 2019.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.
ejp/1550653273.

Lubbers:2019:SCB

[LM19b] Jan-Erik Lübbers and Matthias Meiners. The speed of critically biased random walk in a one-dimensional percolation model. Electronic Journal of Probability, 24(??):23:1–23:29, ???? 2019. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1553306439.

Leo:2023:BSN

[LM23] Gayral Léo and Sablik Mathieu. On the Besicovitch-stability of noisy random tilings. Electronic Journal of Probability, 28(??):1–38, 2023. CODEN ????

ISSN 1083-6489. URL https://projecteuclid.org/journals/electronic-
journal-of-probability/volume-28/issue-none/On-the-Besicovitch-stability-
of-noisy-random-tilings/10.1214/23-EJP917.full.

Leon:2012:ALS

[LMCV12] Jorge Leon, David Márquez-Carreras, and Josep Vives. Anticipating linear stochastic differential equations driven by a Lévy process. Electronic Journal of Probability, 17:89:1–89:26, 2012. CODEN ???? ISSN 1083-6489. URL http://

ejp.ejpecp.org/article/view/1910.

Loewe:2004:RMR

[LMM04] Matthias Loewe, Heinrich Matzinger, and Franz Merkl. Reconstructing a multicolor random scenery seen along a random walk path with bounded jumps. Electronic Journal of Probability, 9:15:436–15:507, 2004. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/206.

Lyons:2008:EUS

[LMS08] Russell Lyons, Benjamin Morris, and Oded Schramm. Ends in uniform spanning forests. Electronic Journal of Probability, 13:58:1702–58:1725, 2008. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/566.

Little:2022:NRE

[LMS22] Alex Little, Francesco Mezzadri, and Nick Simm. On the number of real eigenvalues of a product of truncated orthogonal random matrices. Electronic Journal of Probability, 27(??):1–32, 2022. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-27/issue-none/On-the-number-of-real-eigenvalues-
of-a-product-of/10.1214/21-EJP732.full.

Lamacz:2015:MBC

[LNO15] Agnes Lamacz, Stefan Neukamm, and Felix Otto. Moment bounds for the corrector in stochastic homogenization of a percolation model. Electronic Journal of Probability, 20(??):106:1–106:30, ???? 2015. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/3618.

Lodewijks:2020:PTP

[LO20] Bas Lodewijks and Marcel Ortgiese. A phase transition for preferential attachment models with additive fitness. Electronic Journal of Probability, 25(??):1– 54, 2020. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

journals/electronic-journal-of-probability/volume-25/issue-none/A-
phase-transition-for-preferential-attachment-models-with-additive-
fitness/10.1214/20-EJP550.full.

Luh:2021:ECN

[LO21] Kyle Luh and Sean O’Rourke. Eigenvectors and controllability of non-Hermitian random matrices and directed graphs. Electronic Journal of Probability, 26(??): 1–43, 2021. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.

org/journals/electronic-journal-of-probability/volume-26/issue-none/
Eigenvectors-and-controllability-of-non-Hermitian-random-matrices-
and-directed/10.1214/21-EJP588.full.

Lodewijks:2022:JPV

[Lod22] Bas Lodewijks. On joint properties of vertices with a given degree or label in the random recursive tree. Electronic Journal of Probability, 27(??):1–45, 2022.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-27/issue-none/On-joint-properties-
of-vertices-with-a-given-degree-or/10.1214/22-EJP877.full.

Lochowski:2021:LTT

[ŁOPS21] Rafał M. Łochowski, Jan Obłój, David J. Prömel, and Pietro Siorpaes. Local times and Tanaka–Meyer formulae for càdlàg paths. Electronic Journal of Probability, 26(18):1–29, February 2021. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-26/issue-none/Local-times-and-TanakaMeyer-formulae-
for-c%c3%a0dl%c3%a0g-paths/10.1214/21-EJP638.full.

Lototsky:2001:LSP

[Lot01] Sergey Lototsky. Linear stochastic parabolic equations, degenerating on the boundary of a domain. Electronic Journal of Probability, 6:24:1–24:14, 2001.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/

Lou:2022:DAB

[Lou22] Shuwen Lou. Discrete approximation to Brownian motion with varying dimension in unbounded domains. Electronic Journal of Probability, 27(??):1–33, 2022.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-27/issue-none/Discrete-approximation-
to-Brownian-motion-with-varying-dimension-in-unbounded/10.1214/22-
EJP829.full.

Lamberton:2008:PBA

[LP08] Damien Lamberton and Gilles Pagès. A penalized bandit algorithm. Electronic Journal of Probability, 13:13:341–13:373, 2008. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/489.

Luschgy:2009:EGP

[LP09] Harald Luschgy and Gilles Pagès. Expansions for Gaussian processes and Parseval frames. Electronic Journal of Probability, 14:42:1198–42:1221, 2009. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/649.

Lopker:2013:TRP

[LP13] Andreas Löpker and Zbigniew Palmowski. On time reversal of piecewise deterministic Markov processes. Electronic Journal of Probability, 18:13:1–13:29, 2013. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/

view/1958.

Luschgy:2014:CQF

[LP14] Harald Luschgy and Gilles Pagès. Constructive quadratic functional quantization and critical dimension. Electronic Journal of Probability, 19:50:1–50:19, 2014. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/

view/3010.

Louidor:2015:LDE

[LP15] Oren Louidor and Will Perkins. Large deviations for the empirical distribution in the branching random walk. Electronic Journal of Probability, 20(??):18:1– 18:19, ???? 2015. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.

org/article/view/2147.

Laruelle:2019:NRU

[LP19a] Sophie Laruelle and Gilles Pagès. Nonlinear randomized urn models: a stochastic approximation viewpoint. Electronic Journal of Probability, 24(??): 98:1–98:47, ???? 2019. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1568793792.

Lugosi:2019:FSU

[LP19b] Gábor Lugosi and Alan S. Pereira. Finding the seed of uniform attachment trees. Electronic Journal of Probability, 24(??):18:1–18:15, ???? 2019. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.ejp/
1551323285.

Lubinsky:2022:VRZ

[LP22] Doron S. Lubinsky and Igor E. Pritsker. Variance of real zeros of random orthogonal polynomials for varying and exponential weights. Electronic Journal of Probability, 27(??):1–32, 2022. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-27/issue-none/Variance-of-real-zeros-of-random-
orthogonal-polynomials-for-varying/10.1214/22-EJP802.full.

Levajkovic:2015:SEE

[LPSŽ15] Tijana Levajković, Stevan Pilipović, Dora Seleši, and Milica Žigić. Stochastic evolution equations with multiplicative noise. Electronic Journal of Probability, 20(??):19:1–19:23, ???? 2015. CODEN ???? ISSN 1083-6489. URL http://

ejp.ejpecp.org/article/view/3696.

Levajkovic:2018:SEE

[LPSŽ18] Tijana Levajković, Stevan Pilipović, Dora Seleši, and Milica Žigić. Stochastic evolution equations with Wick-polynomial nonlinearities. Electronic Journal of Probability, 23(??):116:1–116:25, ???? 2018. CODEN ???? ISSN 1083-6489.

URL https://projecteuclid.org/euclid.ejp/1543028704.

Lupu:2015:VTB

[LPT15] Titus Lupu, Jim Pitman, and Wenpin Tang. The Vervaat transform of Brownian bridges and Brownian motion. Electronic Journal of Probability, 20(??):51:1– 51:31, ???? 2015. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.

org/article/view/3744.

Lim:2011:EUM

[LQ11] Thomas Lim and Marie-Claire Quenez. Exponential utility maximization in an incomplete market with defaults. Electronic Journal of Probability, 16:53:1434– 53:1464, 2011. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/918.

Ledoux:2010:SDB

[LR10] Michel Ledoux and Brian Rider. Small deviations for beta ensembles. Electronic Journal of Probability, 15:41:1319–41:1343, 2010. CODEN ???? ISSN 1083-

6489. URL http://ejp.ejpecp.org/article/view/798.

Louhichi:2011:FCS

[LR11] Sana Louhichi and Emmanuel Rio. Functional convergence to stable Lévy motions for iterated random Lipschitz mappings. Electronic Journal of Probability, 16:89:2452–89:2480, 2011. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/965.

LeGoff:2018:VRN

[LR18] Line C. Le Goff and Olivier Raimond. Vertex reinforced non-backtracking random walks: an example of path formation. Electronic Journal of Probability, 23(??):39:1–39:38, ???? 2018. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1525852816.

Linker:2020:CPD

[LR20] Amitai Linker and Daniel Remenik. The contact process with dynamic edges on Z. Electronic Journal of Probability, 25(??):1–21, 2020. CODEN ????

ISSN 1083-6489. URL https://projecteuclid.org/journals/electronic-
journal-of-probability/volume-25/issue-none/The-contact-process-with-
dynamic-edges-on-mathbb-Z/10.1214/20-EJP480.full.

Lachieze-Rey:2022:DGE

[LR22] Raphaël Lachièze-Rey. Diophantine Gaussian excursions and random walks. Electronic Journal of Probability, 27(??):1–33, 2022. CODEN ????

ISSN 1083-6489. URL https://projecteuclid.org/journals/electronic-
journal-of-probability/volume-27/issue-none/Diophantine-Gaussian-
excursions-and-random-walks/10.1214/22-EJP854.full.

Leonenko:2011:FEH

[LRMT11] Nikolai Leonenko, Maria D. Ruiz-Medina, and Murad S. Taqqu. Fractional elliptic, hyperbolic and parabolic random fields. Electronic Journal of Probability, 16:40:1134–40:1172, 2011. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/891.

Lachieze-Rey:2013:FGF

[LRP13] Raphael Lachieze-Rey and Giovanni Peccati. Fine Gaussian fluctuations on the Poisson space, I: contractions, cumulants and geometric random graphs. Electronic Journal of Probability, 18:32:1–32:32, 2013. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/2104.

Lin:2020:SOB

[LRTY20] Yiqing Lin, Zhenjie Ren, Nizar Touzi, and Junjian Yang. Second order backward SDE with random terminal time. Electronic Journal of Probability, 25(??):1– 43, 2020. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

journals/electronic-journal-of-probability/volume-25/issue-none/Second-
order-backward-SDE-with-random-terminal-time/10.1214/20-EJP498.full.

Liptser:1999:MDT

[LS99] R. Liptser and V. Spokoiny. Moderate deviations type evaluation for integral functionals of diffusion processes. Electronic Journal of Probability, 4:17:1–17:25, 1999. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/

view/54.

Limic:2006:SC

[LS06] Vlada Limic and Anja Sturm. The spatial Λ-coalescent. Electronic Journal of Probability, 11:15:363–15:393, 2006. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/319.

Liu:2011:HFK

[LS11a] Gi-Ren Liu and Narn-Rueih Shieh. Homogenization of fractional kinetic equations with random initial data. Electronic Journal of Probability, 16:32:962– 32:980, 2011. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/896.

Lubetzky:2011:EEC

[LS11b] Eyal Lubetzky and Allan Sly. Explicit expanders with cutoff phenomena. Electronic Journal of Probability, 16:15:419–15:435, 2011. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/869.

Li:2014:LBD

[LS14] Xinyi Li and Alain-Sol Sznitman. A lower bound for disconnection by random interlacements. Electronic Journal of Probability, 19:17:1–17:26, 2014. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/3067.

Li:2019:OPF

[LS19] Xinyi Li and Daisuke Shiraishi. One-point function estimates for loop-erased random walk in three dimensions. Electronic Journal of Probability, 24(??): 111:1–111:46, ???? 2019. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1570586691.

Ledger:2021:MCN

[LS21a] Sean Ledger and Andreas Søjmark. At the mercy of the common noise: blow-ups in a conditional McKean–Vlasov problem. Electronic Journal of Probability, 26(??):1–39, 2021. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-26/issue-none/At-the-mercy-of-the-common-noise-
-blow-ups/10.1214/21-EJP597.full.

Lember:2021:EFS

[LS21b] Jüri Lember and Joonas Sova. Exponential forgetting of smoothing distributions for pairwise Markov models. Electronic Journal of Probability, 26(18):1–30, February 2021. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.

org/journals/electronic-journal-of-probability/volume-26/issue-none/
Exponential-forgetting-of-smoothing-distributions-for-pairwise-Markov-
models/10.1214/21-EJP628.full.

Lata:2022:NRC

[LŚ22a] Rafał Lata and Witold Świa̧tkowski. Norms of randomized circulant matrices. Electronic Journal of Probability, 27(??):1–23, 2022. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-27/issue-none/Norms-of-randomized-circulant-
matrices/10.1214/22-EJP799.full.

Li:2022:HCS

[LS22b] Xinyi Li and Daisuke Shiraishi. The Hölder continuity of the scaling limit of three-dimensional loop-erased random walk. Electronic Journal of Probability, 27(??):1–37, 2022. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-27/issue-none/The-H%c3%b6lder-continuity-of-
the-scaling-limit-of-three-dimensional/10.1214/22-EJP869.full.

Larsson:2020:EPM

[LSF20] Martin Larsson and Sara Svaluto-Ferro. Existence of probability measure valued jump-diffusions in generalized Wasserstein spaces. Electronic Journal of Probability, 25(??):1–25, 2020. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-25/issue-none/Existence-of-probability-measure-
valued-jump-diffusions-in-generalized-Wasserstein/10.1214/20-EJP562.
full.

Lupu:2019:ICS

[LST19] Titus Lupu, Christophe Sabot, and Pierre Tarrès. Inverting the coupling of the signed Gaussian free field with a loop-soup. Electronic Journal of Probability, 24(??):70:1–70:28, ???? 2019. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1561687603.

Lashari:2021:DS

[LST21a] Abid Ali Lashari, Ana Serafimović, and Pieter Trapman. The duration of a supercritical. Electronic Journal of Probability, 26(18):1–49, February 2021.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-26/issue-none/The-duration-
of-a-supercritical-SIR-epidemic-on-a-configuration/10.1214/21-EJP679.
full.

Lupu:2021:IRK

[LST21b] Titus Lupu, Christophe Sabot, and Pierre Tarrès. Inverting the Ray–Knight identity on the line. Electronic Journal of Probability, 26(18):1–25, February 2021. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

journals/electronic-journal-of-probability/volume-26/issue-none/Inverting-
the-Ray-Knight-identity-on-the-line/10.1214/21-EJP657.full.

Lawler:2001:OAE

[LSW01] Gregory Lawler, Oded Schramm, and Wendelin Werner. One-arm exponent for critical 2D percolation. Electronic Journal of Probability, 7:2:1–2:13, 2001.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
101.

Liu:2012:FER

[LSW12] Dangzheng Liu, Xin Sun, and Zhengdong Wang. Fluctuations of eigenvalues for random Toeplitz and related matrices. Electronic Journal of Probability, 17: 95:1–95:22, 2012. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.

org/article/view/2006.

Liu:2022:AAH

[LSW22] Gi-Ren Liu, Yuan-Chung Sheu, and Hau-Tieng Wu. Asymptotic analysis of higher-order scattering transform of Gaussian processes. Electronic Journal of Probability, 27(??):1–27, 2022. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-27/issue-none/Asymptotic-analysis-of-higher-
order-scattering-transform-of-Gaussian-processes/10.1214/22-EJP766.
full.

Lacoin:2015:MPM

[LT15a] Hubert Lacoin and Augusto Teixeira. A mathematical perspective on metastable wetting. Electronic Journal of Probability, 20(??):17:1–17:23, ???? 2015. CO-

DEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
3241.

Latala:2015:NSC

[LT15b] Rafał Latała and Tomasz Tkocz. A note on suprema of canonical processes based on random variables with regular moments. Electronic Journal of Probability, 20(??):36:1–36:17, ???? 2015. CODEN ???? ISSN 1083-6489. URL http://

ejp.ejpecp.org/article/view/3625.

Limic:2015:DLS

[LT15c] Vlada Limic and Anna Talarczyk. Diffusion limits at small times for Λcoalescents with a Kingman component. Electronic Journal of Probability, 20 (??):45:1–45:20, ???? 2015. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/3818.

Lauriere:2022:BPC

[LT22] Mathieu Laurière and Ludovic Tangpi. Backward propagation of chaos. Electronic Journal of Probability, 27(??):1–30, 2022. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-27/issue-none/Backward-propagation-of-chaos/
10.1214/22-EJP777.full.

Lucon:2011:QLF

[Luç11] Eric Luçon. Quenched limits and fluctuations of the empirical measure for plane rotators in random media. Electronic Journal of Probability, 16:26:792–26:829, 2011. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/

view/874.

Luo:2020:TGS

[Luo20] Peng Luo. A type of globally solvable BSDEs with triangularly quadratic generators. Electronic Journal of Probability, 25(??):1–23, 2020. CODEN

????  ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-25/issue-none/A-type-of-
globally-solvable-BSDEs-with-triangularly-quadratic-generators/10.
1214/20-EJP504.full.

Lupu:2021:ITD

[Lup21] Titus Lupu. Isomorphisms of β-Dyson’s Brownian motion with Brownian local time. Electronic Journal of Probability, 26(18):1–31, February 2021. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-26/issue-none/Isomorphisms-
of-%ce%b2-Dysons-Brownian-motion-with-Brownian-local-time/10.1214/
21-EJP697.full.

Loubaton:2011:ASL

[LV11] Philippe Loubaton and Pascal Vallet. Almost sure localization of the eigenvalues in a Gaussian information plus noise model. Application to the spiked models. Electronic Journal of Probability, 16:70:1934–70:1959, 2011. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/943.

Luo:2001:STP

[LW01a] S. Luo and John Walsh. A stochastic two-point boundary value problem. Electronic Journal of Probability, 7:12:1–12:32, 2001. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/111.

Lyne:2001:WSS

[LW01b] Owen Lyne and David Williams. Weak solutions for a simple hyperbolic system. Electronic Journal of Probability, 6:20:1–20:21, 2001. CODEN ???? ISSN 1083-

6489. URL http://ejp.ejpecp.org/article/view/93.

Li:2020:EEG

[LW20a] Pei-Sen Li and Jian Wang. Exponential ergodicity for general continuous-state nonlinear branching processes. Electronic Journal of Probability, 25(??):1–25, 2020. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

journals/electronic-journal-of-probability/volume-25/issue-none/Exponential-
ergodicity-for-general-continuous-state-nonlinear-branching-processes/
10.1214/20-EJP528.full.

Lun:2020:CSP

[LW20b] Chin Hang Lun and Jon Warren. Continuity and strict positivity of the multi-layer extension of the stochastic heat equation. Electronic Journal of Probability, 25(??):1–41, 2020. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-25/issue-none/Continuity-and-strict-positivity-
of-the-multi-layer-extension-of/10.1214/20-EJP511.full.

Liu:2021:SLC

[LW21] Mingchang Liu and Hao Wu. Scaling limits of crossing probabilities in metric graph GFF. Electronic Journal of Probability, 26(??):1–46, 2021. CODEN ????

ISSN 1083-6489. URL https://projecteuclid.org/journals/electronic-
journal-of-probability/volume-26/issue-none/Scaling-limits-of-crossing-
probabilities-in-metric-graph-GFF/10.1214/21-EJP598.full.

Lyne:2000:TWC

[Lyn00] Owen Lyne. Travelling waves for a certain first-order coupled PDE system. Electronic Journal of Probability, 5:14:1–14:40, 2000. CODEN ???? ISSN 1083-

6489. URL http://ejp.ejpecp.org/article/view/70.

Li:2010:ELM

[LZ10] Wenbo Li and Xinyi Zhang. Expected lengths of minimum spanning trees for non-identical edge distributions. Electronic Journal of Probability, 15:5:110– 5:141, 2010. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/735.

Liu:2012:CSP

[LZ12] Huili Liu and Xiaowen Zhou. The compact support property for the Λ-Fleming– Viot process with underlying Brownian motion. Electronic Journal of Probability, 17:73:1–73:20, 2012. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/1928.

Lamberton:2013:OSO

[LZ13] Damien Lamberton and Mihail Zervos. On the optimal stopping of a onedimensional diffusion. Electronic Journal of Probability, 18:34:1–34:49, 2013.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
2182.

Lalley:2015:CBB

[LZ15] Steven P. Lalley and Bowei Zheng. Critical branching Brownian motion with killing. Electronic Journal of Probability, 20(??):118:1–118:29, ???? 2015. CO-

DEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
4466.

Li:2021:ECC

[LZ21] Bo Li and Xiaowen Zhou. On the explosion of a class of continuous-state nonlinear branching processes. Electronic Journal of Probability, 26(18):1–25, February 2021. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

journals/electronic-journal-of-probability/volume-26/issue-none/On-
the-explosion-of-a-class-of-continuous-state-nonlinear/10.1214/21-
EJP715.full.

Lacker:2023:SSL

[LZ23] Daniel Lacker and Jiacheng Zhang. Stationary solutions and local equations for interacting diffusions on regular trees. Electronic Journal of Probability, 28(??): 1–37, 2023. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.

org/journals/electronic-journal-of-probability/volume-28/issue-none/
Stationary-solutions-and-local-equations-for-interacting-diffusions-
on-regular/10.1214/22-EJP889.full.

Maffucci:2020:RAL

[Maf20] Riccardo W. Maffucci. Restriction of 3D arithmetic Laplace eigenfunctions to a plane. Electronic Journal of Probability, 25(??):60:1–60:17, ???? 2020.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.
ejp/1588924817.

Maida:2007:LDL

[Mai07] Mylène Maida. Large deviations for the largest eigenvalue of rank one deformations of Gaussian ensembles. Electronic Journal of Probability, 12:41:1131– 41:1150, 2007. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/438.

Major:2007:MVB

[Maj07] Peter Major. On a multivariate version of Bernstein’s inequality. Electronic Journal of Probability, 12:34:966–34:988, 2007. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/430.

Mallein:2015:MDB

[Mal15a] Bastien Mallein. Maximal displacement in a branching random walk through interfaces. Electronic Journal of Probability, 20(??):68:1–68:40, ???? 2015. CO-

DEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
2828.

Maller:2015:SLZ

[Mal15b] Ross A. Maller. Strong laws at zero for trimmed Lévy processes. Electronic Journal of Probability, 20(??):88:1–88:24, ???? 2015. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/3839.

Marx:2018:NAC

[Mar18] Victor Marx. A new approach for the construction of a Wasserstein diffusion. Electronic Journal of Probability, 23(??):124:1–124:54, ???? 2018. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.ejp/
1545188691.

Martin:2020:SDM

[Mar20] James B. Martin. Stationary distributions of the multi-type ASEP. Electronic Journal of Probability, 25(??):43:1–43:41, ???? 2020. CODEN ???? ISSN

1083-6489. URL https://projecteuclid.org/euclid.ejp/1585879251.

Masiero:2007:RPT

[Mas07] Federica Masiero. Regularizing properties for transition semigroups and semilinear parabolic equations in Banach spaces. Electronic Journal of Probability, 12:13:387–13:419, 2007. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/401.

Masson:2009:GEP

[Mas09] Robert Masson. The growth exponent for planar loop-erased random walk. Electronic Journal of Probability, 14:36:1012–36:1073, 2009. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/651.

Mastrostefano:2022:ASU

[Mas22] Daniele Mastrostefano. An almost sure upper bound for random multiplicative functions on integers with a large prime factor. Electronic Journal of Probability, 27(??):1–21, 2022. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-27/issue-none/An-almost-sure-upper-bound-for-
random-multiplicative-functions-on/10.1214/22-EJP751.full.

Matic:2011:LDP

[Mat11] Ivan Matic. Large deviations for processes in random environments with jumps. Electronic Journal of Probability, 16:87:2406–87:2438, 2011. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/962.

Matsui:2023:SDI

[Mat23] Muneya Matsui. Subexponentialiy of densities of infinitely divisible distributions. Electronic Journal of Probability, 28(??):1–29, 2023. CODEN ????

ISSN 1083-6489. URL https://projecteuclid.org/journals/electronic-
journal-of-probability/volume-28/issue-none/Subexponentialiy-of-densities-
of-infinitely-divisible-distributions/10.1214/23-EJP928.full.

Maximilian:2006:EMD

[Max06] Duerre Maximilian. Existence of multi-dimensional infinite volume selforganized critical forest-fire models. Electronic Journal of Probability, 11: 21:513–21:539, 2006. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.

org/article/view/333.

Monter:2014:IDF

[MB14] Sergio Almada Monter and Amarjit Budhiraja. Infinite dimensional forward– backward stochastic differential equations and the KPZ equation. Electronic Journal of Probability, 19:40:1–40:21, 2014. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/2709.

McKenna:2021:LDE

[McK21] Benjamin McKenna. Large deviations for extreme eigenvalues of deformed Wigner random matrices. Electronic Journal of Probability, 26(??):1–37, 2021.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-26/issue-none/Large-deviations-
for-extreme-eigenvalues-of-deformed-Wigner-random-matrices/10.1214/
20-EJP571.full.

Marquez-Carreras:2003:LDP

[MCS03] David Marquez-Carreras and Monica Sarra. Large deviation principle for a stochastic heat equation with spatially correlated noise. Electronic Journal of Probability, 8:12:1–12:39, 2003. CODEN ???? ISSN 1083-6489. URL http://

ejp.ejpecp.org/article/view/146.

Meester:2009:USD

[MdBL09] Ronald Meester, Anne Fey den Boer, and Haiyan Liu. Uniqueness of the stationary distribution and stabilizability in Zhang’s sandpile model. Electronic Journal of Probability, 14:32:895–32:911, 2009. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/640.

Meliot:2019:ART

[Mél19] Pierre-Loïc Méliot. Asymptotic representation theory and the spectrum of a random geometric graph on a compact Lie group. Electronic Journal of Probability, 24(??):43:1–43:85, ???? 2019. CODEN ???? ISSN 1083-6489. URL

https://projecteuclid.org/euclid.ejp/1555466613.

Mendez:2007:ETS

[Men07] Pedro Mendez. Exit times of symmetric stable processes from unbounded convex domains. Electronic Journal of Probability, 12:4:100–4:121, 2007. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/393.

Menz:2014:AOR

[Men14a] Georg Menz. The approach of Otto–Reznikoff revisited. Electronic Journal of Probability, 19(??):107:1–107:27, ???? 2014. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/3418.

Menz:2014:BLT

[Men14b] Georg Menz. A Brascamp–Lieb type covariance estimate. Electronic Journal of Probability, 19:78:1–78:15, 2014. CODEN ???? ISSN 1083-6489. URL http://

ejp.ejpecp.org/article/view/2997.

Merlet:2008:CTS

[Mer08] Glenn Merlet. Cycle time of stochastic max-plus linear systems. Electronic Journal of Probability, 13:12:322–12:340, 2008. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/488.

Miclo:2017:DHO

[Mic17] Laurent Miclo. Duality and hypoellipticity: one-dimensional case studies. Electronic Journal of Probability, 22(??):91:1–91:32, ???? 2017. CODEN ???? ISSN

1083-6489. URL https://projecteuclid.org/euclid.ejp/1508464837.

Miclo:2020:CMV

[Mic20] Laurent Miclo. On the construction of measure-valued dual processes. Electronic Journal of Probability, 25(??):6:1–6:64, ???? 2020. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/euclid.ejp/1580202285.

Miermont:2001:OAC

[Mie01] Grégory Miermont. Ordered additive coalescent and fragmentations associated to Lévy processes with no positive jumps. Electronic Journal of Probability, 6: 14:1–14:33, 2001. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.

org/article/view/87.

Mimica:2012:HIS

[MK12] Ante Mimica and Panki Kim. Harnack inequalities for subordinate Brownian motions. Electronic Journal of Probability, 17:37:1–37:23, 2012. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/1930.

Menozzi:2010:SNA

[ML10] Stéphane Menozzi and Vincent Lemaire. On some non asymptotic bounds for the Euler scheme. Electronic Journal of Probability, 15:53:1645–53:1681, 2010.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
814.

Mountford:2015:LER

[MM15] Thomas Mountford and Jean-Christophe Mourrat. Lyapunov exponents of random walks in small random potential: the upper bound. Electronic Journal of Probability, 20(??):49:1–49:18, ???? 2015. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/3489.

Mailler:2017:MVP

[MM17] Cécile Mailler and Jean-François Marckert. Measure-valued Pólya urn processes. Electronic Journal of Probability, 22(??):26:1–26:33, ???? 2017. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.ejp/
1490061796.

Maller:2018:MNS

[MM18] Ross A. Maller and David M. Mason. Matrix normalised stochastic compactness for a Lévy process at zero. Electronic Journal of Probability, 23(??):69:1–69:37, ???? 2018. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.

org/euclid.ejp/1532570597.

Maller:2020:CCP

[MM20] Ross A. Maller and David M. Mason. Compactness and continuity properties for a Lévy process at a two-sided exit time. Electronic Journal of Probability, 25(??):51:1–51:26, ???? 2020. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1588125886.

Mijatovic:2021:LPS

[MM21] Aleksandar Mijatović and Veno Mramor. Lévy processes on smooth manifolds with a connection. Electronic Journal of Probability, 26(18):1–39, February 2021. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

journals/electronic-journal-of-probability/volume-26/issue-none/L%
c3%a9vy-processes-on-smooth-manifolds-with-a-connection/10.1214/21-
EJP702.full.

Mueller:2008:RDS

[MN08] Carl Mueller and David Nualart. Regularity of the density for the stochastic heat equation. Electronic Journal of Probability, 13:74:2248–74:2258, 2008. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/589.

Menard:2014:PUI

[MN14] Laurent Ménard and Pierre Nolin. Percolation on uniform infinite planar maps. Electronic Journal of Probability, 19(??):79:1–79:27, ???? 2014. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/2675.

Mossel:2018:NSC

[MN18a] Elchanan Mossel and Joe Neeman. Noise stability and correlation with half spaces. Electronic Journal of Probability, 23(??):16:1–16:17, ???? 2018. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.ejp/
1519354945.

Muller:2018:RAC

[MN18b] Noela Müller and Ralph Neininger. Refined asymptotics for the composition of cyclic urns. Electronic Journal of Probability, 23(??):117:1–117:20, ???? 2018.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.
ejp/1543028707.

Morrison:2021:STB

[MN21] Natasha Morrison and Jonathan A. Noel. A sharp threshold for bootstrap percolation in a random hypergraph. Electronic Journal of Probability, 26(18):1–85, February 2021. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.

org/journals/electronic-journal-of-probability/volume-26/issue-none/
A-sharp-threshold-for-bootstrap-percolation-in-a-random-hypergraph/
10.1214/21-EJP650.full.

Mossel:2016:CTP

[MNS16] Elchanan Mossel, Joe Neeman, and Allan Sly. Consistency thresholds for the planted bisection model. Electronic Journal of Probability, 21(??):21:1–21:24, ???? 2016. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.

org/euclid.ejp/1457706457.

Meerschaert:2011:FPP

[MNV11] Mark Meerschaert, Erkan Nane, and P. Vellaisamy. The fractional Poisson process and the inverse stable subordinator. Electronic Journal of Probability, 16:59:1600–59:1620, 2011. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/920.

Monmarche:2015:ECC

[Mon15] Pierre Monmarché. On H^{1} and entropic convergence for contractive PDMP. Electronic Journal of Probability, 20(??):128:1–128:30, ???? 2015. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/3581.

Mountford:2001:NLB

[Mou01] Thomas Mountford. A note on limiting behaviour of disastrous environment exponents. Electronic Journal of Probability, 6:1:1–1:10, 2001. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/74.

Mourrat:2012:QCL

[Mou12] Jean-Christophe Mourrat. A quantitative central limit theorem for the random walk among random conductances. Electronic Journal of Probability, 17:97:1– 97:17, 2012. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/2414.

Mikulevicius:2005:CDP

[MP05] Remigijus Mikulevicius and Henrikas Pragarauskas. On Cauchy–Dirichlet problem in half-space for linear integro-differential equations in weighted Hölder spaces. Electronic Journal of Probability, 10:42:1398–42:1416, 2005. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/292.

Matzinger:2007:DLP

[MP07] Heinrich Matzinger and Serguei Popov. Detecting a local perturbation in a continuous scenery. Electronic Journal of Probability, 12:22:637–22:660, 2007.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
409.

Mourrat:2020:EPF

[MP20] Jean-Christophe Mourrat and Dmitry Panchenko. Extending the Parisi formula along a Hamilton–Jacobi equation. Electronic Journal of Probability, 25 (??):23:1–23:17, ???? 2020. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1581735875.

Michelen:2019:IPG

[MPR19] Marcus Michelen, Robin Pemantle, and Josh Rosenberg. Invasion percolation on Galton–Watson trees. Electronic Journal of Probability, 24(??):31:1–31:35, ???? 2019. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.

org/euclid.ejp/1554256913.

Menezes:2022:VSS

[MPX22] Otávio Menezes, Jonathon Peterson, and Yongjia Xie. Variable speed symmetric random walk driven by the simple symmetric exclusion process. Electronic Journal of Probability, 27(??):1–14, 2022. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-27/issue-none/Variable-speed-symmetric-random-
walk-driven-by-the-simple-symmetric/10.1214/21-EJP735.full.

Mikulevicius:2001:NKT

[MR01] R. Mikulevicius and B. Rozovskii. A note on Krylov’s L_{p}-theory for systems of SPDEs. Electronic Journal of Probability, 6:12:1–12:35, 2001. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/85.

Merkl:2008:BRE

[MR08] Franz Merkl and Silke Rolles. Bounding a random environment bounding a random environment for two-dimensional edge-reinforced random walk. Electronic Journal of Probability, 13:19:530–19:565, 2008. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/495.

Merkl:2009:SBC

[MR09] Franz Merkl and Silke Rolles. Spontaneous breaking of continuous rotational symmetry in two dimensions. Electronic Journal of Probability, 14:57:1705– 57:1726, 2009. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/671.

Marinelli:2010:WPA

[MR10] Carlo Marinelli and Michael Roeckner. Well posedness and asymptotic behavior for stochastic reaction–diffusion equations with multiplicative Poisson noise. Electronic Journal of Probability, 15:49:1529–49:1555, 2010. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/818.

Marcus:2012:CLT

[MR12] Michael Marcus and Jay Rosen. Central limit theorems for the L^{2} norm of increments of local times of Lévy processes. Electronic Journal of Probability, 17:7:1–7:111, 2012. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.

org/article/view/1740.

Merkl:2013:PAV

[MR13] Franz Merkl and Silke Rolles. Perturbation analysis of the van den Berg Kesten inequality for determinantal probability measures. Electronic Journal of Probability, 18:92:1–92:20, 2013. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/2339.

Merlevede:2015:SAA

[MR15] Florence Merlevède and Emmanuel Rio. Strong approximation for additive functionals of geometrically ergodic Markov chains. Electronic Journal of Probability, 20(??):14:1–14:27, ???? 2015. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/3746.

Martin:2017:RRM

[MR17a] James B. Martin and Balázs Ráth. Rigid representations of the multiplicative coalescent with linear deletion. Electronic Journal of Probability, 22(??): 83:1–83:47, ???? 2017. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1507946758.

Morris:2017:MTF

[MR17b] Ben Morris and Anastasia Raymer. Mixing time of the fifteen puzzle. Electronic Journal of Probability, 22(??):9:1–9:29, ???? 2017. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/euclid.ejp/1485831705.

Marcus:2018:SPP

[MR18] Michael B. Marcus and Jay Rosen. Sample path properties of permanental processes. Electronic Journal of Probability, 23(??):58:1–58:47, ???? 2018. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.ejp/
1528704076.

Michelen:2020:FMN

[MR20] Marcus Michelen and Josh Rosenberg. The frog model on non-amenable trees. Electronic Journal of Probability, 25(??):49:1–49:16, ???? 2020. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.ejp/
1588039468.

Matzavinos:2016:RWS

[MRS16] Anastasios Matzavinos, Alexander Roitershtein, and Youngsoo Seol. Random walks in a sparse random environment. Electronic Journal of Probability, 21 (??):72:1–72:20, ???? 2016. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1480993226.

Montgomery-Smith:1998:CRM

[MS98] Stephen Montgomery-Smith. Concrete representation of martingales. Electronic Journal of Probability, 3:15:1–15:15, 1998. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/37.

Morters:1999:SSL

[MS99] Peter Mörters and Narn-Rueih Shieh. Small scale limit theorems for the intersection local times of Brownian motion. Electronic Journal of Probability, 4: 9:1–9:23, 1999. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/46.

Matsui:2009:EFO

[MS09] Muneya Matsui and Narn-Rueih Shieh. On the exponentials of fractional Ornstein–Uhlenbeck processes. Electronic Journal of Probability, 14:23:594– 23:611, 2009. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/628.

Masse:2012:RNS

[MS12] Bruno Massé and Dominique Schneider. Random number sequences and the first digit phenomenon. Electronic Journal of Probability, 17:86:1–86:17, 2012.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
1900.

Matsumoto:2013:CFZ

[MS13] Sho Matsumoto and Tomoyuki Shirai. Correlation functions for zeros of a Gaussian power series and Pfaffians. Electronic Journal of Probability, 18:49:1–49:18, 2013. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/

view/2545.

Matic:2015:EDW

[MS15] Ivan Matic and David Sivakoff. Excited deterministic walk in a random environment. Electronic Journal of Probability, 20(??):44:1–44:19, ???? 2015. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/3874.

Milos:2019:EPT

[MŞ19] Piotr Miłoś and Batı Şengül. Existence of a phase transition of the interchange process on the Hamming graph. Electronic Journal of Probability, 24 (??):64:1–64:21, ???? 2019. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1561169148.

Marguet:2021:LTB

[MS21] Aline Marguet and Charline Smadi. Long time behaviour of continuousstate nonlinear branching processes with catastrophes. Electronic Journal of Probability, 26(18):1–32, February 2021. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-26/issue-none/Long-time-behaviour-of-continuous-
state-nonlinear-branching-processes-with/10.1214/21-EJP664.full.

Murugan:2015:ATB

[MSC15] Mathav Kishore Murugan and Laurent Saloff-Coste. Anomalous threshold behavior of long range random walks. Electronic Journal of Probability, 20(??): 74:1–74:21, ???? 2015. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/3989.

Mach:2020:RTP

[MSS20] Tibor Mach, Anja Sturm, and Jan M. Swart. Recursive tree processes and the mean-field limit of stochastic flows. Electronic Journal of Probability, 25 (??):61:1–61:63, ???? 2020. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1589335470.

Marcovici:2019:ESC

[MST19] Irène Marcovici, Mathieu Sablik, and Siamak Taati. Ergodicity of some classes of cellular automata subject to noise. Electronic Journal of Probability, 24 (??):41:1–41:44, ???? 2019. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1555034440.

Mueller:1997:FWR

[MT97] Carl Mueller and Roger Tribe. Finite width for a random stationary interface. Electronic Journal of Probability, 2:7:1–7:27, 1997. CODEN ???? ISSN 1083-

6489. URL http://ejp.ejpecp.org/article/view/21.

Matsumoto:2002:WFS

[MT02a] Hiroyuki Matsumoto and Setsuo Taniguchi. Wiener functionals of second order and their Lévy measures. Electronic Journal of Probability, 7:14:1–14:30, 2002.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
113.

Mueller:2002:HPR

[MT02b] Carl Mueller and Roger Tribe. Hitting properties of a random string. Electronic Journal of Probability, 7:10:1–10:29, 2002. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/109.

Mueller:2004:SPA

[MT04] Carl Mueller and Roger Tribe. A singular parabolic Anderson model. Electronic Journal of Probability, 9:5:98–5:144, 2004. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/189.

Martinez:2012:ODP

[MT12a] Miguel Martinez and Denis Talay. One-dimensional parabolic diffraction equations: pointwise estimates and discretization of related stochastic differential equations with weighted local times. Electronic Journal of Probability, 17:27:1– 27:30, 2012. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/1905.

Meleard:2012:NHS

[MT12b] Sylvie Méléard and Viet Chi Tran. Nonlinear historical superprocess approximations for population models with past dependence. Electronic Journal of Probability, 17:47:1–47:32, 2012. CODEN ???? ISSN 1083-6489. URL http://

ejp.ejpecp.org/article/view/2093.

Melbourne:2022:APM

[MT22] Ian Melbourne and Dalia Terhesiu. Analytic proof of multivariate stable local large deviations and application to deterministic dynamical systems. Electronic Journal of Probability, 27(??):1–17, 2022. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-27/issue-none/Analytic-proof-of-multivariate-
stable-local-large-deviations-and-application/10.1214/22-EJP750.full.

Mucha:2021:STO

[Muc21] Jacek Mucha. Spectral theory for one-dimensional (non-symmetric) stable processes killed upon hitting the origin. Electronic Journal of Probability, 26(??):1– 33, 2021. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

journals/electronic-journal-of-probability/volume-26/issue-none/Spectral-
theory-for-one-dimensional-non-symmetric-stable-processes-killed/
10.1214/21-EJP594.full.

Mukherjee:2016:FPC

[Muk16] Sumit Mukherjee. Fixed points and cycle structure of random permutations. Electronic Journal of Probability, 21(??):40:1–40:18, ???? 2016. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.ejp/
1465991838.

Muller:2008:CTM

[Mül08] Sebastian Müller. A criterion for transience of multidimensional branching random walk in random environment. Electronic Journal of Probability, 13:41:1189– 41:1202, 2008. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/517.

Muller:2017:PLD

[Mül17] Patrick E. Müller. Path large deviations for interacting diffusions with local mean-field interactions in random environment. Electronic Journal of Probability, 22(??):76:1–76:56, ???? 2017. CODEN ???? ISSN 1083-6489. URL

https://projecteuclid.org/euclid.ejp/1505527232.

Munsonius:2011:AIP

[Mun11] Goetz Olaf Munsonius. On the asymptotic internal path length and the asymptotic Wiener index of random split trees. Electronic Journal of Probability, 16: 35:1020–35:1047, 2011. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/889.

Mikulevicius:2000:SEE

[MV00] R. Mikulevicius and G. Valiukevicius. On stochastic Euler equation in R^{d}. Electronic Journal of Probability, 5:6:1–6:20, 2000. CODEN ???? ISSN 1083-

6489. URL http://ejp.ejpecp.org/article/view/62.

Menshikov:2008:URR

[MV08] Mikhail Menshikov and Stanislav Volkov. Urn-related random walk with drift ρx^{α}/t^{β}. Electronic Journal of Probability, 13:31:944–31:960, 2008. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/508.

Mountford:2015:RWG

[MV15] Thomas S. Mountford and Maria Eulalia Vares. Random walks generated by equilibrium contact processes. Electronic Journal of Probability, 20(??):3:1–3:17, ???? 2015. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/3439.

Mourrat:2016:PTC

[MV16] Jean-Christophe Mourrat and Daniel Valesin. Phase transition of the contact process on random regular graphs. Electronic Journal of Probability, 21 (??):31:1–31:17, ???? 2016. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1460652932.

Mijatovic:2020:SOZ

[MV20] Aleksandar Mijatović and Vladislav Vysotsky. Stability of overshoots of zero mean random walks. Electronic Journal of Probability, 25(??):63:1–63:22, ???? 2020. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

euclid.ejp/1591668284.

Mijatovic:2014:MCA

[MVJ14] Aleksandar Mijatovic, Matija Vidmar, and Saul Jacka. Markov chain approximations for transition densities of Lévy processes. Electronic Journal of Probability, 19:7:1–7:37, 2014. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/2208.

Mountford:2020:CSA

[MVX20] Thomas Mountford, Maria Eulália Vares, and Hao Xue. Critical scaling for an anisotropic percolation system on. Electronic Journal of Probability, 25(??):1– 44, 2020. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

journals/electronic-journal-of-probability/volume-25/issue-none/Critical-
scaling-for-an-anisotropic-percolation-system-on-Z-2/10.1214/20-EJP533.
full.

Mountford:2013:MDC

[MVY13] Thomas Mountford, Daniel Valesin, and Qiang Yao. Metastable densities for the contact process on power law random graphs. Electronic Journal of Probability, 18:103:1–103:36, 2013. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/2512.

Mountford:2005:TCN

[MW05] Thomas Mountford and Li-Chau Wu. The time for a critical nearest particle system to reach equilibrium starting with a large gap. Electronic Journal of Probability, 10:13:436–13:498, 2005. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/242.

Miermont:2008:RPR

[MW08] Grégory Miermont and Mathilde Weill. Radius and profile of random planar maps with faces of arbitrary degrees. Electronic Journal of Probability, 13: 4:79–4:106, 2008. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.

org/article/view/478.

Madras:2010:TAP

[MW10] Neal Madras and C. Wu. Trees, animals, and percolation on hyperbolic lattices. Electronic Journal of Probability, 15:66:2019–66:2040, 2010. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/837.

McRedmond:2018:CHP

[MW18] James McRedmond and Andrew R. Wade. The convex hull of a planar random walk: perimeter, diameter, and shape. Electronic Journal of Probability, 23 (??):131:1–131:24, ???? 2018. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1545447916.

Minsker:2019:MIM

[MW19] Stanislav Minsker and Xiaohan Wei. Moment inequalities for matrix-valued U-statistics of order 2. Electronic Journal of Probability, 24(??):133:1–133:32, ???? 2019. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.

org/euclid.ejp/1573614085.

Moinat:2020:LBS

[MW20a] Augustin Moinat and Hendrik Weber. Local bounds for stochastic reaction diffusion equations. Electronic Journal of Probability, 25(??):17:1–17:26, ????

2020. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

euclid.ejp/1580871683.

Muller:2020:TFG

[MW20b] Sebastian Müller and Gundelinde Maria Wiegel. On transience of frogs on Galton–Watson trees. Electronic Journal of Probability, 25(??):1–30, 2020. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-25/issue-none/On-transience-
of-frogs-on-GaltonWatson-trees/10.1214/20-EJP558.full.

Mayer-Wolf:2002:ACC

[MWZZ02] Eddy Mayer-Wolf, Ofer Zeitouni, and Martin Zerner. Asymptotics of certain coagulation–fragmentation processes and invariant Poisson–Dirichlet measures. Electronic Journal of Probability, 7:8:1–8:25, 2002. CODEN ???? ISSN 1083-

6489. URL http://ejp.ejpecp.org/article/view/107.

Mytnik:2007:LES

[MX07] Leonid Mytnik and Jie Xiong. Local extinction for superprocesses in random environments. Electronic Journal of Probability, 12:50:1349–50:1378, 2007. CO-

DEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
457.

Matoussi:2008:SSS

[MX08] Anis Matoussi and Mingyu Xu. Sobolev solution for semilinear PDE with obstacle under monotonicity condition. Electronic Journal of Probability, 13:35:1035– 35:1067, 2008. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/522.

Maejima:2003:LMS

[MY03] Makoto Maejima and Kenji Yamamoto. Long-memory stable Ornstein– Uhlenbeck processes. Electronic Journal of Probability, 8:19:1–19:18, 2003. CO-

DEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
168.

Nagahata:2012:LBE

[Nag12] Yukio Nagahata. Lower bound estimate of the spectral gap for simple exclusion process with degenerate rates. Electronic Journal of Probability, 17:92:1–92:19, 2012. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/

view/1916.

Najim:2001:CTT

[Naj01] Jamal Najim. A Cramér type theorem for weighted random variables. Electronic Journal of Probability, 7:4:1–4:32, 2001. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/103.

Najnudel:2020:CVR

[Naj20] Joseph Najnudel. On consecutive values of random completely multiplicative functions. Electronic Journal of Probability, 25(??):59:1–59:28, ???? 2020. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.ejp/
1588644042.

Nakajima:2022:MET

[Nak22] Shuta Nakajima. Maximal edge-traversal time in first-passage percolation. Electronic Journal of Probability, 27(??):1–32, 2022. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-27/issue-none/Maximal-edge-traversal-time-in-
First-passage-percolation/10.1214/22-EJP746.full.

Nane:2006:LIL

[Nan06] Erkan Nane. Laws of the iterated logarithm for α-time Brownian motion. Electronic Journal of Probability, 11:18:434–18:459, 2006. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/327.

Nandan:2023:SPS

[Nan23] Shubhamoy Nandan. Spatial populations with seed-banks in random environment: III. Convergence towards mono-type equilibrium. Electronic Journal of Probability, 28(??):1–36, 2023. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-28/issue-none/Spatial-populations-with-seed-
banks-in-random-environment--III/10.1214/23-EJP922.full.

Nassif:2022:ZRS

[Nas22] Michel Nassif. Zooming in at the root of the stable tree. Electronic Journal of Probability, 27(??):1–38, 2022. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-27/issue-none/Zooming-in-at-the-root-of-the-
stable-tree/10.1214/22-EJP764.full.

Nejjar:2021:DPT

[Nej21] Peter Nejjar. Dynamical phase transition of ASEP in the KPZ regime. Electronic Journal of Probability, 26(18):1–20, February 2021. CODEN

????  ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-26/issue-none/Dynamical-
phase-transition-of-ASEP-in-the-KPZ-regime/10.1214/21-EJP642.full.

Nemish:2017:LLP

[Nem17] Yuriy Nemish. Local law for the product of independent non-Hermitian random matrices with independent entries. Electronic Journal of Probability, 22 (??):22:1–22:35, ???? 2017. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1487991681.

Neunhauserer:2007:RWI

[Neu07] Jörg Neunhäuserer. Random walks on infinite self-similar graphs. Electronic Journal of Probability, 12:46:1258–46:1275, 2007. CODEN ???? ISSN 1083-

6489. URL http://ejp.ejpecp.org/article/view/448.

Nguyen:2012:LSV

[Ngu12] Hoi Nguyen. On the least singular value of random symmetric matrices. Electronic Journal of Probability, 17:53:1–53:19, 2012. CODEN ???? ISSN 1083-

6489. URL http://ejp.ejpecp.org/article/view/2165.

Nica:2021:IDL

[Nic21] Mihai Nica. Intermediate disorder limits for multi-layer semi-discrete directed polymers. Electronic Journal of Probability, 26(18):1–50, February 2021. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-26/issue-none/Intermediate-
disorder-limits-for-multi-layer-semi-discrete-directed-polymers/10.
1214/21-EJP614.full.

Nilssen:2020:RLP

[Nil20] Torstein Nilssen. Rough linear PDE’s with discontinuous coefficients — existence of solutions via regularization by fractional Brownian motion. Electronic Journal of Probability, 25(??):34:1–34:33, ???? 2020. CODEN ???? ISSN

1083-6489. URL https://projecteuclid.org/euclid.ejp/1584669820.

Nishioka:2001:BCO

[Nis01] Kunio Nishioka. Boundary conditions for one-dimensional biharmonic pseudo process. Electronic Journal of Probability, 6:13:1–13:27, 2001. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/86.

Nitzschner:2018:DLS

[Nit18] Maximilian Nitzschner. Disconnection by level sets of the discrete Gaussian free field and entropic repulsion. Electronic Journal of Probability, 23(??): 105:1–105:21, ???? 2018. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1540260051.

Nualart:2004:LSM

[NM04] Eulalia Nualart and Thomas Mountford. Level sets of multiparameter Brownian motions. Electronic Journal of Probability, 9:20:594–20:614, 2004. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/169.

Neufeld:2013:SUV

[NN13] Ariel Neufeld and Marcel Nutz. Superreplication under volatility uncertainty for measurable claims. Electronic Journal of Probability, 18:48:1–48:14, 2013.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
2358.

Nestoridi:2021:FSR

[NN21] Evita Nestoridi and Oanh Nguyen. The full spectrum of random walks on complete finite d-ary trees. Electronic Journal of Probability, 26(18):1–17, February 2021. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

journals/electronic-journal-of-probability/volume-26/issue-none/The-
full-spectrum-of-random-walks-on-complete-finite-d/10.1214/21-EJP608.
full.

Nakajima:2023:FTD

[NN23] Shuta Nakajima and Makoto Nakashima. Fluctuations of two-dimensional stochastic heat equation and KPZ equation in subcritical regime for general initial conditions. Electronic Journal of Probability, 28(??):1–38, 2023. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-28/issue-none/Fluctuations-
of-two-dimensional-stochastic-heat-equation-and-KPZ-equation/10.1214/
22-EJP885.full.

Nourdin:2013:ACC

[NNP13] Ivan Nourdin, David Nualart, and Guillaume Poly. Absolute continuity and convergence of densities for random vectors on Wiener chaos. Electronic Journal of Probability, 18:22:1–22:19, 2013. CODEN ???? ISSN 1083-6489. URL http:

//ejp.ejpecp.org/article/view/2181.

Neuenkirch:2008:DED

[NNT08] Andreas Neuenkirch, Ivan Nourdin, and Samy Tindel. Delay equations driven by rough paths. Electronic Journal of Probability, 13:67:2031–67:2068, 2008.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
575.

Nolin:2008:NCP

[Nol08] Pierre Nolin. Near-critical percolation in two dimensions. Electronic Journal of Probability, 13:55:1562–55:1623, 2008. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/565.

Nordenstam:2010:SAD

[Nor10] Eric Nordenstam. On the shuffling algorithm for domino tilings. Electronic Journal of Probability, 15:3:75–3:95, 2010. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/730.

Nourdin:2008:WPV

[NP08] Ivan Nourdin and Giovanni Peccati. Weighted power variations of iterated Brownian motion. Electronic Journal of Probability, 13:43:1229–43:1256, 2008. CO-

DEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
534.

Nguyen:2017:ESN

[NR17] Gia Bao Nguyen and Daniel Remenik. Extreme statistics of non-intersecting Brownian paths. Electronic Journal of Probability, 22(??):102:1–102:40, ???? 2017. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

euclid.ejp/1511773232.

Newman:2005:CCN

[NRS05] Charles Newman, Krishnamurthi Ravishankar, and Rongfeng Sun. Convergence of coalescing nonsimple random walks to the Brownian Web. Electronic Journal of Probability, 10:2:21–2:60, 2005. CODEN ???? ISSN 1083-6489. URL http:

//ejp.ejpecp.org/article/view/235.

Nourdin:2010:WSI

[NRS10] Ivan Nourdin, Anthony Réveillac, and Jason Swanson. The weak Stratonovich integral with respect to fractional Brownian motion with Hurst parameter 1/6. Electronic Journal of Probability, 15:70:2117–70:2162, 2010. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/843.

Newman:2017:PVM

[NRS17] C. M. Newman, K. Ravishankar, and E. Schertzer. Perturbations of Voter model in one-dimension. Electronic Journal of Probability, 22(??):34:1–34:42, ???? 2017. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.

org/euclid.ejp/1492502428.

Nutz:2012:QSA

[Nut12] Marcel Nutz. A quasi-sure approach to the control of non-Markovian stochastic differential equations. Electronic Journal of Probability, 17:23:1–23:23, 2012.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
1892.

Nourdin:2009:DFC

[NV09] Ivan Nourdin and Frederi Viens. Density formula and concentration inequalities with Malliavin calculus. Electronic Journal of Probability, 14:78:2287–78:2309, 2009. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/

view/707.

Normand:2013:RWV

[NV13] Raoul Normand and Bálint Virág. Random walks veering left. Electronic Journal of Probability, 18:89:1–89:25, 2013. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/2523.

Nualart:2022:QCL

[NXZ22] David Nualart, Panqiu Xia, and Guangqu Zheng. Quantitative central limit theorems for the parabolic Anderson model driven by colored noises. Electronic Journal of Probability, 27(??):1–43, 2022. CODEN ????

ISSN 1083-6489. URL https://projecteuclid.org/journals/electronic-
journal-of-probability/volume-27/issue-none/Quantitative-central-
limit-theorems-for-the-parabolic-Anderson-model-driven/10.1214/22-
EJP847.full.

Nagahata:2009:CLT

[NY09] Yukio Nagahata and Nobuo Yoshida. Central Limit Theorem for a class of linear systems. Electronic Journal of Probability, 14:34:960–34:977, 2009. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/644.

Nagahata:2010:LCL

[NY10] Yukio Nagahata and Nobuo Yoshida. Localization for a class of linear systems. Electronic Journal of Probability, 15:20:636–20:653, 2010. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/757.

Nualart:2019:AES

[NY19] David Nualart and Nakahiro Yoshida. Asymptotic expansion of Skorohod integrals. Electronic Journal of Probability, 24(??):119:1–119:64, ???? 2019. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.ejp/
1572508843.

Nourdin:2014:ITF

[NZ14] Ivan Nourdin and Raghid Zeineddine. An Itô type formula for the fractional Brownian motion in Brownian time. Electronic Journal of Probability, 19(??): 99:1–99:15, ???? 2014. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/3184.

Nualart:2020:AGF

[NZ20] David Nualart and Guangqu Zheng. Averaging Gaussian functionals. Electronic Journal of Probability, 25(??):48:1–48:54, ???? 2020. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/euclid.ejp/1588039467.

Olvera-Cravioto:2019:CPD

[OC19] Mariana Olvera-Cravioto. Convergence of the population dynamics algorithm in the Wasserstein metric. Electronic Journal of Probability, 24(??):61:1–61:27, ???? 2019. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.

org/euclid.ejp/1561082670.

Ocafrain:2021:CQS

[Oça21] William Oçafrain. Convergence to quasi-stationarity through Poincaré inequalities and Bakry–Émery criteria. Electronic Journal of Probability, 26(18):1–30, February 2021. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.

org/journals/electronic-journal-of-probability/volume-26/issue-none/
Convergence-to-quasi-stationarity-through-Poincar%c3%a9-inequalities-
and-Bakry-%c3%89mery/10.1214/21-EJP644.full.

Odasso:2006:SSS

[Oda06] Cyril Odasso. Spatial smoothness of the stationary solutions of the 3D Navier– Stokes equations. Electronic Journal of Probability, 11:27:686–27:699, 2006.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
336.

Oliveira:2012:MHT

[Oli12] Roberto Oliveira. Mixing and hitting times for finite Markov chains. Electronic Journal of Probability, 17:70:1–70:12, 2012. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/2274.

Ondrejat:2010:SNL

[Ond10] Martin Ondrejat. Stochastic non-linear wave equations in local Sobolev spaces. Electronic Journal of Probability, 15:33:1041–33:1091, 2010. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/789.

OConnell:2015:TWA

[OO15] Neil O’Connell and Janosch Ortmann. Tracy–Widom asymptotics for a random polymer model with gamma-distributed weights. Electronic Journal of Probability, 20(??):25:1–25:18, ???? 2015. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/3787.

Oh:2021:CSN

[OO21] Tadahiro Oh and Mamoru Okamoto. Comparing the stochastic nonlinear wave and heat equations: a case study. Electronic Journal of Probability, 26(??):1– 44, 2021. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

journals/electronic-journal-of-probability/volume-26/issue-none/Comparing-
the-stochastic-nonlinear-wave-and-heat-equations--a/10.1214/20-EJP575.
full.

OConnell:2013:WVR

[OP13] Neil O’Connell and Yuchen Pei. A q-weighted version of the Robinson–Schensted algorithm. Electronic Journal of Probability, 18:95:1–95:25, 2013. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/2930.

Ouaki:2022:MSC

[OP22] Mehdi Ouaki and Jim Pitman. Markovian structure in the concave majorant of Brownian motion. Electronic Journal of Probability, 27(??):1–21, 2022. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-27/issue-none/Markovian-
structure-in-the-concave-majorant-of-Brownian-motion/10.1214/22-EJP769.
full.

Oviedo:2022:SOC

[OPR22] Giancarlos Oviedo, Gonzalo Panizo, and Alejandro F. Ramírez. Second order cubic corrections of large deviations for perturbed random walks. Electronic Journal of Probability, 27(??):1–45, 2022. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-27/issue-none/Second-order-cubic-corrections-
of-large-deviations-for-perturbed-random/10.1214/22-EJP786.full.

ORourke:2014:LRP

[OR14] Sean O’Rourke and David Renfrew. Low rank perturbations of large elliptic random matrices. Electronic Journal of Probability, 19:43:1–43:65, 2014. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/3057.

Ortgiese:2017:OPL

[OR17] Marcel Ortgiese and Matthew I. Roberts. One-point localization for branching random walk in Pareto environment. Electronic Journal of Probability, 22 (??):6:1–6:20, ???? 2017. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1484622023.

Orrieri:2020:LDI

[Orr20] Carlo Orrieri. Large deviations for interacting particle systems: joint meanfield and small-noise limit. Electronic Journal of Probability, 25(??):1–44, 2020.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-25/issue-none/Large-deviations-
for-interacting-particle-systems--joint-mean-field/10.1214/20-EJP516.
full.

Oliveira:2020:IDS

[ORS20] Roberto I. Oliveira, Guilherme H. Reis, and Lucas M. Stolerman. Interacting diffusions on sparse graphs: hydrodynamics from local weak limits. Electronic Journal of Probability, 25(??):1–35, 2020. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-25/issue-none/Interacting-diffusions-on-sparse-
graphs--hydrodynamics-from-local-weak/10.1214/20-EJP505.full.

Ortmann:2012:LDN

[Ort12] Janosch Ortmann. Large deviations for non-crossing partitions. Electronic Journal of Probability, 17:34:1–34:25, 2012. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/2007.

ORourke:2011:PIN

[OS11] Sean O’Rourke and Alexander Soshnikov. Products of independent non-Hermitian random matrices. Electronic Journal of Probability, 16:81:2219– 81:2245, 2011. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.

org/article/view/954.

Orenshtein:2020:RWR

[OS20] Tal Orenshtein and Christophe Sabot. Random walks in random hypergeometric environment. Electronic Journal of Probability, 25(??):33:1–33:21, ???? 2020.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.
ejp/1583805862.

Osekowski:2011:SSI

[Ose11] Adam Osekowski. Sharp and strict L^{p}-inequalities for Hilbert-space-valued orthogonal martingales. Electronic Journal of Probability, 16:19:531–19:551, 2011.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
865.

Osekowski:2013:SIM

[Osȩ13] Adam Osȩkowski. Sharp inequalities for martingales with values in ℓ_{∞}^{N}. Electronic Journal of Probability, 18:73:1–73:19, 2013. CODEN ???? ISSN 1083-

6489. URL http://ejp.ejpecp.org/article/view/2667.

Olla:2019:ELD

[OT19] Stefano Olla and Li-Cheng Tsai. Exceedingly large deviations of the totally asymmetric exclusion process. Electronic Journal of Probability, 24(??): 16:1–16:71, ???? 2019. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1550826099.

Ott:2022:EPC

[Ott22] Sébastien Ott. Existence and properties of connections decay rate for high temperature percolation models. Electronic Journal of Probability, 27(??):1–19, 2022. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

journals/electronic-journal-of-probability/volume-27/issue-none/Existence-
and-properties-of-connections-decay-rate-for-high-temperature/10.
1214/22-EJP822.full.

Ocone:2004:DVC

[OW04] Daniel Ocone and Ananda Weerasinghe. Degenerate variance control in the one-dimensional stationary case. Electronic Journal of Probability, 8:24:27, 2004.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
181.

ORourke:2020:LPB

[OW20] Sean O’Rourke and Noah Williams. On the local pairing behavior of critical points and roots of random polynomials. Electronic Journal of Probability, 25(??):1–68, 2020. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-25/issue-none/On-the-local-pairing-behavior-
of-critical-points-and-roots/10.1214/20-EJP499.full.

Owada:2017:FCL

[Owa17] Takashi Owada. Functional central limit theorem for subgraph counting processes. Electronic Journal of Probability, 22(??):17:1–17:38, ???? 2017. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.ejp/
1487127645.

Otto:2014:IMS

[OWW14] Felix Otto, Hendrik Weber, and Maria Westdickenberg. Invariant measure of the stochastic Allen–Cahn equation: the regime of small noise and large system size. Electronic Journal of Probability, 19:23:1–23:76, 2014. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/2813.

Osekowski:2021:SMB

[OZ21] Adam Osȩkowski and Yahui Zuo. Sharp maximal L^{p}-bounds for continuous martingales and their differential subordinates. Electronic Journal of Probability, 26(??):1–22, 2021. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-26/issue-none/Sharp-maximal-Lp-bounds-for-continuous-
martingales-and-their-differential/10.1214/21-EJP596.full.

Pagnard:2017:LLM

[Pag17] Camille Pagnard. Local limits of Markov branching trees and their volume growth. Electronic Journal of Probability, 22(??):95:1–95:53, ???? 2017. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.ejp/
1510110478.

Pain:2016:VBB

[Pai16] Michel Pain. Velocity of the L-branching Brownian motion. Electronic Journal of Probability, 21(??):28:1–28:28, ???? 2016. CODEN ???? ISSN 1083-6489.

URL https://projecteuclid.org/euclid.ejp/1460652929.

Pak:1998:RWF

[Pak98] Igor Pak. Random walks on finite groups with few random generators. Electronic Journal of Probability, 4:1:1–1:11, 1998. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/38.

Pak:1999:RWF

[Pak99] Igor Pak. Random walks on finite groups with few random generators. Electronic Journal of Probability, 4:1:1–1:11, 1999. CODEN ???? ISSN 1083-6489. URL

http://www.math.washington.edu/~ejpecp/EjpVol4/paper1.abs.html.

Panchenko:2005:CLT

[Pan05] Dmitry Panchenko. A Central Limit Theorem for weighted averages of spins in the high temperature region of the Sherrington–Kirkpatrick model. Electronic Journal of Probability, 10:14:499–14:524, 2005. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/258.

Panchenko:2014:RSS

[Pan14] Dmitry Panchenko. On the replica symmetric solution of the K-sat model. Electronic Journal of Probability, 19:67:1–67:17, 2014. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/2963.

Pardoux:2020:MDE

[Par20] Etienne Pardoux. Moderate deviations and extinction of an epidemic. Electronic Journal of Probability, 25(??):25:1–25:27, ???? 2020. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/euclid.ejp/1581994992.

Parekh:2022:PRW

[Par22] Shalin Parekh. Positive random walks and an identity for half-space SPDEs. Electronic Journal of Probability, 27(??):1–47, 2022. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-27/issue-none/Positive-random-walks-and-an-
identity-for-half-space-SPDEs/10.1214/22-EJP775.full.

Pascu:2011:MCR

[Pas11] Mihai Pascu. Mirror coupling of reflecting Brownian motion and an application to Chavel’s conjecture. Electronic Journal of Probability, 16:18:504–18:530, 2011. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/

view/859.

Paulin:2014:CDI

[Pau14] Daniel Paulin. The convex distance inequality for dependent random variables, with applications to the stochastic travelling salesman and other problems. Electronic Journal of Probability, 19:68:1–68:34, 2014. CODEN ???? ISSN 1083-

6489. URL http://ejp.ejpecp.org/article/view/3261.

Paulin:2015:CIM

[Pau15] Daniel Paulin. Concentration inequalities for Markov chains by Marton couplings and spectral methods. Electronic Journal of Probability, 20(??):79:1– 79:32, ???? 2015. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.

org/article/view/4039.

Peligrad:1999:CSS

[Pel99] Magda Peligrad. Convergence of stopped sums of weakly dependent random variables. Electronic Journal of Probability, 4:13:1–13:13, 1999. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/50.

Penrose:2007:GLR

[Pen07] Mathew Penrose. Gaussian limts for random geometric measures. Electronic Journal of Probability, 12:35:989–35:1035, 2007. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/429.

Penrose:2009:NAI

[Pen09] Mathew Penrose. Normal approximation for isolated balls in an urn allocation model. Electronic Journal of Probability, 14:74:2155–74:2181, 2009. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/699.

Penrose:2020:LLP

[Pen20] Mathew D. Penrose. Leaves on the line and in the plane. Electronic Journal of Probability, 25(??):53:1–53:40, ???? 2020. CODEN ???? ISSN 1083-6489. URL

https://projecteuclid.org/euclid.ejp/1588644036.

Pene:2021:LTA

[Pèn21] Françoise Pène. Limit theorems for additive functionals of random walks in random scenery. Electronic Journal of Probability, 26(18):1–46, February 2021.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-26/issue-none/Limit-theorems-
for-additive-functionals-of-random-walks-in-random/10.1214/21-EJP696.
full.

Perlman:2020:RPS

[Per20a] Michael D. Perlman. Are random permutations spherically uniform? Electronic Journal of Probability, 25(??):9:1–9:26, ???? 2020. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/euclid.ejp/1580267009.

Perruchaud:2020:HAK

[Per20b] Pierre Perruchaud. Homogenisation for anisotropic kinetic random motions. Electronic Journal of Probability, 25(??):39:1–39:26, ???? 2020. CODEN

???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.ejp/
1585620094.

Peretz:2021:MDS

[Per21] Tal Peretz. Moderate deviations for the self-normalized random walk in random scenery. Electronic Journal of Probability, 26(18):1–16, February 2021.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-26/issue-none/Moderate-deviations-
for-the-self-normalized-random-walk-in-random/10.1214/21-EJP607.full.

Peski:2022:TPD

[Pes22] Roger Van Peski. q-TASEP with position-dependent slowing. Electronic Journal of Probability, 27(??):1–35, 2022. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-27/issue-none/q-TASEP-with-position-dependent-
slowing/10.1214/22-EJP876.full.

Petrou:2008:MCL

[Pet08] Evangelia Petrou. Malliavin calculus in Lévy spaces and applications to finance. Electronic Journal of Probability, 13:27:852–27:879, 2008. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/502.

Peterson:2010:SOD

[Pet10] Jonathon Peterson. Systems of one-dimensional random walks in a common random environment. Electronic Journal of Probability, 15:32:1024–32:1040, 2010. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/

view/784.

Petrov:2011:PSD

[Pet11] Leonid Petrov. Pfaffian stochastic dynamics of strict partitions. Electronic Journal of Probability, 16:82:2246–82:2295, 2011. CODEN ???? ISSN 1083-

6489. URL http://ejp.ejpecp.org/article/view/956.

Peterson:2012:LDS

[Pet12] Jonathon Peterson. Large deviations and slowdown asymptotics for onedimensional excited random walks. Electronic Journal of Probability, 17:48:1– 48:24, 2012. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/1726.

Petrov:2020:PIS

[Pet20] Leonid Petrov. PushTASEP in inhomogeneous space. Electronic Journal of Probability, 25(??):1–25, 2020. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-25/issue-none/PushTASEP-in-inhomogeneous-space/
10.1214/20-EJP517.full.

Pham:2015:MRS

[Pha15] Cong Dan Pham. Monotonicity and regularity of the speed for excited random walks in higher dimensions. Electronic Journal of Probability, 20(??):72:1–72:25, ???? 2015. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/3522.

Picard:2006:BES

[Pic06] Jean Picard. Brownian excursions, stochastic integrals, and representation of Wiener functionals. Electronic Journal of Probability, 11:8:199–8:248, 2006.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
310.

Pinelis:2006:NDS

[Pin06] Iosif Pinelis. On normal domination of (super)martingales. Electronic Journal of Probability, 11:39:1049–39:1070, 2006. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/371.

Pinelis:2009:OTV

[Pin09] Iosif Pinelis. Optimal two-value zero-mean disintegration of zero-mean random variables. Electronic Journal of Probability, 14:26:663–26:727, 2009. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/633.

Pinelis:2012:AGB

[Pin12] Iosif Pinelis. An asymptotically Gaussian bound on the Rademacher tails. Electronic Journal of Probability, 17:35:1–35:22, 2012. CODEN ???? ISSN 1083-

6489. URL http://ejp.ejpecp.org/article/view/2026.

Pinsky:2013:DTR

[Pin13] Ross Pinsky. Detecting tampering in a random hypercube. Electronic Journal of Probability, 18:28:1–28:12, 2013. CODEN ???? ISSN 1083-6489. URL http:

//ejp.ejpecp.org/article/view/2290.

Pinelis:2016:OBP

[Pin16a] Iosif Pinelis. Optimal binomial, Poisson, and normal left-tail domination for sums of nonnegative random variables. Electronic Journal of Probability, 21 (??):20:1–20:19, ???? 2016. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1457706456.

Pinsky:2016:TRG

[Pin16b] Ross G. Pinsky. Transience/recurrence and growth rates for diffusion processes in time-dependent regions. Electronic Journal of Probability, 21(??):46:1–46:24, ???? 2016. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.

org/euclid.ejp/1469557139.

Pinsky:2018:SDA

[Pin18] Ross G. Pinsky. On the strange domain of attraction to generalized Dickman distributions for sums of independent random variables. Electronic Journal of Probability, 23(??):3:1–3:17, ???? 2018. CODEN ???? ISSN 1083-6489. URL

https://projecteuclid.org/euclid.ejp/1516093311.

Pinsky:2019:ODD

[Pin19] Ross G. Pinsky. Optimizing the drift in a diffusive search for a random stationary target. Electronic Journal of Probability, 24(??):82:1–82:22, ???? 2019.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.
ejp/1568080861.

Pinsky:2022:CCN

[Pin22] Ross G. Pinsky. Clustering of consecutive numbers in permutations under Mallows distributions and super-clustering under general p-shifted distributions. Electronic Journal of Probability, 27(??):1–20, 2022. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-27/issue-none/Clustering-of-consecutive-numbers-
in-permutations-under-Mallows-distributions-and/10.1214/22-EJP812.
full.

Pitman:1999:BMB

[Pit99] Jim Pitman. Brownian motion, bridge, excursion, and meander characterized by sampling at independent uniform times. Electronic Journal of Probability, 4:11:1–11:33, 1999. CODEN ???? ISSN 1083-6489. URL http://

ejp.ejpecp.org/article/view/48;   http://www.math.washington.edu/
~ejpecp/EjpVol4/paper11.abs.html.

Possamai:2015:QBJ

[PKTZ15] Dylan Possamai, Nabil Kazi-Tani, and Chao Zhou. Quadratic BSDEs with jumps: a fixed-point approach. Electronic Journal of Probability, 20(??):66:1– 66:28, ???? 2015. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.

org/article/view/3363.

Piaggio:2016:EZE

[PL16] Matías Carrasco Piaggio and Pablo Lessa. Equivalence of zero entropy and the Liouville property for stationary random graphs. Electronic Journal of Probability, 21(??):55:1–55:24, ???? 2016. CODEN ???? ISSN 1083-6489. URL

https://projecteuclid.org/euclid.ejp/1473188082.

Piera:2008:CRR

[PM08] Francisco Piera and Ravi Mazumdar. Comparison results for reflected jumpdiffusions in the orthant with variable reflection directions and stability applications. Electronic Journal of Probability, 13:61:1886–61:1908, 2008. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/569.

Picco:2009:ODR

[PO09] Pierre Picco and Enza Orlandi. One-dimensional random field Kac’s model: weak large deviations principle. Electronic Journal of Probability, 14:47:1372– 47:1416, 2009. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/662.

Posta:2005:EFO

[Pos05] Gustavo Posta. Equilibrium fluctuations for a one-dimensional interface in the solid on solid approximation. Electronic Journal of Probability, 10:29:962– 29:987, 2005. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/270.

Powell:2018:CGC

[Pow18] Ellen Powell. Critical Gaussian chaos: convergence and uniqueness in the derivative normalisation. Electronic Journal of Probability, 23(??):31:1–31:26, ???? 2018. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

euclid.ejp/1522375271.

Penrose:2011:LCL

[PP11] Mathew Penrose and Yuval Peres. Local Central Limit Theorems in stochastic geometry. Electronic Journal of Probability, 16:91:2509–91:2544, 2011. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/968.

Perkowski:2015:LTT

[PP15] Nicolas Perkowski and David J. Prömel. Local times for typical price paths and pathwise Tanaka formulas. Electronic Journal of Probability, 20(??):46:1–46:15, ???? 2015. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/3534.

Pemantle:2001:WDB

[PPPY01] Robin Pemantle, Yuval Peres, Jim Pitman, and Marc Yor. Where did the Brownian particle go? Electronic Journal of Probability, 6:10:1–10:22, 2001.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
83.

Papapantoleon:2018:EUR

[PPS18] Antonis Papapantoleon, Dylan Possamaï, and Alexandros Saplaouras. Existence and uniqueness results for BSDE with jumps: the whole nine yards. Electronic Journal of Probability, 23(??):121:1–121:68, ???? 2018. CODEN ???? ISSN

1083-6489. URL https://projecteuclid.org/euclid.ejp/1545102139.

Pakkanen:2021:LTT

[PPSV21] Mikko S. Pakkanen, Riccardo Passeggeri, Orimar Sauri, and Almut E. D. Veraart. Limit theorems for trawl processes. Electronic Journal of Probability, 26(18):1–36, February 2021. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-26/issue-none/Limit-theorems-for-trawl-processes/
10.1214/21-EJP652.full.

Peres:2004:MTR

[PR04] Yuval Peres and David Revelle. Mixing times for random walks on finite lamplighter groups. Electronic Journal of Probability, 9:26:825–26:845, 2004. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/198.

Puchala:2005:EAT

[PR05] Zbigniew Puchala and Tomasz Rolski. The exact asymptotic of the time to collision. Electronic Journal of Probability, 10:40:1359–40:1380, 2005. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/291.

Pekoz:2011:EAN

[PR11] Erol Peköz and Adrian Röllin. Exponential approximation for the nearly critical Galton–Watson process and occupation times of Markov chains. Electronic Journal of Probability, 16:51:1381–51:1393, 2011. CODEN ???? ISSN 1083-

6489. URL http://ejp.ejpecp.org/article/view/914.

Procaccia:2012:NSM

[PR12] Eviatar Procaccia and Ron Rosenthal. The need for speed: maximizing the speed of random walk in fixed environments. Electronic Journal of Probability, 17:13:1–13:19, 2012. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.

org/article/view/1800.

Perkowski:2019:KER

[PR19] Nicolas Perkowski and Tommaso Cornelis Rosati. The KPZ equation on the real line. Electronic Journal of Probability, 24(??):117:1–117:56, ???? 2019.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.
ejp/1572314777.

Panloup:2020:SEC

[PR20] Fabien Panloup and Alexandre Richard. Sub-exponential convergence to equilibrium for Gaussian driven stochastic differential equations with semi-contractive drift. Electronic Journal of Probability, 25(??):62:1–62:43, ???? 2020. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.ejp/
1591084854.

Peskir:2023:SFD

[PR23] Goran Peskir and David Roodman. Sticky Feller diffusions. Electronic Journal of Probability, 28(??):1–28, 2023. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-28/issue-none/Sticky-Feller-diffusions/10.1214/
23-EJP909.full.

Profeta:2012:PNR

[Pro12] Christophe Profeta. Penalizing null recurrent diffusions. Electronic Journal of Probability, 17:69:1–69:23, 2012. CODEN ???? ISSN 1083-6489. URL http://

ejp.ejpecp.org/article/view/2267.

Profeta:2021:AUS

[Pro21] Christophe Profeta. The area under a spectrally positive stable excursion and other related processes. Electronic Journal of Probability, 26(18):1–21, February 2021. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

journals/electronic-journal-of-probability/volume-26/issue-none/The-
area-under-a-spectrally-positive-stable-excursion-and-other/10.1214/
21-EJP618.full.

Pekoz:2020:ELA

[PRR20] Erol A. Peköz, Adrian Röllin, and Nathan Ross. Exponential and Laplace approximation for occupation statistics of branching random walk. Electronic Journal of Probability, 25(??):55:1–55:22, ???? 2020. CODEN ???? ISSN

1083-6489. URL https://projecteuclid.org/euclid.ejp/1588644038.

Penington:2022:GSD

[PRT22] Sarah Penington, Matthew I. Roberts, and Zsófia Talyigás. Genealogy and spatial distribution of the N-particle branching random walk with polynomial tails. Electronic Journal of Probability, 27(??):1–65, 2022. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-27/issue-none/Genealogy-
and-spatial-distribution-of-the-N-particle-branching-random/10.1214/
22-EJP806.full.

Popov:2020:TCW

[PRU20] Serguei Popov, Leonardo T. Rolla, and Daniel Ungaretti. Transience of conditioned walks on the plane: encounters and speed of escape. Electronic Journal of Probability, 25(??):52:1–52:23, ???? 2020. CODEN ???? ISSN 1083-6489.

URL https://projecteuclid.org/euclid.ejp/1588125887.

Pitman:2014:RTG

[PRW14] Jim Pitman, Douglas Rizzolo, and Matthias Winkel. Regenerative tree growth: structural results and convergence. Electronic Journal of Probability, 19:70:1– 70:27, 2014. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/3040.

Procaccia:2020:SDW

[PRZ20] Eviatar B. Procaccia, Ron Rosenthal, and Yuan Zhang. Stabilization of DLA in a wedge. Electronic Journal of Probability, 25(??):42:1–42:22, ???? 2020.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.
ejp/1585879250.

Patie:2012:EFE

[PS12] Pierre Patie and Mladen Savov. Extended factorizations of exponential functionals of Lévy processes. Electronic Journal of Probability, 17:38:1–38:22, 2012.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
2057.

Patie:2018:BGF

[PS18a] Pierre Patie and Mladen Savov. Bernstein-gamma functions and exponential functionals of Lévy processes. Electronic Journal of Probability, 23(??): 75:1–75:101, ???? 2018. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1532678638.

Privault:2018:SAF

[PS18b] Nicolas Privault and Grzegorz Serafin. Stein approximation for functionals of independent random sequences. Electronic Journal of Probability, 23(??):4:1–4:34, ???? 2018. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.

org/euclid.ejp/1517367680.

Profeta:2019:CES

[PS19] Christophe Profeta and Thomas Simon. Cramér’s estimate for stable processes with power drift. Electronic Journal of Probability, 24(??):17:1–17:21, ???? 2019.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.
ejp/1551150461.

Pianoforte:2021:PAA

[PS21] Federico Pianoforte and Matthias Schulte. Poisson approximation with applications to stochastic geometry. Electronic Journal of Probability, 26(18):1–36, February 2021. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.

org/journals/electronic-journal-of-probability/volume-26/issue-none/
Poisson-approximation-with-applications-to-stochastic-geometry/10.
1214/21-EJP723.full.

Park:2022:SHC

[PS22a] Hyunchul Park and Renming Song. Spectral heat content for α-stable processes in C^{1,1} open sets. Electronic Journal of Probability, 27(??):1–19, 2022.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-27/issue-none/Spectral-heat-
content-for-%ce%b1-stable-processes-in-C11-open/10.1214/22-EJP752.
full.

Privault:2022:BEB

[PS22b] Nicolas Privault and Grzegorz Serafin. Berry–Esseen bounds for functionals of independent random variables. Electronic Journal of Probability, 27(??):1–37, 2022. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

journals/electronic-journal-of-probability/volume-27/issue-none/Berry-
-Esseen-bounds-for-functionals-of-independent-random-variables/10.
1214/22-EJP795.full.

Peres:2017:IMT

[PSSS17] Yuval Peres, Thomas Sauerwald, Perla Sousi, and Alexandre Stauffer. Intersection and mixing times for reversible chains. Electronic Journal of Probability, 22(??):12:1–12:16, ???? 2017. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1486090892.

Petrelis:2017:SLU

[PST17] Nicolas Pétrélis, Rongfeng Sun, and Niccolò Torri. Scaling limit of the uniform prudent walk. Electronic Journal of Probability, 22(??):66:1–66:19, ???? 2017.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.
ejp/1504749661.

Pipiras:2007:IRP

[PT07] Vladas Pipiras and Murad Taqqu. Integral representations of periodic and cyclic fractional stable motions. Electronic Journal of Probability, 12:7:181–7:206, 2007. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/

view/395.

Privault:2013:PAC

[PT13] Nicolas Privault and Giovanni Luca Torrisi. Probability approximation by Clark–Ocone covariance representation. Electronic Journal of Probability, 18: 91:1–91:25, 2013. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.

org/article/view/2787.

Pitman:2015:SZS

[PT15] Jim Pitman and Wenpin Tang. The Slepian zero set, and Brownian bridge embedded in Brownian motion by a spacetime shift. Electronic Journal of Probability, 20(??):61:1–61:28, ???? 2015. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/3911.

Pitman:2018:APR

[PT18] Jim Pitman and Wenpin Tang. The argmin process of random walks, Brownian motion and Lévy processes. Electronic Journal of Probability, 23(??):60:1–60:35, ???? 2018. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.

org/euclid.ejp/1529460158.

Puhalskii:2004:SDL

[Puh04] Anatolii Puhalskii. On some degenerate large deviation problems. Electronic Journal of Probability, 9:28:862–28:886, 2004. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/232.

Pfaffelhuber:2023:DPM

[PW23] Peter Pfaffelhuber and Anton Wakolbinger. A diploid population model for copy number variation of genetic elements. Electronic Journal of Probability, 28(??): 1–15, 2023. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.

org/journals/electronic-journal-of-probability/volume-28/issue-none/
A-diploid-population-model-for-copy-number-variation-of-genetic/10.
1214/23-EJP934.full.

Pitman:1996:RDD

[PY96] Jim Pitman and Marc Yor. Random discrete distributions derived from selfsimilar random sets. Electronic Journal of Probability, 1:4:1–4:28, 1996. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/4.

Pitman:1999:LMB

[PY99] Jim Pitman and Marc Yor. The law of the maximum of a Bessel bridge. Electronic Journal of Probability, 4:15:1–15:35, 1999. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/52;   http:
//www.math.washington.edu/~ejpecp/EjpVol4/paper15.abs.html.

Pitman:2017:EGS

[PY17] Jim Pitman and Yuri Yakubovich. Extremes and gaps in sampling from a GEM random discrete distribution. Electronic Journal of Probability, 22(??): 44:1–44:26, ???? 2017. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1493777020.

Peng:2022:WPS

[PYZ22] Xuhui Peng, Juan Yang, and Jianliang Zhai. Well-posedness of stochastic 2D hydrodynamics type systems with multiplicative Lévy noises. Electronic Journal of Probability, 27(??):1–31, 2022. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-27/issue-none/Well-posedness-of-stochastic-
2D-hydrodynamics-type-systems-with-multiplicative/10.1214/22-EJP779.
full.

Pham:2013:SNE

[PZ13] Triet Pham and Jianfeng Zhang. Some norm estimates for semimartingales. Electronic Journal of Probability, 18:109:1–109:25, 2013. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/2406.

Procaccia:2020:CMP

[PZ20] Eviatar B. Procaccia and Yuan Zhang. On covering monotonic paths with simple random walk. Electronic Journal of Probability, 25(??):1–39, 2020. CODEN ????

ISSN 1083-6489. URL https://projecteuclid.org/journals/electronic-
journal-of-probability/volume-25/issue-none/On-covering-monotonic-
paths-with-simple-random-walk/10.1214/20-EJP545.full.

Pirogov:2022:CPG

[PZ22] Sergey Pirogov and Elena Zhizhina. Contact processes on general spaces. models on graphs and on manifolds. Electronic Journal of Probability, 27(??):1–14, 2022.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-27/issue-none/Contact-processes-
on-general-spaces-Models-on-graphs-and-on/10.1214/22-EJP765.full.

Qian:2018:LPP

[QW18] Wei Qian and Wendelin Werner. The law of a point process of Brownian excursions in a domain is determined by the law of its trace. Electronic Journal of Probability, 23(??):128:1–128:23, ???? 2018. CODEN ???? ISSN 1083-6489.

URL https://projecteuclid.org/euclid.ejp/1545210235.

Qinwen:2014:JCS

[QZJ14] Wang Qinwen, Su Zhonggen, and Yao Jianfeng. Joint CLT for several random sesquilinear forms with applications to large-dimensional spiked population models. Electronic Journal of Probability, 19(??):103:1–103:28, ???? 2014.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
3339.

Reynolds:2008:DRS

[RA08] David Reynolds and John Appleby. Decay rates of solutions of linear stochastic Volterra equations. Electronic Journal of Probability, 13:30:922–30:943, 2008.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
507.

Ramanan:2006:RDD

[Ram06] Kavita Ramanan. Reflected diffusions defined via the extended Skorokhod map. Electronic Journal of Probability, 11:36:934–36:992, 2006. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/360.

Ramil:2022:QSD

[Ram22] Mouad Ramil. Quasi-stationary distribution for the Langevin process in cylindrical domains, part II: overdamped limit. Electronic Journal of Probability, 27(??):1–18, 2022. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-27/issue-none/Quasi-stationary-distribution-
for-the-Langevin-process-in-cylindrical-domains/10.1214/22-EJP789.
full.

Rapenne:2023:IMC

[Rap23] Valentin Rapenne. Invariant measures of critical branching random walks in high dimension. Electronic Journal of Probability, 28(??):1–38, 2023. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-28/issue-none/Invariant-
measures-of-critical-branching-random-walks-in-high-dimension/10.
1214/23-EJP906.full.

Rassoul-Agha:2017:AVQ

[RASY17] Firas Rassoul-Agha, Timo Seppäläinen, and Atilla Yilmaz. Averaged vs. quenched large deviations and entropy for random walk in a dynamic random environment. Electronic Journal of Probability, 22(??):57:1–57:47, ???? 2017.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.
ejp/1499306456.

Ray:2014:GPH

[Ray14] Gourab Ray. Geometry and percolation on half planar triangulations. Electronic Journal of Probability, 19:47:1–47:28, 2014. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/3238.

Redig:2009:CIM

[RC09] Frank Redig and Jean Rene Chazottes. Concentration inequalities for Markov processes via coupling. Electronic Journal of Probability, 14:40:1162–40:1180, 2009. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/

view/657.

Reddy:2017:LED

[Red17] Tulasi Ram Reddy. Limiting empirical distribution of zeros and critical points of random polynomials agree in general. Electronic Journal of Probability, 22 (??):74:1–74:18, ???? 2017. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1505268105.

Rhodes:2014:HKD

[RGV14] Rémi Rhodes, Christophe Garban, and Vincent Vargas. On the heat kernel and the Dirichlet form of Liouville Brownian motion. Electronic Journal of Probability, 19(??):96:1–96:25, ???? 2014. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/2950.

Rhodes:2010:SHR

[Rho10] Remi Rhodes. Stochastic homogenization of reflected stochastic differential equations. Electronic Journal of Probability, 15:31:989–31:1023, 2010. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/776.

Richier:2015:UAC

[Ric15] Loïc Richier. Universal aspects of critical percolation on random half-planar maps. Electronic Journal of Probability, 20(??):129:1–129:45, ???? 2015. CO-

DEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/

Richier:2018:IIC

[Ric18] Loïc Richier. The incipient infinite cluster of the uniform infinite half-planar triangulation. Electronic Journal of Probability, 23(??):89:1–89:38, ???? 2018.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.
ejp/1536717748.

Riedel:2017:TCI

[Rie17] Sebastian Riedel. Transportation-cost inequalities for diffusions driven by Gaussian processes. Electronic Journal of Probability, 22(??):24:1–24:26, ???? 2017.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.
ejp/1488596710.

Rivera:2021:TIP

[Riv21] Alejandro Rivera. Talagrand’s inequality in planar Gaussian field percolation. Electronic Journal of Probability, 26(??):1–25, 2021. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-26/issue-none/Talagrands-inequality-in-planar-
Gaussian-field-percolation/10.1214/21-EJP585.full.

Rivera-Lopez:2022:LCO

[RLR22] Kelvin Rivera-Lopez and Douglas Rizzolo. The leftmost column of ordered Chinese restaurant process up-down chains: intertwining and convergence. Electronic Journal of Probability, 27(??):1–22, 2022. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-27/issue-none/The-leftmost-column-of-ordered-
Chinese-restaurant-process-up-down/10.1214/22-EJP843.full.

Rosen:2009:ECP

[RM09] Jay Rosen and Michael Marcus. Existence of a critical point for the infinite divisibility of squares of Gaussian vectors in R^{2} with non–zero mean. Electronic Journal of Probability, 14:48:1417–48:1455, 2009. CODEN ???? ISSN 1083-

6489. URL http://ejp.ejpecp.org/article/view/669.

Richou:2013:NES

[RM13] Adrien Richou and Federica Masiero. A note on the existence of solutions to Markovian superquadratic BSDEs with an unbounded terminal condition. Electronic Journal of Probability, 18:50:1–50:15, 2013. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/2124.

Reinert:2010:SMS

[RNP10] Gesine Reinert, Ivan Nourdin, and Giovanni Peccati. Stein’s method and stochastic analysis of Rademacher functionals. Electronic Journal of Probability, 15:55:1703–55:1742, 2010. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/823.

Roberts:2015:FAC

[Rob15] Matthew Iain Roberts. Fine asymptotics for the consistent maximal displacement of branching Brownian motion. Electronic Journal of Probability, 20(??): 28:1–28:26, ???? 2015. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/2912.

Roellin:2008:SCB

[Roe08] Adrian Roellin. Symmetric and centered binomial approximation of sums of locally dependent random variables. Electronic Journal of Probability, 13:24:756– 24:776, 2008. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/503.

Romito:2018:SME

[Rom18] Marco Romito. A simple method for the existence of a density for stochastic evolutions with rough coefficients. Electronic Journal of Probability, 23(??): 113:1–113:43, ???? 2018. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1542942364.

Rosen:1997:LIL

[Ros97] Jay Rosen. Laws of the iterated logarithm for triple intersections of three dimensional random walks. Electronic Journal of Probability, 2:2:1–2:32, 1997.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
16.

Rosen:2023:TTP

[Ros23] Jay Rosen. Tightness for thick points in two dimensions. Electronic Journal of Probability, 28(??):1–45, 2023. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-28/issue-none/Tightness-for-thick-points-in-
two-dimensions/10.1214/23-EJP910.full.

Rousselle:2015:QIP

[Rou15] Arnaud Rousselle. Quenched invariance principle for random walks on Delaunay triangulations. Electronic Journal of Probability, 20(??):33:1–33:32, ???? 2015.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
4006.

Rousselin:2018:IMH

[Rou18] Pierre Rousselin. Invariant measures, Hausdorff dimension and dimension drop of some harmonic measures on Galton–Watson trees. Electronic Journal of Probability, 23(??):46:1–46:31, ???? 2018. CODEN ???? ISSN 1083-6489. URL

https://projecteuclid.org/euclid.ejp/1527213727.

Rossignol:2012:GPF

[RP12] Raphaël Rossignol and Leandro Pimentel. Greedy polyominoes and first-passage times on random Voronoi tilings. Electronic Journal of Probability, 17:12:1– 12:31, 2012. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/1788.

Roberts:1996:QBC

[RR96] Gareth O. Roberts and Jeffrey S. Rosenthal. Quantitative bounds for convergence rates of continuous time Markov processes. Electronic Journal of Probability, 1:9:1–9:21, 1996. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/9.

Rohan:2013:GEA

[RR13] Neelabh Rohan and T. V. Ramanathan. Geometric ergodicity of asymmetric volatility models with stochastic parameters. Electronic Journal of Probability, 18:90:1–90:12, 2013. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.

org/article/view/1871.

Ramirez:2022:CCB

[RR22] Alejandro F. Ramírez and Rodrigo Ribeiro. Computable criteria for ballisticity of random walks in elliptic random environment. Electronic Journal of Probability, 27(??):1–38, 2022. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-27/issue-none/Computable-criteria-for-ballisticity-
of-random-walks-in-elliptic-random/10.1214/22-EJP856.full.

Raimond:2009:SGR

[RS09] Olivier Raimond and Bruno Schapira. On some generalized reinforced random walk on integers. Electronic Journal of Probability, 14:60:1770–60:1789, 2009.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
685.

Rozkosz:2012:SRE

[RS12a] Andrzej Rozkosz and Leszek Slominski. Stochastic representation of entropy solutions of semilinear elliptic obstacle problems with measure data. Electronic Journal of Probability, 17:40:1–40:27, 2012. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/2062.

Ruschendorf:2012:OSC

[RS12b] Ludger Rüschendorf and Tomonari Sei. On optimal stationary couplings between stationary processes. Electronic Journal of Probability, 17:17:1–17:20, 2012. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/

view/1797.

Rath:2013:ESQ

[RS13a] Balázs Ráth and Artëm Sapozhnikov. The effect of small quenched noise on connectivity properties of random interlacements. Electronic Journal of Probability, 18:4:1–4:20, 2013. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.

org/article/view/2122.

Rippl:2013:NRP

[RS13b] Thomas Rippl and Anja Sturm. New results on pathwise uniqueness for the heat equation with colored noise. Electronic Journal of Probability, 18:77:1– 77:46, 2013. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/2506.

Ramirez:2019:NEB

[RS19] Alejandro F. Ramírez and Santiago Saglietti. New examples of ballistic RWRE in the low disorder regime. Electronic Journal of Probability, 24(??): 127:1–127:20, ???? 2019. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1573268591.

Rackauskas:2021:AMW

[RS21a] Alfredas Račkauskas and Charles Suquet. On the asymptotic of the maximal weighted increment of a random walk with regularly varying jumps: the boundary case. Electronic Journal of Probability, 26(18):1–31, February 2021.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-26/issue-none/On-the-asymptotic-
of-the-maximal-weighted-increment-of-a/10.1214/21-EJP691.full.

Roberts:2021:GPD

[RS21b] Matthew I. Roberts and Jason Schweinsberg. A Gaussian particle distribution for branching Brownian motion with an inhomogeneous branching rate. Electronic Journal of Probability, 26(18):1–76, February 2021. CODEN ????

ISSN 1083-6489. URL https://projecteuclid.org/journals/electronic-
journal-of-probability/volume-26/issue-none/A-Gaussian-particle-distribution-
for-branching-Brownian-motion-with-an/10.1214/21-EJP673.full.

Redig:2020:SSE

[RSS20] Frank Redig, Ellen Saada, and Federico Sau. Symmetric simple exclusion process in dynamic environment: hydrodynamics. Electronic Journal of Probability, 25(??):1–47, 2020. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-

of-probability/volume-25/issue-none/Symmetric-simple-exclusion-process-
in-dynamic-environment-hydrodynamics/10.1214/20-EJP536.full.

Rath:2022:PTB

[RSS22] Balázs Ráth, Jan M. Swart, and Márton Szőke. A phase transition between endogeny and nonendogeny. Electronic Journal of Probability, 27(??):1–43, 2022.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-27/issue-none/A-phase-transition-
between-endogeny-and-nonendogeny/10.1214/22-EJP872.full.

Ren:2019:SCL

[RSSZ19] Yan-Xia Ren, Renming Song, Zhenyao Sun, and Jianjie Zhao. Stable central limit theorems for super Ornstein–Uhlenbeck processes. Electronic Journal of Probability, 24(??):141:1–141:42, ???? 2019. CODEN ???? ISSN 1083-6489.

URL https://projecteuclid.org/euclid.ejp/1576638110.

Ren:2018:WDS

[RSZ18] Yan-Xia Ren, Renming Song, and Rui Zhang. Williams decomposition for superprocesses. Electronic Journal of Probability, 23(??):23:1–23:33, ???? 2018.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.
ejp/1519722152.

Rath:2009:ERR

[RT09] Balazs Rath and Balint Toth. Erdős–Rényi random graphs + forest fires = self-organized criticality. Electronic Journal of Probability, 14:45:1290–45:1327, 2009. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/

view/653.

Revelle:2010:CCR

[RT10] David Revelle and Russ Thompson. Critical constants for recurrence on groups of polynomial growth. Electronic Journal of Probability, 15:23:710–23:722, 2010.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
773.

Rossignol:2018:ECF

[RT18] Raphaël Rossignol and Marie Théret. Existence and continuity of the flow constant in first passage percolation. Electronic Journal of Probability, 23 (??):99:1–99:42, ???? 2018. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1537927580.

Riedler:2012:LTI

[RTW12] Martin Riedler, Michèle Thieullen, and Gilles Wainrib. Limit theorems for infinite-dimensional piecewise deterministic Markov processes. Applications to stochastic excitable membrane models. Electronic Journal of Probability, 17: 55:1–55:48, 2012. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.

org/article/view/1946.

Rider:2007:CDP

[RV07] Brian Rider and Balint Virag. Complex determinantal processes and H1 noise. Electronic Journal of Probability, 12:45:1238–45:1257, 2007. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/446.

Rhodes:2010:MMR

[RV10] Rémi Rhodes and Vincent Vargas. Multidimensional multifractal random measures. Electronic Journal of Probability, 15:9:241–9:258, 2010. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/746.

Rebeschini:2015:PTN

[RvH15] Patrick Rebeschini and Ramon van Handel. Phase transitions in nonlinear filtering. Electronic Journal of Probability, 20(??):7:1–7:46, ???? 2015. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/3281.

Ross:2018:SLS

[RW18] Nathan Ross and Yuting Wen. Scaling limits for some random trees constructed inhomogeneously. Electronic Journal of Probability, 23(??):5:1–5:35, ???? 2018.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.
ejp/1517626965.

Riedel:2013:SPD

[RX13] Sebastian Riedel and Weijun Xu. A simple proof of distance bounds for Gaussian rough paths. Electronic Journal of Probability, 18:108:1–108:18, 2013. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/2387.

Rios-Zertuche:2014:PDN

[RZ14] Rodolfo Rios-Zertuche. The pillowcase distribution and near-involutions. Electronic Journal of Probability, 19(??):116:1–116:22, ???? 2014. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/3626.

Sakagawa:2009:CTD

[Sak09] Hironobu Sakagawa. Confinement of the two dimensional discrete Gaussian free field between two hard walls. Electronic Journal of Probability, 14:79:2310– 79:2327, 2009. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/711.

Salins:2022:GSS

[Sal22] Michael Salins. Global solutions for the stochastic reaction-diffusion equation with super-linear multiplicative noise and strong dissipativity. Electronic Journal of Probability, 27(??):1–17, 2022. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-27/issue-none/Global-solutions-for-the-stochastic-
reaction-diffusion-equation-with-super/10.1214/22-EJP740.full.

Sami:2009:LER

[Sam09] Mustapha Sami. Lower estimates for random walks on a class of amenable p-adic groups. Electronic Journal of Probability, 14:51:1513–51:1531, 2009. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/667.

Samson:2017:TEI

[Sam17] Paul-Marie Samson. Transport-entropy inequalities on locally acting groups of permutations. Electronic Journal of Probability, 22(??):62:1–62:33, ???? 2017.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.
ejp/1502244025.

Sarantsev:2015:TSC

[Sar15] Andrey Sarantsev. Triple and simultaneous collisions of competing Brownian particles. Electronic Journal of Probability, 20(??):29:1–29:28, ???? 2015. CO-

DEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
3279.

Savov:2008:CCR

[Sav08] Mladen Savov. Curve crossing for the reflected Lévy process at zero and infinity. Electronic Journal of Probability, 13:7:157–7:172, 2008. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/483.

Stivanello:2021:LTL

[SBB^{+}21] Samuele Stivanello, Gianmarco Bet, Alessandra Bianchi, Marco Lenci, and Elena Magnanini. Limit theorems for Lévy flights on a 1D Lévy random medium. Electronic Journal of Probability, 26(18):1–25, February 2021. CODEN ????

ISSN 1083-6489. URL https://projecteuclid.org/journals/electronic-
journal-of-probability/volume-26/issue-none/Limit-theorems-for-L%
c3%a9vy-flights-on-a-1D-L%c3%a9vy-random/10.1214/21-EJP626.full.

Schweinsberg:2000:CSM

[Sch00] Jason Schweinsberg. Coalescents with simultaneous multiple collisions. Electronic Journal of Probability, 5:12:1–12:50, 2000. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/68.

Schuhmacher:2005:DEP

[Sch05] Dominic Schuhmacher. Distance estimates for Poisson process approximations of dependent thinnings. Electronic Journal of Probability, 10:5:165–5:201, 2005.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
237.

Schmitz:2006:ECD

[Sch06] Tom Schmitz. Examples of condition (T) for diffusions in a random environment. Electronic Journal of Probability, 11:22:540–22:562, 2006. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/337.

Schweinsberg:2008:WM

[Sch08] Jason Schweinsberg. Waiting for m mutations. Electronic Journal of Probability, 13:52:1442–52:1478, 2008. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/540.

Schuhmacher:2009:DED

[Sch09] Dominic Schuhmacher. Distance estimates for dependent thinnings of point processes with densities. Electronic Journal of Probability, 14:38:1080–38:1116, 2009. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/

view/643.

Schweinsberg:2012:DEB

[Sch12] Jason Schweinsberg. Dynamics of the evolving Bolthausen–Sznitman coalescent. Electronic Journal of Probability, 17:91:1–91:50, 2012. CODEN ???? ISSN 1083-

6489. URL http://ejp.ejpecp.org/article/view/2378.

Schweinsberg:2017:RRPa

[Sch17a] Jason Schweinsberg. Rigorous results for a population model with selection I: evolution of the fitness distribution. Electronic Journal of Probability, 22 (??):37:1–37:94, ???? 2017. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1493258436.

Schweinsberg:2017:RRPb

[Sch17b] Jason Schweinsberg. Rigorous results for a population model with selection II: genealogy of the population. Electronic Journal of Probability, 22(??): 38:1–38:54, ???? 2017. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1493258437.

Schiavo:2019:CFD

[Sch19a] Lorenzo Dello Schiavo. Characteristic functionals of Dirichlet measures. Electronic Journal of Probability, 24(??):115:1–115:38, ???? 2019. CODEN

???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.ejp/
1570759240.

Schmid:2019:MTS

[Sch19b] Dominik Schmid. Mixing times for the simple exclusion process in ballistic random environment. Electronic Journal of Probability, 24(??):22:1–22:25, ???? 2019. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

euclid.ejp/1553155302.

Saloff-Coste:2022:RWF

[SCW22] Laurent Saloff-Coste and Yuwen Wang. Random walks on finite nilpotent groups driven by long-jump measures. Electronic Journal of Probability, 27(??):1–31, 2022. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

journals/electronic-journal-of-probability/volume-27/issue-none/Random-
walks-on-finite-nilpotent-groups-driven-by-long-jump/10.1214/22-EJP745.
full.

Saloff-Coste:2009:MTI

[SCZ09] Laurent Saloff-Coste and Jessica Zuniga. Merging for time inhomogeneous finite Markov chains, Part I: Singular values and stability. Electronic Journal of Probability, 14:49:1456–49:1494, 2009. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/656.

Saloff-Coste:2013:LDS

[SCZ13] Laurent Saloff-Coste and Tianyi Zheng. Large deviations for stable like random walks on Z^{d} with applications to random walks on wreath products. Electronic Journal of Probability, 18:93:1–93:35, 2013. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/2439.

Sabot:2006:RWD

[SE06] Christophe Sabot and Nathanaël Enriquez. Random walks in a Dirichlet environment. Electronic Journal of Probability, 11:31:802–31:816, 2006. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/350.

Sega:2009:LRC

[Seg09] Gregor Sega. Large-range constant threshold growth model in one dimension. Electronic Journal of Probability, 14:6:119–6:138, 2009. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/598.

Seidler:2010:EES

[Sei10] Jan Seidler. Exponential estimates for stochastic convolutions in 2-smooth Banach spaces. Electronic Journal of Probability, 15:50:1556–50:1573, 2010. CO-

DEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/

Sellke:2006:RRR

[Sel06] Thomas Sellke. Recurrence of reinforced random walk on a ladder. Electronic Journal of Probability, 11:11:301–11:310, 2006. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/313.

Senizergues:2021:GWR

[Sén21] Delphin Sénizergues. Geometry of weighted recursive and affine preferential attachment trees. Electronic Journal of Probability, 26(18):1–56, February 2021.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-26/issue-none/Geometry-of-
weighted-recursive-and-affine-preferential-attachment-trees/10.1214/
21-EJP640.full.

Seppalainen:1996:MMB

[Sep96] Timo Seppäläinen. A microscopic model for the Burgers equation and longest increasing subsequences. Electronic Journal of Probability, 1:5:1–5:51, 1996.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
5.

Swart:2006:RAC

[SF06] Jan Swart and Klaus Fleischmann. Renormalization analysis of catalytic Wright–Fisher diffusions. Electronic Journal of Probability, 11:24:585–24:654, 2006. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/

view/341.

Sharpe:1999:MRS

[Sha99] Michael Sharpe. Martingales on random sets and the strong martingale property. Electronic Journal of Probability, 5:1:1–1:17, 1999. CODEN ???? ISSN 1083-

6489. URL http://ejp.ejpecp.org/article/view/57.

Shao:2011:NPM

[Sha11] Jinghai Shao. A new probability measure-valued stochastic process with Ferguson–Dirichlet process as reversible measure. Electronic Journal of Probability, 16:9:271–9:292, 2011. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/844.

Shao:2015:CTR

[Sha15] Jinghai Shao. Criteria for transience and recurrence of regime-switching diffusion processes. Electronic Journal of Probability, 20(??):63:1–63:15, ???? 2015.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/

Shcherbina:2022:STM

[Shc22] Tatyana Shcherbina. SUSY transfer matrix approach for the real symmetric 1d random band matrices. Electronic Journal of Probability, 27(??):1–29, 2022.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-27/issue-none/SUSY-transfer-
matrix-approach-for-the-real-symmetric-1d-random/10.1214/22-EJP747.
full.

Shellef:2010:ISP

[She10] Eric Shellef. IDLA on the supercritical percolation cluster. Electronic Journal of Probability, 15:24:723–24:740, 2010. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/775.

Shiraishi:2012:TSR

[Shi12] Daisuke Shiraishi. Two-sided random walks conditioned to have no intersections. Electronic Journal of Probability, 17:18:1–18:24, 2012. CODEN ???? ISSN 1083-

6489. URL http://ejp.ejpecp.org/article/view/1742.

Shi:2017:GFP

[Shi17] Quan Shi. Growth-fragmentation processes and bifurcators. Electronic Journal of Probability, 22(??):15:1–15:25, ???? 2017. CODEN ???? ISSN 1083-6489.

URL https://projecteuclid.org/euclid.ejp/1487127643.

Shkolnikov:2009:CPS

[Shk09] Mykhaylo Shkolnikov. Competing particle systems evolving by I.I.D. increments. Electronic Journal of Probability, 14:27:728–27:751, 2009. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/635.

Shkolnikov:2013:SUE

[Shk13] Mykhaylo Shkolnikov. Some universal estimates for reversible Markov chains. Electronic Journal of Probability, 18:11:1–11:17, 2013. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/1749.

Shen:2020:EID

[SHW20] Yandi Shen, Fang Han, and Daniela Witten. Exponential inequalities for dependent V -statistics via random Fourier features. Electronic Journal of Probability, 25(??):7:1–7:18, ???? 2020. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1580267007.

Simon:2014:CFP

[Sim14] Thomas Simon. Comparing Fréchet and positive stable laws. Electronic Journal of Probability, 19:16:1–16:25, 2014. CODEN ???? ISSN 1083-6489. URL http:

//ejp.ejpecp.org/article/view/3058.

Slominski:2014:WZT

[Slo14] Leszek Slominski. On Wong–Zakai type approximations of reflected diffusions. Electronic Journal of Probability, 19(??):118:1–118:15, ???? 2014. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/3425.

Sulzbach:2014:GLP

[SND14] Henning Sulzbach, Ralph Neininger, and Michael Drmota. A Gaussian limit process for optimal FIND algorithms. Electronic Journal of Probability, 19: 3:1–3:28, 2014. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.

org/article/view/2933.

Sobieczky:2012:BAR

[Sob12] Florian Sobieczky. Bounds for the annealed return probability on large finite percolation graphs. Electronic Journal of Probability, 17:79:1–79:17, 2012. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/2329.

Song:2006:TSE

[Son06] Renming Song. Two-sided estimates on the density of the Feynman–Kac semigroups of stable-like processes. Electronic Journal of Probability, 11:6:146–6:161, 2006. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/

view/308.

Song:2012:URM

[Son12] Yongsheng Song. Uniqueness of the representation for G-martingales with finite variation. Electronic Journal of Probability, 17:24:1–24:15, 2012. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/1890.

Spinka:2014:RWL

[SP14] Yinon Spinka and Ron Peled. Random walk with long-range constraints. Electronic Journal of Probability, 19:52:1–52:54, 2014. CODEN ???? ISSN 1083-

6489. URL http://ejp.ejpecp.org/article/view/3060.

Sun:2019:ALU

[SP19] Wen Sun and Robert Philippe. Analysis of large urn models with local meanfield interactions. Electronic Journal of Probability, 24(??):45:1–45:33, ???? 2019. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

euclid.ejp/1557453644.

Spiliopoulos:2009:WPR

[Spi09] Konstantinos Spiliopoulos. Wiener process with reflection in non-smooth narrow tubes. Electronic Journal of Probability, 14:69:2011–69:2037, 2009. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/691.

Spiliopoulos:2015:QLD

[Spi15] Konstantinos Spiliopoulos. Quenched large deviations for multiscale diffusion processes in random environments. Electronic Journal of Probability, 20(??): 15:1–15:29, ???? 2015. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/3729.

Spinka:2020:FCS

[Spi20] Yinon Spinka. Finitary coding for the sub-critical Ising model with finite expected coding volume. Electronic Journal of Probability, 25(??):8:1–8:27, ???? 2020. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

euclid.ejp/1580267008.

Spiro:2022:OCG

[Spi22] Sam Spiro. Online card games. Electronic Journal of Probability, 27(??):1–15, 2022. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

journals/electronic-journal-of-probability/volume-27/issue-none/Online-
card-games/10.1214/22-EJP768.full.

Schuett:2013:ENR

[SR13] Carsten Schuett and Stiene Riemer. On the expectation of the norm of random matrices with non-identically distributed entries. Electronic Journal of Probability, 18:29:1–29:13, 2013. CODEN ???? ISSN 1083-6489. URL http://

ejp.ejpecp.org/article/view/2103.

Sengupta:2000:FPD

[SS00] Arindam Sengupta and Anish Sarkar. Finitely polynomially determined Lévy processes. Electronic Journal of Probability, 6:7:1–7:22, 2000. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/80.

Schilling:2010:SAS

[SS10] Rene Schilling and Alexander Schnurr. The symbol associated with the solution of a stochastic differential equation. Electronic Journal of Probability, 15: 43:1369–43:1393, 2010. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/807.

Samorodnitsky:2012:DSL

[SS12] Gennady Samorodnitsky and Yi Shen. Distribution of the supremum location of stationary processes. Electronic Journal of Probability, 17:42:1–42:17, 2012.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/

Sarkar:2013:BWS

[SS13] Anish Sarkar and Rongfeng Sun. Brownian web in the scaling limit of supercritical oriented percolation in dimension 1 + 1. Electronic Journal of Probability, 18:21:1–21:23, 2013. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.

org/article/view/2019.

Sturm:2014:SCP

[SS14] Anja Sturm and Jan Swart. Subcritical contact processes seen from a typical infected site. Electronic Journal of Probability, 19:53:1–53:46, 2014. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/2904.

Spohn:2015:PIB

[SS15] Herbert Spohn and Tomohiro Sasamoto. Point-interacting Brownian motions in the KPZ universality class. Electronic Journal of Probability, 20(??):87:1–87:28, ???? 2015. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/3926.

Schertzer:2018:HCP

[SS18a] Emmanuel Schertzer and Florian Simatos. Height and contour processes of Crump-Mode-Jagers forests (I): general distribution and scaling limits in the case of short edges. Electronic Journal of Probability, 23(??):67:1–67:43, ???? 2018. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

euclid.ejp/1532570595.

Sethuraman:2018:HLL

[SS18b] Sunder Sethuraman and Doron Shahar. Hydrodynamic limits for long-range asymmetric interacting particle systems. Electronic Journal of Probability, 23 (??):130:1–130:54, ???? 2018. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1545361594.

Sakai:2019:SMH

[SS19a] Akira Sakai and Gordon Slade. Spatial moments for high-dimensional critical contact process, oriented percolation and lattice trees. Electronic Journal of Probability, 24(??):65:1–65:18, ???? 2019. CODEN ???? ISSN 1083-6489. URL

https://projecteuclid.org/euclid.ejp/1561169149.

Schertzer:2019:HCP

[SS19b] Emmanuel Schertzer and Florian Simatos. Height and contour processes of Crump–Mode–Jagers forests (II): the Bellman–Harris universality class. Electronic Journal of Probability, 24(??):47:1–47:38, ???? 2019. CODEN ???? ISSN

1083-6489. URL https://projecteuclid.org/euclid.ejp/1558145015.

Seppalainen:2020:CEC

[SS20] Timo Seppäläinen and Xiao Shen. Coalescence estimates for the corner growth model with exponential weights. Electronic Journal of Probability, 25(??):1–31, 2020. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

journals/electronic-journal-of-probability/volume-25/issue-none/Coalescence-
estimates-for-the-corner-growth-model-with-exponential-weights/10.
1214/20-EJP489.full. See erratum [SS21].

Seppalainen:2021:ECE

[SS21] Timo Seppäläinen and Xiao Shen. Erratum to: Coalescence estimates for the corner growth model with exponential weights. Electronic Journal of Probability, 26(18):1–4, February 2021. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-26/issue-none/Erratum-to--Coalescence-estimates-
for-the-corner-growth-model/10.1214/21-EJP714.full. See [SS20].

Sapatinas:2011:SNA

[SSG11] Theofanis Sapatinas, Damodar Shanbhag, and Arjun Gupta. Some new approaches to infinite divisibility. Electronic Journal of Probability, 16:85:2359– 85:2374, 2011. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/961.

Shen:2022:TFI

[SSH22] Jinqi Shen, Stilian Stoev, and Tailen Hsing. Tangent fields, intrinsic stationarity, and self similarity. Electronic Journal of Probability, 27(??):1–56, 2022.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-27/issue-none/Tangent-fields-
intrinsic-stationarity-and-self-similarity/10.1214/22-EJP754.full.

Schertzer:2009:SPB

[SSS09] Emmanuel Schertzer, Rongfeng Sun, and Jan Swart. Special points of the Brownian net. Electronic Journal of Probability, 14:30:805–30:864, 2009. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/641.

Sanz-Sole:2013:SWE

[SSS13] Marta Sanz-Solé and André Süss. The stochastic wave equation in high dimensions: Malliavin differentiability and absolute continuity. Electronic Journal of Probability, 18:64:1–64:28, 2013. CODEN ???? ISSN 1083-6489. URL http://

ejp.ejpecp.org/article/view/2341.

Shao:2004:ADB

[SSW04] Qi-Man Shao, Chun Su, and Gang Wei. Asymptotic distributions and Berry– Esséen bounds for sums of record values. Electronic Journal of Probability, 9:17:544–17:559, 2004. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/210.

Shiga:2006:IDR

[ST06] Tokuzo Shiga and Hiroshi Tanaka. Infinitely divisible random probability distributions with an application to a random motion in a random environment. Electronic Journal of Probability, 11:44:1144–44:1183, 2006. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/380.

Sarantsev:2017:SGD

[ST17a] Andrey Sarantsev and Li-Cheng Tsai. Stationary gap distributions for infinite systems of competing Brownian particles. Electronic Journal of Probability, 22 (??):56:1–56:20, ???? 2017. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1499220068.

Sidoravicius:2017:AST

[ST17b] Vladas Sidoravicius and Augusto Teixeira. Absorbing-state transition for Stochastic Sandpiles and Activated Random Walks. Electronic Journal of Probability, 22(??):33:1–33:35, ???? 2017. CODEN ???? ISSN 1083-6489. URL

https://projecteuclid.org/euclid.ejp/1492070448.

Soloveychik:2022:LDC

[ST22] Ilya Soloveychik and Vahid Tarokh. Large deviations of convex polyominoes. Electronic Journal of Probability, 27(??):1–19, 2022. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-27/issue-none/Large-deviations-of-convex-polyominoes/
10.1214/22-EJP835.full.

Steinsaltz:1999:RTC

[Ste99] David Steinsaltz. Random time changes for sock-sorting and other stochastic process limit theorems. Electronic Journal of Probability, 4:14:1–14:25, 1999.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
51.

Stephenson:2013:GFT

[Ste13] Robin Stephenson. General fragmentation trees. Electronic Journal of Probability, 18:101:1–101:45, 2013. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/2703.

Steiner:2021:FKA

[Ste21] Clément Steiner. A Feynman–Kac approach for logarithmic Sobolev inequalities. Electronic Journal of Probability, 26(18):1–19, February 2021. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/

electronic-journal-of-probability/volume-26/issue-none/A-Feynman-
-Kac-approach-for-logarithmic-Sobolev-inequalities/10.1214/21-EJP656.
full.

Strzelecka:2019:ENL

[Str19] Marta Strzelecka. Estimates of norms of log-concave random matrices with dependent entries. Electronic Journal of Probability, 24(??):107:1–107:15, ???? 2019. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

euclid.ejp/1569981822.

Sturm:2003:CPP

[Stu03] Anja Sturm. On convergence of population processes in random environments to the stochastic heat equation with colored noise. Electronic Journal of Probability, 8:6:1–6:39, 2003. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/129.

Stufler:2020:MOS

[Stu20] Benedikt Stufler. On the maximal offspring in a subcritical branching process. Electronic Journal of Probability, 25(??):1–62, 2020. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-25/issue-none/On-the-maximal-offspring-in-a-
subcritical-branching-process/10.1214/20-EJP506.full.

Scoppola:2022:SDC

[STV22] Benedetto Scoppola, Alessio Troiani, and Matteo Veglianti. Shaken dynamics on the 3d cubic lattice. Electronic Journal of Probability, 27(??):1–26, 2022.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-27/issue-none/Shaken-dynamics-
on-the-3d-cubic-lattice/10.1214/22-EJP803.full.

Soner:2011:QSS

[STZ11] Mete Soner, Nizar Touzi, and Jianfeng Zhang. Quasi-sure stochastic analysis through aggregation. Electronic Journal of Probability, 16:67:1844–67:1879, 2011. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/

view/950.

Su:2001:DCD

[Su01] Francis Su. Discrepancy convergence for the drunkard’s walk on the sphere. Electronic Journal of Probability, 6:2:1–2:20, 2001. CODEN ???? ISSN 1083-

6489. URL http://ejp.ejpecp.org/article/view/75.

Su:2006:GFC

[Su06] Zhonggen Su. Gaussian fluctuations in complex sample covariance matrices. Electronic Journal of Probability, 11:48:1284–48:1320, 2006. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/378.

Su:2014:BRW

[Su14] Wei Su. Branching random walks and contact processes on Galton–Watson trees. Electronic Journal of Probability, 19:41:1–41:12, 2014. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/3118.

Su:2020:PSH

[Su20] Weicong Su. On the peaks of a stochastic heat equation on a sphere with a large radius. Electronic Journal of Probability, 25(??):5:1–5:38, ???? 2020.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.
ejp/1579835021.

Subag:2013:LBM

[Sub13] Eliran Subag. A lower bound for the mixing time of the random-to-random insertions shuffle. Electronic Journal of Probability, 18:20:1–20:20, 2013. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/1950.

Suzuki:2019:CBM

[Suz19] Kohei Suzuki. Convergence of Brownian motions on metric measure spaces under Riemannian curvature–dimension conditions. Electronic Journal of Probability, 24(??):102:1–102:36, ???? 2019. CODEN ???? ISSN 1083-6489. URL

https://projecteuclid.org/euclid.ejp/1569463328.

Sethuraman:2005:MPD

[SV05] Sunder Sethuraman and Srinivasa Varadhan. A martingale proof of Dobrushin’s theorem for non-homogeneous Markov chains. Electronic Journal of Probability, 10:36:1221–36:1235, 2005. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/283.

Salminen:2009:SLM

[SV09] Paavo Salminen and Pierre Vallois. On subexponentiality of the Lévy measure of the inverse local time; with applications to penalizations. Electronic Journal of Probability, 14:67:1963–67:1991, 2009. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/686.

Stadje:2011:TKG

[SW11] Wolfgang Stadje and Achim Wübker. Three kinds of geometric convergence for Markov chains and the spectral gap property. Electronic Journal of Probability, 16:34:1001–34:1019, 2011. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/900.

Sosoe:2014:CED

[SW14] Philippe Sosoe and Percy Wong. Convergence of the eigenvalue density for Laguerre beta ensembles on short scales. Electronic Journal of Probability, 19: 34:1–34:18, 2014. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.

org/article/view/2638.

Shi:2019:PTC

[SW19] Quan Shi and Alexander R. Watson. Probability tilting of compensated fragmentations. Electronic Journal of Probability, 24(??):78:1–78:39, ???? 2019.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.
ejp/1565057003.

Shao:2012:HIS

[SWY12] Jinghai Shao, Feng-Yu Wang, and Chenggui Yuan. Harnack inequalities for stochastic (functional) differential equations with non-Lipschitzian coefficients. Electronic Journal of Probability, 17:100:1–100:18, 2012. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/2140.

Safikhani:2019:SCE

[SX19] Abolfazl Safikhani and Yimin Xiao. Spectral conditions for equivalence of Gaussian random fields with stationary increments. Electronic Journal of Probability, 24(??):8:1–8:19, ???? 2019. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1550199786.

Salminen:2005:PIF

[SY05] Paavo Salminen and Marc Yor. Perpetual integral functionals as hitting and occupation times. Electronic Journal of Probability, 10:11:371–11:419, 2005.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
256.

Sheffield:2014:TPR

[SY14] Scott Sheffield and Ariel Yadin. Tricolor percolation and random paths in 3D. Electronic Journal of Probability, 19:4:1–4:23, 2014. CODEN ???? ISSN 1083-

6489. URL http://ejp.ejpecp.org/article/view/3073.

Seuret:2017:MAO

[SY17] Stéphane Seuret and Xiaochuan Yang. Multifractal analysis for the occupation measure of stable-like processes. Electronic Journal of Probability, 22 (??):47:1–47:36, ???? 2017. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1496109646.

Schulte:2019:MSO

[SY19] Matthias Schulte and J. E. Yukich. Multivariate second order Poincaré inequalities for Poisson functionals. Electronic Journal of Probability, 24(??): 130:1–130:42, ???? 2019. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1573527859.

Song:2021:HDC

[SYY21] Jian Song, Jianfeng Yao, and Wangjun Yuan. High-dimensional central limit theorems for a class of particle systems. Electronic Journal of Probability, 26(18):1–33, February 2021. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-26/issue-none/High-dimensional-central-limit-
theorems-for-a-class-of-particle/10.1214/21-EJP646.full.

Song:2015:RDS

[SZ15a] Yulin Song and Xicheng Zhang. Regularity of density for SDEs driven by degenerate Lévy noises. Electronic Journal of Probability, 20(??):21:1–21:27, ???? 2015. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/

view/3287.

Storm:2015:OLR

[SZ15b] Julia Storm and Dirk Zeindler. The order of large random permutations with cycle weights. Electronic Journal of Probability, 20(??):126:1–126:34, ???? 2015.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
4331.

Shang:2019:TCI

[SZ19] Shijie Shang and Tusheng Zhang. Talagrand concentration inequalities for stochastic heat-type equations under uniform distance. Electronic Journal of Probability, 24(??):129:1–129:15, ???? 2019. CODEN ???? ISSN 1083-6489.

URL https://projecteuclid.org/euclid.ejp/1573527858.

Shen:2021:DTS

[SZ21] Yi Shen and Zhenyuan Zhang. On discrete-time self-similar processes with stationary increments. Electronic Journal of Probability, 26(18):1–24, February 2021. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

journals/electronic-journal-of-probability/volume-26/issue-none/On-
discrete-time-self-similar-processes-with-stationary-increments/10.
1214/21-EJP689.full.

Szablowski:2010:MNR

[Sza10] Pawel Szablowski. Multidimensional q-normal and related distributions — Markov case. Electronic Journal of Probability, 15:40:1296–40:1318, 2010. CO-

DEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
796.

Sznitman:2009:DRW

[Szn09] Alain-Sol Sznitman. On the domination of a random walk on a discrete cylinder by random interlacements. Electronic Journal of Probability, 14:56:1670– 56:1704, 2009. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/679.

Sznitman:2016:CAL

[Szn16] Alain-Sol Sznitman. Coupling and an application to level-set percolation of the Gaussian free field. Electronic Journal of Probability, 21(??):35:1–35:26, ???? 2016. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

euclid.ejp/1461332876.

Taggi:2016:ASP

[Tag16] Lorenzo Taggi. Absorbing-state phase transition in biased activated random walk. Electronic Journal of Probability, 21(??):13:1–13:15, ???? 2016. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.ejp/
1456246244.

Takacs:1997:RWP

[Tak97] Christiane Takacs. Random walk on periodic trees. Electronic Journal of Probability, 2:1:1–1:16, 1997. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/15.

Takei:2021:ASB

[Tak21] Masato Takei. Almost sure behavior of linearly edge-reinforced random walks on the half-line. Electronic Journal of Probability, 26(18):1–18, February 2021.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-26/issue-none/Almost-sure-
behavior-of-linearly-edge-reinforced-random-walks-on/10.1214/21-EJP674.
full.

Tang:2006:IND

[Tan06] Qihe Tang. Insensitivity to negative dependence of the asymptotic behavior of precise large deviations. Electronic Journal of Probability, 11:4:107–4:120, 2006.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
304.

Tan:2013:SMF

[Tan13] Xiaolu Tan. A splitting method for fully nonlinear degenerate parabolic PDEs. Electronic Journal of Probability, 18:15:1–15:24, 2013. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/1967.

Tanguy:2019:NAV

[Tan19] Kevin Tanguy. Non asymptotic variance bounds and deviation inequalities by optimal transport. Electronic Journal of Probability, 24(??):13:1–13:18, ???? 2019. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

euclid.ejp/1550653272.

Tang:2021:WUS

[Tan21] Pengfei Tang. Weights of uniform spanning forests on nonunimodular transitive graphs. Electronic Journal of Probability, 26(18):1–62, February 2021.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-26/issue-none/Weights-of-
uniform-spanning-forests-on-nonunimodular-transitive-graphs/10.1214/
21-EJP709.full.

Tang:2022:RPN

[Tan22] Pengfei Tang. Return probabilities on nonunimodular transitive graphs. Electronic Journal of Probability, 27(??):1–27, 2022. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-27/issue-none/Return-probabilities-on-nonunimodular-
transitive-graphs/10.1214/22-EJP859.full.

Tang:2023:NSC

[Tan23] Pengfei Tang. A note on some critical thresholds of Bernoulli percolation. Electronic Journal of Probability, 28(??):1–22, 2023. CODEN

????  ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-28/issue-none/A-note-on-
some-critical-thresholds-of-Bernoulli-percolation/10.1214/23-EJP926.
full.

Tarrago:2015:AIL

[Tar15] Pierre Tarrago. Asymptotic independence in large random permutations with fixed descent set. Electronic Journal of Probability, 20(??):103:1–103:33, ???? 2015. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/

view/4196.

Taylor:2007:CAP

[Tay07] Jesse Taylor. The common ancestor process for a Wright–Fisher diffusion. Electronic Journal of Probability, 12:28:808–28:847, 2007. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/418.

Tropp:2014:SMP

[TC14] Joel Tropp and Richard Chen. Subadditivity of matrix phi-entropy and concentration of random matrices. Electronic Journal of Probability, 19:27:1–27:30, 2014. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/

view/2964.

Teixeira:2009:IPT

[Tei09] Augusto Teixeira. Interlacement percolation on transient weighted graphs. Electronic Journal of Probability, 14:54:1604–54:1627, 2009. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/670.

Telcs:2001:LSG

[Tel01] Andras Telcs. Local sub-Gaussian estimates on graphs: The strongly recurrent case. Electronic Journal of Probability, 6:22:1–22:33, 2001. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/95.

Thevenin:2020:VFO

[Thé20] Paul Thévenin. Vertices with fixed outdegrees in large Galton–Watson trees. Electronic Journal of Probability, 25(??):64:1–64:25, ???? 2020. CODEN

???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.ejp/
1592445678.

Tough:2022:FVP

[TN22] Oliver Tough and James Nolen. The Fleming–Viot process with McKean–Vlasov dynamics. Electronic Journal of Probability, 27(??):1–72, 2022. CODEN ????

ISSN 1083-6489. URL https://projecteuclid.org/journals/electronic-
journal-of-probability/volume-27/issue-none/The-Fleming-Viot-process-
with-McKean--Vlasov-dynamics/10.1214/22-EJP820.full.

Toninelli:2007:CLR

[Ton07] Fabio Lucio Toninelli. Correlation lengths for random polymer models and for some renewal sequences. Electronic Journal of Probability, 12:21:613–21:636, 2007. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/

view/414.

Toninelli:2009:CGF

[Ton09] Fabio Toninelli. Coarse graining, fractional moments and the critical slope of random copolymers. Electronic Journal of Probability, 14:20:531–20:547, 2009.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
612.

Trutnau:2015:CSB

[TOR15] Gerald Trutnau, Youssef Ouknine, and Francesco Russo. On countably skewed Brownian motion with accumulation point. Electronic Journal of Probability, 20(??):82:1–82:27, ???? 2015. CODEN ???? ISSN 1083-6489. URL http://

ejp.ejpecp.org/article/view/3640.

Tournier:2009:IET

[Tou09] Laurent Tournier. Integrability of exit times and ballisticity for random walks in Dirichlet environment. Electronic Journal of Probability, 14:16:431–16:451, 2009. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/

view/609.

Towsner:2015:LSM

[Tow15] Henry Piers Towsner. Limits of sequences of Markov chains. Electronic Journal of Probability, 20(??):77:1–77:23, ???? 2015. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/4188.

Travers:2015:ILI

[Tra15] Nicholas Travers. Inversions and longest increasing subsequence for k-card-minimum random permutations. Electronic Journal of Probability, 20(??):11:1– 11:27, ???? 2015. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.

org/article/view/3602.

Travers:2018:ERW

[Tra18] Nicholas F. Travers. Excited random walk in a Markovian environment. Electronic Journal of Probability, 23(??):43:1–43:60, ???? 2018. CODEN ???? ISSN

1083-6489. URL https://projecteuclid.org/euclid.ejp/1525852820.

Trevisan:2016:WPM

[Tre16] Dario Trevisan. Well-posedness of multidimensional diffusion processes with weakly differentiable coefficients. Electronic Journal of Probability, 21(??): 22:1–22:41, ???? 2016. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1458325000.

Tsirelson:2006:BLM

[Tsi06] Boris Tsirelson. Brownian local minima, random dense countable sets and random equivalence classes. Electronic Journal of Probability, 11:7:162–7:198, 2006.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
309.

Talarczyk:2020:LTI

[TT20] Anna Talarczyk and Łukasz Treszczotko. Limit theorems for integrated trawl processes with symmetric Lévy bases. Electronic Journal of Probability, 25(??): 1–24, 2020. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.

org/journals/electronic-journal-of-probability/volume-25/issue-none/
Limit-theorems-for-integrated-trawl-processes-with-symmetric-L%c3%
a9vy-bases/10.1214/20-EJP509.full.

Tanaka:2022:GED

[TT22] Ryokichi Tanaka and Kenkichi Tsunoda. Glauber-exclusion dynamics: rapid mixing regime. Electronic Journal of Probability, 27(??):1–26, 2022. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-27/issue-none/Glauber-Exclusion-
dynamics-rapid-mixing-regime/10.1214/22-EJP865.full.

Torres:2014:QVF

[TTV14] Soledad Torres, Ciprian Tudor, and Frederi Viens. Quadratic variations for the fractional-colored stochastic heat equation. Electronic Journal of Probability, 19: 76:1–76:51, 2014. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.

org/article/view/2698.

Tugaut:2012:EPM

[Tug12] Julian Tugaut. Exit problem of McKean–Vlasov diffusions in convex landscapes. Electronic Journal of Probability, 17:76:1–76:26, 2012. CODEN ???? ISSN 1083-

6489. URL http://ejp.ejpecp.org/article/view/1914.

Turova:2007:CPT

[Tur07] Tatyana Turova. Continuity of the percolation threshold in randomly grown graphs. Electronic Journal of Probability, 12:36:1036–36:1047, 2007. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/436.

Turkedjiev:2015:TAD

[Tur15] Plamen Turkedjiev. Two algorithms for the discrete time approximation of Markovian backward stochastic differential equations under local conditions. Electronic Journal of Probability, 20(??):50:1–50:49, ???? 2015. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/3022.

Tudor:2003:IFL

[TV03] Ciprian Tudor and Frederi Viens. Itô formula and local time for the fractional Brownian sheet. Electronic Journal of Probability, 8:14:1–14:31, 2003. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/155.

Taylor:2009:CPS

[TV09] Jesse Taylor and Amandine Véber. Coalescent processes in subdivided populations subject to recurrent mass extinctions. Electronic Journal of Probability, 14: 9:242–9:288, 2009. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.

org/article/view/595.

Tao:2011:WDM

[TV11] Terence Tao and Van Vu. The Wigner–Dyson–Mehta bulk universality conjecture for Wigner matrices. Electronic Journal of Probability, 16:77:2104–77:2121, 2011. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/

view/944.

Tribe:2011:SOM

[TW11] Roger Tribe and Nicholas Woodward. Stochastic order methods applied to stochastic travelling waves. Electronic Journal of Probability, 16:16:436–16:469, 2011. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/

view/868.

Tran:2020:STS

[TY20] Huy Tran and Yizheng Yuan. A support theorem for SLE curves. Electronic Journal of Probability, 25(??):18:1–18:18, ???? 2020. CODEN ???? ISSN

1083-6489. URL https://projecteuclid.org/euclid.ejp/1580958251.

Takeda:2023:LTP

[TY23] Shosei Takeda and Kouji Yano. Local time penalizations with various clocks for Lévy processes. Electronic Journal of Probability, 28(??):1–35, 2023. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-28/issue-none/Local-time-
penalizations-with-various-clocks-for-L%c3%a9vy-processes/10.1214/
23-EJP903.full.

Tykesson:2007:NUC

[Tyk07] Johan Tykesson. The number of unbounded components in the Poisson Boolean model of continuum percolation in hyperbolic space. Electronic Journal of Probability, 12:51:1379–51:1401, 2007. CODEN ???? ISSN 1083-6489. URL http:/

/ejp.ejpecp.org/article/view/460.

Tribe:2011:PFO

[TZ11] Roger Tribe and Oleg Zaboronski. Pfaffian formulae for one dimensional coalescing and annihilating systems. Electronic Journal of Probability, 16:76:2080– 76:2103, 2011. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/942.

Uchiyama:2007:AEG

[Uch07] Kouhei Uchiyama. Asymptotic estimates of the Green functions and transition probabilities for Markov additive processes. Electronic Journal of Probability, 12:6:138–6:180, 2007. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/396.

Uchiyama:2010:GFT

[Uch10] Kohei Uchiyama. The green functions of two dimensional random walks killed on a line and their higher dimensional analogues. Electronic Journal of Probability, 15:37:1161–37:1189, 2010. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/793.

Uchiyama:2011:FHT

[Uch11] Kohei Uchiyama. The first hitting time of a single point for random walks. Electronic Journal of Probability, 16:71:1960–71:2000, 2011. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/931.

Uchiyama:2020:PFL

[Uch20] Kohei Uchiyama. The potential function and ladder heights of a recurrent random walk on Z with infinite variance. Electronic Journal of Probability, 25(??):1–24, 2020. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-25/issue-none/The-potential-function-and-ladder-
heights-of-a-recurrent-random/10.1214/20-EJP553.full.

Unterberger:2019:GFL

[Unt19] Jeremie Unterberger. Global fluctuations for 1D log-gas dynamics. covariance kernel and support. Electronic Journal of Probability, 24(??):21:1–21:28, ???? 2019. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

euclid.ejp/1553155301.

Valle:2007:EIT

[Val07] Glauco Valle. Evolution of the interfaces in a two dimensional Potts model. Electronic Journal of Probability, 12:12:354–12:386, 2007. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/346.

Valesin:2010:MCP

[Val10] Daniel Valesin. Multitype contact process on z: Extinction and interface. Electronic Journal of Probability, 15:73:2220–73:2260, 2010. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/836.

Vanneuville:2019:ASR

[Van19] Hugo Vanneuville. Annealed scaling relations for Voronoi percolation. Electronic Journal of Probability, 24(??):39:1–39:71, ???? 2019. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/euclid.ejp/1554861841.

Varvenne:2019:CIS

[Var19] Maylis Varvenne. Concentration inequalities for stochastic differential equations with additive fractional noise. Electronic Journal of Probability, 24(??): 124:1–124:22, ???? 2019. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1573268588.

vanBargen:2009:AGS

[vB09] Holger van Bargen. Asymptotic growth of spatial derivatives of isotropic flows. Electronic Journal of Probability, 14:80:2328–80:2351, 2009. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/704.

vandeBrug:2018:SSL

[vdBCL18] Tim van de Brug, Federico Camia, and Marcin Lis. Spin systems from loop soups. Electronic Journal of Probability, 23(??):81:1–81:17, ???? 2018. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.ejp/
1536717740.

vanderHofstad:2007:DRG

[vdHHZ07] Remco van der Hofstad, Gerard Hooghiemstra, and Dmitri Znamenski. Distances in random graphs with finite mean and infinite variance degrees. Electronic Journal of Probability, 12:25:703–25:766, 2007. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/420.

vanderHofstad:2004:GSC

[vdHS04] Remco van der Hofstad and Akira Sakai. Gaussian scaling for the critical spreadout contact process above the upper critical dimension. Electronic Journal of Probability, 9:24:710–24:769, 2004. CODEN ???? ISSN 1083-6489. URL http:

//ejp.ejpecp.org/article/view/224.

vanderHofstad:2010:CCF

[vdHS10] Remco van der Hofstad and Akira Sakai. Convergence of the critical finite-range contact process to super-Brownian motion above the upper critical dimension: The higher-point functions. Electronic Journal of Probability, 15:27:801–27:894, 2010. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/

view/783.

Velicu:2022:LSI

[Vel22] Andrei Velicu. Logarithmic Sobolev inequalities for Dunkl operators with applications to functional inequalities for singular Boltzmann–Gibbs measures. Electronic Journal of Probability, 27(??):1–25, 2022. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-27/issue-none/Logarithmic-Sobolev-inequalities-
for-Dunkl-operators-with-applications-to-functional/10.1214/22-EJP810.
full.

Venet:2019:NFB

[Ven19] Nil Venet. Nonexistence of fractional Brownian fields indexed by cylinders. Electronic Journal of Probability, 24(??):75:1–75:26, ???? 2019. CODEN ???? ISSN

1083-6489. URL https://projecteuclid.org/euclid.ejp/1562119475.

Versendaal:2019:LDG

[Ver19] Rik Versendaal. Large deviations for geodesic random walks. Electronic Journal of Probability, 24(??):93:1–93:39, ???? 2019. CODEN ???? ISSN 1083-6489.

URL https://projecteuclid.org/euclid.ejp/1568361634.

Veto:2015:TWL

[Vet15] Bálint Vető. Tracy–Widom limit of q-Hahn TASEP. Electronic Journal of Probability, 20(??):102:1–102:22, ???? 2015. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/4241.

Valle:2015:SLR

[VFV15] Glauco Valle, Luiz Renato Fontes, and Leon Alexander Valencia. Scaling limit of the radial Poissonian web. Electronic Journal of Probability, 20(??):31:1–31:40, ???? 2015. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/3395.

Vihola:2011:CAM

[Vih11] Matti Vihola. Can the adaptive Metropolis algorithm collapse without the covariance lower bound? Electronic Journal of Probability, 16:2:45–2:75, 2011.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
840.

Villemonais:2011:IPS

[Vil11] Denis Villemonais. Interacting particle systems and Yaglom limit approximation of diffusions with unbounded drift. Electronic Journal of Probability, 16:61:1663– 61:1692, 2011. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/925.

Villemonais:2015:MQS

[Vil15] Denis Villemonais. Minimal quasi-stationary distribution approximation for a birth and death process. Electronic Journal of Probability, 20(??):30:1–30:18, ???? 2015. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/3482.

Volkov:2003:ERW

[Vol03] Stanislav Volkov. Excited random walk on trees. Electronic Journal of Probability, 8:23:1–23:15, 2003. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/180.

Vollering:2014:VIG

[Völ14] Florian Völlering. A variance inequality for Glauber dynamics applicable to high and low temperature regimes. Electronic Journal of Probability, 19:46:1–46:21, 2014. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/

view/2791.

Voss:2008:LDO

[Vos08] Jochen Voss. Large deviations for one dimensional diffusions with a strong drift. Electronic Journal of Probability, 13:53:1479–53:1528, 2008. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/564.

vonRenesse:2004:ICR

[vR04] Max-K. von Renesse. Intrinsic coupling on Riemannian manifolds and polyhedra. Electronic Journal of Probability, 9:14:411–14:435, 2004. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/205.

Venker:2015:ESU

[VS15] Martin Venker and Kristina Schubert. Empirical spacings of unfolded eigenvalues. Electronic Journal of Probability, 20(??):120:1–120:37, ???? 2015. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/4436.

vonSoosten:2018:PTU

[vSW18] Per von Soosten and Simone Warzel. The phase transition in the ultrametric ensemble and local stability of Dyson Brownian motion. Electronic Journal of Probability, 23(??):70:1–70:24, ???? 2018. CODEN ???? ISSN 1083-6489. URL

https://projecteuclid.org/euclid.ejp/1532570598.

Veto:2008:SRR

[VT08] Balint Veto and Balint Toth. Self-repelling random walk with directed edges on Z. Electronic Journal of Probability, 13:62:1909–62:1926, 2008. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/570.

Veraar:2016:CCM

[VY16] Mark Veraar and Ivan Yaroslavtsev. Cylindrical continuous martingales and stochastic integration in infinite dimensions. Electronic Journal of Probability, 21(??):59:1–59:53, ???? 2016. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1475266507.

Wagner:2006:PGB

[Wag06] Wolfgang Wagner. Post-gelation behavior of a spatial coagulation model. Electronic Journal of Probability, 11:35:893–35:933, 2006. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/359.

Wang:2005:LTS

[Wan05] Qiying Wang. Limit theorems for self-normalized large deviation. Electronic Journal of Probability, 10:38:1260–38:1285, 2005. CODEN ???? ISSN 1083-

6489. URL http://ejp.ejpecp.org/article/view/289.

Wang:2014:SAD

[Wan14] Ruodu Wang. Sum of arbitrarily dependent random variables. Electronic Journal of Probability, 19(??):84:1–84:18, ???? 2014. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/3373.

Wang:2020:FIW

[Wan20] Feng-Yu Wang. Functional inequalities for weighted gamma distribution on the space of finite measures. Electronic Journal of Probability, 25(??):19:1–19:27, ???? 2020. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.

org/euclid.ejp/1580958255.

Wang:2022:DTP

[Wan22] Zhe Wang. A driven tagged particle in asymmetric exclusion processes. Electronic Journal of Probability, 27(??):1–46, 2022. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-27/issue-none/A-driven-tagged-particle-in-asymmetric-
exclusion-processes/10.1214/22-EJP760.full.

Warren:2002:NMP

[War02] Jon Warren. The noise made by a Poisson snake. Electronic Journal of Probability, 7:21:1–21:21, 2002. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/120.

Warren:2007:DBM

[War07] Jon Warren. Dyson’s Brownian motions, intertwining and interlacing. Electronic Journal of Probability, 12:19:573–19:590, 2007. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/406.

Warfheimer:2010:SDI

[War10] Marcus Warfheimer. Stochastic domination for the Ising and fuzzy Potts models. Electronic Journal of Probability, 15:58:1802–58:1824, 2010. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/820.

Webb:2015:CPR

[Web15] Christian Webb. The characteristic polynomial of a random unitary matrix and Gaussian multiplicative chaos — the L^{2}-phase. Electronic Journal of Probability, 20(??):104:1–104:21, ???? 2015. CODEN ???? ISSN 1083-6489. URL http://

ejp.ejpecp.org/article/view/4296.

Webb:2016:LSC

[Web16] Christian Webb. Linear statistics of the circular β-ensemble, Stein’s method, and circular Dyson Brownian motion. Electronic Journal of Probability, 21 (??):25:1–25:16, ???? 2016. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1459960919.

Weber:2021:EII

[Web21] Frederic Weber. Entropy-information inequalities under curvature-dimension conditions for continuous-time Markov chains. Electronic Journal of Probability, 26(18):1–31, February 2021. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-26/issue-none/Entropy-information-inequalities-
under-curvature-dimension-conditions-for-continuous-time/10.1214/
21-EJP627.full.

Weill:2007:ARB

[Wei07] Mathilde Weill. Asymptotics for rooted bipartite planar maps and scaling limits of two-type spatial trees. Electronic Journal of Probability, 12:31:862–31:925, 2007. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/

view/425.

Weiss:2009:EBS

[Wei09] Alexander Weiss. Escaping the Brownian stalkers. Electronic Journal of Probability, 14:7:139–7:160, 2009. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/594.

Werness:2012:RSL

[Wer12] Brent Werness. Regularity of Schramm–Loewner evolutions, annular crossings, and rough path theory. Electronic Journal of Probability, 17:81:1–81:21, 2012.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
2331.

Werner:2021:CPR

[Wer21] Florian Werner. Concatenation and pasting of right processes. Electronic Journal of Probability, 26(18):1–21, February 2021. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-26/issue-none/Concatenation-and-pasting-of-
right-processes/10.1214/21-EJP611.full.

White:2007:PID

[Whi07] David White. Processes with inert drift. Electronic Journal of Probability, 12: 55:1509–55:1546, 2007. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/465.

Wieczorek:2015:SPM

[Wie15] Radosław Wieczorek. A stochastic particles model of fragmentation process with shattering. Electronic Journal of Probability, 20(??):86:1–86:17, ???? 2015.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
4060.

Winter:2002:MSA

[Win02] Anita Winter. Multiple scale analysis of spatial branching processes under the palm distribution. Electronic Journal of Probability, 7:13:1–13:74, 2002. CO-

DEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
112.

Windisch:2008:LCV

[Win08] David Windisch. Logarithmic components of the vacant set for random walk on a discrete torus. Electronic Journal of Probability, 13:28:880–28:897, 2008.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
506.

Windisch:2010:ERW

[Win10] David Windisch. Entropy of random walk range on uniformly transient and on uniformly recurrent graphs. Electronic Journal of Probability, 15:36:1143– 36:1160, 2010. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/787.

Wintenberger:2015:WTI

[Win15] Olivier Wintenberger. Weak transport inequalities and applications to exponential and oracle inequalities. Electronic Journal of Probability, 20(??):114:1– 114:27, ???? 2015. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.

org/article/view/3558.

Worms:1999:MDS

[Wor99] Julien Worms. Moderate deviations for stable Markov chains and regression models. Electronic Journal of Probability, 4:8:1–8:28, 1999. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/45.

Woodard:2009:SCT

[WSH09] Dawn Woodard, Scott Schmidler, and Mark Huber. Sufficient conditions for torpid mixing of parallel and simulated tempering. Electronic Journal of Probability, 14:29:780–29:804, 2009. CODEN ???? ISSN 1083-6489. URL http://

ejp.ejpecp.org/article/view/638.

Warren:2009:SED

[WW09] Jon Warren and Peter Windridge. Some examples of dynamics for Gelfand– Tsetlin patterns. Electronic Journal of Probability, 14:59:1745–59:1769, 2009.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
682.

Werner:2013:CS

[WW13] Wendelin Werner and Hao Wu. From cle(κ) to sle(κ, ρ)’s. Electronic Journal of Probability, 18:36:1–36:20, 2013. CODEN ???? ISSN 1083-6489. URL http://

ejp.ejpecp.org/article/view/2376.

Watanabe:2010:RTI

[WY10] Toshiro Watanabe and Kouji Yamamuro. Ratio of the tail of an infinitely divisible distribution on the line to that of its Lévy measure. Electronic Journal of Probability, 15:2:44–2:74, 2010. CODEN ???? ISSN 1083-6489. URL http:

//ejp.ejpecp.org/article/view/732.

Wong:2020:SPI

[WYZ20] Chi Hong Wong, Xue Yang, and Jing Zhang. Stochastic partial integraldifferential equations with divergence terms. Electronic Journal of Probability, 25(??):50:1–50:22, ???? 2020. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1588039469.

Wu:2017:BAE

[WZ17] Hao Wu and Dapeng Zhan. Boundary arm exponents for SLE. Electronic Journal of Probability, 22(??):89:1–89:26, ???? 2017. CODEN ???? ISSN

1083-6489. URL https://projecteuclid.org/euclid.ejp/1508292260.

Xiao:2006:SLN

[XKW06] Yimin Xiao, Davar Khoshnevisan, and Dongsheng Wu. Sectorial local nondeterminism and the geometry of the Brownian sheet. Electronic Journal of Probability, 11:32:817–32:843, 2006. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/353.

Xu:2020:HSL

[Xu20] Lu Xu. Hyperbolic scaling limit of non-equilibrium fluctuations for a weakly anharmonic chain. Electronic Journal of Probability, 25(??):1–40, 2020. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-25/issue-none/Hyperbolic-

scaling-limit-of-non-equilibrium-fluctuations-for-a-weakly/10.1214/
20-EJP488.full.

Xi:2017:LCL

[XYY17] Haokai Xi, Fan Yang, and Jun Yin. Local circular law for the product of a deterministic matrix with a random matrix. Electronic Journal of Probability, 22(??):60:1–60:77, ???? 2017. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1500602612.

Xu:2010:MPC

[XZ10a] Jing Xu and Bo Zhang. Martingale property and capacity under G-framework. Electronic Journal of Probability, 15:67:2041–67:2068, 2010. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/832.

Xu:2010:EEM

[XZ10b] Lihu Xu and Bogusław Zegarliński. Existence and exponential mixing of infinite white α-stable systems with unbounded interactions. Electronic Journal of Probability, 15:65:1994–65:2018, 2010. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/831.

Xu:2023:EPS

[XZ23] Lu Xu and Linjie Zhao. Equilibrium perturbations for stochastic interacting systems. Electronic Journal of Probability, 28(??):1–30, 2023. CODEN ????

ISSN 1083-6489. URL https://projecteuclid.org/journals/electronic-
journal-of-probability/volume-28/issue-none/Equilibrium-perturbations-
for-stochastic-interacting-systems/10.1214/22-EJP900.full.

Yamazaki:2023:TDM

[Yam23] Kazuo Yamazaki. Three-dimensional magnetohydrodynamics system forced by space-time white noise. Electronic Journal of Probability, 28(??):1–66, 2023.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-28/issue-none/Three-dimensional-
magnetohydrodynamics-system-forced-by-space-time-white-noise/10.1214/
23-EJP929.full.

Yang:2019:EUS

[Yan19] Fan Yang. Edge universality of separable covariance matrices. Electronic Journal of Probability, 24(??):123:1–123:57, ???? 2019. CODEN ???? ISSN 1083-6489.

URL https://projecteuclid.org/euclid.ejp/1573030844.

Yang:2022:LDS

[Yan22] Fan Yang. Limiting distribution of the sample canonical correlation coefficients of high-dimensional random vectors. Electronic Journal of Probability, 27(??):1– 71, 2022. CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/

journals/electronic-journal-of-probability/volume-27/issue-none/Limiting-
distribution-of-the-sample-canonical-correlation-coefficients-of-
high/10.1214/22-EJP814.full.

Yang:2023:HQU

[Yan23] Kevin Yang. Hairer–Quastel universality in non-stationarity via energy solution theory. Electronic Journal of Probability, 28(??):1–26, 2023. CODEN ????

ISSN 1083-6489. URL https://projecteuclid.org/journals/electronic-
journal-of-probability/volume-28/issue-none/Hairer-Quastel-universality-
in-non-stationarity-via-energy-solution-theory/10.1214/23-EJP908.
full.

Yao:2013:CWA

[Yao13] Changlong Yao. A CLT for winding angles of the arms for critical planar percolation. Electronic Journal of Probability, 18:85:1–85:20, 2013. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/2285.

Yearwood:2022:TS

[Yea22] Stephen Yearwood. The topology of SLE. Electronic Journal of Probability, 27(??):1–14, 2022. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/journals/electronic-journal-of-probability/volume-
27/issue-none/The-topology-of-SLE%ce%ba-is-random-for-%ce%ba4/10.
1214/22-EJP871.full.

Yoo:1998:USS

[Yoo98] Hyek Yoo. On the unique solvability of some nonlinear stochastic PDEs. Electronic Journal of Probability, 3:11:1–11:22, 1998. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/33.

Youssef:2013:ECR

[You13] Pierre Youssef. Estimating the covariance of random matrices. Electronic Journal of Probability, 18:107:1–107:26, 2013. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/2579.

Yang:2012:CED

[YP12] Yanrong Yang and Guangming Pan. The convergence of the empirical distribution of canonical correlation coefficients. Electronic Journal of Probability, 17: 64:1–64:13, 2012. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.

org/article/view/2239.

Zahle:2001:RBR

[Zäh01] Iljana Zähle. Renormalizations of branching random walks in equilibrium. Electronic Journal of Probability, 7:7:1–7:57, 2001. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/106.

Zambotti:2008:CEB

[Zam08] Lorenzo Zambotti. A conservative evolution of the Brownian excursion. Electronic Journal of Probability, 13:37:1096–37:1119, 2008. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/525.

Zeindler:2010:PMM

[Zei10] Dirk Zeindler. Permutation matrices and the moments of their characteristics polynomials. Electronic Journal of Probability, 15:34:1092–34:1118, 2010.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
781.

Zemlys:2008:HFS

[Zem08] Vaidotas Zemlys. A Hölderian FCLT for some multiparameter summation process of independent non-identically distributed random variables. Electronic Journal of Probability, 13:75:2259–75:2282, 2008. CODEN ???? ISSN 1083-

6489. URL http://ejp.ejpecp.org/article/view/590.

Zerner:2011:IP

[Zer11] Martin Zerner. Interpolation percolation. Electronic Journal of Probability, 16:33:981–33:1000, 2011. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/895.

Zerner:2018:RTC

[Zer18] Martin P. W. Zerner. Recurrence and transience of contractive autoregressive processes and related Markov chains. Electronic Journal of Probability, 23 (??):27:1–27:24, ???? 2018. CODEN ???? ISSN 1083-6489. URL https://

projecteuclid.org/euclid.ejp/1521079340.

Zhan:2006:SPA

[Zha06] Dapeng Zhan. Some properties of annulus SLE. Electronic Journal of Probability, 11:41:1069–41:1093, 2006. CODEN ???? ISSN 1083-6489. URL http://

ejp.ejpecp.org/article/view/338.

Zhang:2009:TCV

[Zha09] Yu Zhang. The time constant vanishes only on the percolation cone in directed first passage percolation. Electronic Journal of Probability, 14:77:2264–77:2286, 2009. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/

view/706.

Zhang:2011:SHF

[Zha11] Xicheng Zhang. Stochastic homeomorphism flows of SDEs with singular drifts and Sobolev diffusion coefficients. Electronic Journal of Probability, 16:38:1096– 38:1116, 2011. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/887.

Zhang:2013:HDL

[Zha13a] Liang Zhang. Hausdorff dimension of limsup random fractals. Electronic Journal of Probability, 18:39:1–39:26, 2013. CODEN ???? ISSN 1083-6489. URL http:

//ejp.ejpecp.org/article/view/2273.

Zhang:2013:DIS

[Zha13b] Xicheng Zhang. Degenerate irregular SDEs with jumps and application to integro-differential equations of Fokker–Planck type. Electronic Journal of Probability, 18:55:1–55:25, 2013. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/2820.

Zhang:2014:GPA

[Zha14] Lixin Zhang. A Gaussian process approximation for two-color randomly reinforced urns. Electronic Journal of Probability, 19(??):86:1–86:19, ???? 2014.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
3432.

Zhai:2018:ECC

[Zha18] Alex Zhai. Exponential concentration of cover times. Electronic Journal of Probability, 23(??):32:1–32:22, ???? 2018. CODEN ???? ISSN 1083-6489. URL

https://projecteuclid.org/euclid.ejp/1523325625.

Zhang:2019:DVK

[Zha19] Xicheng Zhang. A discretized version of Krylov’s estimate and its applications. Electronic Journal of Probability, 24(??):131:1–131:17, ???? 2019. CO-

DEN ???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.ejp/
1573527860.

Zhang:2020:LDS

[Zha20] Rangrang Zhang. Large deviations for stochastic porous media equations. Electronic Journal of Probability, 25(??):1–42, 2020. CODEN ????

ISSN 1083-6489. URL https://projecteuclid.org/journals/electronic-
journal-of-probability/volume-25/issue-none/Large-deviations-for-
stochastic-porous-media-equations/10.1214/20-EJP556.full.

Zhan:2021:TCG

[Zha21] Dapeng Zhan. Two-curve Green’s function for 2-SLE: the boundary case. Electronic Journal of Probability, 26(??):1–58, 2021. CODEN

????  ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-26/issue-none/Two-curve-
Greens-function-for-2-SLE--the-boundary/10.1214/21-EJP592.full.

Zhang:2022:BEB

[Zha22] Zhuo-Song Zhang. Berry–Esseen bounds for generalized U-statistics. Electronic Journal of Probability, 27(??):1–36, 2022. CODEN ???? ISSN 1083-

6489. URL https://projecteuclid.org/journals/electronic-journal-
of-probability/volume-27/issue-none/BerryEsseen-bounds-for-generalized-
U-statistics/10.1214/22-EJP860.full.

Zhou:2003:CBC

[Zho03] Xiaowen Zhou. Clustering behavior of a continuous-sites stepping-stone model with Brownian migration. Electronic Journal of Probability, 8:11:1–11:15, 2003.

CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/
141.

Zhu:2022:DND

[Zhu22] Theodore Zhu. The distribution of the number of distinct values in a finite exchangeable sequence. Electronic Journal of Probability, 27(??):1–25, 2022.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/journals/
electronic-journal-of-probability/volume-27/issue-none/The-distribution-
of-the-number-of-distinct-values-in-a/10.1214/22-EJP815.full.

Zervos:2019:DSS

[ZRLB19] Mihail Zervos, Neofytos Rodosthenous, Pui Chan Lon, and Thomas Bernhardt. Discretionary stopping of stochastic differential equations with generalised drift. Electronic Journal of Probability, 24(??):140:1–140:39, ???? 2019.

CODEN ???? ISSN 1083-6489. URL https://projecteuclid.org/euclid.
ejp/1575514915.

Zhao:2013:MLA

[ZZ13] Minzhi Zhao and Huizeng Zhang. On the maximal length of arithmetic progressions. Electronic Journal of Probability, 18:79:1–79:21, 2013. CODEN ????

ISSN 1083-6489. URL http://ejp.ejpecp.org/article/view/2018.

Zanella:2015:BSP

[ZZ15] Giacomo Zanella and Sergei Zuyev. Branching-stable point processes. Electronic Journal of Probability, 20(??):119:1–119:26, ???? 2015. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/4158.